A Bibliography of Publications in The International Journal of Quantum Chemistry:

1980–1989

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

 E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
            WWW URL: https://www.math.utah.edu/~beebe/

08 November 2023 Version 2.19

Title word cross-reference

(e, 2e) [AM83b, TMC84]. (n−) [MOT^{+}83]. (N − 1) [Mor89]. + [AM82, BNK86, Bow86, ES81, GT86, GT87a, GM86a, KU80, KS89b, KM89b, LB83, RBL84, SBE87, SS84b, SS87, YKZ^{+}89]. 1/Z [PS84]. $104.75 [Pic83b]. 12 [KU80]. 14 [SST84]. $15.00 [Mår80]. 1s [NMK84, Ten88b]. 1s2p_{J}^{2}p′_{j}J [ST84]. 1s2s [DS80a]. 1s2s2p_{j}J [ST84]. 1s2s^{2} [ST84]. 1s2s^{3} [DS81c]. 2 [BW85, MOT^{+}83, MK83, ddF86]. $26.25 [Wol81]. $28.50 [Pic83a]. 2′ [LMG85, SP82a]. 2: 1 [GP83a]. 3 [BW85, ddF86]. 3′ [KG80]. 3d [WS80b]. $49.00 [Sie83a]. 4L [FEKC84]. 4L + 2 [FEKC84]. 4 × 4 [Lin81a]. $59.00 [Gos80]. 5d [CPGP83]. $85.00 [LM81]. $88.00 [Mår82]. $9.95 [Lun82]. < p_{u}|∂/∂v|s > [JL85]. < r_{12}^{−1} > [BB81a]. < r_{1}^{−1} > [BB81a]. = [Dat82, KU80, PXLC86, PFLB83, WBK86]. ^{+} [CEB86, CLM80, CD89a, CC82b, Con87, DF89, DLVÖ84, EY82, ESKG83, GLLOB88, GP82, JŻ85, Kog84, Kun86, LPS83b, NMK84, PC83b, Sch80a, Sch81, Sch83a, ZJŻ88, vDTBN80]. _{3}^{+} [NPL89a]. ^{−} [AM83a, BN84, CS83c, HL80, LS86, SOS84a, SOS84b, SJ85, SHP^{+}80, vDTBN80, NS82b]. ^{−1} [Del87]. 0 [GLLOB88, Ing80, LC82a]. ^{1} [DS80a, LR89, MASF88, NS82b, SG81a, Vog86]. ^{1+} [TL82]. ^{1,3} [Mül81, DMB86]. ^{13} [FGM87]. ^{14} [JMS^{+}85]. ^{17} [LCCW87]. ^{19} [KGEP89, KG89, RDG^{+}86]. ^{1}Σ^{+} [DG85]. ^{2} [AM84, Kun86, Sim81, Var87]. ^{2+} [GM86a, GP83a, KD80b, LG84, TL82]. _{2}^{2+} [VBS85]. ^{2}Δ [KG86a, KG86b]. ^{2}p [Yar87]. ^{2}Π [LLS83]. ^{2}S_{1/2} [ST84]. ^{2}Σ^{+} [KG86a, KG86b, LLS83]. ^{3} [Bri87, GT86, KS89b, LCLC88, Lee83]. ^{3+} [CPS81]. ^{31} [Gup80, MG80]. ^{35} [KG89]. ^{3}Σ [NP80]. ^{4} [Bha80, Bri87, LC82a, PKP^{+}81, SS80b]. ^{5} [VARPU84]. 6 [PKP^{+}81, Tew87, Tew88a, Tew88b]. ^{N} [KG85, KG86c, SMJ83]. ^{N−1} [Win83]. ^{P he} [LCP81]. ′ [Gre82]. ′′_{2} [Mül81]. _{1} [ES81, FCPL84, KD80b, MASF88, PDMC80]. _{1−x} [CE86, CE87, CE88, GOLC87, LOGC86]. _{10} [KKTE80]. _{1u} [LR89]. _{2} [AYBP83, AJY80, AM82, ADT^{+}89, BBH87, BLR81, BMMS81, BF83a, BDK86, Bow86, CBV85, CMDBVMA89, CEB86, CKM81, CV82a, CHKB83, CL83b, CF81b, CH82, CP85b, Dat82, DPC82, DS81b, Del87, DF89, DLVÖ84, DYD89, Don82, DWQ86, ESKG83, ES81, Enz80, EHK83, ESCC85, FD89b, GVVDB89, GT86, GG82a, GDBRV87, GJMLRGSA85, GLLOB86a, GLLOB86b, GLLOB88, GA82, Gub82, HDFL84, HL80, Ish86, Ish88b, INO88, JTST87, JŻ85, JRK84, JR85, KH85, Kat80, KHCH85, KU80, KBT80, KTS86, KM89b, LPS83b, LJC83, LHC80, LB83, LFL^{+}81, LL89, LC88b, LGL^{+}84, MS87, MTWO83, MSI80, MK80, NT80, Nov83a, NP80, PZ84, PMOW89, PR83, PDC80, PDMC82, PDKC84, PSBB83, RS86, Rér89, RBL84, RS88, RKHK88, RKO^{+}89, SD80, SOS84a, SOS84b, SJ85, SJ88, SBE87, Sch80a]. _{2} [Sch81, Sch83a, SAS86, SBB87, SNNG86, Sie83b, Sok81, SSC83, SRL^{+}83, SR87, SS87, STG86, SS88, TNY86, TST89, TL80, VP84, VŻH82, Wea83, WHT86, Wój86a, Wój86b, YOJ81, YKZ^{+}89, ZW83, ZJŻ88]. _{2−x} [MS88]. _{2}^{+} [CEB86, CS80b, SBE87]. _{2}^{−} [AM83a]. _{2n} [GBO86a]. _{2n+2} [GBO86a]. _{2V} [PSBB83]. _{3} [BLR81, BMMS81, BR80, BFOP83, CBV85, CPS89, CD87a, CV82a, CHKB83, CF81b, DK87, DS81b, DSB84, Del87, ES81, FD89b, Ish86, Ish88b, KHCH85, KBT80, KKTE80, KTS86, LJC83, LS86, LWZ88, LGL^{+}84, Mül81, PXLC86, RKHK88, RKO^{+}89, SJ88, Sch83a, SRL^{+}83, SS87, SR80, SS88, TMO86, Wój86a, Wój86b]. _{3}^{+} [LWZ88]. _{4} [BLR81, BR80, FN84, FD89b, GA82, KBT80, KKTE80, KTS86, LH85, LLAG89, LGL^{+}84, MS88, NC84, PDC81b, SM85, TST89, VŻH82, WHT86]. _{4}^{2−} [LOR84]. _{5} [BLR81, BR80, BDK86, KBT80, LH85, MMB84, Sch80a, SJRP88, SHP^{+}80, TWL82]. _{6} [BR82, Dat82, GP86b, GVB80, HDTR83, KFD85, KKTE80, MOT^{+}83, PZ84, PG89, SHP^{+}80, TWL82]. _{6}^{n−} [MTO^{+}81, MK83]. _{7} [BFOP83, MMB84, Sch81]. _{8} [CH82]. _{a} [Cal84, PO85, XYSJ81]. _{α} [CPGP83, GLP83, GBO86b, IST89, OM81, PXLC86, RSGK88, RLS87]. _{b} [Cal84]. _{c} [LSW89, Sch88, SS88]. _{e} [DYD89]. _{eff} [FN83]. _{f} [YKZ^{+}89]. _{II} [MMB84]. _{K} [Nas83]. _{m} [Che80]. _{N} [Ser84, CBV85, Che80, Del87, INFU84, PFLB83, SOS84a, SOS84b, SJ85, Sin89, TL80, Wor81]. _{n−1} [SMD83, Del87]. _{n}^{2+} [GLLOB86a]. _{π} [HDFL84]. _{rs} [BAG83a]. _{σ} [HDFL84]. _{x} [CE86, CE87, CE88, GOLC87, KKA80b, LOGC86, MS88]. A + BC → AC + B [Var85]. α [Blo83b, BF81, BS80c, CY80, DG86, DC86, EHK83, GN87, HKK^{+}82, Jon84, JBW87, JR85, KT80a, LHC80, LHK81, Mof81, Mof82, MP84a, Nag87, SA80, SRP81, SC89, SPSG85, SNSM82, TMO86, TH81b, Vol88, vDTBN80]. α_{R} [MMB84]. B [Lee83]. β [Cio88b, DS80b, DM83a, EHK83, HKK^{+}82, LCP^{+}88, MSC^{+}83b, PKH^{+}81, SHH^{+}83a, Tho80, VU85]. β(1 → 3) [YL83]. β(1 → 4) [Yad83]. β^{−} [IHI80]. c [AWCL89]. L^{2} [YKZ^{+}89]. · [CLP83, Del87, LG84, LP80, SHP^{+}80, TTRL88, VŻH82, Wój86a]. d [BSZ81, FFF88, JN88, LÖ85b, Mod81, Mor83, RFST85, SIJ85, Yan85]. δ [HKK^{+}82]. Δ^{2} [Tew88a]. δE/δR [Eps87]. ΔE_{SCF} [PZ84]. E [JBW87]. E ⊗ ϵ [Sin81]. η^{5} [XHR86]. exp(−r) [Ten88a]. f [FF80, LÖ85b]. g [LHC80]. G(1, 2) [SL86]. γ [GM86c]. K [MTLL82, Ish83, Kut80]. k · p [DDP81a, DDP81b]. m [AW82]. μ [NDBE82]. μ^{+} [KB80]. N [BFC80, ERB82, GM86b, Gru85, HDO88, Kry82c, Mor89, ST89, WB89, BB82a, BKJ88, GOLC87, Gre81b, MMN87, OG86, Sch83b, SV82, TV83a, TV83b, TV84, TV89, Van85, VT85, Van86]. n = 0, 2, 3, 4 [SOS84a]. n = 1, 2 [SOS84b, SJ85]. n = 1−−4 [TL80]. n = 3−−7 [GBO86a]. n = 4, 3 [MOT^{+}83, MK83]. n^{ième} [SV82]. nl4f [JMS88]. ns^{2} [SB83a]. Ω [WS82]. O ⊃ T ⊃ D_{2} [Her84]. p [FC80a, FC81a, LÖ85b]. π [BB82a, CGT86b, Cio88b, DMB86, FMF80, GTB80, Gre87, Has82a, Has82b, Has82c, KSSH85, PMT84, TN80, WK83, YZ85, Živ86, Živ87c]. R [EKBE86, Lin89]. r_{12}^{−1} [LKB89]. ρ_{2} [CL85]. → [AM82, ES81, GT86, GT87a, HDFL84, KS89b, LB83, NS82b, SBE87, SS87, YKZ^{+}89]. ⇀ ↽ [TOK80]. S [Ave84c, CJZ^{+}86, LÖ85b, RSST86, STS83]. S_{N1+N2} ⊃ S_{N1} ⊗ s_{n2} [Wen86]. σ [Bha82c, Kle83, PMT84]. SO(3) [ACCYCC80]. SU(2) [ZS85]. su(n) [TV83a, TV83b, TV84, TV89, Van85, VT85, Van86]. T [WS82, Lin81b]. T_{d} ⊃ T ⊃ D_{2} [Her84]. U(n) [Bor82]. U(n = n_{1} + n_{2}): U(n_{1}) × u(n_{2}) [GP86a]. U(n_{1} + n_{2}) ⊃ u(n_{1}) ⊗ u(n_{2}) [Wen86]. V [ACCYCC80]. φ = A exp(−αr^{ν}) [Tay81]. φ_{4} [ECM89]. xy [LB86]. Z [BVGV89, Kin89, Mez83d].

* [BB82a].

-5 [Gre82]. -adrenergic [PKH^{+}81]. -adrenoceptor [DS80b]. -aminobutyric [GM86c]. -ary [OG86]. -band [Mor83]. -bonded [Kle83]. -carbolines [MSC^{+}83b]. -catalyzed [SRC84]. -chymotrypsin [SNSM82]. -classes [FEKC84]. -coupled [ERB82]. -coupling [ACCYCC80]. -cyano [Mof82]. -cyanocarbonium [Mof81]. -cyclodextrin [MP84a]. -cyclopentadienyl [XHR86]. -D [ddF86]. -Decay [IHI80]. -deoxy- [LMG85]. -deoxyfuranose [LMG85]. -deoxyuridines [SP82a]. -dimensional [AW82, GM86b]. -donor [Bha82c]. -doped [BKJ88]. -electron [Cio88b, Sch83b, Živ86, Živ87c]. -end [KG80]. -endorphins [LCP^{+}88]. -fluorofuranose [LMG85]. -function [Jon84]. -guanosine [Gre82]. -guanylhistamine [MTWO83]. -H [GLLOB88]. -helix [Vol88, vDTBN80]. -HMX [HKK^{+}82]. -hydroxycytosine [PKP^{+}81]. -isocyano [Mof82]. -isocyanocarbonium [Mof81]. -isopentenyl [Tew88a]. -Lactam [DM83a, SHH^{+}83a]. -Local [Gre81b]. -matrix [Lin81b, Ave84c]. -methylguanine [PKP^{+}81]. -network [KSSH85]. -O [EHK83]. -OH [LHK81]. -Orbital [GTB80]. -orbitals [RSST86, STS83]. -particle [Kut80, BFC80, HDO88, Mor89]. -Point [MMN87]. -propiolactone [Tho80]. -quartz [SC89]. -receptor [MTWO83]. -receptors [PMOW89]. -representability [Gru85, Kry82c]. -spiral [VU85]. -split [WK83]. -strictly [Ing80]. -SW [RLS87]. -symmetry-invariant [SM85]. -systems [Gre87]. -th [SV82]. -time-dependent [AM84]. -tocopherol [SPSG85]. -triazenes [SJH89]. -Type [GOLC87]. -U [Wor81]. -weighted [Kin89]. -well [WB89].

/2 [GjC86]. /2H [MH82]. /3 [DA80, MH82]. /3H [MH82]. /Boyd [LWK84]. /INDO [RB88a, FF80]. /LUMO [EIMS88]. /MERGE [HKKM83, HKLM85, LHK81]. /Na [RSGK88]. /Propagator [Bak85]. /S [Deá83]. /VRDDO [LHK81, HKKM83, HKLM85, SHP^{+}80].

1 [Boč87, Boč88b, GSRG80, KRST83, PM86, SJH89]. 1- [Gre82]. 1-dimethyl-3- [SJH89]. 1-hydrogen [BBLR84]. 1-nitropropene [DT83]. 1-Reduced [Tho86a]. 1/Z [PM86]. 10-electon [KTS86]. 102 [BV86]. 12-dimethylbenz [Sey80]. 13- [LL82a]. 13-carbon [KL80]. 14- [LL82a]. 14-electron [LM86]. 164 [SBE^{+}83]. 175-ASP [LCA89]. 1977 [ADL78, Kel78]. 1978 [Cal81a, Cal81c, NB79]. 1979 [Ågr80, Cal81b, Cal82b, Dun80, Pul79, Sie81]. 1980 [Cal82a, PTG80]. 1981 [Ågr82, WJ82]. 1983 [Ler85]. 1984 [Ågr85, Cal86d, Tap86]. 1985 [Ågr87a, Ågr87c, Bra86b, Brä86a, Cal86c, Cal88, Muj86, Ågr88]. 1986 [Ågr87b, Ågr88, Fro88a, Fro88b, Han86, Kum87, Mic87, Nat86, Rat87, Tap88]. 1987 [Fax87, Mou89, Nat88, Rat88]. 1988 [Gee89, LGJLCB89, Lun89b]. 1s [Kog84, Har87].

2 [CV82a, DS81b, DBS82, KN80, LM81, Lás82, Lip83, Mar81a, MBKH83, MH82]. 2- [LOB85, POS81, SMD83]. 2-AHP [BB85]. 2-aminomethyl-3-carboxymethyl [SRS81]. 2-anilino-6-naphthalenesulfonate [WS80a]. 2-benzosemiquinone [SL80]. 2-dioxethane [Hil81a, Hil81b, Hil81c]. 2-electron [SvG89]. 2-fermion [Cas88]. 2-HP [BB85]. 2-methylthio-N [Tew88a]. 2-naphthyl [ST89]. 2-naphthylnitrenium [Gre82]. 2-orbital [HM89b]. 2-oxetanone [Tho80]. 2-oxopyridine [DKLP81]. 20-oxopregnane [VT82]. 20th [Urr82]. 20th-century [Urr82]. 21G [RR87b]. 22- [LL82a]. 235-TRP [LCA89]. 25 [Pic83a]. 2H [MH82]. 2P [Har87].

3 [SP81, Fro81, RR87b, Tak86]. 3- [BVDA88]. 3-21G [RR87b]. 3-540-09738-4 [Fro83a]. 3-540-10005-9 [Fro81]. 3’-amino-3’-deoxyadenosine [SP81]. 3-deazapurine [SC84a]. 3’-deoxyadenosine [SP81]. 3-dimensional [YKZ^{+}89]. 3-methyl-fentanyl [CL83a]. 3-methylcholanthrene [HKP81, HPK81]. 3-sigmatropic [RBR80]. 3-tetrathiofulvalenes [POS81]. 30 [SST83]. 3000eV [SST83]. 300K [CP83b]. 32 [NSP80]. 3D [Har87]. 3G [WK83, Zie82]. 3H [MH82]. 3s [Mül81].

4 [KRST83]. 4-Benzodiazepin-2-ones [KRST83]. 4-cyclohexadiene [SAM82]. 4-diamino [PBM85]. 4-electron [HM89b]. 4-hydroxy-2-pyridinone [Sch82]. 4-hydroxytautomeric [Zie82]. 4-hydroxyuracil [SZR80]. 4-oxopyrimidine [LOB85]. 4-tetrahydroacridine [PBKB89]. 450 [CLL86, LL85]. 4F [Har87]. 4H [FA86]. 4H-pyran-4-one [FA86].

5 [RWOT81]. 5-dinitrobenzoyl [ST89]. 5-ethynyl- [SP82a]. 5-fluorouracil [Zie82]. 5-HT [RWOT81]. 5-HT/LSD [RWOT81]. 5-hydroxytryptamine [MWO^{+}84]. 5-Isoxazolol [BVDA88]. 5-methoxyuracil [Zie82]. 5-methylchrysene [Sey80]. 5-substituted [BCD^{+}80, GJMLRGSA85, SP86b]. 5-vinyl- [SP82a]. 500 [GSRG80]. 500-eV [GSRG80].

6 [Sme84]. 6-dimethyl-N-nitroso [LHK81]. 6-methano- [BBGS84]. 64asp [NSP80]. 692 [Cal86b]. 6R [Dat82]. 6th [Löw89].

70 [BH89a].

80 [Pic83b].

9 [Fro81]. 9-amino-1 [PBKB89]. 9-methylguanine [PSS86]. 90 [CP83b, LM81, Mar81a, Sme84]. 90-277-1039-2 [LM81]. 90-277-1171-2 [Mar81a]. 90-277-1488-6 [Sme84]. 90-years [LGJLCB89].

= [GVVDB89].

A&M [Pic83a, WJ82]. A-G [SRSR85]. A. [Mas87a]. A/T [GUSS89]. A1 [CE87]. A23187 [GP83a]. AAF [HB82]. Ab-Initio [RKO^{+}89, Abd80a, CHKB83, GAL86, INO88, Kau83, LS80, LT81, RB81, RKHK88, Tho80, Ågr87c, GM81c, JN88, LHK81, PKC85, SRT81, SRL^{+}83]. ability [Bha82c]. abortifacient [KD80b]. abortificient [Kot81]. absence [SCER88]. Absorption [BCD^{+}80, HHL80, BSD82, DPZ80, GF88, GC84f, Hir80c, LaF89, Suh86, VPK80]. Abstract [Ano85a]. Abstracts [Ano80a, Ano81a, Ano82a]. Academic [Cal86d, Sie81, Sie83a, Tap86].

Academy [Mår82, Ano84a]. accelerators [WGS85]. acceptable [Gos85]. acceptor [CE87, DBCJ86]. access [BVGV89]. accessibility [LCP81, LPP81]. accessible [FCPL84]. according [GBO86b]. account [SP82b]. accumulation [PZPL86]. accuracy [Blo83a, Dea89, Gre81b, GA82, LRT88, Moh80b, Moh80a, MW80, PWL89]. Accurate [EB87, KD86, SD86, DL83, Fli84, PABD80, Umr89, AM83a]. acetal [TK83]. Acetaldehyde [YG86]. acetamide [FGC84]. acetic [LGL^{+}84, TH88]. acetone [LM82a, SB81, SB83b]. acetone-related [LM82a]. acetonitrile [LP80, MBR80, SA86b]. acetyl [Yad83]. acetylcholine [RTM81, YHG89]. acetylene [Bak85, CPMP86, GG82b, HSR86, LRZ84a, LWZ88, SJ88, SS84b, ZCH89]. ache [LHKK82]. achiral [CT87]. acid [AN86, BMB87, CA86, DDFA87, DC82b, DBC84, FW84, FCPL84, GLP80, GM86c, GSS85b, HM84, HP81, HKG81, LCCB81, LT81, LGL^{+}84, MR85, OLB^{+}82, PMCA88, PSS^{+}87, PS81, SA86b, Shi85, SHK87b, ŠHH83b, SM80, SG86, Tew87, Tew88a, Tew88b, TLC87, VCB83, WHT86, Yad83, Yon85]. acid-base [CA86]. Acidic [RS82]. acidity [HL80]. acidium [RT85]. acids [AEHM85, BK84b, BVGV89, CGT86b, CC82a, CAML87, DPZ80, DF89, Dos84, FW84, GGM81, JM81, MY81, Mat82, MG89, MLA80, MGBH81, PMCA88, PP80, SJ86, SMHK89, SK81, TH88]. ACS [Ano80b]. Acta [Zer87]. action [Ben81, KD80b, Kot81, MG89, NSP80, Sey81]. activated [Sla83b, Yam81]. activation [CSB86a, LČ84, MTWO83, PMOW89, SG81b, CPN89]. Active [Pan81, vDTBN80, AWCL89, BJR86, CLA83b, DOW87, GM81c, GB81, KTK80, KRST83, KRBM85, KNS^{+}86, KKSM87, KNSM88, KMPTK89, KK82b, KK83, KKD80, LER87, Pin88, RSP^{+}80, RJBR^{+}87, Roo80, TLT^{+}84, YL83]. Active-site [vDTBN80]. activities [FHH^{+}87, XYSJ81]. Activity [RP81b, AABW80, AAW82a, AAW82b, CL83a, CGP87, Cow82, GDR85, GP80, LPW83, LP81b, LPSK83, Mar89, MSC^{+}83b, PP86, PL84a, Ran84, RJBR^{+}87, SRO81, Sar89, Sey81, Sok81, SSC83, Vol88, Vol89b, YGK^{+}89, YKTA80]. activity-linked [GP80]. Adamowicz [FSP87]. Adamson [Lun82]. adaptation [Cal85, GCP87, Mat84, MNS85, PP89, RSD82, SPÖ89, Tai89]. adapted [CK82, CS80b, FS80, GC87, KC82a, LER87, Pan82a, Pan82b, Pan83, Pan84, PP89, Voj85]. adapting [Muk84]. adaption [GC82a]. adaptive [Koh86]. adatom [TW88]. adatoms [LK81, MM80]. Addenda [Sab80a]. Addendum [CT81, Löw81a]. addition [JL84, LY83, LPCF80, SS86c, SS86d, Wea84, Wea88, Nov83b, Pan82a]. Additional [KHP81, LK87]. additions [ZCH89]. additive [GCP82, GPC85, GCP86]. Additivity [EB81, KS86c, YG86]. address [Fax86, Wij83a]. adduct [CLL86]. adenine [BCD^{+}80, Gre82, MGN82, SJ86, SS84a, Tew87, Tew88a, Tew88b]. adenine-amino [SJ86]. Adiabatic [DS88, GM86b, SG85a, KL81, SST83, DS88]. Adions [LK81]. adjacency [Kia80, YS80b, YS80c]. Adjoint [GC84d]. adjusting [SS86b]. adjustment [ZNS80]. adrenergic [PKH^{+}81]. adrenoceptor [DS80b]. adsorbate [MBD84]. adsorbate-substrate [MBD84]. adsorbates [GM80]. adsorbed [TW88]. Adsorption [Sla83a, Sla84, Bou84, GM81a, JA82, LHH^{+}83, MB89, PH87, WD81]. Adsorption-complex [Sla83a, Sla84]. Advanced [ADL78, Cal82a, Cal86b, Dun80, Kel78, NB79, Cal81c]. Advances [Cal89a, RC86, Cal86b, Csi82, Don79, Her83, Fro88a]. AF [HB82]. affecting [Che82]. affinities [BCA87, Ort87, WSD80]. affinity [KSK86, LC88b, ÖHD89, YGK^{+}89]. aflatoxin [FCPL84]. aflatoxins [LP81b]. Ag [MB84, PC83b, SD80, SAS86, TW88]. against [SHK87a]. agarose [CFC83]. ageing [Vas82]. agent [RT85]. agents [DS80b, ECM89, KKSM87, Pul86, SS80f]. aggregates [NSN89, Oga80]. agonism [BVDA88, DS80b, MTWO83]. agonist [CZM83, TLT^{+}84]. agonists [YHG89]. AGP [Gos82a, Öhr86, ESKG83, JWOÖ82, JWÖ83, KEGS81, KWJ83, Lar86a, LWÖ81, Öhr85, OWÖ81, SGEK81]. agp-based [Öhr86, Öhr85]. agreement [Ran80d]. AGS [KM89b]. AH [Del87, Del87, PFLB83]. Aharonov [Nag85]. AHP [BB85]. aid [Var86]. aided [FP84, NS89]. aids [Mez84e]. Al [CE86, CE88, Ågr88, GOLC87, LGL^{+}84]. Ala [MMB84]. alaninate [ST89]. alanine [HGHGF^{+}89, NSL88, YJ85]. alanylcarbonyl [Tew88b]. Albert [Fox87, Lad87]. albumin [Gas80]. albumin-methylglyoxal [Gas80]. alcohol [CD89b, CF81b, LBV86]. aldehyde [Abd80a]. Alder [Bur86]. algebra [ACCCBF80, DH84, GCP87, LP89, Löw81a, Löw82b, PS88, Pan85, SPRS89, SLT81, SLT84a, TSL80, TV83a, TV83b, TV84, TV89, Van85, VT85, Van86, KG85, KG86c, KG86d]. Algebraic [Cio87a, DK81a, DK81b, IL83, Lin81a, MBT81, MPB84, AČP82a, BS82b, Dem86, DYD89, HL88, Mez84a, RB89, Wei88]. algèbre [TV89]. algègbre [TV83a, TV83b, TV84, Van85, VT85, Van86]. Alger [Cal81a]. algorithm [BD88, BEO80, CNT89, Duc86, GHGP89, Kar84b, Lin89, NKV^{+}89, Pau81, TS80]. algorithms [CY80, HV83, LRT88]. AlH [BLR81]. aliphatic [BNS^{+}87, CBV85, DBC84]. alkali [BKJ88, GEdP81, GP85, PD84a, PGM80, RSSS81, SS86a, TLDK89, SSF^{+}84]. alkaline [GEdP81, LBP84, SS81a]. alkaline-earth [GEdP81]. alkane [ADF^{+}88]. alkanes [Ran83]. alkyl [HMF82, Mof81]. alkyl-substituted [HMF82]. alkylated [FPDA89]. alkylating [TR86]. Alkylation [DD89]. all-electron [SG89, WKJ88]. all-trans-polyenes [GBO86a]. all-valence [Lip88]. allene [DBD^{+}85]. allomorphic [CLP82]. allowance [ZV85]. allowing [SV82]. alloying [SR84a]. alloys [FK83, Zun85]. allyl [PCA89]. along [Hil81a]. alpha [MTWO83, SSYR87, VB85]. also [Dun88]. alteration [NBE87]. alternance [ZRY83]. Alternant [Živ87a, ČPV81, Fer88, Has82b, JT85, Živ87b, Živ87d, ddF86, PP83]. alternating [Poh84]. alternation [GK80a, PC83a, PCG83, PT84, PP83, SI89b, TP84, TP85]. Alternative [GB85b, Tim86, DH85, FDAC89, MT87]. Altmann [Rös87]. aluminum [Boe88, IST89]. Amaldi [DCF82b]. American [Bry80]. Ames [LSH88]. amide [OLB^{+}82]. amides [DA80, KNS^{+}86]. amidogen [AYBP83]. amine [Abd80a, BPCT83, Gas81, KS87b]. amines [PBKB89, RT84]. amino [Dos84, FW84, GGM81, JM81, Mat82, PMCA88, PBKB89, SJ86, SP81, Shi85, SMHK89, SK81, SG86, VCB83, Yon85]. aminoacyladenylate [BCV81]. aminobutyric [GM86c]. aminomethyl [SRS81]. aminopyrone [MBKH83]. aminotetralins [KŠHH83]. aminothiol [VV87]. ammonia [BFOP83, EV87, NL89, OLB^{+}82, PMCA88, SS82, SSR^{+}87, OLB^{+}82]. ammonia-formic [PMCA88, OLB^{+}82]. ammonia-glycine [OLB^{+}82]. ammoniac [OF80]. ammonium [GP85, OF80]. AMO [ABD^{+}80b]. among [Lar86a, YGK^{+}89]. amorphous [CJ89, MTI80]. amount [LCD82]. amplitude [Suc81, YKZ^{+}89]. amplitudes [GM86a, HT82]. Amsterdam [Ågr80, Cal86a, Tap88]. Amsterdem [Ågr87c]. amu [CPS81]. analgesia [CL83a]. analgesic [CL83a, KKSM87]. analgesic-antipyretic [KKSM87]. analog [SRS81, VARPU84]. analogs [BCD^{+}80, CL83a, FZZ87, GT86, HGHGF^{+}89, MWO^{+}84, PKP^{+}81, SRO81, SC84a]. analogue [EKBE86]. analogues [AIJW82, CPZS86, FA86, RR87a, RR88, SWA82]. analogy [SIN80a]. Analyses [Toy80]. Analysis [DF83a, FGM87, FM86, Gre87, Hib81, Hur82, Ojh89, dB84, ADF^{+}88, AWCL89, BLR81, Big86, BH82, BR81b, BPCT83, BDD81, CG80, DPS87, FN84, FZZ87, FC80b, FC81b, FC81c, FC81d, FC82e, JI88, Kar84b, KB80, Kry82c, Low82d, LP87, MH80b, MH89, Mal91, Mal02, May84, May85, Mez83a, MSF86, NPL89b, NC84, PB82, RW80a, RM84, RDG^{+}86, SGEK81, SL81, SKEDR81, SS80e, Shi84, SST84, Sjö87, SH87a, SWA82, SM80, Sto80, SZJS82, SSZJ82, SNSM82, Tva89, TWW81, VARPU84, WS80a, Wol81, ALH86, Has82a, Sch81, Tay86a, Tay86b]. Analytic [GRS^{+}82, KK82a, SSR^{+}87, Tab82, WW89, BFJ81, GSRG80, JBW87, Ori85b, SM85, WJ80, Wea88]. Analytical [ALB84, GLLOB88, JEW88, Kar84a, LB86, Sel86a, Sel86b, AMFC83, CTV80, CT81, FB83, Gus87, HA84]. analyzer [SL81]. andalusite [HHL80]. Anderson [MMN87, MA80a, MBD84]. André [Cal81c]. Andrew [Ano87s, Buc89]. Anesthetic [DBCJ86]. anesthetics [RS82]. angle [CL83b, POT80, SKIK82, TNT^{+}86]. angles [ZLH87]. Angular [AW82, MNS85, Ols86, UB85, AØ80a, AL86, Mic81b, YKZ^{+}89, ZLH87]. Angular-Momentum [Ols86]. anharmonic [Bha80, CV85, ČV86, Gra82, GC82a, Sil82]. anharmonicity [VM81]. anhydrase [CLA83b, GP80, KR81, KBSG81, LL89]. anhydrase-B [KR81]. anhydrase-catalyzed [LL89]. aniline [SG81a]. aniline-binding [SG81a]. anilino [WS80a]. anilinonaphthalenesulfonate [WS80a]. anion [AEM86, PBA88, SS86a, SL80]. anionic [GMP^{+}84a, RJ88, ŠKH81]. anions [HL80, Sim82b]. Anisotropic [TTS84, VI84]. anisotropics [PGM80]. anisotropies [ST86a]. Ann [Mår82]. annealing [DBF89, MSWC88]. annihilation [Ori82a, Pan82b]. anniversary [LGJLCB89]. Announcement [Ano81b, Ano82b, Ano82c, Ano83a, Ano84a, Ano85b, Ano86a, Ano86b, Ano88a, Ano88b, Ano88c, Ano88d, Ano88e, Ano88f, Ano88g, Ano89a, Ano89b, Ano89c, Ano89d, Ano89e, Ano89f, Ano89g, Ano89h, Ano89i, Ano89j, Ano89k, Cha85, Jor86, Lan86, LB82, Löw86a, Ano88z, Ano88y]. Announcements [Ano81c, Ano82d]. annulene [BBGS84]. Anomalous [GSG84, Kar80, PFC88, BMMS81]. Anomeric [KVR85]. antagonism [BVDA88]. antagonist [CZM83]. antagonists [WWD88]. Anthony [Pic83a]. anthracene [MTI80, PZSL83, Sey80]. anthracenes [Hir80a]. anthraquinone [PBM85]. anti [KKSM87, HB82]. anti-inflammatory [KKSM87]. antibacterial [DM83a]. antibiotic [KMRK86]. antibiotics [CS81, SC80a, Sar81, SP81, SP82a, SP86b, Sar89]. antibodies [SBH^{+}85, SBH^{+}86]. antibody [BR81b, CPZS86]. antibond [TWW81]. anticancer [BC80, BV86]. anticholinesterases [HLKK82, LHKK82]. anticonvulsant [AD80]. antiepileptic [JW86]. antiferromagnet [KBT80]. antiferromagnetically [MK80]. antigen [BR81b, YL83]. antigen/antibody [BR81b]. antiinflammatory [MG89]. antiparticle [Jar89]. antipyretic [KKSM87]. Antisymmetric [Røe81a, Gru85, KE82]. Antisymmetrized [Gos82a, DP81, EÖWZ84, LER87, SPÖ89, WK85]. antitumor [AAW82b, CGP87, KMRK86, Pul86, TH88]. antitumour [AAW82a]. any [Kob82, RB88c]. aperiodicity [IA87, Lad83]. apparatus [Tac89]. Applicability [JP80, WJP83, WJP85]. Application [ACCCBF80, ACDLH84, BS80a, CSK87, ESKG83, FD86, JR85, KCK83, Khe81, KTK80, KRST83, KRBM85, LM82a, LAGS89, LLC83, OS89b, SL81, SI89b, Tac89, TR82, Zhe83, ADF^{+}88, BVAB81, Cal86b, CS80a, De 80, Don79, DSG80, FEKC84, JD85, JSSS80, JT85, Kim81, KWJ83, Loh87, MSWC88, Ott86, Sil86, SRSR85, Tai89, WOT89, AN85, ANH81, AM88b, BR81a, BPCT83, CPS81, CD83, Chu86, ČV86, DDP81b, DM85, DG85, DG83, Dov84, GM86a, GA82, HH89, Jun80, MK81, MRJ85b, Pan85, Rér89, RSGK88, STH80, Shi87, SB81, SB83b, SSR86, SSRR87, VT85, YOJ81]. Applications [Fro81, ACGT89, Bal82d, Bor82, Bor85, ČU80, Cio87b, DMB^{+}84, DMB89, DL82, DBES80, Dur86, Fle86, Fuk81b, HG86, HR83, JJZ^{+}88b, Kha85, LFL^{+}81, PBKB89, Rat87, SS89b, WGS85, Bal88, BMRJ86, Fau85, GCP82, HL83, Kin86, KWBC80, TR80, Zer87, Mår82]. applied [ERB82, Kal85, LGS87, NT80, Ols86, Rab84, RVK80, Sie83b]. Applying [Ort87]. Appraisal [SMS84, RFST85, STS83]. Approach [GT83, KB87b, AEM86, AØ80a, AØ80b, AØM81, AØC82, AA89, BIN86, BT84, BM89, BS82b, BL81c, Csa86a, Csa86b, Dem86, DÖL87, DYD89, DBF89, Dov84, DK82b, EB82a, EB82b, EB82c, EN87, GCP87, GVVDB89, GV88, Gla85, GjC86, GC84b, GC84c, GC84d, GC84e, GC85b, GP86a, Gou86, GM81b, Gro85, GDR85, GLP83, HA84, HYZ89, IL83, IRP81, IA87, KS82, KS84, KS86a, KN84, KN86, KL81, KPS87b, KPS88, KB87a, Kva83, LL86, Lar86b, Leu83, LZ89, LP89, Lin81b, Lin85, LSBS81, LY83, Loh87, Mar88c, Mat87b, Mez83a, Mic83a, MT87, MPB84, PT84, Pay82, PSG83, PNA86, PNA88, RB88a, Ran84, RZ86, RB89, SBNK89, See84, Ser84, Sim86, SS89b, SHPK80, SM87a, TP84, TP85, TT81, VTT81, VVD80a, VV87, VM81, Was89, WJ80]. approach [Wei82, WG82, Yam80b, YY83, Yon85, Zhe83, Živ86, Živ87c].

Approaches [FMF80, AM89, KSSH85, PTC84, SCER88, WB89, WP80, ZA89]. approximants [Bha82b, CB88, Cio88a, ČVV87]. Approximate [Flu85, GKAP81, KM89a, Löw85a, SMMM87, SK80, Sen80a, AGB82, AMFC83, BH82, CTV80, CT81, DCF82a, Don81, FC82c, Fli84, Gee88, Gre81b, GA82, HO85b, Mar88c, Noa80a, Noa80b]. approximating [Has82c, BK84a]. Approximation [Chu88, BNK86, Böh83, BFC80, BLV88, CMDBVMA89, Cof83, DS87, DLVÖ84, DÖ89, EC83, Eng83b, Gee88, Kir89, Lar83, Lef83, MÖ80, Nag85, Nas83, NPL89b, OK80, PPB87, PGG80, PMT84, Sha85, ŠKR84, WG80, WK85, Zho88, KD80a]. approximations [Bec85, GD89, HS89, Löw82c, MPP83, Noa80a, Noa80b, PLL86, Pul83, WG82, Wei85, ZA89]. April [Pul79, PTG80]. APSG [DP81]. APW [DDP81a, DDP81b]. aqueous [BCD89, CD89a, CAML87, GjC86, JRK84, MB86, New80, PV89]. aquo [Bha82c]. arbitrarily [BL81a]. arbitrary [Csa83, DAPS81, DS88, FW80a, FW80b, Hel82, Ish83, KS86a, LW81, Wea82, ZLH87]. Archimedean [KGB87]. architect [Urr82]. architecture [DF83b]. areas [Hay82]. arene [LPCF80]. argon [KWBC80]. ArH [CLM80]. arising [Ish83, Ish84a, Ish85, MTLL82]. ArN [GDBRV87]. ArO [GDBRV87]. aromatic [GK80a, JI88, NTM80, SS84b, SVS81, Sey83b, Sey83c, SL87]. aromaticity [Gre87, RHST88]. arrangement [GB81, KKA80b]. array [UV83]. arsenide [PPB87, RZWL80]. art [Bar85, Löw86c, Ågr87a]. Arthur [Lun82]. artifact [SHPK80]. artificial [Lip83, VB85]. ary [OG86]. aryl [Che82]. Aryloxy [Sey80]. As/GaAs/Ga [CE87, CE88]. Ascorbate [MPP^{+}81]. ascorbic [LT81]. ASI [Fro88a]. ASIF [FCPL84]. ASP [LCA89]. aspartate [NS83]. Aspects [Muk86, ABD81, Blo83a, BT82, BCD89, CGG81, ČV82b, DSG86, DFH84, DC87, FG82, JMD88, KB87a, Löw84, NB79, PGL^{+}86, Pau81, PBBK89, PZSL83, SR81, SA86a, SMJ83, UB85, Umr89, Voj85]. assemblies [Sch80b]. Assessing [Mcd81a]. Assignment [SL80, SS86b, YMO80]. assignments [CHB88]. assistance [HR84, MR87]. associated [Ban82b, KEB87, Poh84, SS89a, UT80]. associates [DS89a]. Association [ENH^{+}87, Del88]. associative [Koh86]. assumption [YG86]. asymmetric [Mez87a, RB88c]. asymmetry [FGS85]. Asymptotic [Bli80, Lev89, Shi84, VM83, YLM80, Bha82b, Cha82, Emc80, Ing80, Khu82, Sil86]. ATM [PFBF81]. atmosphere [SOS84a, SOS84b]. atmospheric [DSW80]. atom [AB87, AC85, AGB82, AMFC83, BS80a, CV83b, DM83b, DBS80, DD87a, EB87, Ess82, FC85, Fro83a, GMV83, HT85, IL84, JSSS80, KEGS81, KWJ83, Lin89, MMS^{+}86, Mcc83, Mic83a, Min80, MG85, NP81, Ols86, Par84, PD84b, PNA86, PNA88, RI88, SD86, SI89a, SIN80a, SDBDSMG86, Suc81, TH81a, Tay81, TB81, TTRL88, UB86, Var81b, VM81, VV83, Vrs87a, Vrs87b, YUKK84]. atom-diatom [BS80a, NP81]. atom-in-molecule [PNA86]. atom-molecule [VM81]. atom-surface [IL84]. Atomic [Ano80c, DB81, FSK79, FSS^{+}89, RLPFA86, See88, Ten86a, AN85, ADF^{+}88, BIN86, CS80a, CPS81, ČV82b, Csa85, Csa87, Csa89, Del83, DIV86, DCF82a, Dov86, ES80, Ela87, Fra84, Fra85, Fri86, GB85a, GSMRM85, Gla85, Gra84, Gre81a, HR83, Jun80, KD80a, Lás85, LP87, Mar85b, MC88, ML81, Mic83b, MS80a, MB80a, MK81, MPPW82, Niu84a, Niu84b, Nov83b, Oga80, OS89b, Pia83, Pra83, PC86, RA81, Rei82, RACPL86, SM85, Sen80a, SHK87b, SMHK89, SF80b, TH81a, TS80, Ten86b, Ten88b, Tru83b, WOT89, YHG89, YLM80, Fli86, Ågr80]. atomization [Živ83]. Atoms [Bra86b, Ågr88, AEHNEA85, ACP84, AM83b, ABD^{+}80b, AA89, BN80, Blo89, Bor85, CADL86, CTV80, Csa80, CT81, Csa81, Csa82, CP85b, Das80, Das82, DG83, DH84, Dur86, FDAC89, Gás86b, GTB80, GBZ89, HF86, HSSW83, HBC86, Hic83, HS80b, HM89b, Ish83, Ish86, Ish88b, JJZ^{+}88a, JSS80, JA82, Kag83, Kal86, KS85, KK86, Kau82b, Kun86, LJW87, LBP84, LS86, Lib84, Loh84, Mar88c, NSS86, Nes87, NB79, NA83, NC87, Oga80, Par84, PGPRN86, Pyy86, Pyy93, Pyy00, Rud84, SS81a, Sen84, SF81a, ŠKR84, SBW86, STG86, Sza85, Var87, VGG88, WBK86]. ATP [MSM81]. attached [NH81, SBE^{+}83]. attachment [BMG82]. attack [CSB86a, VŻH82]. attraction [Wea88]. aufbau [Coo87]. Auger [Nes85, SD81]. August [ADL78, WJ82]. autoassociation [GLP80]. autocorrelation [GD82a]. autoionization [Chu86, ST84, Win85]. autoionizing [Gub82, NMK84]. Automatic [BSZ81, CMB84, Mar88c]. aux [VT85]. auxins [RP81b]. Average [CDG80b, BB85, BBO83, CS82, NSS86, Sen80a, VVD80b]. averaged [Zer89]. Avoided [BKP83, LC83a]. Avoiding [JT85]. axioms [Löw82b]. aza [PD87]. azabiphenyls [BLR86]. azines [BH89b]. aziridine [Sap80]. azobenzene [PD85].

B [LM81, LR89, Mar81a, Nat88, Zer87, NP80, AL87, BMMS81, CC82b, DG85, FCPL84, GP86b, Ish86, Ish88b, KR81, LCDP80, LP81a, LPP81, Nas83, PZ84, PG89, SRSR85, VC83]. B-DNA [CC82b, LCDP80, LP81a, LPP81, SRSR85, VC83]. B.V [Ågr87c]. B43 [Cal86b]. B57 [Cal82a]. BA [DBS82]. BaC [HDTR83]. backbone [HP81, HKG81, LJHU89]. bacteria [Sey81]. bacterial [Yad83]. bacteriorhodopsin [ES88, KPTG88, KPT88, San84]. balance [GT87b]. banana [PMT84]. Band [LTLF87, ABTP83, BS83a, Böh84, BDD81, BC84, BC85a, BC85b, Bur85, Chr84, CE88, DC87, FDAC89, HDTR83, KS87a, Koe86, KWBC80, Kun81, LS80, LK87, Mod81, Mor83, MB80b, Ori81b, Per85, Per86, RB88b, VVD80b, VI84, YRS85, LH87]. band-structure [KS87a]. banding [TT81]. bands [BW85, FL83, PSIT87, Sto82a, Sto82b, TP80, ZY85]. bare [MC88]. barrier [CPS89, ENH^{+}87, JRK84, KAH80, LMG85, Mül81, NBE87, RBL84, SSR^{+}87, SS86c, SS86d]. barriers [DC86, SB88, TWW81]. Bartlett [Ågr87a]. Base [HB82, BKP83, CLS81, CS83b, CCD^{+}84, CA86, GK80c, H^{+}90, HP81, KJL87, LGPP^{+}80, MGN82, OM82a, OM82b, OM83b, OM83a, OM84, OM85, OM86, OMH89, OYH^{+}80, Pic83b, PKP^{+}81, SS84a, SS85, Tew87, Tew88b]. based [Böh83, Böh84, DMB^{+}84, GV88, Gre81a, HOT87a, HOT87b, HR87, HPKW81, HL88, KKH^{+}83, KHK87, KWJ83, MC80, Mez84e, MRJ85a, MRJ85b, NSL88, Öhr85, Öhr86, OK80, PPB87, PLL86, PNGL83, RA81, RFST85, Roo80, RSMM86, SH89, SDS83, Ten86b, Was89]. Bases [GC87, AN86, Del88, DF89, DD89, LCCB81, PSS^{+}87, PKP^{+}81, SHK87b, SM80, SGP^{+}80, Tew88a, Voj85]. basicity [GC85a]. Basis [Ada81, Bec89, BYBB84, Del87, Fin87, FSP87, GG85, Ish84b, Kut84, Lin85, OI86, Pan85, STS83, SHKP80, SR87, Ågr87c, AM83a, AN85, AGCT87a, AGCT87b, Bag83b, BSHK86, BVC80, BPCT83, BS86, CFS86, CFS87, DPC82, DA81, FPT81, FEKC84, GHGP89, GT87b, Gou86, GM81c, GPC85, GSO^{+}87, GCV84, Har83a, HS89, HA84, HL83, INFU84, IO89, IBB85, IBB87, IQ87, IS88, KGEP89, KG89, KO81, LER87, LC81, MK85, McD80, Mor88, MNS85, ÖHD89, Pan83, PDMC80, PDC80, PDC81a, PDC81b, PDMC82, PKC85, RB88a, RB88d, RLPFA86, RR87b, RST87, RBR80, RVK80, RC80b, RS88, STH80, SR81, STS82, ŠKR84, SRHK83, SHK87a, SC80b, TS80, TP85, TH80, TPN86, WK83, WH87, YG86, CS83c, RB88a, RDG84b]. Basis-Set [OI86, BYBB84, Ish84b]. Basis-set-free [Bec89]. BaTiO [BMMS81]. bay [KSJT83, Low82d, Low82e, Sey83b]. bay-region [Low82d, Low82e]. BCL [TMO86, MPD87]. BCS [TF83a]. Be [WBK86, Ada82, AJ83, BN84, CS83c, GY87, Kal85, KEGS81, LB83, Mor88]. be-isoelectronic [AJ83]. BE2 [CPZS86, SBH^{+}85, SBH^{+}86]. BeAl [CH82]. beam [Fut87]. BeBH [BR80]. Beck [Gos80]. become [Mis84]. BeH [LB83, GG82a, PSBB83]. Behavior [Ada80b, MS80b, BLR86, BB81a, BMBK82, BV80, DK82a, KK83, LOGC86, OYN80, PFC88, RGW88, Sla85, VGG88, ZAB85]. Belgium [ADL78]. benchmark [SBB87]. Bender [Avr82]. bending [GUSS89]. Benjamin [Lun82]. Benjamin/Cummings [Lun82]. benz [PZSL83]. Benzene [KNS^{+}86, EB82b, PČP82, TOK80, WK83, Živ86, Živ87c, SS84b]. Benzene-containing [KNS^{+}86]. benzenes [Ågr83, Tak86]. benzenoid [BK84b, Cio87d, EB81, EBEE87, KSJT83, Ran80a, EB82b]. Benzinger [Chu88]. benzo [AEHNEA85, PZPL86, SG80a]. benzo-derivatives [AEHNEA85]. Benzodiazepin [KRST83]. benzodiazepine [KTGR89, VL89, YGK^{+}89]. benzodiazepines [YGK^{+}89]. benzohydroxamic [AEHM85]. benzoic [BK84b, MG89]. benzomorphan [CZM83]. Benzophenone [KNSM88]. benzoquinone [CS83b]. benzosemiquinone [SSUF80, SKS83, SS86b, SS86a, SL80]. Berlin [Ågr88, Fro81, Fro83a, GSC89]. Bernd [Ano81d]. Berthier [Ågr88, Fro83a]. beryllium [BSR80, CP85b, EB87, MBR80, Ori81a, Ori81b, Ori82c, Ori82a]. betacarboline [KTGR89]. betacarbolines [YGK^{+}89]. Between [BCV80, ACP84, AN86, AGB82, AM89, BZA84, Bar81, BSL81, BMMS81, BJR86, CLA80a, CL80, Cof83, DF86, DCF82a, EY82, Fis84, Fis82, FSP87, GSM83, GJMLRGSA85, GP85, KAD81, KPP85, KPP86, LCCB81, LS86, LC80, LWZ88, LZY83, Löw84, Low82e, LLM86, MZ85, MJ89, MBSC81, MDT85, MGBH81, MA80b, Nag82a, NL89, NSL88, Odd82, Par84, Pie84, Pie85, POK85, PO85, RM87, SS82, SFC86, Sar89, Sch89, SVS81, Ska84, Smi89, Sti82, TPN86, UB86, VBS85, Wes81, Wes82, WBK86, Wor85, YHG89, YS80a, ZA89]. beyond [PHGR88, Sha85, Tay87b, Zho88, RBH^{+}86]. beyond-Hartree [Tay87b]. BFGS [Kar84b]. BH [BLR81, IS83, WKH80]. Bibliography [Ågr88, OM83a, H^{+}90, OM82a, OM82b, OM83b, OM84, OM85, OM86, OMH89, Pic83b, Pyy86, Pyy93, Pyy00]. Bicine [SK81]. bifluorenylidene [SAP81]. bifurans [AEHH81]. bilayer [NTM80]. bimolecular [BCT86, Chr80a, Chr80b]. binary [CF81b, EB86, Lim82, SBWW80, TP87]. Binding [CPMP86, Cia81a, Del81a, JG81, PS80a, SBWW80, AHB^{+}80, ABD81, BD81, BKS87, Böh84, Bri86, CGP87, CZM83, CMM83, Csa85, Csa87, CE88, Csa89, Das82, Fra84, FCPL84, GM81a, GEdP81, HH89, HKFW87, Jac89, KEIG86, KMRK86, Laf81, LCP81, Luk81, MMM81, Mod81, PZSL83, PZPL86, RB88d, RWOT81, SG81a, SAB^{+}80c, SS80f, Sto82a, Sto82b, SRDG85, VL89, Zho88]. bindings [Sin80b]. binuclear [Böh83, BDK86, CEZ83]. bioanode [PSG83]. biochemical [FCPL84, HZL84, PSG83, PP80]. Biochemistry [Mar81a, PTG80, Khe81, LM81, Pul79, LM81]. Biological [DS80c, SRO81, SS80a, SG82, AL80, Ave84a, Bro88, CTP83, Chi81, DS89a, DA80, Dos84, Frö83b, GSMSG82, HJ80, KCP80, Lar82, Mår80, Mår82, NS86, Pet79, PL84a, Sar89, Sey81, SS81b, Ste84, Tap88, The84, TR80, XYSJ81]. biologically [GSG84, GM81c, KTK80, KRST83, KRBM85, KNS^{+}86, KKSM87, KNSM88, KMPTK89, Pin88]. biology [Dav87, FHHHE85, Kar80, Lan84]. biomolecular [SRSR85]. biomolecules [Blo83a, Mez87a, RMM83, RTM86, RC80a, Rom81, SRH86, ZA89]. biophysical [Moo81]. biophysics [Khe81, Lad87]. biopolymer [Sin80b]. biopolymers [BL81b, CP80b, LOBS86, RSP^{+}80]. biosciences [CCC^{+}89]. Biquadratic [Ori80, Ori82b]. bis [XHR86]. bisantrene [CGP87]. bivariational [LFM89b, Wei82]. Bloch [BOW80, GM83]. block [FF80]. blocking [DS80b, PKH^{+}81]. BMV [BVD81]. BO [ANH81]. body [Ade87, Ano81s, AZY84, Ave86, ALH86, BK89, Ble82, CD87a, CC83, CC84b, DB81, Fla80, FH81, FM86, HBSC83, HB86, Hub80, IS85, IS80, Kal85, KCK83, KS82, Kha85, KB88, LB83, LC88a, MK85, Mat87c, Mat87d, McD81b, MJ89, Muk86, MP84b, Muk84, MPD87, NCKL81, Nov83a, PS80b, PM84b, SF81a, SH87b, UI80, Var87, WJP83, WJP85, Win85]. Boer [RM82]. Bohm [Bel80b, Nag85]. Bohr [Har87]. Boltzmann [Ulm85]. Bonchev [Ler85]. Bond [BVD81, EB86, GMM81, May84, May85, May86a, PC83a, PCG83, PT84, PP83, TP84, TP85, TWW81, ZLH87, ÅFL81, Ber81, Bes88, BF83a, BBL81, CPN89, CCD^{+}84, CL83b, CZ82b, CLZ84, Cul88, Cul89, Del87, DS89b, DF83a, Eng83a, Eng83b, FGM87, Fli84, FEKC84, FH83b, GMM84, Her83, Hib81, HO85a, HO85b, HG86, Hur82, Jul84, Ker81, Laf81, LZ87, LW84b, LS82, LGL^{+}84, MS87, MY81, MGN82, May86b, McW88, MK83, NÁNSP87, NP80, OLB^{+}82, OYN80, PNGL83, RB81, RS84, RKHK88, RKO^{+}89, SS85, SHH^{+}83a, SRT81, SWM82, SNSM82, SI89b, TF83b, TS81, TH81b, Ver80, VMMV80, Wed81, Živ86, Živ87c]. bond-effective [SWM82].

Bonded [LZ89, AGCT87a, AGCT87b, CLS81, Cze87, DDFA87, GB85b, GBW88, GBW89, GCP82, IHI80, Kle83, LCCB81, LP80, MZ85, Mil86, MC89, MB82, Nov83a, RSBR85, Sch80a, Sch81, SD88, Wor85]. Bonding [CKM81, PD84a, ABD81, BP84, Bau86, BS83a, Boč89, BDK86, BLOW89, CJZ^{+}86, Cia81b, CEZ83, CF81b, Del88, DF83a, EIKG^{+}88, FFF88, FD89b, GMS81, GP86b, HC80, JM81, KBS81, KAH80, Kin86, KD86, Lal81, MH82, PGL^{+}86, PMCA88, PP81, PG89, SRS81, SFC86, SS84a, ST86a, SWM82, TP80, TWL82, XHR86, ZSB80, Hil85]. bonds [BLR81, BDP^{+}81, DF86, GBO86b, KEIG86, LW84b, PMT84, SD86, Ser81, Yam81]. Book [Ågr80, Ågr82, Ågr85, Ågr87a, Ågr87b, Ågr87c, Ågr88, Bra86b, Brä86a, Cal81a, Cal81c, Cal81b, Cal82a, Cal82b, Cal86b, Cal86a, Cal86c, Cal86d, Cal88, Cal89a, Dun80, Fro81, Fro83a, Fro88a, Fro88b, Gee89, Gos80, Kum87, LM81, Ler85, Lun82, Mår80, Mår82, Mar81a, Mic87, Mou89, Muj86, Nat86, Nat88, Pic83b, Pic83a, Rat87, Rat88, Rös87, Sie81, Sie83a, Sme84, Tap86, Tap88, Wol81, Zer87]. Books [Ano80d]. boranes [TL86]. Borel [YS80a]. born [HRL88, Nag85, Röm83]. boron [DPR80, LL86, LÖ85b, MG83, YZ85]. Bose [CTP83, MB83, OT80]. bosons [MBSC81]. Boston [Cal81c, Cal81b]. both [RL88, RFST85]. bound [BN84, Csa86a, ENH^{+}87, FAC84, GDBRV87, Lev89, LO81, MSM81, NSS86, NBE87, Suc84]. bound-state [FAC84]. boundaries [Ker81, Mez83d]. boundary [FC81b, FC81c, FC81d, FC82c, Swe83]. bounded [AMFC83, Swe82]. bounding [Cas88]. Bounds [MPPW82, BH86, ČV86, CV83b, Gol80]. bovine [Gas80, LSS83]. box [LAGS89, TS82]. Boyd [LWK84]. Br [PXLC86]. brain [Koh86]. brain-like [Koh86]. branching [BS82a, Rut81]. Brandow [SH87b]. Bravais [RB86]. Breakdown [Nag85, GC83]. breaking [Böh83, Cow82, RKHK88, RKO^{+}89, Tru83a]. Breit [BG88, FWS^{+}85]. bridge [CGK82, FHHE80, HR84, MR87, San84, SD86, TTRL88]. bridged [BR80, MBT81, OKZ85, VG80]. brief [LL89, Löw85b]. Bright [Ano80e]. Brillouin [Ori84a, RB86, MW87, Ori80, Ori82b, Ori85a, Ori85b]. Brillouin-zone [Ori84a, RB86]. broadening [LCCW87]. Broglie [Bel80b]. broken [BBH87, Gos85, HM89a, Lef83, Pra83]. BSSE [MSV89, MS89]. BSSE-free [MSV89]. buffalo [Fin89]. build [Pin88]. build-up [Pin88]. bulk [CB88]. bundles [Cho80]. bungarotoxin [SSR86]. buried [NS83]. Butadiene [Živ86, Živ87c, BKP83]. Buthus [Gri81].

C [Ågr82, BLR81, Cal81c, Cal86b, Fro88a, KBT80, Mas87a, Mic87, RKO^{+}89, Sie83a, SHP^{+}80, SS87, LCA89, MPP^{+}81, Ano87s, CS83b, COS82, Deá83, FGM87, GBO86a, Ish86, Ish88b, LWZ88, LH85, LCCW87, Mak80, MMB84, PDMC80, PDC80, PSBB83, SRDG85, TST89, VŻH82]. C-state [COS82]. C. [ČD87c, Hal88]. C3 [KTGR89]. C3-substituted [KTGR89]. CA [KD80b, GP83a]. cadmium [Das82]. caffeine [DS89a]. calcium [BL81c, HKFW87, SRDG85]. Calcul [OF80, SvG89]. calculate [DVV80]. Calculated [HLR80, LHZ80, PP82, WCLZ86, BYBB84, FK80, Luk81, OL82, Ran80d, Ver80]. calculating [JSS80, Mül81]. Calculation [BL89b, Das82, DCF82a, DCF82b, FVZ^{+}88, FAC84, FGS85, Fuk83, GOS87a, Ish83, KS86b, KM89b, LC82a, LL84, LW82, LW84a, LW84b, MTLL82, Mal86, Mod81, RP81a, RTM81, RTM83, RMM83, RTM86, SMD83, SB83a, Ská83, VU85, WKH80, Wen86, XYSJ81, YS80b, YS80c, Ada81, AB87, AJY80, BS82a, BR81a, BN80, Blo89, BYBZB83, Bou80, BFOP83, Bur85, CLM80, Che80, CL86, CA86, Csa83, Csa89, DS80a, DP87, DG86, DG83, DBVMB85, Don82, Don85, DS81c, EÖWZ84, Ess84, FP84, FC80a, FC81a, FC85, HK86, HTD80, HHL80, HC81, Hir80b, HYZ89, Hub80, IBB87, IS80, INO88, JN88, JJČ87, Kal86, KO81, Kau82a, KE82, LM82a, LdMH80, LHC80, Löw85a, Mes85, MMI^{+}82, MV87, MSI80, NKM81, OF80, PPB87, PZ84, PABD80, PS84, RNW86]. calculation [RL86, RL88, Rér89, RM84, Sad89, ST84, SS84a, Sel86a, Sel86b, SG87, SP82b, Sve88, TPN86, TL82, TH81b, WSD80, WHT86, Yon85, GP86a]. Calculations [ADF^{+}88, CMM83, Fro81, SA80, Tat81b, ÅCFRJ85, AEHA87, AB85, AEM86, ACP84, AM82, AM87a, ARZ^{+}89, AØ80a, AØ80b, AØM81, AØC82, Bal81a, Bal88, BB82b, BG81, BSHK86, BVD81, Bar86, BIN86, BG88, BD88, Bha81b, Bha82c, BS83a, BBPS83, Blo83b, BVB88, BVC80, BC80, BDK86, BHVR83a, BHVR83b, BH89b, BCC82, Bro88, BVAB81, Bue86, Bur83, ČU80, CY80, CHKB83, Cha80, CPGP83, Chr84, CR81, CFS86, CAML87, CHZ82, CF81b, DMB^{+}84, DPC82, DB81, DL82, Deá83, DA80, DFP80, Del84b, DFH84, DH85, DC87, DOW87, DPR80, DC86, EG84, FC81a, Fli86, FSK79, FSS^{+}89, FDAC89, GT86, GG81, GG82b, GG82a, GG83, Gás86b, GPLP85, GAL86, GLLOB86b, GLLOB88, GT87b, Gos85, GM81c, GK80c, GSS85b, GCV84, GBO86b, HPK81, HKK^{+}82, HLKK82]. calculations [HKKM83, HKLM85, Har80, HM85a, HKV85, Ho81, HAB83, H^{+}90, HAmGS86, INFU84, Ish86, Ish88b, Ish84b, IBB85, IQ87, IS88, IAAA89, JWÖ83, JŻ85, Kag83, KSO84, KG89, Kar81, KHNT86, KT80a, KHP81, KHCH85, KHRK87, KHK87, Kau87, KU80, KKA80c, KNM83, KST84, KD80b, KEB86, LSP85, LCULC87, LRT88, LJK86, Lás82, LT81, LGS87, LMG85, LHKK82, Lew88, LKB89, Lib84, LJ82, LC88b, LC88a, Lin89, LSBS81, LHP80, LN89a, LHK81, LP87, LÖ85b, LN89b, MS87, MM87a, Mar89, MMS^{+}86, McW84, Mes80, MW87, MMC83, Mül82, ME81, Nas83, NS84, NSL88, Noa80a, Noa80b, NKV^{+}89, OM82a, OM82b, OM83b, OM83a, OM84, OM85, OM86, OMH89, Ort89, PWL89, PL80, PLL86, PNGL83, PMCA88, PR83, PBFL80, Pay82, Pic83b, PD80, Pit84, PDMC80, PDC80]. calculations [PDC81a, PDC81b, PDMC82, PDKC84, PKC85, PM81, RS86, RT84, RT85, RBH^{+}86, RKHK88, RKO^{+}89, RSMM86, STH80, SD80, SR81, SRS81, SS83, SR84b, ST80b, SC86b, SI89a, SBB87, SBE^{+}83, SF81a, SNNG86, SC89, STS82, SMS84, ST82, SHL82, SHL86, SHP^{+}80, SRL^{+}83, SHK87a, SS86c, SS86d, SS87, SC80b, SM80, STG86, SM87b, SSF^{+}84, ST86c, SSN^{+}84, TS81, Tay87b, TS82, Tew88a, Tho85, Tho80, TN80, Tin80, THKM81, TMO86, TL80, VARPU84, VP84, Voj85, Vol89b, WKJ88, Wes82, Win83, Win86, WS80b, ZW83, ZJŻ88, GG82b, LFL^{+}81, RDH84, SY82, TW88, WK83, Yad83, Ågr87c, Sie83a]. calmodulin [MR89]. Cambridge [Ano87t]. Can [Ada82, Kal85]. cancellation [Hub80]. cancer [JL87, SG80b, SG82, SG85b]. Candidate [LPCF80]. Canonical [Mat84, KN84, Leu83, Per82, GBW88]. capacity [MV87]. Capellos [Fro88a]. capture [PGPRN86]. carbanions [BKJ88, Mof82]. carbene [JM83]. carbides [PT89a]. carbohydrates [JM81, KVR85]. carbolines [MSC^{+}83b]. carbon [CPN89, DGSO85, KL80, KGB87, LCF89, LS86, SKIK82, TKY89, VMMV80, Ran83]. Carbon-13 [Ran83]. carbon-hydrogen [CPN89]. carbonate [BLOW89]. carbonic [CLA83b, GP80, KR81, KBSG81, LL89]. carbonium [RT87]. carbonyl [Mar89]. carbonylheme [HLR80]. carbonyls [DMB86, GSMSG82]. carboxylate [SR84b]. carboxylic [DBC84]. carboxymethyl [SRS81]. carboxypeptidase [VWCC82]. carcinogenesis [Lad86, LOBS86, Löw86b, PP80, RT84, RT85, Sey80, SL87, ST82, TR86, Tho89, Var86, PTG80, Mar81a]. carcinogenic [DD89, HKP81, KSJT83, LP84, LP81b, LPSK83, HPK81]. carcinogenicities [LPP84]. Carcinogenicity [Gre82, Low82d, Low82e, MP87, SG80a, SVS81, Sey83b, Sey83c]. carcinogens [BMG82, LČ84, LPCF80]. cardiac [DS80b]. Carlo [CC83, CFC83, CC84b, DKLP81, DS89a, DBF89, LG84, MB86, Mcd81a, MMB84, MK81, RBL84, RBH^{+}86, RC80a, Rom81, RJK83, Umr89, VR86]. carriers [LOGC86]. Cársky [Fro81]. case [CF81a, DS88, Eng83a, FD86, GPLP85, KG85, KG86c, MJ89, Ska84, Sri86, VCV87]. cases [Cia81a, FN84, SRL^{+}83, Yon85]. CASSCF [ANH81, Bal88, KHCH85]. CASSCF/CI [Bal88]. catacondensed [EB82a]. catalase [AWCL89, LCCW87]. catalysis [Cia81a, CLA83b, FG82, Jac89, Sim86, Sla83a, Sla84, TF80, Pul79, LM81]. catalysts [Sok81]. Catalytic [SSC83, BBPS83, NS82a, NS83, Sok81, ST88]. catalyzed [CD89b, CD89c, LL89, OLB^{+}82, SRC84]. cataractogenesis [Var81a]. catechol [SS80e]. catechol-O-methyltransferase [SS80e]. Catecholamine [ŠKH81, SS80a]. cathedrals [Urr82]. cathode [LK81]. Cation [BKS87, GPC85, RSP^{+}80, Big86, BL81c, CKM84, GEdP81, MR80c, OL80, UTP82, BL81c]. cationic [KHK87, RJ88, THKM81]. cations [Bre80, GMM84, GP85, LCDP80, MBR80, RGW88, WBK86]. caused [HG86, SWW89]. CBO [GC83, GC84f]. CC [Yam81]. CC1 [SBH^{+}85, SBH^{+}86]. CCI [KG86a]. CCl [KG86b]. CCSD [SSR^{+}87]. CD [KM80]. CdS [CCS84, Fle81, KWBC80]. CdTe [HLPV83]. cell [DAPS81, Lad86, LS85, SL81, SAB^{+}80c]. cell-selfregulation [LS85]. cells [BS83b, GMP^{+}84a, Gup80, PSG83, Poh80, Poh83, PPR84, Poh84]. Cellular [Poh83, PPP87, FL83, FD86, Gas85, KFD85, LFL^{+}81, LTLF87, Poh82]. center [BBR82, Cof83, CE88, CLZ84, GKMV80, Gus87, Hur82, Jon81b, Jon82, Jon83, Jon84, Jon86, KK82b, KK83, LCULC87, MPB84, PSM87, RLPFA86, ST80a, Tin80, Wea88]. centers [CB83, LPP81, MBKH83, PC83b, RLR88]. Centrifugal [Loh87]. century [Urr82]. cephalosporin [PBBK89]. cerium [LFM89a]. certain [JL87, LC83a]. CF [CF81b]. CFP [MPBN81]. CH [CF81b, INFU84, KKA80b, Wój86a, Fro88a, Wój86b, AYBP83, CHKB83, ESKG83, GA82, HL80, Ish86, Ish88b, KHCH85, KTS86, PDC81b, RS86, SRL^{+}83, TST89]. Chain [LBV86, BLR83, FY82, FDAC89, Kir89, LJHU89, MW83, Nag82b, Pin88, SKL80, SRC84, SNSM82, TNY83, Yar82]. chains [ABD^{+}80b, ABD81, ABTP83, BSL81, BS80c, DL82, Del84b, DFH84, DH85, DC87, FDA83, FM86, Her84, HJ80, Kib83, Kve82, MW87, Min87, PS81, Ran80c, SRC83, Tus86, Vol88]. Chairman [Löw89]. Chairman’S [Gás86b]. challenge [Cal82a, Woo80]. Challenges [Ben81]. change [DS81a, DQT80, DQ80, DQ81, SH89, ZV85]. Changes [Dat82, GMP^{+}84a, Lim82, Sen80a, VU85]. channel [Lin85, UVP^{+}81, UTP82, UTVP85, VARPU84]. channels [Ovc81, Var81b]. chaos [Tay87a]. character [MGN82]. Characteristic [ACH83, BL88, Dem83, KT86, KJT85, Kry82b, MKVP80, RBK85]. characteristics [Hil85, KTGR89, SS80a, TK86]. Characterization [FH81, FH83b, AM87b, AM88a, BT84, BMT80, Bon83, GYJ83, Ler85, MBT81, OMC82, PBM85, PBA88, RL86, Ran81, Ran88]. Charge [CC84a, CC82a, FZB86, Fla80, Fli84, Fli86, Has82a, Ish89, MGN82, MDO^{+}82, MBD84, Ågr83, Bec82, BQS87, CEB86, CL80, Cla83a, Csa83, ECM89, Fut87, GSMRM85, GCP82, GCP86, Has82b, Hil81b, HDTR83, Ish86, Ish88b, JN88, JBW87, LS80, LCCB81, LT81, Mez81a, Mez82, MB85, NW82, OL80, Par84, PNA86, RBSR84, RSBR85, RR87a, RR88, SH89, ST80a, SKEDR81, SST83, Suh84b, YZ85, Hil81a, PDC81a]. charge-mass [SH89]. Charge-transfer [MGN82, Ågr83, GCP82, GCP86, NW82, YZ85]. charged [Dem83, PDC81b, Rab84, RGW88, SMHK89]. charges [CB82, Fli86, Mez86b, YHG89]. Charles [Sme84]. Chatelier [Mez84b]. CHCH [LS86]. Cheetham [Rat88]. chelates [FHHE80]. chelating [SS80f]. Chemica [Zer87]. Chemical [BT82, BWEW89, Cal89a, FD89b, Fle86, Gro85, Kle86, KT80b, KLDD84, May83, MBT80, PD83, RM84, ST83, Ade87, AM89, Bal82b, Bal82c, Bal82d, BZA84, BB83, BD81, BLR81, BTK83, BT84, BW80, Ber81, BDP^{+}81, BS83a, Bon83, Cio88a, ČPV81, Cla86, CA86, CAML87, DD87b, DC86, FFF88, FHHE80, Fli84, Fli86, FGG83, Fro88a, Fuk81a, Fuk83, Fuk81b, GTB80, GF83, GHF84, GJMLRGSA85, GSS85b, HLKK82, HKLM85, HC80, JL84, Jul84, KPH88, KKZ86, Kau81, KJMS85, Kot82, KJT85, KPSK80, Lar86b, Ler83, Ler85, LHKK82, LP81b, MSC^{+}83a, MSC^{+}83b, May86b, Mez83b, Mez84a, Mez84b, Mez84e, Mez85b, Mic87, MMI^{+}82, NHKY83, Nat88, PNGL83, PSG83, PP80, RK82, Ran83, RD84, RC86, RT84, RT85, RKHK88, RKO^{+}89, ST89, SZ84, Ser84, Sey80, ST86a]. chemical [SBE^{+}83, STS82, SMS84, Sin88, Sla83a, Sla84, Sla85, SPSG85, SMKT86, Tac87, Tew88a, Tho89, TKR86, Urr88, Var86, Var85, Vas82, VGG88, VMMV80, Vog86, XHR86, Yad83, Rat87]. chemically [SMVV86b, SMVV86a]. chemisorbed [MB85, NA83]. Chemisorption [AS86, SAS86, BASB86, FS86c, GPN80, IRP81, JR85, Ker81, RSGK88, SR84a, See81, Wed81]. chemisorptive [TH81b]. Chemistry [Ågr88, Cal81b, Cal86a, Cal89a, Fro81, Fro83a, KS81, Mar81a, PTG80, Ågr85, Ågr87a, Ada64, Ada69, Ada80a, Bar85, Bec89, Ben81, Ber81, Cal82c, ČU80, CD92, CB85, CCG^{+}82, CCD^{+}84, CCC^{+}89, CP86b, De 80, DD87b, Dun87, FC83, HZ84, Hay82, HM81, HZ83, H^{+}90, HYS89, Jor89, Kel78, Kib83, KST84, Kou80, KPS87a, KPS87b, KPS88, Laf81, LM81, Law87, Löw86c, Löw89, MH80b, MH89, Mal91, Mal02, Mas87a, MN81, Mat87b, Mat87c, NKZ88, NS86, Nat88, NB79, OM82a, OM82b, OM83b, OM83a, OM84, OM85, OM86, OMH89, OYH^{+}80, Pau89b, PSK86, Pic83b, PBKB89, Pul79, Ran80b, RZ86, Rat88, RBH^{+}86, Sad85, San81, Sey83a, Sim86, Tap88, Tri86, Zer87, Fro81, Brä86a, Dun80, Gos80, Lun82, Muj86, Wol81]. chemists [Kun83]. chemodynamics [FTY81]. chemomechanical [Urr88]. Chen [Ban82a]. Chen-Ning [Ban82a]. CHF [EC83, FC81a]. Chichester [Cal88, Nat86]. Chimie [Cal81a]. Chiral [CT87, ST89]. Chiral-achiral [CT87]. chirality [SS80a]. chiroptical [BHVR83a, BHVR83b]. chloride [GjC86, Hil85, LCC84, LPP84, LP84, MP87, PP82]. chlorinated [PH82]. chlorine [LdMH80, RS82]. chlorooxirane [LP84]. chlorophyll [EZ83, KHNT86, OMC82]. choice [Bel80b, MM87b]. Cholesky [OS89b]. Cholesky-like [OS89b]. Cholesterol [Lal81]. cholinergic [WMP^{+}83]. choosing [RR87a, RR88]. chromatography [PH87]. chromium [FBM80, SG89, MK80]. chromium-group [SG89]. chromophore [San84, SG81b]. chromophores [HMF82, NTM80]. chromosome [TT81]. chymotrypsin [SKEDR81, SNSM82, VB85]. CI [Big86, GK80c, Yam80b, Bal88, BL81a, BEO80, BEP87a, BEP87b, BBPS83, BBO83, BC83, Bue86, Bur83, CHKB83, CR81, CP80a, Has82a, HZBR89, HC81, KGEP89, KHCH85, KHRK87, KHK87, KS81, LCLC86, LCULC87, LCLC88, LJ82, MR81, MR80b, NH81, Nak83, Nas83, RW80b, RW81, RJK83, Roo80, RS80, RKHK88, RKO^{+}89, SP86a, SNNG86, SG89, Sie83b, WBK86, Yam80a]. circulant [CP85a]. cis [KL80, RB81, EIMS87, HET86]. cis-trans} [KL80]. cis/trans} [RB81]. CISDT [SSR^{+}87]. CISDTQ [SSR^{+}87]. Cl [KKTE80, PXLC86, CV82a, DS81b, KG89, KS89b, LJC83, SJ85, STG86, Wój86a, Wój86b]. Clarendon [Ågr82, Ågr85, Gee89, Rat88, Rös87]. Clary [Mic87]. class [Gra82, Kat89, Ojh89, PSIT87, TH88]. classes [FEKC84]. Classical [DBF89, Hal80, HC80, SHL86, Tho83, ACP82b, Avr82, BM89, BQS87, Dun87, KB87a, Rec80, RBL84, SHH^{+}83a, Sil84, SM87b]. Classification [Kry82a, Fra85, FEKC84, Kva82, Mar88c]. Claverie [Cze87]. clay [LGL^{+}84]. Cleantes [Gos80]. cleavage [Yam81]. Clebsch [Alt89, Her84, Kib83]. ClF [CV82a, DS81b]. Clifford [GCP87, LP89]. ClN [LJC83]. ClNCO [LJC83]. close [DBVMB85, PABL88]. close-coupling [DBVMB85]. Closed [PSM87, DG83, KC83, KS85, MT85, Mal82, MC88, PL80, WS80b]. closed-shell [DG83, KC83, KS85, Mal82, PL80, WS80b]. closest [MT87]. closure [BBGS84]. cloth [LM81, Mår80]. clupeine [PAFW84]. Cluster [DL82, GLP83, NH81, Nak83, PB82, PR83, Van82, ALB84, AB85, AB87, AS86, BS81a, BANF86, Dun88, Fin85, GOS87a, Gee88, HK86, HB86, HLPV83, JR85, Kal86, Kin86, KSGB^{+}83, Kle86, KS89a, Kva83, LPB82, Lau88, LLAG89, LSR^{+}84, MS88, MJW88, Muk86, PTC84, PP89, PSBB83, RSGK88, RSMM86, SSR^{+}87, See84, See88, SB84, SB87, SAS86, SM84, TP85, TH81b, HHL80, Küm83, MB85]. Clustered [CB83]. clusters [AH86, BL88, BASB86, Bor81, BCB85, BAN88b, CJZ^{+}86, CP85b, DBS80, Ell84, EG84, FS86c, GPN80, GVVDB89, GLLOB86a, KBS81, Kau87, Kin86, KJL87, KGB87, LBC^{+}87, LNN84, Loh84, Mar81b, MBSC81, MTN83, PP89, RJ88, SR84a, SOS84a, SOS84b, SJ85, See88, Fro83a]. CN [MA80b]. CNDO [BH82, Boč87, Bou80, CV82a, Che80, DS81b, DBS82, Deá83, FF80, GjC86, GRP^{+}80, KN80, Lás82, Lip83, LWK84, MBKH83, MH82, RB88a, RM84, SMD83]. CNDO/1 [Boč87]. CNDO/2 [CV82a, DS81b, DBS82, KN80, Lás82, Lip83, MBKH83, MH82, SMD83]. CNDO/2- [SMD83]. CNDO/S [Big86, SMD83]. CNDO/SW [GRP^{+}80]. CNO [RKHK88]. CO [SOS84b, Gos80, AS86, BF83a, BBL81, CD87a, CL83b, DF89, DFH84, Don85, HDFL84, JTST87, JR85, KU80, LL89, MB85, RW80b, RW81, Sok81, SSC83, SRL^{+}83, ZW83, MK83]. CO/Cu [MB85]. coal [GMP84b]. Code [FM80, Bed84, FF82, JKL85]. codes [EBEE87, FF84]. coding [FF84]. coefficients [Alt89, ACCYCC80, AØ80b, BD85, EY82, GP86a, GM81b, Her84, Kar84b, Kib83, LSH88, Löw83, Pau89a, Wen86, ZS85]. coenzyme [PSG83]. coenzyme-immobilized [PSG83]. coexistence [BB81b, Kve82]. coexisting [Sti82]. CoF [MTO^{+}81, MOT^{+}83, MK83]. cognate [Shi85, Yon85]. Cohen [DF83a]. Cohen-type [DF83a]. coherence [Van85, Van85]. Coherent [BSD82, CD89a, CFG^{+}83, Frö83b, GM83, Ská83, KSGB^{+}83]. coil [MT85, MY81]. coincide [GBZ86]. coli [FPDA89, JG81, MLC^{+}81]. collagen [VCB83]. collagenic [CBV85]. collapse [Kut84]. Colle [CL85]. collective [Kum87, Sew86, Toy80]. College [Pic83a]. collinear [KS89b, Nal80, NP81]. Collision [Chr80a, Chr80b, FR84, Ano80c, Gir86, Ols83, PL89, ZJŻ88]. Collisional [GPLP85, Hic83, Mic85, Mic86, SM87a, YUKK84, CMDBVMA89, Mic81a, Mic83a, YKZ^{+}89]. collisions [All87, CPS81, Del83, Dun87, GM86a, Gir80, GSR83, KH85, Köh89, Mcc83, Mic83b, Pia83, RB89, SG85a, Tru83b, VM81, VM83, WB81]. colloquium [Vas82]. colored [Cox87]. COLUMBUS [SSP^{+}88]. combination [Dov86, HAB83]. Combinatorial [Pau81]. combined [Sav88]. Combining [Bue86]. combustion [GSR83]. Comment [Mak80, Swe82, Hal88, SL80]. Comments [Ano86c, BB83, Ber81, CTP83, Hub80, LSZ83a, Löw86b, Tus86, CF81a, FC82a, FC82e, LSZ83d, LSZ83b, LSZ83c, Löw86e, NBE87]. commonly [Kun83]. commutative [Eng83b, Eng83a]. Compact [STH80, TH80, POT80]. Company [Ågr80, Ågr87a, Cal86a, Dun80, Mic87, Ågr87b]. comparability [EBEE87]. Comparative [DS81b, Luk81, Wag83, AWCL89, CV82a, EIKG^{+}88, Kar84b, SKEDR81, SMD83]. Comparison [Bar85, CL80, Csa88, DS89b, DWQ86, LC80, SKER83, SHK87a, YHG89, FD89b, LC83b, LC88a, MR89, MJ89, Mez85a, SSR^{+}87, See84, Tel88, VBS85, BB82b, BB85, DQ81, GY87, Hal80, LH85, PP89, SH87b, YKZ^{+}89, Ågr87a]. compendium [Ågr87c, PKC85]. competition [MB84]. competitive [DS80b, Shu84]. complementarity [NSN89]. complementary [CLS81, CS83b, PKP^{+}81, Živ88]. complete [Roo80, SF80b]. Complex [BS82a, Edw88, KEB86, LGS87, MGBH81, REB83a, AL87, BF83b, Brä86c, Chu86, CCG^{+}82, Cof83, CH82, EKBE86, FSW87, Jun80, LdMH80, LG84, LLAG89, LHZ80, Löw86e, LFM89b, LP80, Moi81, MBR80, NKM81, NMK84, Ori84b, RM82, Rud84, SMSA80, SAP81, SBB87, Sim80, Sla83a, Sla84, SHP^{+}80, ŠHH83b, SR80, VTT81, VŻH82, Win86, Wój86b, Yon85]. complexed [GM81a]. complexes [AAW82b, Bha82c, Boč88a, BPCT83, BC80, Chi81, CEZ83, Del87, DF89, FHHHE85, Gas80, GAL86, GP83a, HLR80, Hil85, KN80, KS87b, Lip88, LNN84, MDO^{+}82, MTO^{+}81, MOT^{+}83, MK83, MK80, NSN89, NW82, Pee83, PS80a, RK82, RD84, RBSR84, ST89, SJ88, Sla83b, WCLZ86, WK83, WH87, Wój86a, YZ85]. complexity [Sey81, Sil86]. complexons [KKD80]. Compliant [MR85]. component [Kah84, YL83]. components [BCT86, DPZ80, JKL85, RS88, SR87, SHK87a]. composed [Sie80, ZLH87]. composés [OG86, VT85]. Composite [Sin86, Gir80, Gou86, Mat81b, Sto80, VT85]. composites [OG86]. composition [DBS82, SGP^{+}80]. compositions [FW84]. compound [DT83]. compounds [BSHK86, BDP^{+}81, BH82, Böh82, Böh83, BJR86, CD92, CPGP83, CNK87, DGSO85, FZZ87, HKK^{+}82, HKKM83, HKLM85, JN88, KM80, LCF89, LmJQHXZ83, MG83, Mar89, MBT81, MFM89, PDMC82, PDKC84, SA80, SY82, SRP81, SCER88, SSF^{+}84, TR82, THKM81, TL83, VMMV80, WBM88]. compressed [CTV80, Csa80, CT81, Csa82]. compression [FNC86]. Compton [AM83b, BLOW89, CBC^{+}89, DT81, PGM80, Smi89, Tha83]. computation [BAG83a, BEO80, ČVV87, ČV88, Duc85, KJT85, PD84b]. Computational [Kau82a, ST89, AM87a, Ano81s, Brä86a, CB85, CCG^{+}82, CP86b, DFH84, DH85, DR84, JI88, KHP81, Kau87, MP87, PP84, RDH84, Tri86]. Computations [GCP82, GPC85, KS87a, Kun83, SD86]. computed [Del87, MHAS86]. Computer [ACP82b, GD82a, Jon80, Jon81b, Jon81a, Jon83, Jon85, Jon86, KSJT83, SG86, Bri84, FP84, GSS85b, Jon84, Kau81, Mez84e, NS89, SMS84]. computer-aided [FP84, NS89]. computer-based [Mez84e]. Computer-generated [Jon80, Jon81b, Jon81a, Jon83, Jon85, Jon86, Kau81]. Computers [KST83, Cla80b, Kun83]. computing [Con88b, MKVP80, Pin88]. concentrations [CC82b]. Concept [ÅFL81, OL80, Tac83, AM89, BBR82, Brä87, Cio87b, ČPV81, DAPS81, HS88, Jul84, Laf81, Par84, PNA86, PNA88, Sch83b]. concepts [Brä86a, DSG80, GB85b, GBW88, GBW89, PNGL83, Tri86, Fli84]. Conceptual [Zer87]. concerned [Cal82c]. concerning [Cla80b, Tat81b, Tho86b]. concerted [HSR86]. condensate [SS80b]. condensation [CTP83, Dos84]. condensed [Bar86, Fri87, KH80, Lar86b, ZV85]. condition [BD85]. conditions [FC81b, FC81c, FC81d, FC82c, Gru85, KC82b, Kar86, Khu82, Kry82c, oLCC81, LW81, Swe83, TS82]. Condon [HYZ89, Kat80, KDA81, PM83, PS88, SPRS89]. conductance [BCD89]. conducting [BCBS81, Ker86, SRC83, VARPU84]. Conduction [UVP^{+}81, BBMP82, CE88, VI84]. conductors [DR80]. Conference [WJ82]. Configuration [DG85, DH84, KG86a, Mül82, SBB87, WP80, Bal81b, Bur83, CS83c, CMB84, DBES80, DK82b, FSS^{+}89, GB85a, GRBP87, Ish89, KC82c, KC82d, KPM83, LER87, LSBS81, Mez84d, Mez87b, OY83, OL82, PB82, Pau81, Pay82, PS81, RP81a, Sav88, SSR^{+}87, Sie80, TR82, TB81, Was89, Yam80b, Zer89, Fuk83, KG86b]. configuration-averaged [Zer89]. configurational [BCC82, LJHU89]. configurations [DIV86, DRS88, FW80a, FW80b, KPM83, KHRK87, KS82, KNM83, LCCB81, MPBN81, Nas83, PŻ85, RSD82, Rud84, SB83a, Sie80, Tai89, Tay83, Tay84, Tay86a, Tay86b]. confirm [CHKB83]. confluent [Niu80]. Conformation [BV86, KG80, Luo87, PD85, PD87, Sar81, AN86, Csi82, GM86c, Gup81, JPAB87, KAH80, KD80b, KMRK86, Mez83a, Pac82, PS81, RTM81, RMM83, RTM86, SP81, Sar89, SS80a, SS86b, TK83, VT82, WWD88, CS81, SC80a, SP82a, SC84a, SP86b, Cal86a]. Conformational [BLR86, BR81b, ITKD80, KŠHH83, Kot82, LCP^{+}88, LN89b, PT89b, RB88d, SWA82, SNSM82, Tew88a, Tva89, WS80a, YL83, AM88b, BMRJ86, BH82, CBV85, CZM83, KTGR89, LM82a, SHL86, Tew87, Tew88b, Yad83]. conformationally [AM88b]. conformations [AM87b, AM88a, BCV81, DPS87, JY81, MY81, MMB84, RB81, Ran80c, RTM83, RMM83, RTM86, SS86a, Sin80b, LHK81]. conformer [DS80b]. conformer-receptor [DS80b]. conformers [LCP^{+}88]. Conformons [VCB83]. congress [Jor89]. conjectured [TV84]. conjoncturelles [TV84]. conjugated [ABTP83, BH82, Bou80, ČPV81, GMM84, Has82a, HO85a, HO85b, KT80b, MS84, OG86, RN86, RSG^{+}87, SF80a, YS80b, YS80c, Živ83, dSdMB86]. Conjugation [RHST88, Vog86]. conjugués [OG86]. connected [Cio87c]. connection [Sil86, Ulm85]. connection-formula [Sil86]. Connections [Sch89, Wes81, Wes82, Lar83]. consequences [BB87, Sla85, TV83a, TV83a]. Conservation [ACCC80]. Considerations [MT88, DOW87, EN87, FN83, KC82c, LSW89, Pul86]. considering [CF81a]. consistency [BCC82]. Consistent [DDP81a, DDP81b, MM81, Wei85, BDK86, Cio89, DS80a, Den83, Den85, DS81c, EG84, Eng83a, Eng83b, FPT81, Lef83, Lib84, LTLF87, LLC83, MM82, MN81, Mic83b, Oga80, PMCA88, SRS81, SSSH82, SM82, Toy88, WG80]. constant [CP85a, CZ82a, Cul88, Cul89, GT86, MA80b, Sel86a, Sel86b, SS89a]. constant-denominator [CP85a]. constantes [TV84]. constants [BYBZB83, EC83, FC80a, FC81a, FGC84, GOS87a, Jon81c, KGEP89, KG89, Kau82b, Kob82, Luk81, Mil86, POK85, RDG^{+}86, SC86b, SS86b, SL80, TV84, WHT86]. constituent [KTK80, KRST83, KRBM85]. constituents [Rab84]. constrained [PM81, RSST86, ZW83]. constraints [Mez86a, Sil80, Swe83]. Constructing [AC89, Bal82a, KPH87]. Construction [AL86, BK82, Cal82b, Cio89, HPKW81, Mey88, Pau79, BH86, Bha82b, DBF89, GBZ86, Kry84b, Mat82, Pan82a, Pan82b, Pan83, Pan84, PKH87, ST86c, Voj85]. constructive [May86c]. contact [FGC84, Ish85]. containing [BR81b, Blo89, BCV82, CNK87, GG81, Gás86b, GTB80, GC84f, KNS^{+}86, MGBH81, PDMC82, PDKC84, RR87b, SD88, SS86a, Živ83]. contains [RL88]. content [BEP81, Fra84, Kin89]. context [ČV82b, Røe81a]. continua [AM88a]. Continued [ČV83a, Mas87b]. continuing [CS87]. continuous [PM88]. Continuum [Ras88, TNB86, Ban82b, Blo83b, CA85, CA86, GjC86, Löw85a, TT83]. contour [HKP81, HLKK82, HKLM85, LHK81, SHKP80, SHP^{+}80, THKM81]. Contracted [BBPS83, STH80, Sie83b, TH80]. Contraction [Bag83b, Pan89, Ben81, GM81b, RS88]. contradiction [Kve82]. contrasting [LPP84]. contribution [DS89b, GCP82, GCP86, HBSC83, KJMS85, Luk81, Mas87a]. contributions [Big86, FC81a, FGM87, FNC86, GTB80, LCC80, TT83, Cul88, Cul89, PHGR88]. Control [Mar88c, Gas85]. Controlled [Pip85, Van82]. conventionally [Ban82b]. Convergence [BMBK82, DMB^{+}84, Del84b, HČP88, Kar81, NS84, IS85, OJY83, PČP82, WGS85, CS80b]. convergent [Ada82, HGPF89]. Converging [CHZ82]. conversion [TLC87]. converters [KB87b]. Convolution [Niu84b]. Cooley [Tel88]. Cooper [Ågr82, Muj86]. cooperation [Shu84]. Cooperative [Jac89, DS80c, OPSO86, OPSO87, DSG80]. cooperativity [SS89a]. Coordinate [Bro80, CS83c, Don82, NS82b, Ade87, BS80a, BS82a, BLV88, DL83, DLVÖ84, FSW87, GM86b, KRT83, KMW80, Lev89, Mez84c, MT88, Moi81, SK87, Sim80, Tac87, YY83]. Coordinate-rotated [Don82]. coordinated [CSB86a]. coordinates [AW83, ACDLH84, AZY84, ALH86, BB83, Ban88a, Ble82, Csa86b, LÖ85a, LV87, NPL89a, NKM81, NMK84, SB89, Win86]. coordination [SCER88, VMMV80]. copolymerizations [FW84]. copper [Bag83b, LdMH80, MTN83]. cordycepin [SP81]. Core [Ågr83, Dur86, LP87, MB84, PC88, AHB^{+}80, BBH87, Boč89, CMM83, Kah84, KK87, LJC83, LLS83, PLL86, SAD81, SSF^{+}84, WKJ88, Woo82]. core-electron [CMM83]. core-exciton [SAD81]. core-valence [PLL86, SSF^{+}84]. Core/valence [PC88]. cores [ALL83]. Corfu [Kel78]. corrected [AGCT87a, GBZ89, SRHK83]. correction [CV83b, Lud83, Rom81, VV83, Ten88a]. Corrections [PHGR88, AGCT87b, BSZ81, BCT86, BCC82, BBO83, CPGP83, Csa81, FSS^{+}89, GV88, Ish89, LJC83, McD80, MMC83, Noa80a, Noa80b, PW86, RLS87, SNSM82, VI84]. Correlated [Bor85, LC88b, MÖ80, Gos82b, HO85b, JD83, JD85, KC82c, KC82d, Luk82, MSF86, NP80, PŻ85, TB81]. Correlation [Big86, CF83, CFLM85, CL86, Ess86b, GT87b, II82, LCLC88, MK85, MA80b, Sad83, Sar89, SJH89, AB84, ABTP83, AYCD81, ACDLH84, Bec81, Bec85, Böh85, BC83, CR81, Cio87c, CSF80, CP80a, CP83a, Dav81, Del87, DS89b, DRS83, ESS86a, FD89a, FDA83, FH81, FH83b, GBZ86, HBSC83, Has82a, Has82b, HS80b, IS83, JMS^{+}85, JMS88, KU80, KKA80c, KFN80, Lad83, LCLC86, LC80, LT89, Lev89, LZ89, LSH88, Mar85a, Mar88b, Mic81a, Mic83a, Mic85, Mic86, MHAS86, MMC83, MRJ85a, MRJ85b, Nag87, Nes87, Ori84b, PW86, RBR80, Røe81b, Røe82, Røe87, SSB83, SF81a, SL86, SMJ83, SM87a, SSF^{+}84, Suh83, Suh84a, TPČ83, UB85, UB86, VM81, VM83, WSD80, WBB80, YMO80, BR81b]. Correlation-induced [II82]. Correlations [BMMS81, Dur86, GR89, HMF82, Hub80, JŻ85, MK81, NH81, PHGR88, SS80b]. corresponding [Mof81, RC80a, Ská83]. coset [QX86]. Coulomb [ALH86, BD88, Bli80, Bli84, Bot81, Bri87, CC84a, CP80a, CP83a, ENH^{+}87, GBK86, Gus87, Jon81b, Jon82, LO81, Lie83, MC88, MM87b, NBE87, Noa80a, Noa80b, WW89]. Coulombic [KM89a]. Coulson [ČD87c, FEKC84, Hal88]. coumarin [BLY^{+}86]. count [Cio87a]. countercation [CLP82]. counterion [LCP81]. counterions [BKJ88, CC82b]. counterpart [Loh85]. Counterpoise [BCT86, AGCT87b]. Coupled [DP81, SB87, SM84, ALB84, AB85, AB87, BAG83a, BS81a, BL81a, BS83b, BHS85, DL82, ERB82, GOS87a, Gee88, HK86, HB86, HZBR89, HY89, JP80, Kal86, Kva83, LPB82, Lef83, LAGS89, Lin85, MJW88, MR87, MK80, MMC83, Muk86, PTC84, PABL88, PP89, Pul83, PSBB83, RSMM86, SSR^{+}87, SB84, Küm83, TP85]. Coupled-cluster [SB87, SM84, ALB84, BS81a, HK86, Kal86, Kva83, MJW88, Muk86, PTC84, PP89, PSBB83, RSMM86, SB84, Küm83]. coupled-pair [JP80]. Coupling [WW82, Ågr82, ACCYCC80, BS84, BD85, BLV88, CKD81, CT87, DBVMB85, DIV86, EC83, Eng83a, Eng83b, FC80a, FC81a, FGC84, FW80a, FW80b, Fis84, GOS87a, Gre87, GC82b, GC83, GC84f, HS80a, HLPV83, Kar84a, KGEP89, KG89, KL89, RTC81, Röm83, RDG^{+}86, SC86b, SS86b, SS89b, SDS83, WHT86, YA87, AØ80b, Cal86d]. couplings [RDG^{+}86]. covalency [JL85, MOT^{+}83]. covalent [Has82c, Hur82, Ols83, PNGL83, SAB^{+}80c]. covalently [MT85]. coverage [RLH86]. CoXF [Dat82]. CPA [DL82]. Cr [CEZ83]. Craig [Tap86]. CrC [SG89]. creatine [GB81]. creation [Pan82b]. CrGe [SG89]. Crick [SS85]. cristal [OF80]. Criteria [Lit81, MW80, Gas86a]. criterion [GA82]. Critical [BH82, LM86, Mou89, DBC84, EB86, LSR^{+}84, Sta71, Sta87]. criticality [Fis82]. Croatica [Zer87]. cross [Blo83b, BF83b, DSG86, FSW87, FK80, Gan81a, Gir86, HDFL84, LGS87, MM87a, MW83, MÖ80, NC87, SM83]. cross-section [LGS87]. Crossed [Fut87]. Crossed-Molecular [Fut87]. crossing [BKP83, LC83a]. Crowther [Buc89]. CrSi [SG89]. Crystal [Cal81b, SRL^{+}83, ALL83, AØ80a, AØ80b, AØM81, AØC82, Ban82b, Böh84, BFOP83, Cal85, DP87, INFU84, Kar81, KS84, OF80, PD89, RB88c, RKHK88, RKO^{+}89, SP82b, WS79, ZY85, Zhe83, AØM81, AØC82, RB88a]. crystallization [LM89]. crystallographic [GC87, MR89]. crystallography [Fra89]. crystals [AS89, Ano87a, Bou84, BLOW89, BTC^{+}83, BOW80, FM83, KBS81, Kau87, KS87a, LSR^{+}84, MK83, OPSO86, OPSO87, PS80a, ST80a, ST80b, See81, Sin86, ST86b, Sto82a, Sto82b, TP87, UT80]. CS [DFH84, Mül82, Kat80]. CSFs [BL81a]. Csizmadia [Ågr87c, Cal86a]. CsO [LLS83]. Cu [SS88, BCA87, MB84, MB85, PC83b, SD80, Sch88]. cubic [BOW80, Mod81, WHT86]. cubical [KG85, KG86c]. CuCl [CCS84, KWBC80, WKJ88]. CuF [LHC80]. Cummings [Lun82]. Cumulant [Mic81a]. Cumulative [Nas83, SHK87b, SMHK89]. CuO [MS88]. current [CC84a, FH83b, LRZ84a, LRZ84b, Wal88]. currents [GMOBA80, TKY89]. curvature [Sil84]. curve [BSR80]. curves [BH89a, CLS81, KT84, MS80b, PPB87, SG85a, Tat81a]. cyanide [TF83b]. cyanides [Mof81]. cyano [Hir80a, LH85, Mof82]. cyano-substituted [Hir80a]. cyanocarbonium [Mof81]. CYBER [AM87a]. cycle [Fox81, PSG83, SD82]. cyclic [BMT80, Bou80, DDFA87, KHRK87, PB82, PC83a, PCG83, PT84, PP83, RSP^{+}80, ST83, TP84, TP85, TPČ83, VB85]. cycloaddition [GDX89]. cycloadditions [SZJS82, HSR86]. cyclobutadiene [ÅCFRJ85]. cyclodextrin [MP84a]. Cyclodextrins [BP83]. cyclodisiloxane [BS86]. cyclohexadiene [SAM82]. cyclooctatetraene [Wol88]. cyclopentadienyl [XHR86]. cyclopentadienylidene [SS80c]. cyclophosphamide [Ulm81]. cyclopropane [PGL^{+}86]. cysteamine [BCV80]. cysteine [GSMSG82]. cytidyl [Gre82]. cytidyl- [Gre82]. Cytochrome [MPP^{+}81, AWCL89, CLL86, LCCW87, LCA89, LL85, PL83, PL84b, WCLZ86]. cytosine [BCD^{+}80, KPSS87, LOB86, RBSR84, SS85]. cytotoxic [Gas86a].

D [Ågr82, Ågr87a, Ågr87b, Cal81b, Dat82, Dun80, Gos80, Mic87, Sie81, Sme84, Tap86, Wol81, BW85, ESCC85, GT87a, WHT86, Wój86a, ddF86, KG85]. damped [SCH84]. Damping [vZ81, Dun87]. Danail [Ler85]. Dancoff [TF83a]. dans [TV83a, TV83b, TV84, TV89, Van85, VT85, Van86]. data [FSK79, FF89, H^{+}90, KPS87a, MR89, OM82a, OM82b, OM83b, OM83a, OM84, OM85, OM86, OMH89, OYH^{+}80, Pic83b]. DATB [HKKM83]. Daudel [Fro88b]. Davydov [Cal86c, Tus86]. Day [Rat88]. dc [HS84b]. DCl [Wój86a]. DCN [Cha88]. dCpdG [HB82]. deactivating [Sey80]. deaminase [SRS81]. deazapurine [SC84a]. Debye [Fra89]. decay [Emc80, MGBH81, IHI80]. decaying [Kau82b]. decomposed [ML81]. decomposition [AGCT87a, ACH83, BCT86, CA80, DT83, KHCH85, LY83, OS89b, RKHK88, RKO^{+}89, SRHK83, SLT84a, SLT84b, ZS86]. decoupling [Dal80, Shi87]. Dedication [Rav82]. deduced [FS80]. deductions [Sin88]. deep [RZWL80]. deexcitation [YUKK84]. defect [Ell84, GAL86]. defect-pair [GAL86]. defects [BS84, BF81, Cal80, CS80a, EGP86, KWM86, SF89, SC89]. defense [LP82]. deficient [Vog86]. defining [Ori84b]. Definition [DQT80, DQ80, DQ81, CM85, Gre87, SMVV86b, SMVV86a, TV84, TV84]. definitive [GC83]. deformation [AØ80a, AØ80b, AØM81, AØC82, BSCR80, FZB86, GKP83, Eis88]. deformations [BST86, HG86]. deformed [AS89, GKAP81]. degeneracies [Avr82]. Degeneracy [LL86, PTC84, Bal81b, CDL82, Ojh89]. Degenerate [CS80b, BB81b, DS88, Gol84, ML81, WJP83, WJP85]. degradation [DSG86, IDK87]. degree [MOT^{+}83]. degrees [HL88]. dehydrogenase [KK82b, KK83]. Del [Fro83a]. delay [WB81, YKZ^{+}89]. delayed [Bel80b]. delayed-choice [Bel80b]. Delhalle [Cal81c]. delocalization [BC83, PSG83, PM88, Pip89, SNSM82]. delocalized [LW84b]. Dence [Nat88]. dendritic [Ran81]. denominator [CP85a, CZ82a, Cul88, Cul89]. dense [Fla81]. densities [AØM81, BS83a, DT81, Dea89, Eis88, Fla80, HS84a, Ish86, Ish88b, KAD81, POK85, PO85, SMVV86b, SMVV86a, SST84, TM81, TTS84, TP89, WB81, YLM80, HLPV83]. Density [Bel80a, BK84a, BOW80, Cio89, CFG^{+}83, FM83, Har83a, KL87, KGO89, Lie83, LCG82a, Mor88, NPL89b, Per85, Per86, Sha85, Tac87, Ada80b, Ada86, AØ80a, AØ80b, AØM81, AØC82, BPT88, Bec82, Bec89, Bel80b, Cal84, CLA80a, CEB86, CH86, Cas88, CSF80, CFLM85, Coh86a, CG80, Csa81, Csa85, Csa86a, Csa87, Csa89, Dal80, DG83, Del84b, Don81, DK81a, DK81b, EIMS88, Ell84, EB87, FCB80, GD82a, GV88, Gos82b, Gre80, Har80, Hil81a, HG86, HM89b, Ish89, JBW87, JMD88, KPH88, Kau82b, KMW80, Kog84, Kon83, KN86, KM89a, Kry84a, KPS87a, KPS87b, KPS88, KEB86, KPH87, Küm83, KWM86, Lar83, LRZ84a, LRZ84b, LP82, LP85, Lev89, LCCW87, Löw86d, Luk82, LW81, Mal86, Mar86a, MC80, Mes80, Mes85, MS80a, MKW85, Min87, Mod81]. density [Mor89, NSS86, NDBE82, Noa80a, Noa80b, PFBF81, PSK86, PW86, PC86, SKIK82, Sav88, Sch80b, Sen80a, SL86, SY86, SHL86, Tac88, Tac89, VCG86, VBS85, YKZ^{+}89, ZW83, ZAB85]. density-functional [Bec89, Csa89, Lev89, PSK86]. density-matrix [Mes85]. deoxy [LMG85]. deoxyadenosine [SP81]. deoxyfuranose [LMG85]. deoxyuridines [SP82a]. dependant [STV80, SV85a, SV85b]. Dependence [PM88, AGCT87a, BS86, Buf82, CS83c, Dal80, Fra85, HS84b, Has85, MC88, MTI80, OI86, ÖHD89, RSP^{+}80, Sad89, TSA86, ZP82, AN86]. dependences [LM82a]. Dependency [BSD82]. dependent [AM84, AJY80, BP86, Bot81, DÖ89, FC82d, GM86a, GC84b, GC84e, GY87, HA84, Jac80, Lar83, MZ85, Mar85b, MLC^{+}81, McW83, MB80a, PGM80, QX89, RB88d, Sla85, SV85a, SV85b, SvG89, SHS84, Tac82, Tru83a, SvG89]. depending [Ste81]. deperturbation [LGS87]. dephasing [BP86]. deposited [MB89]. depth [BH89a, CE86]. depth-dose [BH89a]. Derivation [LL82b, Fol81, Hel82, KC82b, KSGB^{+}83, Kut80, oLCC81, Pau84, Pau89a, Sil82, Tim86, Toy88]. derivative [She87]. derivatives [Ågr87b, AEHNEA85, AT85, BJ87, BSL^{+}80, BCD^{+}80, BLY^{+}86, CGP87, Cio87b, JG81, JD85, JS86, KNSM88, KTGR89, LOB85, LOB86, Luk81, NHKY83, OS89b, PD85, SB87, Sel88, Sey80, SS80e, SKS83, SJ84, WMP^{+}83, WS80a, WHT86]. Derived [RSBR85, BS83a, Jon82, RBSR84]. deriving [Wes81]. describe [Eng83a, Eng83b]. described [Røe81b, Røe82, Røe87]. described* [GS89]. describing [Ish85, RGW88]. description [AM88b, BFOP83, Fli84, FGG83, Gee89, Gol88, GC82b, HO85a, HO85b, HJ80, Koe86, Kva82, MM80, PBGT80, SCH84, Sch89, MA80a]. descriptions [Mar88c, OWÖ81, RACPL86]. deshielding [FNC86]. Design [AIJW82, Lit81, Mez85b, NS89, PSG83, Sok81]. designed [LRT88]. desirable [Kau87]. detection [CD89a, SDS83]. Determinant [Kia80, BL81a, CFLM85, Fri86, GS80, KHCH85, Lim88, LC82b, Pau89b, YS80b, YS80c]. determinantal [PL80]. determinants [Duc86, Gal82, Löw83, RWOT81, Sey83a, Was89, WMP^{+}83, MWO^{+}84]. Determination [FF89, GSS85b, Pac83, RVK80, SV85a, BL81a, BF83a, BF83b, BV86, CC82b, GM81b, Kar84b, KJMS85, LLS83, LAGS89, MMB84, Mic83b, SK80, SV82, TV84, Umr89, DT81, Umr89, SV85a]. determinations [SB83b]. determined [KB86, Qia83, RSST86]. déterminer [SV82]. determining [BB84, Bar86, DS89b, LT89, Poh83]. detonation [BWEW89]. detonations [OPSO86, OPSO87]. detoxification [LLM86]. deuterium [BBMP82]. deuterium-hydrogen [BBMP82]. Development [JIS87, OMCD81, Pic83a, RL88, Ågr85, RDG84b, RR88, Str82, Var86]. Developments [HC81, Enz80, Jor81, Kar80, KEB86, PP84, PP81, Tho86b]. devices [Ber88]. d’évolution [TV83b, Van86]. Dewar [TOK80]. d’hyperpolarisabilite [SvG89, SV85b]. Di- [RTM83]. Diabatic [Hal87, KPH87, PKH87, PABL88]. diabetes [RB88d]. diacetamide [BDR^{+}88]. diacridines [AAW82a]. diagnostic [LT89]. diagonal [CF81a, FC82b]. diagonalization [CS89]. Diagram [BS82b, Rut81, YMO80, MP84b]. Diagrammatic [RS84, BK89, KU80, SI89a]. diagrams [ESS86a, Hub80, Kva82, MRJ85a, MRJ85b]. dialkylnitrosamines [LPSK83]. diallene [DBD^{+}85]. Diamagnetic [Kau82b, Bou80, DCF82b, SG87]. Diamagnetism [GMS81]. diamino [PBM85]. diaminodichloroplatinum [EIMS87]. diamond [Lás82]. Diatom [SB89, BS80a, Lin89, NP81, Suc81, Var81b]. Diatom-diatom [SB89]. diatomic [AB84, AS89, CKM81, GSC89, JEW88, KG86a, KG86b, KK82a, Kob82, Kob86, Kry84b, LSP85, LPC^{+}80, LL82a, LM86, Leu83, Moh80b, Moh80a, PL80, PGM80, PPMG89, RZP84, Röm83, WS80b, ZNS80, ST86a]. diatomics [BF83a, BVAB81, Kun86, PD84a, RSSS81, SST84, TTS84, Voj85, ZNS80]. diatomics-in [ZNS80]. diatomics-in-molecules [Kun86, Voj85]. dibenzo [Che82]. dibenzo-p-dioxins [Che82]. dibenzofurans [Che82]. dibenztetrathiofulvalen [SDS83]. dichloro [VG80]. dichloro-bridged [VG80]. dichloroethylen [LPP84]. dichlorooxirane [LPP84]. Dielectric [CE86, CE87, Pet79, CA85, GGDG86, GGDG87, MR87, Poh83, VVD80a, Mår80]. dielectronic [Win85]. Diels [Bur86]. Diels-Alder [Bur86]. Diestler [Nat88]. difference [GSS85b, IPV81, SMVV86b, SMVV86a, SHKP80]. Differences [KHRK87]. different [AGCT87b, CC82b, Coh86a, EIKG^{+}88, FNC86, GLLOB86a, GC84f, Hil81a, LC88a, PP89, Sey81, SBWW80, TPN86, YGK^{+}89]. differentiable [Cho80, Mez83b]. Differential [Mar86a, Dem86, ERB82, WSD80]. differentials [EIMS88]. difficult [Ada64, ČD87c]. diffraction [CFG^{+}83, GM83, KPS87a]. diffuseness [LR89]. diffusion [BASB86, Ulm83, Ulm85, Van82, VR86]. diffusion-controlled [Van82]. difluoroethylenes [HET86]. digital [Cla80b]. dihydrides [Bal88]. dihydrofolate [GC85a]. Dilated [ERB82, FB83]. dilemma [PMT84]. dilemmas [VFA81]. dilithioacetylene [JPAB87]. dilute [ESCC85, Fla81]. DIM [Voj85]. Dimaprit [PMOW89]. Dimensional [LPS83b, PBMR^{+}84, AM88b, AW82, Ave86, ALH86, BVB88, CC84a, DFP80, DR80, FF85a, FF85b, GD82b, GM86b, HM89a, II82, Ker81, Khu82, KS87a, KBT80, Kry84a, Kve82, LPS83a, LSP85, MR80a, Mar86b, MC89, MST^{+}80, MBD84, Niu84a, PS88, PD89, RB86, RB88a, RB88b, Ran88, See84, Sri86, SRDG85, TIYH86, TNY83, Tel88, Tus86, YKZ^{+}89]. Dimensionality [MC88, Mar85b]. dimensionally [MS80a]. dimensions [Avr82]. dimer [BSR80, CSB86b, HB86, KNM83, Lip83, MR85, OMC82, OL80, SHK87a, WHT86, XHR86, ZCH89]. dimers [AGCT87a, AGCT87b, Cze87, HFSB86, Mil86, OKZ85, Pee83, SF81b, VG80, WS80b]. dimethyl [LHK81, SJH89, TR86]. dimethylamino [PD87]. dimethylaminoaniline [PD85]. dimethylaminobenzylidene [PD85]. dimethylbenz [Sey80]. Dimethylnitrosamine [ST82]. dinitrobenzoyl [ST89]. dinucleoside [Gre82, GK80c]. dinucleotide [JY81]. diol [LPCF80]. dioxethane [Hil81a, Hil81b, Hil81c]. dioxide [Blo89, SKIK82]. dioxins [Che82]. dipalmitoyl [DS81a]. dipeptide [HGHGF^{+}89, MMB84, YJ85]. dipeptides [HH89]. dipolar [FC81a, HHT82]. dipole [JD85, LM82a, Mak80, MT87, MPP83, SSR^{+}87, SS81a, SJ84, SMJ83, VU85, WKH80]. dipoles [KS86c, PPP87]. DIPS [ECM89]. Dirac [AN85, Csa81, Csa88, EG84, Ish84b, IBB85, IBB87, IRA87, IQ87, IS88, KK86, KPP85, KPP86, Kut84, VV83]. Dirac-Fock [IBB85]. Direct [ÅCFRJ85, AB85, BR81a, Ess84, FPT81, GM81b, JJČ87, Kal86, Ret86, Sie80, SP82b, DG83, DH85, DK82b, GRBP87, Iga85, LSBS81, MJ89, Mes80, Mes85, RS80, Sin80b, Win82, Kar86]. directly [Sin88]. Dirichlet [FC81b]. disaccharide [YL83, Yad83]. discharge [KKCM85, KKC^{+}86, KKCM87]. Discrete [BJ87, Ban88a, TT83, AM88a, DR88, IBB85, IBB87, SA80]. discrimination [Shi85]. Discussion [AMFC83, Jas80, Löw84]. disilene [GZWYCHC88]. disilene-silylsilylene [GZWYCHC88]. disilyne [KSO84, SWM82]. dislocation [ALL83]. dislocations [Ban82b, De 80]. disorder [DR80, RM82, SKL80]. disordered [Fau85, Sri86]. Dispersion [VTT81, Cla83a, Kar80, Kau83, Obc87, PPB87, YS80a, BSL81]. Dispersion-equation [VTT81]. displaced [RLR88]. displacement [HB82, SK87]. displays [Fin89]. dissipated [LCD82]. dissipative [CD83, CD89a, SCH84]. dissociation [ADT^{+}89, Bec85, BF83a, Eng83a, Eng83b, Hil81a, Hil81b, Hil81c, LAGS89, PABL88, Pit84, RSSS81]. distance [AGCT87b, LBKV88, MT87, MK83]. distorted [MW87]. distortion [KKA80b]. distortions [Loh87]. distribution [ČD87c, Csa82, Csa83, HOT87b, Hil81a, HDTR83, KKZ86, LCCB81, LCCW87, Pac83, ST80a, SHKM80, UB86]. distributions [Bow86, CHA87, Fli84, Fli86, Hil81b, OI86, SSK84, SHL86, TMC84]. disubstituted [Ågr83, KKA^{+}83, WCE89]. divalent [Pac83, UTP82]. divergence [SM87b]. divergent [Bha82b]. divided [Pes83]. division [SAB^{+}80c]. DMN [ST82, ST82]. DNA [BG81, BST86, BC80, Bre80, BCV80, CC82b, CCG^{+}82, CCD^{+}84, CLP82, CP83b, Del88, DD89, Fin89, FPDA89, GM81a, GUSS89, HH89, KRMK86, KJL87, KBR83, KMRK86, KMKC88, KPSK80, LS80, LOF83, LCDP80, LP81a, LPP81, MT85, MLC^{+}81, MLA80, MSGPP80, Ott86, Pac82, Pac83, PWL89, POT80, PZSL83, PZPL86, SMSA80, Sar82, SKER83, SRSR85, Suh84b, SR83, VPK80, VC83, WMS81, vZ81]. DNA-dependent [MLC^{+}81]. DNA-lucanthone [SMSA80]. Do [Dun88, Cas85, ST86a]. Does [Mis84]. domain [MB83, PKG86, Van85]. domaine [Van85]. domains [Ori84b]. dominant [KHRK87]. Donini [Cal86b]. donor [Bha82c, CM81, CE86, PS80a]. donors [CE88, DAL89, GF88]. doped [BKJ88, Fer88, Ker86, MTI80, PD89, ST80a, ST80b]. doping [ECM89]. Dordrccht [Dun80]. d’ordre [SvG89]. Dordrecht [Ågr87b, Cal81c, Cal81b, Fro88a, Fro88b, Gos80, LM81, Mar81a, Mic87, Sme84]. Dordrect [Cal86c]. dose [BH89a]. Double [DP82, Lim86, QX86, Alt86, Alt05, Bel80b, BST86, CFC83, Cul88, DBF89, GG82b, GP83b, GSS85b, ITKD80, JD83, JM83, KO81, Lás85, Lim88, Mey88, MMI^{+}82, Ori84a, Ori85a, Ott86, Pau89b, PHGR88, PSBB83, Rös87, Sar82, SSR^{+}87, VPK80, Var87, Yam81]. double-ended [DBF89]. double-helical [ITKD80, VPK80]. double-point [Mey88]. double-slit [Bel80b]. double-stranded [GP83b, Ott86]. double-zeta [GG82b, Lás85]. doubles [Gee88, LPB82]. doublet [Don85]. doubly [Gub82, MS86, PP89, Vol89b]. downward [Pau81]. downward-Robb [Pau81]. Dr. [Ban82a]. dressed [SH89]. driven [Cio89, FC87, YG87]. Drug [Dav87, BV86, CL83a, JW86, Luk81, MWO^{+}84, Mez85b, NS89, PKH^{+}81]. drug-receptor [Luk81]. drugs [AEHA87, AD80, BC80, TH88]. dT [CLP83]. dual [ACCC80]. duality [Wor81]. due [Shi87, Swe83, VU85, ZRY83]. d’un [SV85a, SvG89]. duplexes [BG81]. During [Dat82, ES88, PSG83]. DV [TH81b]. DVM [GBO86b]. DWBA [Suc81]. dyes [PD87]. Dynamic [ES80, Sin81, TNY86, BEO80, Rér89, Sve88, VU85]. Dynamical [Fol81, MSF86, SA86a, CD85, DÖL87, FK80, PW86, Sch80b, Tab82, Toy80]. dynamically [MR87]. Dynamics [Gup81, MLL83, MM80, RN84, STG86, Ade87, ABCV86, Bes88, BS83b, Cal82a, Cla86, CJ89, D^{+}85, DÖ89, DBF89, Emc80, Fro88b, Fut87, GHF84, GD82b, GDBRV87, KH80, KH86, Luo87, MS84, OF80, Pal89, PABL88, RC86, Sch89, SB81, Woo80, YY83, Yar82, ZH86, Fro88a, Mic87]. Dynamique [OF80].

E. [Ano80e]. each [Kau82a, TPN86]. earlier [CHKB83, MR81, MR80b]. earth [ADS83, GEdP81, LBP84, SS81a, SS86a]. Ebb [SHRH82]. Eco [Fin89]. economical [RR87b]. Ed [Cal86a, Cal86b]. edge [DA88]. edges [LGL^{+}84, VI84]. Edited [Ågr87a, Ågr87b, Ågr88, Brä86a, Cal81c, Cal81b, Cal82a, Dun80, Fro88a, Fro88b, Gos80, LM81, Mår82, Mar81a, Mic87, Sie81, Sme84, Tap88, Zer87]. Editor [Ano89m, LCÖ84, BL89a, Cal89a, Swe83]. Editorial [Ano86c, Ano89l, LSZ83d, LSZ83a, LSZ83b, LSZ83c, Sab80b]. Editors [Rat88]. Education [Bry80]. Effect [BBMP82, BSL^{+}80, CP83a, CE88, DKLP81, DS81a, FS86c, FCPL84, KL80, KAM85, KBSG81, LCDP80, MT85, MG80, MB80b, NSP80, SMJ83, SKL80, YA87, YRS85, AGCT87b, AL81, Aus84, BBLR84, BKS87, BP84, BEL89, BP86, BK84b, CV85, CLP82, CT87, CP80b, DS89a, DOW89, DQ80, DQ81, FHHE80, GSC89, HET86, Hir80a, IHI80, KVR85, KFD84, LK87, LY83, MHAS86, NS82a, NS83, PNA88, PP80, SS84a, See84, SL87, SS89a, Sil82, Sin81, SHKP80, SRHK83, VWCC82, Bak85, Boč89, CBV85, RBR80, SS85, SGP^{+}80]. Effective [GMM84, GR89, GCV84, JSS80, OY83, Var81b, BK82, CF81a, DB81, GBZ86, Kah84, KK87, KSGB^{+}83, LJC83, LLS83, MT87, PC88, PNA86, SWM82, WKJ88, Wes81]. Effects [CZ81, NW82, RS88, SR84a, SAD81, SD88, AJ83, BSHK86, Bec81, BMRJ86, Boč87, Boč88a, Boč88b, Boč89, Böh82, BYBB84, BC83, CKD81, CZM83, Cla83a, CG80, CA85, CA86, CP80a, Dav81, Del87, DSG80, DS80c, DR80, Dur86, EV87, FGC84, Fli84, Fli86, Fuk83, FM86, GPN80, GjC86, Gre87, GKP83, HF86, HBC86, HM86, HL80, Ish89, IS83, JMS^{+}85, JMS88, KKZ86, KM80, KFN80, KKD80, KS86c, LCLC86, LCLC88, LP87, LLM86, Lun89a, MK85, MFM89, MB84, MR87, MR88, Min80, MB85, NH84, Nov83a, PC88, Pac82, PCA89, Pee89, Pee83, PD84b, PP81, PZPL86, PBMR^{+}84, PTG80, Rab84, RKV^{+}80, Sad83, SSB83, Sey83c, Sme86, SR87, SST83, STG86, SP82b, Suh83, Suh84a, SI89b, Toy88, Tru83a, YG86, Yar87]. effects [ZSB80, SO85, Wol88, Mar81a]. efficiency [HZL84, OY83]. Efficient [Duc85, AB84, CY80, GHGP89, NKV^{+}89]. EHF [MA80a, PS87]. EHMO [TW88]. EHT [BSZ81]. eigenenergies [Moh80b, Moh80a]. Eigenfunctions [Cal82b, Pau79, AL86, Duc86, Lim86, Lim87, Rei80, Sil86]. Eigenlevel [BD88]. eigenstate [BHM87]. Eigenvalue [CS89, DM83a, Dem86, FCO89, Kob82, Kob86, Sen80b, Sim82a]. eigenvalues [AMFC83, BEP87a, Cas88, ČV86, ČD87c, MW80, Rei82, Sil80, Sil86, Ská83, Swe83, TL83]. eigenvectors [Ská83]. eikonal [SM82]. Einstein [Gre81a, MB83]. Einstein-like [MB83]. ejected [DT81, MÖ80]. ejection [CJ89, MM87a, KPT88]. Elaboration [GCP86]. Elastic [Kon83, Blo83b, Chi81, GSRG80, GVB80, KS86b]. elastin [VU85]. elastomeric [LJHU89, Urr87, Urr88]. elastomers [UV83]. electon [KTS86]. Electric [MB85, Rér89, ST80a, ADF^{+}88, BB87, DBD^{+}85, FC85, HS84b, HR87, IS80, MA80b, Pac83, Rab84, RPS82, ST80b, Shi85, SMJ83, WKH80, MPPW82]. electric-dipole [SMJ83]. electric-field-gradient [ST80b]. electric-field-induced [HS84b]. electric-field-variant [ADF^{+}88]. Electrical [BEP81, KKTE80, BBMP82, DDFA87, Poh82]. electrically [Dem83]. electrochemical [GSG84]. electrocyclic [FEKC84, SZJS82]. electrode [MV87]. electrode/liquid [MV87]. Electrodynamics [Tap86, CT84, CT98, FGMM80]. electrofusion [PPR84]. electrolyte [BQS87, MV87]. electrolytes [Lim82]. electromagnetic [Lun89a]. electromagnetism [HM85b]. Electron [Ada86, BCA87, BS83a, CP80a, DSG86, DRS83, FD89a, FDA83, FBM80, Gan81a, Gas80, IL84, IS83, JMS^{+}85, KCP80, LBC^{+}87, Lar82, Lar86b, LVL^{+}85, MM87a, Mar85a, Mar86b, MR87, Nes87, Ort89, PSG83, PPMG89, PŻ85, RJ88, RJ87, Røe81b, Røe82, Røe87, SCER88, SHKM80, Sim82b, SAB^{+}80c, SDS83, Suh84a, TNY83, WBB80, Wea84, Ada80b, AN85, All87, AL80, BBH87, BMG82, BB82b, BS84, BLG85, BN80, Bec83, Bha81b, BSL^{+}80, Blo83b, Böh85, BS81b, Bor82, Bor85, Bri86, BL89b, BOW80, CLA80a, CHA87, CMM83, Cio88b, CFLM85, Csa82, Dat83, Dav81, De 83, DG83, DÖL87, EIMS88, EB87, Ess84, FPT81, FZB86, FHHHE85, FH81, FH83b, Gas81, GSM83, GSG84, GR89, GTB80, GD82b, GHGP89, Gla85, GM86b, GC84b, GC84c, GC84d, GC84e, GC85b, GSRG80, GBZ86]. electron [GBK86, Gus85, HF86, HR84, HS84a, HOT87b, Han88, Has82b, HS89, HS80b, HT85, Hun86, HM89b, IDK87, IPV81, II82, JMS88, JŻ85, JSS80, Jor81, KPH88, KKZ86, Kau82b, KB86, KS87a, KJMS85, Kon83, KPS87a, KEB87, LJW87, LPS83a, LCC80, LCULC87, Lar83, LBKV88, LL82a, LM86, LRZ84b, LM89, LC83b, LT89, LZ89, LC88b, MZ85, Mar86a, MR88, MÖ80, MDT85, MPP^{+}81, NH81, Nes85, New80, NMK84, NC87, Nor83, OI86, OY83, Ori82d, Ort87, Ort88, POK85, PO85, PDKC84, Pul83, RP81a, RBR80, Sad83, SD80, SKIK82, SSB83, Sch83b, SMVV86b, SMVV86a, Sen80a, SNNG86, SG89, SST84, SMJ83, SvG89, SB88, Sri86, SST83, Sto82a, Sto82b, Suc84, Suh83, TH81a, Tai89, TS80, TPČ83, TNT^{+}86, Tat81a, Tay83, Tay84, Tay86a, Tay86b, Ten86a]. electron [Ten88b, TP89, TPN86, TN80, TMO86, UB86, VTT81, WSD80, WJ80, Wea83, WOT89, WKJ88, ZV85, ZV86, ZW83, Živ86, Živ87c, NH81, YMO80, GVB80]. electron-exchange [New80]. electron-H [Wea83]. Electron-hole [IL84]. electron-lattice [BS84]. electron-molecule [All87, Blo83b, Jor81, Nor83, WJ80]. electron-oxygen [GSRG80]. electron-pair [UB86]. electron-repulsion [Gus85]. electron-SF [GVB80]. Electron-Transfer [SCER88, MPP^{+}81]. electronegativities [GBZ89]. Electronegativity [Loh84, Loh85, SM80]. Electronic [AEHM85, AYBP83, ADL78, BCB85, BS80c, CEB86, CZM83, CLA83b, DPZ80, DS82, De 80, DBS80, EZ83, ECM89, Fau85, FG82, FFF88, HBSC83, Her85, HM84, IST89, Kah84, KN80, Ker86, KPSK80, LOR84, LLAG89, MTN83, OM82c, Ori82c, PL89, PC83b, Sap80, Sch88, SNNG86, SM80, TKY89, TM81, TWL82, TH81b, Ulm81, VG80, WBM88, XHR86, YZ85, Ågr85, AGB82, AD80, ADS83, Ano87a, AL87, BZA84, BL81b, BVB88, BCD^{+}80, BDD81, BCBS81, BLV88, BC83, Bur83, CHKB83, CZ81, CPGP83, CG86, Csa86a, CLZ84, DFH84, DCF82a, DF83b, EB86, EV87, FMF80, FD89b, Fis84, FA86, FD89c, GSMRM85, GAL86, GJMLRGSA85, GT87b, GC82b, GC83, GC84f, GBO86a, Hel82, Hil85, HC81, HG86, HL88, HS88, JP80, JD83, Kal85, KC83, KT80a, KKA80a, KKA80c]. electronic [KLS89, KL89, Kry82b, KWBC80, LmJQHXZ83, LKB89, LZY83, LNN84, Loh84, LS82, Luk82, Mår80, MGN82, MPBN81, MS88, Mcc83, Mez82, Mez86a, NSS86, Nec83, NPL89b, NB79, NP80, OF80, OWÖ81, PG80, Pan85, PKG86, Pan89, Pet79, PBM85, POK85, PBGT80, PBMR^{+}84, RDG84b, RDG84a, RGDF85, SP86a, SS80a, Sch81, Sch83a, SS80c, SNNO86, Sin88, ST82, SKL80, TF83a, TTS84, TPN86, TP87, Toy80, Tru83b, WCLZ86, Wal88, WJP83, WJP85, YOJ81, ZY85, dSdMB86, BDK86, Zun85, Ho81, Fro83a]. electronically [GC84f]. electronique [OF80]. Electrons [Str82, WS79, BW85, DT81, DRS88, JTST87, KM85, Kon83, KTS86, Mar88b, TST89, SvG89, Cal81b, Pic83a]. electrophilic [JL84, LY83, VŻH82]. Electrostatic [BCV82, FW80a, HKP81, NS82a, NSN89, Pee83, PP80, RS82, TH88, WWD88, BSZ81, BEL89, BV86, CLP82, CPZS86, Eis88, GM86c, GKAP81, HLKK82, HKLM85, HR87, HPKW81, JW86, JEW88, LCDP80, LCP81, LP81a, LPW83, LPW81, LHK81, Mez85b, NÁNSP87, NS83, PP86, RNW86, RL86, RL88, RBSR84, RSBR85, Sen85, SHKP80, SHP^{+}80, THKM81, ZLP81, ZP82]. Electrostatics [LN83, NS89, Ric88, FD89c, Ras88, RP89]. electrotatic [CLP83]. element [FSW87, GC84c, Gou86, Pan84, Pay82]. elementary [AL80, Ben82, DBL84]. elements [AW83, BPLB87, Boč89, BFJ81, Cha82, CMB84, DOW89, DH84, Ess82, FF80, GB85b, GC84e, GP86a, JIS87, Kau82a, KS82, Lim89, MZ85, Niu80, Pan85, Pur88, QX89, Ret86, Tay85, WP80, Wor85, Cal81a]. Elimination [BL81a, BBLR84]. Ellis [Muj86]. Elsevier [Ågr80, Ågr87c, Cal86a, Pic83b, Tap88]. embedded [EG84]. emission [DPZ80, GC84f, Nes85, UT80]. emitting [ST86b]. Empirical [Sen85, HH89, OMCD81, Sad85, SL87, dB84, Muj86]. EMS [FPDA89]. enamines [Yam81]. enclosed [FC81b, FC81c, FC81d, FC82e, Swe83]. end [KG80, SI89b]. ended [DBF89]. endoplasmic [SAB^{+}80c]. endorphins [LCP^{+}88]. energetic [HKK^{+}82, HKKM83, HKLM85, OPSO86, OPSO87, THKM81]. energetical [DBES80]. Energetics [LS86, Sch83a, KS86c]. énergie [OF80]. Energies [GBO86b, Hub85, Ada82, AJY80, All87, AHB^{+}80, Bec85, BF83a, BS86, Bri86, CPMP86, CMM83, Cla83a, Coo87, Das80, Das82, Del87, DS89b, DCF82a, FAC84, FC82d, Fli84, FF89, GSMRM85, Gol80, GOLC87, GY87, Gre81b, GKAP81, Has85, Has86, HČP88, Ho81, Kal86, KS85, KEIG86, Kau83, Kry82b, LCC84, Lás85, LHP80, LH87, MMM81, McD80, MMB84, Mez86b, MTO^{+}81, Mof81, Mof82, Nas83, NS82b, Ort89, PG80, Pee89, PD84b, PS80a, Pit84, PDMC82, PFLB83, PHGR88, RSG^{+}87, RBSR84, RSBR85, SAD81, SF81a, SJH89, SBWW80, TL83, XYSJ81, YOJ81, Har87]. Energy [Bal81a, BH89a, DIV86, GLP80, HDTR83, KG85, KG86c, KG86d, LCP^{+}88, RPS80, Sen80b, UTVP85, Ågr80, Ågr87b, ACP84, AGCT87a, AGCT87b, AT85, ADT^{+}89, AGB82, AYCD81, AM89, Ave84c, Bag83b, BZA84, BB83, BJ87, BPT88, Bec83, BSL81, BH86, BMRJ86, BSL^{+}80, BS83a, BP80, BPCT83, BCT86, BBL81, CEB86, CPS81, Cas85, CLVC84, CG86, CLS81, COS82, CSK87, Cio87c, CB88, Cio88b, ČD87c, CSF80, CFLM85, CSB86b, Csa81, Csa85, Csa86a, Csa87, CE88, Csa89, CZ82b, DP87, DSG86, Dea89, Don81, DC86, EB81, ES88, FPT81, FN84, FL83, FZB86, Fli86, Fra84, Fra85, FSK79, FSS^{+}89, Fut87, GS89, GTB80, GPLP85, GA82, GCP82, GCV84, HF86, Hal88, Han88, Hel82, HZBR89, Hir80b, HMF82, HG86, HDO88, HSB88, Hub80, JS88, JŻ85]. energy [JJČ87, JSS80, JS86, KH86, Kat80, Kau83, KU80, KNM83, KB86, Kon83, KD80b, Kry84b, Kry84a, KPS87b, KPS88, KTS86, KEGS81, LB86, Lev89, LO81, Lim82, LÖ85a, LCD82, LY83, LN89a, LVL^{+}85, Mar85b, MC88, MC80, Mez81a, Mez82, Mez83d, Mez86a, Mic81a, Mic83a, MHAS86, MW87, MS80b, Mül81, NHKY83, Nal80, NSS86, Nom80, NCKL81, NC84, Obc87, OF80, Ori82d, PKG86, PČP82, PPMG89, PD89, RSSS81, Ran80a, RTM81, RTM83, RMM83, RTM86, RB89, RBH^{+}86, RVK80, RS88, SO85, SKR81, SR81, SSR^{+}87, SB84, Sel88, ST86a, She87, SHKM80, SS86a, SHRH82, Sil82, Sla85, SB83b, SRL^{+}83, SRHK83, SR87, SHK87a, SG85a, ST86c, Swe83, SR80, TNY86, Tay86a, Tay86b, Ten86b, Tho86a, TP80, TOK80, UB86, Umr89, Urr88, Var87, VM83, WSD80, YRS85]. energy [YS80a, ZY85, ZAB85, ZHB^{+}87]. Energy-band [HDTR83]. Energy-Conformational [LCP^{+}88]. Energy-eigenvalue [Sen80b]. energy-partitioned [Kau83]. engineering [CCC^{+}89, GSS85a]. England [Ågr85]. enhanced [TW88]. enhancement [NS82a]. enhancements [ES81]. enneagonaux [OG86]. enol [PSS86]. ensemble [Per82, RJK83, SV85a]. ensure [Kau82a]. entering [FN83]. enthalpy [PL83, PL84b]. entrapment [GP85]. Entropic [Urr87, LJHU89]. entropy [Fra85, GB85a, LJHU89, Tac89, UV83]. enumeration [Ran80c]. envelopes [LP81a]. environment [NSP80, PWL89, RKHK88, RKO^{+}89, SAD81, SRSR85]. Environmental [Mar81a, CG80, MFM89, PTG80]. enzymatic [Ben81, FG82, Jac89, LPW83, NSP80, SHH^{+}83a]. enzyme [AABW80, Gas85, RN84, SS80e, TB83, VWCC82]. enzyme-inhibitory [AABW80]. enzymes [Sey81, Sok81, VB85]. EOM [Bak85, GY87]. epoxidations [CD89c]. epoxide [BP84]. epoxides [KHRK87, LPCF80, PL84a]. equal [GC84f, Jon81c]. equalization [Loh84]. equation [AN85, Au82, ASL88, BH86, Bot81, CP86a, ČV86, CTV80, Csa80, CT81, Csa86b, Dat85, Den83, Den85, FCO89, FCB80, FM83, Ham82, IRA87, KK86, LPS83b, Lim80, LL82b, Mar86a, Mas87b, MM80, Ori84b, RI88, SCH84, SC86a, SC87, Ská83, SV82, Tac82, Tel88, Toy80, Ulm85, VTT81, WB89, SV82]. Equations [STV80, AB84, BEP87b, BHM87, CSF80, Coh86a, DSB84, DL83, ERB82, HH87, IPV81, Kob82, Kre81, LPS83a, LPS83b, Mar88a, MC80, MDT86, MDT87, PS87, PP89, RB88a, Sch89, SIJ85, SF80a, SV85a, SV85b, SvG89, SC84b, Tat81b, Tru83a, Van86, FB83, SV85a, SV85b, SvG89, Van86]. equations-of-motion [SF80a]. equidistant [KKA80b]. Equilbrium [DDFA87]. equilibria [CA86, MP87]. Equilibrium [DBD^{+}85, AGCT87b, CHKB83, DKLP81, DC82a, GRP^{+}80, KL80, KB87a, KB87b, MK83, Tho80, HM84]. equivalence [FF82]. ergodography [ZCH89]. Ermler [Sie83a]. errata [Sab80a]. Erratum [Ano83b, Ano86d, Ano88h, BC85a, CGT86a, Con88a, FF85a, Ish88a, KG86a, LB81, Mes81, Nal81, RM88, VU88, Vol89a]. error [Cas85, Fra88, LC81]. errors [AGCT87a]. ESCA [LdMH80, MPP83, PGG80]. Escherichia [FPDA89]. espace [TV83b]. ESR [LCCW87]. ester [ST88]. estimate [TL83]. estimates [PČP82, Sen84, Sen85]. Estimating [Har82, EB82a, EB82b, EB82c]. Estimation [GM81a, MASF88, RSSS81, FM86, KHNT86, LCD82, WHT86, LC81]. estrogens [DA81]. etats [TV89, TV83a, TV83b, TV84, Van85, VT85, Van86]. ethane [Bak85]. ethanolamine [ŠHH83b]. ethers [BNS^{+}87, KHRK87, ÖHD89]. ethylene [Bak85, CLL86, CKM84, GG82b, GDX89, HSR86, JL84, LRZ84b, LR89, SS86c, SS86d, WBK86]. ethylenes [PH82]. ethynyl [SP82a]. ETO [WJ82, Sme84]. étude [TV83a, TV83b, TV84, TV89, Van85, VT85, Van86, OG86, SvG89]. Eugene [Mou89]. Euler [Bha82b]. eupeus [Gri81]. eutectics [Mat80]. eV [GSRG80]. evaluate [DFP80, DAPS81]. evaluated [AW83]. evaluating [PFBF81, Ran80d]. Evaluation [DA80, DFH84, GM81c, Gus85, Har83b, Jon82, Luk82, Mic81b, Ori82a, Ori82d, SF80b, BL81a, DD87a, ECM89, Gla85, GB85b, GC84c, Gou86, Gus87, HV83, HA84, JL85, JEW88, KT86, MPBN81, OS89b, Pan84, Pan85, Pul83, Ret86, RR88, SSR^{+}87, SB87, She87, YOJ81, AØ80a, Pan83]. even [Big86, CHKB83]. Evidence [Frö83b, Ker81, SRS81, Wed81]. evolution [FM81, FHH^{+}87, MS84, SM83, TV83b, Van86, Wil89]. evolutionary [Fox86]. evolved [Fox84]. EWMO [DL81]. Exact [DPR80, Has86, HL81, Jon81c, KS89a, MM87b, PD80, BHM87, HT85, Hun86, LP85, TH81a]. Exact-exchange [DPR80, PD80]. exactly [Cio87c, CF83, CL85, Has85]. examination [Ada64, AGCT87a, BH82, BCT86, CZ88, LC80, LLM86, SL87]. example [CCD^{+}84, SBB87, Sim80, TP80]. examples [ACGT89, ČV88, Ela87, GDBRV87]. Excess [Lim82, Sto82b]. Exchange [ACP84, ABTP83, Kle87, Bec83, Bec85, BBMP82, Cof83, Csa88, Csa89, Das80, DPR80, GKAP81, GKP83, GCV84, Hal80, JJČ87, Jon84, Jon85, Lev89, Nag87, New80, PD80, RBL84, Shi84, UB85, VV83]. exchange-correlation [Bec85, Lev89, Nag87]. exchange-energy [Csa89]. exchange-repulsion [GCV84]. Excimer [UT80]. exciplexes [Hir80a]. excitable [LN83]. Excitation [GY87, Woo82, AYBP83, AJY80, BP86, DP87, Frö83b, Gan81a, GSO^{+}87, Has86, IL84, Kal86, KCP80, LH87, LS82, MDT85, MTO^{+}81, Nak83, NS82b, Pee89, PD84b, PD89, PSBB83, SSR^{+}87, SB84, Shi87, Tay83, Tay84, Tay86a, Tay86b, TL83, VP84, WK85, YOJ81]. Excitations [BFC80, Cul88, Cul89, DL81, EIMS88, FY82, GBO86b, KMRK86, KAM85, MC89, MR81, MR80b, Nes85, PW86]. Excited [BPT83, EIMS88, KK87, AM83a, AB85, AEM86, ANH81, BB85, BKP83, Bue86, CHKB83, CS80b, Cha80, CHZ82, CJ89, CKM84, DMB89, DBVMB85, DF83a, DWQ86, Flu85, Gub82, KH85, Kat83, Kau82a, KS81, LFL^{+}81, Lip88, MGN82, MMS^{+}86, MK80, Mül81, NH81, NB79, PP89, SMD83, SBE87, SNNO86, TOK80, Vol89b, YUKK84, BEP87a, MS84, NB79, Gos80]. excited-state [SMD83, MS84]. Exciton [KBT80, ST86b, AHB^{+}80, AH86, PS80a, SAD81, Sin86, Suh84b]. exciton-phonon [Sin86]. excitonic [CCS84, FK83, Lit81, Suh86]. excitons [AHB^{+}80]. exclusion [Hub80]. exclusion-principle-violating [Hub80]. exemplified [FCPL84]. EXGEM [Røe82, Røe87]. existence [BD85]. existing [SM80]. expanded [JBW87, MT88]. expansion [BEO80, Cha82, ČV83a, Csa81, DVV80, DC82a, Ish84b, IBB85, IBB87, IS88, JJČ87, KS89a, KN86, Kut84, Löw83, MM87b, Mic81a, Muk86, NH81, Nak83, Nov83b, OY83, Pie82, PM86, RNW86, RL86, RLR88, SF80b, Tin80, Var87, VV83, GCP82]. expansions [BEL89, CS80b, Cio87c, GS80, HČP88, KSGB^{+}83, Kle86, MR81, MR80b, PČP82, Sil80, Sil86, Ten80a, TP89, Tim86, Van82, VM83, Vrs87a, Vrs87b]. Expectation [Küm83]. experiment [Ågr87a, Bar85, Bel80b, DPZ80, DWQ86, LM81, Pul79, Ska84]. Experimental [Ker81, KPSS87, MP84a, PSS^{+}87, TNT^{+}86, Wed81, Bri86, Cal82a, Chr80a, Chr80b, CFG^{+}83, FNPS82, FF89, MWO^{+}84, NB79, Sin80b, TLT^{+}84, Woo80, SCER88, SSN^{+}84]. experiments [AM87a, All87, CD89a, DD87a]. Explaining [Rec80]. Explanation [SHPK80]. Explicit [HM89b, HH87, KDA81, LP89, Lim89, ČVV87, GCP86, KPP85, KPP86, Mor89, Ade87]. Explicitly [KC82c, KC82d, TB81]. exploration [Sok81, You84]. Exploratory [KC82d]. Exponential [BVAB81, Dem83, PS80b, Wit81]. exponential-type [PS80b]. exponentially [Kau82b]. expression [KDA81, KK82a, SM85, SV85b, SV85b]. expressions [AMFC83, DIV86]. Extended [AEM86, Bak85, BDK86, Del84b, DFH84, DH85, Gos82a, GPC85, HLPV83, INFU84, IQ87, KT84, MD87, May86c, MM81, RSSS81, Røe81b, Røe82, Røe87, TPČ83, Kun86, LHP80, VP84]. extended-Hückel [LHP80, VP84]. Extended-Koopmans-theorem [AEM86]. Extension [DL81, FC82d, Kar85, SR87, GM83, MBD84]. extensivity [BBO83, MJ89]. exterior [OI86, Sim80]. External [NH84, KM89a, KS86c, PV89, SS85, SD88]. externally [Rab84, Sad83, Sie83b]. extracting [Bha82b]. extracule [TM81]. Extrapolation [CB88, Bur83, FSW87, SBB87]. Extraterrestrial [You84]. extreme [Lar86a]. eye [Var81a]. Eyring [Urr82].

F [Ågr85, Che80, CH82, FD89b, KU80, Mår82, PFLB83, RDG^{+}86, JMS^{+}85, KEIG86, BF83a, Ish86, Ish88b, KGEP89, KG89, KU80, LFL^{+}81, MK83, RDG^{+}86, SJ85, STG86, YKZ^{+}89, KG86c, SMJ83]. face [JA82]. Factor [MH80b, MH89, Mal91, Mal02, BG81, Dem83, Fra89, Gru85, JL85, KB86, Wol81]. Factoring [RBK85]. factorization [Moh80a, Pay82]. Factors [LNN84, SRH86, Che82, CFG^{+}83, FGS85, Fri86, GR89, Gla85, HYZ89, KDA81, KAA85, LHC80, PM83, Pia83, RB88c, SPRS89, SIN80a, TT83, YRS85, YLM80, PL84a]. FAD [PSG83]. Faddeev [KM89b]. fails [TWW81]. Failure [Khu82, BH82]. FAKE [HTD80]. Faraday [DQ80, DQ81]. fast [JTST87, TST89]. fatty [DC82b]. fcc [Bag83b]. FCN [CP83a]. Fe [AL87, CJZ^{+}86, DK87, GLLOB86a, GLLOB86b, GLLOB88, KO81, SK87]. Fe-ligand [SK87]. Fe-porphine [KO81]. Fe-porphyrin [SK87]. features [CZM83, DMB86, DMB89, HDFL84, Mil86]. FeCO [BVD81]. Feenberg [Bha82a]. Feinberg [Ish85, Har80, Ish86, Ish88b, Ish89]. fentanyl [CL83a]. Fermi [CT81, Csa81, DCF82b, BPT88, CP80a, CTV80, Csa80, Ess82, FGC84, FD89a, Ish85, Lud83, LC82b, Ols86, Ori81a, Ten86a, Ten86b, UB85, WHT86]. Fermi-contact [Ish85]. fermion [Cas88, DK81a, DK81b, Lar86a, Mat87d, MK81, Wei88]. fermions [SBW86]. ferric [WCLZ86]. ferrocyanide [SA86a]. few [LRT88, MM87b, Pul86, Ten80a]. few-particle [MM87b]. Feynman [ESS86a, FC82b, GN87, Kva82, Sil84]. FH [Del87, DF89, ZW83]. Fiber [Cho80]. Field [KFN80, Var85, ADF^{+}88, ACCCBF80, AØM81, BB87, Cio89, CHB88, DG86, Den83, Den85, Eng83a, Eng83b, FCPL84, FGMM80, HS84b, HC80, HR87, IS80, Jac80, KKZ86, KS84, Khu82, KG86d, LRT88, LRZ84b, Lef83, Lib84, LLC83, MC88, MN81, MMC83, MB85, MA80b, NH84, NP81, NS83, Noa80a, Noa80b, PMCA88, Rab84, RPS82, Rér89, Rom81, RVK80, SRS81, SSSH82, ST80b, SCH84, SC86a, SC87, SL87, SP82b, SHLL83, Zhe83]. field-induced [MMC83]. field-theoretic [Eng83a, Eng83b]. fields [Ban82b, DS80c, FVZ^{+}88, FC85, HCCB86, LM89, LOGC86, MR85, NÁNSP87, Pac83, Poh80, Poh84, Ran81, SH89]. Fifth [Cul88, Cul89, CZ88, Nec83]. Fifth-order [Cul88, Cul89]. Fifty [Coh83]. filing [UTVP85]. filling [Ten86a]. films [MTI80, WD81]. filtering [TN89]. Fim [Kle83]. Final [HM86, KS81]. find [Löw86e]. Fine [BN84, SBH^{+}85, SBH^{+}86]. finely [Pes83]. Finite [Kir89, MMC83, BD85, CE86, Gos82a, HČP88, HL81, IPV81, Mey88, RNW86, RS84, TP84, TNY83, TK86, PT84]. finite-difference [IPV81]. Finite-field [MMC83]. finite-order [TP84, PT84]. First [JD85, Suh84b, AEM86, BN84, Bec85, BF83a, Boč89, BST86, DBS82, FN84, GTB80, Ham82, Kau82b, Löw86d, LL85, MB84, Mül81, NC84, SSC83, Suh86, NHKY83]. first- [DBS82, FN84, NC84]. first-group [Boč89]. first-order [Ham82, Löw86d]. first-row [Bec85, BF83a, Kau82b]. Fischer [Cal86d]. fit [ZW83]. fitted [CFLM85]. Fitting [HS84a, JS88, LB86, RSST86, ST86c]. five [Ano87t, Löw85b, SK87]. five-coordinate [SK87]. Fixed [MT87, GBZ86]. flavin [TF80, TF82]. flavins [HOT87a, HOT87b]. flavodoxin [TF80]. flavone [TH88]. flavones [BLT^{+}84]. Fleming [Rat87]. flexibility [FCPL84, KG80]. flexible [AM88b, OMCD81]. Floating [BR80, Hur82, SBE^{+}83]. Florida [Fax87, Han86, WJ82]. flow [Car82, SHRH82]. Fluctuation [Kub82]. fluctuations [CD83, CD85, PV89, SB88]. Fluid [Fis82, GD82b, MD82, Sti82]. fluidity [HJ80]. fluids [Per82]. Fluorescence [MTI80, CD83, Kat80, PZPL86, MGBH81]. fluoride [GjC86, PCA89]. Fluorine [CS87, RS82]. fluorofuranose [LMG85]. fluorouracil [Zie82]. fluxes [MD82]. Fock [FVZ^{+}88, Has82c, LSP85, NS84, SP85, TL80, TL82, TL83, Wag83, AN85, AJY80, Bec83, Böh83, BANF86, CDG80b, CK82, DMB^{+}84, DSB84, Del84b, DH85, DC87, DD87a, Den83, Den85, DPR80, Dov84, Dov86, ESS86a, Fin87, FCB80, Fuk81b, GM86a, Gol80, GA82, Hir80b, HM89a, IBB85, IBB87, IQ87, IS88, JJČ87, JSYO83, KC82a, KC82b, KC83, Kar86, KMW80, KT84, LHC80, Lás85, oLCC81, LFM89b, LÖ85b, MD87, McW83, Mes80, Mes85, MDT86, MDT87, MMC83, Noa80a, Noa80b, OJY83, Ort89, PPB87, PC83a, PCG83, Pan85, Pan89, Pau84, PABD80, PD80, PBA88, RI88, Roo80, SIJ85, STV80, SV82, SV85a, SV85b, SvG89, Tay87b, Ten88b, VVD80a, VBS85, Wag83, Yam80a, Zer89]. Fock-Dirac [AN85]. Fock-matrix-based [Roo80]. folate [SWA82]. folates [SWA82]. folded [LCP^{+}88]. folding [SRC84]. folds [SRDG85]. followed [CMM83]. FOOF [CS87]. forbidden [GC84f, Min80]. Force [NHKY83, CHB88, FVZ^{+}88, HCCB86, Kog84, LJHU89, MBSC81, MA80b, Sel86a, Sel86b, SL87, SS89a, Tho80, Urr87, WHT86]. forced [SMJ83]. forces [HFSB86, Ish83, Ish84a, MTLL82, Obc87, Tha83]. form [ČVV87, Fri86, Gla85, McW88]. formaldehyde [DF86, GDX89, SS86c, SS86d]. formalism [Böh84, DR84, DS88, DQT80, Ess84, FF80, KL87, PM84b, Tim86, Wor85]. Formalisms [New80, KS81]. formamide [DDFA87, OL80, WHT86]. Formation [BW85, BBL81, CLL86, Emi83, Eng83a, Eng83b, Kun86, LWZ88, LGL^{+}84, MGBH81, OLB^{+}82, Par84, PK88, SBE87, Sey80, TF83b, Wil89]. formhydroxamic [HM84]. formic [DDFA87, MR85, OLB^{+}82, PMCA88, ŠHH83b, WHT86]. forms [Bli80, BCD^{+}80, CLP82, FD89a, HPKW81, LWZ88, POT80, YLM80, Zie82]. formula [BL81a, KKTE80, Pau89a, Sil86]. Formulas [Jon84, Avr82, FC82b, GS80, Jon80, Jon81b, Jon81a, Jon81c, Jon83, Jon85, Jon86, Lim89, LL82b, PSM87, Sin88, Ish83]. formulated [Löw82b]. Formulation [RC80b, AZY84, Bel80a, BS81b, BVAB81, Gre80, Kel86, Mar88a, Mat87d, Mez84b, MP84b, PM84b, Pul83, Roo80, RM84, Tin80]. formulations [GF83]. Forssman [YL83]. forward [Kon83]. FOTOS [NT80]. Foundations [Suc84, DBS80]. Four [Kry84a, CC84b, EB82b, Jon86, Khu82, LCULC87, MBKH83, Niu84a, GKMV80]. four-body [CC84b]. four-center [Jon86, GKMV80]. Four-dimensional [Kry84a, Khu82, Niu84a]. four-electron [LCULC87]. Fourier [GKMV80, MPP83, Niu84a, Niu84b]. Fourier-transform [GKMV80]. Fourth [CZ82a, KU80, Wil80b, CZ82b, MPD87, Nec83, Ort88, Ort89, PHGR88, Vas82]. Fourth-Order [CZ82a, KU80, Wil80b, MPD87, Ort89]. FPT [FC80a].

Fractal [Fin89]. fraction [SS80b]. fractional [FW80b]. fractions [ČV83a, Mas87b]. Fraga [Ågr80]. fragment [GSC89, GG81, GG83, SS82, SC80b]. fragmentation [RGW88]. fragments [BVAB81, GG81, Hib81, Kle83, KKD80, LM82b, NÁNSP87, Ott86, SFC86, SC80b]. fragments-in-molecules [Kle83]. frame [DM85, FGM87, Wea83]. frame-transformation [Wea83]. frames [KST84]. framework [BB82a, Böh82, DP81, GCP82, GC83, KC83, Kry82a, LW82, LW84a, LW84b, LWK84, RM84]. Franck [HYZ89, Kat80, KDA81, PM83, PS88, SPRS89]. Frank [HYZ89]. Free [BMRJ86, GSMSG82, MN81, Obc87, Urr88, BD88, Bec89, CSK87, DRS88, GMP^{+}84a, Gas86a, Lim82, LN89a, Mat87b, Mat87c, Mat87d, MSV89, McW88, RSD82, Sen84, SST83, SAB^{+}80c]. freedom [HL88]. freeon [Mat87a, Mat87c]. French [OF80, OG86, SV82, SV85a, SV85b, SvG89, TV83a, TV83b, TV84, TV89, Van85, VT85, Van86]. Frenkel [AH86, SAD81]. frequencies [Blo89, CHB88, DVV80, FVZ^{+}88, MHAS86, PSK^{+}81, SSR^{+}87, SZR80, ZV85]. Frequency [HS84b, SHS84, ESCC85, GY87, MB85]. Frequency-dependent [SHS84]. Frictionally [SCH84]. Frozen [PD84b, LP87]. frozen-core [LP87]. FSGO [PL80]. fuel [PSG83]. full [AGCT87b, Bur83, Has82a, LER87, PPB87, PB82]. Fully [RI88, DD87a, HO85b, LPS83a, LPS83b, LSP85]. fulminic [SA86b]. function [AL80, Bha80, Böh82, BEP81, Bri87, CHA87, Chu88, Cla83a, DS80a, Dem83, DS88, DA81, FC84, Fra84, FH81, FH83b, GD82a, Gas85, GLLOB86b, Gri81, GBZ86, GBK86, Gru85, Hel82, HSB88, Hun80, IRP81, IO89, Jon84, JBW87, KR81, Kau82b, KMW80, KK82a, KPS87a, Leu83, Löw86d, Mic83a, MKW85, NH81, Nak83, OMCD81, OY83, Ori84b, PAFW84, Pau84, PKH^{+}81, SS80a, SSR^{+}87, See84, SL86, SDBDSMG86, SM87a, Tay81, UI80, UB86, Urr87, VM81, Win83]. functional [BPT88, Bec89, BK84a, Cio87b, CL85, Csa81, Csa85, Csa86a, Csa87, Csa89, Don81, FAC84, GV88, Har83a, JMD88, KL87, KGO89, Kry84a, KPS87a, KPS87b, KPS88, KWM86, LP82, Lev89, LÖ85a, LCG82a, LM82b, Mat82, Mez82, MS80a, Min87, Mor89, NDBE82, Per85, Per86, PSK86, Sav88, SL86, Smi89, Tac87, Tac88, Tac89, VCG86, VBS85, ZAB85]. functionals [CSF80, Csa89, KN86, LP85, Lie83, Sha85]. functioning [KPTG88, KPT88]. functions [AØ80b, Ave86, ALH86, Bli84, Boč89, BVC80, CFS86, CFS87, CG80, CP80a, DG86, DM83b, Dea89, DD87a, EI81, ESKG83, EN87, Eng83a, Eng83b, FC82c, Fra88, GM86a, GHGP89, GB85b, GBW88, GBW89, GJ83, Gos82a, Gre81b, GA82, HKP81, Has82a, Has82b, Has82c, Hib81, HČ83, Hun81, Hur82, ITKD80, IS88, JD83, JD85, JSSS80, JSSS82, JSYO83, Kah84, KPM83, Kau83, KT88, Kry84b, Lás85, Lib84, LPW81, Luk82, LLC83, MM82, MMN87, MZ85, Mar88c, Mat81b, May86a, MM87b, Mic81a, Mic85, Mic86, MT88, Moh80b, Moh80a, Mor88, MMC83, Niu80, Obc87, PB82, Pan82a, Pan82b, Pan83, Pan84, Pos83, RBL84, RDH84, Rut81, Sad89, SPÖ89, SM85, SIJ85, SD86, She87, Ská83, Sla85, SRL^{+}83, SP85, SM87b, Ten80a, Ten80b, TB81, TTRL88]. functions [VM83, WP80, Wor85]. Fundamental [Mar81a, Cia81a, PTG80]. Further [FC80b, RR88, SHL82, Hal88, DMB^{+}84, DMB89, FC82e, ZNS80]. fused [EB82b, Kin86, LK87].

G [Ågr88, Cal82a, Cal86a, Cal86d, Kum87, Tap88, Wol81, SRSR85]. Ga [CE86, CE87, CE88, GOLC87, LOGC86, CE86, CE87, CE88]. GaAs [CE86, CE87, CE88, BW85, CB82, CE88, GOLC87, LOGC86, SF89]. GaAs-A [LOGC86]. GaAs-Al [GOLC87]. GaAs-Ga [CE86]. GABA [AIJW82, BL81c, BVDA88, GM86c, LG84, PK88]. GABA-A [BVDA88]. GABA-transaminase [AIJW82]. GAGP [KE82]. gallium [PPB87, RZWL80]. Galois [HH87]. Gap [SC84b, Per85, Per86, VVD80b]. gaps [ABTP83, LK87, LH87]. Gas [BBLR84, Sie81, Bec83, CC84a, CADL86, Fla80, Han88, KKA^{+}84, KS87b, Ori82d, Set79, SS81b]. GaSb [BF81]. gaseous [GLB81, Wój86a, Wój86b]. gases [TSA86]. Gaston [Fro83a]. gating [LN83]. gauge [BYBZB83, SG87]. gauge-variational [SG87]. Gaussian [Ågr87c, AN85, ABCV86, Bag83b, BR80, BYBZB83, CFS86, CFS87, DPC82, GHGP89, GM81c, Har83a, IBB85, IBB87, IQ87, IS88, MT88, Niu80, PDMC80, PDC80, PDC81a, PDC81b, PKC85, RFST85, RC80b, STH80, TH80, Ten80a, Ten80b]. Gaussian-type [IS88, STH80, TH80]. Gaussians [KC82c, KC82d, RL88, TB81]. Ge [SNNG86]. Gegenbauer [Wea84, Wea88]. Gelboin [Mar81a]. Gel’fand [HR83, LP89, Mat81a]. Gell [DS88]. Gell-Mann [DS88]. geminal [EÖWZ84, ESS86a, Gos82a, KE82, Røe81b, Røe82, Røe87, SPÖ89, WK85]. geminals [DP81, JJZ^{+}88a, JJZ^{+}88b, Lar86a, OT80, Røe81a]. gene [FPDA89]. General [DÖ89, Kag83, Mil86, Özk84, Wei82, BR85, Chr89, CCC^{+}89, Cof83, GP86a, Gou86, JS88, Kau81, Kim81, Lan84, Mez84b, Muk86, SM85, TV83a, BVAB81, DR84, DQT80, Ess84]. générales [TV83a]. Generalization [RM82, Cho80, Kut80, Wea88, ZNS80, SC80b]. Generalized [Ade87, Bha82b, FM80, GN87, Gol80, KPP85, KPP86, MPBN81, Mor83, OFB81, PM86, SBNK89, Sil84, YKZ^{+}89, Bal82d, BFC80, BC84, BC85a, BC85b, CB88, Cio88a, FF82, HF86, KEB87, KE82, Lar83, MPP83, NP80, Rut81, SWM82]. generated [Gru85, HKP81, Jon80, Jon81b, Jon81a, Jon83, Jon85, Jon86, Kau81]. Generating [SM87b, KPM83]. Generation [KH85, LER87, Rut81, Sie81, Ada81, AM83a, CMB84, FSP87, GHGP89, HS84b, KSJT83, LRT88, ML81, MB80a, RB86, Set79]. Generator [BLV88, KRT83, DL83, DLVÖ84, Mat84, Mat87b, MT88]. generator-state [Mat87b]. generators [Duc85]. Genetic [FM80, FF84, Bed84, FF82, JKL85]. genomic [Fin89]. genotoxic [MFM89]. geometric [HG86, RB86]. Geometrical [She87, SJ84, LRT88, PBMR^{+}84, SB87, SB88, JS86, Ågr87b]. Geometrie [Van85]. Geometries [BDP^{+}81, HOT87a, HOT87b, Mil86, PDMC82, MG83]. Geometry [CH86, FM81, JMD88, WH87, BSZ81, BP80, CHKB83, DBD^{+}85, DDFA87, HOT87a, HGPF89, HM84, JM81, Mez83d, ÖHD89, PG80, Sey83b, SJRP88, SKS83, SS86b, Tho80, Van85]. Gerloch [Nat86]. Ghost [LC81]. Giuseppe [Fro83a]. given [Cio89, MM80]. glassy [BMMS81]. Global [Sto80, DR88, Mez86a, RTM81, RTM83, RMM83, RTM86, Tho86a, YG87]. globes [Mez86b]. globular [CGG81]. glow [Tat81a]. Glucocorticoids [KMKC88]. glucosamine [Yad83]. glutathione [LLM86]. gly [CBV85]. gly-pro-R [CBV85]. gly-R [CBV85]. glycine [Bur85, LGL^{+}84, OLB^{+}82, RTM83]. glycosides [TK83]. glycylcarbonyl [Tew87]. glyoxal [OL80, TWW81]. glyoxalase [BBAT82, BLT^{+}84, BLY^{+}86, TB83]. glyoxalase-I [TB83]. Goldman [Gee89]. Goldstone [Kva82]. good [Ran80d]. Gordan [Alt89, Her84, Kib83]. governing [LNN84]. Gradient [VV83, Csa81, DMB89, Hel82, IS80, KN86, MA80b, Pul83, ST80b, PDC81b]. gradients [ALB84, PC86, SSR^{+}87]. Graham [Rat87]. gramicidin [UVP^{+}81, UTVP85, VARPU84]. grand [Buf82, Per82]. Graph [Kin86, RW80a, Ran81, Zho88, BBR82, Cio88a, EB82a, EB82b, EB82c, Gro85, GDR85, GS80, Kia81, Kle86, KT80b, KN86, KJT85, MBT80, Ori82d, RZ86, RSNT88, ST83, TKR86, VV87, Wen86]. graph-theoretic [Kle86]. Graph-theoretical [RW80a, EB82a, EB82b, EB82c, RSNT88]. Graphical [KS86a, Ran80c, Was89, DK82b, KT86, LSBS81]. graphics [Bri84, RP89]. graphite [DPR80, LL86, NOMK80, TR82]. graphs [Bal82c, BBR82, Flu84, Löw85a, Muk84, Ojh89, Ran80b, RK82, RD84]. Grassmann [Gru85]. gravitational [Rec80]. gravity [ST80a]. greater [DK82a]. Greece [Kel78, NB79]. green [KKTE80, Bli84, Böh82, Bri87, EI81, EN87, HSB88, IRP81, MMN87, See84, UI80, Win83]. Green’s [Cla83a]. Green’s-function-type [Cla83a]. grids [IAAA89]. Griffin [MDT86, MDT87]. groove [KMRK86]. Ground [AL87, EHK83, KS85, AM83a, AEM86, ACP84, Bec82, BH86, Bha80, BSR80, CK82, CHKB83, CV85, ČV86, DF83a, ESKG83, HČP88, HLR80, Ho81, HT85, JJZ^{+}88a, JSS80, JSSS80, KC82a, KAD81, Kat83, Kau82a, KS89a, KEGS81, Lar86a, LFL^{+}81, LC88a, Lip88, MG83, RBH^{+}86, TH81a, Tay81, Tay86a, Tay86b, Ten88b, Toy88, Yam82, Živ86, Živ87c, TS81]. Ground-state [EHK83, KS85, ACP84, BSR80, CV85, HČP88, HLR80, RBH^{+}86, Tay86b]. Group [Kva82, Mez85b, Mez87a, Abd80a, ACCYCC80, BDP^{+}81, BD85, BS82b, Boč89, BS81b, Cal86b, CNK87, DH84, Don79, DK82b, Duc85, FS80, GCP87, GC84b, GC84c, GC84d, GC84e, GC85b, GP86a, Gou86, GC87, GRBP87, Her84, HDO88, HL80, Kat89, KAH80, KS82, KS84, KS86a, LZ89, LP89, LSBS81, LHP80, Loh84, Luk81, Mat87a, Mat87d, Mez85a, Mez85c, MP84b, Nec83, Pay82, PM84b, Qia83, RSD82, Ret86, SS80d, SG89, Tai89, Wei82, WP80, Zhe83]. group-IV [LHP80]. Groups [Flu84, Alt86, Alt05, Bal82c, BR82, FW80a, FW80b, FS80, Kib83, KAA85, Kry82a, Mey88, SHLL83, Rös87]. growing [Poh80]. GTO [Har83b, MPP83]. guanidine [MTWO83, SSS81]. guanidines [SSS81]. guanidinium [SR84b]. Guanine [SJ86, FS87, Gre82, GP85, HB82, PSS86, SS85]. guanosine [Gre82]. guanylhistamine [MTWO83]. Guessing [Eri89]. guide [HS89].

GVB [CHKB83]. Györgyi [Fox87, Lad87].

H [Ågr82, BLR81, BDK86, Bow86, CEB86, CF81b, Dat82, DBS82, DSB84, Dun80, ES81, GT86, GLLOB86a, GLLOB86b, GLLOB88, GBO86a, KU80, Kog84, KBT80, KM89b, LS86, LB83, LWZ88, LH85, Mar81a, Mou89, PZ84, SOS84a, SJ85, Sch80a, Sch81, Sch83a, SBB87, SHP^{+}80, SS87, TST89, VŻH82, YKZ^{+}89, JY81, AM82, BNK86, BD81, Bow86, CS80b, CD89a, CS83c, Con87, CP85b, DPC82, DBS82, DF89, DLVÖ84, DYD89, DWQ86, EY82, ES81, ESCC85, GT86, GT87a, GA82, Gub82, HKV85, HDFL84, Ing80, Ish86, Ish88b, INO88, JŻ85, JRK84, KH85, Kog84, KTS86, KM89b, LPS83b, LB86, LGL^{+}84, MTWO83, NPL89a, NT80, Nov83a, NP80, PMOW89, PR83, PDMC82, PDKC84, Rér89, RBL84, RI88, RKHK88, RKO^{+}89, SJ88, SS84a, SS85, SBE87, SD88, SG81a]. H [SS87, TL80, Vog86, Wea83, Wój86a, Wój86b, YKZ^{+}89, ZJŻ88, CPS81, Con87]. H-bond [SS85]. H-bonded [SD88]. H-bonding [SS84a]. H-chemical [Vog86]. half [DL82, Gir86]. half-collision [Gir86]. halide [KS87b, PGM80]. halide-amine [KS87b]. halides [BP80, DF86, EIKG^{+}88, KEIG86]. Hall [CP80b, Ten88a]. haloacetic [CAML87]. Halogenated [PL84a, LL85, LLM86, RS82]. HAM [ÅFL81, Tak86]. HAM/3 [Tak86]. Hamiltonian [AW83, BB82b, BG88, Bha81b, BK82, Böh82, CMB84, CA80, FWS^{+}85, FN83, GMM84, GF83, GHF84, Has82a, HL88, HS88, Ish84b, Kat83, KL89, Kva83, Lim89, May83, Tus86]. Hamiltonians [GC84e, KSGB^{+}83, Löw82a, PL89, RS84, RB88a, SMMM87, Wes81, AČP82a]. Handbook [PKC85, Ågr87c]. handed [JY81, Vol88]. handedness [SW81]. handling [OY83]. hard [HHT82]. hardness [GBZ89, NKZ88]. hardware [SMS84]. Harmonic [PM83, AMFC83, AL86, Blo89, Bri87, DÖ89, FVZ^{+}88, HS84b, MB80a, MPB84, MPS86, Özk84, PSM87, PD83, SSR^{+}87, SV82, TS82, Wes82, CF81a, GC84f]. harmonically [PD83]. harmonics [AW82, AØC82, KK82a, LE80, Niu84a]. harmonique [SV82]. Harriman [Ish85]. Hartree [Has82c, PC83a, PCG83, Pau84, AN85, AJY80, Bec83, Böh83, BANF86, CK82, DSB84, Del84b, DH85, DC87, DD87a, Den83, Den85, DPR80, Dov84, Dov86, FVZ^{+}88, Fin87, FCB80, Fuk81b, GM86a, Gol80, GA82, Hir80b, HM89a, JJČ87, JSYO83, KC82a, KC82b, KC83, Kar86, KT84, LSP85, LHC80, Lás85, LJ82, Lim87, oLCC81, LFM89b, LÖ85b, McW83, Mes80, Mes85, MDT86, MDT87, MMC83, NS84, Noa80a, Noa80b, OJY83, Ort89, PPB87, Pau89b, PABD80, PD80, PBA88, RI88, STV80, SV82, SV85a, SV85b, SvG89, SP85, Tay87b, Ten88b, TL80, TL82, TL83, VVD80a, VBS85, Wag83, Yam80a, Zer89]. Hartree-Slater [Gol80]. Hasbrouck [Wil81]. having [EB82b, Kin86]. HC [LC88b]. HCl [KS89b, SBWW80, Wój86a, Yar87]. HCN [Cha88, COS82, CP80a, TTRL88]. HCOOH [SD88]. HD [GT87a, ESCC85]. HDMN [ST82]. HDO [Shi87]. headpiece [KRMK86]. Heats [Živ83, GTB80]. Heavy [SBW86, BDK86, Das80, Dur86, Gás86b, HBC86, Köh89, Mar85b, MMM81, Min80]. heavy-atom [Min80]. heavy-ion [Köh89]. heavy-metal [BDK86]. heavy-molecule [Gás86b]. HeH [LPS83b]. HeHe [JJČ87]. height [RBL84, SS86c, SS86d]. Heisenberg [KSGB^{+}83, LL82b, SMMM87]. held [ADL78, Kel78, NB79, Pul79, PTG80, WJ82]. helical [Bal81a, ITKD80, JY81, UV83, VPK80]. helices [GSS85b, Ott86, PABD80, SW81, SSYR87]. helium [BB81a, DS80a, DM83b, Dea89, DS81c, Flu85, Har87, HT85, IST89, JSSS80, KWJ83, LCLC88, Lás85, SB83a, TH81a]. helium-trapped [IST89]. heliumlike [DS81c]. helix [BST86, CLP83, CFC83, KBR83, LCDP80, MT85, Sar82, Vol88, vDTBN80]. Hellmann [FC82b, GN87, Sil84]. heme [LCCW87]. hemoglobin [SG81a]. Henry [Ågr85, Urr82]. heptamer [BFOP83]. Herbert [Sme84]. Hermitean [Röm83]. Hermitian [ACH83, CA80, Kva83, RSMM86, SLT84a, SLT84b, ZS86]. Herndon [KS89a]. hero [Fra89]. hetero [AEHNEA85]. hetero-atoms [AEHNEA85]. heteroatoms [TIS^{+}88]. heterobicyclic [Gre87]. heteroboranes [TL86]. heteroconjugated [LK87]. heterocycles [AEHNEA85, DA80, HPKW81, KKA^{+}83]. heterocyclic [PBKB89]. heterogeneous [RN84, Sim86, Sla83a, Sla84]. heteronuclear [PM86]. heterostructures [GOLC87]. hexachloroiridates [CPGP83]. hexagonal [DPR80, MB89]. hexanitride [TWL82]. HF [INFU84, DG85, GLB81, HKV85, Hil81a, KU80, KTS86, MMC83, RSMM86, SJ88, SNNG86, SG89, VŻH82, YKZ^{+}89]. HgTe [HLPV83]. Hierarchy [Coh86a]. Higgins [BR82]. High [CKK86, Fle81, Moh80b, Moh80a, YGK^{+}89, Csa80, Csa82, FZB86, Gee89, Gol88, GBW88, HZL84, KB86, Kon83, KTS86, LSW89, LRT88, McM86, Pac82, Sch88, SHKM80, SS88, ZV85, Kel78, Dun80]. High-affinity [YGK^{+}89]. high-energy [FZB86, Kon83, KTS86, SHKM80]. High-resolution [CKK86, Gee89, Gol88]. high-salt [Pac82]. High-temperature [Fle81]. Higher [Bry80, Bha82a, OJY83, Toy88, Avr82, GG82a]. higher-order [Bha82a, OJY83]. Highly [DL83, BCBS81, DBVMB85, Ker86, SBE87, YUKK84]. Hijmans [RM82]. Hilbert [Nag82a]. Hill [MDT86, MDT87]. Hiller [Har80, Ish85, Ish86, Ish88b, Ish89]. Hinchliffe [Cal88]. Hindered [BTC^{+}83]. His-64asp-32 [NSP80]. histamine [MTWO83, PMOW89, RMM83]. histamines [XYSJ81]. histidine [MSM81]. historic [Löw85b]. HLi [ZCH89]. HMX [HKK^{+}82]. HNO [CHKB83, Nom80]. HNSi [SRP81]. Hoffmann [CD89c]. Hohenberg [Kel86, Kry80, LP82, LM82b]. Hole [BBH87, Ågr83, BB85, CF83, CL86, DL81, GBK86, IL84, KD80a, Lar83, LC82b, Mat81a, NMK84, Odd82, Sto82a, UB85, Wor85, Živ88]. hole-particle [Wor85]. holes [CR81, CP80a, CP83a]. Holland [Dun80, Mic87, Sme84]. HOMO [EIMS88]. homoallylic [MR80c]. Homogeneity [Kry82b]. homogeneous [Han88, Ori82d]. homology [SRDG85]. homonuclear [PD84a, PPMG89, SM85, WS80b]. homopolar [Has86]. Homotetrahedryl [Ran80b]. homotopy [Mez83c]. hormone [RTM86]. hormones [KMKC88, VT82]. Horwitz [Gre81a]. Horwood [Muj86]. Howery [Wol81]. HP [BB85]. HS [TL80]. HSO [KKTE80]. HT/LSD [RWOT81]. Hubbard [FN83, Has85, Has86, HČP88, KAM85]. hückel [BV80, BDK86, Cio88b, HLPV83, KT86, KSSH85, LHP80, MM81, VP84]. Hückel- [KSSH85]. Human [And87, KR81, Löw84, Sey81, SG81a]. humidities [CC82b]. Hund [Bon81, BC83, WBB80]. Hunting [BL81b]. Hurley [Ano87s, Buc89, Mas87a]. Husimi [CHA87]. hybrid [Gus87, Her83, Jon81a, PM81, RW80b, RW81, Yan85, ZLH87]. hybridization [KGB87]. Hydrated [RT87, LN89b, MP84a, Nov83a]. Hydration [CFC83, LG84, Bar81, BEP81, CGG81, HRL88, HPKW81, LL89, MMB84, Ras88, Sok81, SSC83, TK83]. hydride [CPN89, Ver80]. hydrides [EB86, LFM89a, ST86a]. hydrocarbon [CPMP86, Che82, HJ80, PL84a, RN86, SL87, Ver80]. hydrocarbons [Cio87d, ČPV81, EB81, EB82c, EBEE87, GMM84, GBO86b, Has82b, KSJT83, LSH88, PDMC80, PDC80, PBMR^{+}84, Ran80a, RSG^{+}87, SVS81, Sey83b, Sey83c, SF80a, Sin88, Vog86]. hydrodynamic [SIN80a]. Hydrogen [BASB86, BLOW89, DAL89, JM81, KBS81, Nov83b, PG89, ACP84, AC85, AGCT87a, AGCT87b, ABD^{+}80b, ADF^{+}88, AMFC83, Aus84, ALH86, BBLR84, BP84, BBMP82, Bes88, BR80, Blo89, CPN89, CLS81, CV85, CF81b, Cze87, DK87, Del87, Del88, DS89b, DF86, DDFA87, FC80a, FC81a, FC85, FF85a, FF85b, FS86c, FDA83, FDAC89, GSR83, GP86b, GCP82, GMV83, Hur82, IHI80, IS80, KAH80, KD86, KS87b, Lal81, LCCB81, LS82, LP80, MGN82, MMS^{+}86, Mil86, MS80a, MC89, MH82, MB80a, MB82, Nov83a, PMCA88, PP81, PGPRN86, RSBR85, RS82, RLH86, SRS81, SFC86, Sch80a, Sch81, SD86, SIN80a, Sil82, SS86c, SS86d, TF83b, TT83, TNB86, TTRL88, UI80, Vrs87a, Vrs87b, WOT89, WD81, YS80a, ddF86, CCD^{+}84]. hydrogen-bonded [DDFA87, GCP82, IHI80, LP80, MC89, RSBR85, Sch80a, Sch81]. Hydrogen-bonding [JM81]. hydrogen-bridged [BR80]. Hydrogen-like [Nov83b, ALH86]. Hydrogenic [BVC80, AČP82a, BPT88, CB83, CE88, Tay81, Ten80a, Ten80b]. hydrogenic-type [AČP82a]. hydrogenlike [Ave86]. hydroloysis [ST88]. hydrolysis [SRC84]. hydrolytic [BLR83]. Hydrophile [GRP^{+}80]. hydrophobe [GRP^{+}80]. Hydrophobic [DS89a, YVK^{+}80, BMRJ86]. hydrophobicity [LSH88, SJH89]. hydroxy [BCD^{+}80, KŠHH83, Sch82, SK81]. hydroxy-2-aminotetralins [KŠHH83]. hydroxyacetone [KAH80]. hydroxycytosine [PKP^{+}81]. Hydroxydimethylnitrosamine [ST82]. hydroxyl [BDP^{+}81]. hydroxylase [Che82]. hydroxylazoaromatic [FZZ87]. hydroxytautomeric [Zie82]. hydroxytryptamine [MWO^{+}84]. hydroxyuracil [SZR80]. Hylleraas [Ho81, LCULC87, LC82a]. hyper [RDG84a]. hypercylindrical [SB89]. hyperfine [BN84, KGEP89, KG89, LCCW87, SS86b, SL80]. hypergeometric [Niu80]. Hyperpolarizability [GD89]. hyperpolarizabilty [SV85b]. hypersensitive [SMJ83]. Hyperspherical [Mez84c, NPL89a, AW82, AW83, ACDLH84, AZY84, ALH86, LÖ85a, LV87]. hypersurface [JŻ85, Mez81a, NHKY83]. hypersurfaces [Mez86a]. hypersusceptibilities [GMOBA80]. hyperthermal [VM81]. Hyperviral [FC82d]. Hypervirial [FC81b, FC81c, FC81d, FC83, FC82e, MPS86, ZHB^{+}87, CF81a, FC80b, Swe82, TS82]. hypervirial-scaling [FC80b]. Hypevirial [Sch80b]. hyponitrous [BMB87]. hypothesis [DBES80]. hysteresis [BB83].

I-102 [BV86]. Ian [Muj86]. ice [Kry86]. ICQC [Löw89]. Ideal [Var85]. idempotency [CFG^{+}83]. identical [KPP85, KPP86]. identification [BNS^{+}87, KSJT83, SMKT86, LCP^{+}88]. identifications [Dia86]. identities [Ish86, Ish88b, KPP85, KPP86, LE80]. identity [Har80, Ish85, Ish89]. IGLO [FGM87]. ignition [Bar86]. II [AAW82b, Böh84, EIMS87, GLLOB86b, RB88b, RGDF85, SRC84, VG80, AEHM85, AN85, Ano85c, Bal81a, BLR81, BEP87b, BL81c, BP80, Boč88a, BDK86, BVAB81, CK82, CD83, CR81, Cio89, CA86, CAML87, DDP81b, DPR80, DQ80, DK81b, EB82b, Eng83a, Faz87, FN84, FC81c, Fra85, Gan81a, GPLP85, GBW88, GC84d, GC85b, Gou86, GKMV80, GC82b, HOT87b, Har87, HKK^{+}82, HO85b, Hil81b, Hil85, Ish83, KC82d, KPM83, KHRK87, Kle83, KT80b, KD80b, KPT88, KPS87a, Kun86, LPS83b, LCULC87, Lar83, LHKK82, LW84a, LHP80, MMC83, MRJ85b, Niu84b, Noa80b, ÖL83, PCG83, Pan82b, Pan89, PDC80, PDKC84, PNA88, Pyy93, RDG84a, RT85, RDH84, RKO^{+}89, SOS84a, SS85, SCH84, Sey83c, Sil80, STS83, SRL^{+}83, SMHK89, Sto82b, SLT81, Tay86a]. II [Tay86b, TV83b, WS82, WJP85, WS80b, YVK^{+}80, Cal89a, Fro88b]. III [MK80, BSZ81, Boč88b, CPGP83, CA80, DS88, DQ81, EB82c, Eng83b, FC81a, Fau85, Faz87, FC81d, Fuk83, GBW89, GC84e, GC85b, GC83, HKKM83, Hil81c, Kib83, KG85, KPS87b, KPS88, LÖ85b, Pan83, PP83, PDC81a, Pyy00, Ran80b, RGDF85, SOS84b, SC80a, SIN80a, SMJ83, SV85a, TV84, TK86, WKH80, WP80, ZJŻ88]. Illustration [FDAC89, PD80]. Illustrative [Wes82]. images [And87]. imaginary [Rei82, Sil82]. ImH [vDTBN80]. imidazole [BKS87]. imine [OYN80]. Iminosilicon [SRP81]. immobilized [PSG83]. immunologically [YL83]. Impact [PS81, DSG86, Gan81a, Lad87, Lar83, MM87a]. implantation [RZWL80]. Implementation [FSW87, HV83, Jon84, LSBS81, New80, OS89a]. Implications [BCV81, Con88b, Fin85, LOBS86, PT89b, SS81b, TW88]. Importance [GKP83, Laf81, EB82a, EB82b, EB82c, JN88, SS86c, SS86d, See81]. important [SR81]. Improved [BL81a, Ho81, LO81, MS89, QX89, SRHK83, TS80, GM86a, GM81b, KK86, Kau82a, OY83]. Improvement [OK80, Mar88c]. impulse [Lar83]. impure [CB83]. Impurities [EGP86, CADL86, ESCC85, Faz87, MF87, SF89, ST86b]. impurity [Ada80b, Ada86, Fer88, PC83b, ST80a, ST80b, See84]. impurity-doped [ST80a, ST80b]. Imre [Ågr87c]. in-plane [LOGC86]. inactivation [SG81b]. inadequacies [dB84]. Incidence [CPGP83]. included [CC84b]. including [Cla83a, CG80, GjC86, GCP86, HF86, HKKM83, JŻ85, Mey88, MFM89, Nov83a, SSB83, SSF^{+}84]. Inclusion [Boč87, Boč88a, Boč88b, Boč89, FC80a, FC81a, Toy88, Yar82]. incoherent [KB86]. incommensurate [MB89]. Incomplete [McD80]. incorporation [Bec81, KBR83]. increasing [FGS85]. Independent [Živ87b, Ess86b, Lim86, Mey88, Mor88]. index [Kin89, MB80a]. indexing [DRS88]. indicator [Dea89]. indicators [Low82d]. indices [Bon83, Ler85, LSH88, Low82e, RSNT88, RR87a, RR88, SVS81]. indirect [LO81]. individual [EB82a, EB82b, EB82c]. INDO [RB88a, BB85, BH82, Boč88b, Böh82, CV82a, DS81b, DBS82, DBES80, FC80a, FC81a, FF80, LmJQHXZ83, Lip88, SIJ85, SSUF80, SS86b, SL80, TWW81]. INDO/1 [Boč88b]. INDO/2 [BB85]. INDO/2-AHP [BB85]. INDO/2-HP [BB85]. INDO/spd [Lip88]. indoles [Cal84, Yam81]. induced [Ågr83, Bal81b, HS84b, II82, Kat80, KPTG88, LRZ84a, LRZ84b, MDO^{+}82, MMC83, OPSO86, OPSO87, PPP87, Wal88]. Induction [Cla83a, Che82]. inelastic [All87, GLB81, SST83, ZJŻ88]. Inequalities [Mez83c, Löw88, Sil80]. infinite [BNK86, BSL81, CK82, DL82, KC82a, Suh83]. inflammatory [KKSM87]. Influence [KMRK86, WD81, DK87, GC84f, IO89, KAH80, LCP81, Nes85, PP86, SB88, FC80a]. influences [Sey83b, Sey83c]. influencing [SRH86]. Information [Bon81, Bon83, Bha82b, BT82, Dos84, Fra84, FEKC84, GB85a, Kin89, Koh86, Mar88c, RFST85, Ste84, Vas82, STS83, Ler85]. information-entropy [GB85a]. Infrared [Hir80c, KPSS87, SSR^{+}87, SDS83, SSN^{+}84, Wój86a, Wój86b]. Inhibition [BLT^{+}84, BLY^{+}86, CLA83b, KBSG81, LHKK82]. inhibitor [LSS83]. inhibitors [BBAT82, GC85a, GM86c, MSC^{+}83a, MSC^{+}83b, TB83]. inhibitory [AABW80, BR81b, Lew88]. inhomogeneous [Bec83, RJ87]. initial [BHS85, IDK87, Lar83]. initiation [Lad86, THKM81]. Initio [Ågr87a, Cal89a, Fro81, RKO^{+}89, Ågr85, Ågr87c, Abd80a, AEM86, AN86, AM82, AMM83, BSHK86, Bar85, BF83a, Blo89, BK84b, BBL81, BYBZB83, Bou84, BFOP83, BSCR80, BMB87, ČU80, CHKB83, CC83, CHB88, CF81b, CH82, Del88, DF89, DBD^{+}85, Dov84, Dov86, Duk82, DO84, FDAC89, GG81, GG82b, GG82a, GG83, GAL86, GLLOB86b, GLLOB88, GM81c, GP86b, GPC85, HKP81, HPK81, HLKK82, HKKM83, HKLM85, HCCB86, HM85a, HHL80, HKV85, HAB83, H^{+}90, INFU84, INO88, JN88, JŻ85, KSO84, KVR85, KGEP89, KG89, Kar81, KO81, KHNT86, KHP81, Kau83, KHRK87, KHK87, Kau87, KS87b, LS80, LdMH80, LHC80, LJK86, LT81, Law87, LMG85, LY83, LBV86, Loh87, LHK81, LP80, LN89b, MM87a, MS88, MMS^{+}86, May86b, May86a, MSI80, ME81, Nag87]. initio [NL89, New80, NSL88, OF80, OM82a, OM82b, OM83b, OM83a, OM84, OM85, OM86, OMH89, OLB^{+}82, OL82, PNGL83, PBM85, Pic83b, PDMC82, PDKC84, PKC85, RB81, RT84, RT85, RT87, RR87b, RW80b, RW81, RBR80, RJK83, RKHK88, SK87, SP86a, SSS81, SS82, SR84b, SOS84a, SOS84b, SJ85, SJ86, SFC86, SJ88, SS84a, SY82, SSB83, Sch82, SC86b, SNNG86, SG89, Sie83a, SRT81, SY86, ST82, SHP^{+}80, SRL^{+}83, SHK87a, ŠHH83b, SS86c, SS86d, SS87, SC80b, ST88, SSN^{+}84, TF83b, TS81, Tho80, THKM81, TLC87, TL80, TL82, VŻH82, WK83, WKJ88, WHT86, YHG89, Yam80a, vDTBN80, De 83, DF86, HKK^{+}82]. injection [KPT88]. Inner [CV85, BN80, ČV86, ČVV87, EC83, FNC86, Mül82, VCV87]. inner-electron [BN80]. insertion [JM83, PSBB83]. Instabilities [CK82, KC82a, Car82, Fin87, Kve82, PV89]. instability [oLCC81, OBMU82, SHH^{+}83a]. Institute [ADL78, Cal82a, Cal86b, Dun80, Kel78, NB79]. Institutes [Cal81c]. instructed [Dos84]. insulators [Kun81, LVL^{+}85]. insulin [MG80, Moo81, RB88d]. intact [Gup80]. integrability [KT88, Tab82]. Integral [HV83, Tay87b, Cof83, HS89, IRA87, JL85, Mar88a, OY83, OS89b, Ori84b, OK80, PS88, PS80b, YS80a]. Integrals [BAG83a, WJ82, AØ80a, AØ80b, AØM81, AØC82, FP84, GHGP89, GKMV80, Gus85, Gus87, Har83b, Ish83, Ish84a, Ish85, Jon80, Jon81b, Jon81a, Jon81c, Jon82, Jon83, Jon84, Jon85, Jon86, LCULC87, LL82b, MTLL82, Mic81b, MPB84, MPS86, MG85, Nov83b, OS89b, Ori82b, Ori84a, Ori85a, Özk84, PSM87, Pan83, PŻ85, SF80b, TS80, Wea88, CFS87, Sme84]. integrated [FCPL84]. integration [BD88, BVB88, Lim80, MS84, Ori80, RB88c]. integration-free [BD88]. integrations [AW82, Dem86, Ori85b, RB86, RB88c]. intense [SH89]. Intensities [Min80, Blo89, CPMP86, LdMH80, MPP83, PGG80, Sad89, SSR^{+}87]. intensity [MB80a]. intensity-dependent [MB80a]. inter [BGT84, Gre87]. inter- [BGT84]. inter-ring [Gre87]. Interacting [GF88, SRC83, GKP83, HL81, MM80, PD83, Var87, YS80a]. Interaction [BSL81, BCV80, CS83b, ENA89, GDM^{+}81, HDO88, KRCN80, MG89, MDT85, PKP^{+}81, SKR81, ACP84, AN86, AGCT87a, AL80, Bal81b, BPCT83, BC80, BVDA88, BCC82, CMB84, CKLR80, CP83b, CP85b, DH84, DBES80, DK82b, EY82, FW80a, Fis84, FSS^{+}89, Fuk83, GM80, Gas81, GSMSG82, GBZ89, Gre82, GP82, GP83b, GRBP87, GKAP81, HP81, HKG81, Ish85, Ish89, JJČ87, KG86a, KG86b, KC82c, KC82d, KPM83, KM89a, KAM85, LCC84, LSBS81, Mül82, NL89, NPL89b, NSL88, OL82, PB82, Pau81, Pay82, PPMG89, PH87, RP81a, RBSR84, RSBR85, RS88, ST89, SR84b, SS84a, SS85, Sav88, SSR^{+}87, SBB87, Sie80, SRHK83, SR87, SHK87a, ŠKH81, SPSG85, SM87a, SR80, TIYH86, TB81, Was89, WBK86, WP80, YA87, Yam80b, Wal88]. Interactions [LCCB81, MLA80, SS84b, SGP^{+}80, ACGT89, ALH86, BR85, Bar81, BMRJ86, BR81b, BL81c, BPCT83, BGT84, BCT86, Bot81, Bre80, CC84b, CT84, CT98, Cze87, DM85, DAPS81, GR89, GCP82, GPC85, GCP86, GLP83, GK80c, HH89, HG86, HAB83, HKG81, KPSK80, LGPP^{+}80, LT81, MK85, MSV89, MWO^{+}84, Mic83b, Mic85, Mic86, MG85, MR85, NTM80, NW82, Ols83, OL82, PLL86, PV89, PP86, Pie84, Pie85, PABD80, RSP^{+}80, RA81, Rec80, RM87, SBNK89, SJ86, SS85, See81, SKER83, Shi84, SF81b, TF80, TLT^{+}84, VCG86, VWCC82, WWD88, Yar82, ZH86, ZA89, Tap86]. interactive [RP89]. Interatomic [Tha83, VCG86, SBNK89]. intercalated [SMSA80]. intercalation [TR82, MLA80]. Intercalative [CGP87, PZSL83]. interconstituent [SKER83]. interdisciplinary [ČV82b]. interest [DA80, Kib83, Ran80b]. interface [Con87, LVL^{+}85, MV87, Sti82, Tru83b]. interfaces [AHB^{+}80, BQS87, Her85]. Interfering [Sla83b]. interhalogens [DS82]. Intermediate [DD87b, YJ85, Fut87, LC83a, SD82, ST82, Sla83b, Var81b, Nat88]. Intermediates [Sie81, CLL86, Set79]. Intermolecular [Cze87, GCP86, HFSB86, OL82, SF81b, Ada82, Bar81, GLLOB86b, GCP82, HAB83, JM83, KPH88, Kau83, MGN82, MSV89, MS89, OMCD81, Pee83, RA81, SR81, SFC86, WHT86, YA87, ZH86]. Internal [BD81, SB81, BTC^{+}83, FGS85, GBW88, KAH80, LS80, MMB84, SHL82, TWW81, WW82]. International [WJ82, Jor89, Shu84, Vas82, Ano84a]. interplay [Ska84]. interpolation [KPH87]. Interpretation [AAW82a, HCCB86, ACGT89, Ban82b, BN80, Bel80a, BGT84, Bon81, Dun87, FEKC84, Jas80, JJ83, Kog84, KBR83, PFC88, Pra83, RNW86, Wój86a, Wój86b, KS81]. interpreter [CP86b]. interpreting [FMF80]. Interscience [Cal89a, Wol81]. Intershell [JMS88]. interstitial [BANF86, SF89]. Intersubband [GOLC87]. Intertwining [Mel87]. interval [LBP84]. intervalence [CL80]. intervention [BLR83, LBV86]. Interventions [Jas80]. intra [OMCD81, WHT86]. intra- [OMCD81]. intracellular [MG80]. intracule [TTS84]. Intramolecular [COS82, HR84, Bar81, BGT84, OL82, SRS81, TLC87]. intrinsic [ACCC80, BWEW89, YY83, YGK^{+}89]. Introduction [Ano81e, Ano84b, Ano84c, Ano87b, Ano87c, Ban82a, Brä87, Csa81, Del81b, Del84a, FH83a, Fre80, Gos87b, LÖ82, LZ83, LÖSZ85a, LÖSZ85b, LÖSZ86a, LÖSZ86b, LÖSZ87, LÖSZ88a, LÖSZ88b, ML83, Mie89, ÖL83, Öhr87, ÖSZ89a, ÖSZ89b, Pyy84, Sta71, Sta87, Tap86, Cox87, CT84, CT98, Ela87, Löw84, AM87a, CA85, Mou89]. Introductory [And85, Fax84, DBL84, Löw89, Buc89]. Invalidity [FS86a]. Invariance [LWÖ81, Pie82, CH86, DS87, SIJ85]. invariant [ACH83, BYBZB83, CA80, DQT80, DQ80, DQ81, Mez83d, SM85, SLT84a, SLT84b, Wil80b, ZS86]. invariants [Bha82a, Cas88, Mic81b]. inverse [BERE85, DMB^{+}84, HR83, Mel87, TV89, Urr88]. inversion [BB82a, CPS89, DA88, PPP87, SSR^{+}87]. Invertibility [WS81]. Investigation [Csa86a, DF86, POT80, RA81, BB82a, Boe88, Böh83, Böh85, Bri86, CPZS86, CD89b, Fle81, Gas81, HP81, JŻ85, Kar80, KS87b, LOF83, LCA89, LGL^{+}84, LL85, MLC^{+}81, PAFW84, SAM82, SBH^{+}85, SBH^{+}86, ŠHH83b, Suh83, TLDK89, VMMV80, VŻH82, WS80b, ZJŻ88, OL80]. investigations [Bre80, HOT87a, Nec83, RDG84a, RGDF85, TB83, Wol88, HKG81]. involve [KHRK87]. involved [Fli84]. involving [BP84, BBPS83, FC81a, GTB80, Hic83, LT81, LE80, LW84b, MR81, MR80b, PP89, SS84b, Ten80a]. invstigations [RDG84b]. iodide [DM85]. iodine [JL84]. Ion [Fla81, BNK86, BLG85, BDK86, Car82, CS80b, CAML87, CM81, CE86, CE87, DBCJ86, Del87, Del88, Fla80, GjC86, Gup80, HRL88, Hur82, Köh89, KD80b, MM87a, MK85, MDO^{+}82, Pac82, PPP87, RT85, RI88, RZWL80, SSSH82, SOS84a, SOS84b, SDS83, SSYR87, UTP82, UTVP85, vDTBN80]. ion-induced [PPP87]. Ion-ion [Fla81, Fla80]. ion-molecule [BNK86, Del88, SOS84a, SOS84b]. ion-selective [UTP82]. ion-surface [BLG85]. ion-water-water [MK85]. Ionic [Ols83, Bec81, BCD^{+}80, Bou84, BCD89, CC82b, LWZ88, Lim82, LSR^{+}84, MB86, ST80a, ST80b, UTP82]. Ionization [GSMRM85, HMF82, KM85, BBH87, Bak85, BN84, BSL^{+}80, GP80, GBO86b, HK86, HDFL84, LJW87, LJC83, Lar83, MM87a, MB84, PZ84, Pee83, SAM82, SGP^{+}80, TL83, UI80, WSD80, YUKK84, Nak83]. ionized [NH81, PS80a, TS81]. ionizing [IDK87]. ionophore [GP83a]. Ionophores [Ovc81]. ions [BZA84, Bec82, Bha82c, CD89a, DCF82b, DS81c, ENA89, Gan81b, GM86b, Gre82, GKAP81, GKP83, KG85, KG86c, KG86d, KO80, Kun86, KS86c, LK81, LC82a, LOR84, LWZ88, Mar85b, MMM81, Mof81, MPPW82, Nov83a, Pac83, RN86, RT87, Rud84, SD88, SF81a, Sin89, The84, TL80, TL82, Yan85]. IP [RSMM86]. IPPP [SC86b]. IR [Bri87, Dat82]. iriallene [DBD^{+}85]. iridium [EIKG^{+}88]. iron [PD84b, SS80f, FHHHE85]. irreducible [GC87, ZS86]. irregular [SRC83]. Irreversibility [CD83, Emc80, SC86a, SC87]. irreversible [MD82, Sch89]. ISBN [Fro81, Fro83a, LM81, Mar81a, Sme84]. ISO [Ho81]. isobaric [RJK83]. isocyanide [TF83b]. isocyanides [Mof81]. isocyano [LH85, Mof82]. isocyanocarbonium [Mof81]. isocytosine [JKL85, LOB86]. isodensities [PC86]. Isoelectronic [LL82a, AJ83, Fra85, Gan81b, Lás85, LM86, MS87, Mez81a, MB84, MPPW82, NSS86, PS84, SB83a, Sen80a]. isoguanine [JKL85]. isolated [HCCB86, Lip88, Mar81b]. Isomeric [RW80a]. isomerism [PBBK89, Sla83a, Sla83b, Sla84]. isomerization [Bal82a, GZWYCHC88, Mof81, Mof82, MP87, TMON80, TOK80]. Isomerizations [RD84]. isomers [Dun88, KSO84, LH85, LHK81, SWM82]. isomorphism/order [FF82]. isopentenyl [Tew88a]. isopotential [MT87]. isoprotonic [Mez81a]. Isoscalar [KAA85]. Isospin [ŠKR84]. isostructural [VMMV80]. isotachysterol [BP86]. isothermal [RJK83]. isotope [BBLR84, STG86]. isotopic [GT86]. isotopomers [CHB88, NPL89a]. isoxazole [TMON80]. Isoxazolol [BVDA88]. Israel [Pul79, PTG80]. Issues [Zer87]. Iteration [ASL88]. Iterative [Gol84, BEP87a, BEP87b, FC80b, Mül81]. Itinerant [SS88, Enz80]. IV [AØ80a, CPGP83, FC82e, GC84f, HKLM85, KG86c, LSP85, LHP80, McD80, PT84, Pan84, PDC81b, RSP^{+}80, RK82, RMM83, SP81, SC84a, ST83, SG89, SV85b, Van85]. Ivar [Ano89-28, LGJLCB89, Ano89-29, Lun89b]. IX [KRST83, VŻH82].

J [Ågr87a, Cal81c, Cal81b, Fro88b, Nat88, BAG83a]. J.-M [Cal81c]. Jacek [Ågr80]. Jack [Ågr87b]. Jahn [Sin81]. Jerusalem [LM81, Mar81a, Pul79, PTG80]. Joanna [Muj86]. John [Cal86b, Cal88, Mår80, Nat88, Wil81]. Jones [Sme84]. Jørgensen [Ågr87b]. Joseph [Nat88, Rav82, Yan82]. Josiane [Fro83a]. Julg [Cal81a]. junction [GUSS89]. June [NB79, LGJLCB89, Lun89b]. JWKB [Sil86].

K-region [Low82d, Low82e]. Karadakov [May86c]. Kari [Ågr87c]. Karmowski [Ågr80]. Kato [Tat81b]. Kekulé [EB82a, EB82b, EB82c, EBO83, Kia80]. Kelm [Dun80]. kernel [Suc81]. ket [Rér89]. Keto [PSS86]. Keto-enol [PSS86]. ketonization [LBV86]. keV [CPS81]. keV/amu [CPS81]. Kikuchi [RM82]. kinase [GB81, MSM81]. kind [PH87]. Kinetic [CGK82, Kre81, LÖ85a, CEB86, Csa86a, Csa89, GZWYCHC88, Kat80, Kha85, NP81, OS89a, ST86a, STG86, Ulm83, ZAB85]. kinetic-energy [Csa89, ST86a]. Kinetics [BQS87, GF83, New80, Ulm85]. Kirtman [SH87b]. Knight [KB80]. known [Mor89, TN89]. Kobeissi [Tel88]. Kohn [Kel86, HF86, Kry80, LP82, LM82b, Wea82]. Koopman [Coo87]. Koopmans [AEM86, LJC83, MD87]. Korb [Fro88b]. Kos [NB79]. Kotani [Ohn84]. Kr [ESCC85]. Kramers [MM80]. KVV [SD81].

L [Ågr82, Kum87, Muj86, Rös87, Cal84, DIV86, LN89b, NL89, Sim81]. L-serine [LN89b, NL89]. laboratory [Bri86]. lac [KRMK86]. lac-repressor [KRMK86]. laci [FPDA89]. Lactam [DM83a, SHH^{+}83a]. lactate [KK82b, KK83]. lactic [PS81]. ladder [BPLB87, Hub85, MPS86, PM83]. Ladik [Cal81c]. Lagrange [NDBE82]. LAMMA [SL81]. Lanczos [HL83]. Landé [LBP84]. Langevin [Ade87]. lanthanoid [LmJQHXZ83]. Laplace [DFP80]. Large [BHVR83a, BHVR83b, ČV82b, Sim82a, Wu82, Ano82s, AH86, BEO80, BEP87a, BEP87b, Ben82, BAN88b, Cow82, Gás86b, GM86b, HH89, HAmGS86, KHP81, Kau87, KD86, Mar86a, MC88, MD82, McW84, Mez84c, MS80b, NS83, OMCD81, OY83, Ott86, Ran80a, RSG^{+}87, Rec80, SHK87a, SC80b, Tho83]. large-molecule [Gás86b, McW84]. Large-Order [Wu82]. Large-scale [BHVR83a, BHVR83b, BEO80]. Laser [Hir86, Kat80, DR84, KKCM85, KKC^{+}86, KKCM87, RDH84, SL81, MK80]. Laser-excited [MK80]. Laser-induced [Kat80]. Lattice [ST80b, Ban82b, BS84, DFP80, Fin89, Mod81, SP82b]. lattices [RM82, RB86]. Lawley [Cal89a]. Layered [Cal81b, WS79]. LCAO [DFH84, BB82b, Bha81b, BH82, BK84b, CD87b, Del84b, DC86, Mal82, MSI80, PNA86, PNA88, RZWL80]. LCAO-MO [Mal82]. LCAO-X [DC86]. LCAS [RC80b]. LCGTO [SAS86, AS86, JR85, RS86, RSGK88]. lcgto-lsd [AS86, RS86]. Lead [MB89]. Least [Tho85, GM81b, RSST86]. Least-squares [Tho85, GM81b, RSST86]. Lecture [Ågr88, Fro81, Fro83a]. lectures [GC84a]. left [JY81]. left- [JY81]. Legends [ZA89]. Lemaistre [Fro88b]. Length [Has85, PC83a, PCG83, PT84, PP83, SHH^{+}83a, SI89b, TP84, TP85]. lengths [Ver80]. lens [Var81a]. Letter [BL89a, Swe83]. Leu [VARPU84]. leucine [ST89]. leucomycin [JG81]. Level [Mez82, BL81a, BP86, De 83, HO85a, LCLC86, LCLC88, Mez83d, PABD80, SC80b]. levels [FSK79, FSS^{+}89, FF89, HZBR89, KG85, KG86c, KG86d, Mül82, NPL89a, PČP82, RPS80, RZP84, RZWL80, ST84, Woo82, ZHB^{+}87, Ågr80]. Levente [Bra86b]. Levitan [HR83]. Lewis [CGT86b, ESS86a]. Li [PFLB83, AM83a, BBH87, CEB86, CPS81, DF89, EY82, ESKG83, GG82a, NS82b, SD81, TH80, Yar87]. LiBeH [BR80]. LiBH [BR80]. Library [SHK87b, SMHK89]. librational [UV83]. librations [VU85]. LiCH [BR80]. Lie [DYD89, SHLL83, SLT84a, ACCCBF80, Pan85, SLT81, TSL80, TV83a, TV83b, TV84, TV89, Van85, VT85, Van86, Wei88]. Lieb [Csa86a]. Life [Wal84, Fox84, Löw84, You84]. lifetimes [Löw85a]. lifting [LL86]. ligand [ACCCBF80, BL81c, DG86, GPC85, NSN89, SK87, SHLL83, VL89]. ligand-field [ACCCBF80, DG86]. liganded [LCCW87]. ligands [KTGR89, YGK^{+}89]. light [CPMP86, Kar89, KFN80, KPTG88, Pan81, SM87a]. light-induced [KPTG88]. LiH [SR80, AM83a, AEM86, CEB86, FD89a, GG82a, GPLP85, KC82d, ZW83]. like [ALH86, ECM89, Koh86, Nov83b, OS89b, MB83]. likely [LP84]. limit [BPT88, CB88, GM86b, Hal80, Ori82b, Tho83]. Limited [Muj86, AGCT87b, Eri89, RP81a]. Limiting [BB81a, Sla85]. Linderberg [RSSS81]. line [PS84]. linéaires [Van86]. Linear [BH89b, KT88, ABD^{+}80b, ABTP83, BPT83, BR81a, BSL81, BEP87b, Big86, Dal80, Dem86, Dia86, Dov86, DO84, FC82a, FDA83, FM86, GTB80, Has85, Has86, Ish84a, Kar84b, Kir89, Kub82, OJY83, Ori85b, RSMM86, SB84, TN89, TP89, Van86, Wei82, DSB84]. linear-combination-of-atomic-orbitals [Dov86]. linearly [FC87, Lim86, Mey88, Mor88]. lines [LRT88]. link [BPT88]. linkage [AEHM85, BVGV89, HM84]. linked [GP80, Muk86, YL83, Yad83]. Liouvillians [Löw82a]. lipid [HJ80, MH80a, SAB^{+}80c]. liposomes [DS81a]. liquid [BMRJ86, CC83, CC84b, Hal83, MV87]. liquids [Gol88, LCD82, Gee89]. List [Ano85c, Ano86e, Ano86f, Ano87d, Ano88i, Ano88j, CS83c, Ano82s]. literature [H^{+}90, OM82a, OM82b, OM83b, OM83a, OM84, OM85, OM86, OMH89, OYH^{+}80, Pic83b]. lithiation [JI88]. lithium [CEB86, Gan80, GjC86, Ho81, MMS^{+}86, MPPW82, OL80, SDBDSMG86]. lithium-chloride [GjC86]. lithium-fluoride [GjC86]. lithiumlike [LC82a]. Littleton [SC89]. lived [Gas86a]. living [MB83, Mis84, Poh83, Poh84, SG80b]. LMO [HAmGS86]. lobe [Ten80b]. Local [Bec85, BV80, Dea89, Ell84, MS80a, SHRH82, ZAB85, BPT88, Bec81, Cas85, CA85, CP83a, FGM87, GA82, JMD88, KAD81, KD86, KMRK86, KPS87a, KPS87b, KPS88, KEB86, LCLC86, LCLC88, MC80, MKW85, Min87, PSK86, PW86, See81, Sha85, Tho86a, Yar82, Gre81b, KWM86, Pan82b, TP84]. local-density [JMD88, MKW85, Min87, KWM86]. local-energy [Tho86a]. local-scaling [KPS87a, KPS87b, KPS88, PSK86]. local-space [KD86]. Localization [Pip89, BBH87, Cal80, CS83a, LBC^{+}87, LNN84, RJ88, Sie87, TNY83]. Localized [CZ82b, LPB82, LS82, PMT84, AL80, Csa83, Cul88, Cul89, DBES80, GK80b, Gos85, KCK83, KHK87, KSGB^{+}83, LL84, MBSC81, Nes85, OL82, Pip85, PSIT87, Hil81c]. localized-site [KSGB^{+}83]. localizing [Koe86]. location [CBV85]. Locations [Dia86]. London [Cal82a, Cal82b, Cal86d, Tap86]. lone [PS81]. Long [EY82, LBKV88, PABD80, RM87, DAPS81, Gas86a, Kog84, Pie84, Pie85, YA87]. long-lived [Gas86a]. Long-Range [EY82, PABD80, RM87, DAPS81, Kog84, Pie84, Pie85]. longest [GC84f]. longest-wavelength [GC84f]. longitudinal [DBD^{+}85]. Longuet [BR82]. Longuet-Higgins [BR82]. look [Bha81a, MK81]. Looking [Bri86]. loss [LVL^{+}85]. low [All87, DMB89, DS88, EV87, Fla80, Fut87, GD82a, HF86, Has82c, Mar86b, Nom80, SO85, SNNO86, SS81b, ZV86, RGDF85]. low-dimensional [Mar86b]. low-energy [HF86]. low-lying [DMB89, Has82c, Nom80, SNNO86, RGDF85]. low-temperature [SS81b]. Löwdin [AEG80, BM86, ČVV87, GC84a, Jon84, JBW87, Pan82a, Pan82b, Pan83, Pan84, WS81, WS82, HL83]. Lower [ČV86, CV83b, BH86, Csa86a, Dia86, LO81, TOK80, Gol80]. lowering [MB80b]. lowest [BN84, BHM87, KG86a, KG86b, WS80b]. lsd [AS86, RS86, RWOT81, SAS86]. lucanthone [SMSA80]. luminescence [MK80]. luminescent [PC83b]. LUMO [EIMS88]. Lyapunov [KB87b]. lying [DMB89, Has82c, Nom80, RGDF85, SNNO86]. lysozyme [BBMP82, GDM^{+}81, MP84a, PP86].

M [Ågr80, Cal81c, Cal81b, Fro88a, Mår82]. Maclaurin [TP89]. macrocyclic [RHST88]. macromolecular [Kar80, PP80]. macromolecules [LP81a, Luo87, PFC88]. macroscopic [SC86a, SC87]. made [NTM80]. Magnesium [BL81c, Gup80, MBR80]. Magnetic [KSS81, LRZ84a, And87, BAG83a, BYBZB83, BYBB84, BH89b, DS80c, EHK83, FC85, FNC86, Gla85, GLP83, Hir86, IS83, JL87, KL80, LRZ84b, LM89, LOGC86, MSGPP80, NH84, PAFW84, SCH84, SPSG85, TK86]. magnetic-field [LRZ84b]. magnetism [DK87, Enz80, FS86c, SR84a, SS88, Cal88]. Magnus [KKTE80]. main [KST84, Loh84, Yar82]. main-group [Loh84]. Maksić [Zer87]. Malcom [Nat86]. Malinowski [Wol81]. man [NTM80]. man-made [NTM80]. manganian [HHL80]. Manifold [Mez81b, Bak85, Mez83b]. manifolds [Bor82, Cho80]. Mann [DS88]. mannose [BR81b]. mannose-containing [BR81b]. Many [Ave86, CD87a, IS80, NMK84, PŻ85, Ade87, AN85, Ano81s, AW83, AZY84, AL86, ALH86, ASL88, BK89, DG83, DRS83, Ess84, FH81, FH83b, FM86, GR89, GD82b, Gla85, GC84b, GC84c, GC84d, GC84e, GC85b, GBK86, HB86, Hub80, IS85, IPV81, Kal85, KCK83, KPM83, KM85, Kha85, KPS87a, KB88, LJW87, LCULC87, LB83, LZ89, LC88a, LFM89b, Mat87c, Mat87d, McD81b, MJ89, Mic83a, Muk86, MP84b, Muk84, MPD87, PM84b, Sad83, SF81a, Sie80, Sto82b, Suc84, SH87b, Tay83, Tay84, Tay86a, Tay86b, UI80, Var87, WJP83, WJP85, Win85, Sto82a]. many-atom [Mic83a]. Many-body [CD87a, IS80, Ade87, Ano81s, AZY84, Ave86, ALH86, BK89, FM86, HB86, Hub80, IS85, Kal85, KCK83, Kha85, KB88, LB83, LC88a, Mat87c, Mat87d, McD81b, MJ89, Muk86, MP84b, Muk84, MPD87, PM84b, SF81a, UI80, Var87, WJP83, WJP85, Win85]. many-center [LCULC87]. Many-dimensional [Ave86, ALH86]. Many-Electron [PŻ85, NMK84, AN85, DG83, DRS83, Ess84, FH81, FH83b, GR89, GD82b, Gla85, GC84b, GC84c, GC84d, GC84e, GC85b, GBK86, KPS87a, LJW87, LZ89, Sad83, Sto82b, Suc84, Tay83, Tay84, Tay86a, Tay86b, Sto82a]. many-particle [AW83, AL86, ASL88, KPM83, LFM89b]. MAO [MSC^{+}83a]. map [PC86]. Mapping [KAD81, MT87, KPH88]. maps [Cas85, FZB86, GM86c, HKP81, HLKK82, HKLM85, LHK81, RMM83, SHKP80, SHP^{+}80, TH88, THKM81]. March [Fax87, Han86]. Marcus [Bre80]. marginal [HT82]. Marineland [Fax87, Han86]. Maruani [Fro88b]. Masao [Ohn84]. mass [FWS^{+}85, FS86a, HYS89, KJMS85, PBFL80, SH89, SL81]. mass-velocity [FWS^{+}85, FS86a]. Masthead [Ano80f, Ano80g, Ano80h, Ano80i, Ano80j, Ano80k, Ano80l, Ano80m, Ano80n, Ano80o, Ano80p, Ano80q, Ano80r, Ano80t, Ano80s, Ano81f, Ano81g, Ano81h, Ano81i, Ano81j, Ano81k, Ano81l, Ano81m, Ano81n, Ano81o, Ano81p, Ano81q, Ano81r, Ano82e, Ano82f, Ano82g, Ano82h, Ano82i, Ano82j, Ano82k, Ano82l, Ano82m, Ano82n, Ano82o, Ano82p, Ano82q, Ano82r, Ano83c, Ano83d, Ano83e, Ano83f, Ano83g, Ano83h, Ano83i, Ano83j, Ano83k, Ano83l, Ano83m, Ano83n, Ano83o, Ano83p, Ano84d, Ano84e, Ano84f, Ano84g, Ano84h, Ano84i, Ano84j, Ano84k, Ano84l, Ano84m, Ano84n, Ano84o, Ano84p, Ano84q, Ano85d, Ano85e, Ano85f, Ano85g, Ano85h, Ano85i, Ano85j, Ano85k, Ano85l, Ano85m, Ano85n, Ano85o, Ano85p, Ano85q, Ano86g, Ano86h, Ano86i, Ano86j, Ano86k, Ano86l, Ano86m, Ano86n]. Masthead [Ano86o, Ano86p, Ano86q, Ano86r, Ano86s, Ano86t, Ano87e, Ano87f, Ano87g, Ano87h, Ano87i, Ano87j, Ano87k, Ano87l, Ano87m, Ano87n, Ano87o, Ano87p, Ano87q, Ano87r, Ano88x, Ano88k, Ano88l, Ano88m, Ano88n, Ano88o, Ano88p, Ano88q, Ano88r, Ano88s, Ano88t, Ano88u, Ano88v, Ano88w, Ano89n, Ano89o, Ano89p, Ano89q, Ano89r, Ano89s, Ano89t, Ano89u, Ano89v, Ano89w, Ano89x, Ano89y, Ano89z, Ano89-27]. MATCH [KKH^{+}83]. matching [SB83a]. Materials [Cal81b, BMMS81, Cal82c, Coh86b, Fau85, Faz87, Fri87, GMM81, Hay82, KH80, Koe86, KWBC80, McM86, MST^{+}80, Pet79, Zun85, Mår80]. Mathematical [GC82b, Tri86, Ban82b, Sim82a]. Mathematics [Brä86a]. matrices [BEP87a, BEP87b, CH86, Coh86a, Del84b, Duc85, DK81a, DK81b, Gos82b, KMW80, Küm83, Lar83, Löw86d, LW81, Mal86, Ret86, SS81b]. matricielle [SV82, SV85b]. Matrix [AW83, BPLB87, BFJ81, Kry82c, MZ85, Mar81b, Pay82, Wor85, Ave84c, Bel80a, Bel80b, Cas88, Cha82, Cha88, CMB84, CFG^{+}83, CSF80, DH84, Don81, EKBE86, FCB80, FM83, FSW87, GB85b, GP86a, Gou86, Gre80, Har83a, HCCB86, JD85, JBW87, Kau82a, KS82, Kia80, Kim81, KEB87, KPH87, LP85, Lim89, Lin81b, Lin89, Luk82, MC80, Mes80, Mes85, Mor88, Mor89, NPL89b, Niu80, OS89b, Pan85, QX89, Roo80, Sch80b, SY86, SV82, SSN^{+}84, Tay85, Tho86a, WP80, YS80b, YS80c, GC84c, GC84e, Pan84, SV85b, TNB86]. matrix-isolated [HCCB86]. matrix-oriented [JD85]. Matsen [Wor81]. matter [BB81b, ENA89, Fri87, Lar86b, Mat87a, Mis84]. Matthias [Ano81d]. Maurice [Gee89]. Maximization [GB85a]. Maximum [LZ87, FSP87, Pip89]. May [PTG80]. Mayer [Rav82, Yan82]. MB [KU80]. MBPT [AB87, RSMM86, SHK87a]. MC [CL83b, Iga85]. MC-SCF [Iga85]. MCD [KM80, Mar89]. MCF [Ott86]. MCRPA [YOJ81]. MCSCF [ÅCFRJ85, CHKB83, DG85, DMB86, GYJ83, GSO^{+}87, HC81, MR81, MR80b, OY83, PKG86]. MCSCF/CI [CHKB83]. MCTDHF [YOJ81, GY87]. McWeeny [BAG83a]. Me [WBK86]. mean [CEB86, Pee89]. Meaning [Gra82]. means [BAG83a, BBO83, FSW87, HČ83, Mak80, PZ84, Toy80]. measure [CLA80a, Fra88, PM88, Pip89]. measured [AØM81]. measurement [Dun87, Jas80]. measurements [Bro80, CP80b]. Measures [BHS85, BJR86, CD87b, Gre81b]. Mechanical [BCV80, AZY84, Ave86, ALH86, CLL86, DAPS81, Dew88, FC81b, FC81c, FC81d, FC82e, GSS85a, HCCB86, HM85a, JRK84, Kim81, LOF83, LOBS86, LM89, LFL^{+}81, Mar89, PS80b, PBM85, Pur88, SBH^{+}85, SBH^{+}86, SHL86, SF81b, Suh84a, SSN^{+}84, Tac83, VCB83]. mechanically [GS89, Ran80d]. mechanics [AA89, CSK87, ČV88, Con88b, DC82a, FPDA89, Gre80, Hal80, Jas80, JJ83, MSWC88, SIN80a, Sil84, Tho83, VARPU84, Yan82]. Mechanism [GP80, SS80f, TT81, Abd80a, BBLR84, CSB86a, CL83a, HZ84, Kot81, KCP80, KPTG88, KPT88, LWZ88, MTWO83, OPSO86, OPSO87, PMOW89, RT84, RT85, Shi85, SMJ83, ST88, UV83, ZV85, ZV86, THKM81]. Mechanisms [Mar81a, DBL84, Dew88, ES88, EC83, FGC84, HKFW87, LČ84, Mez84a, Mez85a, Mez85c, PTG80, Tho86b, UTP82]. Mechanistic [LPCF80, LPSK83, PL83, PL84b, MFM89]. media [PCA89, RN84]. mediated [Gas85]. medium [ACGT89, FGC84, Kun83, KS87b, MR87, RJ87, ZV85]. medium-scale [Kun83]. mellitus [RB88d]. Melting [Nag82b]. membrane [Con87, DBCJ86, DK82a, Hos80, KKD80, Rab84, RSP^{+}80, YVK^{+}80]. membrane-active [KKD80]. membranes [HJ80, IMTG80, Lal81, LN83, NTM80, RKV^{+}80, Tat81a]. memorization [Koh86]. memory [Fox87, Ská83]. mercaptan [SG81b]. MERGE [HKP81, HPK81, HKK^{+}82, HKKM83, HKLM85, LHK81, THKM81]. Mermin [Tac89]. MESFET [BW85]. Meta [BK84b]. Meta- [BK84b]. metabolic [LL85]. Metabolism [SG85b, PZPL86, SG80a, LP81b]. metabolite [LP84, PZSL83, ST82]. metabolites [DD89, HKP81, HPK81, LT81, PP82, RT87, TR86, Ulm81]. Metal [Bor81, FHHE80, GM80, LFM89a, Ser81, The84, Bal88, BSHK86, Bau86, Bha82c, BQS87, BP80, Boč88a, Böh82, Böh83, BDK86, BKJ88, CKM81, Ell84, Faz87, GPN80, Hil85, JN88, KN80, Kin86, Lau88, LNN84, LHH^{+}83, MM80, MB85, MG85, Nec83, Nov83a, PD84a, PT89a, SF89, SSF^{+}84, TLDK89, TP80, WBK86, WS80b]. Metal-semiconductor [LFM89a]. metallic [DC87, HM81, Suh83]. Metals [FHHHE85, BBPS83, BANF86, ČD87c, De 80, EGP86, FK83, Hal88, Kar89, Mar88b, MM80, Mor83, MF87, Nes85, PC88, SS86a, Wed81, WS80b]. metastable [BS80a, PBA88]. methane [KRCN80, NOMK80, NCKL81, CPMP86]. methanes [CKNG80, LL85, LLM86]. methano [BBGS84]. methanol [SA86b, SS89a, TF83b]. Method [Bal82a, KPS87a, KPS87b, KPS88, MKVP80, Pan82a, Pan82b, Pan83, Pan84, PSK86, ÅFL81, ALB84, AB84, ANH81, AM87b, AØ80a, BAG83a, BS80a, BPT83, BS81a, BB82b, BB85, BSZ81, BHM87, BP80, Brä86c, BFJ81, BS80c, CLVC84, CS89, CMM83, Chu86, Cio88b, Cio89, Coo87, Cze87, DDP81a, DDP81b, DM85, DG85, DMB89, DG83, Dem83, DVV80, DF86, DK82b, Duc85, DO84, Dun87, Ess84, FVZ^{+}88, FC80b, FAC84, FD86, Fer88, FSP87, FSW87, GG85, GN87, GOS87a, HK86, HTD80, HHL80, Her84, HO85b, Hir80b, HR83, HAmGS86, IRP81, ITKD80, IPV81, IST89, JD85, JS88, Jon84, JT85, JR85, Jun80, Kag83, KN80, Kal86, KC83, Kar86, KPM83, KT80a, KFD85, Kle83, KMW80, KT84, KS81, KPH87, KM89b, LC82a, LC80, LFL^{+}81, LTLF87, Lip83]. method [Lip88, LZ87, LW84b, Löw86e, LLC83, Mal82, MG83, MPBN81, MC80, MJW88, Mes85, MSGPP80, Moh80a, Moi81, MK81, Mül81, Nag87, Nas83, Noa80a, Noa80b, OM81, OL82, Ott86, Özk84, PC83a, PCG83, PXLC86, Pau81, PP83, PKP87, PBFL80, PS80b, PABD80, PD80, Poh83, POT80, PŻ85, RNW86, RM82, RSG^{+}87, Røe81a, Roo80, RS80, RSGK88, RC80b, RLS87, SA80, Sav88, SC86b, SS80d, SF80a, Sie83b, Sim80, SP82b, Sto82a, Sto82b, SM84, SBWW80, SI89b, Sve88, TF83a, TNB86, TPN86, Tho80, Tin80, UI80, Wag83, WSD80, Wea88, Win82, Živ83, ddF86, LW82]. Methods [Cal89a, Fro81, Mat81b, Ågr85, AČP82a, Ada64, Ano81s, Ber81, BH82, BS82b, Boč87, Boč88a, Boč88b, Boč89, BCT86, ČU80, CV82a, CSK87, CS87, Cla83a, DS81b, DS89b, DBS80, DR84, Duk82, DQ81, Eng83a, Eng83b, FGG83, GKMV80, JMD88, KK85, KS86b, KT84, Lar86b, Law87, LT89, LC81, MSV89, Mel87, MS80a, Muj86, MM81, OS89b, Ori80, PG80, RSST86, RDH84, Sad85, SMD83, SIJ85, SNNG86, SS86b, Sim81, Tel88, Tho85, Wes81, dB84]. methoxy [BCD^{+}80]. methoxyl [GSG84]. methoxyl-substituted [GSG84]. methoxyuracil [Zie82]. methyl [CL83a, DM85, HL80, JSDK89, KAH80, KPSK80, ST89, TF83b, XYSJ81, XHR86]. methylacetamide [KL80]. methylamine [GSM83]. methylation [SS80e]. methylcholanthrene [HKP81, HPK81]. methylchrysene [Sey80]. methylcopper [CPN89]. methylene [AYBP83, AMM83]. methylenedioxytryptamines [RWOT81]. methylgiyoxal [GSM83]. methylglyoxal [Gas80, Gas81, KL80]. methylguanine [PSS86, PKP^{+}81]. methylpyrazines [PO85]. methylsilane [KSK86]. methylthio [Tew88a]. methyltransferase [SS80e]. metmyoglobin [AWCL89, LCCW87]. metric [Mez84d]. MeV [BH89a]. Mg [WBK86, CS83c, GP83a, Kal86]. microcluster [TKY89]. microcomputers [KST84]. microcrystals [Mar81b]. microlocal [Sjö87]. microphysical [DSG80]. microprobe [SL81]. microscope [HRL88]. Microscopic [Köh89, VVD80b, Fol81, PBGT80, SC86a, SC87]. microscopically [RJ87]. Microwave [CP80b, CP83b, VPK80, KAH80, MMI^{+}82]. milk [BR81b]. mimetics [VB85]. mind [Fox84, Löw84, Mis84, Wal84]. MINDO [DA80, MH82]. minima [GC84f, MASF88]. minimal [SS86a]. minimization [BB83, FPT81, PG80, Sel86a, Sel86b, Tho86a]. minimum [BVC80, RTM81, RTM83, RMM83, RTM86, SC80b, Tho85]. minimum-basis [BVC80]. minimum-variance [Tho85]. minor [KMRK86]. Miroslav [Fro81]. mirror [Jar89]. mispair [SRSR85]. mispairs [KBR83]. mitochondria [YKTA80]. mitochondrial [FF84]. mitogenesis [Moo81]. mitomycin [CS83b, SS82]. mixed [CFS86, CFS87, DBS82, Gos82b, KS82, Lim82, MPBN81, OKZ85]. mixed-valence [OKZ85]. mixing [ZNS80]. mixtures [CF81b, TSA86]. Mn [LLAG89, SJRP88]. MnCl [KBT80]. MNDO [DS87, FS86b, FS87, MG83, MBKH83, OG86, PL83, PL84b, SA86b]. MO [BB82b, Bha81b, MSI80, ÅFL81, BVD81, BS80b, BH82, BK84b, BPCT83, BC80, CD87b, CKM81, Che80, GAL86, GK80c, Kar84a, KG89, KHNT86, KŠHH83, KK83, LY83, Mal82, OYN80, PM88, PNA86, PNA88, RZWL80, SK87, SD80, SSSH82, ŠHH83b, TMON80, TF80, TF82, WK83, dB84, CJZ^{+}86]. MO-LCAO [PNA86]. Mobility [LC82b]. Möbius [RZ86]. Mode [Sey81, ES81, KMKC88, SZR80, SHRH82, SS89b]. Model [GBK86, KT80a, MH80a, PWL89, SBE^{+}83, Ave84a, Ave84b, BB82a, BB82b, BS80b, Bes88, BR80, Bha81b, Böh83, BST86, BP83, BVDA88, Bou84, BWEW89, Bur85, CBV85, CLS81, CR81, CFG^{+}83, CL86, CCS84, CZ82b, CLZ84, Cul88, Cul89, DP87, DT83, Del88, DFP80, DAPS81, Del84b, DFH84, DH85, DK82a, DL81, DOW87, EZ83, FY82, FL83, FNPS82, GMM81, GM83, GSRG80, GMP84b, Gre81a, GC82b, GC83, Has85, Has86, HLR80, HRL88, HL88, Hos80, HT85, Hun86, IMTG80, JP80, JSS80, KRT83, KKH^{+}83, KS89a, Köh89, KEB87, Kva83, LPB82, Lar86a, LMG85, LSR^{+}84, LHZ80, LCCW87, MR80a, MMN87, MS88, MA80a, Mcc83, MG89, MDO^{+}82, Mez83b, MC89, Moo81, Muk86, MB85, NKV^{+}89, NP80, OLB^{+}82, OMC82, Pac82, PMOW89, PMCA88, Pul83, PSBB83, Rab84]. model [RL88, RA81, RS84, RBSR84, RSBR85, RM84, Røe81b, RSGK88, SS82, SY82, See84, Shi87, SNNO86, SvG89, Sok81, ŠKH81, SRC83, SST83, Ste84, SLT81, SH87b, TH81a, TPČ83, TSL80, Tew88b, Tin80, WCLZ86, WJP83, WJP85, Yar82, ZLP81, ZP82, GG83, LW84a]. modèle [SvG89]. modeling [FHHHE85, GSS85b, LCC84, MR89, Nal80, SRSR85, SSR86, SSRR87, MBT80]. Models [KPH88, OKZ85, Zer87, BBPS83, CV83b, Dew88, Gas81, HR84, HČP88, HLPV83, HM89a, JY81, KG86d, Kun86, Löw86b, McW84, Mez85a, NW82, PB82, PČP82, PZSL83, PL83, PL84b, RGW88, Røe82, Røe87, ST89, See88, Sla85, TKY89, TS82, Var86, VV83, VWCC82]. modes [LSS83, Shi87, vZ81]. Modification [BS80b, FM80]. modifications [KPSK80]. Modified [BG88, BP80, Hir80b, Lip88, RR87a, BDK86, CFS87, CP83a, DFH84, Den83, Den85, HB82, LW84b, LWK84, Tew87, Tew88a, Tew88b]. MODPOT [HKP81, HPK81, HKK^{+}82, HLKK82, HKKM83, HKLM85, LHK81, SHP^{+}80, SRL^{+}83, SHK87a, THKM81]. MODPOT/VRDDO [HPK81, HLKK82, SRL^{+}83]. MODPOT/VRDDO/MERGE [HKP81, HKK^{+}82, THKM81]. modular [CB85]. modulation [Urr88]. modulators [VL89]. moduli [KS86b]. moieties [KVR85]. moléculaire [OF80]. Molecular [AEHH81, AEHNEA85, AM83a, AT85, AM88a, BB82a, CGG81, CPS81, CS81, CLP83, CT84, CT98, Csi82, DBC84, FPDA89, Fox84, Fox86, Fut87, GC85a, HPK81, HM85b, HR87, Ish84a, Ish85, JW86, KH86, KK82b, LCDP80, LP81a, Lew88, LZY83, LPW81, LHK81, MR89, MBR80, NÁNSP87, NKZ88, OF80, OPSO86, OPSO87, PG80, PLL86, PV89, Pee89, PSK^{+}81, RWOT81, RSD82, SC80a, SP81, SP82a, SC84a, SP86b, Shi85, SKS83, SHP^{+}80, Sto82a, Sto82b, SM83, Tap86, UTP82, VV87, VB85, VL89, WJ82, WMP^{+}83, Ågr85, Ågr87b, Ågr87c, AEHM85, Ade87, AN85, AN86, ADF^{+}88, Ano80c, Ano84a, Ano87a, AM87b, AM89, ACCC80, Bec85, BEL89, Bha82c, BK82, BPCT83, BKP83, BYBB84, BH89b, Bri84, BTC^{+}83, Bri86, BVGV89, BSCR80, BVAB81, Bue86, Cal86c]. molecular [CLA80a, CDL82, CD83, CL80, CSK87, ČV82b, CFS86, Con88b, CLP82, CT87, CJ89, CZ82b, CLZ84, CD89b, CF81b, CH82, DP87, DPC82, D^{+}85, Dav87, Dav85, Dav91, Del88, DÖ89, DF83a, DCF82a, Duk82, DF83b, DO84, DC82b, Dun87, DQT80, DQ80, DQ81, Dur86, EÖWZ84, Edw88, EBEE87, EIMS87, ES88, Ela87, FP84, Fer88, FZB86, FS86b, FS87, FHH^{+}87, Fro83a, Fro88b, FH81, FH83b, GM80, GG81, GG82b, GG82a, GG83, GDBRV87, Gos85, GKMV80, GMP84b, Gre81a, GGM81, HKP81, HLKK82, HKLM85, HTD80, Har83b, HDFL84, Hib81, Hil81c, HR83, HAmGS86, Hun86, IL83, IDK87, JD85, Jon83, JT85, JBW87, JSDK89, JS86, KBS81, KO81, KFD84, KM85, KB86, Kia81, Kon83, KTGR89, Kot84, Kot81, KCP80, Kry82a, Kry82b, KBSG81]. molecular [LPS83a, LPS83b, LSP85, LCCB81, LCULC87, LCP81, LL84, Lin81b, Lip88, LCD82, LW82, LW84a, LSH88, LP80, LHH^{+}83, MR80a, MSC^{+}83a, MMS^{+}86, Mat81b, MLL83, Mcc83, Mez83d, Mez86b, Mic85, Mic86, Min80, MS80b, MK81, MSWC88, MMC83, MR80c, MR85, NSN89, NS84, OI86, OMCD81, Ojh89, OM82c, Pal89, PMT84, PP83, PBFL80, Pee83, PABL88, Pip89, PDMC80, PDC80, PDC81a, PDC81b, PKC85, PH87, PD89, PBGT80, PC86, RNW86, RL86, RL88, RB81, RW80a, Ran83, RB89, RR87a, RR88, Ric88, RP89, STH80, Sch82, See81, SK80, SG80a, SS86b, SJH89, Sim82b, Sin80b, Sin86, ST86b, SG87, Ska84, Sme84, SHKP80, SC80b, SG85a, SLT81, ST86c, SC84b, TNY86, TF83a, TSL80, Tho85, TN80, THKM81, TMC84, Tru83b, Tus86, VM81]. molecular [VM83, VŻH82, WKH80, WB81, ZP82, ZY85, Zho88, ddF86, vDTBN80, AEHA87, HOT87a, HOT87b, KD80b, VARPU84, Cal86a, Cal88, Tap86].

Molecular-orbital [AEHNEA85, MBR80, Hil81c, MR80c]. molecule [AC85, All87, ADF^{+}88, BBH87, BNK86, BD81, Blo83b, CLA80a, CKM81, CG80, CT84, CT98, DSB84, Del88, DWQ86, ECM89, FGC84, FD89a, FF85a, FF85b, GS89, GSC89, Gás86b, GVB80, Hur82, IS80, JPAB87, Jor81, KK82a, KEB87, LRZ84b, McW84, MSI80, MASF88, MPD87, Mül82, NSL88, Nor83, Par84, PČP82, PGPRN86, PNA86, RKHK88, RKO^{+}89, SH89, Sap80, SOS84a, SOS84b, TNY86, TNT^{+}86, TNB86, TW88, TWL82, UI80, VM81, WKH80, WJ80, Wea84, ddF86]. Molecules [BB87, Ågr87a, Ågr88, AB84, AM83b, Ano87a, AM88b, AH86, AA89, Bar85, BSZ81, Blo89, Bor85, Bou80, BHVR83a, BHVR83b, BTC^{+}83, Bro88, CBV85, Cal82a, CPMP86, CKD81, Che80, CZM83, Chi81, CJ89, DM85, DS89a, DMB89, DBS80, DBVMB85, DSW80, DVV80, DA88, EG84, FGS85, Fuk81b, GTB80, GM81c, Gro85, GRBP87, GSO^{+}87, HR84, HS84a, HBC86, HO85a, HO85b, HC81, Hir80c, HZ83, HH89, HS88, Hub85, Ish86, Ish88b, JTST87, JEW88, Kag83, Kah84, KKZ86, KG86a, KG86b, KH85, KHP81, Kau87, KFD85, KTK80, KRST83, KRBM85, KNS^{+}86, KKSM87, KNSM88, KMPTK89, Kle83, Kob82, KRCN80, KCC80, KL81, KTS86, Kun86, LPS83a, LSP85, LM82a, LPC^{+}80, LL82a, LFL^{+}81, Leu83, Lip88, Loh87, Mal82, Mar86a, Mar86b, Mar88c, MBT80, MS84, Mez84c, MLA80, Moh80b]. molecules [Moh80a, ME81, Nes87, NB79, NT80, NC87, Nov83a, PL80, PGG80, PXLC86, PGM80, PPMG89, Pie84, Pie85, PFLB83, PM86, Pyy86, Pyy93, Pyy00, RB88d, RTC81, RACPL86, RR87b, Röm83, RVK80, SMD83, SWW89, SM85, Sen80b, Sen85, SNNG86, SG89, Sie83a, SSK84, SB81, SB83b, SHL86, SC80b, Sza85, TLDK89, TLT^{+}84, TP89, TPN86, TST89, TWW81, Umr89, VVD80a, Ver80, Voj85, WW82, Woo80, YUKK84, YS80b, YS80c, ZNS80, Živ83, ZHB^{+}87, Bra86b, Ågr82]. Möller [PHGR88, PFLB83, SS87]. Moment [DVV80, LM82a, SSR^{+}87, VU85]. momenta [Mar85a, Mar86b, SBWW80]. moments [AJY80, CLM80, Cio87c, FZB86, JD85, Mak80, MRJ85a, MRJ85b, RL88, Røe81a, SJ84, SHK87b, SMHK89, TPN86, WKH80]. Momentum [BIN86, HSSW83, Kog84, Ols86, STS82, AL86, ASL88, Bri86, Cal89b, DT81, DSB84, DD87a, GB85a, IRA87, MNS85, RFST85, RI88, SSK84, SST84, TP89, TMC84, YKZ^{+}89, YLM80, ZLH87]. monitoring [Set79, Sie81]. monoamine [MSC^{+}83b]. monocarbides [Nec83]. monochalcogenides [ADS83]. monoclinic [KAA85]. monocompounds [TP80]. monolayer [MH80a, MS80a, ZLP81]. monolayer-phase [MH80a]. monolayers [DC82b, DBC84, MB89, ZP82]. monomer [UT80, WHT86]. monomers [PSS86]. mononitrides [Nec83]. Mononitro [HKLM85]. monophosphate [Gre82]. monophosphates [GK80c]. monopole [MT87]. monoradical [EBEE87]. monothioformic [TLC87]. monovalent [LCDP80, Pac83, UTP82]. monoxides [Nec83]. Monte [DKLP81, DS89a, LG84, CC83, CFC83, CC84b, DBF89, MB86, Mcd81a, MMB84, MK81, RBL84, RBH^{+}86, RC80a, Rom81, RJK83, Umr89, VR86]. Monte-Carlo [Rom81]. Morokuma [Pic83b]. morphine [CZ81]. morpholine [LHK81]. Morse [BPLB87, Cha82, LAGS89, RPS80, YG87]. motifs [SW81]. Motion [KN84, Ade87, DL83, Emi83, GS89, GSC89, HS80a, Hir80c, HJ80, Lin81b, LL82b, MB86, SCH84, SF80a, Wal88, Yar82]. motions [Fis84, LJK86, Shi87, Toy80]. Mott [Kve82, SC89]. Mott-type [Kve82]. mouvement [TV84]. movement [IHI80, TV84]. moving [KO80]. MP2 [MHAS86]. MP4 [PHGR88]. MR [Yam80b]. MRCI [Ess84, She87, SSP^{+}88]. MRD [KGEP89, KHRK87, KHK87, RKHK88, RKO^{+}89]. MRD-CI [KGEP89, KHRK87, RKHK88, RKO^{+}89]. MS [RC80b, BF81, CPGP83, OM81]. MSM [SRSR85]. MSW [SRP81]. muffin [BC84, BC85a, BC85b, RLS87]. Mulliken [Sie83a, May84, May85]. Multibody [FN84, NC84]. Multicenter [WJ82, Gus85, PŻ85, Sme84, SF80b]. Multichannel [AL81, GSL86]. multicomponent [Hos80]. multiconfiguration [BS81a, CHKB83, Mat81b, McW83, MN81]. Multiconfigurational [AJY80, JSYO83, SS80d, IS85, JSDK89, MRJ85a, MRJ85b, OJY83, RS80]. multidimensional [BHM87, CP86b, FC84, Mez81b, Mez83a]. Multimode [CKD81]. Multiphoton [LJW87, MS86, BSD82, Chu86, SAM82]. Multiple [KEIG86, Lin81b, PBFL80, Blo83b, CY80, CEZ83, DG86, EI81, Gra82, GBO86b, IST89, LHC80, PMT84, TMO86, FF84]. multiple-scattering [Blo83b, DG86, LHC80]. multiple-well [Gra82]. multiplet [CDG80a, DG86]. multiplicity [VARPU84, Wol88]. multiplier [NDBE82]. multiply [RGW88]. Multipole [RL86, SIN80a, BEL89, ES80, Pie82, RL88, Sen84, SHK87b, SMHK89, TT83]. multipotential [KHRK87]. Multireference [KHCH85, KB88, Yam80b, BBO83, Bur83, CHKB83, LB83, LSBS81, MJ89, RSMM86, SBB87]. multishell [Tai89]. Multistep [SRSR85, SSR86, SSRR87]. multistreak [BEL89]. Mum [Cal88]. muon [TSA86]. muonium [STG86]. muramic [Yad83]. muscarinic [WMP^{+}83]. Muscle [Ben81, MSM81]. mutagenic [PKP^{+}81]. mutagenicity [LSH88, SVS81, SJH89]. mutant [RB88d]. mutations [KBR83]. Mutual [Con88b]. my [Kot84].

N [CMDBVMA89, Pic83a, Sch81, Tew87, TWL82, BH89b, DMB86, GC84d, Gou86, JY81, KL80, ST89, Wor81, TR86, AM84, AYBP83, ADT^{+}89, BF83a, Che80, CZ81, CZM83, DF89, Don82, DD89, Gre87, GT83, HDFL84, Ish86, Ish88b, JSDK89, KU80, KPSK80, LH85, LHK81, MS87, MTWO83, MK80, MA80b, NC84, PKP^{+}81, Rér89, RT84, Sie83b, SRL^{+}83, Tew87, Tew88a, Tew88b, TR86, Yad83, ZW83]. N* [RR87b]. N-acetyl [Yad83]. N-acetyl-glucosamine [Yad83]. N-alanylcarbonyl [Tew88b]. n-azines [BH89b]. N’-dimethyl [TR86]. N-glycylcarbonyl [Tew87]. N-heterobicyclic [Gre87]. N-methyl [JSDK89]. N-methyl-N-nitrosourea [KPSK80]. n-methylacetamide [KL80]. N-nitrosamine [DD89, TR86]. N-nitrosamines [RT84]. n-propylamide [ST89]. N-protonated [CZM83]. N-protonation [CZ81]. N-substituent [CZM83]. Na* [Kun86]. NaCl [BEP81, DDP81b, PC83b]. NAD [PSG83]. naloxone [CZ81]. Namur [ADL78]. naphthacene [MTI80]. naphthalene [WK83]. naphthalenesulfonate [WS80a]. naphthyl [ST89]. naphthylamines [Gre82]. naphthylnitrenium [Gre82]. Náray [Tap88]. Náray-Szabó [Tap88]. narrow [KWBC80, Kun81]. narrow-band [KWBC80]. Native [SF89, BF81, LSS83]. NATO [ADL78, Cal81c, Cal82a, Cal86b, Dun80, Fro88a, Kel78, NB79]. Natural [HSB88, Poh82, Poh84, RW80b, RW81, Lar86a, Löw84, MSC^{+}83b, Mül81, Yam80a]. Nature [Bre80, Bry80, CBV85, ES88, PNGL83, PNA88, ST89, Suh86]. NbC [TP80]. NDO [BS80b]. Near [Dat82, Ada80b, Ada86, AHB^{+}80, De 80, Fis82, JBW87, Kun86, LK81, YVK^{+}80]. necessary [BD85, Kry82c]. Need [KKA80c]. negative [Bec82, CB82, Del87, MF87, SOS84a, SOS84b, Sin89]. negatively [Sil82]. Nenad [Brä86a]. neocarzinostatin [SG81b]. neon [AB87, Bec82, PS84, SI89a, YUKK84]. neopentane [YUKK84]. nephelauxetic [PNA88]. nerve [DBCJ86]. nervous [Cow82]. nested [Mez86b]. net [YHG89]. Network [Mez84e, KSSH85]. networks [Koh86, Mez83b]. Neumann [FC81c]. neurotoxins [Gri81]. Neutral [CGT86b, SMHK89, Bec82, Csa81, KL80, VGG88]. neutrino [KJMS85]. Neutron [Chi81]. Newns [MBD84]. Newton [Sch89, YKZ^{+}89]. Newton- [Sch89]. NH [BFOP83, Del87, KBT80, KKTE80, CPS89, GP82, KTS86, Mül81, SJ88]. NHOH [Sch83a]. Ni [RB88b, JR85, RSGK88, RLH86]. nickel [FS86c, SR84a, WD81]. NiCO [RW80b, RW81]. Nicolaides [Gos80]. nicotinic [YHG89]. NIDDM [RB88d]. Ning [Ban82a]. nitride [DPR80, LL86]. nitrides [PT89a]. nitriles [MA80b]. nitroaniline [PD85]. nitrobenzylidene [PD85]. Nitrocubanes [HKLM85]. Nitroexplosives [HKK^{+}82, HKKM83]. nitrogen [Gan81b, GG81, GC85a, KB86, KKC^{+}86, MG83, RR87b, RJK83, SF81a, TL83, Yam81, Živ83]. nitrogen-activated [Yam81]. nitrogen-containing [RR87b]. Nitromethane [CSB86b, Bar86, KHCH85, RKHK88, RKO^{+}89, SHK87a, SRL^{+}83]. nitropolycyclic [LSH88]. nitropropene [DT83]. nitrosamine [DD89, RT87, TR86]. nitrosamines [RT84]. nitrosating [RT85]. nitrosation [RT84]. nitroso [LHK81]. nitrosourea [KPSK80]. nitrosoureas [Lew88, SS83]. nitrous [FD89c, RT85]. NMP [ZY85]. NMP-TCNQ [ZY85]. NMR [Gee89, BH89b, FC80a, FC81a, Fuk83, Gol88, Gup80, GB81, GSS85b, GUSS89, Khe81, MS86, MSM81, MG80]. NO [CHKB83, RKO^{+}89, SRL^{+}83, Fro81, Fro83a, BF83a, MSI80, Nak83, KHCH85]. noble [CADL86, SSF^{+}84, TSA86]. noble-metal [SSF^{+}84]. nodal [SSK84]. Nodeless [Hun80, Hun81]. NOE [GSS85b]. Non [RLS87, BC84, BC85a, BC85b, BOW80, KHRK87, WJP83, SS85]. non-Bloch [BOW80]. non-degenerate [WJP83]. Non-muffin [RLS87]. non-muffin-tin [BC84, BC85a, BC85b]. non-protonated [KHRK87]. non-Watson [SS85]. nonabelian [PS88]. nonactin [GP82]. Nonadditive [GPN80, Nov83a]. Nonadditivity [CKLR80, SR80, FM86]. Nonadiabatic [KL89, Pos83, DM85, Dun87, MR80a, Röm83, Sme86, Yar87]. nonadjacent [EBO83]. nonalternant [PSIT87]. Nonalternate [EB82c]. nonanes [RW80a]. nonbenzenoid [EB82c, EBEE87]. nonbijective [KN84]. Nonbonded [RSP^{+}80]. nonbonding [PSIT87]. noncanonical [McD81b]. nonclassical [MR80c, PSIT87, TK86, TIS^{+}88]. Noncovalent [ACGT89]. noncrossing [ZRY83]. nondegenerate [Kal85, MJ89, WJP85]. nondiagonal [Ort89]. Nonempirical [BLR81, BDK86, BCBS81, Ran84, WS80b, HPK81, LHK81, SHP^{+}80]. Nonenzymatic [SRC84]. nonequilibrium [DC82a, KCP80]. noninsulin [RB88d]. nonintercalating [KMRK86]. noninterfering [Sla83b]. noninvasive [Gup80]. Nonlinear [Bes88, CB82, MD82, SCH84, Dat85, DR84, HC80, Kir89, Lun89a, OJY83, RDH84, SC86a, SC87, Sin81, Toy80, Tru83a, WGS85, WBM88]. nonlinearly [BS83b]. Nonlocal [Lun89a, GR89, SL86, Ulm85]. nonmetallic [ADS83, Woo82]. nonparabolicity [CE88]. nonplanarity [BKJ88, SY82]. nonpolar [Hir80c]. nonrelativistic [Bli80, Bli84, Köh89]. nonrelaxed [SB87]. nonresonance [SHS84]. nonrigid [LM82a, SB81, SB83b, TNY86]. Nonrigidity [KKD80]. nonseparable [Hal80]. Nonsinglet [CK82]. nonstationary [BHS85]. Nonuniform [Per82, IAAA89]. nonunitary [Mal86]. nonvariational [HK86]. nonzero [YKZ^{+}89]. Normal [SZR80, GMP^{+}84a, LSS83, RB88d, RDG84a]. Normal-mode [SZR80]. normalization [MMRY88]. Normalized [NP81]. Normally [HYZ89, CS87]. Nos [Zer87]. Note [TS82, Wój86a, Wor81, FF82, IS85, Kau87, Sab80b, LCÖ84]. Notes [Ågr88, Fro81, Fro83a, Ing80, Ten80b]. notion [Del81a]. Novel [EB82a, EB82b, EB82c, GK80a, BT84, CNT89, DM83a, Lar86b, Živ83]. NP [FD89b]. NPF [FD89b]. NS [KG86a, KG86b]. Nuclear [KKA80b, Mez86b, AC89, And87, BLV88, BB81b, DL83, Ela87, Fis84, GSMRM85, GS89, GSC89, HL88, HLPV83, KL80, Mez81a, Mez82, Mez84d, Mez87b, MSGPP80, MS80b, PG80, PAFW84, SH89, SC86b, SG81a, Wal88, Wea88]. nuclei [GS89, Sen85]. Nucleic [HKG81, AN86, BVGV89, CC82a, DPZ80, FCPL84, GSS85b, HP81, LCCB81, MY81, MLA80, MGBH81, PSS^{+}87, PP80, SM80, SGP^{+}80, Tew87, Tew88a, Tew88b, SHK87b]. nucleophile [BP84]. nucleophiles [FS86b, LPCF80]. nucleophilic [CSB86a]. nucleoside [CS81, SRO81, SC80a, Sar81, SP81, SP82a, SC84a, SP86b, Sar89]. nucleosides [SC84a]. nucleotide [PT89b]. nucleotides [JY81, SGP^{+}80, Yon85]. nucleotidyl [SW81]. nucleus [AGB82, BZA84, IQ87, MA80b]. number [ČD87c, KPM83, MC88, ZLH87]. numbers [BNS^{+}87, CS82, EBO83, KEB87, Rec80, SMKT86]. Numerical [Bot81, DSB84, IAAA89, MS84, Mes85, PČP82, PBA88, Tel88, AB84, AB87, AM87a, BVB88, CP86a, DD87a, GKMV80, HC81, LPS83a, LPS83b, LSP85, LP87, RB88c, RI88, Tho85, WB89, Yon85, DD87a, JMD88]. NY [Mår82].

O [AM82, BMMS81, BDK86, CF81b, Dat82, ES81, EHK83, GLLOB86a, JRK84, Mar81a, PXLC86, PR83, SOS84a, SJ85, Sch80a, Sin89, SS86b, SS86a, AJY80, BMMS81, BF83a, BBL81, CP85b, DPC82, Deá83, DF89, FN84, Gan81a, GT86, GT87a, GLLOB86b, GLLOB88, GC87, GA82, Ish86, Ish88b, INO88, KS89b, KTS86, LCCW87, NT80, PKP^{+}81, Rér89, SJ88, SAS86, SS80e, Wój86a, YOJ81]. o-benzosemiquinone [SS86b, SS86a]. O-methylation [SS80e]. O/Pt [PR83]. observables [Ran80d, TV83b]. observations [Eri89, NCH^{+}89]. observed [Gup80, MG80]. obtained [Cze87, GSMRM85]. Obtaining [MM82, CTV80, CT81, VTT81]. obtainment [DQ81]. obtention [TV84]. OC [SHP^{+}80]. occupancy [DS80b]. occupation [CS82]. occurring [ECM89, JL85]. octahedral [ACCYCC80, Bha82c, RD84]. octanitro [HKLM85]. octic [VCV87]. October [Kel78]. OD [GT87a]. off [CF81a, CE88, FC82b]. off-center [CE88]. off-diagonal [CF81a, FC82b]. Office [Cal81a]. OH [AM82, CF81b, ES81, GT86, GT87a, KS89b, LGL^{+}84, SHP^{+}80, Bau86, Del87, Kry86, LHK81, SOS84a, SOS84b, SS89a]. OH-transition [Bau86]. OHCl [Wój86b]. Ohno [Pic83b, Pau89a]. olefins [CSB86a, FZZ87, LY83]. oligomers [DPS87]. oligonucleotides [GP83b, WMS81]. oligosaccharides [BR81b, Tva89]. omega [SDBDSMG86]. Omission [CV82a]. oncogenes [LČ84]. oncoproteins [LS85]. One [BLG85, De 83, FW80b, Lás85, PŻ85, Tin80, BB82b, Bha81b, CC84a, Cio89, DB81, DFP80, DR80, EB87, FA86, Gru85, Hal87, HSB88, HM89a, HT85, Hun86, Kah84, Ker81, KS87a, KPS87a, Kve82, LC83b, LP85, MR80a, MZ85, MC89, MST^{+}80, MPB84, PSM87, RP81a, See84, TIYH86, TH81a, TNY83, Tel88, TPN86, Tus86, Wei88, ZW83, Hur82, LPS83a, PDKC84, Sto82b]. One- [BLG85, FW80b, PŻ85, MPB84, PSM87, Sto82b]. one-body [DB81]. One-center [Tin80]. one-component [Kah84]. one-dimensional [CC84a, DFP80, DR80, HM89a, Ker81, KS87a, Kve82, MR80a, MC89, TIYH86, TNY83, Tel88]. One-electron [De 83, BB82b, Bha81b, EB87, HT85, Hun86, KS87a, KPS87a, LC83b, MZ85, RP81a, TH81a, TPN86, ZW83, LPS83a, PDKC84]. one-matrix [LP85]. One-Parameter [Lás85]. one-particle [Cio89, Gru85, HSB88, Wei88]. ones [KRST83]. only [GTB80, Gla85, LRT88]. onto [KPH88]. OP [FD89b]. Open [HK86, Mes80, BB82b, Bha81b, Big86, BK82, BS81b, Bor85, CS82, HČ83, Mes85, Obc84, PL80, Pan81, Sch80b, FCB80]. open- [PL80]. Open-Shell [HK86, Mes80, BB82b, Bha81b, Big86, BK82, BS81b, Bor85, HČ83, Mes85, FCB80]. Opening [Fri82, Man81, Wij83a, Wij83b, KHRK87]. Operator [SPRS89, Bak85, Böh83, Bor82, CMM83, DMB^{+}84, FS86a, HYZ89, HYS89, KS82, KS84, Kut84, Lar83, MD87, MS84, Niu80, Özk84, Pan89, PBFL80, Tay85, Tim86, WS81]. operators [Alt89, ACH83, Bal82d, BPLB87, Bro80, CDG80b, CA80, Dal80, Dem86, GC84c, GC84d, Hub85, KPP85, KPP86, Kat89, Lar83, Löw82a, LFM89b, MZ85, ML81, MPS86, PM83, Pan82a, Pan82b, Pan85, QX89, Ste81, SLT84a, SLT84b, Tai89, Wei88, Wit81, ZS85, ZS86, WS82]. opioid [CZM83]. Oppenheimer [Röm83]. Optical [AH86, Gan80, Gan81b, SST83, TM83, Vol88, BMMS81, Big86, BH89b, GSL86, GF88, GBO86b, Has85, HLR80, LHZ80, Suh84a, VTT81, Vol89b]. Optimal [SH87a, VR86]. optimization [BSZ81, BMBK82, DPC82, Dem83, DR88, HGPF89, HWZ88, JMD88, Kar86, Kar84b, KEGS81, LWÖ81, PG80, PM81, RTM81, RTM83, RMM83, RTM86, SS80d, Ten86b, WH87, WBM88]. optimized [ČV86, GSO^{+}87, HOT87a, HOT87b, Mor88, PG80, PDC81a, PDC81b, SM84, ZNS80]. optimizing [CP86b, JWOÖ82]. optimum [Kar84b]. optogalvanic [KKCM85, KKC^{+}86, KKCM87]. orbit [Ågr82, Har87, HS80a, HBC86, RTC81, VP84]. Orbital [Coo87, GTB80, LCC80, MRJ85a, MRJ85b, Yam82, Ågr87c, AEHH81, AEHM85, AEHNEA85, AEHA87, AN86, AØ80a, AØ80b, AØM81, AL86, Bal81b, BB82a, BR80, BMBK82, Bha82c, Böh84, BSCR80, CS81, CD89b, CF81b, CH82, Das82, Del88, DCF82a, Duk82, DO84, Edw88, EIMS87, FC81a, Fer88, FS86b, FS87, Gal82, GC85a, GKMV80, HOT87a, HOT87b, HTD80, Hib81, Hil81c, HM89b, INFU84, IAAA89, JT85, JSDK89, Kar81, KO81, KS82, KTGR89, KD80b, KK82b, LER87, Lew88, Lib84, LZY83, LSH88, LP80, Mal86, MBR80, MR81, MR80b, Mül81, MR80c, OS89a, OS89b, PP83, PKP87, PBFL80, PD84b, PKC85, PSK^{+}81, PM81, Qia83, RB81, SC80a, SP81, SP82a, SC84a, SP86b, Sch82, SG80a, SJH89, STS82, TS80, TP85, Ten88b, TN80, TMC84]. orbital [Yam80a, ZLH87, Živ86, Živ87c, ddF86, vDTBN80, LHP80, RB88a]. orbitales [SV85a]. Orbitals [Ger86, AB87, AM87a, ADF^{+}88, ACCC80, BSZ81, BVC80, BYBZB83, Bri86, BVAB81, Cal85, CEB86, CP85a, CA80, CP83a, DBES80, Dov86, Ess86b, FC80a, FC81a, FFF88, GM83, GKMV80, Gus85, Gus87, Her83, HDFL84, Hir80b, HSB88, Jon80, Jon81b, Jon81a, Jon81c, Jon82, Jon83, Jon85, Jon86, KCK83, Kau82a, KHK87, KS87a, Koe86, LPB82, LL84, LC81, McD81b, Mey88, Niu84a, Niu84b, Nov83b, OI86, OK80, OL82, PMT84, Pip85, QX86, RFST85, RLR88, RW80b, RW81, RSST86, SB87, STS83, SV85a, SF80b, SNSM82, SBWW80, TMO86, Yan85, ZLH87, Zho88, Nat86]. Order [CZ82a, Wu82, Ano82s, BNK86, BEO80, Ben82, Bha82a, BP80, ČV82b, CZ82b, CZ88, Cul88, Cul89, DF83a, Don85, ERB82, FF82, GV88, GY87, Ham82, Han88, Hir80b, HG86, KK86, KU80, Löw86d, Luk82, LW81, May84, May85, MMC83, MPD87, OJY83, Ort88, Ort89, PT84, PHGR88, RM82, SF81a, Sve88, TP84, Tay87a, Wil80b, YS80a]. order-disorder [RM82]. order-equivalence [FF82]. order-isomorphism [FF82]. order-isomorphism/order-equivalence [FF82]. Ordered [Wit81, HYZ89, MY81, SGP^{+}80, Zun85]. Ordering [EBO83, Gro85, MB83, Ret86, EBEE87]. orders [FEKC84, May86b, May86a, Sim82a, Tay83, Tay84, Tay86a, Tay86b]. Ordinary [Eng83a, Eng83b]. Organic [Cal86a, BHVR83a, BHVR83b, CGT86b, DR80, PL83, PL84b, RZ86, See81, Sey83a, Yam82, dSdMB86]. organization [Koh86, SGP^{+}80, ZP82]. Organizer [Jor89]. organometallic [Böh85, SCER88, WH87]. organophosphorus [HLKK82, LHKK82]. Orientational [MR87]. oriented [JD85, Khe81, NTM80, TM83]. Origin [HZL84, Bed84, Dos84, DQT80, DQ80, DQ81, JBW87, Löw84, Ste84, You84]. Origins [Fox81]. ortho [JI88, KKA^{+}83]. Ortho-disubstituted [KKA^{+}83]. ortho-lithiation [JI88]. Orthogonal [Lim87, DMB89, FPT81, Iga85]. orthogonality [JT85, Kau82a, PLL86]. Orthogonalization [Cal89b, AEG80, BM86, Pip85, Rut81, GBW89]. orthogonalized [MÖ80]. Orthogonally [PP89]. orthonormal [Cio89]. orthonormality [PM81]. Orthonormalization [Cal85]. orthorhombic [KAA85]. Oscillating [Poh80]. oscillator [AMFC83, Bha80, Bri87, CF81a, CV85, FC87, Gan80, Gan81b, GY87, GC81, GC82a, HZBR89, MPB84, MPS86, Özk84, PM83, PSM87, Rei80, Sil82, TS82, VCV87, Wes82, YOJ81, YG87, NT80]. oscillators [BS83b, BHS85, ČV86, Gra82, HY89, Lef83, LAGS89, RPS80]. oscillatory [VGG88]. other [ES88, HBC86, Kau82a, OWÖ81, PP81, RKHK88, RKO^{+}89]. our [GC83]. out-of-plane [SK87]. Overlap [MG85, Cla83a, FSP87, Jon80, Jon81c, Jon82, LZ87, Pan83, Qia83, ST80a]. Overlapping [Blo83b]. overlayer [Boe88]. Overview [KSSH85, Sim82a]. oxatriazole [JSDK89]. oxene [PL83, PL84b]. oxetane [KHRK87]. oxetanes [KHK87]. oxetanone [Tho80]. Oxford [Ågr82, Ågr85, Gee89, Kum87, Mou89, Rat87, Rat88, Rös87]. oxidase [MSC^{+}83b]. oxidation [PSG83, SG82, Wol88]. oxidation-reduction [PSG83]. oxidations [CD89b, PL83, PL84b]. oxidative [Yam81, YKTA80]. oxide [Ell84, FD89c, LJK86, SA86b]. oxides [EGP86, LPCF80, PT89a, Sch88, SS88, WBK86]. oxidized [KKTE80]. oxirane [FS86b, FS87, KHRK87]. oxo [CD89b, CD89c]. oxo-catalyzed [CD89b]. oxocarbons [VFA81]. oxopregnane [VT82]. oxopyridine [DKLP81]. oxopyrimidine [LOB85]. oxy [Yan85]. oxy-ions [Yan85]. oxyanions [SA80]. oxygen [GG81, GSRG80, GSR83, LCCW87, Min80, RZWL80, SF81a, Sin89, SG85a, Živ83]. oxygen-liganded [LCCW87]. oxyheme [LHZ80]. ozone [ARZ^{+}89, CKK86, MASF88, Yam80a].

P [Ågr82, Ågr88, BDK86, Cal89a, Fro88a, Fro88b, KS89b, Kun86, LC82a, Mar81a, Rat88, Sme84, Tap86, Che82, CS83b, GT86, PD85, PD87, PO85, CLL86, Gup80, LL85, MMB84, MG80, NS82b, NC84]. P-450 [CLL86, LL85]. p-benzoquinone [CS83b]. p-dimethylamino [PD87]. p-dimethylaminobenzylidene-p-nitroaniline [PD85]. p-nitrobenzylidene-p-dimethylaminoaniline [PD85]. P450 [LCCW87, PL83, PL84b, WCLZ86]. P700 [OMC82]. pA [GJMLRGSA85]. packet [ABCV86]. packets [BS83b]. Padé [BEO80, Bha80, Bha82b, CB88, Cio88a, JSSS80, Rei82]. pages [Zer87, Fro83a]. Pair [CG80, GSO^{+}87, SS80b, AB84, AL87, FPT81, GAL86, GBZ86, HS80b, IL84, ITKD80, IPV81, JP80, KHNT86, MGN82, McD80, Nes87, OMC82, Pul83, SS84a, UB86, vDTBN80]. Pair-excitation [GSO^{+}87]. pair-wise [ITKD80]. paired [PKP87, Rut81]. pairing [AØ80a, Kar85, May86c, Yam82]. pairons [Lar86a]. pairs [CLS81, CS83b, CCD^{+}84, CAML87, DDFA87, GjC86, HP81, KJL87, Lar86a, PS81, SS84a, SS85, SSYR87, YS80a]. Pairwise [NP80]. palladium [BASB86, GPN80]. pancreatic [LSS83]. panel [Jas80, Löw84]. papain [DOW87, DOW89, LPW83, vDTBN80]. para [BK84b]. para-substituted [BK84b]. Paradoxes [Gás86b]. Parallelism [CCD^{+}84]. paramagnetic [FBM80, KG85, KG86c, KG86d]. Parameter [Lás85, DS87, FC82d, Nag87, ZNS80]. parameter-dependent [FC82d]. parameterized [VP84]. Parameters [HKFW87, BV80, Deá83, FC80a, FC81a, NKZ88, PV89, SK81, Tat81a, Ten80a]. parametric [FC87, FSK79]. Parametrization [Kar84b, SF80a, JIS87, LWK84, KN80]. parametrized [LHP80]. parentage [FW80b]. Part [Cal89a, Ano85c, ML83, ÖL83]. Partial [DS80b, Ort88, SST84, MTWO83, NMK84, Ort89]. Partial-wave [SST84]. partially [BB81b, KKTE80, NTM80, TN89, TM83]. Participants [Ano88i, Ano88j, Ano82s, Ano85c, Ano86f, Ano87d, Ano86e]. Particle [Mat81a, Živ88, AW83, AL86, ASL88, BFC80, Cio89, DL81, Ess86b, FW80b, Gru85, HS80a, HG86, HDO88, HSB88, Jar89, Kar89, KPM83, KN84, Kut80, Lar83, LWZ88, LFM89b, MM87b, Mor89, Odd82, OT80, PW86, Pos83, Wei88, Wor85]. Particle-hole [Mat81a, Živ88, DL81, Odd82]. Particlelike [FY82]. particles [Dem83, Gir80, KPP85, KPP86, PD83, RM87, Sto80]. particular [CCC^{+}89]. Partition [Kia81, GJ83, LSH88, PC88, Sla85]. partitioned [Kau83, SRL^{+}83]. Partitioning [Löw82c, GP86a, Gou86, HL83, RLPFA86, Tay83, Tay84, Tay86a, Tay86b, WS81, WS82, HS88]. Partitions [Sch83b, LCLC86, LCLC88]. partner [CT87]. parts [KTK80, KRST83, KRBM85, Rei82, Sil82]. passage [UTVP85]. passivation [DAL89]. Path [Mar88a, BM89, EBEE87, Hil81a, MPP^{+}81]. Pathological [DC87]. pathophysiology [RB88d]. paths [BB84, Mez87b, PL83, PL84b]. pathway [BDR^{+}88, HSR86, KHCH85, PSBB83, SS82]. pathways [KHRK87]. patterns [JM81, MSC^{+}83a]. Pauli [FWS^{+}85, BG88, Bon81, KM89a]. Pauncz [Cal82b, Wor81]. Pb [SRC84]. PBG [SRS81]. PbTe [HLPV83]. PCILO [BP80, LPW81, LW82, LW84a, LW84b, LWK84, Tew88a, VT82, Yad83]. Pd [AS86, DK87]. PdCl [LOR84]. peak [CPMP86, PGG80, Suh86, MPP83]. Peierls [Kve82, MW87, OBMU82]. Peierls- [Kve82]. penetration [BEL89]. pentasaccharide [YL83]. pentasulfur [TWL82]. peptide [AEHM85, GG83, HM84, HKG81, LGL^{+}84, MDO^{+}82, OLB^{+}82, RB81, SFC86, SY82, UV83, VU85, Vol88]. peptides [HKFW87, MGBH81]. peptidoglycan [Yad83]. Perdew [Csa88]. perfect [BOW80]. performance [KST84]. perhalogenated [CKNG80]. peri [SL87]. period [Blo89]. Periodic [Car82, Ess82, Blo89, DPR80, Dov86, FC81c, KRT83, KKA80a, LW82, LW84a, MS80a, PABD80, PD80, Pur88]. periodicity [LW82, LW84a]. permettant [SV82]. Permutation [Kva82, DH84, Gal82, HDO88, Pan83, Ret86]. Permutation-Group [Kva82, Ret86]. Permutational [FS80, KPP85, KPP86]. permute [Hal87]. peroxidase [AWCL89, LCCW87, LCA89]. peroxidation [SAB^{+}80c]. peroxide [CS87]. personal [Wil80a]. Perturbation [Au82, Aus84, Ben82, CLM80, Cha80, FC84, FC85, HHT82, HČ83, LS85, Mal82, RPS82, Suh83, TPČ83, Ada82, AČP82a, ACP84, AT85, Ano82s, BAG83a, Bha81a, Bha82a, Bha82b, Bue86, CS80b, CP85a, CMM83, CSK87, ČV82b, CV85, CZ82a, CZ82b, CZ88, Cul88, Cul89, DBCJ86, DP81, DS88, Gol84, Gra82, Ham82, HB86, HČP88, HA84, Hir80b, Hub80, IS85, IPV81, JJČ87, Kal85, KCK83, Kle87, KB88, LJC83, LC80, LB83, LGS87, LC88a, Löw82c, LN89a, Mak80, Mat81a, McD81b, MJ89, MMC83, Muk86, Muk84, MPD87, NBE87, PT84, PČP82, PS84, PHGR88, Sch83b, SI89a, Sel88, SF81a, Sil80, Sil82, Sim82a, SS87, SH87b, TNY86, TP84, Tay83, Tay84, Tay86a, Tay86b, Tim86, Vrs87a, Vrs87b, Wes81, Wes82, Wil80b, WJP83]. perturbation [WJP85, Wu82]. perturbation-dependent [HA84]. Perturbational [IA87, LL86, Sil80]. Perturbations [SWW89, Ing80, MT85]. Perturbative [DBES80, HZBR89, BPT88, OL82, Ish89, McD80]. perturbed [AS89, CTV80, CT81, ENH^{+}87, HČ83, LW82, LW84a, RPS80, Sad83, Sil80, Wes82]. Pethig [Mår80]. Petr [Fro81]. PGA [KD80b]. pH [MG80]. PHARM [KKH^{+}83]. PHARM-MATCH [KKH^{+}83]. Pharmacological [FHH^{+}87]. pharmacologically [BJR86]. pharmacology [GJ83, Kau81, KKH^{+}83, Pal89]. pharmacophores [KKH^{+}83]. Phase [BB81b, Dat82, DK82a, Mou89, Sie81, TN89, ACP82b, BBLR84, BL81a, BSD82, BBO83, BFC80, BS83b, DS81a, DC82b, GRS^{+}82, GLP83, Hal83, Hir80a, Hos80, II82, KKA^{+}84, KS87b, MH80a, SO85, ST89, Set79, SS81b, Sta71, Sta87, UT80, WBM88]. Phase-coexistence [BB81b]. phases [Sti82]. phencyclidine [WMP^{+}83]. phenol [SHL82, WCE89]. phenolic [BDP^{+}81, Sey80]. phenols [LaF89, MG89, SG80a]. phenomena [Ber88, DR84, DS80c, DBC84, Har82, Hic83, LM86, Ras88, RDH84, Sew86, Sta71, Sta87, Mou89, Kum87]. phenomenological [Dun87, TT81]. phenotypic [GMP^{+}84a]. phenyl [SJ88, SJH89]. phenylalanine [KM80]. phenylene [LJK86]. phenylethylamine [AEHA87, Luk81]. phenylfurans [AEHH81]. phonon [Sin86, SDS83]. Phonons [Cal81b, CTP83, LVL^{+}85, ST86b, WS79]. phosphate [KVR85, LN89b, NL89]. phosphatidylcholine [DS81a]. phosphodiester [BVGV89]. phospholipid [ZLP81, ZP82]. phospholipids [Lal81, SPSG85]. phosphonitrilic [FFF88, FD89b]. phosphorescence [MMI^{+}82]. phosphorus [RDG84a]. phosphorylation [LHKK82, YKTA80]. photo [Yam81]. photo-oxidative [Yam81]. photoaddition [FZZ87]. photochemical [LC83a, NH84, OYN80, San84, TMON80, YMO80]. photochemistry [BKP83, CPN89, Mår82, TR80, YG86]. photocycle [ES88]. photodissociation [KL81, SM82, SM83]. Photoelectron [CKNG80, CNK87, KNS^{+}86, KKSM87, KNSM88, KMPTK89, KKA^{+}83, KKA^{+}84, NA83, Big86, Böh82, CKK86, KTK80, KRST83, KRBM85, KS81, LaF89, MW83, MKW85, PXLC86]. photoemission [BN80, CD85]. photoexcitation [HDFL84]. photoexcited [NTM80]. photofragmentation [LGS87]. Photoionization [FK80, VC83, FSW87, HM86, LPC^{+}80, Lar83, MLL83, MÖ80]. photoisomerization [Sme86]. photon [BP86, SH89]. photons [MDT85]. photoreaction [SD82]. Photoreactions [Pes83]. photosynthesis [Ave84b]. photosynthesizing [YVK^{+}80]. photosynthetic [ES88, RKV^{+}80, Tat81a]. photosystem [YVK^{+}80]. Physical [Ban82b, IRP81, Pic83b, Ada64, Ada69, Ada80a, Bel80a, ČPV81, CL86, DD87b, HZ83, LČ84, Laf81, Löw86e, Lun82, Nat88, Vas82]. Physicochemical [YKTA80, DBL84]. Physics [Cal81b, Cal89a, Bal82d, Bri84, Cho80, ČV82b, Ela87, HR83, See88, Var86]. physisorption [GK80b, Kre81]. pi [Don85, ST89]. Picture [Ano87t, Ano87s, AH86, DR80, GC83]. piecewise [Kau82b]. pigments [San84]. pilgrimage [Kot84]. Piron [Gre81a]. pK [XYSJ81]. planar [GLP80, LR89, TKY89]. Planck [Chu88]. plane [LOGC86, MÖ80, SK87]. planning [Mez84e]. Plasma [KO80]. plasmas [Fla81]. plates [Fis82]. platinum [BC80, PGPRN86, AAW82b]. plausible [LCP^{+}88]. Plenum [Cal82a, Cal82b, Cal86b]. Plesset [PHGR88, PFLB83, SS87]. plots [JY81]. plus [HSR86, Kau83]. PN [FD89b]. pocket [YVK^{+}80]. Point [BS84, GCP87, SBE^{+}83, BFOP83, CB82, FW80a, FW80b, FS80, GC87, Kib83, KBR83, KAA85, LLC83, Mey88, RBSR84, RSBR85, RSD82, Tai89, MMN87]. point-charge [RSBR85]. points [BB84, Ban88a, EB86, GYJ83, GBZ86, LRT88, RB86]. Poirier [Ågr87c]. Poisson [Csa86b, LPS83b]. polar [BSL^{+}80, GGDG86, GGDG87, PCA89, PS80a, SWW89]. polarizabilities [ADF^{+}88, ES80, GY87, GGM81, Kir89, Mak80, MMC83, MPPW82, Rér89, SK80, SS81a, Sen84, SIN80a, SJ84, TT83, WKH80, YOJ81]. polarizability [DBD^{+}85, DDFA87, MMC83]. polarizable [KS87b]. Polarization [RKV^{+}80, BVD81, Boč89, EC83, FNC86, GOS87a, Gee88, KWJ83, MDO^{+}82, MB84, MMC83, Odd82, Öhr85, Öhr86, RDG^{+}86, SD86, SST83, TTRL88, TSA86, WG80]. polarized [BPCT83, RR87b, YG86]. polaron [Emi83]. polarons [KSS81]. poles [KEB87]. poly [Bur85, CLP83, GLP80, JY81, PS81, CLP83]. polyacene [KLS89]. polyacetylene [KKA80b, KAM85, KWM86, Mar85a, MB80b, OBMU82, SI89b]. polyacetylenes [Dov84]. polyamino [GGM81]. Polyaminopolynitrobenzenes [HKKM83]. polyaniline [DPS87]. Polyatomic [AM87a, ME81, CR81, DM85, DSB84, GSC89, JS88, KFD85, KL81, LPC^{+}80, SH89, SM87a, YY83, Sie83a]. polybutatriene [Kir89]. polychloroethylenes [DO84]. polycyclic [BBR82, MBT81, PBMR^{+}84, SVS81, Sin88]. polycytidine [LS80]. polydiacetylene [Kir89]. polyelectrolyte [PP84]. polyene [PB82, Suh86, TPČ83]. polyenes [BPT83, Big86, BKJ88, CK82, Dia86, FEKC84, GBO86a, Has85, Has86, KC82a, PC83a, PCG83, PT84, PP83, TP84, TP85]. polyethylene [BSCR80]. polyfluoroethylenes [DO84]. polyglycine [Bal81a, KKA80a]. polyhedra [Kin86]. polyhedral [BL88, Loh84, TL86]. polyhedron [Mez85c]. polymal [Pos83]. polymer [BDD81, DL82, FDAC89, IA87, KS86b, TIYH86, Yar82]. polymerase [MLC^{+}81]. polymeric [ABD81]. polymerization [KHK87, THKM81]. polymers [ADL78, BSL81, Böh85, BCBS81, CB88, DFP80, DAPS81, Dov84, Duk82, DF83b, HM81, KKA80c, Ker86, KSSH85, Lad83, LK87, MW83, Nag82b, PSIT87, See81, SRC83, Suh83, Suh84a, TK86, TIS^{+}88, dSdMB86, TIS^{+}88, Cal81c]. polymorphic [BDR^{+}88, SR83]. polynomial [CP86a, KJT85, MKVP80, RBK85, ST83, Vrs87a, Vrs87b]. polynomials [BR85, BL88, GC82a, KT86, KT80b]. polynucleotide [SW81, SRC84]. polynucleotides [HKG81, ITKD80, SGP^{+}80]. polypentapeptide [VU85]. polypeptide [BCV81, FY82]. polypeptides [Bal81a, Gup81, LJHU89, Pin88, Urr88]. polyphenylene [LJK86]. polyphthalocyaninato [RB88b]. polysequential [JY81]. polysilane [Min87]. polysulfanes [TL80]. population [May84, May85, SH87a]. populations [Gas86a]. porphin [SF81b]. porphinatoiron [RGDF85]. porphine [KO81]. porphyrin [SK87]. porphyrinatonickel [Böh84]. porphyrins [RDG84b, RDG84a, RGDF85, WCE89]. Portrait [Ano89-28]. position [NMK84, RFST85]. positions [GBZ86, GC84f, SRH86]. positive [BZA84, CB82, DBCJ86, MMM81, Van86, Van86]. Positron [Ori81b, FBM80, Ori82a]. possibility [MF87]. Possible [KBR83, ZNS80, BPT88, BDR^{+}88, DS87, DA81, JY81, KMKC88, MP87, MPP^{+}81, PP82, Pra83, RT87, ST82, FCPL84, LOBS86, TR86]. posteriori [BDD81]. potassium [BLOW89, SA86a]. potency [DM83a, DBCJ86, RB88d]. potent [YKTA80]. Potential [BBL81, BBAT82, CLS81, DAPS81, GC82a, KNM83, Nom80, RBSR84, RSBR85, SB83b, TOK80, AGB82, AM89, BZA84, BB85, BD88, BPLB87, BSR80, Buf82, Bur85, CP86a, CSB86b, CLP82, CLP83, CM81, Csa83, DB81, ENH^{+}87, FN84, FF89, GSL86, GLLOB86b, GLLOB88, GM86c, GC81, GC83, HKP81, HLKK82, HKLM85, HS80a, HL81, Hel82, HH89, HR87, HM89b, ITKD80, JS88, JW86, JŻ85, JSS80, JEW88, KT80a, Kau83, KN84, KT84, Kob82, KM89a, Kry84b, LM82a, LB86, LCDP80, LCP81, LP81a, LHK81, Mez81b, Mez83a, Mez84c, Mez84e, Mez85a, Mil86, MT87, MS80b, MSI80, NHKY83, Nal80, NSL88, NCH^{+}89, NCKL81, NC84, OMCD81, OM82c, PD83, RJK83, RVK80, RS82, SH89, San84, SMD83, SM85, SWM82, SHKP80, SHP^{+}80, SRL^{+}83, SS80f, SG85a, ST86c]. potential [TS82, TB83, TH88, THKM81, UI80, Umr89, Urr88, Var87, Var81b, VGG88, WKJ88, ZLP81]. Potential-Derived [RSBR85]. Potential-energy [Nom80, TOK80, RVK80]. potentially [PKP^{+}81]. Potentials [FDAC89, AØM81, Bak85, BEL89, CC83, CPZS86, Eis88, FC81c, GRS^{+}82, GSS85a, GC82a, GC84f, GBO86b, HBSC83, HK86, HPKW81, Hun81, Kah84, KPH88, KAD81, KM85, KK87, KDA81, LJC83, LLS83, LPW81, MM87b, Mez85b, MB84, NP81, PC88, PZ84, PS80b, RNW86, RL86, RL88, RZP84, RLPFA86, Sch80a, SWW89, Sen85, Tin80, VTT81, WSD80, WW89, WJ80, Wea84, Win83, Yon85]. Poul [Ågr87b]. power [EÖWZ84, Gos82a, KE82, NS83, WK85]. powerful [JWOÖ82, RT85]. powers [FS80, SO85, Smi89]. pp [Cal86b, Cal89a, Gos80, Mår80, Wol81, Cal86a, Rös87]. PPP [Has82a]. PPT [Ott86]. PPT-MCF [Ott86]. Practical [DOW87, MMRY88]. prebiotic [BCV81, LGL^{+}84]. predicted [Bak85, RDG84b]. predicting [HLKK82, LHKK82]. Prediction [Coh86b, KKH^{+}83, SHH^{+}83a, Kau81, Sin80b, TN89]. predictions [SHPK80, Zun85, LCCW87]. predissociation [Chu86, COS82, DBVMB85, GDBRV87, Moi81, Mül81, TNY86]. Preface [Löw80, Löw81b, LCG82b, LCG83, SSS^{+}86, Ano89m, Jor89]. preferences [Tew87, Tew88a, Tew88b]. Preferred [RMM83, RTM86]. Preliminary [Ano88y, Bou84, Min87, NCH^{+}89, Wol88, Aus84, CHKB83, NW82, JD85]. preparations [BHS85]. presence [Var81b]. present [Löw89, SOS84a, SOS84b]. Press [Ågr82, Cal86b, Gee89, Kum87, Mou89, Pic83a, Rös87, Sie83a, Tap86, Ågr85]. pressure [Dun80, ESCC85, Kel78, McM86, TSA86]. Preuss [GSO^{+}87]. Price [Gos80, KST84, LM81, Lun82, Mår80, Mår82, Pic83b, Pic83a, Sie83a, Wol81]. Prigogine [CDL82]. primary [Abd80a, Ave84b, DOW89, San84]. principal [Odd82]. principle [Bon81, Coo87, CTV80, CT81, Dal80, Hub80, Jac80, Kat80, Mat81b, Mez84b, Tac89, Wea82, WOT89, YKZ^{+}89]. principles [Fuk81a, Suh84b]. pro [CBV85, CBV85]. probabilities [SMJ83, YKZ^{+}89]. probability [Dea89, SHL86]. Probable [LČ84]. probe [GMP^{+}84a, MY81, NPL89b]. probes [GB81, RM87]. Probing [JY81, Mår82, TR80]. Problem [Lin81a, ADF^{+}88, ACDLH84, AZY84, Ave86, ALH86, Bli84, BERE85, Cio87c, ČD87c, CS87, CF83, CL85, DM83a, DS87, Dun87, GBW89, GC84c, GC84d, GC84e, GC85b, Hal80, Kar81, Kob86, Lal81, LZ89, MMRY88, MS89, McD80, MBKH83, MR80c, Pan84, Pan89, Per85, Per86, PS80b, PSBB83, Sel86a, Sel86b, Sil86, TV89, WW89]. problème [TV89]. problèmes [TV83b]. problems [Ade87, AM88b, Ble82, DMB^{+}84, Dem86, DBES80, Edw88, FCO89, Gás86b, GJ83, HS88, JT85, Kim81, Kou80, Lan84, MSWC88, RM82, SRT81, Ska84, Tho89, TV83b, TKR86, Wei82, Woo82]. procedure [BB84, BBO83, GCP82, GPC85, GCP86, Iga85, JWOÖ82, KPS87a, Mat87a, MR81, MR80b, OM81, Pin88, Zer89]. procedures [BMBK82, Dal80, HPK81, HWZ88, LGS87, LHK81, PZ84, PP89, RB88c, SHP^{+}80, Ten86b]. Proceedings [WJ82, ADL78, Don79, Jor89, Kel78, NB79, Pul79, PTG80]. process [AL80, Ave84b, CGG81, CS82, HM86, Jas80, KU80, Sey80, TOK80]. processes [BLG85, CGK82, CSK87, Cla80b, DBL84, FK83, Kha85, LC83b, LCD82, MS86, MLL83, See81, Ulm83, UVP^{+}81, Var85, ZJŻ88]. processing [Tay87b]. processor [SBE^{+}83]. produce [Dun88]. produced [GSMSG82, IDK87, RZWL80]. Product [Ori82b, DP81, JJZ^{+}88a, JJZ^{+}88b, Mor88, Røe81a, SPÖ89]. production [Gas86a, SAB^{+}80c]. Products [Kat89, Bal82d, CP86b, LER87, Wit81]. professor [Ano87s, Ano89-28, Fox87, LGJLCB89, Ohn84]. profile [AM83b, GM81a, PGM80]. profiles [DT81, Smi89, Tha83, UTVP85]. program [BH89b, GRBP87, KKH^{+}83, SSP^{+}88, BHVR83a, BHVR83b]. Progress [Cal86a, Faz87, Kar89, SSP^{+}88, Tru83b]. Projected [PS87, TF83a, TP80, Gal82, KC82b, KC83, Kar86, oLCC81, Löw83, SGEK81]. Projection [Alt89, AL86, BHM87, CV85, ČV86, ČVV87, LP87, SS86c, SS86d, SS87, Ste81, Tai89, Tim86, VCV87, ZS85, KS84]. projections [EC83, FNC86]. projective [Alt89, Her84]. projectors [Nag82a]. prolate [Csa86b]. promotion [Tho86b]. proof [May86c, Wor81]. propagation [KHK87]. Propagator [Win82, Bli80, BS81b, Bor82, Bor85, Don82, EC83, FNC86, GOS87a, Gee88, KWJ83, Odd82, Öhr85, Öhr86, Ort87, Ort88, Ort89, RDG^{+}86, Sim82b, TR82, WG80, WG82, WK85, Wei85, Bak85, JWÖ83]. Propagators [Yar82, Bli84, Dal80, Hir86, Lar83, PM84a]. propanelike [WW82]. Proper [GYJ83]. Properties [ADL78, Kun81, LP85, MPD87, Wal88, AT85, ADS83, ARZ^{+}89, Ano87a, AL87, AØC82, AH86, AWCL89, BSHK86, BEO80, BMMS81, BR81b, BEP81, BYBB84, BHVR83a, BHVR83b, BH89b, BCBS81, BCV82, BV86, CV82a, CS80b, CPGP83, CJZ^{+}86, ČPV81, CL86, Coh86b, CA86, DS81b, De 83, DD87b, DF83a, DRS83, ECM89, Eps87, EB87, EHK83, FB83, GSG84, Gos82b, GEdP81, HOT87a, HOT87b, HSSW83, HA84, HLR80, HM84, HZ83, IS83, JL87, KC83, KK87, KKTE80, LCC80, LPP84, LP84, LPW83, LRZ84a, Lip88, LFM89b, Mår80, MR80a, MBT80, Mez84d, Mez87b, MPPW82, NS84, Nat88, NDBE82, NB79, PAFW84, PC83b, Pet79, Pie82, Pit84, Poh83, PDKC84, PSIT87, PP82, PH82, RP81a, RW80a, Ran80b, RK82, Ran83, RD84, RFST85, SMD83, SS83, SAP81, SB87]. properties [SZR80, STS82, Suh84a, TK86, TIS^{+}88, WD81, YKTA80, ZP82, ZW83, Živ87a, JS86, Ågr87b]. property [GMS81, Gro85, Mez83d, DPC82]. propiolactone [Tho80]. proportional [CG86]. Proposed [CD89a, CM85, Csa89, KEB87, LCA89]. propranolol [PKH^{+}81]. propylamide [ST89]. prospects [DBS80]. prostaglandins [GRP^{+}80, KD80b, Kot81, Kot82]. protamines [PAFW84]. proteases [ST88]. Protein [HP81, HKG81, CGG81, Chu88, Fox81, Fra89, Gas81, LSS83, SSR86, SSRR87, TF80, Urr87, NSN89]. Protein-ligand [NSN89]. proteinases [NS82a, NS83]. proteinoid [FNPS82]. proteins [BEP81, BL89b, CCG^{+}82, FHH^{+}87, HKFW87, LCCW87, LJHU89, PP81, Pin88, SS84a, SS85, SB88, SKL80, YRS85]. protobiological [Mat82]. protocells [Mat82]. Protometabolism [SG84]. Proton [MSM81, MC89, MB82, PK88, Sch80a, Sch81, BP83, BCD89, DBCJ86, DOW87, DOW89, IHI80, KSK86, KJL87, KPTG88, KPT88, KS86c, LS86, LCA89, MP84a, NSP80, ÖHD89, PMCA88, SO85, San84, Sch83a, SSB83, SRH86, SD88, SWW89, SS86b, SG85a, FNC86]. protonated [CZM83, FS86b, FS87, KHRK87, KSK86]. Protonation [MBKH83, CZ81, KHRK87, LCF89, XYSJ81, KHRK87]. protonic [Loh85]. protons [BH89a, FC81a]. prototypes [DC86]. pruning [BD88]. pseudobinary [Zun85]. pseudoconcept [Del81a]. pseudocrossings [OM82c]. Pseudodegenerate [EB82b]. Pseudopotential [SSF^{+}84, Sza85, TLDK89, BL81c, DF86, Gás83, GG85, KSO84, Mor83, WBK86, GPN80, Bra86b]. pseudopotential-CI [WBK86]. pseudopotentials [Coh83, GG81, GG82b, GG82a, GG83]. pseudoresonances [WOT89]. pseudorotation [LMG85]. Pt [PR83, BDK86, KKTE80]. PtCl [LOR84]. Publications [Ano89-29, Cal81a]. publishers [Ågr87c]. Publishing [Ågr80, Ågr87a, Cal86a, Dun80, Gos80, Mic87, Ågr87b]. puckering [Wol88]. Pullman [LM81, Mar81a]. Pulsed [KKCM85, KKC^{+}86, KKCM87]. pulses [BSD82]. pure [Gos82b, KS82, Was89]. Purine [LGPP^{+}80, KBR83]. purines [Cal84, Eis88, SS81b]. purity [LW81]. putative [CLL86]. pyramidal [RK82]. pyran [FA86]. pyrazofurins [CS81]. pyrazolones [KKA^{+}84]. pyrene [PZPL86, SG80a, UT80]. pyridine [BH89b, IHI80, TW88, YZ85, TS81]. pyridines [POK85]. pyridinone [Sch82]. pyrimidines [BCD^{+}80, Eis88, SS81b]. pyrolysis [GMP84b]. pyrrole [IHI80, LP80, RDG84b, SRS81]. pyrrolidine [Sap80]. pyruvate [KK83, MSM81]. Pyykkö [Ågr88].

QCLDB [OYH^{+}80]. QSAR [BNS^{+}87, CNT89, MFM89, RSNT88]. QSARs [Lew88]. quadratic [CV85, Niu84a]. Quadratically [HGPF89, Sla85]. quadruple [Cul89]. quadrupole [JD85, Mak80, SS81a]. qualitative [BW80, SL81, TP87]. Qualitatively [Dav81]. quality [DPC82, GG82b, GG82a, LT89, Mcd81a, PDMC80, PDC80, PDC81a, PDC81b, PDMC82, RSST86, RST87, STS83, SP85]. Quantal [GVVDB89, CMDBVMA89, DBVMB85, YKZ^{+}89]. quantique [Cal81a, SV85a, SV85b]. quantiques [TV83a, TV83b, TV84, TV89, Van85, VT85, Van86]. Quantitative [AABW80, AAW82b, BJR86, GDR85, SL81, Sey81]. quantities [BAG83a]. quantization [AL81, GSL86, Gir80, HA84, LAGS89, Ols86]. Quantum [BNK86, Ber88, BS83b, BCV80, BVGV89, CBV85, Cal89a, CLL86, Con88b, CAML87, Del83, FGMM80, GGDG86, GGDG87, GD82b, GM83, Gol88, GJMLRGSA85, GK80b, GSS85a, HLKK82, H^{+}90, IMTG80, JL84, JRK84, KJMS85, LOBS86, LM89, LHKK82, LP81b, MG83, Mar89, Mar81a, MSC^{+}83a, MSC^{+}83b, MMI^{+}82, OM82a, OM82b, OM83b, OM83a, OM84, OM85, OM86, OMH89, Pul79, PTG80, RBL84, RBH^{+}86, SZ84, Sew86, SBE^{+}83, SPSG85, Tap86, VMMV80, Woo80, Yad83, Ågr87a, ACP82b, Ano84a, AZY84, Ave86, ALH86, AA89, Bar85, Bec89, Bel80a, Ben81, Ber81, BQS87, BERE85, BV86, Cal82c, CDL82, CD83, CSK87, CB85, ČV88, CCD^{+}84, CT84, CT98, CE86, CE87, CE88, De 80, DFP80, DAPS81, Dew88, Dun87, Eri89, FC81b, FC81c, FC81d, FC83, FC82e, GS89]. quantum [GPLP85, GOLC87, GJ83, Gos80, Gre80, Hal80, HKLM85, HCCB86, Hay82, HM85a, HM81, HYS89, Hun80, Jac80, Jas80, Jor89, JJ83, Kau81, Kib83, Kim81, Kot84, Kot82, Kou80, KLDD84, KPS87a, KPS87b, KPS88, KEB87, KB87a, Kun83, LOF83, Lan84, LM81, Lar86b, Law87, LFL^{+}81, LOGC86, LCD82, Löw82b, Löw86c, Löw89, LN89a, MN81, Mat87b, Mat87c, Mat87d, May86b, Mcd81a, Mez83b, Mez84a, Mez84b, Mez84e, Mez85b, Muj86, NB79, Obc84, Pal89, PNGL83, Pau89b, PS80b, PP81, PBM85, PSK86, PSG83, Pur88, Ran80d, Sad85, SBH^{+}85, SBH^{+}86, SIN80a, Sil84, STS82, SMS84, Sin88, Sla83a, Sla84, SV85a, SV85b, SvG89, SHL86, SF81b, SSN^{+}84, Tac83, Tew88a, Tho83, TV83a, TV83b, TV84, TV89, Umr89, Van85, VT85, Van86, VCB83, WS81, WS82]. quantum [Zer87, ZLH87, Živ88, ADL78, CA86, FHHE80, Hil81b, HZ83, Löw81a, OYH^{+}80, Ågr85, Cal81c, Cal82a, Gee89, Kum87, Pic83b]. Quantum-chemical [GJMLRGSA85, JL84, VMMV80, Mez83b, Mez84a, Mez84e, PNGL83, SMS84, Sla83a, Sla84]. Quantum-field [FGMM80]. Quantum-mechanical [JRK84, LM89, AZY84, HM85a, PBM85, SSN^{+}84]. quantum-mechanically [Ran80d]. quartet [LÖ85b]. quartic [ČV86, Rei80]. quartz [SC89]. quasi [Boč87, FF89, MST^{+}80, See84, SST83, Tus86, WJP85, Boč88b]. quasi-degenerate [WJP85]. quasi-free [SST83]. quasi-one-dimensional [MST^{+}80, See84, Tus86]. quasi-relativistic [Boč87, Boč88b]. quasi-stationary [FF89]. quasianalytic [Ori80]. quasiclassical [ES81]. quasidegeneracy [AJ83]. Quasidegenerate [Kva83, JP80, Kal85, Mat81a, MJW88, SF81a, SH87b, WJP83, Yam80b]. Quasifree [BMG82]. quasiparticle [BK84a, HS88, LH87, PPB87]. Quasirelativistic [BC80, KK85, FS86a, KK86]. quasispin [ČPV81]. quasistationary [Tac82]. quaternions [Alt86, Alt05, Rös87]. Quelques [TV83a]. Quenching [GMP^{+}84a, MMS^{+}86, Yar87]. questions [Ada64, DPZ80]. quinazolines [AABW80]. quinizarin [PBM85]. quinones [GSG84, Pul86]. quinoxalines [KKA^{+}83]. quotient [Cio87b].

R [Cal82a, Cal88, Dat82, Fro88b, Gos80, Mår80, Rat87, Wol81, CBV85, TNB86]. R-Matrix [TNB86]. rabbit [MSM81]. Racah [MPBN81]. radial [IPV81]. Radiation [Tap86, CT84, CT98, IDK87, MST^{+}80, MGBH81, SM82, SM83, Wal88]. Radiation-Molecular [Tap86]. radiation-molecule [CT84, CT98]. radical [Big86, CKM84, Nak83, SP86a, Sey80, SS86b, SAB^{+}80c, SL80, SDS83]. radicals [CV82a, DS81b, GSMSG82, GMP^{+}84a, Gas86a, SSUF80, SKS83, SS86a, SL80, SKS83]. radii [HČP88]. Radiobiology [Var86]. radiofrequency [KKC^{+}86]. radiological [Var86]. radiolysis [GSR83]. Radioprotector [BCV80, BV86]. radioprotectors [BCV82, VV87]. radius [Wea82]. Raman [BMMS81, SHS84, TW88]. Randic [VV87]. Random [CS83a, BBO83, BFC80, Lar83, MY81, PV89, RWG83]. random-phase [BBO83]. Random-walk [CS83a]. Range [EY82, BW80, CPS81, CP83b, DAPS81, GSRG80, GSS85a, Kog84, NPL89b, Pie84, Pie85, PABD80, RM87, YA87]. rare [ADS83]. Raster [Bri84]. rat [Fin89, Sey81]. rate [Bar86, Chr89, ES81, GT86, Kha85, MD82, NS82a]. rate-determining [Bar86]. rates [BL89b, GT87a, LAGS89]. ratio [BB81a, SH89]. rational [ČVV87, JSSS82, Löw82c, RSNT88]. rationale [Tac87]. ratios [BS82a]. rattling [Hir80c]. ray [CFG^{+}83, GM83, KB86, KPS87a, POT80]. Rayleigh [Tay83, Tay84, Tay86a, Tay86b, Bha81a, Bha82a, CMM83, Cio87b, Hub80, PČP82, Tho85, Vrs87a, Vrs87b, Wil80b]. Raymond [Ågr87c]. rays [JTST87, KTS86, TST89]. RDX [HKK^{+}82]. Re [Fro83a, KEIG86, Toy88]. Reaction [GT87a, KKZ86, ZCH89, Abd80a, BB84, BBGS84, BP84, BDR^{+}88, Bur86, Chr89, Cla86, Dew88, ES81, FGG83, Fro88a, Fuk81a, Fut87, GSM83, GTB80, GDX89, HZ84, Hil81a, JL84, JM83, KH80, KU80, KLDD84, KM89b, LLM86, Mez83b, Mez83c, Mez84a, Mez85a, Mez85c, Mez86b, Mez87b, Mic87, NHKY83, PMCA88, RC86, RBL84, Rom81, SS82, SBE87, Ser84, SA86b, SS80e, Sok81, SSC83, Tac87, YY83, YKZ^{+}89, BBLR84, PL83, PL84b]. reaction-field [Rom81]. Reactions [Hal83, Ågr83, BNK86, Bal82a, BLR83, BBPS83, BCD89, Chr80a, Chr80b, DÖL87, DC86, FZZ87, FS86b, FS87, FEKC84, Fuk81b, GHF84, HZL84, LC83a, MR87, NH84, Nal80, NP81, OLB^{+}82, PL83, PL84b, RN84, SD88, SCER88, SAB^{+}80c, Sme86, STG86, SZJS82, Van82, Var85, YMO80, Yam81, Yam82, GTB80, LWZ88]. Reactive [GSR83, LPP84, LV87, Set79, BR81a, BM89, Gir80, LP84, LPP81, Lin85, Lin89, PH82, Suc81, SB89, Var81b, Sie81]. Reactivity [LC83b, BW80, FCPL84, HZ83, KPH88, KKZ86, PP80, Ser84, SVS81, Sla85, dB84]. Real [Lin81a, BD85, Löw85a, Rei82, Sil82, Swe83]. Reality [Sil86, ZA89]. Realization [AL80]. rearrangement [PCA89, Ran80b, RK82]. rearrangements [RBR80, Sch83a, SSZJ82]. reasonable [BFOP83]. Reasons [BH82]. recall [Koh86]. Received [Ano80d]. recently [Csa89]. receptor [BVDA88, CZM83, DS80b, KTGR89, Luk81, MTWO83, MWO^{+}84, PMOW89, RMM83, RB88d, TLT^{+}84, VL89, YHG89, YGK^{+}89]. receptor-binding [RB88d]. receptors [Dav87, PMOW89, RWOT81, WMP^{+}83]. reciprocal [RB88c]. recognition [KRMK86, MWO^{+}84, Ric88, VB85, WMP^{+}83, Yon85]. recombination [Fla80, Fla81, Win85]. reconsideration [HET86]. reconstruction [Lás82]. recurrence [KPM83, SPRS89, MPS86]. recurrent [Pan82a, Pan82b, Pan83, Pan84]. recursion [SI89b]. red [Kry86]. redistribution [BSL^{+}80]. redistributions [Sch81]. redox [Gas85]. redox-mediated [Gas85]. Reduced [CSF80, CA80, GA82, PABL88, Ban88a, Cas88, Coh86a, Gos82b, GP86a, KPH87, Lar83, Löw86d, Luk82, LW81, MC80, ML81, Mez84d, Mez87b, Tho86a]. Reduced-diabatic [PABL88]. reductase [GC85a]. Reduction [Kat83, JL85, LL85, MPP^{+}81, PSG83, RS88, TF82, Niu80, Niu84a]. Reductions [Lin81a]. Redundant [Ban88a]. reference [BS81a, LT89, MJ89, Ort89, RS80, Sie80, SM84]. refinement [KR81, SSRR87, TN89]. Reflection [Mez87b]. refolding [Chu88]. reformulation [Röm83]. refraction [DQ81, DQ80]. refractive [MB80a]. regarding [Sla85]. region [Kry86, Low82d, Low82e, Sey83b]. regions [KSJT83]. Regression [SS80e]. regular [Pur88]. regularities [Ran83]. regulation [KMKC88, Vas82]. Reidel [Ågr87a, Ågr87b, Cal81b, Cal86c, Dun80, Fro88a, Fro88b, Gos80, LM81, Mar81a, Mic87, Sme84]. Reidle [Cal81c]. related [Ano81s, CFG^{+}83, FZZ87, GMS81, GS89, Hay82, KM80, LCC84, LM82a, Pit84, SA80, SS88, TWW81, VARPU84]. relating [BVDA88, WP80]. Relation [BZA84, Has82b, AN86, AGB82, Bri87, CBV85, CF81a, Dal80, DCF82a, EKBE86, GJMLRGSA85, HM89b, KM89a, LPP84, LPW83, Löw84, MP87, Nag82a, OWÖ81, RSSS81, YKTA80, DF83b, FSP87, Sim81]. Relations [Bar81, CF81a, Chr89, FF82, KPM83, KPP85, KPP86, Low82e, Mez83c, Mez84e, MPS86, OFB81, Odd82, RVK80, SPRS89, Shu84, Swe82, TS82, AM89, MA80a, May84, May85]. Relationship [AYCD81, CL83a, DA81, GDR85, KST84, LZY83, NC87, POK85, PO85, Sen80a, Sen80b, Sey81, Smi89, UB86, XYSJ81]. Relationships [SVS81, AAW82b, MSC^{+}83b, RP81b, FHH^{+}87, Lar86a]. relative [BSD82, CZM83, DBS82, EB82a, EB82b, EB82c, Fra84]. Relativistic [Bal88, Chr84, Das80, Gra84, HLPV83, Mar85b, MMM81, Pit84, Pyy86, Pyy93, Pyy00, Rud84, ST84, SI89a, VI84, ZSB80, Ave84c, BIN86, Boč87, Boč88a, Boč88b, Boč89, CPGP83, Dur86, Ess84, FSS^{+}89, Gre81a, HF86, HBC86, Kag83, Kah84, KFD84, LLS83, Mey88, MB84, PC88, ŠKR84, Suc84, Ågr88]. Relativistically [LHP80, VP84]. relaxation [Böh85, CG86, GGDG86, GGDG87, GLB81, KH85, Kub82, Lás82, LCD82, Pee83, ST80b, SG81a]. relaxed [Das82, JN88]. releasing [RTM86]. Relevance [Hay82, GSMSG82, RB88d, Lar83]. relevant [GSG84, Pal89, RT84, RT85, SG85a]. reliability [Ran80d, Cze87]. reliable [CS87]. Remarks [Cia81b, CS82, Fax87, Han86, Löw86d, Par84, And85, Buc89, Cla80b, Fax84, Fri82, Gás86b, GMV83, JL87, JSSS82, Löw89, LN89a, McW83, Wij83b]. reminiscence [Kot84]. reminiscences [Bat83, Wil80a]. removal [CDL82]. Renormalized [ČVV87, VCV87]. Rentzepis [Fro88a]. Reorganization [Böh82]. repair [FPDA89]. repeated [FW84]. repeating [UV83]. report [DBL84, Min87, NW82, SSP^{+}88]. representability [Gru85, GT83, Kry82c, Löw86d, Pan89]. Representation [AČP82a, CA85, CA86, Duc85, Fra84, Fra85, LP89, Mez84c, OT80, Ori82b, Pan82b, Ret86, RP89, Sch80b, Var85]. representations [Alt89, Coh86a, FS80, GC87, Her84, KS86a, OM82c, RFST85]. represented [Iga85]. repressor [KRMK86]. Reproducibility [FW84]. reproducing [KPS87a]. reproduction [Vas82, GPC85]. repulsion [DCF82a, GHGP89, Gus85, GCV84]. requirement [Pur88]. Research [HAmGS86, Cia81a, Var86]. residues [CBV85, CLA83b, MSM81]. resolution [CKK86, Gee89, Gol88, Lim88]. resolved [PZPL86]. resolvents [HH87]. Resonance [Ran80a, RSG^{+}87, And87, Ave84c, Don82, Don85, EB81, GSL86, Gas80, Gas81, GSG84, HDFL84, Hir86, INO88, KL80, KSSH85, KS89a, LL82b, Löw85a, Löw86e, MMI^{+}82, MSGPP80, MW80, MR80c, PAFW84, Poh83, PPP87, Sim81, SPSG85, WW89, Win86, WHT86, Živ86, Živ87c]. resonance-theoretic [KSSH85]. Resonances [Ela87, GVB80, Sjö87, BS80a, Bow86, Brä87, CMDBVMA89, Chu86, CS83c, ENH^{+}87, FSW87, Jun80, KS89b, KEB87, KM89b, Moi81, NKM81, Poh82, Sie87, Wea83]. resonant [BR81a, CM85, KFN80]. resorcinol [SKS83]. Response [Obc84, BH89b, CE86, CE87, DSG80, HA84, Ish84a, JSYO83, Kub82, Obc87, Pra83, RSMM86, SS80a, SB84]. resting [AWCL89]. restricted [ÅCFRJ85, BB82b, DC87, Fin87, GM86c, KC83, KPS87b, KPS88, Mes80, Mes85, KC82a, PC83a, RBL84, CK82]. Restricted-Hartree [KC82a, CK82]. restricted-step [ÅCFRJ85]. resulting [Csa83]. Results [Sie83a, Aus84, Bou84, CHKB83, Cze87, FK83, GBO86b, Has86, KWM86, LPS83b, LSP85, LOF83, MKW85, ME81, Nag82a, REB83a, RW80b, RW81, SSR^{+}87, SCER88, SM80, Ten88a, DPR80]. reticulum [SAB^{+}80c]. retinal [KPTG88, SNNO86]. Reuse [Gre81a]. revealed [PH87]. reversal [Mey88]. reversibility [SC86a, SC87]. reversible [PZPL86, Sch89]. Review [Ågr80, Ågr82, Ågr85, Ågr87a, Ågr87b, Ågr87c, Ågr88, Bra86b, Brä86a, Cal81a, Cal81c, Cal81b, Cal82a, Cal82b, Cal86b, Cal86c, Cal86d, Cal88, Cal89a, Dun80, Fro81, Fro83a, Fro88a, Fro88b, Gee89, Gos80, Kum87, LM81, Ler85, Lun82, Mår80, Mår82, Mar81a, Mic87, Mou89, Muj86, Nat86, Nat88, Nor83, Pic83b, Pic83a, Rat87, Rat88, Rös87, Sie81, Sie83a, Sme84, Tap86, Tap88, Wol81, Zer87, Cal86a, CCS84, KEB86, LL89, REB83a]. revised [ACGT89]. revisit [EBEE87]. revisited [ABD^{+}80b, ČV83a, CV85, Har87, Kub82]. RF [KKCM85, KKCM87]. rhenium [KEIG86]. RHF [Big86, HČ83, WS80b, ZW83]. RHF-CNDO [Big86]. RHF-CNDO/S [Big86]. rhodopsin [SD82]. RI [Fin89]. ribonuclease [Chu88]. ribosomes [JG81]. riboxanthylic [GLP80]. Ricatti [Au82, FCO89]. Richards [Ågr82, Mår82]. right [JY81, Vol88]. right-handed [JY81, Vol88]. rigid [RPS82]. Rigorous [BF83b]. Ring [GMOBA80, BBGS84, DM83a, GC85a, Gre87, HS80a, Has82c, KN84, LJK86, Sap80, SHH^{+}83a, TKY89]. ring-shaped [HS80a, KN84]. rings [EB82b, KHRK87, LK87, LMG85, Sin89]. Ritz [Tho85]. RKR [RZP84]. RNA [MLC^{+}81, SR81, SR83, SRC84]. Robb [Pau81]. Robert [Sie83a]. rock [TP80]. rock-salt [TP80]. Rodney [Ågr87a]. Role [KD80b, TTRL88, Vol89b, Dov86, FFF88, Fli86, Kat80, LCCB81, LSH88, LGL^{+}84, MTWO83, New80, PV89, SRC84, Yan85, SD86]. roles [CLA83b, GUSS89]. Roothaan [SP85, BB82b, Mar88a]. roots [Fox84]. rotagraphs [Cio87a]. rotated [Don82, EKBE86, NS82b, REB83a]. rotating [Hub85, MS86]. rotation [BS80a, BS82a, BK89, CS83c, FGS85, IL83, KAH80, KK82a, Kob86, Moh80b, Moh80a, Sim80, SHL82, ST86c, TWW81, SNSM82]. Rotational [Bow86, GDBRV87, NOMK80, Shi87, ARZ^{+}89, DS87, DC86, ESCC85, Kob82, Mic81b, PGM80, RZP84, SIJ85, HJ80]. Rotations [Alt86, Alt05, BTC^{+}83, Iga85, WW82, Rös87]. rotator [RPS82]. rotor [SHL86]. Routes [Bur83]. rovibrational [GPLP85, NPL89a]. row [Bec85, BF83a, DBS82, FN84, GTB80, JIS87, Kau82b, LSP85, NC84]. Roy [Ågr87c]. RPA [BPT83, BHVR83a, BHVR83b, Toy88]. RPAC [BHVR83a, BHVR83b, BH89b]. RS [vDTBN80]. RSPT [KU80]. Ru [CD89b, CD89c]. Ruben [Cal82b]. rule [Bon81, GK80a, LBP84, Mez86b, TL86, WBB80, ZRY83]. rules [BC83, CD89c, HAB83, SH89, Sin88, TP87, Win86, Yam82, Živ87d]. ruthenium [VG80]. Rydberg [Hic83, SAM82, San81].

S [Ågr80, Cal86c, PXLC86, Ser84, Sie83a, TL80, TWL82, SOS84a, SOS84b, Big86, DS80a, Deá83, DIV86, DS81c, FN84, HM89b, KEB87, LCLC88, LC82a, NS82b, Nov83a, PDMC82, PDKC84, SMD83, SM85, TL80, Var87, Wor81]. SAC [NH81, Nak83]. saccharides [TK83]. saddle [BB84, LLC83]. saddle-point [LLC83]. Sadlej [Muj86]. salicylic [MG89]. salt [KKTE80, Pac82, SGP^{+}80, TP80]. salts [KL80, SDS83]. Salvetti [CL85]. SAMO [DO84]. sample [SO85]. samples [TM83]. sangivamycin [SC80a]. Sanibel [Ano87u, Fax87, Han86, Löw83, Löw85b, Pau89a, ZS85]. Sasaki [Pau89a]. satellite [LdMH80]. satisfy [MM82]. saturated [Ver80]. Saxena [Ågr80]. scale [BEO80, BHVR83a, BHVR83b, Cow82, Kun83]. Scaled [BB82b, Bha81b, CHB88, HCCB86, KEB86, Lib84]. Scaling [Gás83, Bha82a, Brä86c, Chu86, CZ88, FC80b, KD80a, KPS87a, KPS87b, KPS88, Lev89, Löw86e, Mar85b, PSK86, Sim80]. scalings [GSMRM85]. scattered [CS89]. Scattering [CS80a, JTST87, KTS86, REB83b, TST89, Ano81s, BS80a, BR81a, BLG85, BM89, Blo83b, BERE85, BLOW89, CY80, Chi81, DG86, DYD89, EI81, FZB86, GLB81, GSRG80, GVB80, HF86, HL83, IL84, IST89, Jor81, Kau83, KB86, Kon83, KFN80, KEB87, LHC80, Lin85, LV87, Lin89, Mel87, Nag85, Nes85, NC87, Nor83, PBFL80, POT80, SKIK82, SHKM80, Sim81, SST83, SHS84, Suc81, SB89, TNT^{+}86, TMO86, VTT81, Var81b, Wea83, WOT89, YLM80]. SCF [DFH84, SRP81, Ada81, AM84, AGCT87a, AM87a, BPT83, BB82b, BSHK86, BVD81, Bha81b, BVC80, BS80c, BC83, CDG80a, Cha80, CS82, CL83b, CA85, CA86, CP80a, CHZ82, DMB^{+}84, De 83, DCF82a, Eis88, FB83, GjC86, GPC85, GK80c, HO85a, HZBR89, HC81, Iga85, Kau83, KHRK87, Kut80, LN89b, Mat81b, MSV89, MS89, MSI80, NL89, NHKY83, NS84, Nov83a, OF80, PXLC86, PM81, RW80b, RW81, RJK83, Roo80, RC80b, RLS87, SP86a, SSS81, SD86, SHPK80, SRL^{+}83, SRHK83, SP85, SG85a, SR80, Živ86, Živ87c, ZHB^{+}87]. SCF-CNDO [GjC86]. SCF-type [Kut80]. SCF-X [PXLC86]. Schaeffer [Ågr85]. scheme [AM84, DIV86, DRS88, FDAC89, LFM89b, Ori85b, OK80, SRHK83]. schemes [FW80a, FW80b, Kar86, oLCC81]. Schiff [BKP83]. Schluter [Cal81b]. Schmidt [Rut81]. Schrödinger [Tay83, Tay84, Tay86a, Tay86b, ASL88, BH86, BHM87, Bha81a, Bha82a, Bot81, CP86a, CMM83, ČV86, Dat85, Hub80, LPS83a, LPS83b, Lim80, Mas87b, PČP82, SCH84, SC86a, SC87, Sch89, Ská83, Tac82, Tel88, Tru83a, Vrs87a, Vrs87b, WB89, Wil80b]. Schrödinger-type [SCH84, SC86a, SC87, Sch89]. Schwinger [WOT89].

Science [Shu84, Ågr87c, Cal82c, CCC^{+}89, Fau85, Hay82, Urr82]. Sciences [Mår82, Pic83b, Ano84a, Kot84, Löw84]. Scientific [Ågr80, Bat83, Cal86a, Kot84, Löw85b, Wil80a]. scorpion [Gri81]. screened [BD88, WW89]. screening [BYBZB83, CLP82, CB82, Jon81c, LCDP80, VVD80a]. search [Ban88a, CS89, DR88, RJBR^{+}87]. Second [Ano88z, Don85, Gir80, OJY83, YS80a, BEO80, DBS82, ERB82, FN84, GV88, GY87, HS84b, Han88, HA84, Hir80b, JIS87, KK86, LSP85, Luk82, NHKY83, NC84, BP80, MMC83]. Second- [OJY83, MMC83]. second-harmonic [HS84b]. Second-order [Don85, YS80a, BEO80, ERB82, Hir80b, KK86, Luk82, BP80]. second-quantization [HA84]. second-row [DBS82, FN84, JIS87, LSP85, NC84]. secondary [DOW89]. section [FK80, LGS87, NC87]. sections [Blo83b, BF83b, DSG86, FSW87, Gan81a, Gir86, HDFL84, MM87a, MW83, MÖ80, SM83]. Secular [Lin81a]. segment [TK83]. select [Nas83]. Selected [SD80, GT86]. Selection [Živ87d, DG85, Fin85, Fox84, Fox86, RSNT88, SBB87]. selective [GEdP81, GP85, Min80, Nes85, UTP82]. selectivity [UTP82]. selenium [BS80c]. Self [DDP81a, DDP81b, DS80a, Dos84, DS81c, EG84, GBZ89, Koh86, MM81, PMCA88, SRS81, SSSH82, WG80, Wil89, BDK86, CG86, Cio89, Den83, Den85, Eng83a, Eng83b, FPT81, Lef83, Lib84, LTLF87, LLC83, MM82, MN81, Mic83b, Oga80, Ori82d, Ort89, SGP^{+}80, SM82]. Self-Consistent [DDP81a, DDP81b, MM81, DS80a, DS81c, EG84, PMCA88, SRS81, WG80, BDK86, Cio89, Den83, Den85, Eng83a, Eng83b, FPT81, Lef83, Lib84, LTLF87, MM82, MN81, Mic83b, Oga80, SM82]. self-consistent-field [LLC83]. self-energies [Ort89]. self-energy [Ori82d]. Self-formation [Wil89]. Self-instructed [Dos84]. Self-interaction-corrected [GBZ89]. Self-organization [Koh86, SGP^{+}80]. selfregulation [LS85]. Semi [Sad85, dB84, Muj86]. Semi-empirical [Sad85, dB84, Muj86]. Semiclassical [BP86, Mic83a, BGT84]. Semiconducting [HM81, GMM81, Suh83]. semiconductor [Ber88, Fin85, Her85, LFM89a]. semiconductors [BS84, CB83, FD86, Fer88, GMS81, GF88, KSS81, LVL^{+}85, MF87, Pes83, PW86, VVD80b, VI84, Zun85]. Semidiffuse [KG86b, KG86a]. Semiempirical [Bha82c, EÖWZ84, KK83, MR80c, NC87, POS81, Sen84, TN80, BB82a, BS80b, Böh84, CG80, DBS80, FC80a, FC81a, KS86b, KS81, LC81, MM87a, PG80, SHPK80, YHG89, Yon85, Živ83, BVD81]. Semigroups [Gre80]. semiorthogonal [HS89]. semiorthogonalized [OK80]. sensory [DSG80, Koh86]. Separate [Kob82]. separated [DP81]. separation [GM86b, KMW80, Tru83a]. separations [MS80b]. September [ADL78, Kel78]. Sequence [KRMK86, AN86, BB81a, FPDA89, Fox86, Ho81, Lás85, LM86, MPPW82, PS84]. Sequence-specific [KRMK86, FPDA89]. sequences [CBV85, GUSS89, LL82a, LM86, MB84, PT89b, UV83, VCB83]. Sequential [Sla83b, JJZ^{+}88a, JJZ^{+}88b]. ser [SNSM82]. ser-195 [SNSM82]. Serafin [Ågr80]. Series [Cal82a, DM83b, Dun80, Fro88a, AJ83, ADF^{+}88, BDP^{+}81, Bha82b, Cal81c, Cal86b, CZM83, DBS82, Fra85, GJMLRGSA85, Ing80, JBW87, KTS86, Kun83, LKB89, MG83, NSS86, ÖHD89, PO85, POT80, SB83a, Sil82]. serine [LN89b, NL89, NS82a, NS83, ST88, RC80a]. Serines [Rom81]. Serre [Fro83a]. serum [Gas80]. session [Gás86b]. Set [OI86, Ada81, AM83a, Ada64, AGCT87a, BSHK86, Bec89, BYBB84, BS86, CS83c, CFS86, CFS87, Del87, ERB82, Eri89, GG85, GBW88, GPC85, GSO^{+}87, Ish84b, IQ87, IS88, KGEP89, KG89, KO81, Kut84, LC81, MK85, McD80, Mez82, Mez83d, ÖHD89, RDG84b, RR87b, RST87, RBR80, RC80b, RS88, STS83, SV85a, SHKP80, SRHK83, SR87, SHK87a, SC80b, TPN86, WK83, Wei88, YG86]. sets [Ågr87c, AM83a, AN85, AGCT87b, Bag83b, DPC82, Fin87, FSP87, GT87b, GM81c, GCV84, HS89, HA84, INFU84, IO89, PDMC80, PDC80, PDC81a, PDC81b, PDMC82, PKC85, STH80, STS82, SF80b, TH80, WH87, ZLH87]. Setser [Sie81]. seven [EB82b]. several [AMFC83, GLLOB86a, GM86c, ČV88]. Sewell [Kum87]. sextic [ČV86, HH87]. shallow [DAL89]. Sham [HF86, Kel86]. Shape [AM88b, Buf82, KEB87, Tac88, AM88a, AM89, BR81b, Don82, Don85]. shaped [HS80a, KN84]. shapes [Mez87a, PM88]. Shaping [ZW83]. sheaf [Cho80]. shear [MD82]. Shell [HK86, NKV^{+}89, Bak85, BB82b, Bha81b, Big86, BK82, BS81b, Bor85, CS82, Csa81, CV83b, DG83, FCB80, HČ83, JMS^{+}85, KC83, KS85, Mal82, Mes80, Mes85, PL80, SD80, SSC83, SLT81, TSL80, VV83, WS80b]. shells [HPKW81, MC88]. Shellwise [KD80a]. shielding [BH89b, FGM87, FNC86, Kau82b, SIN80a, TT83]. shift [GC84c, GSS85b, JSS80, KB80, Kry86, PM84a, PM84b]. shifts [ESCC85, Fuk83, GRS^{+}82, HG86, MP84b, Ran83, RM84, Vog86]. Shock [KH80, OPSO86, OPSO87]. shocked [Bar86, KH86]. Shockley [TP80]. Short [FGG83, GSS85a, Nag82b]. short-chain [Nag82b]. Short-time [FGG83]. shortcomings [Mcc83]. Shubnikov [KAA85]. Si [Dat82, Deá83, LGL^{+}84, SNNG86, TH81b]. SIBFA [GPC85]. side [LJHU89, SKL80, SNSM82]. side-chain [LJHU89, SKL80]. Siegert [Sim81, Win82]. SiF [KG86a, KG86b]. sigmatropic [RBR80, SSZJ82]. sign [GK80a]. sign-alternation [GK80a]. Significance [Tat81a, FCPL84, JD83, PP80]. significant [Dav81]. silatranes [LZY83]. silica [SZ84]. silicon [ALL83, AL87, BANF86, BAN88b, CADL86, CM81, DAL89, DS87, DGSO85, FL83, GAL86, JA82, LCF89, Lás82, LTLF87, LLAG89, SRP81, Ver80, VMMV80]. silicon-carbon [DGSO85]. silver [JA82, MB89]. Silverstone [Ten88b]. silylsilylene [GZWYCHC88]. similar [CLA80a]. Similarity [SY86, BTK83, BJR86, CLA80a, CD87b, EBEE87, HR87, RR87a, RR88, VV87, Wit81]. Simon [Rös87]. Simons [Ågr87b]. Simple [Hel82, oLCC81, May86c, McW84, RGW88, RB86, TP87, Ada64, BB85, BST86, ČV88, Cox87, DA80, Duc86, DR80, Edw88, FC83, Flu84, FS86b, GHF84, Her83, HO85b, HAB83, JM83, Mod81, NSS86, Sim80, EB82a, SH89]. simplest [Lud83]. Simplification [MR81, MR80b, Tai89]. simplifications [Bal81b]. Simplified [RLR88]. Simpson [KS89a]. Simulated [Duk82, DBF89, DO84, MSWC88]. Simulation [BCV80, ACP82b, Cla80b, CFC83, DKLP81, DS89a, GD82a, LG84, PK88, RC80a, Rom81, TLT^{+}84, HČ83]. simulations [BMRJ86, CC84b, KH86, MB86, RJK83, SG86]. simultaneous [HGPF89]. SINDO1 [JIS87]. Singer [Pos83]. Single [JM83, PL80, ACCC80, BLOW89, Cul88, KS81, LT89, LC82b, MJ89, Ott86, PW86, PHGR88, PSBB83, ST80b, SSR^{+}87, SL81, UTVP85, VARPU84, Ver80]. single- [Ott86]. single-determinant [LC82b]. Single-determinantal [PL80]. single-filing [UTVP85]. single-reference [LT89, MJ89]. singles [LPB82]. singlet [AMM83, DWQ86, Has86, KSO84, Kar84a, LPS83b, MGN82, NH81, SS80c, KC82a]. singletons [Wil89]. singly [Flu85, PP89]. singular [BHM87, ERB82, Ing80, BEP87b]. Singularities [LRZ84b, Ban82b]. SiS [LC88a]. site [AWCL89, CLA83b, DOW87, GB81, KSGB^{+}83, MPP^{+}81, SG81a, SRC84, TLT^{+}84, vDTBN80]. sites [DBCJ86, GM81a, SF89, ŠKH81]. situation [Löw89]. sixth [Jor89]. Size [BCC82, BBO83, MJ89, AYCD81, KBSG81, RS88]. Size-extensivity [MJ89]. sizes [Fin85]. Slater [GKMV80, AM83a, CFLM85, CFS86, CFS87, EG84, FVZ^{+}88, Fri86, Gal82, Gol80, GSO^{+}87, Gus85, Gus87, Jon80, Jon81b, Jon81a, Jon81c, Jon82, Jon83, Jon85, Jon86, LSP85, Löw83, RFST85, RLPFA86, RLR88, SF80b, TL80, TL82, TL83, Wag83, Was89]. Slater-orbital [GKMV80]. Slater-type [Gus85, Gus87, Jon80, Jon81b, Jon81a, Jon81c, Jon82, Jon83, Jon85, Jon86, SF80b]. slit [Bel80b]. slow [Del83, ENA89]. Small [Emi83, SKIK82, Ågr87a, AA89, Bar85, Ber88, BCB85, CP85b, Cze87, DMB86, DMB89, GM81c, GSO^{+}87, HDO88, Kun83, MTN83, PGG80, POT80, SMD83, SFC86, TNT^{+}86, Vol88]. small- [Kun83]. Small-angle [SKIK82, POT80, TNT^{+}86]. Small-polaron [Emi83]. sodium [Boe88, BCB85, LCC84, UTVP85]. Soft [TLT^{+}84]. softness [NKZ88]. software [SMS84]. solar [ES88]. Solavation [Boč88a]. Solid [CD92, CLVC84, CJ89, Deá83, EHK83, ESCC85, HLPV83, Kun83, MS86, RJK83, RKHK88, RKO^{+}89, See88, Rat88]. solid-state [Kun83]. solids [BN80, Bec81, BL81b, CS80a, CLVC84, Coh86b, DVV80, Fau85, GGDG86, GGDG87, GLP83, JT85, Kau87, Kha85, KWBC80, LOBS86, Laf81, LH87, Mar86b, NB79, PGM80, RB88a, Woo82]. Soliton [KWM86]. solitonic [KAM85, MC89]. Solitons [Dav85, Dav91, KO80, Cal86c]. soluble [Cio87c, CL85]. solute [Bar81]. Solution [IRA87, MSGPP80, Ori84b, Ada64, Bot81, CP86a, CA86, CAML87, Csa80, Csa86b, DS87, DSB84, GjC86, GSS85b, GUSS89, JRK84, MMRY88, New80, PMCA88, RN84, SG86, Tac82, Tva89, WW89, WB89, YJ85, LPS83b]. solutions [BCD89, CD89a, CTV80, CT81, DL83, Edw88, Gos85, Ham82, HL81, Hil81a, HM89a, KC82b, KC83, LPS83a, LPS83b, MB86, PV89, RI88, SMMM87, Sin80b]. solvable [CF83, Has85, Lar86a]. solvation [MR88, PK88, SSC83, CCG^{+}82]. solvational [LJHU89]. solve [AN85]. solved [TKR86]. Solvent [DA88, KM80, Pal89, BMRJ86, CA85, CA86, GjC86, PCA89, RC80a, Rom81, SCER88, SGP^{+}80, VWCC82]. solvents [Ade87, BSL^{+}80, SS86a, vZ81]. Solving [Lim80, AB84, IPV81, PS80b, Tel88]. solvophobic [Sin80b]. soman [CPZS86, SBH^{+}85, SBH^{+}86]. Some [All87, ABD81, BEP87a, BEP87b, BCD89, CF81a, Cla80b, EN87, FC82a, FN83, FB83, Gos82b, GMV83, HR83, JL87, JSSS82, Kar80, Kou80, KB87a, LOF83, LRT88, LP84, Löw84, Löw86e, LFM89b, LN89a, McW83, Nag82a, NBE87, NDBE82, PNGL83, PGL^{+}86, PP81, PGPRN86, SS89b, SK81, Tho86b, WGS85, Wil80a, AABW80, AAW82a, AAW82b, AEHNEA85, AEHA87, BB82a, Ben82, BR80, BMBK82, BR81b, BCD^{+}80, BCV82, CA86, CHB88, CKNG80, CNK87, DMB86, Ela87, FC83, FHHE80, GSG84, GJ83, GRP^{+}80, GGM81, JJZ^{+}88a, Jon83, KEIG86, KRT83, KKA^{+}84, KCC80, KTGR89, KGB87, LCC80, LOB85, LOB86, Löw86b, Löw88, MB84, MTO^{+}81, MOT^{+}83, Mof81, MA80b, MP87, PL80, PXLC86, PFC88, PP82, PT89b, PBKB89, Rab84, RBSR84, RSBR85, REB83a, SMD83, SR81, SS83, SRP81]. some [SS81b, TLDK89, Tho89, TP87, TKR86, WWD88, YHG89, YRS85, Živ83, dB84, BB87, Cia81a, DPZ80, GG81, Gra84, KKSM87, Ska84, SMJ83, TV83a, Woo82]. Sons [Cal88, Mår80, Nat88]. soot [LWZ88]. sound [Lee83]. source [LJHU89, NS83]. sources [Dun88]. space [ASL88, Bal81b, BIN86, BS83b, Cal89b, DSB84, DH85, DD87a, ESS86a, GD82b, HSSW83, Ish83, IRA87, KD86, LER87, MTLL82, Mez81a, Mez82, Mez83d, Mez84d, Mez87b, Muk86, Nag82a, Pan85, Pan89, RB88c, RFST85, RI88, Roo80, RDG^{+}86, SC86b, STS82, SH87b, TV83b, Var85, BD88]. spaces [AW82, Ban88a, GB85a]. spacing [VR86]. Spatial [Mar88b, BJ87, Ban88a, DSG86, GLP80, GB81, Pau84]. spd [Lip88]. Special [Zer87, KHNT86, OMC82, RB86, ZRY83]. special-pair [OMC82]. species [AM89, IDK87, KSK86, Ler83, MS87, Nov83a, SD82, SS89a, TR86]. Specific [MR88, ES81, FPDA89, KRMK86, KTGR89, SRC84]. Specificity [Kot81, Ave84a, CPZS86, HZL84, SKR81, SBH^{+}85, SBH^{+}86]. Spectra [Bal82b, BL88, Dat82, AEHM85, Bak85, BN80, BMMS81, BH89a, Big86, BK89, Böh82, BDK86, CD87a, CKNG80, CNK87, DPZ80, DGSO85, EÖWZ84, FGS85, FK83, GOS87a, GC84f, HCCB86, HDFL84, IL83, JN88, KS81, KKA^{+}84, KM80, KPSS87, LaF89, LCCW87, LH87, Mar89, MTI80, MMI^{+}82, MKW85, Nak83, OKZ85, OBMU82, PXLC86, SA80, Sla85, SM80, SDS83, SSN^{+}84, VVD80a, VC83, WSD80, WCLZ86, Wój86a, Wój86b]. Spectral [BR85, Gas80]. spectrometer [CBC^{+}89]. Spectroscopic [AEHA87, DMB89, DPS87, SAP81, SG81b, Tay87a, BB87, Bre80]. spectroscopie [Cal81a]. Spectroscopy [Lee83, BP86, BCD^{+}80, Bri86, CD83, DR84, Don79, Fle86, Gup80, GSS85b, KTK80, KRST83, KRBM85, KNS^{+}86, KKSM87, KNSM88, KMPTK89, KKA^{+}83, KKCM85, KKC^{+}86, KKCM87, LVL^{+}85, MST^{+}80, MSGPP80, MG80, NA83, RDH84, TM83, TW88, TMC84, TMO86, Cal86b, Rat87]. spectrum [ÅCFRJ85, AYBP83, Bec82, BP86, CKK86, DM85, EV87, HHL80, HLR80, KAH80, LC82a, LHZ80, Ojh89, RDG84b, SD81, SHL82, Suh84b]. sphere [Wea82]. spheres [Blo83b, HHT82]. Spherical [Pie84, BR80, KC82c, KC82d, LE80, Pie85, SvG89]. spherically [Csa83]. sphérique [SvG89]. spheroidal [Csa86b]. spiky [Hun81]. Spin [HS80a, Ish86, Ish88b, LCCW87, MN81, Mat87b, Mat87c, Mat87d, Pau79, PS87, PM84a, RTC81, SPÖ89, VP84, Živ87b, DRS88, Duc86, EC83, FC80a, FC81a, FGC84, Gal82, Gas80, Gas81, GSG84, GOS87a, Gla85, GC84b, GC84e, Har80, Has82a, Has82b, HBC86, HLPV83, Kar86, KPP85, KPP86, KS86a, KMW80, LER87, Lar86a, Lim86, Lim87, Löw83, LW81, MZ85, McW88, Min80, MP84b, Muk84, Nes85, Pan82a, Pan82b, Pan83, Pan84, Pan85, PFBF81, PM84b, PP89, Poh83, PPP87, QX89, RSD82, Ret86, RDG^{+}86, Rut81, SC86b, SV85a, SS86c, SS86d, SS87, TSA86, Voj85, Wen86, Wol88, Wor85, KT84, KPS87b, KPS88, WP80, Ågr82, Cal82b]. spin-adaptation [PP89]. spin-adapted [LER87, Pan82a, Pan82b, Pan83, Pan84, PP89, Voj85]. spin-bonded [Wor85]. spin-density [Har80, PFBF81]. spin-dependent [GC84b, GC84e, MZ85, QX89]. Spin-extended [KT84]. spin-forbidden [Min80]. Spin-Free [MN81, Mat87b, Mat87c, Mat87d, DRS88, McW88]. Spin-Independent [Živ87b]. spin-only [Gla85]. Spin-orbit [RTC81, VP84, HBC86]. spin-paired [Rut81]. Spin-Projected [PS87, Kar86, Löw83]. spin-purity [LW81]. Spin-restricted [KPS87b, KPS88]. spin-selective [Nes85]. Spin-shift [PM84a, PM84b]. spin-spin [RDG^{+}86]. spinor [MNS85]. spiral [VU85]. splines [AC89]. split [SD80, WK83]. split-shell [SD80]. Spontaneous [Cow82, GSM83]. spontaneously [Pra83]. Springer [Ågr88, Fro81, Fro83a]. Springer-Verlag [Ågr88, Fro81]. square [GC81, Lau88]. squares [GM81b, RSST86, Tho85]. Sr [MS88, KB80]. stabilities [DF89, KPSS87, LK87, MG83, Mof81, Mof82]. Stability [KC83, Ler83, SWM82, BMRJ86, Cio87d, CH82, DM83a, DBS82, KC82b, Kar86, Ott86, RN86, SS84a, SS85, SKEDR81, Tac83, Yam82, YS80b, YS80c, vDTBN80]. Stabilization [SSYR87, DA88, Jun80, Löw85a]. Stable [CY80, Tac82, HWZ88]. stacked [LCCB81]. stacking [AN86, DS89a, Gre82, LGPP^{+}80]. stages [BST86]. Stanley [Mou89]. Stark [AL81, Rei82, Sil82]. starting [BFOP83]. State [LK81, WB81, Ågr87a, AM83a, AEM86, ACP84, AIJW82, AL87, BS81a, Bar85, Bec82, BH86, Bha80, BSR80, BBAT82, CK82, CD92, COS82, CV85, ČV86, DG85, DS80a, Deá83, DA88, DS81c, EIMS88, Ell84, Eri89, EHK83, FAC84, Fut87, GT86, GK80a, Gup80, HČP88, HLR80, Ho81, HM86, HT85, JSS80, JSSS80, KC82a, KS85, KAD81, Kat83, Kau82a, KK87, KSGB^{+}83, KS89a, KS81, KS89b, Kun83, KEGS81, KWJ83, Lar83, Lar86a, LC88a, LR89, Lip88, Löw86c, MG83, Mat87b, MS84, Mis84, NMK84, NP80, OM81, OY83, OM82c, Rat88, RBH^{+}86, RS80, RDH84, SMD83, See88, SF81a, Sie80, SM84, SM83, SG80b, Tac82, Tac83, TH81a, Tay81, Tay86a, Tay86b, Ten88b, TPN86, Toy88, WJP85, Wol88, YMO80, Yam82, YKZ^{+}89, ZNS80, LCLC88]. State [MB83]. state-selected [GT86]. state-to-state [Fut87, SM83, YKZ^{+}89]. statement [Man81]. states [AB85, ANH81, AS89, AWCL89, BPT83, BB82a, BB85, BR81a, BN84, BCA87, BL81b, Big86, BKP83, BS86, Bue86, BOW80, Cal84, CDG80a, CHKB83, CS80b, Cha80, CZM83, CM85, CHZ82, CKM84, DMB86, DMB89, DG86, De 80, DBS80, Dia86, DF83a, DWQ86, EB82b, ESKG83, ENH^{+}87, FMF80, Fau85, Fer88, Flu85, Fro83a, FBM80, Ger86, Gos80, Gos82b, Gub82, GC82b, GC83, GC84f, Har87, Has82c, HWZ88, HL83, HLPV83, HČ83, HM89b, INO88, Jac89, JP80, JN88, JWOÖ82, JJZ^{+}88a, KSO84, Kal85, KG86a, KG86b, KD80a, Kau82a, KS81, KDA81, KPH87, KWBC80, LPS83b, LC83a, LFL^{+}81, LP89, LLAG89, Löw85a, Löw86e, LÖ85b, MGN82, Mat81a, Mat84, ML81, MJW88, Mod81, Moh80b, Moh80a, MASF88, Mül81, NTM80, NH81, NBE87, NB79, NMK84, Nom80]. states [Ort89, PKH87, PBM85, PT89a, PBA88, RDG84b, RDG84a, RGDF85, San81, SNNO86, Sin86, Suc84, TS81, TNB86, TPN86, TV83a, TV83b, TV84, TV89, TP80, TOK80, Van85, VT85, Van86, VARPU84, Vol89b, WJP83, WJP85, WS80b, Živ86, Živ87c, Nat86]. Static [SS81a, AL81, BEO80, GMOBA80, SK80, Sen84]. Statics [ZH86]. Station [Pic83a]. Stationary [Eps87, Nat88, Ban88a, DD87b, FF89, GYJ83, ST89, Ská83]. stationary-phase [ST89]. Statistical [FP84, Fra84, Fra85, CDL82, CD83, CV83b, DC82a, GK80b, JRK84, LCD82, RA81, RSG^{+}87, SBNK89, VV83, Yan82]. statistics [BQS87, PS81]. status [Ber81, SSP^{+}88]. steady [RDH84]. steady-state [RDH84]. step [ÅCFRJ85, GC81, KHK87, LHKK82, LL85, San84]. steps [Bar86, VARPU84]. stepwise [PLL86]. stereo [SRC83]. stereo-irregular [SRC83]. Stereochemistry [AD80]. Stereoelectronic [PZSL83]. Steric [LPP81, FNC86, LCP81, NW82]. steroid [KMKC88]. steroids [KMPTK89, KRBM85]. Stieltjes [HDFL84]. stilbene [PD85]. Stirling [ČV83a]. STO [WK83, Zie82]. STO-3G [WK83, Zie82]. stochastic [Chr89, Cla80b, FGG83]. stochasticity [BHS85, YG87]. stopping [OS89a, SO85, Smi89]. storage [ES88]. STOS [Cof83, WJ80]. strain [Ban82b]. stranded [CLP83, GP83b, Ott86]. strange [BL81b]. Stranges [Pic83a]. strategies [KHP81, Kau87]. Strategy [Kau81, DFH84, DH85, Kau82a]. streamlines [SM87b]. strength [ACGT89, LZ87, ZLH87]. strengths [Gan80, Gan81b, GY87, GC84f, Lim82, NT80, PS84, YOJ81]. stress [Fol81]. stresses [BD81]. stretching [Kry86, MA80b, Shi87, SS89a, Wój86b]. strictly [Ing80]. Strong [Sme86, BS84, Khu82, LW84b, Rec80]. strongly [OL82]. Structural [BST86, Che82, DMB86, OMC82, SW81, BB87, Cal81a, CC82b, DMB89, EHK83, Gre87, GUSS89, HO85b, Kar80, MT85, Ran83, SA86a, Sin88, SRDG85, TL86, BT82]. Structure [ALL83, ADL78, BG81, BSHK86, Cal86a, CL83a, Dat83, D^{+}85, Gri81, KR81, KLS89, LW81, PKG86, PAFW84, PKH^{+}81, PSIT87, RP81b, SD82, SMSA80, Sti82, Toy88, TK86, TIS^{+}88, WMS81, ZY85, Ågr85, ÅCFRJ85, AABW80, AAW82b, AEHM85, AYBP83, AD80, ADS83, Ano87a, BIN86, BN84, BMMS81, BVB88, Böh84, BCD^{+}80, BDK86, BCB85, BS80c, BC84, BC85a, BC85b, Bur85, CD87a, CV82a, CZ81, Chr84, Cio87a, Csa81, CV83b, Csi82, CLZ84, CH82, DS81a, DS81b, DBS82, DS82, DFH84, DOW89, DF83b, DA81, DK81a, DK81b, EZ83, ECM89, FFF88, Fra84, Fra85, FA86, FDAC89, FD89c, GMS81, GVVDB89, Gas85, GR89, GLP80, GAL86, GJMLRGSA85, GT87b, Gra84, Gro85, GDR85, GSS85b, GBO86a, HOT87a, HOT87b, Her85, HM84, Hil85, HDTR83, Hun86, IST89]. structure [Kah84, KN80, KHNT86, KT80a, KKA80a, Ker86, KS87a, KSK86, KTGR89, KS87b, LS80, LOR84, LmJQHXZ83, LKB89, LTLF87, Lip83, LZY83, LPSK83, Loh84, MSC^{+}83b, MS88, Mat87a, Mez84a, MSGPP80, MTN83, MB80b, Nec83, NPL89b, NB79, NP80, Ori82c, OWÖ81, Pal89, PC83b, POT80, PBMR^{+}84, PP80, PC86, RB88b, Ran84, RJBR^{+}87, SP86a, SC84a, Sch88, Sey81, SS80c, SNNG86, SSUF80, SNNO86, SSK84, Ska84, SHPK80, SRL^{+}83, SL80, SRSR85, SSR86, SSRR87, SKL80, SRDG85, Tab82, TH81a, TN89, TB83, TP80, TP87, TWL82, Tru83b, TH81b, Ulm81, UTP82, Urr87, VG80, VV83, Vol88, Wei88, WBM88, XHR86, Yad83, YRS85, Zun85, dB84, dSdMB86, Bal81a, CAML87, HL80, MLC^{+}81, Fro88b]. structure-activity [GDR85, RJBR^{+}87]. structure-property [Gro85]. structure/function [Urr87]. Structures [Cal81b, Mof81, Mof82, Sar82, Vog86, BTK83, BT84, Ber88, BL81b, BMT80, Bon83, BDD81, BS86, Bur83, CBV85, CLA80a, CD89a, CKM84, DF89, DGSO85, EB86, EB82a, EB82b, EB82c, EBO83, ESS86a, FD89b, FH81, FH83b, Has82a, Has82b, HWZ88, Kia80, KT80b, NC84, Ori81b, Pin88, Ran88, SHH^{+}83a, ST82, SGP^{+}80, SS88, SMKT86, TF83a, TKY89, WS79, YZ85, YJ85, LGL^{+}84, Cal81b, Ler85]. studied [BST86, BLOW89, MPD87, Nak83]. Studies [KPM83, KJL87, Pau81, PKP87, SR81, TMC84, TMO86, WS81, WS82, AEHA87, AM83b, Aus84, BCBS81, BSCR80, BMB87, CBV85, CD87a, Cha88, CS81, Chi81, CS83c, CSB86b, CP83b, CH82, CP85b, DBS82, DAPS81, DSW80, Dew88, DR84, Ell84, FHHE80, FPDA89, Fut87, Gas80, GG83, Gro85, GDX89, GRBP87, GUSS89, HZ84, HLPV83, ITKD80, JW86, JRK84, JR85, KG86a, KG86b, KC82c, KC82d, KŠHH83, Kot82, KPSS87, LPC^{+}80, LPCF80, LP81b, LPSK83, LCP^{+}88, Luk81, LHH^{+}83, LN89b, MTWO83, MWO^{+}84, MSM81, MFM89, MGBH81, MP84a, MB85, MR80c, PMOW89, POS81, PSS^{+}87, PSK^{+}81, PGPRN86, PZPL86, PL83, PL84b, Ran84, RJBR^{+}87, RR87b, RBR80, RDH84, Rud84, ST89, SKR81, Sap80, SSSH82, SOS84a, SOS84b, SJ85, SS84a, SS85, SC80a, SP81, SP82a, SC84a, SP86b]. studies [SZ84, SY82, SRP81, SG81b, Sim82b, Ska84, Sla83a, Sla84, SRL^{+}83, SL80, SPSG85, ST88, SSN^{+}84, TNT^{+}86, TF83b, Tew87, Tew88b, Tho89, TK83, WCE89, YL83, Yad83, Yam80a, Yam81, ZNS80, BC83, Kau82b, LmJQHXZ83, SSUF80]. Study [AGB82, ADL78, BDR^{+}88, Cal82a, Cal86b, CR81, CL83b, Dun80, GGM81, HS80b, JPAB87, Kel78, KT84, Moi81, NB79, PXLC86, AABW80, Abd80a, AJ83, AN86, AS86, BBH87, BSHK86, BMRJ86, BR80, BNS^{+}87, Böh82, Böh84, BK84b, BBL81, BP83, Bou84, BVGV89, BANF86, Bue86, Cal81c, CMDBVMA89, CPN89, CV82a, CSK87, Cio87c, ČPV81, CC83, CCD^{+}84, CL85, CLL86, CJ89, DKLP81, DMB86, DPC82, DT83, DS81b, Del88, DF89, DBD^{+}85, DOW89, Dov84, DO84, DD89, EIMS87, EIKG^{+}88, ES81, EC83, FGC84, Faz87, FNC86, FS86b, FS87, GPN80, GVVDB89, GG85, GC85a, GJMLRGSA85, GjC86, GLLOB86a, GEdP81, Gre82, GP82, GP83b, GP83a, GP85, GRP^{+}80, GZWYCHC88, GM86c, HOT87a, HOT87b, HB86, HGHGF^{+}89, Hil81c, HL80, JP80, JL87, JL84, JSDK89, JA82, KVR85, Kar84a]. study [KGEP89, KG89, KL80, Kin86, KCC80, KTGR89, KK82b, KK83, LCF89, LRZ84b, LB83, LG84, LOB85, LL89, LH85, LR89, LZY83, LY83, LBV86, LSR^{+}84, Low82e, LP80, MY81, MG83, MSC^{+}83a, MBKH83, Min87, MTO^{+}81, MOT^{+}83, MK83, MP87, NL89, NPL89a, Nov83a, OLB^{+}82, ÖHD89, OG86, PCA89, PH82, PBKB89, PFLB83, Pul83, RB81, RBSR84, RSBR85, RBL84, RT87, RI88, RLH86, SK87, SMD83, SP86a, SBNK89, SRO81, SSS81, SS82, SJ86, SFC86, SJ88, SS80a, SSB83, Sch82, SAS86, SG80a, SA86b, SG81a, SJRP88, SG89, Sla85, SWM82, SHP^{+}80, ŠKH81, SF81b, SI89b, Tak86, TF82, TR86, TV83a, TV83b, TV84, TV89, TP87, TLC87, Van85, VT85, Van86, VCB83, VT82, VCV87, VŻH82, Wag83, WOT89, WH87, WJP83, WBK86, YMO80]. study [Zie82, vDTBN80, CFG^{+}83, CG80, OYN80, SvG89, TMON80, TF80]. studying [EIMS88]. Sturmian [EN87]. Sturmians [Bli84]. styryl [PD87]. sub [BW85]. sub-bands [BW85]. subduction [Wen86]. subdynamics [OFB81]. Subject [Zer87]. subshell [Ten86a]. subspaces [BB83]. substituent [BK84b, CZM83, HL80, LLM86, POK85, Sey83c]. substituents [BSL^{+}80, LH85, LY83]. substituted [BK84b, BCD^{+}80, GSG84, GJMLRGSA85, Hir80a, HMF82, KTGR89, LaF89, POK85, SSS81, SP86b, Tak86]. substitution [MTWO83]. substitutions [JD83, PHGR88]. substrate [CLL86, Con87, KK83, MB89, MBD84, SRS81, VWCC82]. substrates [GB81, MSC^{+}83a, MBD84, PL83, PL84b]. substructures [RJBR^{+}87]. Subsymposium [Ano81s]. subsystem [AL80]. subtilisin [NSP80]. subtraction [Pan82a]. Successes [Mcc83]. successive [BSD82]. Sucher [Har80, Ish85, Ish86, Ish88b, Ish89]. sudden [BNK86, Lar83, MPP83, PGG80]. sufficient [BD85, TS82]. suicide [CLL86]. Suitability [GC81, PWL89]. suitable [BG88, DQ81]. sulfide [Blo89]. sulfides [ÖHD89]. sulfonamides [KBSG81, Sey81]. sulfur [Blo89, BCV82, FA86, PDMC82, PDKC84, TL83]. sulfur-containing [BCV82, PDMC82, PDKC84]. Sum [Win86, Cio89, GB85a, GS89, Iga85, Sen80b]. Sum-rules [Win86]. summability [Sim82a]. summations [DFP80, Rei82]. sums [SP82b]. super [Kat83, Roo80]. super-CI [Roo80]. super-Hamiltonian [Kat83]. supercoiled [MT85]. supercomputers [CCC^{+}89]. Supercomputing [CCC^{+}89]. superconducting [FGMM80]. Superconductivity [Mat80, CCS84, Fle81, LSW89, Sch88]. superconductor [Lit81]. superconductors [KGO89, SS88]. superexchange [MA80a]. superflow [Con87]. superfluid [Lee83, SS80b]. superlattices [LOGC86, LVL^{+}85]. supermolecular [SHPK80]. Superoperator [WG82, CD85]. superoperators [Löw82a, CH86]. Superoxide [Var81a]. superposition [AGCT87a, LC81, MK85, PŻ85, SHKP80, SRHK83]. Supersecondary [SRDG85]. Supplement [H^{+}90, OM83b, OM83a, OM84, OM85, Pie85, OM82b, OM86]. Surface [AS89, BLG85, Boe88, BBL81, ČD87c, CSB86b, FBM80, FCPL84, Hal88, IL84, JS88, Kou80, Kun86, LP81a, Lin81b, LVL^{+}85, MSI80, MB85, MG85, NOMK80, Ori81a, PT89a, RVK80, RLH86, Sim86, SB83b, TW88, TP80, TP87, TH81b]. surfaces [Ågr87b, AMM83, AM88a, Bor81, CLVC84, Ell84, FN84, Fin85, GM80, GSC89, Hal87, JS86, KPH88, KHRK87, LB86, Lás82, LHH^{+}83, Lun89a, MM80, Mez81b, Mez83a, Mez84c, Mez84e, Mez85a, MB85, Nal80, NP81, Nom80, NC84, PT89a, PL83, PL84b, RM87, ST86c, TOK80, Umr89, Var87]. Survey [Ran83, HWZ88, Löw85b]. susceptibilities [BYBZB83, Bou80, DCF82b, Sve88]. susceptibility [SG87]. SW [BS80c, GRP^{+}80, PXLC86, RLS87]. sweet [TLT^{+}84]. switching [DS88]. Symbolic [ČV88, ČVV87]. Symmetric [DK82b, HM89a, Lin81a, Csa83, GRBP87, Kat89, LFM89b, WP80]. symmetries [ACCC80, BJ87, Ban88a, Flu85]. Symmetrization [Tay85, Lim86]. Symmetry [Bal81b, Bal82c, Bal82d, Böh83, EV87, Fin85, KS87a, Lef83, ML81, Qia83, Ran80b, RK82, RD84, SAM82, SZJS82, SSZJ82, VFA81, Voj85, AØ80b, BBH87, BDD81, Cal85, CK82, CS80b, Cow82, Cox87, Dov86, FS80, GCP87, GBW88, Gos85, GC87, HM89a, Jar89, KC82a, KPP85, KPP86, Kib83, KG85, KG86c, Kry82a, LM82a, LER87, LC80, Mat84, Mey88, MB80b, Pan83, Pra83, QX86, RB88c, SM85, Tai89, TS80, Yam82, ZRY83, Živ88, Tru83a]. symmetry- [LER87]. symmetry-adapted [CK82, CS80b, FS80, KC82a]. symmetry-broken [HM89a]. symmetry-correlation [LC80]. Symmetry-decomposed [ML81]. Symposia [Mar81a, LM81, Löw85b]. Symposium [Fax87, Han86, Kau87, Pul79, PTG80, Ano80b, Ano87u]. syn} [HB82]. Synchronism [Bur86]. Synchrotron [MST^{+}80, MGBH81]. synthesis [Blo83a, BCV81, Fox81, Mat82, Mez84e, HAB83]. synthetase [SSRR87]. system [BNK86, BVD81, BBL81, BFC80, CMDBVMA89, CPS81, CLM80, CB85, DYD89, ES88, Eri89, FF85b, GLLOB86b, GP86a, Gou86, Hos80, IHI80, JŻ85, KHNT86, LAGS89, LCA89, MS86, Moo81, Mor89, PCA89, PR83, SSP^{+}88, SM87a, The84, VP84, ZJŻ88, FF85a]. Systematic [BH86, BSHK86, Mar88c]. systematics [MP84b]. systèmes [TV83a, TV83b, TV84, TV89, Van85, VT85, Van86]. Systems [SBE^{+}83, AN85, ACP82b, Ano81s, AL86, ALH86, BR85, BB82a, BR81a, BD88, BEP87b, Ben82, BR80, BK82, BK84b, BP83, BDK86, CTP83, CDL82, CD83, CLS81, Chu88, CR81, CCG^{+}82, CS82, D^{+}85, Dav85, Dav91, DBS82, DG83, Del88, DD87b, DRS83, DPR80, Dov86, ES80, EZ83, EB82a, EB82b, FC81b, FC81c, FC81d, FC83, FC84, FC82e, FC82d, FD89b, Fra84, Fra85, FR84, FCB80, Fro88b, Frö83b, FGMM80, GSMRM85, Gás86b, GR89, GD82b, GDBRV87, Gla85, GK80a, GLLOB88, Gos82a, GC84b, Gre87, GCP82, GBK86, Has82a, Has82c, HZBR89, Hil85, HDO88, HČ83, IA87, II82, JI88, KC83, Kar80, KH86, KPP85, KPP86, KRT83, KM85, Khe81, KD86, Kle83, KCP80, KM89a, KB87a, LCC84, Lar82, LB86]. systems [Lie83, LW82, LW84a, LW84b, LN89a, MM87b, Mic83a, MB84, MS80a, MC89, MB82, MBD84, MK81, MF87, NHKY83, NS86, Obc84, OMCD81, OL80, PWL89, Pan81, PMCA88, Pip89, PD80, Pos83, PBGT80, RZ86, RHST88, RSBR85, Sad83, Sch80a, Sch81, SS84b, SD88, Sch83b, SCH84, See84, Sey81, SNNO86, SMJ83, Sri86, Swe82, Swe83, SC84b, Tab82, TNY86, TPČ83, TF83b, Tay83, Tay84, Tay86a, Tay86b, Ten88b, TN80, TV83a, TV83b, TV84, TV89, Van85, VT85, Van86, Yam80b, Živ87a, Živ87b, Živ87d, Živ88, Cal86c, Nat88, Tap88]. Szabó [Tap88]. Szasz [Bra86b]. Szent [Fox87, Lad87]. Szent-Györgyi [Fox87, Lad87].

T [Ano81d, Cal81b, Tap86, GUSS89, LSW89, Sch88, SS88]. table [Blo89, Ess82, Pur88]. tableau [LZ89]. tails [IO89]. Tallahassee [WJ82]. Tamm [TF83a]. Tamm-Dancoff [TF83a]. targets [TR80, Mår82]. TATB [HKKM83]. tautomeric [BLR83, BCD^{+}80, DKLP81, HPKW81, LWZ88]. Tautomerism [LOB86, SS81b, BKS87, CL80, KKA^{+}84, LOB85, MB82, PSS86, Sch82]. tautomerization [XYSJ81]. tautomers [KPSS87, PSS^{+}87]. taxonomy [CD87b]. Taylor [JBW87, LKB89]. TCNE [WK83]. TCNQ [ECM89, SAP81, ZY85]. TDHF [GY87, NS82b]. technique [BD88, Bur85, DMB^{+}84, DR88, EIMS88, HPK81, KB80, Kia81, Löw82c, PM81, RTM81, RTM83, RMM83, RTM86, WS81, WS82]. techniques [BEP87a, BEP87b, Bue86, Hay82, New80, RL86, RP81a, VBS85, Was89, Rat88]. technology [Cia81a, Cia81b, De 80, HM81]. tell [Cas85, ST86a]. Teller [Sin81]. Temperature [Hir80a, CP83b, CTV80, CT81, EV87, Fle81, MTI80, MB83, PGM80, SS81b, Urr88]. temperature-dependent [PGM80]. temperature-perturbed [CTV80, CT81]. temperatures [CC82b, Csa80, Csa82, GLLOB86a, TK86, ZV85, ZV86]. temps [STV80, SV85a, SV85b, SvG89]. ten [Yon85]. tenseur [SV85b, SvG89]. tensor [FC82a, FNC86, GC84d, Pan85, Pie84, Pie85, SV85b, ZS86]. tensors [BH89b, DQT80, DQ80, DQ81, FGM87, Fol81]. term [FGC84, FWS^{+}85, Lud83, VV83]. terms [AZY84, ALH86, Dun87, GS89, Ger86, Hib81, KDA81, KEB87, LER87, Nat86, NCKL81, Röm83, SF80b]. Ternary [Zun85, BMMS81]. tertiary [POT80, SRC84]. test [Chr80a, Chr80b, LSH88, PSBB83, SRL^{+}83, SP85, Tin80, Ish86, Ish88b, SST83]. Testing [NP81, YG86]. tests [Kun83, LCULC87, RST87, Sin80b]. tetra [SM85]. tetra-atomic [SM85]. tetragonal [KAA85, RK82]. tetragonal-pyramidal [RK82]. tetrahalides [LHP80]. tetrahedral [DA88, NC84, Ori85b, Yan85]. tetrahedron [BFJ81]. tetrahydroacridine [PBKB89]. tetramers [FN84, GP85, NC84]. tetramethylammonium [GP83b]. tetramethyls [LHP80]. tetramethylsilane [YUKK84]. tetrapeptides [RTM83]. tetraphenyldithiapyranylidene [ECM89]. Tetrathiafulvalene [TL82]. tetrathiofulvalenes [POS81]. tetryl [HKKM83]. Texas [Pic83a]. th [SV82]. Thalidomide [KTK80]. Their [ML81, AEHNEA85, Ada64, AØC82, Bal82d, BN84, BR81b, BSL^{+}80, BC80, Cal86b, DPZ80, Don79, DDFA87, EGP86, Fli86, HV83, HG86, KC83, KTK80, KRST83, KRBM85, LOB85, LOB86, MTI80, Mat84, Mof81, MP87, NKZ88, New80, PD85, Sar89, SF81a, Sil80, Sin89, ST86c, VMMV80, WBK86, WHT86, XYSJ81, ZP82, Zie82, Živ87a, EC83]. theorem [AEM86, Coo87, FC82a, FC82c, GN87, Kar85, Kry80, LJC83, LP82, MM82, MD87, May86c, MPS86, Nal80, Sch80b, Wea84, Wea88]. theorems [FC83, Niu84b, Nov83b, Sil84]. theoretic [Bon83, Eng83a, Eng83b, KSSH85, Kle86, KS89a, Ler85, Mar88c, RFST85, STS83]. Theoretical [Bar86, Boe88, Cal86a, CCG^{+}82, CP85b, DSW80, Gas86a, Gla85, GjC86, GLLOB86a, GP82, GP83b, GP83a, GZWYCHC88, GDX89, HZ84, HL80, Hos80, HJ80, JŻ85, JA82, KCC80, LCC84, LSW89, LPC^{+}80, LLS83, LL89, LY83, LCA89, Low82d, Low82e, LGL^{+}84, LHH^{+}83, MTWO83, MWO^{+}84, MW83, MKW85, MTO^{+}81, MOT^{+}83, MK83, MV87, NS86, PCA89, PZ84, PMOW89, PFC88, PGM80, PBBK89, PBKB89, RDG84b, RDG84a, RGDF85, SO85, SS83, SS84a, SS85, Ser84, SJRP88, Sok81, Tew87, Tew88b, TB83, Tho89, TK83, VŻH82, WCE89, Wój86b, YMO80, Yam81, ZA89, ZJŻ88, BBH87, BB82a, BN80, BH82, Böh83, Böh85, Bou84, CPN89, CS87, CCS84, DT83, DOW89, DD89, EB82a, EB82b, EB82c, ECM89, FGC84, FZZ87, Fle81, GEdP81, Gre82, GP85, Gro85, GDR85, GS80, HGHGF^{+}89]. theoretical [HP81, Kau81, KT80b, KST84, KPSS87, Lad87, LCF89, LOB85, LH85, LR89, LL82b, LL85, LLM86, Mar88c, Mas87a, Mat82, Mez85a, NB79, PWL89, PSS^{+}87, PGPRN86, RW80a, Ran81, RZ86, RSNT88, RLH86, SRO81, SK80, SVS81, SCER88, SKER83, Sim86, Sla85, Suh83, TNT^{+}86, TR86, Wój86a, Woo82, KC82c, SHP^{+}80, TLT^{+}84, Tap88]. théorie [SV85a, SV85b, SvG89]. theories [FS86a, JP80, Khu82, Kle87, Kut80, MJ89, MM81, NH81, dB84, NHKY83]. théorique [Cal81a]. theorists [Kun83, Woo80, Cal82a]. Theory [AHB^{+}80, Bra86b, Cal81c, CM81, Csa82, DR84, Fri87, Gir86, IDK87, KWBC80, LM81, LH87, MB83, NKM81, OT80, Pac82, RDH84, ZV85, ZV86, Ågr88, Ada82, AČP82a, ACP84, Ade87, AT85, ADL78, Ano80c, Ano82s, ACCCBF80, Au82, Aus84, AØM81, BPT88, BS80b, BW80, Bel80a, Ben82, BK84a, Bha81a, Bha82a, BP86, BK89, Bon81, BT82, BS81b, Bor85, BERE85, Bre80, BC84, BC85a, BC85b, Bue86, Cal86b, CS80a, CS80b, CS83a, CDL82, CD83, CP85a, CMM83, Cho80, CGK82, Chr80a, Chr80b, Chr89, Cio87d, Cio88a, ČV82b, CV85, ČD87c, Cla86, CA85, CA86, Csa85, Csa87, CZ82a, CZ82b, CZ88, Cul89, DPZ80, Del83, DP81, DS88, Don79, DCF82b, DWQ86, DC82b, DBC84, EI81, ERB82, EKBE86, FC84, FF84, FEKC84]. theory [FK80, Fuk81b, FGMM80, Gás83, GV88, GGDG86, GGDG87, Gol80, GB85b, GBW88, GBW89, Gol84, GK80b, Gra82, Gra84, GP86b, Gre81a, HF86, Har87, Har83a, HB86, HC80, HHT82, Her83, Hub80, HČ83, Hun80, IS85, IPV81, Ish83, Ish84a, Jac80, Kal85, KCK83, KK86, Kel86, KL87, KS84, Kha85, Kia81, Kim81, Kin86, KT80b, KGO89, KPS87a, KPS87b, KPS88, Kub82, KB88, Kum87, Küm83, LM82a, LB83, LP82, Lev89, LC88a, Löw81a, Löw82b, LCG82a, Löw82c, Mak80, MTLL82, Mal82, MMM81, MS88, Mat81a, MN81, Mat87c, Mat87d, Mat81b, May86b, MS89, McD81b, McW83, McW88, MBT80, Mel87, Mez81b, Mez85b, Mez85c, Mez87a, Mic87, Mor83, Muk86, MP84b, Muk84, MPD87, Nak83, NDBE82, NMK84, OFB81, Obc84, OS89a, OPSO86, OPSO87, Oga80, OJY83]. theory [Ort87, Ort88, PP84, PL89, PT84, Per85, Per86, PP81, PSK86, PW86, Pic83a, PM84b, Pie84, Pie85, PS84, PHGR88, Pul79, Pyy86, Pyy93, Pyy00, RS84, Rec80, REB83a, REB83b, RSMM86, Sch83b, SD81, ST83, SB84, SI89a, Sel88, Sew86, SF81a, SL86, Sim82a, SRT81, Sin80b, Ska84, SV85a, SV85b, SvG89, SS87, Str82, Suc84, Suh84b, SHLL83, SH87b, Sza85, TNY86, Tac88, Tac89, TPČ83, TP84, TR82, Tay83, Tay84, Tay86a, Tay86b, Ten86a, TKR86, Tru83b, VVD80b, Var81a, Var81b, Var85, VCG86, Wea83, Wes81, Wes82, Wil80b, WS81, WS82, WJP83, WJP85, Win85, Wu82, ZNS80, Zhe83, Zho88, Živ86, Živ87c, ZAB85, CS80a, CDG80a, DDP81a, DBES80, Hib81, KWJ83, MRJ85a, Ågr85]. therapeutic [SS80f]. there [Ada82, KFD84]. Thermal [GT86, Hil81a, Hil81b, Hil81c, MB86, Chu88, FHH^{+}87, KKTE80]. Thermodynamic [Blo83a, Bro88, CA86, GZWYCHC88, MBT80, PV89, SZR80, SKER83, SK81]. thermodynamical [Löw88]. thermodynamics [Pan81]. therory [Cul88]. these [TS82]. thin [WD81]. thio [LOB86]. thio-derivatives [LOB86]. thiocarbonyl [DMB89]. thiocarbonyls [Fin87]. thione [FZZ87]. Thiophilic [PH87]. thioxanthenones [MLA80]. Third [MB80a, SF81a, Blo89, MMC83, Sve88]. Third-harmonic [MB80a]. Third-order [SF81a, MMC83, Sve88]. Thirring [Csa86a]. thirteenth [PTG80]. Thirunamachandran [Tap86]. Thomas [CT81, BPT88, CTV80, Csa80, Csa81, DCF82b, Ess82, Lud83, Ols86, Ten86a, Ten86b]. Three [BVB88, NCKL81, Nov83a, AEHNEA85, Ble82, BCT86, CC83, CLZ84, FGS85, Fla80, GTB80, GD82b, HBSC83, HS80b, HY89, Jon83, Jon84, MK85, MBD84, PS80b, Pos83, PD89, RB88a, Ran88, SRDG85, Wea88, LCULC87]. Three-and [LCULC87]. Three-body [NCKL81, Nov83a, Ble82, CC83, Fla80, HBSC83, PS80b]. three-center [CLZ84, Jon83, Jon84, Wea88]. Three-dimensional [BVB88, GD82b, MBD84, PD89, RB88a, Ran88, SRDG85]. three-electron [HS80b]. three-particle [Pos83]. thresholds [BN84]. thromboxane [WWD88]. through-space [RDG^{+}86, SC86b]. throw [KPTG88]. throw-over [KPTG88]. thymine [MGN82, RBSR84, Zie82]. thyrotropin [RTM86]. thyrotropin-releasing [RTM86]. Ti [Dat82]. TiBe [Enz80]. TiC [BS83a]. tight [Böh84, Mod81, Sto82a, Sto82b, Zho88]. tight-binding [Böh84, Mod81, Sto82a, Sto82b, Zho88]. Time [Jac80, Lar83, PZPL86, Sad89, Tru83a, VM81, AM84, AJY80, BP86, Bot81, CD83, CD85, Dal80, DÖ89, FGG83, GM86a, McW83, MS84, Mey88, Mic81a, Mic83a, Mic85, Mic86, MB83, SV85a, SV85b, SvG89, SM87a, Tac82, VM83, WB81, SM83]. Time-correlation [VM81, Mic81a, Mic83a, Mic85, Mic86, VM83]. Time-dependence [Sad89]. Time-dependent [Jac80, Lar83, Tru83a, AJY80, BP86, Bot81, DÖ89, GM86a, McW83, SV85a, SV85b, SvG89, Tac82]. Time-resolved [PZPL86]. times [GLB81, YKZ^{+}89]. tin [BC84, BC85a, BC85b, RLS87, BS83a]. TiO [BS83a]. Titanium [BAN88b, Hil85]. Titchmarsh [EKBE86, REB83a, REB83b]. TNT [DT83]. tocopherol [SPSG85]. TOM [CPMP86]. Tometsko [Mår82]. tomorrow [KST83]. tool [Mez85b]. Topological [BMT80, Sey83a, Sey83b, Sey83c, Sin88, PNA86, PNA88]. topology [Cal84, Hil81b, Kin86, Mez81a, Mez82, Mez83c]. torsional [JRK84, LM82a]. Total [CLVC84, Cio88b, Csa85, Csa87, Bag83b, BZA84, CB88, Csa89, DCF82a, Fra85, Mar85b, MC88, Mez86b, MW87, NMK84, NC87, YKZ^{+}89, Fra84, PDMC82]. totally [HOT87a, HOT87b]. TOX [KKH^{+}83]. TOX-MATCH [KKH^{+}83]. toxicity [BNS^{+}87, HLKK82, Kau81, LHKK82, LZY83]. toxicology [Kau81, KKH^{+}83]. toxicophores [KKH^{+}83]. toyocamycin [SC80a]. trace [Löw81a, Löw82b]. track [CJ89]. tracts [GUSS89]. trailing [KO80]. trajectories [Avr82, DBF89, Dun87]. trajectory [ES81, SM87b, ZJŻ88]. trans [GBO86a, EIMS87, KL80, KKA80b, LPP84, RB81]. trans}-dichloroethylen [LPP84]. trans}-dichlorooxirane [LPP84]. trans}-polyacetylene [KKA80b]. transaminase [AIJW82]. transcription [BST86, KMKC88, SKR81, SR81]. transduction [Moo81, Urr88]. Transfer [SCER88, Ågr83, Ave84c, BLG85, BQS87, BCD89, BL89b, CL80, COS82, CC82a, DÖL87, DOW87, DOW89, ECM89, FHHHE85, Fla80, Fut87, GSM83, GPLP85, GCP82, GCP86, HR84, HKV85, JN88, KH86, KBT80, KS86c, LS80, Lar82, LBKV88, LS86, LT81, LCA89, MGN82, MDO^{+}82, Mic81a, Mic83a, MR87, MR88, MC89, MBD84, MPP^{+}81, NSP80, NW82, OL80, Par84, PMCA88, PK88, RB89, RJ87, Sch80a, Sch83a, SD88, SWW89, SHRH82, SAB^{+}80c, Suh84b, SRC84, TNY86, Vas82, VM83, YZ85, ZV85, ZV86]. Transferability [RACPL86, HAmGS86]. Transferable [AØ80a, AØ80b, AØM81, AØC82]. transference [MB82]. transfers [KJL87, Sch81, SSB83]. Transform [GSO^{+}87, DFP80, GKMV80, PS80b]. Transformation [Gal82, BDR^{+}88, Bha82b, BVAB81, DM85, FPT81, HY89, Iga85, KN84, Kim81, SY86, SH87b, TS80, Ten88a, Wea83, Wes81, Wes82, TV83b]. transformations [BF83b, FC82a, KPS87a, KPS87b, KPS88, LWÖ81, Mal86, Niu84a, OY83, PSK86, Ste81, SHLL83, Wit81]. transformed [GMP^{+}84a]. transforms [MPP83, Niu84a, Niu84b]. Transition [Cal84, Dat82, KS89b, YG87, AJY80, AIJW82, Bal88, BSHK86, Bau86, BW85, Bha82c, BBPS83, BP80, Böh82, Böh83, BBAT82, BANF86, CKM81, CLM80, CMM83, DS81a, De 80, DK82a, DA88, Ell84, EGP86, Faz87, GM86a, GOLC87, GC84f, Has85, HWZ88, Hil85, HG86, Hos80, Jac89, JN88, KN80, Kau82a, KMW80, KLDD84, LFM89a, LNN84, Mal86, ML81, MF87, Nas83, Nec83, Nes85, OM81, PC88, PT89a, Sad89, SF89, SMJ83, Suc81, TPN86, TP80, UT80, Urr88, WS80b]. transition-metal [BSHK86, BP80, CKM81, Ell84, Hil85, KN80, TP80, WS80b]. transition-state [AIJW82, BBAT82, OM81]. Transitions [Mou89, ACP82b, DC82b, FMF80, GLP83, II82, MR80a, MH80a, MT85, Mcc83, Min80, SAM82, SMJ83, Sta71, Sta87]. translated [Muj86]. Translation [Pia83]. transmembrane [UTP82, UTVP85]. Transmission [EC83, FGC84, TMO86]. Transport [PD89, BP83, DP87, IMTG80, Lar86b, MP84a, Pan81, WD81]. transversion [KBR83]. transversion-type [KBR83]. trap [Tat81a]. trapped [IST89, MDT85]. trapping [ST86b]. Treatment [ANH81, DM85, Lad83, MH82, AEHH81, AEHM85, AEHNEA85, AČP82a, AT85, Ave84c, BPT88, CPS81, Cha82, Dov86, ESS86a, GSO^{+}87, HBC86, HZBR89, HAB83, JBW87, KRT83, KD86, LOBS86, MW87, MÖ80, MBD84, MBR80, PABL88, QX89, RZWL80, SM82, Tay81, Ulm83, Ulm85, YKZ^{+}89, ZHB^{+}87, dSdMB86, BNK86, BF81]. trees [Bal82b, KT86]. Tri [RTM83]. Tri- [RTM83]. triad [SKEDR81]. trial [OYH^{+}80, RBL84, Tay81]. triatomic [BR81a, DS82, DSW80, DGSO85, GVVDB89, LCF89, Nal80, ZHB^{+}87]. triazenes [SJH89]. trichlorine [SP86a]. tricine [SK81]. triclinic [KAA85]. trifluoromethyl [CNK87]. trifluoromethylamine [HM85a]. trihalides [YZ85]. trihydrate [SA86a]. trimer [FFF88, HKV85]. trimers [CKLR80, HBSC83, NCKL81]. Trinajstić [Brä86a]. trinuclear [MK80]. triol [PK88]. trioxide [CPGP83]. triple [CLP83, Cul89, PHGR88]. triplet [AMM83, GK80a, KSO84, Kar84a, MGN82, NH81]. Trivedi [Ågr82]. Trivial [FC82b]. tRNA [KG80, LCP81, Shi85, SSRR87, Yon85]. troponin [SRDG85]. troponin-C [SRDG85]. Troxidone [JW86]. TRP [LCA89]. truncated [LKB89]. trypsin [LSS83]. tryptamines [GJMLRGSA85, Yam81]. tryptophan [MGBH81]. tryptophan-containing [MGBH81]. Tsetlin [LP89]. Ts’o [Mar81a]. TTF [TL82, TL82]. tumor [Lew88, Tho86b]. tumor-inhibitory [Lew88]. tungsten [CPGP83, Kun86]. tunneling [Har82, Kar89, NOMK80, SB88]. Twelfth [Pul79]. Twenty [Löw85b]. Twenty-five [Löw85b]. Two [AM89, HG86, KS82, LPS83a, LPS83b, LSP85, Lan84, LE80, Umr89, ACP84, AM88b, BLG85, BN84, BSL81, BSD82, BP86, CLA80a, CC83, Cof83, FF85a, FF85b, FW80b, FH81, GHGP89, GM86b, GBZ86, Gus87, HS89, Iga85, II82, JSS80, Jon81b, Jon82, Ker81, KDA81, KBT80, LCC80, LCULC87, LC83b, LAGS89, Mez85a, MS80a, MBSC81, MBD84, MPB84, Nag82a, OY83, PSM87, PS88, Pie84, Pie85, PŻ85, RB88a, RB88b, RLPFA86, SNNO86, ST86b, Sri86, TS80, Ten88b, WJP85, RB86, SHL86]. two- [CC83, Ker81, MBD84, RB88a]. Two-body [KS82, FH81]. two-by-two [Iga85]. two-center [Cof83, Gus87, Jon81b, Jon82, MPB84, PSM87, RLPFA86]. Two-Dimensional [LPS83b, LPS83a, LSP85, AM88b, FF85a, FF85b, II82, KBT80, PS88, RB88b, Sri86, RB86]. two-dimensionally [MS80a]. two-electron [BLG85, GHGP89, GM86b, HS89, JSS80, LCC80, LCULC87, LC83b, OY83, PŻ85, TS80, Ten88b]. two-particle [FW80b]. two-photon [BP86]. Two-rotor [SHL86]. two-state [WJP85]. tylophorinidine [GDM^{+}81]. Type [GOLC87, AČP82a, Bha80, CGT86b, DF83a, Gus85, Gus87, IS88, JSSS80, Jon80, Jon81b, Jon81a, Jon81c, Jon82, Jon83, Jon85, Jon86, KT84, KBR83, Kut80, Kve82, PS80b, STH80, SC86a, SC87, Sch89, Sin81, SF80b, TH80, YZ85, Cla83a, SCH84]. types [Bal81b]. typical [VCB83, Yon85]. tyr [SSRR87]. tyr-tRNA [SSRR87].

U [Wor81, FN83, GC84d, Gou86, MF87]. ubiquitous [FS86a]. UHF [BVAB81, Coo87, HČ83, KT84, Yam80a]. UHF-natural [Yam80a]. UHF-type [KT84]. ultimate [LPCF80]. ultrafast [Fle86, Rat87]. Ultraviolet [LaF89, SM83]. uncatalyzed [OLB^{+}82]. uncorrelated [Gos82b]. uncoupler [YKTA80]. underlying [PL84a]. underpotentially [MB89]. Understanding [Ada64, Ada69, Ada80a, HLKK82, Kar89, LHKK82, Lun82]. Unified [Gus87, Sil82, dSdMB86, Cio87d, Fox86, Oga80, Rec80, Ulm83, Ulm85, Win85, AT85]. Uniform [DPC82, PDMC80, PDC80, PDC81a, PDC81b, PDMC82]. unifying [Del81a]. unimolecular [BBLR84, Chr80a, Chr80b]. unique [SV85a]. unit [DAPS81, SY82, SW81]. unitarily [ACH83, SLT84a, SLT84b, ZS86, CA80]. Unitary [GC84c, GC84d, GC84e, GC85b, GP86a, Gou86, KS84, Ste81, SH87b, BS81b, CH86, Duc85, FS80, GCP87, GC84b, HY89, KS82, KS86a, LZ89, LP89, LSBS81, Mat87a, Mat87d, MP84b, Pay82, PM84b, SS80d, SHLL83, Wes81, WP80, Zhe83]. unitary-group [Mat87d]. units [RB88c]. universal [GBK86, LP85]. Universality [FM80]. universe [Wal84]. universitaires [Cal81a]. University [Kum87, Mou89, Pic83a, WJ82]. unknown [MMRY88]. unoccupied [TMO86]. unprotonated [BKP83]. Unrestricted [Fuk81b, Has82c, Ort89, RBL84, SS87, Yam80a, PCG83]. Unrestricted-Hartree [Has82c]. Unsaturated [KMPTK89, MG83]. unsolved [TKR86]. unstable [Gir80, Sin89]. Unsymmetrical [FC81d]. Unusual [RWG83, BLR81]. Unusually [NS83]. updated [FNPS82]. upon [AEM86, BP84, FCPL84, KHRK87, KS86c, NH84, SD88, Ten86b]. upper [BH86]. Uppsala [Muj86]. uracil [BSL^{+}80, HCCB86, LOB86, RBSR84, SS84a, SZR80, Zie82, MB82]. uracils [SSN^{+}84]. Urban [Fro81]. urcail [CHB88]. urea [VB85]. ureas [DA80]. uridines [SP86b]. USA [WJ82]. Usage [TV83a, TV83b, TV84, TV89, Van85, VT85, Van86]. Use [Ave84c, Cas88, Dew88, Har80, LJC83, LKB89, MR80a, McD81b, Mül81, Nal80, PFBF81, Pau79, RSMM86, Win83, AN85, BNS^{+}87, BPCT83, ČPV81, CS82, DMB^{+}84, DFP80, DAPS81, FNC86, Ish85, Ish89, IQ87, IS88, JSSS82, Lib84, MP84b, Muk84, NKZ88, RB88c, ST86c, TL83, GC82a, TV83a, TV83b, TV84, TV89, Van85, VT85, Van86, Cal82b]. used [BAG83a, Kun83, STS82]. useful [Bha82b, LE80, SMVV86b, SMVV86a]. Using [PC86, AØM81, AW83, BPLB87, BYBZB83, BH89b, CP85a, CB88, CHB88, CPZS86, DH84, DBES80, EBO83, EC83, GOS87a, Gre87, GSS85b, HPK81, Iga85, IAAA89, IST89, JMD88, Kah84, KCK83, KC82c, KC82d, Kau82a, KKH^{+}83, KS86b, KM89b, Lar86b, LLS83, LWK84, Luk81, LN89b, Mcd81a, MKW85, MGBH81, MPP83, MMC83, NKM81, NMK84, Noa80a, Noa80b, OL82, PG80, RP81a, RSSS81, RTM81, RTM83, RMM83, RTM86, RP89, RSGK88, SIJ85, SB87, SS80d, SBE^{+}83, SNNG86, SS86b, SRL^{+}83, SS87, SC80b, TN89, Tay81, TPN86, TB81, WJ80, Wea84, Wen86, Yon85, YS80a]. utility [ML81]. utilization [CF81a, KC83]. utilizing [HAmGS86, KK87]. UV [BCD^{+}80, Bre80, SG81b, Suh84b, Suh86].

V [HDFL84, ES81, SMD83, SBE87, YKZ^{+}89, AØ80b, Boč89, CS81, DPC82, Faz87, FHHE80, KG86d, LCP81, LR89, MLA80, PSIT87, RTM86, Sch81, SvG89, TP84, TK83, VT85, Win83, Yad83]. vacancy [CS80a, IST89, TH81b]. vacuum [Toy88]. Valence [Dur86, MS87, Bak85, Dia86, EB82a, EB82b, EB82c, GMM84, Her83, Hib81, HO85a, HO85b, Lip88, May84, May85, McW88, Mül82, NP80, OKZ85, PC88, PLL86, Pic83a, RS84, SD80, SF81a, SRT81, SWM82, SSF^{+}84, Str82, SBWW80, TS81, VI84, LW84b]. Valence-bond [MS87, Her83, RS84, LW84b]. valence-shell [Bak85]. valences [May86b, May86a]. valency [Pra83]. Validity [FWS^{+}85, BW80, HYS89, LBP84, MM80]. valinomycin [GEdP81]. value [GBZ86]. values [FN83, Küm83, PO85, SKER83, XYSJ81]. vanishing [Lim82]. vapor [DSG86, Hir80a]. vapor-phase [Hir80a]. variable [Böh82, Böh83, Cio88b, SM80]. variables [KMW80, Tru83a]. variance [Tho85, Tho86a]. variant [ADF^{+}88, DMB89, Rér89]. Variants [Bha81a]. Variation [Dal80, Ver80, CL83b, Kag83, LN89a, MRJ85a, MRJ85b, SHH^{+}83a]. Variational [ARZ^{+}89, Bha80, CTV80, Csa80, CT81, Csa83, FL83, Fuk81a, Ham82, HT82, JSSS80, KFD85, KSGB^{+}83, Mal82, Mic83b, PKH87, Pul83, Sel88, Sil80, Tay81, BG88, Bue86, Csa86b, Csa89, CZ88, FAC84, FD86, HBC86, Jac80, JSSS82, Kut84, LFL^{+}81, LTLF87, LLC83, MK81, RSST86, SA80, SG87, Sve88, Ten80a, Wea82, WOT89, YKZ^{+}89, ddF86, HK86, Mat81b]. Variational-Perturbation [Mal82]. variational-X [SA80]. Variationally [GM86a]. variations [RW80a]. various [BG81, BS86, CLP82, GC85a, KNM83, Nov83a, PČP82, STS82, STS83]. VDW [GVVDB89]. vecteurs [Van85]. vector [AØ80a, HL88, LRT88]. vector-pairing [AØ80a]. vectors [Van85]. velocities [BWEW89]. velocity [FWS^{+}85, FS86a, GD82a, HYS89]. Verlag [Ågr88, Fro81, Fro83a]. version [Boč88b, MDT86, MDT87, Win82]. versions [SH87b]. versus [JM83, MBKH83, RBL84, Sla83b, UTP82]. Vertical [UI80, AYBP83, AJY80, PZ84, WSD80, LJC83]. very [DR80, HO85b, Ran80a]. VI [CH86, Faz87, KT84, RD84, SP82a, TP85, Van86]. via [Au82, BD88, BF83b, Cio87c, ČV86, CL85, Csa85, Csa86a, Csa87, Fer88, HG86, JJČ87, KN84, KS89a, Obc87, Rec80, SC86a, SC87]. vibrating [Hub85]. vibration [ARZ^{+}89, BK89, IL83, KK82a, Kob86, KCP80, Moh80b, Moh80a, SK87, ST86c, Wój86b, vZ81]. vibration-rotation [ST86c]. vibration-rotational [ARZ^{+}89]. Vibrational [Cha88, CHB88, CKM84, DBVMB85, ESCC85, RZP84, SSN^{+}84, VVD80a, Wea83, ARZ^{+}89, Blo89, DVV80, HCCB86, HS88, KL89, MHAS86, MB85, OBMU82, PABL88, PSK^{+}81, SSR^{+}87, SZR80, Shi87, YUKK84, ZV85, ZHB^{+}87, PGM80, TNY86]. Vibrationally [GLB81, CJ89, DBVMB85, KH85, ZJŻ88]. vibrationally-excited [KH85]. vibrations [Kry86]. Vibronic [Fis84, GC82b, GC83, GC84f, TIYH86, CKD81, Mar89, MDT85, YA87, Cal86d]. view [REB83b]. Viewing [HRL88]. VII [AØM81, LCDP80, MB83, NP80, TV89, TIS^{+}88]. VIII [DO84, KTK80, PGM80, SP86b]. villain [Fra89]. vinyl [LPP84, LP84, LBV86, MP87, ÖHD89, PP82, SP82a]. vinylguanidinium [SSSH82]. violating [Hub80]. Violations [ZRY83]. Virial [FC82c, LM82b, DC82a, FC82a, KD80a, MM82, Nal80, OFB81, RVK80]. Virtual [MY81, BP86, CP83a, HZBR89, Hir80b, Kau82a]. viscous [vZ81]. visible [SM82, SM83]. visual [San84]. vitamin [GP86b, PG89]. vitro [BLT^{+}84, BLY^{+}86, PZPL86]. Vleck [Wes82, Wil81]. voids [RKO^{+}89]. Vol [Ågr88, Cal86a, Cal89a, Fro88a, LM81, Mår82, Mar81a, Pic83b, Zer87, Cal86b]. Volume [Cal81c, Cal81b]. Volumes [Zer87]. VRDDO [HKP81, HKK^{+}82, LHK81, THKM81, HPK81, HLKK82, HKKM83, HKLM85, SHP^{+}80, SRL^{+}83]. vs [LJHU89, PMT84, RFST85, RS82]. VT [RB89].

W [Ågr82, Cal88, Lun82, Mår82, Sie81, Sme84]. Waals [AM88a, BTC^{+}83, DBVMB85, HZ83, HFSB86, Ish83, Ish84a, KCC80, MTLL82, MBSC81, SBB87, TNY86]. wake [Car82]. walk [CS83a]. walks [RWG83]. wall [KH85]. Waller [Ano89-29, Ano89-28, Fra89, LJ82, Lim87, LGJLCB89, Lun89b, Pau84, Pau89b]. walls [Fis82]. Walter [Sie83a]. Wang [Csa88]. Wannier [TP85]. Warped [Ori85b]. Water [BLR83, LBV86, BMRJ86, Bha82c, CKLR80, CC83, Con87, CG80, CC84b, CP83b, CF81b, DKLP81, DS81a, DSG86, FM86, GG82b, HBSC83, HB86, JL87, KHNT86, KNM83, KJL87, Kry86, LGPP^{+}80, Lip83, Loh87, MK85, NSL88, NCH^{+}89, Nov83a, OL80, RBSR84, RC80a, Rom81, SA86a, SHKM80, TNT^{+}86, YJ85, MR85]. water-alcohol [CF81b]. Water-Chain [LBV86]. water-membrane [Con87]. Watson [SS85]. wave [ABCV86, Ave86, ALH86, BS83b, CS89, CG80, CP80a, DS80a, DG86, DM83b, Dea89, DD87a, ESKG83, Eng83a, Eng83b, FC82c, FC84, Fra88, FH81, FH83b, GM86a, Gos82a, Gre81b, GA82, HKP81, Has82a, Has82b, Has82c, Hel82, Hib81, HČ83, Hun80, Hun81, IO89, JD83, JD85, Kah84, KK82a, KT88, KPS87a, LPW81, Löw86d, Luk82, LLC83, MM82, Mar88c, Mat81b, May86a, MM87b, MÖ80, MT88, Moh80b, Moh80a, NH81, Nak83, PB82, Pan82a, Pan82b, Pan83, Pan84, Pau84, Pos83, RDH84, Sad89, SPÖ89, SSR^{+}87, She87, SST84, SP85, SM87b, Tac88, TB81]. wave-function [Löw86d]. wave-functions [Gos82a]. wave-packet [ABCV86]. wavefunction [AC89, HGPF89, HO85b, KEGS81, LC82b, Mcd81a, OWÖ81, SS80d, Tay86a, Tay86b]. wavefunctions [Bri86, Cio89, Ess84, HC81, KE82, MSF86, SGEK81, Umr89]. wavelength [GC84f]. wavepacket [Har83a]. waves [CC84a, OPSO86, OPSO87]. waves-induced [OPSO86, OPSO87]. way [GB85b]. Weak [Enz80, NTM80, ZH86, KG86d, Pan89]. Weak-field [KG86d]. weakly [GDBRV87]. Weatherford [Sme84]. weighted [BM86, BNS^{+}87, Kin89]. weighting [RB88c]. weights [HO85b, Ori82b, VR86]. Weizsacker [Lud83, CV83b]. Welcome [Fax86]. well [CHKB83, CE86, CE87, CE88, Gra82, JD83, WB89]. well-correlated [JD83]. wells [DP82, GOLC87, HL81, LOGC86]. Weyl [ERB82, EKBE86, REB83a, REB83b]. Wheeler [MDT86, MDT87]. where [GBZ86, TWW81]. which [Eng83a, Eng83b, Hal87, MM82]. while [SR83]. Whitehead [LWK84]. widths [FF89, NMK84, ST84]. Wietling [Cal81b]. Wigner [GP86a, LM89]. Wiley [Bra86b, Brä86a, Cal88, Cal89a, Ler85, Mår80, Nat86, Nat88, Wol81]. Wiley-Interscience [Cal89a, Wol81]. Wilson [Ano80e, Ten88b]. wise [ITKD80]. within [BL81c, Cul88, Cul89, DP81, EC83, GC83, KS81, LW82, LW84a, LW84b, MH82, NSS86, PNA86, PNA88, RC80b, SHPK80]. without [CV85, Dem86, FC84, HR84, Kut84, LM89, Lim80, SB83a]. Woodward [CD89c]. Woolley [Cal82a]. work [Chu88, Gra84]. Workshop [Ano82s]. wreath [Bal82d]. wrong [Lau88]. Wu [Avr82].

X [BF81, BS80c, CPGP83, Dat82, GVVDB89, GBO86b, HL80, JR85, OM81, PXLC86, RSGK88, RLS87, SRP81, SJH89, TH81b, Bha80, KB86, KPS87a, KTS86, BMMS81, Blo83b, CY80, CFG^{+}83, DG86, DC86, GN87, GM83, GLP83, IST89, JTST87, KT80a, LHC80, LLS83, Nag87, Niu80, PXLC86, POT80, SA80, TMO86, TST89, GVVDB89]. X-phenyl [SJH89]. x-ray [KB86, KPS87a, CFG^{+}83, GM83, POT80]. x-rays [KTS86, JTST87, TST89]. xanthine [GLP80]. Xe [ESCC85]. XeF [BR82]. xenon [KKCM85, KKCM87]. XI [KRBM85, KSGB^{+}83, KS81, SHP^{+}80, TS81]. xii [Cal86b, HPK81]. XIII [Mis84]. XIV [LHK81]. XLII [Fli86]. XPS [SA80]. XPY [PXLC86]. XVI [KNSM88]. XXII [MN81]. XXIII [Mat87b]. XXIV [Mat87c]. XXV [Mat87d]. XXXIV [Fli84]. xylylenes [Kar84a].

Yang [Ban82a]. YBa [SS88]. years [Coh83, Löw85b, LGJLCB89, Lun89b]. yields [IDK87]. York [Bra86b, Cal82a, Cal82b, Cal86b, Cal89a, Ler85, Mår80, Mår82, Nat88, Pic83b, Rat87, Sie81, Sie83a, Wol81]. ytterbium [XHR86]. Yukawa [GRS^{+}82].

Z [PM86, VC83]. Z-DNA [VC83]. ZDO [Boč87, Boč88a, Boč88b, Boč89, PMT84, SHPK80]. Zeeman [Aus84, CV85]. zeolites [SZ84]. zero [GBZ86, YKZ^{+}89]. Zeroes [Rei80]. zeta [GG82b, KO81, Lás85]. Zinc [Das82]. Zintl [WBM88]. ZN [LG84, WBK86]. ZnS [FD86]. zone [MW87, Ori80, Ori82b, Ori84a, Ori85a, Ori85b, RB86]. Zvonimier [Zer87]. zwitterion [NSL88, RC80a]. zwitterions [BVDA88, SG86].

References

Avery:1989:NAQ

[AA89] John Avery and Frank Antonsen. A new approach to the quantum mechanics of atoms and small molecules. International Journal of Quantum Chemistry, 36(S23):159–168, April 1–8, 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Abdul-Ahad:1980:QSA

[AABW80] P. G. Abdul-Ahad, T. Blair, and G. A. Webb. Quantitative structure–activity study of some enzyme-inhibitory quinazolines. International Journal of Quantum Chemistry, 17 (5):821–831, May 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Abdul-Ahad:1982:IAA

[AAW82a] P. G. Abdul-Ahad and G. A. Webb. Interpretation of the antitumour activity of some diacridines. International Journal of Quantum Chemistry, 21(6):945–965, June 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Abdul-Ahad:1982:QSA

[AAW82b] P. G. Abdul-Ahad and G. A. Webb. Quantitative structure activity relationships for some antitumor platinum(II) complexes. International Journal of Quantum Chemistry, 21(6): 1105–1115, June 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Adamowicz:1984:NEN

[AB84] Ludwik Adamowicz and Rodney J. Bartlett. New efficient numerical method for solving pair correlation equations for diatomic molecules. International Journal of Quantum Chemistry, 26(2):213–221, August 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Adamowicz:1985:DCC

[AB85] Ludwik Adamowicz and Rodney J. Bartlett. Direct coupled cluster calculations on excited states. International Journal of Quantum Chemistry, 28(S19):217–220, March 18–23, 1985. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Adamowicz:1987:MCC

[AB87] Ludwik Adamowicz and Rodney J. Bartlett. MBPT and coupled cluster calculation on the neon atom with numerical orbitals. International Journal of Quantum Chemistry, 31(1): 173–177, January 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Arickx:1986:GWP

[ABCV86] F. Arickx, J. Broeckhove, W. Coene, and P. Van Leuven. Gaussian wave-packet dynamics. International Journal of Quantum Chemistry, 30(S20):471–481, March 10–15, 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Syposium on Atomic, Molecular, and Solid-State Theory, Scattering Problems, Many Body Phenomena, and Computational Quantum Chemistry.

Abdulnur:1980:ISM

[Abd80a] S. F. Abdulnur. Ab-initio study of the mechanism for the reaction of a primary amine with an aldehyde group. International Journal of Quantum Chemistry, 18(S7):171–176, March 5–8, 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Andre:1980:ALC

[ABD^{+}80b] J.-M. André, J.-L. Brédas, J. Delhalle, Y. Kalenov, L. Piela, and J. L. Calais. AMO in linear chains of hydrogen atoms revisited. International Journal of Quantum Chemistry, 18 (S14):419–429, March 10–20, 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory, Collision Phenomena, Quantum Statistics, and Computational Methods.

Andre:1981:SAB

[ABD81] J. M. André, J. L. Brédas, and J. Delhalle. Some aspects of bonding and binding in polymeric chains. International Journal of Quantum Chemistry, 19(6):1105–1111, June 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Andre:1983:ECB

[ABTP83] J. M. André, J. L. Brédas, B. Thémans, and L. Piela. Exchange, correlation, and band gaps in conjugated linear chains. International Journal of Quantum Chemistry, 23(3):1065– 1072, March 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Adams:1985:HAH

[AC85] William H. Adams and Meredith M. Clayton. The hydrogen atom in the hydrogen molecule. International Journal of Quantum Chemistry, 28(S19):333–348, March 18–23, 1985. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Alexander:1989:CNW

[AC89] S. A. Alexander and R. L. Coldwell. Constructing a nuclear wavefunction with splines. International Journal of Quantum Chemistry, 36(S23):169–181, April 1–8, 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Au-Chin:1980:CID

[ACCC80] Tang Au-Chin and Sun Chia-Chung. Conservation of intrinsic dual and single symmetries of molecular orbitals. International Journal of Quantum Chemistry, 18(1):57–63, July 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Au-Chin:1980:ALA

[ACCCBF80] Tang Au-Chin, Sun Chia-Chung, and Li Bei-Foo. Application of lie algebra to ligand-field theory. International Journal of Quantum Chemistry, 18(2):545–555, August 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Au-Chin:1980:CCO

[ACCYCC80] Tang Au-Chin, Cho Ching-Yu, and Sun Chia-Chung. V - coupling coefficients from SO(3) to the octahedral group. International Journal of Quantum Chemistry, 18(2):557–573, August 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Avery:1984:AHC

[ACDLH84] John Avery, Birte L. Christensen-Dalsgaard, Peter Sommer Larsen, and Shen Hengyi. Application of hyperspherical coordinates to the correlation problem. International Journal of Quantum Chemistry, 26(S18):321–338, March 1–15, 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Atomic, Molecular and Solid-State Theory, and Computational Quantum Chemistry.

Aagren:1985:DRS

[ÅCFRJ85] H. Ågren, N. Correia, A. Flores-Riveros, and H. J. Aa. Jensen. Direct restricted-step MCSCF calculations on the structure and spectrum of cyclobutadiene. International Journal of Quantum Chemistry, 28(S19):237–246, March 18–23, 1985. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Alagona:1989:NIM

[ACGT89] G. Alagona, R. Cammi, C. Ghio, and J. Tomasi. Noncovalent interactions of medium strength. A revised interpretation and examples of its applications. International Journal of Quantum Chemistry, 35(1):223–239, January 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Au-Chin:1983:COU

[ACH83] Tang Au-Chin and Guo Hong. Characteristic operators and unitarily invariant decomposition of Hermitian operators. International Journal of Quantum Chemistry, 23(1):217–226, January 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Adams:1982:RTP

[AČP82a] B. G. Adams, J. Čížek, and J. Paldus. Representation theory of so (4,2) for the perturbation treatment of hydrogenic-type Hamiltonians by algebraic methods. International Journal of Quantum Chemistry, 21(1):153–171, January 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Alder:1982:CSP

[ACP82b] B. J. Alder, D. M. Ceperley, and E. L. Pollock. Computer simulation of phase transitions in classical and quantum systems. International Journal of Quantum Chemistry, 22(S16): 49–61, March 1–13, 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Chemistry, Theory of Condensed Matter, and Propagator Methods in the Quantum Theory of Matter.

Adams:1984:EPT

[ACP84] William H. Adams, Meredith M. Clayton, and E. E. Polymeropoulos. Exchange perturbation theory calculations of the interaction energy between two ground-state hydrogen atoms. International Journal of Quantum Chemistry, 26(S18):393– 406, March 1–15, 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Atomic, Molecular and Solid-State Theory, and Computational Quantum Chemistry.

Andrews:1980:SES

[AD80] P. R. Andrews and J. A. Defina. Stereochemistry and electronic structure of anticonvulsant drugs. International Journal of Quantum Chemistry, 18(S7):297–313, March 5–8, 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Adamson:1964:UPC

[Ada64] Arthur W. Adamson. Understanding physical chemistry; a set of simple yet difficult examination questions and their methods of solution. W. A. Benjamin, Inc., New York, NY, USA, 1964. xix + 489 (two volumes) pp. LCCN QD453 .A26.

Adamson:1969:UPC

[Ada69] Arthur W. Adamson. Understanding physical chemistry. W. A. Benjamin, Inc., New York, NY, USA, 2d edition, 1969. ISBN 0-8053-0127-5. xiii + 516 pp. LCCN QD456 .A3 1969.

Adamson:1980:UPC

[Ada80a] Arthur W. Adamson. Understanding physical chemistry. Benjamin/Cummings Pub. Co., Menlo Park, CA, USA, 3d edition, 1980. ISBN 0-8053-0128-3. xii + 482 pp. LCCN QD456 .A3 1980.

Adawi:1980:BED

[Ada80b] I. Adawi. Behavior of the electron density near an impurity. International Journal of Quantum Chemistry, 17(4):799–801, April 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Adamowicz:1981:BSG

[Ada81] Ludwik Adamowicz. Basis set generation for the SCF calculation. International Journal of Quantum Chemistry, 19 (4):545–551, April 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Adams:1982:CTC

[Ada82] William H. Adams. Can there be a convergent perturbation theory of intermolecular energies? International Journal of Quantum Chemistry, 22(S16):319–330, March 1–13, 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Chemistry, Theory of Condensed Matter, and Propagator Methods in the Quantum Theory of Matter.

Adawi:1986:EDN

[Ada86] I. Adawi. Electron density near an impurity. International Journal of Quantum Chemistry, 30(1):27–29, July 1986. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Adelman:1987:GLT

[Ade87] Steven A. Adelman. Generalized Langevin theory for manybody problems in chemical dynamics: Explicit coordinate motion in molecular solvents. International Journal of Quantum Chemistry, 32(S21):199–215, March 12, 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Chemistry, Solid-State Theory, and Computational Methods.

Andre:1988:CMP

[ADF^{+}88] J. M. Andre, J. Delhalle, J. G. Fripiat, G. Hennico, and L. Piela. Calculations of molecular polarizabilities from electric-field-variant atomic orbitals: An analysis of the problem and its application to the hydrogen molecule and to the alkane series. International Journal of Quantum Chemistry, 34(S22):665–678, March 12–19, 1988. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Syposium on Quantum Chemistry, Solid-State Theory, and Computational Methods.

Andre:1978:QTP

[ADL78] Jean-Marie André, Joseph Delhalle, and János Ladik, editors. Quantum theory of polymers: proceedings of the NATO Advanced Study Institute on Electronic Structure and Properties of Polymers held at Namur, Belgium, 31 August–14 September 1977, volume 39 of NATO advanced study institutes series: Series C, Mathematical and physical sciences. D. Reidel, Dordrecht, Boston, Lancaster, Tokyo, 1978. ISBN 90-277-0870-3. LCCN QD380 .N36 1977.

Anisimov:1983:ESP

[ADS83] F. Anisimov, R. Dagys, and A. Sargautis. The electronic structure and properties of nonmetallic rare earth monochalcogenides. International Journal of Quantum Chemistry, 24(5): 429–433, November 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Almlof:1989:DE

[ADT^{+}89] Jan Almlöf, Bradley J. Deleeuw, Peter R. Taylor, Charles W. Bauschlicher, Jr., and Per Siegbahn. The dissociation energy of N_{2}. International Journal of Quantum Chemistry, 36(S23): 345–354, April 1–8, 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Aiken:1980:LO

[AEG80] John G. Aiken, John A. Erdos, and Jerome A. Goldstein. On Löwdin orthogonalization. International Journal of Quantum Chemistry, 18(4):1101–1108, October 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Abu-Eittah:1987:SSS

[AEHA87] R. H. Abu-Eittah, M. M. Hamed, and M. M. Abdou. Spectroscopic studies on some phenylethylamine drugs: Molecular orbital calculations. International Journal of Quantum Chemistry, 31(6):855–869, June 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Abu-Eittah:1981:MOT

[AEHH81] Rafie Abu-Eittah, Rifaat Hilal, and Maher M. Hamed. Molecular orbital treatment of phenylfurans and bifurans. International Journal of Quantum Chemistry, 19(3):383–399, March 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Abu-Eittah:1985:ESP

[AEHM85] Rafie Abu-Eittah, Rifaat Hilal, and Hussein Moustafa. Electronic structure of the peptide linkage. II. A molecular orbital treatment of the electronic spectra of benzohydroxamic acids. International Journal of Quantum Chemistry, 27(2):115–133, February 1985. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Abu-Eittah:1985:MOT

[AEHNEA85] Rafie H. Abu-Eittah, Maher M. Hamed, Abdullah Nigm, and Adel El-Azhary. Molecular-orbital treatment of some heterocycles with three hetero-atoms and their benzo-derivatives. International Journal of Quantum Chemistry, 28(6):731–740, December 1985. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Adamowicz:1986:EKT

[AEM86] Ludwik Adamowicz, James C. Ellenbogen, and E. A. McCullough, Jr. Extended-Koopmans-theorem approach to ab initio calculations upon the ground state and first excited state of the LiH anion. International Journal of Quantum Chemistry, 30(5):617–623, November 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Aasbrink:1981:CBM

[ÅFL81] L. Åsbrink, C. Fridh, and E. Lindholm. Concept of bond in the MO method HAM. International Journal of Quantum Chemistry, 19(6):1203–1206, June 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Alonso:1982:SAR

[AGB82] J. A. Alonso, D. J. González, and L. C. Balbás. Study of an approximate relation between the energy of an atom and the electronic potential at the nucleus. International Journal of Quantum Chemistry, 22(5):989–997, November 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Alagona:1987:DSI

[AGCT87a] Giuliano Alagona, Caterina Ghio, Roberto Cammi, and Jacopo Tomasi. The decomposition of the SCF interaction energy in hydrogen bonded dimers corrected for basis set superposition errors: An examination of the basis set dependence. International Journal of Quantum Chemistry, 32(2):227–248, August 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Alagona:1987:EFL

[AGCT87b] Giuliano Alagona, Caterina Ghio, Roberto Cammi, and Jacopo Tomasi. The effect of “full” and “limited” counterpoise corrections with different basis sets on the energy and the equilibrium distance of hydrogen bonded dimers. International Journal of Quantum Chemistry, 32(2):207–226, August 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Aagren:1980:BRB

[Ågr80] Hans Ågren. Book review: Atomic energy levels. Serafin Fraga, K. M. S. Saxena, and Jacek Karmowski. Elsevier Scientific Publishing Company, Amsterdam, 1979. International Journal of Quantum Chemistry, 18(3):911, September 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). See [FSK79].

Aagren:1982:BRB

[Ågr82] Hans Ågren. Book review: Spin–orbit coupling in molecules. By W. C. Richards, H. P. Trivedi, and D. L. Cooper Clarendon Press, Oxford, 1981. International Journal of Quantum Chemistry, 22(2):443, August 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). See [RTC81].

Aagren:1983:CHI

[Ågr83] Hans Ågren. Core hole induced charge-transfer reactions in disubstituted benzenes. International Journal of Quantum Chemistry, 23(2):577–585, February 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Aagren:1985:BRB

[Ågr85] Hans Ågren. Book review: Quantum chemistry: The development of ab initio methods in molecular electronic structure theory. By Henry F. Schaeffer, Clarendon press, Oxford, England, 1984. International Journal of Quantum Chemistry, 27 (5):637, May 1985. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Aagren:1987:BRBa

[Ågr87a] Hans Ågren. Book review: Comparison of Ab Initio quantum chemistry with experiment for small molecules. The state of the art. Edited by Rodney J. Bartlett, D. Reidel Publishing Company, 1985. International Journal of Quantum Chemistry, 31(2):325, February 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). See [Bar85].

Aagren:1987:BRBb

[Ågr87b] Hans Ågren. Book review: Geometrical derivatives of energy surfaces and molecular properties. Edited by Poul Jørgensen and Jack Simons, D. Reidel publishing company, Dordrecht, 1986. International Journal of Quantum Chemistry, 32(4): 547, October 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). See [JS86].

Aagren:1987:BRBc

[Ågr87c] Hans Ågren. Book review: Handbook of Gaussian basis sets: A compendium for ab-initio molecular orbital calculations. By Raymond Poirier, Roy Kari, and Imre Csizmadia. Elsevier science publishers B.V., Amsterdem, 1985. International Journal of Quantum Chemistry, 32(4):549, October 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). See [PKC85].

Aagren:1988:BRB

[Ågr88] Hans Ågren. Book review: Relativistic theory of atoms and molecules. A Bibliography 1916–1985. By P. Pyykkö, Lecture Notes in Chemistry, Vol. 41, Edited by G. Berthier, et al., Springer-Verlag, Berlin, 1986. International Journal of Quantum Chemistry, 34(5):487, November 1988. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). See [Pyy86, Pyy93, Pyy00].

Avery:1986:OPL

[AH86] John Avery and Søren Hvidt. Optical properties of large molecules and clusters in the Frenkel exciton picture. International Journal of Quantum Chemistry, 29(3):497–510, March 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Allen:1980:TCE

[AHB^{+}80] Roland E. Allen, Harold P. Hjalmarson, Helmut Büttner, Peter Vogl, Donald J. Wolford, Otto F. Sankey, and John D. Dow. Theory of core exciton binding energies for excitons near interfaces. International Journal of Quantum Chemistry, 18(S14):607–615, March 10–20, 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory, Collision Phenomena, Quantum Statistics, and Computational Methods.

Andrews:1982:DTS

[AIJW82] P. R. Andrews, M. N. Iskander, G. P. Jones, and D. A. Winkler. Design of transition-state analogues for GABAtransaminase. International Journal of Quantum Chemistry, 22(S9):345–353, March 4–6, 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Adams:1983:SQE

[AJ83] B. G. Adams and K. Jankowski. A study of quasidegeneracy effects in the be-isoelectronic series. International Journal of Quantum Chemistry, 24(S17):297–306, March 3–12, 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-state Theory, Collision Phenomena, and Computational Quantum Chemistry.

Albertsen:1980:MTD

[AJY80] Preben Albertsen, Poul Jørgensen, and Danny L. Yeager. Multiconfigurational time-dependent Hartree–Fock calculation of vertical excitation energies and transition moments of O_{2}. International Journal of Quantum Chemistry, 18(S14):249–260, March 10–20, 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory, Collision Phenomena, Quantum Statistics, and Computational Methods.

Ananchenkov:1980:REB

[AL80] V. I. Ananchenkov and V. B. Leontyev. Realization of the elementary biological function as a localized electron subsystem interaction process. International Journal of Quantum Chemistry, 17(2):265–272, February 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Atabek:1981:MQS

[AL81] O. Atabek and R. Lefebvre. Multichannel quantization and the static Stark effect. International Journal of Quantum Chemistry, 19(5):901–906, May 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Avery:1986:COA

[AL86] John Avery and Peter Sommer Larsen. Construction of orbital angular momentum eigenfunctions for many-particle systems by harmonic projection. International Journal of Quantum Chemistry, 29(6):1735–1743, June 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Assali:1987:GSE

[AL87] L. V. C. Assali and J. R. Leite. Ground state electronic properties of Fe–B complex pair in silicon. International Journal of Quantum Chemistry, 32(S21):89–97, March 12, 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Chemistry, Solid-State Theory, and Computational Methods.

Adamowicz:1984:AGC

[ALB84] L. Adamowicz, W. D. Laidig, and R. J. Bartlett. Analytical gradients for the coupled-cluster method. International Journal of Quantum Chemistry, 26(S18):245–254, March 1– 15, 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Atomic, Molecular and Solid-State Theory, and Computational Quantum Chemistry.

Avery:1986:QMM

[ALH86] John Avery, Peter Sommer Larsen, and Shen Hengyi. The quantum mechanical many-body problem in hyperspherical coordinates. Analysis of systems with Coulomb interactions in terms of many-dimensional hydrogen-like wave functions. International Journal of Quantum Chemistry, 29(1):129–147, January 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Altmann:1983:SDC

[ALL83] S. L. Altmann, A. Lapiccirella, and K. W. Lodge. Structure of dislocation cores in the silicon crystal. International Journal of Quantum Chemistry, 23(3):1057–1063, March 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Allan:1987:SRE

[All87] Michael Allan. Some recent experiments on inelastic electronmolecule collisions at low energies. International Journal of Quantum Chemistry, 31(1):161–172, January 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Altmann:1986:RQD

[Alt86] Simon L. Altmann. Rotations, quaternions, and double groups. Oxford science publications. Clarendon Press, New York, NY, USA, 1986. ISBN 0-19-855372-2. xiv + 317 pp. LCCN QC174.17.R65 A48 1986. URL http://www.loc.gov/

catdir/enhancements/fy0639/85013615-d.html.

Altmann:1989:POC

[Alt89] S. L. Altmann. Projection operators and Clebsch–Gordan coefficients for projective representations. International Journal of Quantum Chemistry, 35(3):441–456, March 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Altmann:2005:RQD

[Alt05] Simon L. Altmann. Rotations, quaternions, and double groups. Dover Publications, Inc., New York, NY, USA, 2005. ISBN 0-486-44518-6 (paperback). ???? pp. LCCN QC174.17.R65 A48

2005. URL http://www.loc.gov/catdir/enhancements/
fy0625/2005049630-d.html.

Alexander:1982:ICH

[AM82] S. A. Alexander and F. A. Matsen. Ab initio calculations of H_{2}O → OH + H. International Journal of Quantum Chemistry, 22(S16):445–450, March 1–13, 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Chemistry, Theory of Condensed Matter, and Propagator Methods in the Quantum Theory of Matter.

Adamowicz:1983:MBS

[AM83a] Ludwik Adamowicz and E. A. McCullough, Jr. Molecular basis set generation: Accurate Slater basis sets for LiH^{−} ground and excited state and Li_{2}^{−} ground state. International Journal of Quantum Chemistry, 24(1):19–23, July 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Allan:1983:CPS

[AM83b] N. L. Allan and N. H. March. Compton profile, (e, 2e) studies, and “atoms in molecules”. International Journal of Quantum Chemistry, 24(S17):227–240, March 3–12, 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-state Theory, Collision Phenomena, and Computational Quantum Chemistry.

Adamowicz:1984:TDS

[AM84] Ludwik Adamowicz and E. A. McCullough, Jr. N^{2}-time-dependent SCF scheme. International Journal of Quantum Chemistry, 26(3):373–381, September 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Alexander:1987:PSC

[AM87a] S. A. Alexander and H. J. Monkhorst. Polyatomic SCF calculations with numerical orbitals. I. Introduction and computational experiments on the CYBER 205. International Journal of Quantum Chemistry, 32(3):361–375, September 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Arteca:1987:MCM

[AM87b] Gustavo A. Arteca and Paul G. Mezey. A method for the characterization of molecular conformations. International Journal of Quantum Chemistry, 32(S14):133–147, March 14, 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Arteca:1988:MCM

[AM88a] Gustavo A. Arteca and Paul G. Mezey. Molecular conformations and molecular shape: a discrete characterization of continua of van der Waals surfaces. International Journal of Quantum Chemistry, 34(6):517–526, December 1988. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Arteca:1988:SDC

[AM88b] Gustavo A. Arteca and Paul G. Mezey. Shape description of conformationally flexible molecules: Application to twodimensional conformational problems. International Journal of Quantum Chemistry, 34(S15):33–54, March 12–19, 1988. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Syposium on Quantum Biology and Quantum Pharmacology.

Arteca:1989:TAC

[AM89] Gustavo A. Arteca and Paul G. Mezey. Two approaches to the concept of chemical species: Relations between potential energy and molecular shape. International Journal of Quantum Chemistry, 36(S23):305–320, April 1–8, 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Arteca:1983:DSA

[AMFC83] Gustavo A. Arteca, Sergio A. Maluendes, Francisco M. Fernández, and Eduardo A. Castro. Discussion of several analytical approximate expressions for the eigenvalues of the bounded harmonic oscillator and hydrogen atom. International Journal of Quantum Chemistry, 24(2):169–184, August 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Alexander:1983:ISS

[AMM83] S. A. Alexander, Carol Mcdonald, and F. A. Matsen. Ab initio surfaces of singlet and triplet methylene. International Journal of Quantum Chemistry, 24(S17):407–414, March 3–12, 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-state Theory, Collision Phenomena, and Computational Quantum Chemistry.

Aerts:1985:UGB

[AN85] P. J. C. Aerts and W. C. Nieuwpoort. On the use of Gaussian basis sets to solve the Hartree–Fock-Dirac equation. II. Application to many-electron atomic and molecular systems. International Journal of Quantum Chemistry, 28(S19):267– 277, March 18–23, 1985. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Aida:1986:IMO

[AN86] Misako Aida and Chikayoshi Nagata. An ab initio molecular orbital study on the stacking interaction between nucleic acid bases: Dependence on the sequence and relation to the conformation. International Journal of Quantum Chemistry, 29 (5):1253–1261, May 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Andersson:1985:IR

[And85] Johnny Andersson. Introductory remarks. International Journal of Quantum Chemistry, 28(S19):15–17, March 18–23, 1985. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Andrew:1987:HIN

[And87] E. Raymond Andrew. Human images by nuclear magnetic resonance. International Journal of Quantum Chemistry, 32 (S14):331–339, March 14, 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Almlof:1981:TES

[ANH81] J. Almlöf, A. V. Nemukhin, and A. Heiberg. Treatment of excited states in the CASSCF method: Application to BO. International Journal of Quantum Chemistry, 20(3):655–667, September 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1980:A

[Ano80a] Anonymous. Abstracts. International Journal of Quantum Chemistry, 18(S7):433–434, March 5–8, 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Anonymous:1980:AS

[Ano80b] Anonymous. ACS symposium. International Journal of Quantum Chemistry, 17(1):151, January 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1980:AMC

[Ano80c] Anonymous. Atomic and molecular collision theory. International Journal of Quantum Chemistry, 17(5):1042, May 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1980:BR

[Ano80d] Anonymous. Books Received. International Journal of Quantum Chemistry, 17(3):596, March 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1980:BW

[Ano80e] Anonymous. E. Bright Wilson. International Journal of Quantum Chemistry, 18(S14):9–16, March 10–20, 1980. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory, Collision Phenomena, Quantum Statistics, and Computational Methods.

Anonymous:1980:Ma

[Ano80f] Anonymous. Masthead. International Journal of Quantum Chemistry, 17(1):fmi–fmii, January 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1980:Mb

[Ano80g] Anonymous. Masthead. International Journal of Quantum Chemistry, 17(2):fmi–fmii, February 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1980:Mc

[Ano80h] Anonymous. Masthead. International Journal of Quantum Chemistry, 17(3):fmi–fmii, March 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1980:Md

[Ano80i] Anonymous. Masthead. International Journal of Quantum Chemistry, 17(4):fmi–fmii, April 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1980:Me

[Ano80j] Anonymous. Masthead. International Journal of Quantum Chemistry, 17(5):fmi–fmii, May 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1980:Mf

[Ano80k] Anonymous. Masthead. International Journal of Quantum Chemistry, 17(6):fmi–fmii, June 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1980:Mg

[Ano80l] Anonymous. Masthead. International Journal of Quantum Chemistry, 18(1):fmi–fmii, July 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1980:Mh

[Ano80m] Anonymous. Masthead. International Journal of Quantum Chemistry, 18(2):fmi–fmii, August 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1980:Mi

[Ano80n] Anonymous. Masthead. International Journal of Quantum Chemistry, 18(3):fmi–fmii, September 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1980:Mj

[Ano80o] Anonymous. Masthead. International Journal of Quantum Chemistry, 18(4):fmi–fmii, October 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1980:Mk

[Ano80p] Anonymous. Masthead. International Journal of Quantum Chemistry, 18(5):fmi–fmii, November 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1980:Ml

[Ano80q] Anonymous. Masthead. International Journal of Quantum Chemistry, 18(6):fmi–fmii, December 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1980:Mm

[Ano80r] Anonymous. Masthead. International Journal of Quantum Chemistry, 18(S7):fmi, March 5–8, 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Anonymous:1980:Mo

[Ano80s] Anonymous. Masthead. International Journal of Quantum Chemistry, 18(S7):fmi, March 5–8, 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Anonymous:1980:Mn

[Ano80t] Anonymous. Masthead. International Journal of Quantum Chemistry, 18(S14):fmi, March 10–20, 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory, Collision Phenomena, Quantum Statistics, and Computational Methods.

Anonymous:1981:Ac

[Ano81a] Anonymous. Abstracts. International Journal of Quantum Chemistry, 20(S8):481–482, March 5–7, 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Anonymous:1981:Ab

[Ano81b] Anonymous. Announcement. International Journal of Quantum Chemistry, 20(6):1337, December 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1981:Aa

[Ano81c] Anonymous. Announcements. International Journal of Quantum Chemistry, 20(2):589–590, August 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1981:BM

[Ano81d] Anonymous. Bernd T. Matthias. International Journal of Quantum Chemistry, 20(S15):5–19, March 8–14, 1981. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory, Collision Phenomena, and Computational Quantum Chemistry.

Anonymous:1981:I

[Ano81e] Anonymous. Introduction. International Journal of Quantum Chemistry, 20(S8):vii, March 5–7, 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Anonymous:1981:Ma

[Ano81f] Anonymous. Masthead. International Journal of Quantum Chemistry, 19(2):fmi–fmii, February 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1981:Mb

[Ano81g] Anonymous. Masthead. International Journal of Quantum Chemistry, 19(3):fmi–fmii, March 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1981:Mc

[Ano81h] Anonymous. Masthead. International Journal of Quantum Chemistry, 19(4):fmi–fmii, April 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1981:Md

[Ano81i] Anonymous. Masthead. International Journal of Quantum Chemistry, 19(5):fmi–fmii, May 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1981:Me

[Ano81j] Anonymous. Masthead. International Journal of Quantum Chemistry, 19(6):fmi–fmii, June 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1981:Mf

[Ano81k] Anonymous. Masthead. International Journal of Quantum Chemistry, 20(1):fmi–fmii, July 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1981:Mg

[Ano81l] Anonymous. Masthead. International Journal of Quantum Chemistry, 20(2):fmi–fmii, August 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1981:Mh

[Ano81m] Anonymous. Masthead. International Journal of Quantum Chemistry, 20(3):fmi–fmii, September 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1981:Mi

[Ano81n] Anonymous. Masthead. International Journal of Quantum Chemistry, 20(4):fmi–fmii, October 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1981:Mj

[Ano81o] Anonymous. Masthead. International Journal of Quantum Chemistry, 20(5):fmi–fmii, November 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1981:Mk

[Ano81p] Anonymous. Masthead. International Journal of Quantum Chemistry, 20(6):fmi–fmii, December 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1981:Ml

[Ano81q] Anonymous. Masthead. International Journal of Quantum Chemistry, 20(S8):fmi, March 5–7, 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Anonymous:1981:Mm

[Ano81r] Anonymous. Masthead. International Journal of Quantum Chemistry, 20(S15):fmi, March 8–14, 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory, Collision Phenomena, and Computational Quantum Chemistry.

Anonymous:1981:SSM

[Ano81s] Anonymous. Subsymposium on scattering by many-body systems and related computational methods. International Journal of Quantum Chemistry, 20(S15):557, March 8–14, 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory, Collision Phenomena, and Computational Quantum Chemistry.

Anonymous:1982:Ad

[Ano82a] Anonymous. Abstracts. International Journal of Quantum Chemistry, 22(S9):425–427, March 4–6, 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Anonymous:1982:Aa

[Ano82b] Anonymous. Announcement. International Journal of Quantum Chemistry, 21(4):767–768, April 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1982:Ac

[Ano82c] Anonymous. Announcement. International Journal of Quantum Chemistry, 22(1):219, July 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1982:Ab

[Ano82d] Anonymous. Announcements. International Journal of Quantum Chemistry, 21(5):943, May 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1982:Ma

[Ano82e] Anonymous. Masthead. International Journal of Quantum Chemistry, 21(1):fmi–fmii, January 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1982:Mb

[Ano82f] Anonymous. Masthead. International Journal of Quantum Chemistry, 21(2):fmi–fmii, February 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1982:Mc

[Ano82g] Anonymous. Masthead. International Journal of Quantum Chemistry, 21(3):fmi–fmii, March 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1982:Md

[Ano82h] Anonymous. Masthead. International Journal of Quantum Chemistry, 21(4):fmi–fmii, April 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1982:Me

[Ano82i] Anonymous. Masthead. International Journal of Quantum Chemistry, 21(5):fmi–fmii, May 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1982:Mf

[Ano82j] Anonymous. Masthead. International Journal of Quantum Chemistry, 21(6):fmi–fmii, June 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1982:Mg

[Ano82k] Anonymous. Masthead. International Journal of Quantum Chemistry, 22(1):fmi–fmii, July 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1982:Mh

[Ano82l] Anonymous. Masthead. International Journal of Quantum Chemistry, 22(2):fmi–fmii, August 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1982:Mi

[Ano82m] Anonymous. Masthead. International Journal of Quantum Chemistry, 22(3):fmi, September 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1982:Mj

[Ano82n] Anonymous. Masthead. International Journal of Quantum Chemistry, 22(4):fmi–fmii, October 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1982:Mk

[Ano82o] Anonymous. Masthead. International Journal of Quantum Chemistry, 22(5):fmi–fmii, November 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1982:Ml

[Ano82p] Anonymous. Masthead. International Journal of Quantum Chemistry, 22(6):fmi–fmii, December 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1982:Mm

[Ano82q] Anonymous. Masthead. International Journal of Quantum Chemistry, 22(S9):fmi, March 4–6, 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Anonymous:1982:Mn

[Ano82r] Anonymous. Masthead. International Journal of Quantum Chemistry, 22(S16):fmi, March 1–13, 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Chemistry, Theory of Condensed Matter, and Propagator Methods in the Quantum Theory of Matter.

Anonymous:1982:WPT

[Ano82s] Anonymous. Workshop on perturbation theory at large order: List of participants. International Journal of Quantum Chemistry, 21(1):209–211, January 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1983:A

[Ano83a] Anonymous. Announcement. International Journal of Quantum Chemistry, 24(1):137, July 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1983:E

[Ano83b] Anonymous. Erratum. International Journal of Quantum Chemistry, 24(2):241, August 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1983:Ma

[Ano83c] Anonymous. Masthead. International Journal of Quantum Chemistry, 23(1):fmi–fmii, January 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1983:Mb

[Ano83d] Anonymous. Masthead. International Journal of Quantum Chemistry, 23(2):fmi–fmii, February 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1983:Mc

[Ano83e] Anonymous. Masthead. International Journal of Quantum Chemistry, 23(3):fmi–fmii, March 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1983:Md

[Ano83f] Anonymous. Masthead. International Journal of Quantum Chemistry, 23(4):fmi–fmii, April 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1983:Me

[Ano83g] Anonymous. Masthead. International Journal of Quantum Chemistry, 23(5):fmi–fmii, May 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1983:Mf

[Ano83h] Anonymous. Masthead. International Journal of Quantum Chemistry, 23(6):fmi–fmii, June 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1983:Mg

[Ano83i] Anonymous. Masthead. International Journal of Quantum Chemistry, 24(1):fmi–fmii, July 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1983:Mh

[Ano83j] Anonymous. Masthead. International Journal of Quantum Chemistry, 24(2):fmi–fmii, August 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1983:Mi

[Ano83k] Anonymous. Masthead. International Journal of Quantum Chemistry, 24(3):fmi–fmii, September 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1983:Mj

[Ano83l] Anonymous. Masthead. International Journal of Quantum Chemistry, 24(4):fmi–fmii, October 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1983:Mk

[Ano83m] Anonymous. Masthead. International Journal of Quantum Chemistry, 24(5):fmi–fmii, November 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1983:Ml

[Ano83n] Anonymous. Masthead. International Journal of Quantum Chemistry, 24(6):fmi–fmii, December 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1983:Mm

[Ano83o] Anonymous. Masthead. International Journal of Quantum Chemistry, 24(S10):i, March 14–16, 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Anonymous:1983:Mn

[Ano83p] Anonymous. Masthead. International Journal of Quantum Chemistry, 24(S17):fmi, March 3–12, 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-state Theory, Collision Phenomena, and Computational Quantum Chemistry.

Anonymous:1984:AIA

[Ano84a] Anonymous. Announcement. International academy of quantum molecular sciences. International Journal of Quantum Chemistry, 25(1):273, January 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1984:Ia

[Ano84b] Anonymous. Introduction. International Journal of Quantum Chemistry, 26(S11):vii–viii, March 12–15, 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Anonymous:1984:Ib

[Ano84c] Anonymous. Introduction. International Journal of Quantum Chemistry, 26(S18):iii–iv, March 1–15, 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Atomic, Molecular and Solid-State Theory, and Computational Quantum Chemistry.

Anonymous:1984:Ma

[Ano84d] Anonymous. Masthead. International Journal of Quantum Chemistry, 25(1):fmi, January 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1984:Mb

[Ano84e] Anonymous. Masthead. International Journal of Quantum Chemistry, 25(2):fmi–fmii, February 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1984:Mc

[Ano84f] Anonymous. Masthead. International Journal of Quantum Chemistry, 25(3):fmi–fmii, March 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1984:Md

[Ano84g] Anonymous. Masthead. International Journal of Quantum Chemistry, 25(4):fmi–fmii, April 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1984:Me

[Ano84h] Anonymous. Masthead. International Journal of Quantum Chemistry, 25(5):fmi–fmii, May 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1984:Mf

[Ano84i] Anonymous. Masthead. International Journal of Quantum Chemistry, 25(6):fmi–fmii, June 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1984:Mg

[Ano84j] Anonymous. Masthead. International Journal of Quantum Chemistry, 26(1):fmi–fmii, July 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1984:Mh

[Ano84k] Anonymous. Masthead. International Journal of Quantum Chemistry, 26(2):fmi–fmii, August 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1984:Mi

[Ano84l] Anonymous. Masthead. International Journal of Quantum Chemistry, 26(3):fmi–fmii, September 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1984:Mj

[Ano84m] Anonymous. Masthead. International Journal of Quantum Chemistry, 26(4):fmi–fmii, October 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1984:Mk

[Ano84n] Anonymous. Masthead. International Journal of Quantum Chemistry, 26(5):fmi–fmii, November 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1984:Ml

[Ano84o] Anonymous. Masthead. International Journal of Quantum Chemistry, 26(6):fmi–fmii, December 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1984:Mm

[Ano84p] Anonymous. Masthead. International Journal of Quantum Chemistry, 26(S11):fmi, March 12–15, 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Anonymous:1984:Mn

[Ano84q] Anonymous. Masthead. International Journal of Quantum Chemistry, 26(S18):fmi, March 1–15, 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Atomic, Molecular and Solid-State Theory, and Computational Quantum Chemistry.

Anonymous:1985:Ab

[Ano85a] Anonymous. Abstract. International Journal of Quantum Chemistry, 28(S12):287–288, March 14–16, 1985. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Anonymous:1985:Aa

[Ano85b] Anonymous. Announcement. International Journal of Quantum Chemistry, 28(5):663, November 1985. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1985:LPP

[Ano85c] Anonymous. List of participants: Part II. International Journal of Quantum Chemistry, 28(S19):5–14, March 18–23, 1985. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Anonymous:1985:Ma

[Ano85d] Anonymous. Masthead. International Journal of Quantum Chemistry, 27(1):fmi–fmii, January 1985. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1985:Mb

[Ano85e] Anonymous. Masthead. International Journal of Quantum Chemistry, 27(2):fmi–fmii, February 1985. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1985:Mc

[Ano85f] Anonymous. Masthead. International Journal of Quantum Chemistry, 27(3):fmi–fmii, March 1985. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1985:Md

[Ano85g] Anonymous. Masthead. International Journal of Quantum Chemistry, 27(4):fmi–fmii, April 1985. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1985:Me

[Ano85h] Anonymous. Masthead. International Journal of Quantum Chemistry, 27(5):fmi–fmii, May 1985. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1985:Mf

[Ano85i] Anonymous. Masthead. International Journal of Quantum Chemistry, 27(6):fmi–fmii, June 1985. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1985:Mg

[Ano85j] Anonymous. Masthead. International Journal of Quantum Chemistry, 28(1):fmi–fmii, July 1985. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1985:Mh

[Ano85k] Anonymous. Masthead. International Journal of Quantum Chemistry, 28(2):fmi–fmii, August 1985. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1985:Mi

[Ano85l] Anonymous. Masthead. International Journal of Quantum Chemistry, 28(3):fmi–fmii, September 1985. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1985:Mj

[Ano85m] Anonymous. Masthead. International Journal of Quantum Chemistry, 28(4):fmi, October 1985. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1985:Mk

[Ano85n] Anonymous. Masthead. International Journal of Quantum Chemistry, 28(5):fmi–fmii, November 1985. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1985:Ml

[Ano85o] Anonymous. Masthead. International Journal of Quantum Chemistry, 28(6):fmi–fmii, December 1985. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1985:Mm

[Ano85p] Anonymous. Masthead. International Journal of Quantum Chemistry, 28(S12):fmi, March 14–16, 1985. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Anonymous:1985:Mn

[Ano85q] Anonymous. Masthead. International Journal of Quantum Chemistry, 28(S19):fmi, March 18–23, 1985. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Anonymous:1986:Aa

[Ano86a] Anonymous. Announcement. International Journal of Quantum Chemistry, 29(5):1685, May 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1986:Ab

[Ano86b] Anonymous. Announcement. International Journal of Quantum Chemistry, 30(5):715, November 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1986:EC

[Ano86c] Anonymous. Editorial comments. International Journal of Quantum Chemistry, 29(2):ix, February 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1986:E

[Ano86d] Anonymous. Erratum. International Journal of Quantum Chemistry, 30(S20):772, March 10–15, 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Syposium on Atomic, Molecular, and Solid-State Theory, Scattering Problems, Many Body Phenomena, and Computational Quantum Chemistry.

Anonymous:1986:LPa

[Ano86e] Anonymous. List of Participants. International Journal of Quantum Chemistry, 29(2):xi–xl, February 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1986:LPb

[Ano86f] Anonymous. List of participants. International Journal of Quantum Chemistry, 30(S20):3–12, March 10–15, 1986. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Syposium on Atomic, Molecular, and Solid-State Theory, Scattering Problems, Many Body Phenomena, and Computational Quantum Chemistry.

Anonymous:1986:Ma

[Ano86g] Anonymous. Masthead. International Journal of Quantum Chemistry, 29(1):fmi–fmii, January 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1986:Mb

[Ano86h] Anonymous. Masthead. International Journal of Quantum Chemistry, 29(2):fmi–fmii, February 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1986:Mc

[Ano86i] Anonymous. Masthead. International Journal of Quantum Chemistry, 29(3):fmi–fmii, March 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1986:Md

[Ano86j] Anonymous. Masthead. International Journal of Quantum Chemistry, 29(4):fmi–fmii, April 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1986:Me

[Ano86k] Anonymous. Masthead. International Journal of Quantum Chemistry, 29(5):fmi–fmii, May 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1986:Mf

[Ano86l] Anonymous. Masthead. International Journal of Quantum Chemistry, 29(6):fmi–fmii, June 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1986:Mg

[Ano86m] Anonymous. Masthead. International Journal of Quantum Chemistry, 30(1):fmi–fmii, July 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1986:Mh

[Ano86n] Anonymous. Masthead. International Journal of Quantum Chemistry, 30(2):fmi, August 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1986:Mi

[Ano86o] Anonymous. Masthead. International Journal of Quantum Chemistry, 30(3):fmi–fmii, September 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1986:Mj

[Ano86p] Anonymous. Masthead. International Journal of Quantum Chemistry, 30(4):fmi–fmii, October 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1986:Mk

[Ano86q] Anonymous. Masthead. International Journal of Quantum Chemistry, 30(5):fmi–fmii, November 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1986:Ml

[Ano86r] Anonymous. Masthead. International Journal of Quantum Chemistry, 30(6):fmi–fmii, December 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1986:Mm

[Ano86s] Anonymous. Masthead. International Journal of Quantum Chemistry, 30(S13):fmi, March 6–8, 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Anonymous:1986:Mn

[Ano86t] Anonymous. Masthead. International Journal of Quantum Chemistry, 30(S20):fmi, March 10–15, 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Syposium on Atomic, Molecular, and Solid-State Theory, Scattering Problems, Many Body Phenomena, and Computational Quantum Chemistry.

Anonymous:1987:ESP

[Ano87a] Anonymous. The electronic structure and properties of molecules and molecular crystals. International Journal of Quantum Chemistry, 32(3):403, September 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1987:Ia

[Ano87b] Anonymous. Introduction. International Journal of Quantum Chemistry, 31(1):1, January 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1987:Ib

[Ano87c] Anonymous. Introduction. International Journal of Quantum Chemistry, 32(S14):v, March 14, 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Anonymous:1987:LP

[Ano87d] Anonymous. List of participants. International Journal of Quantum Chemistry, 32(S21):3–11, March 12, 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Chemistry, Solid-State Theory, and Computational Methods.

Anonymous:1987:Ma

[Ano87e] Anonymous. Masthead. International Journal of Quantum Chemistry, 31(1):fmi–fmii, January 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1987:Mb

[Ano87f] Anonymous. Masthead. International Journal of Quantum Chemistry, 31(2):fmi–fmii, February 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1987:Mc

[Ano87g] Anonymous. Masthead. International Journal of Quantum Chemistry, 31(3):fmi–fmii, March 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1987:Md

[Ano87h] Anonymous. Masthead. International Journal of Quantum Chemistry, 31(4):fmi–fmii, April 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1987:Me

[Ano87i] Anonymous. Masthead. International Journal of Quantum Chemistry, 31(5):fmi–fmii, May 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1987:Mf

[Ano87j] Anonymous. Masthead. International Journal of Quantum Chemistry, 31(6):fmi–fmii, June 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1987:Mg

[Ano87k] Anonymous. Masthead. International Journal of Quantum Chemistry, 32(1):fmi–fmii, July 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1987:Mh

[Ano87l] Anonymous. Masthead. International Journal of Quantum Chemistry, 32(2):fmi–fmii, August 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1987:Mi

[Ano87m] Anonymous. Masthead. International Journal of Quantum Chemistry, 32(3):fmi–fmii, September 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1987:Mj

[Ano87n] Anonymous. Masthead. International Journal of Quantum Chemistry, 32(4):fmi–fmii, October 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1987:Mk

[Ano87o] Anonymous. Masthead. International Journal of Quantum Chemistry, 32(5):fmi–fmii, November 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1987:Ml

[Ano87p] Anonymous. Masthead. International Journal of Quantum Chemistry, 32(6):fmi–fmii, December 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1987:Mm

[Ano87q] Anonymous. Masthead. International Journal of Quantum Chemistry, 32(S14):fmi, March 14, 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Anonymous:1987:Mn

[Ano87r] Anonymous. Masthead. International Journal of Quantum Chemistry, 32(S21):fmi, March 12, 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Chemistry, Solid-State Theory, and Computational Methods.

Anonymous:1987:PPA

[Ano87s] Anonymous. Picture of professor Andrew C. Hurley. International Journal of Quantum Chemistry, 31(3):335, March 1987.

CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1987:PCF

[Ano87t] Anonymous. Picture of the “Cambridge five”. International Journal of Quantum Chemistry, 31(3):337, March 1987. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1987:SS

[Ano87u] Anonymous. Sanibel Symposium. International Journal of Quantum Chemistry, 32(6):723, December 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1988:Aa

[Ano88a] Anonymous. Announcement. International Journal of Quantum Chemistry, 33(2):169, February 1988. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1988:Ab

[Ano88b] Anonymous. Announcement. International Journal of Quantum Chemistry, 33(4):369, April 1988. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1988:Ac

[Ano88c] Anonymous. Announcement. International Journal of Quantum Chemistry, 33(4):371, April 1988. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1988:Ad

[Ano88d] Anonymous. Announcement. International Journal of Quantum Chemistry, 33(5):477, May 1988. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1988:Ae

[Ano88e] Anonymous. Announcement. International Journal of Quantum Chemistry, 33(5):479, May 1988. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1988:Af

[Ano88f] Anonymous. Announcement. International Journal of Quantum Chemistry, 34(2):191–192, August 1988. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1988:Ag

[Ano88g] Anonymous. Announcement. International Journal of Quantum Chemistry, 34(3):305, September 1988. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1988:E

[Ano88h] Anonymous. Erratum. International Journal of Quantum Chemistry, 34(S22):741, March 12–19, 1988. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Syposium on Quantum Chemistry, Solid-State Theory, and Computational Methods.

Anonymous:1988:LPa

[Ano88i] Anonymous. List of participants. International Journal of Quantum Chemistry, 34(S15):vii–xxvii, March 12–19, 1988. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Syposium on Quantum Biology and Quantum Pharmacology.

Anonymous:1988:LPb

[Ano88j] Anonymous. List of participants. International Journal of Quantum Chemistry, 34(S22):xi–xxii, March 12–19, 1988. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Syposium on Quantum Chemistry, Solid-State Theory, and Computational Methods.

Anonymous:1988:Ma

[Ano88k] Anonymous. Masthead. International Journal of Quantum Chemistry, 33(1):fmi–fmii, January 1988. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1988:Mb

[Ano88l] Anonymous. Masthead. International Journal of Quantum Chemistry, 33(2):fmi–fmii, February 1988. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1988:Mc

[Ano88m] Anonymous. Masthead. International Journal of Quantum Chemistry, 33(3):fmi–fmii, March 1988. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1988:Md

[Ano88n] Anonymous. Masthead. International Journal of Quantum Chemistry, 33(4):fmi–fmii, April 1988. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1988:Me

[Ano88o] Anonymous. Masthead. International Journal of Quantum Chemistry, 33(5):fmi–fmii, May 1988. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1988:Mf

[Ano88p] Anonymous. Masthead. International Journal of Quantum Chemistry, 33(6):fmi–fmii, June 1988. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1988:Mg

[Ano88q] Anonymous. Masthead. International Journal of Quantum Chemistry, 34(1):fmi–fmii, July 1988. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1988:Mh

[Ano88r] Anonymous. Masthead. International Journal of Quantum Chemistry, 34(2):fmi–fmii, August 1988. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1988:Mi

[Ano88s] Anonymous. Masthead. International Journal of Quantum Chemistry, 34(3):fmi–fmii, September 1988. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1988:Mj

[Ano88t] Anonymous. Masthead. International Journal of Quantum Chemistry, 34(4):fmi–fmii, October 1988. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1988:Mk

[Ano88u] Anonymous. Masthead. International Journal of Quantum Chemistry, 34(5):fmi–fmiii, November 1988. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1988:Ml

[Ano88v] Anonymous. Masthead. International Journal of Quantum Chemistry, 34(6):fmi–fmii, December 1988. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1988:Mm

[Ano88w] Anonymous. Masthead. International Journal of Quantum Chemistry, 34(S15):fmi, March 12–19, 1988. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Syposium on Quantum Biology and Quantum Pharmacology.

Anonymous:1988:M

[Ano88x] Anonymous. Masthead. International Journal of Quantum Chemistry, 34(S22):fmi, March 12–19, 1988. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Syposium on Quantum Chemistry, Solid-State Theory, and Computational Methods.

Anonymous:1988:PA

[Ano88y] Anonymous. Preliminary Announcement. International Journal of Quantum Chemistry, 33(6):581, June 1988. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1988:SA

[Ano88z] Anonymous. Second announcement. International Journal of Quantum Chemistry, 34(2):193, August 1988. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1989:Aa

[Ano89a] Anonymous. Announcement. International Journal of Quantum Chemistry, 35(1):253, January 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1989:Ab

[Ano89b] Anonymous. Announcement. International Journal of Quantum Chemistry, 35(1):255, January 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1989:Ac

[Ano89c] Anonymous. Announcement. International Journal of Quantum Chemistry, 35(2):359, February 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1989:Ad

[Ano89d] Anonymous. Announcement. International Journal of Quantum Chemistry, 35(3):457, March 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1989:Ae

[Ano89e] Anonymous. Announcement. International Journal of Quantum Chemistry, 35(3):459, March 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1989:Af

[Ano89f] Anonymous. Announcement. International Journal of Quantum Chemistry, 36(3):427, September 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1989:Ag

[Ano89g] Anonymous. Announcement. International Journal of Quantum Chemistry, 36(4):535, October 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1989:Ah

[Ano89h] Anonymous. Announcement. International Journal of Quantum Chemistry, 36(4):537, October 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1989:Ai

[Ano89i] Anonymous. Announcement. International Journal of Quantum Chemistry, 36(4):541, October 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1989:Aj

[Ano89j] Anonymous. Announcement. International Journal of Quantum Chemistry, 36(5):671–672, November 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1989:Ak

[Ano89k] Anonymous. Announcement. International Journal of Quantum Chemistry, 36(6):791, December 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1989:E

[Ano89l] Anonymous. Editorial. International Journal of Quantum Chemistry, 36(1):1–3, July 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1989:EP

[Ano89m] Anonymous. Editor’s preface. International Journal of Quantum Chemistry, 36(3):189–190, September 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1989:Ma

[Ano89n] Anonymous. Masthead. International Journal of Quantum Chemistry, 35(1):fmi–fmii, January 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1989:Mb

[Ano89o] Anonymous. Masthead. International Journal of Quantum Chemistry, 35(2):fmi–fmii, February 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1989:Mc

[Ano89p] Anonymous. Masthead. International Journal of Quantum Chemistry, 35(3):fmi–fmii, March 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1989:Md

[Ano89q] Anonymous. Masthead. International Journal of Quantum Chemistry, 35(4):fmi–fmii, April 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1989:Me

[Ano89r] Anonymous. Masthead. International Journal of Quantum Chemistry, 35(5):fmi–fmii, May 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1989:Mf

[Ano89s] Anonymous. Masthead. International Journal of Quantum Chemistry, 35(6):fmi–fmii, June 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1989:Mg

[Ano89t] Anonymous. Masthead. International Journal of Quantum Chemistry, 36(1):fmi–fmii, July 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1989:Mh

[Ano89u] Anonymous. Masthead. International Journal of Quantum Chemistry, 36(2):fmi–fmii, August 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1989:Mi

[Ano89v] Anonymous. Masthead. International Journal of Quantum Chemistry, 36(3):fmi–fmii, September 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1989:Mj

[Ano89w] Anonymous. Masthead. International Journal of Quantum Chemistry, 36(4):fmi–fmii, October 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1989:Mk

[Ano89x] Anonymous. Masthead. International Journal of Quantum Chemistry, 36(5):fmi–fmii, November 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1989:Ml

[Ano89y] Anonymous. Masthead. International Journal of Quantum Chemistry, 36(6):fmi–fmii, December 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1989:Mm

[Ano89z] Anonymous. Masthead. International Journal of Quantum Chemistry, 36(S16):ii, April 1–8, 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Anonymous:1989:Mn

[Ano89-27] Anonymous. Masthead. International Journal of Quantum Chemistry, 36(S23):i, April 1–8, 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Anonymous:1989:PPI

[Ano89-28] Anonymous. Portrait of Professor Ivar Waller. International Journal of Quantum Chemistry, 35(6):701, June 1989. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Anonymous:1989:PIW

[Ano89-29] Anonymous. Publications by Ivar Waller. International Journal of Quantum Chemistry, 35(6):709–710, June 1989. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Avery:1980:TIDa

[AØ80a] John Avery and Per-Johan Ørmen. Transferable integrals in a deformation density approach to crystal orbital calculations. IV. Evaluation of angular integrals by a vector-pairing method. International Journal of Quantum Chemistry, 18(4): 953–971, October 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Avery:1980:TIDb

[AØ80b] John Avery and Per-Johan Ørmen. Transferable integrals in a deformation density approach to crystal orbital calculations. V. Coupling coefficients for crystal symmetry functions. International Journal of Quantum Chemistry, 18(S14):629–636, March 10–20, 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory, Collision Phenomena, Quantum Statistics, and Computational Methods.

Avery:1982:TID

[AØC82] John Avery, Per-Johan Ørmen, and Ibha Chatterjee. Transferable integrals in a deformation density approach to crystal calculations. I. Crystal harmonics and their properties. International Journal of Quantum Chemistry, 21(3):515–542, March 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Avery:1981:TID

[AØM81] John Avery, Per-Johan Ørmen, and Donald Mullen. Transferable integrals in a deformation density approach to crystal orbital calculations. VII. Crystal field theory using measured densities and potentials. International Journal of Quantum Chemistry, 20(S15):477–486, March 8–14, 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory, Collision Phenomena, and Computational Quantum Chemistry.

Anmin:1989:VCV

[ARZ^{+}89] Tian Anmin, Liu Ruifeng, Sun Zemin, Zhou Xuefeng, and Yan Guosen. Variational calculations of vibrational and vibrationrotational properties of ozone. International Journal of Quantum Chemistry, 36(6):765–771, December 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Andzelm:1986:CCP

[AS86] J. Andzelm and D. R. Salahub. Chemisorption of CO on Pd(100): An lcgto-lsd cluster study. International Journal of Quantum Chemistry, 29(5):1091–1104, May 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Andrade:1989:SSD

[AS89] L. Andrade and M. Sánchez. Surface states in deformed and perturbed diatomic crystals. International Journal of Quantum Chemistry, 36(5):553–561, November 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Avery:1988:IMP

[ASL88] John Avery and Peter Sommer-Larsen. Iteration of the manyparticle Schrödinger equation in momentum space. International Journal of Quantum Chemistry, 34(S22):437–444, March 12–19, 1988. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Syposium on Quantum Chemistry, Solid-State Theory, and Computational Methods.

Almlof:1985:MPP

[AT85] Jan Almlöf and Peter R. Taylor. Molecular properties from perturbation theory: A Unified treatment of energy derivatives. International Journal of Quantum Chemistry, 27(6): 743–768, June 1985. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Au:1982:PTR

[Au82] C. K. Au. Perturbation theory via the Ricatti equation. International Journal of Quantum Chemistry, 21(1):179–190, January 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Austin:1984:PTS

[Aus84] E. J. Austin. Perturbation theory studies of the hydrogen Zeeman effect—preliminary results. International Journal of Quantum Chemistry, 26(S18):449–455, March 1–15, 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Atomic, Molecular and Solid-State Theory, and Computational Quantum Chemistry.

Avery:1984:MBS

[Ave84a] John Avery. A model for biological specificity. International Journal of Quantum Chemistry, 26(5):843–855, November 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Avery:1984:MPP

[Ave84b] John Avery. A model of the primary process in photosynthesis. International Journal of Quantum Chemistry, 26(5):917–931, November 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Avery:1984:UMR

[Ave84c] John Avery. Use of the S-matrix in the relativistic treatment of resonance energy transfer. International Journal of Quantum Chemistry, 25(1):79–96, January 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Avery:1986:MDH

[Ave86] John Avery. Many-dimensional hydrogenlike wave functions and the quantum mechanical many-body problem. International Journal of Quantum Chemistry, 30(S20):57–63, March 10–15, 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Syposium on Atomic, Molecular, and Solid-State Theory, Scattering Problems, Many Body Phenomena, and Computational Quantum Chemistry.

Avron:1982:BWF

[Avr82] J. E. Avron. Bender–Wu formulas and classical trajectories: Higher dimensions and degeneracies. International Journal of Quantum Chemistry, 21(1):119–124, January 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Avery:1982:AID

[AW82] John Avery and Zhen-Yi Wen. Angular integrations in mdimensional spaces and hyperspherical harmonics. International Journal of Quantum Chemistry, 22(4):717–738, October 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Avery:1983:MEM

[AW83] John Avery and Zhen-Yi Wen. Matrix elements of a manyparticle Hamiltonian evaluated using hyperspherical coordinates. International Journal of Quantum Chemistry, 24(S17): 201–206, March 3–12, 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-state Theory, Collision Phenomena, and Computational Quantum Chemistry.

Axe:1989:CAA

[AWCL89] Frank U. Axe, Ahmad Waleh, Lek Chantranupong, and Gilda H. Loew. A comparative analysis of the active site properties of the resting states of cytochrome c peroxidase, metmyoglobin, and catalase. International Journal of Quantum Chemistry, 35(1):181–191, January 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Adams:1983:ESV

[AYBP83] George F. Adams, David R. Yarkony, Rodney J. Bartlett, and George D. Purvis. Electronic structure and vertical excitation spectrum of methylene amidogen CH_{2}N. International Journal of Quantum Chemistry, 23(2):437–446, February 1983.

CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Ang:1981:RCE

[AYCD81] M. H. Ang, K. Yates, I. G. Csizmadia, and R. Daudel. Relationship of correlation energy and size. International Journal of Quantum Chemistry, 20(4):793–806, October 1981. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Avery:1984:FQM

[AZY84] John Avery and Wen Zhen-Yi. A formulation of the quantummechanical many-body problem in terms of hyperspherical coordinates. International Journal of Quantum Chemistry, 25 (6):1069–1087, June 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Baba-Ahmed:1983:IJU

[BAG83a] A. Baba-Ahmed and J. Gayoso. Integrals I_{rs} and J_{rs} used in the computation of magnetic quantities by means of the McWeeny coupled perturbation method. International Journal of Quantum Chemistry, 23(1):71–80, January 1983. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Bagayoko:1983:CGB

[Bag83b] D. Bagayoko. Contraction of Gaussian basis sets and the total energy of fcc copper. International Journal of Quantum Chemistry, 24(S17):527–535, March 3–12, 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-state Theory, Collision Phenomena, and Computational Quantum Chemistry.

Baker:1985:ESP

[Bak85] Jon Baker. EOM /Propagator ionization potentials: Effect of an extended operator manifold on the predicted valence-shell ionization spectra of ethane, ethylene, and acetylene. International Journal of Quantum Chemistry, 27(2):145–157, February 1985. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Balaji:1981:ECH

[Bal81a] V. N. Balaji. Energy calculations on helical polypeptides: Structure of polyglycine II. International Journal of Quantum Chemistry, 20(2):347–356, August 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Balasubramanian:1981:SSS

[Bal81b] K. Balasubramanian. Symmetry simplifications of space types in configuration interaction induced by orbital degeneracy. International Journal of Quantum Chemistry, 20(6):1255–1271, December 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Balasubramanian:1982:MCI

[Bal82a] K. Balasubramanian. Method for constructing isomerization reactions. International Journal of Quantum Chemistry, 22 (2):385–395, August 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Balasubramanian:1982:SCT

[Bal82b] K. Balasubramanian. Spectra of chemical trees. International Journal of Quantum Chemistry, 21(3):581–590, March 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Balasubramanian:1982:SGC

[Bal82c] K. Balasubramanian. Symmetry groups of chemical graphs. International Journal of Quantum Chemistry, 21(2):411–418, February 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Balasubramanian:1982:SOG

[Bal82d] K. Balasubramanian. Symmetry operators of generalized wreath products and their applications to chemical physics. International Journal of Quantum Chemistry, 22(5):1013– 1031, November 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Balasubramanian:1988:RCC

[Bal88] K. Balasubramanian. Relativistic CASSCF/CI calculations: Applications to transition metal dihydrides. International Journal of Quantum Chemistry, 34(S22):465–476, March 12– 19, 1988. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Syposium on Quantum Chemistry, Solid-State Theory, and Computational Methods.

Band:1982:IDC

[Ban82a] William Band. Introduction to Dr. Chen-Ning Yang. International Journal of Quantum Chemistry, 22(S16):19, March 1–13, 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Chemistry, Theory of Condensed Matter, and Propagator Methods in the Quantum Theory of Matter.

Band:1982:PIM

[Ban82b] William Band. Physical interpretation of the mathematical singularities in the continuum strain fields conventionally associated with crystal lattice dislocations. International Journal of Quantum Chemistry, 22(S16):125–136, March 1–13, 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Chemistry, Theory of Condensed Matter, and Propagator Methods in the Quantum Theory of Matter.

Banerjee:1988:DSS

[Ban88a] Ajit Banerjee. Discrete spatial symmetries: Redundant coordinates and search for stationary points in reduced spaces. International Journal of Quantum Chemistry, 33(1):27–40, January 1988. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Broer:1988:TLS

[BAN88b] R. Broer, G. Aissing, and W. C. Nieuwpoort. Titanium in large silicon clusters. International Journal of Quantum Chemistry, 34(S22):297–308, March 12–19, 1988. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Syposium on Quantum Chemistry, Solid-State Theory, and Computational Methods.

Broer:1986:HFC

[BANF86] R. Broer, G. Aissing, W. C. Nieuwpoort, and L. F. Feiner. Hartree–Fock cluster study of interstitial transition metals in silicon. International Journal of Quantum Chemistry, 29(5): 1059–1066, May 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Barone:1981:RBH

[Bar81] G. Barone. Relations between the hydration and the intermolecular or intramolecular solute–solute interactions. International Journal of Quantum Chemistry, 19(6):1137–1138, June 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Bartlett:1985:CIQ

[Bar85] Rodney J. Bartlett, editor. Comparison of ab initio quantum chemistry with experiment for small molecules: the state of the art. D. Reidel, Dordrecht, Boston, Lancaster, Tokyo, 1985. ISBN 90-277-2129-7. LCCN QD462.A1 C66 1985.

Bardo:1986:TCR

[Bar86]  Richard D. Bardo.  Theoretical calculations of rate-

determining steps for ignition of shocked, condensed nitromethane. International Journal of Quantum Chemistry, 30 (S20):455–469, March 10–15, 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Syposium on Atomic, Molecular, and Solid-State Theory, Scattering Problems, Many Body Phenomena, and Computational Quantum Chemistry.

Baykara:1986:HCD

[BASB86] N. A. Baykara, J. Andzelm, D. R. Salahub, and S. Z. Baykara. Hydrogen chemisorption on, and diffusion through, palladium clusters. International Journal of Quantum Chemistry, 29(4): 1025–1032, April 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Bates:1983:SR

[Bat83] David R. Bates. Scientific reminiscences. International Journal of Quantum Chemistry, 24(S17):5–32, March 3–12, 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-state Theory, Collision Phenomena, and Computational Quantum Chemistry.

Bauschlicher:1986:OTM

[Bau86] Charles W. Bauschlicher, Jr. OH-transition metal bonding. International Journal of Quantum Chemistry, 30(S20):563– 572, March 10–15, 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Syposium on Atomic, Molecular, and Solid-State Theory, Scattering Problems, Many Body Phenomena, and Computational Quantum Chemistry.

Bhattacharyya:1981:LBR

[BB81a] Kamal Bhattacharyya and Sankar Prasad Bhattacharyya. Limiting behavior of the ratio of < r_{12}^{−1} > and < r_{1}^{−1} > in the helium sequence. International Journal of Quantum Chemistry, 20(4):887–889, October 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Buchler:1981:PCP

[BB81b] J. Robert Buchler and Manuel Barranco. Phase-coexistence in partially degenerate nuclear matter. International Journal of Quantum Chemistry, 20(S15):381–384, March 8–14, 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory, Collision Phenomena, and Computational Quantum Chemistry.

Banerjee:1982:MIS

[BB82a] Manas Banerjee and Sankar Prasad Bhattacharyya. Molecular inversion in the n π* states: a theoretical investigation of some model systems in a semiempirical molecular orbital framework. International Journal of Quantum Chemistry, 22 (4):775–782, October 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Banerjee:1982:SOE

[BB82b] Manas Banerjee and Sankar Prasad Bhattacharyya. Scaled one-electron Hamiltonian model for open-shell LCAO–MO– SCF calculations: Comparison with the restricted open-shell method of Roothaan. International Journal of Quantum Chemistry, 21(5):905–915, May 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Balint:1983:CEM

[BB83] I. Bálint and M. I. Bán. Comments on energy minimization in subspaces of coordinates and on “chemical hysteresis”. International Journal of Quantum Chemistry, 24(2):161–167, August 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Balint:1984:PDR

[BB84] Imre Bálint and Miklós I. Bán. A procedure for determining reaction paths and saddle points. International Journal of Quantum Chemistry, 25(4):667–675, April 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Banerjee:1985:IAA

[BB85] Manas Banerjee and Sankar Prasad Bhattacharyya. The INDO/2-AHP (average hole potential) method for excited states: Comparison with the simple INDO/2-HP (hole potential) method. International Journal of Quantum Chemistry, 28(5):565–571, November 1985. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Banerjee:1987:MEF

[BB87] Manas Banerjee and Sankar Prasad Bhattacharyya. Molecules in an electric field: Some structural and spectroscopic consequences. International Journal of Quantum Chemistry, 32(1): 61–69, July 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Brandt:1982:PTS

[BBAT82] Richard B. Brandt, Mark E. Brandt, Michael E. April, and Colin Thomson. Potential transition-state inhibitors of glyoxalase. I. International Journal of Quantum Chemistry, 22(S9): 335–343, March 4–6, 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Barzaghi:1984:RCR

[BBGS84] M. Barzaghi, R. Bianchi, C. Gatti, and M. Simonetta. The ring closure reaction in 1,6-methano-[10]annulene. International Journal of Quantum Chemistry, 26(S18):433–438, March 1–15, 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Atomic, Molecular and Solid-State Theory, and Computational Quantum Chemistry.

Bacskay:1987:HLB

[BBH87] George B. Bacskay, Glenn Bryant, and Noel S. Hush. Hole localization and broken symmetry: a theoretical study of core electron ionization in the Li_{2} molecule. International Journal of Quantum Chemistry, 31(3):471–487, March 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Bonapasta:1981:PES

[BBL81] A. Amore Bonapasta, M. Bossa, and A. Lapiccirella. Potential energy surface of the system CO + O: An ab initio study on the O CO bond formation. International Journal of Quantum Chemistry, 19(6):1139–1145, June 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Barone:1984:GPU

[BBLR84] Vincenzo Barone, Nadia Bianchi, Francesco Lelj, and Nino Russo. Gas phase unimolecular 1,1-hydrogen elimination: Reaction mechanism and isotope effect. International Journal of Quantum Chemistry, 26(5):621–636, November 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Behi:1982:EDH

[BBMP82] Joy Behi, Stephen Bone, Hywel Morgan, and Ronald Pethig. Effect of deuterium-hydrogen exchange on the electrical conduction in lysozyme. International Journal of Quantum Chemistry, 22(S9):367–374, March 4–6, 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Brandas:1983:SEC

[BBO83] E. J. Brändas, G. L. Bendazzoli, and F. Ortolani. Size extensivity corrections to a multireference CI by means of a randomphase average procedure. International Journal of Quantum Chemistry, 24(S17):321–329, March 3–12, 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-state Theory, Collision Phenomena, and Computational Quantum Chemistry.

Blomberg:1983:CCC

[BBPS83] M. Blomberg, U. Brandemark, L. Pettersson, and P. Siegbahn. Contracted CI calculations of models for catalytic reactions involving transition metals. International Journal of Quantum Chemistry, 23(3):855–863, March 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Bonchev:1982:GCC

[BBR82] Danail Bonchev, Alexandru T. Balaban, and Milan Randić. The graph center concept for polycyclic graphs. International Journal of Quantum Chemistry, 22(2):441, August 1982. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Boudreaux:1980:QMC

[BC80] E. A. Boudreaux and T. P. Carsey. Quasirelativistic MO calculations on platinum complexes (anticancer drugs) and their interaction with DNA. International Journal of Quantum Chemistry, 18(2):469–479, August 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Brown:1983:SCS

[BC83] Richard E. Brown and Johannes Pieter Colpa. SCF and CI Studies of Hund’s rules for the effects of electronic correlation and delocalization. I. International Journal of Quantum Chemistry, 24(6):593–602, December 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Brown:1984:GNM

[BC84] Robert G. Brown and Mikael Ciftan. A generalized nonmuffin-tin theory of band structure. International Journal of Quantum Chemistry, 26(S18):87–104, March 1–15, 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Atomic, Molecular and Solid-State Theory, and Computational Quantum Chemistry.

Brown:1985:EGN

[BC85a] Robert G. Brown and Mikael Ciftan. Erratum: A generalized non-muffin-tin theory of band structure. International Journal of Quantum Chemistry, 28(S19):751–752, March 18–23, 1985. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). See [BC85b]. Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Brown:1985:GNM

[BC85b] Robert G. Brown and Mikael Ciftan. A generalized nonmuffin-tin theory of band structure. International Journal of Quantum Chemistry, 27(6):803–804, June 1985. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). See erratum [BC85a].

Beck:1987:EAS

[BCA87] Donald R. Beck, Ziyong Cai, and George Aspromallis. Electron affinities of states of Sc and Cu. International Journal of Quantum Chemistry, 32(S21):457–468, March 12, 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Chemistry, Solid-State Theory, and Computational Methods.

Brener:1985:ESS

[BCB85] N. E. Brener, J. Callaway, and P. Blaha. Electronic structure of small sodium clusters. International Journal of Quantum Chemistry, 28(S19):603–612, March 18–23, 1985. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Bredas:1981:NSE

[BCBS81] J. L. Brédas, R. R. Chance, R. H. Baughman, and R. Silbey. Nonempirical studies of the electronic properties of highly conducting polymers. International Journal of Quantum Chemistry, 20(S15):231–241, March 8–14, 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory, Collision Phenomena, and Computational Quantum Chemistry.

Brandas:1982:SCC

[BCC82] Erkki J. Brändas, Leon L. Combs, and Nestor S. Correia. Size consistency corrections for configurational interaction calculations. International Journal of Quantum Chemistry, 21(1):

259–267, January 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Borodavkin:1980:AUS

[BCD^{+}80] A. V. Borodavkin, V. O. Chekhov, Yu. S. Dolin, Yu. V. Morozov, F. A. Savin, E. I. Budowsky, and D. Yu. Yakovlev. Absorption UV spectroscopy and electronic structure of ionic and tautomeric forms of hydroxy and methoxy derivatives of cytosine and adenine, and of some 5-substituted analogs of pyrimidines. International Journal of Quantum Chemistry, 17(4):803–811, April 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Brandas:1989:SAI

[BCD89] E. J. Brändas and C. A. Chatzidimitriou-Dreismann. Some aspects on ionic conductance and proton transfer reactions in aqueous solutions. International Journal of Quantum Chemistry, 36(S23):147–151, April 1–8, 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Bonaccorsi:1986:CCC

[BCT86] Rosanna Bonaccorsi, Roberto Cammi, and Jacopo Tomasi. Counterpoise corrections to the components of bimolecular energy interactions: An examination of three methods of decomposition. International Journal of Quantum Chemistry, 29 (3):373–378, March 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Broch:1980:QMS

[BCV80] H. Broch, D. Cabrol, and D. Vasilescu. Quantum mechanical simulation of the interaction between the radioprotector cysteamine and DNA. International Journal of Quantum Chemistry, 18(S7):283–295, March 5–8, 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Broch:1981:IAC

[BCV81] H. Broch, D. Cabrol, and D. Vasilescu. Implications of aminoacyladenylate conformations in the prebiotic polypeptide synthesis. International Journal of Quantum Chemistry, 20(S8):139–148, March 5–7, 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Broch:1982:EPS

[BCV82] H. Broch, D. Cabrol, and D. Vasilescu. Electrostatic properties of some sulfur-containing radioprotectors. International Journal of Quantum Chemistry, 22(S9):111–123, March 4–6, 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Bamzai:1981:ISC

[BD81] Anjuli S. Bamzai and B. M. Deb. Internal stresses and chemical binding in the H molecule. International Journal of Quantum Chemistry, 20(6):1315–1329, December 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Bickerstaff:1985:NSC

[BD85] R. P. Bickerstaff and T. Damhus. A necessary and sufficient condition for the existence of real coupling coefficients for a finite group. International Journal of Quantum Chemistry, 27 (4):381–391, April 1985. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Baykara:1988:ECS

[BD88] N. A. Baykara and M. Demiralp. Eigenlevel calculations for screened Coulomb potential systems via an integrationfree algorithm: Space pruning technique. International Journal of Quantum Chemistry, 34(3):231–245, September 1988. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Bozovic:1981:PSA

[BDD81] I. B. Božović, J. Delhalle, and M. Damnjanović. A posteriori symmetry analysis of polymer electronic band structures. International Journal of Quantum Chemistry, 20(6):1143–1163, December 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Boudreaux:1986:NSC

[BDK86] Edward A. Boudreaux, Stephen P. Doussa, and Mariusz Klobukowski. Nonempirical self-consistent modified extended Hückel calculations on heavy-metal systems. II. Electronic structure, bonding, and spectra of the binuclear Pt_{2}(P_{2}O_{5}H_{2}) ion. International Journal of Quantum Chemistry, 30(S20): 239–252, March 10–15, 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Syposium on Atomic, Molecular, and Solid-State Theory, Scattering Problems, Many Body Phenomena, and Computational Quantum Chemistry.

Bertholon:1981:GCB

[BDP^{+}81] G. Bertholon, C. Decoret, M. Perrin, J. Royer, A. Thozet, and B. Tinland. Geometries of the chemical bonds around the hydroxyl group in a series of phenolic compounds. International Journal of Quantum Chemistry, 19(6):1167–1170, June 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Bayard:1988:SPR

[BDR^{+}88] F. Bayard, C. Decoret, J. Royer, Y. G. Smeyers, and J. J. Randez. Study of a possible reaction pathway in the polymorphic transformation of diacetamide. International Journal of Quantum Chemistry, 33(1):61–68, January 1988. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Beck:1981:ILC

[Bec81] Donald R. Beck. On the incorporation of local correlation effects into ionic solids. International Journal of Quantum Chemistry, 20(S15):521–538, March 8–14, 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory, Collision Phenomena, and Computational Quantum Chemistry.

Beck:1982:SNI

[Bec82] Donald R. Beck. On the spectrum of negative ions and the charge density of the ground state of neutral neon. International Journal of Quantum Chemistry, 22(S16):345–355, March 1–13, 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Chemistry, Theory of Condensed Matter, and Propagator Methods in the Quantum Theory of Matter.

Becke:1983:HFE

[Bec83] A. D. Becke. Hartree–Fock exchange energy of an inhomogeneous electron gas. International Journal of Quantum Chemistry, 23(6):1915–1922, June 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Becke:1985:LEC

[Bec85] A. D. Becke. Local exchange-correlation approximations and first-row molecular dissociation energies. International Journal of Quantum Chemistry, 27(5):585–594, May 1985. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Becke:1989:BSF

[Bec89] Axel D. Becke. Basis-set-free density-functional quantum chemistry. International Journal of Quantum Chemistry, 36 (S23):599–609, April 1–8, 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Bedian:1984:OGC

[Bed84] Vahe Bedian. The origin of the genetic code. International Journal of Quantum Chemistry, 26(S11):87–89, March 12–15, 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Belinfante:1980:DMF

[Bel80a] Frederik J. Belinfante. Density matrix formulation of quantum theory and its physical interpretation. International Journal of Quantum Chemistry, 17(1):1–24, January 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Bell:1980:BBD

[Bel80b] J. S. Bell. De Broglie–Bohm, delayed-choice, double-slit experiment, and density matrix. International Journal of Quantum Chemistry, 18(S14):155–159, March 10–20, 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory, Collision Phenomena, Quantum Statistics, and Computational Methods.

Berrondo:1989:EPM

[BEL89] M. Berrondo, S. W. Eggleston, and E. G. Larson. The effect of penetration on the multipole and multistreak expansions of molecular electrostatic potentials. International Journal of Quantum Chemistry, 36(6):749–764, December 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Bennett:1981:MCE

[Ben81] H. Stanley Bennett. Muscle contraction and enzymatic action: Challenges to quantum chemistry. International Journal of Quantum Chemistry, 20(S8):317–329, March 5–7, 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Bender:1982:PTL

[Ben82] Carl M. Bender. Perturbation theory in large order for some elementary systems. International Journal of Quantum Chemistry, 21(1):93–104, January 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Bendazzoli:1980:PAC

[BEO80] G. L. Bendazzoli, S. Evangelisti, and F. Ortolani. Padé algorithm for the computation of static and dynamic second-order properties by large-scale CI expansion. International Journal of Quantum Chemistry, 18(6):1393–1404, December 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Bone:1981:EPP

[BEP81] Stephen Bone, Joyce Eden, and Ronald Pethig. Electrical properties of proteins as a function of hydration and NaCl content. International Journal of Quantum Chemistry, 20(S8): 307–316, March 5–7, 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Bendazzoli:1987:SITa

[BEP87a] G. L. Bendazzoli, S. Evangelisti, and P. Palmieri. Some iterative techniques for large CI matrices. I. Excited eigenvalues. International Journal of Quantum Chemistry, 31(4):663–672, April 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Bendazzoli:1987:SITb

[BEP87b] G. L. Bendazzoli, S. Evangelisti, and P. Palmieri. Some iterative techniques for large CI matrices. II. Singular systems of linear equations. International Journal of Quantum Chemistry, 31(4):673–684, April 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Berthier:1981:CSC

[Ber81] G. Berthier. Comments on the status of the chemical bond in methods of quantum chemistry. International Journal of Quantum Chemistry, 19(6):985–989, June 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Berggren:1988:QPS

[Ber88] K.-F. Berggren. Quantum phenomena in small semiconductor structures and devices. International Journal of Quantum Chemistry, 33(3):217–245, March 1988. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Brandas:1985:IPQ

[BERE85] Erkki Brändas, Erik Engdahl, Magnus Rittby, and Nils Elander. On the inverse problem in quantum scattering theory. International Journal of Quantum Chemistry, 28(S19): 457–466, March 18–23, 1985. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Besnainou:1988:NDM

[Bes88] S. Besnainou. Nonlinear dynamics of a model hydrogen bond. International Journal of Quantum Chemistry, 34(2):143–160, August 1988. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Brescansin:1981:MXT

[BF81] Luiz M. Brescansin and A. Fazzio. MS–Xα: Treatment for native defects in GaSb. International Journal of Quantum Chemistry, 20(S15):457, March 8–14, 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory, Collision Phenomena, and Computational Quantum Chemistry.

Binkley:1983:IDB

[BF83a] J. S. Binkley and M. J. Frisch. ab initio determination of bond dissociation energies: The first-row diatomics CO, N_{2}, NO, O_{2}, and F_{2}. International Journal of Quantum Chemistry, 24 (S17):331–337, March 3–12, 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-state Theory, Collision Phenomena, and Computational Quantum Chemistry.

Brandas:1983:RDC

[BF83b] Erkki Brändas and Piotr Froelich. Rigorous determination of cross sections via complex transformations. International Journal of Quantum Chemistry, 24(S17):113–116, March 3– 12, 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-state Theory, Collision Phenomena, and Computational Quantum Chemistry.

Brener:1980:EPS

[BFC80] N. E. Brener, J. L. Fry, and T. C. Collins. Excitations of the N-particle system in the generalized random phase approximation. International Journal of Quantum Chemistry, 18(S14):455–460, March 10–20, 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory, Collision Phenomena, Quantum Statistics, and Computational Methods.

Brener:1981:MEA

[BFJ81] N. E. Brener, J. L. Fry, and R. A. Johnson. Matrix elements in the analytic tetrahedron method. International Journal of Quantum Chemistry, 20(S15):459–468, March 8–14, 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory, Collision Phenomena, and Computational Quantum Chemistry.

Bridet:1983:ICH

[BFOP83] Jacques Bridet, Sándor Fliszár, Simone Odiot, and Robert Pick. Ab initio calculation of the heptamer (NH_{3})_{7} as a reasonable starting point for a description of the ammonia crystal. International Journal of Quantum Chemistry, 24(6):687–695, December 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Bansal:1981:SFC

[BG81] Manju Bansal and Goutam Gupta. Structure factor calculations of various DNA duplexes. International Journal of Quantum Chemistry, 20(2):407–417, August 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Baretty:1988:MBP

[BG88] Reinaldo Baretty and Carmelo Garcia. Modified Breit– Pauli Hamiltonian suitable for variational calculations. International Journal of Quantum Chemistry, 34(S22):425–430, March 12–19, 1988. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Syposium on Quantum Chemistry, Solid-State Theory, and Computational Methods.

Bonaccorsi:1984:SII

[BGT84] Rosanna Bonaccorsi, Caterina Ghio, and Jacopo Tomasi. On a semiclassical interpretation of inter- and intramolecular interactions. International Journal of Quantum Chemistry, 26(5): 637–686, November 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Birner:1982:CEA

[BH82] P. Birner and H.-J. Hofmann. Critical examination of approximate LCAO MO methods. I. Reasons for the failure of the CNDO and INDO methods in theoretical conformational analysis of conjugated compounds. International Journal of Quantum Chemistry, 21(5):833–843, May 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Bessis:1986:SCU

[BH86] Daniel Bessis and Carlos R. Handy. Systematic construction of upper and lower bounds to the ground state energy of the Schrödinger equation. International Journal of Quantum Chemistry, 30(S20):21–32, March 10–15, 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Syposium on Atomic, Molecular, and Solid-State Theory, Scattering Problems, Many Body Phenomena, and Computational Quantum Chemistry.

Bichsel:1989:ESD

[BH89a] Hans Bichsel and T. Hiraoka. Energy spectra and depth-dose curves for 70 MeV protons. International Journal of Quantum Chemistry, 36(S23):565–574, April 1–8, 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Bouman:1989:LRC

[BH89b] Thomas D. Bouman and Aage E. Hansen. Linear response calculations of molecular optical and magnetic properties using program RPAC: NMR shielding tensors of pyridine and nazines. International Journal of Quantum Chemistry, 36(S23): 381–396, April 1–8, 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Bhattacharjee:1980:VFP

[Bha80] Ravi Shankar Bhattacharjee. Variational function of the Padé type for the ground state of the x^{4} anharmonic oscillator. International Journal of Quantum Chemistry, 18(6):1485–1487, December 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Bhattacharyya:1981:VRS

[Bha81a] Kamal Bhattacharyya. Variants of Rayleigh–Schrödinger perturbation theory—a new look. International Journal of Quantum Chemistry, 20(6):1273–1284, December 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Bhattacharyya:1981:SOE

[Bha81b] Sankar Prasad Bhattacharyya. A scaled “one-electron Hamiltonian” model for open-shell LCAO–MO–SCF calculations. International Journal of Quantum Chemistry, 19(5):735–743, May 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Bhattacharyya:1982:FSH

[Bha82a] Kamal Bhattacharyya. Feenberg scaling and higher-order invariants in Rayleigh–Schrödinger perturbation theory. International Journal of Quantum Chemistry, 21(5):857–860, May 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Bhattacharyya:1982:GET

[Bha82b] Kamal Bhattacharyya. Generalized Euler transformation in extracting useful information from divergent (asymptotic) perturbation series and the construction of Padé approximants. International Journal of Quantum Chemistry, 22(2):307–330, August 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Bhattacharyya:1982:SMO

[Bha82c] Sankar Prasad Bhattacharyya. Semiempirical molecular orbital calculations on octahedral aquo complexes of transition metal ions: σ-donor ability of water. International Journal of Quantum Chemistry, 21(5):851–855, May 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Bessis:1987:EPM

[BHM87] Daniel Bessis, Carlos R. Handy, and T. Morley. An exact projection method for the lowest eigenstate of singular multidimensional Schrödinger equations. International Journal of Quantum Chemistry, 32(S21):755, March 12, 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Chemistry, Solid-State Theory, and Computational Methods.

Brickmann:1985:MSN

[BHS85] Jürgen Brickmann, Paul Hofmann, and Peter C. Schmidt. Measures of stochasticity for nonstationary initial preparations of coupled oscillators. International Journal of Quantum Chemistry, 28(1):109–134, July 1985. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Bouman:1983:LSRa

[BHVR83a] Thomas D. Bouman, Aage E. Hansen, Bjørn Voigt, and Sten Rettrup. Large-scale RPA calculations of chiroptical properties of organic molecules: Program RPAC. International Journal of Quantum Chemistry, 23(2):595–611, February 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Bouman:1983:LSRb

[BHVR83b] Thomas D. Bouman, Aage E. Hansen, Bjørn Voigt, and Sten Rettrup. Large-scale RPA calculations of chiroptical properties of organic molecules: Program RPAC. International Journal of Quantum Chemistry, 23(4):1693, April 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Bigelow:1986:CCR

[Big86] Richard W. Bigelow. Correlation contributions to the radical cation states of linear even polyenes: An open-shell RHF-CNDO/S(CI) analysis of the optical and photoelectron spectra. International Journal of Quantum Chemistry, 29(1):35– 60, January 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Baretty:1986:MSA

[BIN86] Reinaldo Baretty, Yasuyuki Ishikawa, and Jose F. Nieves. Momentum space approach to the relativistic atomic structure calculations. International Journal of Quantum Chemistry, 30 (S20):109–117, March 10–15, 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Syposium on Atomic, Molecular, and Solid-State Theory, Scattering Problems, Many Body Phenomena, and Computational Quantum Chemistry.

Banerjee:1987:DSS

[BJ87] Ajit Banerjee and James O. Jensen. Discrete spatial symmetries in energy derivatives. International Journal of Quantum Chemistry, 31(1):3–16, January 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Bowen-Jenkins:1986:QMS

[BJR86] Philippa E. Bowen-Jenkins and W. Graham Richards. Quantitative measures of similarity between pharmacologically active compounds. International Journal of Quantum Chemistry, 30 (6):763–768, December 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Biskupic:1982:CEH

[BK82] Stanislav Biskupič and Vladimír Kvasnička. Construction of an effective Hamiltonian for open-shell molecular systems. International Journal of Quantum Chemistry, 21(3):633–645, March 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Beznosjuk:1984:DFT

[BK84a] Serge A. Beznosjuk and Eugene S. Kryachko. Density functional theory: Approximating quasiparticle density functional. International Journal of Quantum Chemistry, 25(4):645–651, April 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Bohm:1984:ILM

[BK84b] Stanislav Böhm and Josef Kuthan. An ab initio LCAO–MO study of the substituent effect in benzenoid systems: Metaand para-substituted benzoic acids. International Journal of Quantum Chemistry, 26(1):21–33, July 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Biskupic:1989:MBD

[BK89] Stanislav Biskupič and Vladimír Kvasnička. A many-body diagrammatic theory of rotation–vibration spectra. International Journal of Quantum Chemistry, 35(4):495–512, April 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Burke:1988:NDP

[BKJ88] Luke A. Burke and Karsten Krogh-Jespersen. The nonplanarity of n-doped polyenes and of carbanions with alkali metal counterions. International Journal of Quantum Chemistry, 34 (S22):51–57, March 12–19, 1988. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Syposium on Quantum Chemistry, Solid-State Theory, and Computational Methods.

Bonacic-Koutecky:1983:ACM

[BKP83] V. Bonačić-Koutecký and M. Persico. Avoided crossing of molecular excited states and photochemistry: Butadiene and unprotonated Schiff base. International Journal of Quantum Chemistry, 23(2):517–533, February 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Basch:1987:CBE

[BKS87] H. Basch, M. Krauss, and W. J. Stevens. Cation binding effect on imidazole tautomerism. International Journal of Quantum Chemistry, 31(3):405–415, March 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Bauschlicher:1981:ICF

[BL81a] Charles W. Bauschlicher, Jr. and Byron H. Lengsfield III. Improved CI formula evaluation: Elimination of the phase determination at the determinant level for arbitrarily coupled CSFs. International Journal of Quantum Chemistry, 19(4): 649–654, April 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Biczo:1981:HSE

[BL81b] G. Biczó and I. Lukovits. Hunting of strange electronic states and structures in solids and biopolymers. International Journal of Quantum Chemistry, 20(S8):347–357, March 5–7, 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Blake:1981:MCC

[BL81c] I. Ortega Blake and A. Leś. Magnesium and calcium cation–ligand interactions within the pseudopotential approach. II. Cation–GABA interactions. International Journal of Quantum Chemistry, 19(3):463–475, March 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Balasubramanian:1988:SCP

[BL88] K. Balasubramanian and Xiaoyu Liu. Spectra and characteristic polynomials of polyhedral clusters. International Journal of Quantum Chemistry, 34(S22):319–328, March 12–19, 1988.

CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Syposium on Quantum Chemistry, Solid-State Theory, and Computational Methods.

Berrondo:1989:LE

[BL89a] M. Berrondo and E. G. Larson. Letter to the editor. International Journal of Quantum Chemistry, 36(S16):343–344, April 1–8, 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Broo:1989:CET

[BL89b] Anders Broo and Sven Larsson. Calculation of electron transfer rates in proteins. International Journal of Quantum Chemistry, 36(S16):185–198, April 1–8, 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Bleick:1982:NCT

[Ble82] W. E. Bleick. New coordinates for three-body problems. International Journal of Quantum Chemistry, 22(S16):241–245, March 1–13, 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Chemistry, Theory of Condensed Matter, and Propagator Methods in the Quantum Theory of Matter.

Battaglia:1985:OTE

[BLG85] Franco Battaglia, K. C. Liu, and Thomas F. George. One- and two-electron transfer processes in ion-surface scattering. International Journal of Quantum Chemistry, 28(S19):477–489, March 18–23, 1985. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Blinder:1980:AFN

[Bli80] S. M. Blinder. Asymptotic forms for the nonrelativistic Coulomb propagator. International Journal of Quantum Chemistry, 18(S14):43–51, March 10–20, 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory, Collision Phenomena, Quantum Statistics, and Computational Methods.

Blinder:1984:GFP

[Bli84] S. M. Blinder. On Green’s functions, propagators, and Sturmians for the nonrelativistic Coulomb problem. International Journal of Quantum Chemistry, 26(S18):293–307, March 1– 15, 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Atomic, Molecular and Solid-State Theory, and Computational Quantum Chemistry.

Blomberg:1983:TAA

[Blo83a] Clas Blomberg. Thermodynamic aspects on accuracy in the synthesis of biomolecules. International Journal of Quantum Chemistry, 23(2):687–707, February 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Bloor:1983:OSC

[Blo83b] J. E. Bloor. Overlapping spheres continuum multiplescattering Xα calculations of elastic cross sections in electronmolecule scattering. International Journal of Quantum Chemistry, 24(S17):101–112, March 3–12, 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-state Theory, Collision Phenomena, and Computational Quantum Chemistry.

Bloor:1989:ICH

[Blo89] John E. Bloor. Ab initio calculation of harmonic vibrational frequencies and intensities of molecules containing atoms of the third period of the periodic table. I. hydrogen sulfide and sulfur dioxide. International Journal of Quantum Chemistry, 36(S23):187–197, April 1–8, 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Brauchler:1989:HBP

[BLOW89] M. Bräuchler, S. Lunell, I. Olovsson, and W. Weyrich. Hydrogen bonding in potassium hydrogen carbonate studied by Compton scattering on single crystals. International Journal of Quantum Chemistry, 35(6):895–917, June 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Barone:1981:NAU

[BLR81] Vincenzo Barone, Francesco Lelj, and Nino Russo. Nonempirical analysis of unusual chemical bonds. II. AlH_{2}BH_{4} and AlH_{2}C_{3}H_{5}. International Journal of Quantum Chemistry, 19 (6):1197–1201, June 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Bertran:1983:WCI

[BLR83] J. Bertrán, A. Lledós, and J. A. Revetllat. Water chain intervention in hydrolytic and tautomeric reactions. International Journal of Quantum Chemistry, 23(2):587–593, February 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Barone:1986:CBA

[BLR86] Vincenzo Barone, Francesco Lelj, and Nino Russo. Conformational behavior of azabiphenyls. International Journal of Quantum Chemistry, 29(3):541–551, March 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Brandt:1984:IGV

[BLT^{+}84] Richard B. Brandt, Jerome Laux, Colin Thomson, Morris A. Johnson, and Marvin Gross. Inhibition of glyoxalase I in vitro by flavones. International Journal of Quantum Chemistry, 26 (S11):195–200, March 12–15, 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Broeckhove:1988:GCA

[BLV88] J. Broeckhove, L. Lathouwers, and G. Vanderheyden. Generator coordinate approximation to nuclear electronic coupling. International Journal of Quantum Chemistry, 34(S22):481– 490, March 12–19, 1988. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Syposium on Quantum Chemistry, Solid-State Theory, and Computational Methods.

Brandt:1986:IGV

[BLY^{+}86] Richard B. Brandt, Jerome E. Laux, Steven W. Yates, Marvin R. Boots, Colin Thomson, and Colin Edge. Inhibition of glyoxalase I in vitro by coumarin and coumarin derivatives. International Journal of Quantum Chemistry, 30(S13):155–165, March 6–8, 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Bhatia:1986:WLO

[BM86] Rajendra Bhatia and Kalyan K. Mukherjea. On weighted Löwdin orthogonalization. International Journal of Quantum Chemistry, 29(6):1775–1778, June 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Billing:1989:CPA

[BM89] G. D. Billing and J. T. Muckerman. A classical path approach to reactive scattering. International Journal of Quantum Chemistry, 36(2):179–186, August 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Brown:1987:ISH

[BMB87] Richard E. Brown, G. David Mendenhall, and Rodney J. Bartlett. Ab initio studies of hyponitrous acid. International Journal of Quantum Chemistry, 32(S21):603–612, March 12, 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Chemistry, Solid-State Theory, and Computational Methods.

Bhattacharyya:1982:CBS

[BMBK82] Sankar Prasad Bhattacharyya, Debashis Mukherjee, Debendra Madhab Bhattacharyya, and Priyotosh Khan. Convergence behavior of some orbital optimization procedures. International Journal of Quantum Chemistry, 22(4):753–759, October 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Bakale:1982:QEA

[BMG82] George Bakale, Richard D. McCreary, and Earle C. Gregg. Quasifree electron attachment to carcinogens. International Journal of Quantum Chemistry, 22(S9):15–25, March 4–6, 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Bertoluzza:1981:CBS

[BMMS81] A. Bertoluzza, A. Marinangeli, M. A. Morelli, and R. Simoni. Correlations between structure and properties of materials. X. Anomalous optical properties and Raman spectra of glassy ternary B_{2}O_{3}BaTiO_{3} − Na_{2}O materials. International Journal of Quantum Chemistry, 19(6):1161–1165, June 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Beveridge:1986:FES

[BMRJ86] D. L. Beveridge, M. Mezei, G. Ravishanker, and B. Jayaram. Free energy simulations: Applications to the study of liquid water, hydrophobic interactions and solvent effects on conformational stability. International Journal of Quantum Chemistry, 29(5):1513–1523, May 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Bonchev:1980:TCC

[BMT80] D. Bonchev, Ov. Mekenyan, and N. Trinajstić. Topological characterization of cyclic structures. International Journal of Quantum Chemistry, 17(5):845–893, May 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Beck:1980:TIC

[BN80] D. R. Beck and C. A. Nicolaides. On the theoretical interpretation and calculation of inner-electron photoemission spectra in atoms and solids. International Journal of Quantum Chemistry, 18(S14):323–333, March 10–20, 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory, Collision Phenomena, Quantum Statistics, and Computational Methods.

Beck:1984:FHS

[BN84] Donald R. Beck and C. A. Nicolaides. Fine and hyperfine structure of the two lowest bound states of Be^{−} and their first two ionization thresholds. International Journal of Quantum Chemistry, 26(S18):467–481, March 1–15, 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Atomic, Molecular and Solid-State Theory, and Computational Quantum Chemistry.

Baer:1986:QIO

[BNK86] Michael Baer, Hiroki Nakamura, and Donald J. Kouri. Quantum infinite order sudden approximation for ion-molecule reactions: Treatment of the He + H system. International Journal of Quantum Chemistry, 30(S20):483–493, March 10–15, 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Syposium on Atomic, Molecular, and Solid-State Theory, Scattering Problems, Many Body Phenomena, and Computational Quantum Chemistry.

Bogdanov:1987:UWI

[BNS^{+}87] B. Bogdanov, S. Nikolić, A. Sabljić, N. Trinajstić, and S. Carter. On the use of the weighted identification numbers in the QSAR study of the toxicity of aliphatic ethers. International Journal of Quantum Chemistry, 32(S14):325–330, March 14, 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Boca:1987:IRE

[Boč87] Roman Boča. Inclusion of relativistic effects into ZDO methods. I. A quasi-relativistic CNDO/1. International Journal of Quantum Chemistry, 31(6):941–950, June 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Boca:1988:IREa

[Boč88a] Roman Boča. Inclusion of relativistic effects into ZDO methods. II. Solavation of metal complexes. International Journal of Quantum Chemistry, 33(2):159–167, February 1988. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Boca:1988:IREb

[Boč88b] Roman Boča. Inclusion of relativistic effects into ZDO methods. III. A. Quasi-relativistic INDO/1 version. International Journal of Quantum Chemistry, 34(4):385–399, October 1988. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Boca:1989:IRE

[Boč89] Roman Boča. Inclusion of relativistic effects into ZDO methods. V. Effect of core and polarization functions in bonding of first-group elements. International Journal of Quantum Chemistry, 36(6):727–739, December 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Boettger:1988:TIS

[Boe88] J. C. Boettger. Theoretical investigation of a sodium overlayer on the (111) surface of aluminum. International Journal of Quantum Chemistry, 34(S22):737–738, March 12–19, 1988. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Syposium on Quantum Chemistry, Solid-State Theory, and Computational Methods.

Bohm:1982:REP

[Böh82] Michael C. Böhm. Reorganization effects in the photoelectron spectra of transition metal compounds: A Green’s function study in the framework of a variable INDO Hamiltonian. International Journal of Quantum Chemistry, 22(5):939–970, November 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Bohm:1983:SBH

[Böh83] Michael C. Böhm. Symmetry breaking in the Hartree–Fock approximation for binuclear transition metal compounds—a theoretical investigation based on a variable model operator. International Journal of Quantum Chemistry, 24(2):185–237, August 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Bohm:1984:BSP

[Böh84] Michael C. Böhm. The band structure of porphyrinatonickel(II). A semiempirical crystal orbital study based on the tight-binding formalism. International Journal of Quantum Chemistry, 25(5):817–851, May 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Bohm:1985:TIE

[Böh85] Michael C. Böhm. A theoretical investigation of electron correlation and relaxation in organometallic polymers. International Journal of Quantum Chemistry, 27(3):323–374, March 1985. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Bonchev:1981:ITI

[Bon81] D. Bonchev. Information theory interpretation of the Pauli principle and Hund rule. International Journal of Quantum Chemistry, 19(4):673–679, April 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Bonchev:1983:ITI

[Bon83] Danail Bonchev. Information theoretic indices for characterization of chemical structures, volume 5 of Chemometrics series. Research Studies Press, Chichester, West Sussex, UK, 1983. ISBN 0-471-90087-7. xiv + 249 pp. LCCN QD471 .B76 1983. US$59.00.

Borello:1981:MCS

[Bor81] E. Borello. Metal clusters and surfaces. International Journal of Quantum Chemistry, 19(6):1065–1081, June 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Born:1982:OME

[Bor82] Gregory Born. U(n) operator manifolds for electron propagator applications. International Journal of Quantum Chemistry, 22(S16):633–639, March 1–13, 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Chemistry, Theory of Condensed Matter, and Propagator Methods in the Quantum Theory of Matter.

Born:1985:CEP

[Bor85] Gregory Born. Correlated electron propagator theory and applications for open-shell atoms and molecules. International Journal of Quantum Chemistry, 28(3):335–348, September 1985. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Bottcher:1981:NST

[Bot81] C. Bottcher. Numerical solution of the time-dependent Schrödinger equation with Coulomb interactions. International Journal of Quantum Chemistry, 20(S15):671–682, March 8–14, 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory, Collision Phenomena, and Computational Quantum Chemistry.

Boucekkine:1980:CCD

[Bou80] A. Boucekkine. CNDO calculation of the diamagnetic susceptibilities of cyclic conjugated molecules. International Journal of Quantum Chemistry, 18(1):193–197, July 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Bourg:1984:PRT

[Bou84] M. Bourg. Preliminary results of a theoretical ab initio model study of adsorption on ionic crystals. International Journal of Quantum Chemistry, 26(5):775–781, November 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Bulski:1980:DNB

[BOW80] T. Bulski, S. Olszewski, and A. Wierzbicki. Density of non-Bloch electron states in perfect cubic crystals. International Journal of Quantum Chemistry, 17(4):697–718, April 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Bowman:1986:RDR

[Bow86] Joel M. Bowman. Rotational distributions from resonances in H + H_{2}. International Journal of Quantum Chemistry, 30 (S20):681–687, March 10–15, 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Syposium on Atomic, Molecular, and Solid-State Theory, Scattering Problems, Many Body Phenomena, and Computational Quantum Chemistry.

Boca:1980:MPM

[BP80] R. Boča and P. Pelikán. Modified PCILO method. II. Secondorder energy and geometry of transition-metal halides. International Journal of Quantum Chemistry, 18(6):1361–1370, December 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Bone:1983:CMS

[BP83] Stephen Bone and Ronald Pethig. Cyclodextrins as model systems for the study of proton transport. International Journal of Quantum Chemistry, 24(S10):133–141, March 14–16, 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Bauer:1984:EEN

[BP84] Joanne Bauer, Sr. and Peter Politzer. The effect of an epoxide– nucleophile reaction upon hydrogen bonding involving the nucleophile. International Journal of Quantum Chemistry, 25 (5):869–879, May 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Birge:1986:STD

[BP86] Robert R. Birge and Brian M. Pierce. Semiclassical timedependent theory of two-photon spectroscopy. The effect of dephasing in the virtual level on the two-photon excitation spectrum of isotachysterol. International Journal of Quantum Chemistry, 29(4):639–656, April 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Bonaccorsi:1983:UMP

[BPCT83] R. Bonaccorsi, P. Palla, R. Cimiraglia, and J. Tomasi. On the use of a MO polarized basis for the analysis of the interaction energy in molecular interactions: Application to amine complexes. International Journal of Quantum Chemistry, 24(3): 307–316, September 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Berrondo:1987:MEM

[BPLB87] M. Berrondo, A. Palma, and J. L. López-Bonilla. Matrix elements for the Morse potential using ladder operators. International Journal of Quantum Chemistry, 31(2):243–249, February 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Baldo:1983:ESL

[BPT83] Marcello Baldo, Renato Pucci, and Pasquale Tomasello. Excited states of linear polyenes in the SCF–RPA method. International Journal of Quantum Chemistry, 23(3):1111–1119, March 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Bartolotti:1988:TFT

[BPT88] Libero J. Bartolotti, Robert G. Parr, and Yoram Tal. Thomas–Fermi theory from a perturbative treatment of the hydrogenic limit energy density functional: a possible link to local density functional theory. International Journal of Quantum Chemistry, 34(6):509–515, December 1988. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Blum:1987:KCT

[BQS87] L. Blum, M. Quijada, and P. P. Schmidt. Kinetics of charge transfer at metal electrolyte interfaces: From classical to quantum statistics. International Journal of Quantum Chemistry, 32(S21):245–249, March 12, 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Chemistry, Solid-State Theory, and Computational Methods.

Bhargava:1980:FSG

[BR80] Shobha Bhargava and N. K. Ray. Floating spherical Gaussian orbital model study of some hydrogen-bridged systems: LiBeH_{3}, LiBH_{4}, LiCH_{5}, and BeBH_{5}. International Journal of Quantum Chemistry, 17(5):907–914, May 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Basilevsky:1981:DCR

[BR81a] M. V. Basilevsky and V. M. Ryaboy. Direct calculation of resonant states in reactive scattering. Application to linear triatomic systems. International Journal of Quantum Chemistry, 19(4):611–635, April 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Biswas:1981:CAS

[BR81b] Margaret Biswas and V. S. R. Rao. Conformational analysis of some mannose-containing and milk oligosaccharides:

Correlation of their shape with inhibitory properties in antigen/antibody interactions. International Journal of Quantum Chemistry, 20(1):99–121, July 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Brocas:1982:LHG

[BR82] J. Brocas and C. Rusu. Longuet-Higgins groups of XeF_{6}. International Journal of Quantum Chemistry, 22(2):331–350, August 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Balasubramanian:1985:SPS

[BR85] K. Balasubramanian and M. Randíc. Spectral polynomials of systems with general interactions. International Journal of Quantum Chemistry, 28(4):481–498, October 1985. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Brandas:1986:BRBb

[Brä86a] Erkki Brändas. Book review: Mathematics and computational concepts in chemistry. Edited by Nenad Trinajstić, Wiley, 1985. International Journal of Quantum Chemistry, 30 (6):867, December 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). See [Tri86].

Brandas:1986:BRBa

[Bra86b] Erkki Brandas. Book review: Pseudopotential Theory of Atoms and Molecules. By Levente Szasz. Wiley, New York, 1985. International Journal of Quantum Chemistry, 29(6): 1841, June 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). See [Sza85].

Brandas:1986:MCS

[Brä86c] Erkki Brändas. The method of complex scaling. International Journal of Quantum Chemistry, 30(S20):119–127, March 10– 15, 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Syposium on Atomic, Molecular, and Solid-State Theory, Scattering Problems, Many Body Phenomena, and Computational Quantum Chemistry.

Brandas:1987:ICR

[Brä87] Erkki Brändas. Introduction to the concept of resonances. International Journal of Quantum Chemistry, 31(5):699–705, May 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Bregadze:1980:NDI

[Bre80] V. G. Bregadze. Nature of DNA interactions with cations: UV Spectroscopic investigations and Marcus theory. International Journal of Quantum Chemistry, 17(6):1213–1219, June 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Brickmann:1984:RCG

[Bri84] Jürgen Brickmann. Raster computer graphics in molecular physics. International Journal of Quantum Chemistry, 26(S18):647–659, March 1–15, 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Atomic, Molecular and Solid-State Theory, and Computational Quantum Chemistry.

Brion:1986:LOL

[Bri86] C. E. Brion. Looking at orbitals in the laboratory: The experimental investigation of molecular wavefunctions and binding energies by electron momentum spectroscopy. International Journal of Quantum Chemistry, 29(5):1397–1428, May 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Briels:1987:CGF

[Bri87] W. J. Briels. The Coulomb Green’s function in IR^{3} and its relation to the harmonic oscillator in IR^{4}. International Journal of Quantum Chemistry, 32(3):349–359, September 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Broyles:1980:CMO

[Bro80] A. A. Broyles. Coordinate measurements and operators. International Journal of Quantum Chemistry, 17(1):47–74, January 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Brooks:1988:TCB

[Bro88] Charles L. Brooks III. Thermodynamic calculations on biological molecules. International Journal of Quantum Chemistry, 34(S15):221–234, March 12–19, 1988. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Syposium on Quantum Biology and Quantum Pharmacology.

Bryan:1980:NAH

[Bry80] Robert A. Bryan. On the nature of American Higher Education. International Journal of Quantum Chemistry, 18(S14): 5–7, March 10–20, 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory, Collision Phenomena, Quantum Statistics, and Computational Methods.

Bacic:1980:ACR

[BS80a] Zlatko Bačić and Jack Simons. Application of the coordinate rotation method to metastable atom-diatom scattering resonances. International Journal of Quantum Chemistry, 18 (S14):467–475, March 10–20, 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory, Collision Phenomena, Quantum Statistics, and Computational Methods.

Basilevsky:1980:MNM

[BS80b] M. V. Basilevsky and A. G. Shamov. Modification of the NDO model in semiempirical MO theory. International Journal of Quantum Chemistry, 17(2):285–292, February 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Broclawik:1980:ESS

[BS80c] Ewa Broclawik and Vedene H. Smith, Jr. Electronic structure of selenium chains by the SCF–SW–Xα method. International Journal of Quantum Chemistry, 18(S14):395–403, March 10–20, 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory, Collision Phenomena, Quantum Statistics, and Computational Methods.

Banerjee:1981:CCM

[BS81a] Ajit Banerjee and Jack Simons. The coupled-cluster method with a multiconfiguration reference state. International Journal of Quantum Chemistry, 19(2):207–216, February 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Born:1981:UGF

[BS81b] Gregory Born and Isaiah Shavitt. A unitary group formulation of open-shell electron propagator theory. International Journal of Quantum Chemistry, 20(S15):63, March 8–14, 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory, Collision Phenomena, and Computational Quantum Chemistry.

Bacic:1982:CCR

[BS82a] Zlatko Bačić and Jack Simons. Complex coordinate rotation calculation of branching ratios. International Journal of Quantum Chemistry, 21(4):727–739, April 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Black:1982:DAG

[BS82b] R. J. Black and G. E. Stedman. Diagram approach to group algebraic methods. International Journal of Quantum Chemistry, 21(5):823–831, May 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Blaha:1983:EDC

[BS83a] Peter Blaha and Karlheinz Schwarz. Electron densities and chemical bonding in TiC, TiN, and TiO derived from energy band calculations. International Journal of Quantum Chemistry, 23(4):1535–1552, April 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Brickmann:1983:QDW

[BS83b] Jürgen Brickmann and Peter C. Schmidt. Quantum dynamics of wave packets on phase space cells in nonlinearly coupled oscillators. International Journal of Quantum Chemistry, 23 (1):47–63, January 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Baraff:1984:PDS

[BS84] G. A. Baraff and M. Schluter. Point defects in semiconductors with strong electron-lattice coupling. International Journal of Quantum Chemistry, 26(S18):737, March 1–15, 1984.

CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Atomic, Molecular and Solid-State Theory, and Computational Quantum Chemistry.

Brenstein:1986:BSD

[BS86] Robert J. Brenstein and Steve Scheiner. The basis set dependence of structures and energies of various states of cyclodisiloxane. International Journal of Quantum Chemistry, 29 (5):1191–1208, May 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Brower:1980:IMO

[BSCR80] A. L. Brower, J. R. Sabin, B. Crist, and M. A. Ratner. Ab initio molecular orbital studies of polyethylene deformation. International Journal of Quantum Chemistry, 18(2):651–654, August 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Besnainou:1982:CMA

[BSD82] Sylvette Besnainou, James Stone, and Jean-Claude Diels. Coherent multiphoton absorption: Dependency on the relative phase of two successive pulses. International Journal of Quantum Chemistry, 22(5):999–1003, November 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Barandiaran:1986:SPT

[BSHK86] Z. Barandiaran, L. Seijo, S. Huzinaga, and M. Klobukowski. Structure and properties of transition-metal compounds. A systematic study of basis set effects in ab initio SCF calculations. International Journal of Quantum Chemistry, 29(5): 1047–1058, May 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Blagoy:1980:ESE

[BSL^{+}80] Y. U. P. Blagoy, G. G. Sheina, A. V. Luzanov, L. K. Silina, V. F. Pedash, Y. U. V. Rubin, and E. A. Leibina. Effect of substituents on electron energy redistribution in uracil derivatives and their ionization in polar solvents. International Journal of Quantum Chemistry, 18(4):913–919, October 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Beleznay:1981:IBP

[BSL81] F. Beleznay, S. Suhai, and J. Ladik. Interaction between polymers: Dispersion energy of two infinite linear chains. International Journal of Quantum Chemistry, 20(3):683–692, September 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Blomberg:1980:GSP

[BSR80] Margareta R. A. Blomberg, Per E. M. Siegbahn, and Björn O. Roos. The ground-state potential curve of the beryllium dimer. International Journal of Quantum Chemistry, 18(S14): 229–247, March 10–20, 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory, Collision Phenomena, Quantum Statistics, and Computational Methods.

Bonaccorsi:1986:SDD

[BST86] R. Bonaccorsi, E. Scrocco, and J. Tomasi. Structural deformations of the DNA double helix in the first stages of DNA transcription studied with a simple model. International Journal of Quantum Chemistry, 29(4):717–735, April 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Beran:1981:AGO

[BSZ81] S. Beran, Z. Slanina, and D. C. Zidarov. Automatic geometry optimization for molecules with d orbitals. III. The EHT method with electrostatic corrections. International Journal of Quantum Chemistry, 19(4):585–591, April 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Bonchev:1982:CIT

[BT82] D. Bonchev and N. Trinajstić. Chemical information theory: Structural aspects. International Journal of Quantum Chemistry, 22(S16):463–480, March 1–13, 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Chemistry, Theory of Condensed Matter, and Propagator Methods in the Quantum Theory of Matter.

Barysz:1984:NAC

[BT84] M. Barysz and N. Trinajstić. A novel approach to the characterization of chemical structures. International Journal of Quantum Chemistry, 26(S18):661–673, March 1–15, 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Atomic, Molecular and Solid-State Theory, and Computational Quantum Chemistry.

Briels:1983:HIR

[BTC^{+}83] W. J. Briels, J. Tennyson, M. Claessens, Th. Van Der Lee, and A. Van Der Avoird. Hindered internal rotations in van der Waals molecules and molecular crystals. International Journal of Quantum Chemistry, 23(3):1091–1100, March 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Barysz:1983:SCS

[BTK83] M. Barysz, N. Trinajstić, and J. V. Knop. On the similarity of chemical structures. International Journal of Quantum Chemistry, 24(S17):441–451, March 3–12, 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-state Theory, Collision Phenomena, and Computational Quantum Chemistry.

Buckingham:1989:ACH

[Buc89] A. D. Buckingham. Andrew Crowther Hurley: Introductory remarks. International Journal of Quantum Chemistry, 36(S23):XXXI–XXXIII, April 1–8, 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Buenker:1986:CPT

[Bue86] Robert J. Buenker. Combining perturbation theory techniques with variational CI calculations to study molecular excited states. International Journal of Quantum Chemistry, 29(3): 435–460, March 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Buff:1982:SDG

[Buf82] Frank P. Buff. Shape dependence of the grand potential. International Journal of Quantum Chemistry, 22(S16):671, March 1–13, 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Chemistry, Theory of Condensed Matter, and Propagator Methods in the Quantum Theory of Matter.

Burton:1983:RFC

[Bur83] Peter G. Burton. Routes to full CI through extrapolation in multireference configuration calculations for electronic structures. International Journal of Quantum Chemistry, 23(2): 613–623, February 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Burke:1985:MPT

[Bur85] L. A. Burke. A model potential technique for the calculation of the band structure of poly(glycine). International Journal of Quantum Chemistry, 28(S12):25–31, March 14–16, 1985. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Burke:1986:SDA

[Bur86] Luke A. Burke. Synchronism in the Diels-Alder reaction. International Journal of Quantum Chemistry, 29(3):511–518, March 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Bruggemann:1980:LBH

[BV80] R. Brüggemann and J. Voitländer. Local behavior of hückel parameters. International Journal of Quantum Chemistry, 18 (3):775–782, September 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Broch:1986:CEP

[BV86] H. Broch and D. Vasilescu. Conformation and electrostatic properties quantum determination of the new radioprotector and anticancer drug I-102. International Journal of Quantum Chemistry, 30(S13):81–94, March 6–8, 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Brucenta:1981:ETM

[BVAB81] A. Bruceñta, P. Vermeulin, P. Archirel, and G. Berthier. Exponential transformation of molecular orbitals. II. General formulation for UHF calculations and application to diatomics and molecular fragments. International Journal of Quantum Chemistry, 20(6):1285–1300, December 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Boerrigter:1988:TDN

[BVB88] P. M. Boerrigter, G. Te Velde, and J. E. Baerends. Threedimensional numerical integration for electronic structure calculations. International Journal of Quantum Chemistry, 33 (2):87–113, February 1988. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Boerth:1980:HOM

[BVC80] Donald W. Boerth and F. A. Van-Catledge. Hydrogenic orbitals as minimum-basis functions for SCF calculations. International Journal of Quantum Chemistry, 18(4):973–993, October 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Barbier:1981:BPF

[BVD81] C. Barbier, C. Vincent, and G. Del Re. Bond polarization in the FeCO system: Semiempirical MO–SCF (BMV) calculations. International Journal of Quantum Chemistry, 19(6): 1057–1063, June 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Boulanger:1988:IZM

[BVDA88] Thierry Boulanger, Daniel P. Vercauteren, François Durant, and Jean-Marie André. 3- and 5-Isoxazolol zwitterions: A model of interaction with the GABA-A receptor relating to agonism and antagonism. International Journal of Quantum Chemistry, 34(S15):149–165, March 12–19, 1988. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Syposium on Quantum Biology and Quantum Pharmacology.

Broch:1989:QMS

[BVGV89] H. Broch, R. Viani, H. Grassi, and D. Vasilescu. Quantum molecular study of the Z access for the phosphodiester linkage in nucleic acids. International Journal of Quantum Chemistry, 36(S16):171–184, April 1–8, 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Basilevsky:1980:VRQ

[BW80] M. V. Basilevsky and N. N. Weinberg. On the validity range of the qualitative theory of chemical reactivity. International Journal of Quantum Chemistry, 17(3):399–413, March 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Berggren:1985:FSB

[BW85] K.-F. Berggren and B. L. Widlund. Formation of sub-bands and the 3D to 2D transition of electrons in a GaAs MESFET. International Journal of Quantum Chemistry, 28(S19):567– 578, March 18–23, 1985. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Brenner:1989:CMI

[BWEW89] Donald W. Brenner, C. T. White, Mark L. Elert, and Franklin E. Walker. Chemical model for intrinsic detonation velocities. International Journal of Quantum Chemistry, 36 (S23):333–337, April 1–8, 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Boucekkine-Yaker:1984:BSE

[BYBB84] Gh. Boucekkine-Yaker, A. Boucekkine, and G. Berthier. Basis-set effects on calculated molecular magnetic properties. International Journal of Quantum Chemistry, 26(S18):369– 374, March 1–15, 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Atomic, Molecular and Solid-State Theory, and Computational Quantum Chemistry.

Boucekkine-Yaker:1983:ICM

[BYBZB83] G. Boucekkine-Yaker, A. Boucekkine, M. Zaucer, and G. Berthier. Ab initio calculation of magnetic susceptibilities and screening constants using gauge invariant Gaussian orbitals. International Journal of Quantum Chemistry, 23(2): 365–367, February 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Balbas:1984:RBT

[BZA84] L. C. Balbás, M. L. Zorita, and J. A. Alonso. Relation between total energy, electronic potential at the nucleus, and chemical potential of positive ions. International Journal of Quantum Chemistry, 26(1):145–149, July 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Coleman:1980:RHO

[CA80] A. J. Coleman and I. Absar. Reduced Hamiltonian orbitals. III. Unitarily invariant decomposition of Hermitian operators. International Journal of Quantum Chemistry, 18(5): 1279–1307, November 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Contreras:1985:STC

[CA85] Renato Contreras and Arie Aizman. On the SCF theory of continuum solvent effects representation: Introduction of local dielectric effects. International Journal of Quantum Chemistry, 27(3):293–301, March 1985. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Contreras:1986:STC

[CA86] Renato Contreras and Arie Aizman. On the SCF theory of continuum solvent effects representation II. Quantum chemical calculation of thermodynamic properties of some acidbase equilibria in solution. International Journal of Quantum Chemistry, 30(S20):573–584, March 10–15, 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Syposium on Atomic, Molecular, and Solid-State Theory, Scattering Problems, Many Body Phenomena, and Computational Quantum Chemistry.

Chacham:1986:NGA

[CADL86] H. Chacham, J. L. A. Alves, M. L. De Siqueira, and J. R. Leite. The noble gas atoms as impurities in silicon. International Journal of Quantum Chemistry, 30(S20):347–351, March 10–15, 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Syposium on Atomic, Molecular, and Solid-State Theory, Scattering Problems, Many Body Phenomena, and Computational Quantum Chemistry.

Calais:1980:LD

[Cal80] Jean-Louis Calais. On the localization of defects. International Journal of Quantum Chemistry, 18(2):607–613, August 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Calais:1981:BRBa

[Cal81a] Jean-Louis Calais. Book review: Chimie quantique structural et éléments de spectroscopie théorique. By A. Julg. Office des publications universitaires, Alger, 1978. International Journal of Quantum Chemistry, 20(6):1331–1332, December 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Calais:1981:BRBc

[Cal81b] Jean-Louis Calais. Book review: Electrons and Phonons in Layered Crystal Structures. Edited by T. J. Wietling and M. Schluter. Physics and Chemistry of Materials with Layered Structures, Volume 3 D. Reidel, Dordrecht and Boston, 1979. International Journal of Quantum Chemistry, 20(6): 1332–1333, December 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). See [WS79].

Calais:1981:BRBb

[Cal81c] Jeans-Louis Calais. Book review: Quantum Theory of Polymers. Edited by J.-M. André, J. Delhalle, and J. Ladik. NATO advanced study Institutes series C, Volume 39. Reidle, Dordrecht and Boston, 1978. International Journal of Quantum Chemistry, 20(6):1332, December 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). See [ADL78].

Calais:1982:BRBa

[Cal82a] Jean-Louis Calais. Book review: Quantum dynamics of molecules: The new experimental challenge to theorists. Edited by R. G. Woolley. NATO Advanced Study Institute Series B57, Plenum, New York and London, 1980. International Journal of Quantum Chemistry, 22(5):1083–1084, November 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). See [Woo80].

Calais:1982:BRBb

[Cal82b] Jean-Louis Calais. Book review: Spin Eigenfunctions: Construction and use. By Ruben Pauncz. Plenum, New York and London, 1979. International Journal of Quantum Chemistry, 22(5):1083–1084, November 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). See [Pau79].

Calais:1982:QCC

[Cal82c] Jean-Louis Calais. Is quantum chemistry concerned with materials science? International Journal of Quantum Chemistry, 21(1):231–237, January 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Callis:1984:TDT

[Cal84] Patrik R. Callis. Transition density topology of the L_{a} and L_{b} states in indoles and purines. International Journal of Quantum Chemistry, 26(S18):579–588, March 1–15, 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Atomic, Molecular and Solid-State Theory, and Computational Quantum Chemistry.

Calais:1985:OSA

[Cal85] Jean-Louis Calais. Orthonormalization and symmetry adaptation of crystal orbitals. International Journal of Quantum Chemistry, 28(S19):655–667, March 18–23, 1985. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Calais:1986:BRBb

[Cal86a] Jean-Louis Calais. Book review: Molecular Structure and Conformation. Progress in Theoretical Organic Chemistry, Vol. 3, I. G. Csizmadia (Ed.). Elsevier Scientific Publishing Company, Amsterdam, 1982, 344 pp. International Journal of Quantum Chemistry, 29(6):1839, June 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). See [Csi82].

Calais:1986:BRBa

[Cal86b] Jean-Louis Calais. Book review: Recent advances in group theory and their application to spectroscopy. By John C. Donini, ed., NATO Advanced Study Institute series, vol. B43. Plenum Press, New York, 1979, pp. xii + 692. International Journal of Quantum Chemistry, 29(1):169, January 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). See [Don79].

Calais:1986:BRBc

[Cal86c] Jean-Louis Calais. Book review: Solitons in molecular systems. By A. S. Davydov, Reidel, Dordrect, 1985. International Journal of Quantum Chemistry, 30(6):869, December 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). See [Dav85, Dav91].

Calais:1986:BRBd

[Cal86d] Jean-Louis Calais. Book review: Vibronic coupling. By G. Fischer, Academic, London, 1984. International Journal of Quantum Chemistry, 30(6):871, December 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). See [Fis84].

Calais:1988:BRB

[Cal88] Jean-Louis Calais. Book review: Molecular magnetism. By A. Hinchliffe and R. W. Mum, John Wiley & Sons, Chichester, 1985. International Journal of Quantum Chemistry, 33(6): 577, June 1988. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). See [HM85b].

Calais:1989:BRB

[Cal89a] Jean-Louis Calais. Book review: Ab Initio Methods in Quantum Chemistry, Part I and Part II. Wiley-Interscience, New York, Advances in Chemical Physics, Vol. 67, 556 pp, and Vol. 69, 588 pp, 1987. K. P. Lawley, Editor. International Journal of Quantum Chemistry, 35(4):581, April 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). See [Law87].

Calais:1989:OMS

[Cal89b] Jean-Louis Calais. Orthogonalization in momentum space. International Journal of Quantum Chemistry, 35(6):735–743, June 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Contreras:1987:QCC

[CAML87] Renato R. Contreras, Arie Aizman, Demian Morales, and Maria E. Llanos. Quantum chemical calculations on ion pairs. II. Structure of haloacetic acids in aqueous solution. International Journal of Quantum Chemistry, 32(S21):753–754, March 12, 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Chemistry, Solid-State Theory, and Computational Methods.

Careri:1982:PFI

[Car82] Giorgio Careri. Periodic flow instabilities in the wake of an ion. International Journal of Quantum Chemistry, 22(S16): 227–236, March 1–13, 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Chemistry, Theory of Condensed Matter, and Propagator Methods in the Quantum Theory of Matter.

Casida:1985:WDL

[Cas85] Mark Casida. What do local energy maps tell about local error? International Journal of Quantum Chemistry, 27(4):451– 460, April 1985. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Casida:1988:UIB

[Cas88] Mark E. Casida. Use of invariants in bounding 2-fermion reduced density matrix eigenvalues. International Journal of Quantum Chemistry, 33(1):53–60, January 1988. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Csavinszky:1982:NSP

[CB82] P. Csavinszky and K. R. Brownstein. Nonlinear screening of positive and negative point charges in GaAs. International Journal of Quantum Chemistry, 22(S16):457–462, March 1– 13, 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Chemistry, Theory of Condensed Matter, and Propagator Methods in the Quantum Theory of Matter.

Chao:1983:CHC

[CB83] K. A. Chao and K.-F. Berggren. Clustered hydrogenic centers in impure semiconductors. International Journal of Quantum Chemistry, 23(4):1155–1163, April 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Cisneros:1985:TMS

[CB85] Gerardo Cisneros and Carlos F. Bunge. Towards a modular system in computational quantum chemistry. International Journal of Quantum Chemistry, 28(S19):193–199, March 18– 23, 1985. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Cioslowski:1988:ETE

[CB88] J. Cioslowski and D. Bessis. Extrapolation of the total energy of polymers to the bulk limit using generalized Padé approximants. International Journal of Quantum Chemistry, 34(3): 225–229, September 1988. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Carlsson:1989:CS

[CBC^{+}89] Gudrun Alm Carlsson, Karl-Fredrik Berggren, Carl A. Carlsson, Georg Matscheko, and Roland Ribberfors. The Compton spectrometer. International Journal of Quantum Chemistry, 35(6):721–734, June 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Cabrol:1985:QCS

[CBV85] D. Cabrol, H. Broch, and D. Vasilescu. Quantum conformational studies of model molecules in relation to collagenic structures: Effect of nature and location of aliphatic residues in (gly-R_{2}pro)_{n} and (gly-pro-R_{3})_{n} sequences. International Journal of Quantum Chemistry, 28(S12):141–152, March 14– 16, 1985. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Clement:1982:CTN

[CC82a] Enrico Clement and Giorgina Corongiu. Charge transfer in nucleic acids. International Journal of Quantum Chemistry, 22(S9):213–221, March 4–6, 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Clementi:1982:BDS

[CC82b] E. Clementi and G. Corongiu. B-DNA structural determination of Na^{+} counterions at different humidities, ionic concentrations, and temperatures. International Journal of Quantum Chemistry, 22(3):595–630, September 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Clementi:1983:MCS

[CC83] E. Clementi and G. Corongiu. Monte Carlo study of liquid water with two- and three-body ab initio potentials. International Journal of Quantum Chemistry, 24(S10):31–41, March 14–16, 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Calais:1984:CCW

[CC84a] Jean-Louis Calais and Juan Antonio Cogordan. Charge current waves in a one-dimensional Coulomb gas. International Journal of Quantum Chemistry, 26(S18):67–76, March 1–15, 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Atomic, Molecular and Solid-State Theory, and Computational Quantum Chemistry.

Corongiu:1984:MCL

[CC84b] J. Detrich G. Corongiu and E. Clementi. Monte Carlo liquid water simulations with four-body interactions included. International Journal of Quantum Chemistry, 26(S18):701–707, March 1–15, 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Atomic, Molecular and Solid-State Theory, and Computational Quantum Chemistry.

Clementi:1989:SSS

[CCC^{+}89] E. Clementi, S. Chin, G. Corongiu, J. H. Detrich, M. Dupuis, D. Folsom, G. C. Lie, D. Logan, and V. Sonnad. Supercomputing and supercomputers for science and engineering in general and for chemistry and biosciences in particular. International Journal of Quantum Chemistry, 35(1):3–89, January 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Clementi:1984:PQC

[CCD^{+}84] E. Clementi, G. Corongiu, J. Detrich, S. Chin, and L. Domingo. Parallelism in quantum chemistry: Hydrogen bond study in DNA base pairs as an example. International Journal of Quantum Chemistry, 26(S18):601–618, March 1– 15, 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Atomic, Molecular and Solid-State Theory, and Computational Quantum Chemistry.

Clementi:1982:TCC

[CCG^{+}82] E. Clementi, G. Corongiu, M. Gratarola, P. Habitz, C. Lupo, P. Otto, and D. Vercauteren. Theoretical and computational chemistry of complex systems: Solvation of DNA and proteins. International Journal of Quantum Chemistry, 22(S16): 409–433, March 1–13, 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Chemistry, Theory of Condensed Matter, and Propagator Methods in the Quantum Theory of Matter.

Collins:1984:RTM

[CCS84] T. C. Collins, M. Chandrasekhar, and M. Seel. A review of a theoretical model of the excitonic superconductivity in CdS and CuCl. International Journal of Quantum Chemistry, 26 (5):831–841, November 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Chatzidimitriou-Dreismann:1983:ITF

[CD83] C. A. Chatzidimitriou-Dreismann. Irreversibility, time, and fluctuations in quantum statistical theory of dissipative molecular systems. II. Application to fluorescence spectroscopy. International Journal of Quantum Chemistry, 23(4):1505–1516, April 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Chatzidimitriou-Dreismann:1985:TSD

[CD85] C. A. Chatzidimitriou-Dreismann. On time superoperator and dynamical fluctuations in photoemission. International Journal of Quantum Chemistry, 28(S19):369–382, March 18–23, 1985. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Canuto:1987:MBS

[CD87a] Sylvio Canuto and Geerd H. F. Diercksen. Many-body studies of the structure and spectra of CO_{3}. International Journal of Quantum Chemistry, 32(S21):759–760, March 12, 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Chemistry, Solid-State Theory, and Computational Methods.

Carbo:1987:LMS

[CD87b] Ramon Carbó and Llorenç Domingo. LCAO–MO similarity measures and taxonomy. International Journal of Quantum Chemistry, 32(4):517–545, October 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Cizek:1987:CCS

[ČD87c] Jiří Čížek and Giuseppe Del Re. C. A. Coulson and the surface energy of metals: The distribution of eigenvalues as a difficult problem in number theory. International Journal of Quantum Chemistry, 31(2):287–293, February 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Chatzidimitriou-Dreismann:1989:PED

[CD89a] C. A. Chatzidimitriou-Dreismann. Proposed experiments for the detection of coherent–dissipative structures around H^{+} ions in aqueous solutions. International Journal of Quantum Chemistry, 36(S23):153–158, April 1–8, 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Cundari:1989:MOI

[CD89b] Thomas R. Cundari and Russell S. Drago. A molecular orbital investigation of Ru–oxo-catalyzed alcohol oxidations. International Journal of Quantum Chemistry, 36(S23):489–499, April 1–8, 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Cundari:1989:ROC

[CD89c] Thomas R. Cundari and Russell S. Drago. Ru–oxo catalyzed epoxidations and the Woodward–Hoffmann rules. International Journal of Quantum Chemistry, 36(6):773–790, December 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Cheetham:1992:SSC

[CD92] A. K. Cheetham and Peter Day, editors. Solid state chemistry: compounds. Oxford science publications. Clarendon Press, New York, NY, USA, 1992. ISBN 0-19-855166-5. xii + 304 pp. LCCN QD478 .S6334 1991. URL http://www.loc.gov/

catdir/enhancements/fy0605/91012173-d.html;   http:
//www.loc.gov/catdir/enhancements/fy0605/91012173-
t.html.

Carbo:1980:STM

[CDG80a] R. Carbó, Ll. Domingo, and J. Gregori. SCF Theory of multiplet states. International Journal of Quantum Chemistry, 17 (4):725–736, April 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Carbo:1980:AFO

[CDG80b] Ramon Carbó, Llorençl Domingo, and Josep Gregori. Average Fock operators. International Journal of Quantum Chemistry, 18(5):1207–1221, November 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Chatzidimitriou-Dreismann:1982:RPQ

[CDL82] C. A. Chatzidimitriou-Dreismann and E. Lippert. On the removal of Prigogine’s quantum statistical “degeneracy” in the theory of molecular systems. International Journal of Quantum Chemistry, 22(S16):195–204, March 1–13, 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Chemistry, Theory of Condensed Matter, and Propagator Methods in the Quantum Theory of Matter.

Csavinszky:1986:DRD

[CE86] P. Csavinszky and A. M. Elabsy. Dielectric response to a donor ion in a Ga_{1−x}Al_{x}As − GaAs-Ga_{1−x}Al_{x}As quantum well of finite depth. International Journal of Quantum Chemistry, 30 (S20):325–333, March 10–15, 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Syposium on Atomic, Molecular, and Solid-State Theory, Scattering Problems, Many Body Phenomena, and Computational Quantum Chemistry.

Csavinszky:1987:DRA

[CE87] P. Csavinszky and A. M. Elabsy. Dielectric response to an acceptor ion in a Ga_{1−x}A1_{x}As/GaAs/Ga_{1−x}A1_{x}As quantum well. International Journal of Quantum Chemistry, 32(S21): 79–87, March 12, 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Chemistry, Solid-State Theory, and Computational Methods.

Csavinszky:1988:ENG

[CE88] P. Csavinszky and A. M. Elabsy. Effect of the nonparabolicity of the GaAs conduction band on the binding energy of off-center hydrogenic donors in a Ga_{1−x}Al_{x}As/GaAs/ Ga_{1−x}Al_{x}As quantum well. International Journal of Quantum Chemistry, 34(S22):25–31, March 12–19, 1988. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Syposium on Quantum Chemistry, Solid-State Theory, and Computational Methods.

Cardelino:1986:ECD

[CEB86] B. H. Cardelino, W. H. Eberhardt, and R. F. Borkman. Electronic charge density and mean kinetic energy of lithium orbitals in LiH, LiH^{+}, Li_{2}, Li_{2}^{+}, LiH_{2}^{+}, and Li_{2}H^{+}. International Journal of Quantum Chemistry, 29(5):1635–1649, May 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

CorreaDeMello:1983:CCM

[CEZ83] Paulo Corrêa De Mello, W. Daniel Edwards, and Michael C. Zerner. Cr Cr multiple bonding in binuclear complexes. International Journal of Quantum Chemistry, 23(2):425–436, February 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Castro:1981:SCA

[CF81a] Eduardo A. Castro and Francisco M. Fernández. Some comments about the effective utilization of off-diagonal hypervirial relations considering the diagonal hypervirial relation: Harmonic oscillator case. International Journal of Quantum Chemistry, 19(5):967–968, May 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Curtiss:1981:IMO

[CF81b] L. A. Curtiss and D. J. Frurip. Ab initio molecular orbital calculations on hydrogen bonding in binary wateralcohol mixtures: (CH_{3}OH)_{2}(H_{2}O), (CH_{3}OH) (H_{2}O)_{2}, and (CF_{3}CH_{2}OH)(H_{2}O). International Journal of Quantum Chemistry, 20(S15):189–197, March 8–14, 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory, Collision Phenomena, and Computational Quantum Chemistry.

Cohen:1983:CHE

[CF83] Leon Cohen and C. Frishberg. Correlation hole for an exactly solvable problem. International Journal of Quantum Chemistry, 23(1):249–255, January 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Corongiu:1983:HAD

[CFC83] Giorgina Corongiu, Sandro L. Fornili, and Enrico Clementi. Hydration of agarose double helix: A Monte Carlo simulation. International Journal of Quantum Chemistry, 24(S10): 277–291, March 14–16, 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Clinton:1983:DMM

[CFG^{+}83] W. L. Clinton, C. A. Frishberg, M. J. Goldberg, L. J. Massa, and P. A. Oldfield. Density matrix model for coherent X-ray diffraction: Study of experimental factors related to idempotency. International Journal of Quantum Chemistry, 24(S17): 517–525, March 3–12, 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-state Theory, Collision Phenomena, and Computational Quantum Chemistry.

Cohen:1985:CES

[CFLM85] L. Cohen, C. Frishberg, Chongmoon Lee, and L. J. Massa. Correlation energy for a Slater determinant fitted to the electron density. International Journal of Quantum Chemistry, 28 (S19):525–533, March 18–23, 1985. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Colle:1986:MBS

[CFS86] Renato Colle, Alessandro Fortunelli, and Oriano Salvetti. A mixed basis set of Slater and Gaussian functions for molecular calculations. International Journal of Quantum Chemistry, 29 (6):1825–1837, June 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Colle:1987:MBS

[CFS87] Renato Colle, Alessandro Fortunelli, and Oriano Salvetti. A mixed basis set of Slater and Gaussian functions. Integrals with modified Gaussian functions. International Journal of Quantum Chemistry, 31(6):893–901, June 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Constanciel:1980:PDA

[CG80] R. Constanciel and H. Gritli. Pair density analysis of semiempirical wave functions: Study of water molecule including environmental effects. International Journal of Quantum Chemistry, 17(6):1047–1068, June 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Chermette:1986:SEP

[CG86] H. Chermette and A. Goursot. Is the self energy proportional to electronic relaxation? International Journal of Quantum Chemistry, 29(5):1277–1282, May 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Careri:1981:MAH

[CGG81] G. Careri, A. Giansanti, and E. Gratton. Molecular aspects of the hydration process in a globular protein. International Journal of Quantum Chemistry, 19(6):1133–1135, June 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Christophorov:1982:KTB

[CGK82] L. N. Christophorov, V. V. Gorbach, and V. N. Kharkyanen. Kinetic theory of bridge processes. International Journal of Quantum Chemistry, 22(3):485–496, September 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Chen:1987:IBA

[CGP87] Kai-Xian Chen, Nohad Gresh, and Bernard Pullman. Intercalative binding and antitumor activity of bisantrene and derivatives. International Journal of Quantum Chemistry, 32 (S14):15–25, March 14, 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Cammi:1986:E

[CGT86a] Roberto Cammi, Caterina Ghio, and Jacopo Tomasi. Erratum. International Journal of Quantum Chemistry, 30(2): 303–304, August 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Cammi:1986:NOL

[CGT86b] Roberto Cammi, Caterina Ghio, and Jacopo Tomasi. Neutral organic Lewis acids of π type. International Journal of Quantum Chemistry, 29(3):527–539, March 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Curtiss:1982:IMO

[CH82] L. A. Curtiss and A. Heinricher. ab initio molecular orbital studies of the structure and stability of the BeAl_{2}F_{8} complex. International Journal of Quantum Chemistry, 22(S16):285– 291, March 1–13, 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Chemistry, Theory of Condensed Matter, and Propagator Methods in the Quantum Theory of Matter.

Casida:1986:GDM

[CH86] Mark E. Casida and John E. Harriman. Geometry of density matrices. VI. Superoperators and unitary invariance. International Journal of Quantum Chemistry, 30(2):161–212, August 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Chang:1980:PSC

[Cha80] Tse-Chiang Chang. Perturbation and SCF calculations for excited states. International Journal of Quantum Chemistry, 18(1):43–48, July 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Chakraborty:1982:MME

[Cha82] B. P. Chakraborty. Morse matrix elements: An asymptotic expansion treatment. International Journal of Quantum Chemistry, 22(1):73–88, July 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Charbonneau:1985:A

[Cha85] K. Charbonneau. Announcement. International Journal of Quantum Chemistry, 27(1):105, January 1985. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Casida:1987:HFE

[CHA87] Mark E. Casida, John E. Harriman, and James L. Anchell. The Husimi function for electron distributions. International Journal of Quantum Chemistry, 32(S21):435–456, March 12, 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Chemistry, Solid-State Theory, and Computational Methods.

Chang:1988:VSH

[Cha88] Man Chai Chang. Vibrational studies of HCN and DCN in matrix. International Journal of Quantum Chemistry, 34(S22): 167–171, March 12–19, 1988. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Syposium on Quantum Chemistry, Solid-State Theory, and Computational Methods.

Csaszar:1988:VFA

[CHB88] Pál Császár, László Harsányi, and James E. Boggs. Vibrational frequencies and assignments for some isotopomers of urcail using a scaled ab initio force field. International Journal of Quantum Chemistry, 33(1):1–17, January 1988. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Chen:1980:CMC

[Che80] Cheng Chen. CNDO–MO calculation of N_{n}F_{m} molecules. International Journal of Quantum Chemistry, 18(1):199–204, July 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Cheney:1982:SFA

[Che82] B. Vernon Cheney. Structural factors affecting aryl hydrocarbon hydroxylase induction by dibenzo-p-dioxins and dibenzofurans. International Journal of Quantum Chemistry, 21(2): 445–463, February 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Chidambaram:1981:NES

[Chi81] R. Chidambaram. Neutron elastic scattering studies of biological molecules and complexes. International Journal of Quantum Chemistry, 20(2):449–459, August 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Chabalowski:1983:IMC

[CHKB83] Cary Chabalowski, P. C. Hariharan, Joyce J. Kaufman, and Robert Buenker. Ab-initio multireference CI calculations on CH_{3}NO_{2} confirm earlier preliminary GVB and MCSCF/CI results that HNO_{2} and CH_{3}NO_{2} have multiconfiguration ground electronic states even at equilibrium geometry as well as multiconfiguration excited states. International Journal of Quantum Chemistry, 24(S17):643–644, March 3–12, 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-state Theory, Collision Phenomena, and Computational Quantum Chemistry.

Chow:1980:FBS

[Cho80] Yutze Chow. Fiber bundles, sheaf theory, and generalization of differentiable manifolds in physics. International Journal of Quantum Chemistry, 17(1):85–97, January 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Christov:1980:CTBa

[Chr80a] S. G. Christov. Collision theory of bimolecular and unimolecular reactions and its experimental test. International Journal of Quantum Chemistry, 18(1):263–273, July 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Christov:1980:CTBb

[Chr80b] S. G. Christov. Collision theory of bimolecular and unimolecular reactions and its experimental test. International Journal of Quantum Chemistry, 18(6):1541–1542, December 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Christensen:1984:RBS

[Chr84] N. Egede Christensen. Relativistic band structure calculations. International Journal of Quantum Chemistry, 25(1): 233–261, January 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Christov:1989:STR

[Chr89] S. G. Christov. The stochastic theory and its relations to the general reaction rate theory. International Journal of Quantum Chemistry, 36(3):391–415, September 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Chu:1986:CSM

[Chu86] Shih-I Chu. The complex scaling method: Application to autoionization, predissociation, and multiphoton resonances. International Journal of Quantum Chemistry, 30(S20):129–146, March 10–15, 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Syposium on Atomic, Molecular, and Solid-State Theory, Scattering Problems, Many Body Phenomena, and Computational Quantum Chemistry.

Chun:1988:APB

[Chu88] Paul W. Chun. Approximation of the Planck–Benzinger thermal work function in protein refolding in ribonuclease systems. International Journal of Quantum Chemistry, 34(S15):247– 258, March 12–19, 1988. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Syposium on Quantum Biology and Quantum Pharmacology.

CorreaDeMello:1982:CSC

[CHZ82] P. Corrêa De Mello, M. Hehenberger, and M. C. Zernert. Converging SCF calculations on excited states. International Journal of Quantum Chemistry, 21(1):251–258, January 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Ciambelli:1981:BTF

[Cia81a] Paolo Ciambelli. “binding” technology and fundamental research: Some cases in catalysis. International Journal of Quantum Chemistry, 19(6):1003–1010, June 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Cianetti:1981:RTB

[Cia81b] Elvio Cianetti. Remarks on technology and bonding. International Journal of Quantum Chemistry, 19(6):1001–1002, June 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Cioslowski:1987:ASC

[Cio87a] J. Cioslowski. Algebraic structure count of rotagraphs. International Journal of Quantum Chemistry, 31(4):605–609, April 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Cioslowski:1987:FDR

[Cio87b] J. Cioslowski. The functional derivatives of the Rayleigh’s quotient—concept and applications. International Journal of Quantum Chemistry, 32(4):451–456, October 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Cioslowski:1987:SCM

[Cio87c] J. Cioslowski. A study of the connected moments expansions for the correlation energy via an exactly soluble problem. International Journal of Quantum Chemistry, 32(S21):563–567, March 12, 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Chemistry, Solid-State Theory, and Computational Methods.

Cioslowski:1987:UTS

[Cio87d] J. Cioslowski. A unified theory of the stability of benzenoid hydrocarbons. International Journal of Quantum Chemistry, 31(4):581–590, April 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Cioslowski:1988:GPA

[Cio88a] J. Cioslowski. The generalized Padé approximants and chemical graph theory. International Journal of Quantum Chemistry, 34(3):217–224, September 1988. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Cioslowski:1988:TEE

[Cio88b] J. Cioslowski. Total π-electron energy in the “variable β” Hückel method. International Journal of Quantum Chemistry, 34(5):417–421, November 1988. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Cioslowski:1989:DDS

[Cio89] J. Cioslowski. Density driven self-consistent field method. II. construction of all one-particle wavefunctions that are orthonormal and sum up to a given density. International Journal of Quantum Chemistry, 36(S23):255–262, April 1–8, 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Cui:1989:MDS

[CJ89] S. T. Cui and R. E. Johnson. A molecular dynamics study of ejection of molecules from a vibrationally excited “track” in an amorphous solid. International Journal of Quantum Chemistry, 36(S23):575–582, April 1–8, 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Chunwan:1986:BPM

[CJZ^{+}86] Liu Chunwan, Hua Jianmin, Chen Zhida, Lin Zhenyang, and Lu Jiaxi. The bonding properties of Mo Fe s clusters. International Journal of Quantum Chemistry, 29(4):701–715, April 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Castano:1982:ISA

[CK82] O. Castaño and P. Karadakov. Instabilities of the symmetryadapted restricted-Hartree–Fock ground state in infinite polyenes. II. Nonsinglet instabilities. International Journal of Quantum Chemistry, 22(2):367–383, August 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Cederbaum:1981:MVC

[CKD81] L. S. Cederbaum, H. Köppel, and W. Domcke. Multimode vibronic coupling effects in molecules. International Journal of Quantum Chemistry, 20(S15):251–267, March 8–14, 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory, Collision Phenomena, and Computational Quantum Chemistry.

Cvitas:1986:HRP

[CKK86] T. Cvitaš, L. Klasinc, and B. Kovač. High-resolution photoelectron spectrum of ozone. International Journal of Quantum Chemistry, 29(4):657–661, April 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Clementi:1980:NIW

[CKLR80] E. Clementi, W. Kołos, G. C. Lie, and G. Ranghino. Nonadditivity of interaction in water trimers. International Journal of Quantum Chemistry, 17(3):377–398, March 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Castro:1981:BMT

[CKM81] Miguel Castro, Jaime Keller, and Pilar Mareca. Bonding in the MO_{2} transition-metal diatomic molecule. International Journal of Quantum Chemistry, 20(S15):429–435, March 8– 14, 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory, Collision Phenomena, and Computational Quantum Chemistry.

Cvitas:1984:VSE

[CKM84] T. Cvitaš, L. Klasinc, and R. McDiarmid. Vibrational structures in the excited states of the ethylene radical cation. International Journal of Quantum Chemistry, 26(S18):537–546, March 1–15, 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Atomic, Molecular and Solid-State Theory, and Computational Quantum Chemistry.

Cvitas:1980:PSS

[CKNG80] T. Cvitaš, L. Klasinc, I. Novak, and H. Güsten. Photoelectron spectra of some perhalogenated methanes. International Journal of Quantum Chemistry, 18(S14):305–313, March 10– 20, 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory, Collision Phenomena, Quantum Statistics, and Computational Methods.

Chu:1980:CBI

[CL80] San-Yan Chu and Shyi-Long Lee. Comparison between intervalence charge transfer and molecular tautomerism. International Journal of Quantum Chemistry, 18(1):307–310, July 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Changying:1983:SAR

[CL83a] Chen Changying and Li Lemin. Structure–activity relationship and the mechanism of analgesia of the analgesic drug 3-methyl-fentanyl and its analogs. International Journal of Quantum Chemistry, 23(4):1597–1608, April 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Combs:1983:SMS

[CL83b] Leon L. Combs and Sten Lunell. SCF and MC–SCF study of CO_{2} with bond angle variation. International Journal of Quantum Chemistry, 23(2):465–472, February 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Cohen:1985:SCS

[CL85] Leon Cohen and Chongmoon Lee. A study of the Colle– Salvetti functional for ρ_{2} via an exactly soluble problem. International Journal of Quantum Chemistry, 28(S19):535–545, March 18–23, 1985. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Cohen:1986:CHP

[CL86] Leon Cohen and Chongmoon Lee. Correlation hole and physical properties: a model calculation. International Journal of Quantum Chemistry, 29(3):407–424, March 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Carbo:1980:HSM

[CLA80a] Ramon Carbó, Luis Leyda, and Mariano Arnau. How similar is a molecule to another? An electron density measure of similarity between two molecular structures. International Journal of Quantum Chemistry, 17(6):1185–1189, June 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Claverie:1980:SRC

[Cla80b] Pierre Claverie. Some remarks concerning the simulation of stochastic processes on digital computers. International Journal of Quantum Chemistry, 17(1):145–148, January 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Claverie:1983:IDE

[Cla83a] Pierre Claverie. Induction and dispersion energies including charge overlap effects: Green’s-function-type methods. International Journal of Quantum Chemistry, 23(4):1687–1688, April 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Cook:1983:ERA

[CLA83b] Charles M. Cook, Richard H. Lee, and Leland C. Allen. Electronic roles of active site residues in the catalysis and inhibition of carbonic anhydrase. International Journal of Quantum Chemistry, 24(S10):263–276, March 14–16, 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Clary:1986:TCR

[Cla86] D. C. Clary, editor. The theory of chemical reaction dynamics, volume 170 of NATO ASI series. Series C, Mathematical and physical sciences. D. Reidel, Dordrecht, Boston, Lancaster, Tokyo, 1986. ISBN 90-277-2202-1. LCCN QD501 .N355 1985.

Collins:1986:QMS

[CLL86] J. R. Collins, B. T. Luke, and G. H. Loew. Quantum mechanical study of putative intermediates in adduct formation by the suicide substrate ethylene with cytochrome P-450. International Journal of Quantum Chemistry, 30(S13):143–154, March 6–8, 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Chambaud:1980:PCT

[CLM80] G. Chambaud, B. Levy, and Ph. Millié. Perturbation calculation of transition moments in the ArH^{+} system. International Journal of Quantum Chemistry, 18(1):49–55, July 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Corbin:1982:MEP

[CLP82] Sylvie Corbin, Richard Lavery, and Bernard Pullman. The molecular electrostatic potential of DNA: The effect of countercation screening on various allomorphic forms. International Journal of Quantum Chemistry, 22(S9):103–110, March 4–6, 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Corbin:1983:MEP

[CLP83] Sylvie Corbin, Richard Lavery, and Bernard Pullman. Molecular electrotatic potential of a triple stranded helix: Poly(dT)·poly(dA)·poly(dT). International Journal of Quantum Chemistry, 23(4):1451–1461, April 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Chojnacki:1981:PEC

[CLS81] Henryk Chojnacki, Józef Lipiński, and W. Andrzej Sokalski. Potential energy curves in complementary base pairs and in model hydrogen bonded systems. International Journal of Quantum Chemistry, 20(2):339–346, August 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Chelikowsky:1984:TEM

[CLVC84] James R. Chelikowsky, Steven G. Louie, David Vanderbilt, and C. T. Chan. Total energy method for solids and solid surfaces. International Journal of Quantum Chemistry, 26(S18): 105–120, March 1–15, 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Atomic, Molecular and Solid-State Theory, and Computational Quantum Chemistry.

Cullen:1984:TCB

[CLZ84] John M. Cullen, William N. Lipscomb, and Michael C. Zerner. The three-center bond model for molecular electronic structure. International Journal of Quantum Chemistry, 26(S18): 439–448, March 1–15, 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Atomic, Molecular and Solid-State Theory, and Computational Quantum Chemistry.

Csavinszky:1981:TPD

[CM81] P. Csavinszky and R. A. Morrow. Theory of the potential of a donor ion in silicon. International Journal of Quantum Chemistry, 19(5):957–966, May 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Cisneros:1985:PDR

[CM85] Arturo Cisneros and Harold V. McIntosh. A proposed definition of resonant states. International Journal of Quantum Chemistry, 28(1):135–159, July 1985. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Cisneros:1984:AGC

[CMB84] Gerardo Cisneros, H. V. McIntosh, and Carlos F. Bunge. Automatic generation of configuration interaction Hamiltonian matrix elements. International Journal of Quantum Chemistry, 26(S18):683–689, March 1–15, 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Atomic, Molecular and Solid-State Theory, and Computational Quantum Chemistry.

Campos-Martinez:1989:QAS

[CMDBVMA89] J. Campos-Martinez, G. Delgado-Barrio, P. Villarreal, and S. Miret-Artes. A quantal approximation to the study of the collisional resonances in the NA–N_{2} system. International Journal of Quantum Chemistry, 36(3):369–377, September 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Chong:1983:CCE

[CMM83] Delano P. Chong, Tsutomu Minato, and Prasanta K. Mukherjee. Calculations of core-electron binding energies by transition operator method followed by Rayleigh–Schrödinger perturbation theory. International Journal of Quantum Chemistry, 23(6):1903–1914, June 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Cvitas:1987:PSS

[CNK87] T. Cvitaš, I. Novak, and L. Klasinc. Photoelectron spectra of some compounds containing the trifluoromethyl group. International Journal of Quantum Chemistry, 32(S21):737–742, March 12, 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Chemistry, Solid-State Theory, and Computational Methods.

Carter:1989:NAQ

[CNT89] Stuart Carter, Sonja Nikolić, and Nenad Trinajstić. A novel algorithm for QSAR. International Journal of Quantum Chemistry, 36(S16):323–330, April 1–8, 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Coffman:1983:GAT

[Cof83] R. E. Coffman. A general approximation to the two-center exchange integral between complex STOS. International Journal of Quantum Chemistry, 24(S17):207–216, March 3–12, 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-state Theory, Collision Phenomena, and Computational Quantum Chemistry.

Cohen:1983:FYP

[Coh83] Marvin L. Cohen. Fifty years of pseudopotentials. International Journal of Quantum Chemistry, 24(S17):583–595, March 3–12, 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-state Theory, Collision Phenomena, and Computational Quantum Chemistry.

Cohen:1986:HER

[Coh86a] Leon Cohen. Hierarchy equations for reduced density matrices in different representations. International Journal of Quantum Chemistry, 29(2):249–260, February 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Cohen:1986:PNM

[Coh86b] Marvin L. Cohen. Prediction of new materials and properties of solids. International Journal of Quantum Chemistry, 29 (4):843–854, April 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Conrad:1987:WMI

[Con87] Michael Conrad. The water-membrane interface as a substrate for H^{+} − H^{+} superflow. International Journal of Quantum Chemistry, 32(S14):167–188, March 14, 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Conrad:1988:E

[Con88a] Michael Conrad. Erratum. International Journal of Quantum Chemistry, 33(6):583–584, June 1988. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Conrad:1988:QMM

[Con88b] Michael Conrad. Quantum mechanics and molecular computing: Mutual implications. International Journal of Quantum Chemistry, 34(S15):287–301, March 12–19, 1988. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Syposium on Quantum Biology and Quantum Pharmacology.

Cook:1987:OEA

[Coo87] David B. Cook. Orbital energies: The aufbau principle and Koopman’s theorem in the UHF method. International Journal of Quantum Chemistry, 31(2):251–268, February 1987.

CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Chuljian:1982:IET

[COS82] D. T. Chuljian, J. Ozment, and J. Simons. Intramolecular energy transfer and predissociation in C-state HCN. International Journal of Quantum Chemistry, 22(S16):435–444, March 1–13, 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Chemistry, Theory of Condensed Matter, and Propagator Methods in the Quantum Theory of Matter.

Cowan:1982:SSB

[Cow82] Jack D. Cowan. Spontaneous symmetry breaking in large scale nervous activity. International Journal of Quantum Chemistry, 22(5):1059–1082, November 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Coxeter:1987:SIC

[Cox87] H. S. M. Coxeter. A simple introduction to colored symmetry. International Journal of Quantum Chemistry, 31(3):455–461, March 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Cooper:1980:ECE

[CP80a] Ian L. Cooper and Christopher N. M. Pounder. Electron correlation effects in SCF + CI wave functions—Fermi and Coulomb correlation holes in HCN. International Journal of Quantum Chemistry, 17(4):759–774, April 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Cross:1980:MHE

[CP80b] Thomas E. Cross and Ronald Pethig. Microwave Hall effect measurements on biopolymers. International Journal of Quantum Chemistry, 18(S7):385–395, March 5–8, 1980. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Cooper:1983:EMV

[CP83a] Ian L. Cooper and Christopher N. M. Pounder. Effect of modified virtual orbitals on local Coulomb correlation holes in FCN. International Journal of Quantum Chemistry, 23(1): 257–270, January 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Cross:1983:MSI

[CP83b] Thomas E. Cross and Ronald Pethig. Microwave studies of the interaction of DNA and water in the temperature range 90–300K. International Journal of Quantum Chemistry, 24 (S10):143–152, March 14–16, 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Chen:1985:CDP

[CP85a] Min-Bo Chen and Robert G. Parr. A constant-denominator perturbation theory using circulant orbitals. International Journal of Quantum Chemistry, 27(6):731–741, June 1985. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Curtiss:1985:TSI

[CP85b] L. A. Curtiss and J. A. Pople. Theoretical studies of the interaction of H_{2}O with small clusters of beryllium atoms. International Journal of Quantum Chemistry, 28(S19):613–628, March 18–23, 1985. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Campoy:1986:NSS

[CP86a] G. Campoy and A. Palma. On the numerical solution of the Schrödinger equation with a polynomial potential. International Journal of Quantum Chemistry, 30(S20):33–43, March 10–15, 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Syposium on Atomic, Molecular, and Solid-State Theory, Scattering Problems, Many Body Phenomena, and Computational Quantum Chemistry.

Cole:1986:OIM

[CP86b] Samuel J. Cole and George D. Purvis III. An optimizing interpreter for multidimensional products in computational chemistry. International Journal of Quantum Chemistry, 30(S20):

665–680, March 10–15, 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Syposium on Atomic, Molecular, and Solid-State Theory, Scattering Problems, Many Body Phenomena, and Computational Quantum Chemistry.

Chermette:1983:IRC

[CPGP83] H. Chermette, P. Pertosa, A. Goursot, and E. Penigault. Incidence of the relativistic corrections in electronic properties of 5d compounds: MS–X_{α} calculations on hexachloroiridates (III and IV) and tungsten trioxide. International Journal of Quantum Chemistry, 23(2):459–464, February 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Campoy:1986:BEP

[CPMP86] G. Campoy, E. Poulain, J. Morales, and A. Palma. Binding energies and peak intensities of light hydrocarbon molecules through TOM I. Methane and acetylene. International Journal of Quantum Chemistry, 30(S20):555–561, March 10–15, 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Syposium on Atomic, Molecular, and Solid-State Theory, Scattering Problems, Many Body Phenomena, and Computational Quantum Chemistry.

Castillo:1989:TSP

[CPN89] S. Castillo, E. Poulain, and O. Novaro. A theoretical study of the photochemistry of methylcopper hydride. Activation of the carbon-hydrogen bond. International Journal of Quantum Chemistry, 36(S23):509–520, April 1–8, 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Casaubon:1981:MTA

[CPS81] J. I. Casaubon, R. D. Piacentini, and A. Salin. Molecular treatment of atomic collisions in the keV/amu energy range: Application to the Li^{3+} − H system. International Journal of Quantum Chemistry, 20(S15):663–670, March 8–14, 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory, Collision Phenomena, and Computational Quantum Chemistry.

Campoy:1989:IBN

[CPS89] G. Campoy, A. Palma, and L. Sandoval. The inversion barrier in NH_{3}. International Journal of Quantum Chemistry, 36(S23):355–361, April 1–8, 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Cizek:1981:UCQ

[ČPV81] J. Čîžek, R. Pauncz, and E. R. Vrscay. The use of the concept of quasispin for the study of physical and chemical properties of alternant conjugated hydrocarbons. International Journal of Quantum Chemistry, 20(S15):61–62, March 8–14, 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory, Collision Phenomena, and Computational Quantum Chemistry.

Culberson:1986:ISS

[CPZS86] John C. Culberson, George D. Purvis III, Michael C. Zerner, and Barbara A. Seiders. An investigation into the specificity of soman analogues toward the antibody BE2 using electrostatic potentials. International Journal of Quantum Chemistry, 30 (S13):267–276, March 6–8, 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Cimiraglia:1981:SCH

[CR81] Renzo Cimiraglia and Raffaele Resta. Study of correlation holes. II. CI calculations on model polyatomic systems. International Journal of Quantum Chemistry, 19(2):301–318, February 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Callaway:1980:STD

[CS80a] Joseph Callaway and S. P. Singhal. Scattering theory of defects in solids: Theory and application to the atomic vacancy. International Journal of Quantum Chemistry, 18(S14):651– 663, March 10–20, 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory, Collision Phenomena, Quantum Statistics, and Computational Methods.

Chalasinaski:1980:DSA

[CS80b] Grezegorz Chałasińaski and Krzysztof Szalewicz. Degenerate symmetry-adapted perturbation theory. Convergence properties of perturbation expansions for excited states of H_{2}^{+} ion. International Journal of Quantum Chemistry, 18(4):1071–1089, October 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Chatterjee:1981:MOS

[CS81] C. L. Chatterjee and Anil Saran. Molecular orbital studies on nucleoside antibiotics V. Conformation of pyrazofurins. International Journal of Quantum Chemistry, 20(S8):129–138, March 5–7, 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Colle:1982:RUA

[CS82] Renato Colle and Oriano Salvetti. Remarks on the use of the average occupation numbers in the SCF process for open shell systems. International Journal of Quantum Chemistry, 22(2): 289–292, August 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Chaturvedi:1983:RWT

[CS83a] Meena Chaturvedi and Vipin Srivastava. Random-walk theory for localization. International Journal of Quantum Chemistry, 23(4):1463–1468, April 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Chojnacki:1983:IPB

[CS83b] H. Chojnacki and A. Sawaryn. Interaction of p-benzoquinone and mitomycin C with complementary base pairs. International Journal of Quantum Chemistry, 23(2):729–738, February 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Chuljian:1983:CRS

[CS83c] David T. Chuljian and Jack Simons. Coordinate rotation studies of H^{−}, He^{−}, Be^{−}, Mg^{−} resonances: Basis set and configuration list dependence. International Journal of Quantum Chemistry, 23(5):1723–1738, May 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Clabo:1987:FPF

[CS87] D. Allen Clabo, Jr. and Henry F. Schaefer 3rd. Fluorine peroxide (FOOF): a continuing problem for normally reliable theoretical methods. International Journal of Quantum Chemistry, 31(3):429–433, March 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Chermette:1989:ESD

[CS89] Henry Chermette and Dennis R. Salahub. Eigenvalue search by diagonalization in the scattered wave method. International Journal of Quantum Chemistry, 35(4):577–580, April 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Csavinszky:1980:VST

[Csa80] P. Csavinszky. Variational solution of the Thomas–Fermi equation for compressed atoms at high temperatures. International Journal of Quantum Chemistry, 18(S14):131–134, March 10–20, 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory, Collision Phenomena, Quantum Statistics, and Computational Methods.

Csavinszky:1981:ISS

[Csa81] P. Csavinszky. Introduction of the shell structure with gradient expansion corrections into the Thomas–Fermi–Dirac energy density functional for neutral atoms. International Journal of Quantum Chemistry, 20(S15):387–391, March 8–14, 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory, Collision Phenomena, and Computational Quantum Chemistry.

Csavinszky:1982:TED

[Csa82] P. Csavinszky. Theory of the electron distribution in compressed atoms at high temperatures. International Journal of Quantum Chemistry, 21(5):899–903, May 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Csavinszky:1983:VCP

[Csa83] P. Csavinszky. Variational calculation of the potential resulting from a localized arbitrary spherically symmetric charge distribution. International Journal of Quantum Chemistry, 24(S17):537–540, March 3–12, 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-state Theory, Collision Phenomena, and Computational Quantum Chemistry.

Csavinszky:1985:TAB

[Csa85] P. Csavinszky. Total atomic binding energy via the density functional theory. International Journal of Quantum Chemistry, 28(S19):559–565, March 18–23, 1985. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Csavinszky:1986:ILT

[Csa86a] P. Csavinszky. Investigation of the Lieb–Thirring lower bound to the electronic kinetic energy via an energy–density functional approach. International Journal of Quantum Chemistry, 30(S20):271–276, March 10–15, 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Syposium on Atomic, Molecular, and Solid-State Theory, Scattering Problems, Many Body Phenomena, and Computational Quantum Chemistry.

Csavinszky:1986:VAS

[Csa86b] P. Csavinszky. A variational approach to the solution of Poisson’s equation in prolate spheroidal coordinates. International Journal of Quantum Chemistry, 30(3):305–309, September 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Csavinszky:1987:TAB

[Csa87] P. Csavinszky. Total atomic binding energy via the density functional theory. International Journal of Quantum Chemistry, 32(3):407–408, September 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Csavinszky:1988:CPW

[Csa88] P. Csavinszky. Comparison of the Perdew–Wang exchange with the Dirac exchange. International Journal of Quantum Chemistry, 33(4):265–269, April 1988. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Csavinszky:1989:VDF

[Csa89] P. Csavinszky. A variational density-functional calculation of the total atomic binding energy with recently proposed kinetic-energy and exchange-energy functionals. International Journal of Quantum Chemistry, 36(S23):591–597, April 1–8, 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Cameron:1986:MAC

[CSB86a] Alan D. Cameron, Vedene H. Smith, Jr., and Michael C. Baird. On the mechanism of activation of coordinated olefins toward nucleophilic attack. International Journal of Quantum Chemistry, 30(S20):657–663, March 10–15, 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Syposium on Atomic, Molecular, and Solid-State Theory, Scattering Problems, Many Body Phenomena, and Computational Quantum Chemistry.

Cole:1986:NDP

[CSB86b] Samuel J. Cole, Krzysztof Szalewicz, and Rodney J. Bartlett. Nitromethane dimer potential energy surface studies. International Journal of Quantum Chemistry, 30(5):695–711, November 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Cohen:1980:RDM

[CSF80] L. Cohen, P. Santhanam, and C. Frishberg. Reduced density matrix equations and functionals for correlation energy. International Journal of Quantum Chemistry, 18(S14):143–154, March 10–20, 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory, Collision Phenomena, Quantum Statistics, and Computational Methods.

Csizmadia:1982:MSC

[Csi82] I. G. Csizmadia, editor. Molecular structure and conformation: recent advances, volume 3 of Progress in theoretical organic chemistry. Elsevier, Amsterdam, The Netherlands, 1982. ISBN 0-444-42089-4. 344 pp. LCCN QD461 .M628 1982. fl 220.00 (est.).

Cieplak:1987:AQM

[CSK87] Piotr Cieplak, U. Chandra Singh, and Peter A. Kollman. Application of the quantum mechanics and free energy perturbation methods to study molecular processes. International Journal of Quantum Chemistry, 32(S14):65–74, March 14, 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Csavinszky:1981:AVP

[CT81] P. Csavinszky and C. E. Tarr. Addendum to “Variational principle for obtaining approximate analytical solutions of the temperature-perturbed Thomas–Fermi equation for compressed atoms”. International Journal of Quantum Chemistry, 19(3):477–481, March 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Craig:1984:MQE

[CT84] D. P. Craig and T. Thirunamachandran. Molecular quantum electrodynamics: an introduction to radiation-molecule interactions, volume 8 of Theoretical chemistry. Academic Press, New York, USA, 1984. ISBN 0-12-195080-8. xi + 324 pp. LCCN QC680 .C73 1984.

Craig:1987:CAM

[CT87] D. P. Craig and T. Thirunamachandran. Chiral-achiral molecular coupling: The effect on the chiral partner. International Journal of Quantum Chemistry, 31(3):417–424, March 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Craig:1998:MQE

[CT98] D. P. Craig and T. Thirunamachandran. Molecular quantum electrodynamics: an introduction to radiation-molecule interactions. Dover Publications, Inc., New York, NY, USA, 1998. ISBN 0-486-40214-2 (paperback). ix + 324 pp. LCCN QC680 .C73 1998. URL http://www.loc.gov/

catdir/description/dover033/98012709.htm.

Chatterjee:1983:CBC

[CTP83] R. Chatterjee, J. A. Tuszyński, and R. Paul. Comments on the Bose condensation of phonons in the biological systems. International Journal of Quantum Chemistry, 23(2):709–712, February 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Csavinszky:1980:VPO

[CTV80] P. Csavinszky, C. E. Tarr, and F. Vosman. Variational principle for obtaining approximate analytical solutions of the temperature-perturbed Thomas–Fermi equation for compressed atoms. International Journal of Quantum Chemistry, 17(4):737–745, April 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Carsky:1980:ICM

[ČU80] Petr Čársky and Miroslav Urban. Ab initio calculations: methods and applications in chemistry, volume 16 of Lecture notes in chemistry. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1980. ISBN 3-540-10005-9, 0-387-10005-9. vi + 247 pp. LCCN QD462.

Cullen:1988:FOC

[Cul88] J. M. Cullen. Fifth-order constant denominator perturbation therory within a localized bond model: Contributions from single and double excitations. International Journal of Quantum Chemistry, 33(6):497–527, June 1988. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Cullen:1989:FOC

[Cul89] J. M. Cullen. Fifth-order constant denominator perturbation theory within a localized bond model: Contributions from triple and quadruple excitations. International Journal of Quantum Chemistry, 35(4):519–549, April 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Castro:1982:OCS

[CV82a] Eduardo A. Castro and Enrique J. Vasini. Omission in a comparative study of the structure and properties of ClF_{2} and Cl_{3} radicals by CNDO/2 and INDO methods. International Journal of Quantum Chemistry, 22(2):433–434, August 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Cizek:1982:LOP

[ČV82b] Jiří Čížek and Edward R. Vrscay. Large order perturbation theory in the context of atomic and molecular physics— interdisciplinary aspects. International Journal of Quantum Chemistry, 21(1):27–68, January 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Cizek:1983:CFS

[ČV83a] J. Čížek and E. R. Vrscay. Continued fractions for the Stirling expansion revisited. International Journal of Quantum Chemistry, 24(5):521–522, November 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Csavinszky:1983:LBW

[CV83b] P. Csavinszky and F. Vosman. Lower bounds to the Weizsäcker correction in statistical models for the Na atom with shell structure. International Journal of Quantum Chemistry, 24(1):61–64, July 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Cizek:1985:IPP

[CV85] J. Cizek and E. R. Vrscay. Inner projection with and without perturbation theory: The anharmonic oscillator revisited and the quadratic Zeeman effect in ground-state hydrogen. International Journal of Quantum Chemistry, 28(6):665–686, December 1985. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Cizek:1986:LBG

[ČV86] J. Čížek and E. R. Vrscay. Lower bounds to ground state eigenvalues of the Schrödinger equation via optimized inner projection: Application to quartic and sextic anharmonic oscillators. International Journal of Quantum Chemistry, 30 (S20):65–72, March 10–15, 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Syposium on Atomic, Molecular, and Solid-State Theory, Scattering Problems, Many Body Phenomena, and Computational Quantum Chemistry.

Cizek:1988:SCQ

[ČV88] J. Čížek and F. Vinette. Symbolic computation in quantum mechanics. Several simple examples. International Journal of Quantum Chemistry, 34(S22):537–548, March 12–19, 1988. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Syposium on Quantum Chemistry, Solid-State Theory, and Computational Methods.

Cizek:1987:RIP

[ČVV87] Jiří Čížek, Francine Vinette, and Edward R. Vrscay. Renormalized inner projection, symbolic computation, and Löwdin rational approximants in explicit form. International Journal of Quantum Chemistry, 32(S21):757–758, March 12, 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Chemistry, Solid-State Theory, and Computational Methods.

Case:1980:SEA

[CY80] David A. Case and Cary Y. Yang. Stable and efficient algorithms for Xα multiple scattering calculations. International Journal of Quantum Chemistry, 18(4):1091–1099, October 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Cheney:1981:EPE

[CZ81] B. Vernon Cheney and Dominic A. Zichi. Effects of Nprotonation on the electronic structure of morphine and naloxone. International Journal of Quantum Chemistry, 20(S8): 201–219, March 5–7, 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Cullen:1982:FOC

[CZ82a] John M. Cullen and Michael C. Zerner. Fourth-order constant denominator perturbation theory. International Journal of Quantum Chemistry, 22(S16):301–317, March 1–13, 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Chemistry, Theory of Condensed Matter, and Propagator Methods in the Quantum Theory of Matter.

Cullen:1982:LBM

[CZ82b] John M. Cullen and Michael C. Zerner. Localized bond model for molecular energy to fourth order in perturbation theory. International Journal of Quantum Chemistry, 22(3):497–535, September 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Cullen:1988:EPV

[CZ88] John M. Cullen and Michael C. Zerner. An examination of perturbation–variational theory and scaling at fifth order. International Journal of Quantum Chemistry, 33(5):373–394, May 1988. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Czerminski:1987:IIR

[Cze87] Ryszard Czermiński. Intermolecular interactions: Reliability of results obtained for small hydrogen bonded dimers by Claverie’s method. International Journal of Quantum Chemistry, 31(4):649–662, April 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Cheney:1983:ECF

[CZM83] B. Vernon Cheney, Dominic A. Zichi, and Andrea B. Miller. Electronic and conformational features of a series of Nprotonated benzomorphan molecules: N-substituent effects on relative binding at the agonist and antagonist states of the opioid receptor. International Journal of Quantum Chemistry, 24(S10):43–59, March 14–16, 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Daudel:1985:SDM

[D^{+}85] Raymond Daudel et al., editors. Structure and dynamics of molecular systems. D. Reidel, Dordrecht, Boston, Lancaster, Tokyo, 1985. ISBN 90-277-1977-2 (vol. 1), 90-277-2246-3 (vol. 2). ???? pp. LCCN QD461 .S924 1985. US$40.00.

Defina:1980:EMS

[DA80] J. A. Defina and P. R. Andrews. Evaluation of MINDO /3 calculations in simple amides, ureas, and heterocycles of biological interest. International Journal of Quantum Chemistry, 18(3):797–810, September 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Durani:1981:PBS

[DA81] S. Durani and Nitya Anand. A possible basis for structure– function relationship of estrogens. International Journal of Quantum Chemistry, 20(1):71–83, July 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Dixon:1988:SSE

[DA88] David A. Dixon and Anthony J. Arduengo III. Solvent stabilization of the edge inversion transition state in tetrahedral molecules. International Journal of Quantum Chemistry, 34 (S22):85–98, March 12–19, 1988. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Syposium on Quantum Chemistry, Solid-State Theory, and Computational Methods.

Dalgaard:1980:VPT

[Dal80] Esper Dalgaard. Variation principle for the time dependence of density operators and its relation to linear decoupling procedures for propagators. International Journal of Quantum Chemistry, 17(3):449–458, March 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

DaSilva:1989:HPS

[DAL89] E. C. F. Da Silva, L. V. C. Assali, and J. R. Leite. Hydrogen passivation of shallow donors in silicon. International Journal of Quantum Chemistry, 36(S23):693–699, April 1–8, 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Delhalle:1981:PUU

[DAPS81] J. Delhalle, J. M. Andre, L. Piela, and L. Z. Stolarczyk. Potential use of the unit cell arbitrary concept to evaluate longrange interactions in quantum mechanical studies of model polymers. International Journal of Quantum Chemistry, 20 (S15):217, March 8–14, 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory, Collision Phenomena, and Computational Quantum Chemistry.

Das:1980:REE

[Das80] M. P. Das. Relativistic exchange energies for heavy atoms. International Journal of Quantum Chemistry, 18(S14):67–70, March 10–20, 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory, Collision Phenomena, Quantum Statistics, and Computational Methods.

Das:1982:CRO

[Das82] M. P. Das. Calculation of relaxed orbital binding energies in atoms: Zinc and cadmium. International Journal of Quantum Chemistry, 21(5):845–849, May 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Datta:1982:CNI

[Dat82] Sunil K. Datta. Changes in the near IR spectra of CoXF_{6}, 6R_{2}O (X = Si, Ti; R = H, D) during phase transition. International Journal of Quantum Chemistry, 22(S16):667, March 1–13, 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Chemistry, Theory of Condensed Matter, and Propagator Methods in the Quantum Theory of Matter.

Datzeff:1983:SE

[Dat83] A. B. Datzeff. Structure of the electron. International Journal of Quantum Chemistry, 23(1):81–84, January 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Datseff:1985:NSE

[Dat85] A. B. Datseff. On the nonlinear Schrödinger equation. International Journal of Quantum Chemistry, 28(S19):739–740, March 18–23, 1985. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Davidson:1981:QSE

[Dav81] Ernest R. Davidson. Qualitatively significant effects of electron correlation. International Journal of Quantum Chemistry, 20(S15):65–68, March 8–14, 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory, Collision Phenomena, and Computational Quantum Chemistry.

Davydov:1985:SMS

[Dav85] A. S. (Aleksandr Sergeevich) Davydov. Solitons in molecular systems. Mathematics and its applications (Soviet series). D. Reidel, Dordrecht, Boston, Lancaster, Tokyo, 1985. ISBN 90-277-1854-7. xviii + 319 pp. LCCN QC174.26.W28 D38 1985.

Davies:1987:DRM

[Dav87] R. H. Davies. Drug and receptors in molecular biology. International Journal of Quantum Chemistry, 32(S14):221–243, March 14, 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Davydov:1991:SMS

[Dav91] A. S. (Aleksandr Sergeevich) Davydov. Solitons in molecular systems, volume 61 of Mathematics and its applications (Soviet series). Kluwer Academic Publishers, Norwell, MA, USA, and Dordrecht, The Netherlands, second edition, 1991. ISBN 0-7923-1029-2. xiii + 413 pp. LCCN QC174.26.W28 D3813 1990.

Daudey:1981:ACE

[DB81] J. P. Daudey and M. Berrondo. Atomic calculations with an effective one-body potential. International Journal of Quantum Chemistry, 19(5):907–927, May 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

deBruijn:1984:AIS

[dB84] Steef de Bruijn. Analysis of the inadequacies of some semiempirical MO methods as theories of structure and reactivity. International Journal of Quantum Chemistry, 25(2):367–390, February 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Dunne:1984:MTC

[DBC84] Lawrence J. Dunne, George M. Bell, and Leon L. Combs. Molecular theory of critical phenomena in aliphatic carboxylic acid monolayers. International Journal of Quantum Chemistry, 25(5):795–808, May 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Davies:1986:APP

[DBCJ86] R. H. Davies, R. D. Bagnall, B. Crooks, and W. G. M. Jones. Anesthetic potency and the perturbation of positive ion or proton acceptor sites in the nerve membrane. International Journal of Quantum Chemistry, 30(S13):53–70, March 6–8, 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Delhalle:1985:EGL

[DBD^{+}85] J. Delhalle, V. P. Bodart, M. Dory, J. M. André, and J. Zyss. Equilibrium geometry and longitudinal electric polarizability of allene, diallene, and iriallene: An ab initio study. International Journal of Quantum Chemistry, 28(S19):313–321, March 18–23, 1985. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Douady:1980:PCI

[DBES80] J. Douady, V. Barone, Y. Ellinger, and R. Subra. Perturbative configuration interaction using localized orbitals in the INDO hypothesis. I. Theory and applications to energetical problems. International Journal of Quantum Chemistry, 17 (2):211–233, February 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Doll:1989:CMC

[DBF89] J. D. Doll, Thomas L. Beck, and David L. Freeman. Classical Monte Carlo dynamics: A simulated annealing approach to the construction of double-ended classical trajectories. International Journal of Quantum Chemistry, 36(S23):73–78, April 1–8, 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

DelRe:1984:MEP

[DBL84] Giuseppe Del Re, Vincenzo Barone, and Francesco Lelj. The mechanisms of elementary physicochemical processes: An introductory report. International Journal of Quantum Chemistry, 26(5):563–591, November 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

DelRe:1980:ESM

[DBS80] Giuseppe Del Re, Gaston Berthier, and Josiane Serre. Electronic states of molecules and atom clusters: foundations and prospects of semiempirical methods, volume 13 of Lecture notes in chemistry. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1980. ISBN 3-540-09738-4, 0-387-09738-4. LCCN QD461.

De:1982:CIS

[DBS82] B. R. De, S. Bhattacharjee, and A. B. Sannigrahi. CNDO/2 and INDO studies of structure and relative stability of a series of H AB–H BA systems of mixed first- and second-row composition. International Journal of Quantum Chemistry, 21 (3):665–669, March 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Delgado-Barrio:1985:VPH

[DBVMB85] G. Delgado-Barrio, P. Villarreal, P. Mareca, and J. A. Beswick. Vibrational predissociation of highly vibrationally excited van der Waals molecules: a quantal close-coupling calculation. International Journal of Quantum Chemistry, 27(2): 173–180, February 1985. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Dorfman:1982:VEE

[DC82a] J. R. Dorfman and E. G. D. Cohen. The virial expansion in equilibrium and nonequilibrium statistical mechanics. International Journal of Quantum Chemistry, 22(S16):63–81, March 1–13, 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Chemistry, Theory of Condensed Matter, and Propagator Methods in the Quantum Theory of Matter.

Dunne:1982:TMT

[DC82b] Lawrence J. Dunne and Leon Combs. Towards a molecular theory of phase transitions in fatty acid monolayers. International Journal of Quantum Chemistry, 21(1):239–250, January 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Dunlap:1986:LXC

[DC86] Brett I. Dunlap and Michael Cook. LCAO-Xα calculations of rotational energy barriers—prototypes of chemical reactions. International Journal of Quantum Chemistry, 29(4):767–777, April 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Delhalle:1987:PAR

[DC87] Joseph Delhalle and Jean-Louis Calais. Pathological aspects of restricted Hartree–Fock band calculations for metallic chains. International Journal of Quantum Chemistry, 32(S21):115– 129, March 12, 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Chemistry, Solid-State Theory, and Computational Methods.

Donati:1982:CTM

[DCF82a] E. Donati, E. A. Castro, and F. M. Fernández. Calculation of total molecular energies from an approximate relation between SCF orbital, atomic electronic repulsion, and total SCF energies. International Journal of Quantum Chemistry, 22 (2):429–431, August 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Donnamaria:1982:CDS

[DCF82b] Maria C. Donnamaria, Eduardo A. Castro, and Francisco M. Fernández. Calculation of diamagnetic susceptibilities from the Thomas–Fermi–Amaldi theory for ions. International Journal of Quantum Chemistry, 22(5):1005–1011, November 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Delhalle:1987:TFN

[DD87a] Joseph Delhalle and Mireille Defranceschi. Toward fully numerical evaluation of momentum space Hartree–Fock wave functions. Numerical experiments on the He atom. International Journal of Quantum Chemistry, 32(S21):425–433, March 12, 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Chemistry, Solid-State Theory, and Computational Methods.

Dence:1987:IPC

[DD87b] Joseph B. Dence and Dennis J. Diestler. Intermediate physical chemistry: stationary properties of chemical systems. John Wiley, New York, NY, USA, 1987. ISBN 0-471-81243-9. xi + 344 pp. LCCN QD453.2 .D46 1987.

Duncan:1989:ADB

[DD89] Richard H. Duncan and G. S. Davies. Alkylation of DNA bases by carcinogenic N-nitrosamine metabolites: a theoretical study. International Journal of Quantum Chemistry, 35 (5):665–677, May 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

deAndradaESilva:1986:VAM

[ddF86] E. A. de Andrada E. Silva, I. C. da Cunha Lima, and A. Ferreira da Silva. The variational alternant molecular orbital method for 3-D and 2-D hydrogen molecule. International Journal of Quantum Chemistry, 30(5):685–694, November 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Dory:1987:EGE

[DDFA87] Magdalena Dory, Joseph Delhalle, Joseph G. Fripiat, and Jean-Marie Andre. Equilbrium geometry and electrical polarizability of formic acid, formamide and their cyclic hydrogenbonded pairs. International Journal of Quantum Chemistry, 32(S14):85–103, March 14, 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

DaCunhaLima:1981:SCAa

[DDP81a] I. C. Da Cunha Lima, A. Ferreira Da Silva, and N. J. Parada. Self-consistent APW–k · p method. I. Theory. International Journal of Quantum Chemistry, 19(4):681–691, April 1981.

CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

DaCunhaLima:1981:SCAb

[DDP81b] I. C. Da Cunha Lima, A. Ferreira Da Silva, and N. J. Parada. Self-consistent APW–k · p method. II. Application to NaCl. International Journal of Quantum Chemistry, 20(4):933–949, October 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

DeHosson:1980:ESN

[De 80] J. Th. M. De Hosson. Electronic states near dislocations in transition metals: An application of quantum chemistry in technology. International Journal of Quantum Chemistry, 18 (2):575–582, August 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

DeBrouckere:1983:OEP

[De 83] G. De Brouckére. One-electron properties at the SCF ab Initio level. International Journal of Quantum Chemistry, 23(4): 1677, April 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Deak:1983:CSP

[Deá83] P. Deák. CNDO /S parameters of C, O, and Si for solid state calculations. International Journal of Quantum Chemistry, 23 (4):1165–1177, April 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Deal:1989:LEI

[Dea89] Walter J. Deal. Local energy as an indicator of the accuracy of wave functions and probability densities for helium. International Journal of Quantum Chemistry, 35(4):513–518, April 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

DelRe:1981:BUN

[Del81a] Giuseppe Del Re. Binding: a unifying notion or a pseudoconcept? International Journal of Quantum Chemistry, 19 (6):979–984, June 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

DelRe:1981:I

[Del81b] Giuseppe Del Re. Introduction. International Journal of Quantum Chemistry, 19(6):975–976, June 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Delos:1983:QTS

[Del83] J. B. Delos. Quantum theory of slow atomic collisions. International Journal of Quantum Chemistry, 24(S17):37–48, March 3–12, 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-state Theory, Collision Phenomena, and Computational Quantum Chemistry.

DelRe:1984:I

[Del84a] Giuseppe Del Re. Introduction. International Journal of Quantum Chemistry, 26(5):v–vi, November 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Delhalle:1984:CLD

[Del84b] J. Delhalle. Convergence of LCAO density matrices in Hartree—Fock calculations on extended model chains. International Journal of Quantum Chemistry, 26(5):717–723, November 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

DelBene:1987:BSC

[Del87] Janet E. Del Bene. Basis set and correlation effects on computed negative ion hydrogen bond energies of the complexes AH_{n} · AH_{n−1}^{−1}: AH_{n}NH_{3}, OH_{2}, and FH. International Journal of Quantum Chemistry, 32(S14):27–35, March 14, 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

DelBene:1988:IMO

[Del88] Janet E. Del Bene. An ab initio molecular orbital study of hydrogen bonding and ion-molecule association in model systems for DNA bases. International Journal of Quantum Chemistry, 34(S15):119–135, March 12–19, 1988. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Syposium on Quantum Biology and Quantum Pharmacology.

Demiralp:1983:EFO

[Dem83] Metin Demiralp. Exponential factor optimization in characteristic function method for electrically charged particles. International Journal of Quantum Chemistry, 23(3):779–787, March 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Demiralp:1986:NAA

[Dem86] Metin Demiralp. A new algebraic approach to the eigenvalue problems of linear differential operators without integrations. International Journal of Quantum Chemistry, 29(2):221–227, February 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Deng:1983:MHF

[Den83] Conghao Deng. The modified Hartree–Fock self-consistent field equation. International Journal of Quantum Chemistry, 23(6):1979–1987, June 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Deng:1985:MHF

[Den85] Conghao Deng. The modified Hartree–Fock self-consistent field equation. International Journal of Quantum Chemistry, 27(2):233–234, February 1985. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Dewar:1988:UQM

[Dew88] Michael J. S. Dewar. Use of quantum mechanical models in studies of reaction mechanisms. International Journal of Quantum Chemistry, 34(S22):557–566, March 12–19, 1988. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Syposium on Quantum Chemistry, Solid-State Theory, and Computational Methods.

Dick:1983:ABP

[DF83a] Bernhard Dick and Hans-Joachim Freund. Analysis of bonding properties in molecular ground and excited states by a Cohen-type bond order. International Journal of Quantum Chemistry, 24(6):747–765, December 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Duke:1983:ESP

[DF83b] C. B. Duke and W. K. Ford. The electronic structure of polymers: Relation to molecular architecture. International Journal of Quantum Chemistry, 24(S17):597–608, March 3– 12, 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-state Theory, Collision Phenomena, and Computational Quantum Chemistry.

Ding:1986:IHB

[DF86] Yanbo Ding and Xiaoyuan Fu. Investigation of the hydrogen bonds between formaldehyde and hydrogen halides by pseudopotential Ab Initio method. International Journal of Quantum Chemistry, 30(1):137–142, July 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

DelBene:1989:ISS

[DF89] Janet E. Del Bene and Michael J. Frisch. An ab initio study of the structures and stabilities of the complexes of the bases N_{2}O, CO_{2}, and CO with the acids FH, H^{+}, and Li^{+}. International Journal of Quantum Chemistry, 36(S23):371–380, April 1–8, 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Delhalle:1984:ECA

[DFH84] Joseph Delhalle, Joseph G. Fripiat, and Frank E. Harris. Evaluation of computational aspects of a modified CS–LCAO– SCF–CO strategy for electronic structure calculations of extended model chains. International Journal of Quantum Chemistry, 26(S18):141–152, March 1–15, 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Atomic, Molecular and Solid-State Theory, and Computational Quantum Chemistry.

Delhalle:1980:ULT

[DFP80] Joseph Delhalle, Joseph G. Fripiat, and Lucjan Piela. On the use of Laplace transform to evaluate one-dimensional lattice summations in quantum calculations of model polymers. International Journal of Quantum Chemistry, 18(S14):431–442, March 10–20, 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory, Collision Phenomena, Quantum Statistics, and Computational Methods.

Deb:1983:NMD

[DG83] B. M. Deb and S. K. Ghosh. New method for the direct calculation of electron density in many-electron systems. I. Application to closed-shell atoms. International Journal of Quantum Chemistry, 23(1):1–26, January 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Darvesh:1985:CSM

[DG85] K. V. Darvesh and F. Grein. Configuration selection in the MCSCF method. I. Application to the B^{1}Σ^{+} state of HF. International Journal of Quantum Chemistry, 28(S19):247–256, March 18–23, 1985. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Daul:1986:CLF

[DG86] C. Daul and A. Goursot. The calculation of ligand-field multiplet states from multiple-scattering Xα wave functions. International Journal of Quantum Chemistry, 29(4):779–792, April 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Diercksen:1985:SST

[DGSO85] G. H. F. Diercksen, N. E. Grüner, John R. Sabin, and Jens Oddershede. structures and spectra of triatomic silicon-carbon compounds. International Journal of Quantum Chemistry, 28 (S19):735–736, March 18–23, 1985. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Dinesha:1984:CIM

[DH84] K. V. Dinesha and Juergen Hinze. Configuration interaction matrix elements for atoms using permutation group algebra. International Journal of Quantum Chemistry, 26(4):507–519, October 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Delhalle:1985:ACS

[DH85] Joseph Delhalle and Frank E. Harris. An alternative computational strategy for direct space Hartree–Fock calculations on extended model chains. International Journal of Quantum Chemistry, 27(2):219–229, February 1985. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Diarmid:1986:LVS

[Dia86] Ruth Mc Diarmid. The lower valence states of linear polyenes: Locations and identifications. International Journal of Quantum Chemistry, 29(4):875–881, April 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Dinesha:1986:EEA

[DIV86] K. V. Dinesha, Suman B. Iyer, and Saraswathi Vishveshwara. Energy expressions for atomic configurations in the L–S coupling scheme. International Journal of Quantum Chemistry, 30(6):783–790, December 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Dyadyusha:1981:ASFa

[DK81a] G. G. Dyadyusha and E. S. Kryachko. Algebraic structure of fermion density matrices. I. International Journal of Quantum Chemistry, 19(2):251–257, February 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Dyadyusha:1981:ASFb

[DK81b] G. G. Dyadyusha and Eugene S. Kryachko. Algebraic structure of fermion density matrices. II. International Journal of Quantum Chemistry, 19(4):505–514, April 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Denner:1982:PTB

[DK82a] V. Denner and F. Kaiser. Phase transition behavior of a greater membrane model. International Journal of Quantum Chemistry, 22(S9):41–57, March 4–6, 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Duch:1982:SGG

[DK82b] Wlodzislaw Duch and Jacek Karwowski. Symmetric group graphical approach to the direct configuration interaction method. International Journal of Quantum Chemistry, 22(4): 783–824, October 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

DaSilva:1987:IHM

[DK87] E. Z. Da Silva and C. A. Kuhnen. The influence of hydrogen in the magnetism of Pd_{3}Fe. International Journal of Quantum Chemistry, 32(S21):99–103, March 12, 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Chemistry, Solid-State Theory, and Computational Methods.

Danilov:1981:EWT

[DKLP81] Victor I. Danilov, Józef S. Kwiatkowski, Bogdan Lesyng, and Valery I. Poltev. Effect of water on the tautomeric equilibrium of 2-oxopyridine. A Monte Carlo simulation study. International Journal of Quantum Chemistry, 20(S8):359–364, March 5–7, 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Deshmukh:1981:EEM

[DL81] Pranawa Deshmukh and Jan Linderberg. Extension of EWMO model to particle-hole excitations. International Journal of Quantum Chemistry, 19(4):575–584, April 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Day:1982:CCC

[DL82] R. S. Day and J. Ladik. Cluster CPA calculations for infinite and half–infinite chains and applications to coupled polymer chains. International Journal of Quantum Chemistry, 21 (5):917–926, May 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Deumens:1983:HAS

[DL83] E. Deumens and L. Lathouwers. Highly accurate solutions of generator coordinate nuclear motion equations. International Journal of Quantum Chemistry, 24(S17):461–469, March 3– 12, 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-state Theory, Collision Phenomena, and Computational Quantum Chemistry.

Deumens:1984:GCA

[DLVÖ84] E. Deumens, L. Lathouwers, P. Van Leuven, and Y. Öhrn. The generator coordinate approximation for H_{2}^{+}. International Journal of Quantum Chemistry, 26(S18):339–346, March 1– 15, 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Atomic, Molecular and Solid-State Theory, and Computational Quantum Chemistry.

Davies:1983:LRS

[DM83a] R. H. Davies and T. R. Morris. β-lactam ring stability and antibacterial potency: a novel eigenvalue problem. International Journal of Quantum Chemistry, 23(4):1385–1405, April 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Davis:1983:SWF

[DM83b] C. L. Davis and E. N. Maslen. Series wave functions for the helium atom. International Journal of Quantum Chemistry, 24(S17):217–225, March 3–12, 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-state Theory, Collision Phenomena, and Computational Quantum Chemistry.

Dagata:1985:TNI

[DM85] J. A. Dagata and S. P. McGlynn. Treatment of nonadiabatic interactions in polyatomic molecules by the frame transformation method: Application to the methyl iodide spectrum. International Journal of Quantum Chemistry, 28(S19):323– 332, March 18–23, 1985. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Das:1984:CPS

[DMB^{+}84] Kalyan Kumar Das, Debashis Mukherjee, Sankar Prasad Bhattacharyya, Debendra Madhab Bhattacharyya, and Priyotosh Khan. Convergence problems in SCF calculations: Further applications of a new technique based on the use of inverse Fock operator. International Journal of Quantum Chemistry, 25(5):809–816, May 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Das:1986:SFS

[DMB86] Kalyan Kumar Das, D. Mukherjee, and S. P. Bhattacharyya. Structural features of some small carbonyls in the^{1,3}nπ states: An MCSCF study. International Journal of Quantum Chemistry, 29(6):1745–1754, June 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Das:1989:SSF

[DMB89] Kalyan K. Das, D. Mukherjee, and S. P. Bhattacharyya. Spectroscopic and structural features of small thiocarbonyl molecules in low-lying excited states: Further applications of a variant of the orthogonal gradient method. International Journal of Quantum Chemistry, 35(4):483–494, April 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Duke:1984:SIM

[DO84] B. J. Duke and Brian O’leary. The simulated ab initio molecular orbital (SAMO) method. VIII. A study of linear polyfluoroethylenes and polychloroethylenes. International Journal of Quantum Chemistry, 26(S18):407–432, March 1–15, 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Atomic, Molecular and Solid-State Theory, and Computational Quantum Chemistry.

Deumens:1989:GHA

[DÖ89] E. Deumens and Y. Öhrn. General harmonic approximation for time-dependent molecular dynamics. International Journal of Quantum Chemistry, 36(S23):31–43, April 1–8, 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Deumens:1987:DAE

[DÖL87] E. Deumens, Y. Öhrn, and L. Lathouwers. A dynamical approach to electron transfer reactions. International Journal of Quantum Chemistry, 32(S21):321–339, March 12, 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Chemistry, Solid-State Theory, and Computational Methods.

Donini:1979:RAG

[Don79] John C. Donini, editor. Recent advances in group theory and their application to spectroscopy: [proceedings], volume 43 of NATO Advanced Study Institutes Series: Series B, Physics. Plenum Press, New York, NY, USA; London, UK, 1979. ISBN 0-306-40172-X. LCCN QC450 .N37 1978.

Donnelly:1981:ADM

[Don81] Robert A. Donnelly. An approximate density matrix energy functional. International Journal of Quantum Chemistry, 20 (S15):409–417, March 8–14, 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory, Collision Phenomena, and Computational Quantum Chemistry.

Donnelly:1982:CRP

[Don82] Robert A. Donnelly. Coordinate-rotated propagator calculation on an N_{2} shape resonance. International Journal of Quantum Chemistry, 22(S16):653–659, March 1–13, 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Chemistry, Theory of Condensed Matter, and Propagator Methods in the Quantum Theory of Matter.

Donnelly:1985:SOC

[Don85] Robert A. Donnelly. Second-order calculation on the doublet pi CO shape resonance. International Journal of Quantum Chemistry, 28(S19):363–368, March 18–23, 1985. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Dose:1984:SIC

[Dos84] Klaus Dose. Self-instructed condensation of amino acids and the origin of biological information. International Journal of Quantum Chemistry, 26(S11):91–101, March 12–15, 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Dovesi:1984:IHF

[Dov84] Roberto Dovesi. Ab initio Hartree–Fock approach to the study of polymers: Application to polyacetylenes. International Journal of Quantum Chemistry, 26(2):197–212, August 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Dovesi:1986:RSI

[Dov86] Roberto Dovesi. On the role of symmetry in the ab initio Hartree–Fock linear-combination-of-atomic-orbitals treatment of periodic systems. International Journal of Quantum Chemistry, 29(6):1755–1774, June 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Dijkman:1987:PCC

[DOW87] Johannes P. Dijkman, Roman Osman, and Harel Weinstein. Practical considerations in calculations of the proton transfer in a model active site of papain. International Journal of Quantum Chemistry, 32(S14):211–219, March 14, 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Dijkman:1989:TSE

[DOW89] Johannes P. Dijkman, Roman Osman, and Harel Weinstein. A theoretical study of the effect of primary and secondary structure elements on the proton transfer in papain. International Journal of Quantum Chemistry, 35(1):241–252, January 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Dmitriev:1981:CPT

[DP81] Yu. Dmitriev and G. Peinel. Coupled perturbation theory within the antisymmetrized product of separated geminals (APSG) framework. International Journal of Quantum Chemistry, 19(5):763–769, May 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Damburg:1982:DW

[DP82] R. Damburg and R. Propin. Double wells. International Journal of Quantum Chemistry, 21(1):191–193, January 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Datta:1987:MCT

[DP87] S. N. Datta and S. Priyadarshy. A model calculation on the transport of excitation energy in a molecular crystal. International Journal of Quantum Chemistry, 32(4):411–426, October 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Daudel:1982:UQG

[DPC82] Raymond Daudel, Raymond A. Poirier, and Imre G. Csizmadia. Uniform quality Gaussian basis sets for molecular calculations. V. Property optimization: a study on H_{2}O. International Journal of Quantum Chemistry, 21(4):699–710, April 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Dovesi:1980:EEH

[DPR80] R. Dovesi, C. Pisani, and C. Roetti. Exact-exchange Hartree– Fock calculations for periodic systems. II. Results for graphite and hexagonal boron nitride. International Journal of Quantum Chemistry, 17(3):517–529, March 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Duke:1987:SAC

[DPS87] C. B. Duke, A. Paton, and W. R. Salaneck. Spectroscopic analysis of the conformations of polyaniline oligomers. International Journal of Quantum Chemistry, 32(S21):153–162, March 12, 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Chemistry, Solid-State Theory, and Computational Methods.

Danilov:1980:EAE

[DPZ80] V. I. Danilov, V. I. Pechenaya, and N. V. Zheltovsky. Electronic absorption and emission spectra of nucleic acids and their components: Some questions of theory and experiment. International Journal of Quantum Chemistry, 17(2):307–320, February 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Duprez-Quesnoit:1980:DIMb

[DQ80] C. Duprez-Quesnoit. Definition of invariant molecular tensors in an origin change. II. Refraction and Faraday effect. International Journal of Quantum Chemistry, 17(2):169–179, February 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Duprez-Quesnoit:1981:DIM

[DQ81] C. Duprez-Quesnoit. Definition of invariant molecular tensors in an origin change. III. Comparison of the methods suitable for the obtainment of invariant tensors for refraction and the Faraday effect. International Journal of Quantum Chemistry, 19(2):191–205, February 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Duprez-Quesnoit:1980:DIMa

[DQT80] C. Duprez-Quesnoit and J. Tillieu. Definition of invariant molecular tensors in an origin change. I. General formalism. International Journal of Quantum Chemistry, 17(2):155–167, February 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Dugay:1980:TVS

[DR80] M. Dugay and J. Roustan. Toward a very simple picture of disorder effects in one-dimensional organic conductors. International Journal of Quantum Chemistry, 18(4):1057–1069, October 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Dmitriev:1984:TCM

[DR84] Yuri Yu. Dmitriev and Björn O. Roos. Theory and computational methods for studies of nonlinear phenomena in laser spectroscopy. I. General formalism. International Journal of Quantum Chemistry, 26(1):35–49, July 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Donnelly:1988:DST

[DR88] Robert A. Donnelly and Jack W. Rogers, Jr. A discrete search technique for global optimization. International Journal of Quantum Chemistry, 34(S22):507–513, March 12–19, 1988. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Syposium on Quantum Chemistry, Solid-State Theory, and Computational Methods.

Diercksen:1983:ECP

[DRS83] Geerd H. F. Diercksen, Björn O. Roos, and Andrzej J. Sadlej. Electron correlation and properties of many-electron systems. International Journal of Quantum Chemistry, 24(S17):265– 288, March 3–12, 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-state Theory, Collision Phenomena, and Computational Quantum Chemistry.

Dinesha:1988:ISS

[DRS88] K. V. Dinesha, S. Rettrup, and C. R. Sarma. An indexing scheme for spin-free configurations of electrons. International Journal of Quantum Chemistry, 34(5):445–455, November 1988. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Datta:1980:SCC

[DS80a] Dilip Kumar Datta and Sankar Sengupta. Self-consistent calculation of 1s2s ^{1}S state wave function of helium. International Journal of Quantum Chemistry, 18(4):995–1003, October 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Davies:1980:PAC

[DS80b] R. H. Davies and L. H. Smith. Partial agonism of cardiac β-adrenoceptor blocking agents and competitive conformerreceptor occupancy. International Journal of Quantum Chemistry, 18(S7):331–345, March 5–8, 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement:

Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Ducla-Soares:1980:BEM

[DS80c] Eduardo Ducla-Soares. Biological effects of magnetic fields and cooperative phenomena. International Journal of Quantum Chemistry, 17(6):1043–1046, June 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Das:1981:ECW

[DS81a] Sudipto Das and Gaurishanker Singhal. Effect of change of water structure on the phase transition of liposomes of dipalmitoyl phosphatidylcholine. International Journal of Quantum Chemistry, 20(2):495–504, August 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

De:1981:CSS

[DS81b] B. R. De and A. B. Sannigrahi. Comparative study of the structure and properties of ClF_{2} and Cl_{3} radicals by CNDO/2 and INDO methods. International Journal of Quantum Chemistry, 19(3):485–488, March 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Dutta:1981:SCC

[DS81c] Dilip K. Dutta and Sankar Sengupta. Self-consistent calculation of 1s2s^{3} S state of helium and heliumlike ions. International Journal of Quantum Chemistry, 19(2):283–291, February 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

De:1982:EST

[DS82] B. R. De and A. B. Sannigrahi. Electronic structure of triatomic interhalogens. International Journal of Quantum Chemistry, 22(2):435–440, August 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Deak:1987:PSR

[DS87] Peter Deák and Lawrence C. Snyder. A possible solution to the rotational invariance and silicon parameter problem in the MNDO approximation. International Journal of Quantum Chemistry, 32(S21):547–554, March 12, 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Supplement: Proceedings of the International Symposium on Quantum Chemistry, Solid-State Theory, and Computational Methods.

Dmitriev:1988:APT

[DS88] Yuri Yu. Dmitriev and Olga V. Solnyshkina. Adiabatic perturbation theory for the degenerate case. III. Adiabatic formalism by Gell-Mann and low with an arbitrary switching function. International Journal of Quantum Chemistry, 33(6):543–553, June 1988. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Danilov:1989:HEB

[DS89a] Victor I. Danilov and Anna V. Shestopalova. Hydrophobic effect in biological associates: A Monte Carlo simulation of caffeine molecules stacking. International Journal of Quantum Chemistry, 35(1):103–112, January 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

DelBene:1989:CMD

[DS89b] Janet E. Del Bene and Isaiah Shavitt. Comparison of methods for determining the correlation contribution to hydrogen bond energies. International Journal of Quantum Chemistry, 36(S23):445–452, April 1–8, 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Defranceschi:1984:NSH

[DSB84] M. Defranceschi, M. Suard, and G. Berthier. Numerical solution of Hartree–Fock equations for a polyatomic molecule: Linear H_{3} in momentum space. International Journal of Quantum Chemistry, 25(5):863–867, May 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

dosSantos:1986:UTE

[dSdMB86] M. C. dos Santos, C. P. de Melo, and H. S. Brandi. Unified treatment of the electronic structure of organic conjugated polymers. International Journal of Quantum Chemistry, 30 (1):109–118, July 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Ducla-Soares:1980:AMC

[DSG80] Eduardo Ducla-Soares and Maria Margarida Godinho. An application of microphysical concepts to the sensory response: Cooperative effects. International Journal of Quantum Chemistry, 18(S7):375–383, March 5–8, 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Dayashankar:1986:EIC

[DSG86] Dayashankar, Suk T. Suh, and A. E. S. Green. Electron impact cross sections and spatial aspects of electron energy degradation in water vapor. International Journal of Quantum Chemistry, 30(S20):547–554, March 10–15, 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Syposium on Atomic, Molecular, and Solid-State Theory, Scattering Problems, Many Body Phenomena, and Computational Quantum Chemistry.

Detrich:1980:TSA

[DSW80] John Detrich, Edward S. Sachs, and Arnold C. Wahl. Theoretical studies of atmospheric triatomic molecules. International Journal of Quantum Chemistry, 18(S14):262, March 10–20, 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory, Collision Phenomena, Quantum Statistics, and Computational Methods.

DalCappello:1981:CPE

[DT81] C. Dal Cappello and C. Tavard. Compton profiles for ejected electrons: Determination of momentum densities. International Journal of Quantum Chemistry, 20(5):1121–1125, November 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Davis:1983:NMC

[DT83] Larry P. Davis and Almon G. Turner. 1-nitropropene: A model compound for the theoretical study of TNT decomposition. International Journal of Quantum Chemistry, 24 (S17):193, March 3–12, 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-state Theory, Collision Phenomena, and Computational Quantum Chemistry.

Duch:1985:EMC

[Duc85] W. Duch. Efficient method for computation of representation matrices of the unitary group generators. International Journal of Quantum Chemistry, 27(1):59–70, January 1985. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Duch:1986:DSE

[Duc86] Włodzisław Duch. From determinants to spin eigenfunctions— a simple algorithm. International Journal of Quantum Chemistry, 30(6):799–807, December 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Duke:1982:SIM

[Duk82] Brian J. Duke. Simulated ab initio molecular orbital methods for polymers. International Journal of Quantum Chemistry, 22(S16):391–407, March 1–13, 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Chemistry, Theory of Condensed Matter, and Propagator Methods in the Quantum Theory of Matter.

Dunne:1980:BRB

[Dun80] Lawrence J. Dunne. Book review: High pressure chemistry. Edited by H. Kelm. NATO Advanced Study Institute Series, D. Reidel Publishing Company, Dordrccht. Holland, 1979. International Journal of Quantum Chemistry, 18(3):911, September 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). See [Kel78].

Dunne:1987:MPQ

[Dun87] Lawrence J. Dunne. The “measurement problem” in quantum chemistry and the interpretation of phenomenological damping terms in the method of classical trajectories for nonadiabatic molecular collisions. International Journal of Quantum Chemistry, 32(1):31–33, July 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Dunlap:1988:DNC

[Dun88] Brett I. Dunlap. Do the new cluster sources also produce isomers? International Journal of Quantum Chemistry, 34(S22): 257–264, March 12–19, 1988. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Syposium on Quantum Chemistry, Solid-State Theory, and Computational Methods.

Durand:1986:CVC

[Dur86] Ph. Durand. Core–valence correlations and relativistic effects in heavy atoms for molecular applications. International Journal of Quantum Chemistry, 29(4):915–935, April 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Devreese:1980:MEM

[DVV80] J. T. Devreese, P. E. Van Camp, and V. E. Van Doren. Moment expansion method to calculate vibrational frequencies of molecules and solids. International Journal of Quantum Chemistry, 18(1):317–321, July 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Dressler:1986:CTE

[DWQ86] K. Dressler, L. Wolniewicz, and P. Quadrelli. Comparison of theory and experiment for excited singlet states of the H_{2} molecule. International Journal of Quantum Chemistry, 29 (2):185–189, February 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Ding:1989:LAA

[DYD89] Shiliang Ding, Xizhang Yi, and Conghao Deng. Lie algebraic approach to the scattering system H_{e} + H_{2}. International Journal of Quantum Chemistry, 36(3):379–389, September 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

El-Basil:1981:ARE

[EB81] Sherif El-Basil. Additivity of resonance energy in benzenoid hydrocarbons. International Journal of Quantum Chemistry, 19(4):593–609, April 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

El-Basil:1982:NGTa

[EB82a] Sherif El-Basil. Novel graph-theoretical approach to estimating the relative importance of individual Kekulé valence structures. I. Simple catacondensed systems. International Journal of Quantum Chemistry, 21(5):771–778, May 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

El-Basil:1982:NGTb

[EB82b] Sherif El-Basil. Novel graph-theoretical approach to estimating the relative importance of individual Kekulé valence structures. II. Benzenoid systems having four to seven fused benzene rings: “Pseudodegenerate” states. International Journal of Quantum Chemistry, 21(5):779–792, May 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

El-Basil:1982:NGTc

[EB82c] Sherif El-Basil. Novel graph-theoretical approach to estimating the relative importance of individual Kekulé valence structures. III. Nonalternate and nonbenzenoid hydrocarbons. International Journal of Quantum Chemistry, 21(5):793–797, May 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Edgecombe:1986:BCP

[EB86] K. E. Edgecombe and R. J. Boyd. Bond critical points in the electronic structures of binary hydrides. International Journal of Quantum Chemistry, 29(4):959–973, April 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Esquivel:1987:AED

[EB87] Rodolfo O. Esquivel and Annik Vivier Bunge. Accurate electron density and one-electron properties for the beryllium atom. International Journal of Quantum Chemistry, 32(3): 295–312, September 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

El-Basil:1987:RMP

[EBEE87] S. El-Basil and M. El-Enany. A revisit to molecular path codes: Ordering, comparability, and similarity of benzenoid, nonbenzenoid, and monoradical hydrocarbons. International Journal of Quantum Chemistry, 32(2):165–179, August 1987.

CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

El-Basil:1983:OKS

[EBO83] Sherif El-Basil and A. N. Osman. Ordering of Kekulé structures using nonadjacent numbers. International Journal of Quantum Chemistry, 24(6):571–592, December 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Engelmann:1983:TMS

[EC83] A. R. Engelmann and R. H. Contreras. Transmission mechanisms of spin–spin coupling constants within the CHF approximation: Their study using inner projections of the polarization propagator. International Journal of Quantum Chemistry, 23(3):1033–1045, March 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

ElKhatib:1989:EST

[ECM89] N. El Khatib, H. Chermette, and M. Maitrot. Electronic structure, theoretical electronic properties of tetraphenyldithiapyranylidene DIPS φ_{4} molecule, and evaluation of the charge transfer occurring with doping agents like TCNQ. International Journal of Quantum Chemistry, 35(2):339–358, February 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Edwards:1988:CSS

[Edw88] W. Daniel Edwards. Complex solutions to simple molecular orbital problems. International Journal of Quantum Chemistry, 34(S22):549–556, March 12–19, 1988. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Syposium on Quantum Chemistry, Solid-State Theory, and Computational Methods.

Ellis:1984:SCD

[EG84] D. E. Ellis and G. L. Goodman. Self-consistent Dirac–Slater calculations for molecules and embedded clusters. International Journal of Quantum Chemistry, 25(1):185–200, January 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Ellis:1986:IDT

[EGP86] D. E. Ellis, Diana Guenzburger, and M. R. Press. Impurities and defects in transition metals and their oxides. International Journal of Quantum Chemistry, 30(S20):353–366, March 10–15, 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Syposium on Atomic, Molecular, and Solid-State Theory, Scattering Problems, Many Body Phenomena, and Computational Quantum Chemistry.

Etters:1983:GSS

[EHK83] R. D. Etters, A. A. Helmy, and K. Kobashi. Ground-state structural and magnetic properties of solid α and β-O_{2}. International Journal of Quantum Chemistry, 24(S17):491–499, March 3–12, 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-state Theory, Collision Phenomena, and Computational Quantum Chemistry.

El-Issa:1981:GFM

[EI81] B. D. El-Issa. Green’s functions in multiple scattering theory. International Journal of Quantum Chemistry, 20(S15): 419–428, March 8–14, 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory, Collision Phenomena, and Computational Quantum Chemistry.

El-Issa:1988:CSB

[EIKG^{+}88] B. D. El-Issa, A. Katrib, R. Ghodsian, B. A. Salsa’, and S. H. Addassi. A comparative study of the bonding in different halides of iridium. International Journal of Quantum Chemistry, 33(3):195–216, March 1988. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

El-Issa:1987:MOS

[EIMS87] B. D. El-Issa, M. A. Makhyoun, and B. A. Salsa’. A molecular orbital study of cis and trans diaminodichloroplatinum(II). International Journal of Quantum Chemistry, 31(2):295–307, February 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

El-Issa:1988:ESE

[EIMS88] B. D. El-Issa, M. A. Makhyoun, and B. A. Salsa’. Excited state electron density differentials: a new technique for studying HOMO /LUMO excitations. International Journal of Quantum Chemistry, 33(2):115–125, February 1988. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Eisenstein:1988:SDD

[Eis88] M. Eisenstein. SCF Deformation densities and electrostatic potentials of purines and pyrimidines. International Journal of Quantum Chemistry, 33(2):127–158, February 1988. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Elander:1986:MRT

[EKBE86] Nils Elander, Peter Krylstedt, Erkki Brändas, and Erik Engdahl. The R matrix: Its relation to Titchmarsh–Weyl theory and its complex rotated analogue. International Journal of Quantum Chemistry, 30(S20):73–80, March 10–15, 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Syposium on Atomic, Molecular, and Solid-State Theory, Scattering Problems, Many Body Phenomena, and Computational Quantum Chemistry.

Elander:1987:RNA

[Ela87] Nils Elander. Resonances in nuclear, atomic, and molecular physics—an introduction with some examples. International Journal of Quantum Chemistry, 31(5):707–731, May 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Ellis:1984:LDS

[Ell84] D. E. Ellis. Local density studies on transition-metal clusters, surfaces, and the oxide defect state. International Journal of Quantum Chemistry, 26(S18):183–193, March 1–15, 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Atomic, Molecular and Solid-State Theory, and Computational Quantum Chemistry.

Emch:1980:IDA

[Emc80] Gérard G. Emch. Irreversibility, decay, and asymptotic dynamics. International Journal of Quantum Chemistry, 17(1):

107–120, January 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Emin:1983:SPF

[Emi83] David Emin. Small-polaron formation and motion. International Journal of Quantum Chemistry, 24(S17):647, March 3–12, 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-state Theory, Collision Phenomena, and Computational Quantum Chemistry.

Engelmann:1987:SCS

[EN87] Alejandro R. Engelmann and Mario A. Natiello. Some considerations on the Sturmian approach to Green’s functions. International Journal of Quantum Chemistry, 32(4):457–466, October 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Echenique:1989:ISI

[ENA89] P. M. Echenique, I. Nagy, and A. Arnau. Interaction of slow ions with matter. International Journal of Quantum Chemistry, 36(S23):521–543, April 1–8, 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

England:1983:OFTa

[Eng83a] Walter B. England. Ordinary field-theoretic methods for selfconsistent wave functions which describe bond formation and dissociation. II. Commutative coupling case. International Journal of Quantum Chemistry, 23(3):905–914, March 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

England:1983:OFTb

[Eng83b] Walter B. England. Ordinary field-theoretic methods for selfconsistent wave functions which describe bond formation and dissociation. III. The commutative coupling approximation. International Journal of Quantum Chemistry, 24(S17):357– 365, March 3–12, 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-state Theory, Collision Phenomena, and Computational Quantum Chemistry.

Engelmann:1987:ABS

[ENH^{+}87] Alejandro R. Engelmann, Mario A. Natiello, Mikael Höghede, Erik Engdahl, and Erkki Brändas. Association of bound states of the Coulomb potential with resonances of the Coulomb potential perturbed by a barrier. International Journal of Quantum Chemistry, 31(5):841–845, May 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Enz:1980:WIM

[Enz80] Charles P. Enz. Weak itinerant magnetism, new developments: TiBe_{2}. International Journal of Quantum Chemistry, 18(S14):723–727, March 10–20, 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory, Collision Phenomena, Quantum Statistics, and Computational Methods.

Edwards:1984:SAG

[EÖWZ84] W. Daniel Edwards, N. Yngve Öhrn, Brian L. Weiner, and Michael C. Zerner. Semiempirical antisymmetrized geminal power calculation for molecular spectra. International Journal of Quantum Chemistry, 26(S18):507–517, March 1–15, 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Atomic, Molecular and Solid-State Theory, and Computational Quantum Chemistry.

Epstein:1987:SP

[Eps87] Saul T. Epstein. Stationary properties of δE/δR. International Journal of Quantum Chemistry, 31(3):425–427, March 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Elander:1982:DWT

[ERB82] Nils Elander, Magnus Rittby, and Erkki Brändas. Dilated Weyl theory applied to a set of N-coupled singular secondorder differential equations. International Journal of Quantum Chemistry, 22(3):445–457, September 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Eriksson:1989:GSQ

[Eri89] K.-E. Eriksson. Guessing the state of a quantum system from a limited set of observations. International Journal of Quantum Chemistry, 35(6):745–749, June 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Easa:1980:DMP

[ES80] S. I. Easa and G. C. Shukla. Dynamic multipole polarizabilities of atomic systems. International Journal of Quantum Chemistry, 17(4):673–678, April 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Elgersma:1981:QTS

[ES81] H. Elgersma and G. C. Schatz. A quasiclassical trajectory study of mode specific reaction rate enhancements in H + H_{2}O (v_{1} v_{2} v_{3}) → OH + H_{2}. International Journal of Quantum Chemistry, 20(S15):611–619, March 8–14, 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory, Collision Phenomena, and Computational Quantum Chemistry.

El-Sayed:1988:MMS

[ES88] M. A. El-Sayed. On the molecular mechanisms of solar energy storage during the photocycle of the other photosynthetic system in nature, bacteriorhodopsin. International Journal of Quantum Chemistry, 34(S22):367–375, March 12–19, 1988. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Syposium on Quantum Chemistry, Solid-State Theory, and Computational Methods.

Etters:1985:VRF

[ESCC85] R. D. Etters, B. Silvi, V. Chandrasekharan, and M. Chergui. Vibrational and rotational frequency shifts of dilute H_{2}, D_{2}, and HD impurities in solid Ar, Kr, and Xe under pressure. International Journal of Quantum Chemistry, 28(S19):675–686, March 18–23, 1985. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Elander:1983:AAW

[ESKG83] Nils Elander, Erland Sangfelt, Henry Kurtz, and Osvaldo Goscinski. Application of AGP wave functions to the ground states of Li_{2} and CH^{+}. International Journal of Quantum Chemistry, 23(3):1047–1056, March 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Essen:1982:PTE

[Ess82] Hanno Essén. Periodic table of the elements and the Thomas– Fermi atom. International Journal of Quantum Chemistry, 21 (4):717–726, April 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Esser:1984:DMM

[Ess84] M. Esser. Direct MRCI method for the calculation of relativistic many-electron wavefunctions. I. General formalism. International Journal of Quantum Chemistry, 26(3):313–338, September 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

England:1986:LSF

[ESS86a] Walter B. England, Thomas E. Sorensen, and David M. Silver. Lewis structures and Feynman diagrams: The treatment of geminal correlation in Fock space. International Journal of Quantum Chemistry, 30(S20):81–94, March 10–15, 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Syposium on Atomic, Molecular, and Solid-State Theory, Scattering Problems, Many Body Phenomena, and Computational Quantum Chemistry.

Essean:1986:CIP

[Ess86b] Hanno Esséan. Correlation with independent particle orbitals. International Journal of Quantum Chemistry, 30(1):89–95, July 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Engelking:1987:SEE

[EV87] P. C. Engelking and V. Vaida. Symmetry effects in the electronic spectrum of ammonia at low temperature. International Journal of Quantum Chemistry, 31(1):73–80, January 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Easa:1982:LRI

[EY82] S. I. Easa and F. N. Yousif. Long-range interaction coefficients between H, He^{+}, He, and Li^{+}. International Journal of Quantum Chemistry, 21(6):1117–1121, June 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Edwards:1983:ESM

[EZ83] W. Daniel Edwards and Michael C. Zerner. Electronic structure of model chlorophyll systems. International Journal of Quantum Chemistry, 23(4):1407–1432, April 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Friedman:1986:ESP

[FA86] Paul Friedman and Leland C. Allen. The electronic structure of 4H-pyran-4-one and its sulfur analogues. International Journal of Quantum Chemistry, 29(5):1503–1512, May 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Fernandez:1984:CBS

[FAC84] Francisco M. Fernández, Gustavo A. Arteca, and Eduardo A. Castro. Calculation of bound-state energies from a variational functional method. International Journal of Quantum Chemistry, 25(6):1023–1033, June 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Faulkner:1985:ESD

[Fau85] J. S. Faulkner. Electronic states in disordered solids. III. Applications to materials science. International Journal of Quantum Chemistry, 28(S19):579–593, March 18–23, 1985. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Faxen:1984:IR

[Fax84] Magnus Faxen. Introductory remarks. International Journal of Quantum Chemistry, 26(S18):21–24, March 1–15, 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Atomic, Molecular and Solid-State Theory, and Computational Quantum Chemistry.

Faxen:1986:WA

[Fax86] Magnus Faxen. Welcome address. International Journal of Quantum Chemistry, 30(S13):1, March 6–8, 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Faxen:1987:RSS

[Fax87] Magnus Faxen. Remarks at the Sanibel Symposium, Marineland, Florida, March 18, 1987. International Journal of Quantum Chemistry, 32(S21):13–14, March 12, 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Chemistry, Solid-State Theory, and Computational Methods.

Fazzio:1987:PST

[Faz87] A. Fazzio. Progress in the study of transition metal impurities in III–V and II–VI materials. International Journal of Quantum Chemistry, 32(S21):65–72, March 12, 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Chemistry, Solid-State Theory, and Computational Methods.

Froelich:1983:SAP

[FB83] Piotr Froelich and Erkki Brändas. Some analytical properties of the dilated SCF Equations. International Journal of Quantum Chemistry, 23(1):91–98, January 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Fry:1980:EPS

[FBM80] J. L. Fry, N. E. Brener, and J. D. Mcnutt. Electron and positron states on the (001) surface of paramagnetic chromium. International Journal of Quantum Chemistry, 18 (S14):645–649, March 10–20, 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory, Collision Phenomena, Quantum Statistics, and Computational Methods.

Facell:1980:IHO

[FC80a] J. C. Facell and R. H. Contreras. Inclusion of hydrogen p orbitals in the semiempirical calculation of NMR parameters. I. Influence on the FPT INDO spin–spin coupling constants. International Journal of Quantum Chemistry, 18(5):1175–1185, November 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Fernandez:1980:FAI

[FC80b] Francisco M. Fernández and Eduardo A. Castro. Further analysis of the iterative hypervirial-scaling method. International Journal of Quantum Chemistry, 17(4):609–618, April 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Facelli:1981:IHO

[FC81a] J. C. Facelli and R. H. Contreras. Inclusion of hydrogen p orbitals in the semiempirical calculation of NMR parameters. III: INDO CHF calculations of orbital and dipolar contributions to spin–spin coupling constants involving protons. International Journal of Quantum Chemistry, 20(4):909–919, October 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Fernandez:1981:HAEa

[FC81b] Francisco M. Fernández and Eduardo A. Castro. Hypervirial analysis of enclosed quantum mechanical systems. I. Dirichlet boundary conditions. International Journal of Quantum Chemistry, 19(4):521–532, April 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Fernandez:1981:HAEb

[FC81c] Francisco M. Fernández and Eduardo A. Castro. Hypervirial analysis of enclosed quantum mechanical systems. II. von Neumann boundary conditions and periodic potentials. International Journal of Quantum Chemistry, 19(4):533–543, April 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Fernandez:1981:HAEc

[FC81d] Francisco M. Fernández and Eduardo A. Castro. Hypervirial analysis of enclosed quantum mechanical systems. III. Unsymmetrical boundary conditions. International Journal of Quantum Chemistry, 20(3):623–632, September 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Fernandez:1982:SCA

[FC82a] Francisco M. Fernandez and Eduardo A. Castro. Some comments about tensor virial theorem and linear transformations. International Journal of Quantum Chemistry, 21(6): 1097–1103, June 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Fernandez:1982:TDH

[FC82b] Francisco M. Fernández and Eduardo A. Castro. Trivial offdiagonal Hellmann–Feynman formulas. International Journal of Quantum Chemistry, 21(3):671–672, March 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Fernandez:1982:VTB

[FC82c] Francisco M. Fernández and Eduardo A. Castro. Virial theorem and boundary conditions for approximate wave functions. International Journal of Quantum Chemistry, 21(4):741–751, April 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Fernndez:1982:HEE

[FC82d] Francisco M. Ferńndez and Eduardo A. Castro. Hyperviral extension for energies of parameter-dependent systems. International Journal of Quantum Chemistry, 21(4):673–676, April 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Fernndez:1982:HAE

[FC82e] Francisco M. Ferńndez and Eduardo A. Castro. Hypervirial analysis of enclosed quantum mechanical systems. IV. Further comments. International Journal of Quantum Chemistry, 21 (5):875–884, May 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Fernandez:1983:HTS

[FC83] Francisco M. Fernández and Eduardo A. Castro. Hypervirial theorems for some simple quantum chemistry systems. International Journal of Quantum Chemistry, 23(3):915–919, March 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Fernandez:1984:PTW

[FC84] Francisco M. Fernández and Eduardo A. Castro. Perturbation theory without wave function for multidimensional systems. International Journal of Quantum Chemistry, 26(4):497–505, October 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Fernandez:1985:PCH

[FC85] Francisco M. Fernández and Eduardo A. Castro. Perturbation calculation on the hydrogen atom in electric and magnetic fields. International Journal of Quantum Chemistry, 28(5): 603–607, November 1985. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Fernandez:1987:LDP

[FC87] Francisco M. Fernández and Eduardo A. Castro. On the linearly driven parametric oscillator. International Journal of Quantum Chemistry, 32(5):591–594, November 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Frishberg:1980:HFD

[FCB80] C. Frishberg, L. Cohen, and P. Blumenau. Hartree–Fock density matrix equation: Open-shell systems. International Journal of Quantum Chemistry, 18(S14):161–165, March 10– 20, 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory, Collision Phenomena, Quantum Statistics, and Computational Methods.

Fernandez:1989:REE

[FCO89] Francisco M. Fernández, Eduardo A. Castro, and John F. Ogilvie. On the Ricatti equation for eigenvalue problems. International Journal of Quantum Chemistry, 36(1):61–65, July 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Furois-Corbin:1984:ENA

[FCPL84] Sylvie Furois-Corbin, Bernard Pullman, and Richard Lavery. Effect of nucleic acid flexibility upon ASIF (accessible surface integrated field). Possible significance for biochemical reactivity exemplified by aflatoxin B_{1} binding. International Journal of Quantum Chemistry, 26(S11):273–286, March 12–15, 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Ferreira:1986:AVC

[FD86] Luiz G. Ferreira and Manoel L. De Siqueira. Application of the variational cellular method to semiconductors: The ZnS case. International Journal of Quantum Chemistry, 30(S20):313– 323, March 10–15, 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Syposium on Atomic, Molecular, and Solid-State Theory, Scattering Problems, Many Body Phenomena, and Computational Quantum Chemistry.

Feng:1989:ECF

[FD89a] Xuejun Feng and Jens Peder Dahl. Electron correlation forms in the LiH molecule — Fermi correlation. International Journal of Quantum Chemistry, 36(6):689–704, December 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Ferris:1989:CBP

[FD89b] Kim F. Ferris and C. B. Duke. Chemical bonding in phosphonitrilic systems — comparison of the electronic structures of (F_{2}PN)_{3}, (F_{2}PN)_{4}, and OP(F_{2})NP(F_{2})NPF_{3}. International Journal of Quantum Chemistry, 36(S23):397–407, April 1–8, 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Frisch:1989:ESE

[FD89c] Michael J. Frisch and Janet E. Del Bene. The electronic structure and electrostatics of nitrous oxide. International Journal of Quantum Chemistry, 36(S23):363–369, April 1–8, 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Fripiat:1983:ECL

[FDA83] J. G. Fripiat, J. Dehalle, and J. M. André. Electron correlation in linear hydrogen chains. International Journal of Quantum Chemistry, 23(4):1179–1189, April 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Fripiat:1989:PAS

[FDAC89] J. G. Fripiat, J. Delhalle, J. M. André, and J. L. Calais. Potentials of an alternative scheme for ab initio polymer band structure calculations. Illustration on the chain of hydrogen atoms. International Journal of Quantum Chemistry, 36(3): 341–351, September 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Fratev:1984:CPC

[FEKC84] F. Fratev, V. Enchev, P. Karadakov, and O. Castaño. A classification of polyenes into 4L + 2- and 4L-classes on the basis of Coulson’s bond orders and information theory and its application to the interpretation of electrocyclic reactions. International Journal of Quantum Chemistry, 26(6):993–1015, December 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

FerreiraDaSilva:1988:ISD

[Fer88] A. Ferreira Da Silva. On the impurity states in doped semiconductors via the alternant molecular orbital method. International Journal of Quantum Chemistry, 34(S22):15–20, March 12–19, 1988. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Syposium on Quantum Chemistry, Solid-State Theory, and Computational Methods.

Fanning:1980:CSI

[FF80] M. O. Fanning and N. J. Fitzpatrick. CNDO /INDO formalism for f block elements. International Journal of Quantum Chemistry, 18(6):1339–1359, December 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Findley:1982:GGC

[FF82] A. M. Findley and G. L. Findley. The generalized genetic code. A note on order-isomorphism/order-equivalence relations. International Journal of Quantum Chemistry, 22(S9):

59–63, March 4–6, 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Findley:1984:GCT

[FF84] A. M. Findley and G. L. Findley. Genetic coding theory: Multiple mitochondrial genetic codes. International Journal of Quantum Chemistry, 26(S11):109–116, March 12–15, 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

FerreiraDaSilva:1985:EHM

[FF85a] A. Ferreira Da Silva and M. Fabbri. Erratum: The hydrogen molecule for a two-dimensional system. International Journal of Quantum Chemistry, 28(5):661, November 1985. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). See [FF85b].

FerreiraDaSilva:1985:HMT

[FF85b] A. Ferreira Da Silva and M. Fabbri. The hydrogen molecule for a two-dimensional system. International Journal of Quantum Chemistry, 27(4):375–379, April 1985. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). See erratum [FF85a].

Froman:1989:DPE

[FF89] Nanny Fröman and Per Olof Fröman. Determination of the potential from experimental data on energies and widths of quasi-stationary levels. International Journal of Quantum Chemistry, 35(6):751–760, June 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Ferris:1988:ESP

[FFF88] Kim F. Ferris, Paul Friedman, and Donald M. Friedrich. Electronic structure of the phosphonitrilic trimer — role of d orbitals in chemical bonding. International Journal of Quantum Chemistry, 34(S22):207–218, March 12–19, 1988. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Syposium on Quantum Chemistry, Solid-State Theory, and Computational Methods.

Ferreira:1982:EAE

[FG82] Ricardo Ferreira and Marcelo A. F. Gomes. Electronic aspects of enzymatic catalysis. International Journal of Quantum Chemistry, 22(3):537–545, September 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Facelli:1984:TSM

[FGC84] J. C. Facelli, C. G. Giribet, and R. H. Contreras. A theoretical study of medium effects on the transmission mechanisms of the Fermi contact term of spin–spin coupling constants in the acetamide molecule. International Journal of Quantum Chemistry, 25(3):515–525, March 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Fonseca:1983:STD

[FGG83] Teresa Fonseca, J. A. N. F. Gomes, and Paolo Grigolini. Shorttime description of chemical reaction by Stochastic methods. International Journal of Quantum Chemistry, 23(2):473–476, February 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Facelli:1987:AIB

[FGM87] Julio C. Facelli, David M. Grant, and Josef Michl. Analysis of the IGLO bond contributions to the ^{13}C shielding tensors in the local bond frame. International Journal of Quantum Chemistry, 31(1):45–55, January 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Fusco-Girard:1980:QFT

[FGMM80] M. Fusco-Girard, F. Mancini, and M. Marinaro. Quantumfield theory and electrodynamics of superconducting systems. International Journal of Quantum Chemistry, 17(1):75–82, January 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Fernandez:1985:CIR

[FGS85] M. Fernández, P. Góamez, and P. Serna. Calculation of internal rotation spectra of three molecules with increasing asymmetry factors. International Journal of Quantum Chemistry, 27(3):235–244, March 1985. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Fukutome:1981:CTB

[FH81] Hideo Fukutome and Katsufumi Hashimoto. Characterization of the two-body correlation structures in a molecular many-electron wave function. International Journal of Quantum Chemistry, 20(S15):33–59, March 8–14, 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory, Collision Phenomena, and Computational Quantum Chemistry.

Fischer-Hjalmars:1983:I

[FH83a] Inga Fischer-Hjalmars. Introduction. International Journal of Quantum Chemistry, 23(4):ix–x, April 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Fukutome:1983:CBC

[FH83b] Hideo Fukutome and Katsufumi Hashimoto. Characterization of bond and current correlation structures in a molecular many-electron wave function. International Journal of Quantum Chemistry, 23(1):27–45, January 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Fox:1987:PAT

[FHH^{+}87] Sidney W. Fox, Franz Hefti, Jukka Hartikka, Emmanuel Junard, Alexander T. Przybylski, and Graham Vaughan. Pharmacological activities in thermal proteins: Relationships in molecular evolution. International Journal of Quantum Chemistry, 32(S14):347–349, March 14, 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Fischer-Hjalmars:1980:MBE

[FHHE80] I. Fischer-Hjalmars and A. Henriksson-Enflo. Metal bridge effect: V. Quantum chemical studies of some metal chelates. International Journal of Quantum Chemistry, 18(2):409–419, August 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Fischer-Hjalmars:1985:MBI

[FHHHE85] I. Fischer-Hjalmars, H. Holmgren, and A. Henriksson-Enflo. Metals in biology: Iron complexes modeling electron transfer.

International Journal of Quantum Chemistry, 28(S12):57–67, March 14–16, 1985. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Fink:1985:SIS

[Fin85] William H. Fink. Symmetry implications for the selection of cluster sizes for semiconductor surfaces. International Journal of Quantum Chemistry, 28(S19):595–602, March 18–23, 1985. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Fink:1987:BSR

[Fin87] William H. Fink. Basis sets and restricted Hartree–Fock instabilities in thiocarbonyls. International Journal of Quantum Chemistry, 32(S21):669–676, March 12, 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Chemistry, Solid-State Theory, and Computational Methods.

Finkel:1989:FDG

[Fin89] David L. Finkel. Fractal displays of genomic DNA. I. Eco RI fractal lattice of buffalo rat. International Journal of Quantum Chemistry, 36(5):575–586, November 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Fisher:1982:FCN

[Fis82] Michael E. Fisher. Fluid criticality near walls and between plates. International Journal of Quantum Chemistry, 22(S16): 237–240, March 1–13, 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Chemistry, Theory of Condensed Matter, and Propagator Methods in the Quantum Theory of Matter.

Fischer:1984:VCI

[Fis84] Gad Fischer. Vibronic coupling: the interaction between the electronic and nuclear motions. Academic Press, London, UK, 1984. ISBN 0-12-257240-8. viii + 222 pp. LCCN QD461 F45

Fujikawa:1980:PCS

[FK80] Takashi Fujikawa and Haruo Kuroda. Photoionization cross section calculated by dynamical theory. International Journal of Quantum Chemistry, 18(1):251–256, July 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Flynn:1983:RRE

[FK83] C. P. Flynn and A. B. Kunz. Recent results for excitonic processes in the spectra of metals and alloys. International Journal of Quantum Chemistry, 24(S17):573–582, March 3– 12, 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-state Theory, Collision Phenomena, and Computational Quantum Chemistry.

Ferraz:1983:VCM

[FL83] Armando C. Ferraz and José R. Leite. Variational cellular model of the energy bands of silicon. International Journal of Quantum Chemistry, 24(S17):483–490, March 3–12, 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-state Theory, Collision Phenomena, and Computational Quantum Chemistry.

Flannery:1980:CTT

[Fla80] M. R. Flannery. Charge transfer in three-body ion-ion recombination at low gas densities. International Journal of Quantum Chemistry, 18(S14):477–482, March 10–20, 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory, Collision Phenomena, Quantum Statistics, and Computational Methods.

Flannery:1981:IIR

[Fla81] M. R. Flannery. Ion-ion recombination in dilute and dense plasmas. International Journal of Quantum Chemistry, 20 (S15):715–727, March 8–14, 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory, Collision Phenomena, and Computational Quantum Chemistry.

Fletcher:1981:HTS

[Fle81] G. Fletcher. High-temperature superconductivity in CdS: A theoretical investigation. International Journal of Quantum Chemistry, 20(S15):539–544, March 8–14, 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory, Collision Phenomena, and Computational Quantum Chemistry.

Fleming:1986:CAU

[Fle86] Graham R. Fleming. Chemical applications of ultrafast spectroscopy, volume 13 of The International series of monographs on chemistry; Oxford science publications. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 1986. ISBN 0-19-503644-1. ix + 262 pp. LCCN QD96.L3 F57 1986.

URL http://www.loc.gov/catdir/enhancements/fy0638/
85004949-d.htm.

Fliszar:1984:CDC

[Fli84] Sándor Fliszár. Charge distributions and chemical effects. XXXIV. Concepts involved in an approximate, but accurate, description of bond energies. International Journal of Quantum Chemistry, 26(5):743–752, November 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Fliszar:1986:CDC

[Fli86] S. Fliszár. Charge distributions and chemical effects. XLII. Atomic charges and their role in energy calculations. International Journal of Quantum Chemistry, 29(3):305–310, March 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Flurry:1984:GSG

[Flu84] R. L. Flurry, Jr. Groups of simple graphs. International Journal of Quantum Chemistry, 25(2):309–320, February 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Flurry:1985:ASS

[Flu85] R. L. Flurry, Jr. Approximate symmetries in the singly excited states of helium. International Journal of Quantum Chemistry, 27(2):195–202, February 1985. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Findley:1980:GGC

[FM80] G. L. Findley and S. P. Mcglynn. The generalized genetic code. A modification of code universality. International Journal of Quantum Chemistry, 18(S7):277–281, March 5–8, 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Findley:1981:GE

[FM81] G. L. Findley and S. P. McGlynn. Geometry and evolution. International Journal of Quantum Chemistry, 20(S8):455–461, March 5–7, 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Frishberg:1983:DME

[FM83] C. Frishberg and L. Massa. Density matrix equation for crystals. International Journal of Quantum Chemistry, 23(1):85– 89, January 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Fulscher:1986:ANE

[FM86] M. P. Fülscher and E. L. Mehler. Analysis of nonadditivity effects and estimation of many-body effects in linear water chains. International Journal of Quantum Chemistry, 29 (4):627–638, April 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Fabian:1980:AIE

[FMF80] J. Fabian, A. Mehlhorn, and F. Fratev. Approaches for interpreting π electronic states and π electronic transitions. International Journal of Quantum Chemistry, 17(2):235–253, February 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Friedel:1983:SCU

[FN83] J. Friedel and C. Noguera. Some considerations on U_{eff} values entering the Hubbard Hamiltonian. International Journal of Quantum Chemistry, 23(4):1209–1221, April 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Feng:1984:MAP

[FN84] W. L. Feng and O. Novaro. Multibody analysis of potential energy surfaces for first- and second-row tetramers. II. The cases of O_{4} and S_{4}. International Journal of Quantum Chemistry, 26(4):521–533, October 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Ferraro:1986:UIP

[FNC86] M. B. Ferraro, M. A. Natielo, and R. H. Contreras. The use of inner projections of the polarization propagator to study different contributions to the magnetic shielding tensor. I. Proton deshielding by steric compression. International Journal of Quantum Chemistry, 30(1):77–88, July 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Fox:1982:UEP

[FNPS82] Sidney W. Fox, Tadayoshi Nakashima, Aleksander Przybylski, and Robert M. Syren. The updated experimental proteinoid model. International Journal of Quantum Chemistry, 22(S9): 195–204, March 4–6, 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Folland:1981:DDM

[Fol81] N. O. Folland. Dynamical derivation of microscopic stress tensors. International Journal of Quantum Chemistry, 20(S15): 369–379, March 8–14, 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory, Collision Phenomena, and Computational Quantum Chemistry.

Fox:1981:OPS

[Fox81] Sidney W. Fox. Origins of the protein synthesis cycle. International Journal of Quantum Chemistry, 20(S8):441–454, March 5–7, 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Fox:1984:MSR

[Fox84] Sidney W. Fox. Molecular selection in the roots of evolved life and mind. International Journal of Quantum Chemistry, 26(S11):17–29, March 12–15, 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Fox:1986:MSU

[Fox86] Sidney W. Fox. Molecular selection in a unified evolutionary sequence. International Journal of Quantum Chemistry, 30(S13):223–235, March 6–8, 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Fox:1987:MPA

[Fox87] Sidney Fox. In memory of Professor Albert Szent-Györgyi. International Journal of Quantum Chemistry, 32(S14):1, March 14, 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Fabbri:1984:SCA

[FP84] M. Fabbri and L. F. Perondi. Statistical computer-aided calculation of molecular integrals. International Journal of Quantum Chemistry, 26(6):1029–1037, December 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Foley:1989:MMS

[FPDA89] C. K. Foley, L. G. Pedersen, T. A. Darden, and M. W. Anderson. Molecular mechanics studies of sequence-specific repair of DNA alkylated by EMS in the laci gene of Escherichia coli. International Journal of Quantum Chemistry, 36(S16):35–44, April 1–8, 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Fantucci:1981:DEM

[FPT81] Piercarlo Fantucci, Stefano Polezzo, and Laura Trombetta. Direct energy minimization of self-consistent electron pair by an orthogonal transformation of the basis. International Journal of Quantum Chemistry, 19(4):493–500, April 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Fricke:1984:CS

[FR84] B. Fricke and A. Rosén. Collision systems. International Journal of Quantum Chemistry, 25(1):97–105, January 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Fraczak:1984:SRA

[Fra84] Marek Fraczak. Statistical representation of atomic systems structure. I. Total binding energy as a function of the relative information content. International Journal of Quantum Chemistry, 25(5):755–766, May 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Fraczak:1985:SRA

[Fra85] Marek Fraczak. Statistical representation of atomic systems structure. II. classification of the representation; entropy dependence of the total energy for isoelectronic series. International Journal of Quantum Chemistry, 27(6):677–689, June 1985. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Fraga:1988:MEW

[Fra88] Serafin Fraga. A measure of the error in wave functions. International Journal of Quantum Chemistry, 34(6):505–507, December 1988. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Frauenfelder:1989:DWF

[Fra89] Hans Frauenfelder. The Debye–Waller factor: From villain to hero in protein crystallography. International Journal of Quantum Chemistry, 35(6):711–715, June 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Freidman:1980:I

[Fre80] Ambassador Bengt Freidman. Introduction. International Journal of Quantum Chemistry, 18(S14):1–4, March 10–20, 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory, Collision Phenomena, Quantum Statistics, and Computational Methods.

Friedman:1982:OR

[Fri82] Bengt Friedman. Opening remarks. International Journal of Quantum Chemistry, 22(S16):15–17, March 1–13, 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Chemistry, Theory of Condensed Matter, and Propagator Methods in the Quantum Theory of Matter.

Frishberg:1986:SDA

[Fri86] C. Frishberg. Slater determinant from atomic form factors. International Journal of Quantum Chemistry, 30(1):1–5, July 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Fritsche:1987:TCM

[Fri87] L. Fritsche. Theory of condensed matter and new materials. International Journal of Quantum Chemistry, 32(S21):15–29, March 12, 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Chemistry, Solid-State Theory, and Computational Methods.

Froelich:1981:BRB

[Fro81] Piotr Froelich. Book review: Ab Initio Calculations — Methods and Applications in Chemistry. By Petr Čársky and Miroslav Urban. Lecture Notes in Chemistry No. 16. Springer-Verlag, Berlin, 1980. ISBN 3-540-10005-9. International Journal of Quantum Chemistry, 20(6):1334–1335, December 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). See [ČU80].

Froelich:1983:BRB

[Fro83a] Piotr Froelich. Book review: Electronic states of molecular and atom clusters. By Giuseppe Del Re, Gaston Berthier, and Josiane Serre. Lecture Notes In Chemistry No. 13, Springer Verlag, Berlin, 1980. ISBN 3-540-09738-4. 177 pages. International Journal of Quantum Chemistry, 24(2):239, August 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). See [DBS80].

Frohlich:1983:ECE

[Frö83b] H. Fröhlich. Evidence for coherent excitation in biological systems. International Journal of Quantum Chemistry, 23(4): 1589–1595, April 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Froelich:1988:BRBa

[Fro88a] P. Froelich. Book review: Advances in chemical reaction dynamics. Edited by P. M. Rentzepis and Ch. Capellos, NATO ASI Series C, Vol. 184, Reidel, Dordrecht, 1986. International Journal of Quantum Chemistry, 34(5):483, November 1988. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). See erratum [RC86].

Froelich:1988:BRBb

[Fro88b] P. Froelich. Book review: Structure and dynamics of molecular systems II. Edited by R. Daudel, J. P. Korb, J. P. Lemaistre, and J. Maruani, Reidel, Dordrecht, 1986. International Journal of Quantum Chemistry, 34(5):485–486, November 1988. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). See [D^{+}85].

Flurry:1980:PSA

[FS80] R. L. Flurry, Jr. and T. H. Siddall III. Permutational– symmetry-adapted powers of representations of the point groups deduced from the unitary group. International Journal of Quantum Chemistry, 18(4):1037–1044, October 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Farazdel:1986:IUM

[FS86a] Abbas Farazdel and Vedene H. Smith, Jr. Invalidity of the ubiquitous mass-velocity operator in quasirelativistic theories. International Journal of Quantum Chemistry, 29(3):311–314, March 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Ford:1986:MMO

[FS86b] George P. Ford and Christopher T. Smith. An MNDO molecular orbital study of the reactions of protonated oxirane with simple nucleophiles. International Journal of Quantum Chemistry, 30(S13):107–119, March 6–8, 1986. CODEN IJQCB2.

ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Fournier:1986:EHC

[FS86c] R. Fournier and D. R. Salahub. Effect of hydrogen chemisorption on the magnetism of nickel clusters. International Journal of Quantum Chemistry, 29(5):1077–1089, May 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Ford:1987:MMO

[FS87] George P. Ford and Christopher T. Smith. An MNDO molecular orbital study of the reactions of protonated oxirane with guanine. International Journal of Quantum Chemistry, 32 (S14):57–64, March 14, 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Fraga:1979:AEL

[FSK79] Serafin Fraga, K. M. S. Saxena, and Jacek Karwowski. Atomic energy levels: data for parametric calculations, volume 4 of Physical sciences data. Elsevier, Amsterdam, The Netherlands, 1979. ISBN 0-444-41838-5. ix + 482 pp. LCCN QC454.A8 F73. fl 95.00.

Francisco:1987:BSG

[FSP87] E. Francisco, L. Seijo, and L. Pueyo. Basis sets generation: Relation between Adamowicz’s and the maximum overlap method. International Journal of Quantum Chemistry, 31 (2):279–285, February 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Fraga:1989:AEL

[FSS^{+}89] S. Fraga, E. San Fabian, J. A. Sordo, M. Campillo, J. D. Climenhaga, and M. Klobukowski. Atomic energy levels from configuration interaction calculations with relativistic corrections. International Journal of Quantum Chemistry, 35(2): 325–330, February 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Froelich:1987:ICC

[FSW87] Piotr Froelich, Erland Sangfelt, and Brian Weiner. Implementation of the complex coordinate method for resonances and photoionization cross sections by means of matrix element extrapolation. International Journal of Quantum Chemistry, 31 (5):823–831, May 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Fukui:1981:TC

[FTY81] Kenichi Fukui, Akitomo Tachibana, and Koichi Yamashita. Toward chemodynamics. International Journal of Quantum Chemistry, 20(S15):621–632, March 8–14, 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory, Collision Phenomena, and Computational Quantum Chemistry.

Fukui:1981:VPC

[Fuk81a] Kenichi Fukui. Variational principles in a chemical reaction. International Journal of Quantum Chemistry, 20(S15):633– 642, March 8–14, 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory, Collision Phenomena, and Computational Quantum Chemistry.

Fukutome:1981:UHF

[Fuk81b] Hideo Fukutome. Unrestricted Hartree–Fock theory and its applications to molecules and chemical reactions. International Journal of Quantum Chemistry, 20(5):955–1065, November 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Fukui:1983:CNC

[Fuk83] H. Fukui. Calculation of NMR chemical shifts. III. Configuration interaction effects. International Journal of Quantum Chemistry, 23(2):633–637, February 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Futrell:1987:CMB

[Fut87] Jean H. Futrell. Crossed-molecular beam studies of the stateto-state reaction dynamics of charge transfer at low and intermediate energy. International Journal of Quantum Chemistry, 31(1):133–159, January 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Fan:1988:CHF

[FVZ^{+}88] Liangyou Fan, Louis Versluis, Tom Ziegler, Evert Jan Baerends, and Walter Ravenek. Calculation of harmonic frequencies and harmonic force fields by the Hartree–Fock–Slater method. International Journal of Quantum Chemistry, 34 (S22):173–181, March 12–19, 1988. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Syposium on Quantum Chemistry, Solid-State Theory, and Computational Methods.

Fieck:1980:EIA

[FW80a] Gerhard Fieck and Joachim Wirsich. Electrostatic interaction for arbitrary point groups, configurations, and coupling schemes. International Journal of Quantum Chemistry, 18(3): 753–773, September 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Fieck:1980:OTP

[FW80b] Gerhard Fieck and Joachim Wirsich. One- and two-particle fractional parentage for arbitrary point groups, configurations, and coupling schemes. International Journal of Quantum Chemistry, 18(3):735–752, September 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Fox:1984:RAA

[FW84] Sidney W. Fox and Charles R. Windsor. Reproducibility of amino acid compositions in repeated copolymerizations of amino acids. International Journal of Quantum Chemistry, 26 (S11):103–108, March 12–15, 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Farazdel:1985:VMV

[FWS^{+}85] Abbas Farazdel, William M. Westgate, Alfredo M. Simas, Robin P. Sagar, and Vedene H. Smith, Jr. Validity of the massvelocity term in the Breit–Pauli Hamiltonian. International Journal of Quantum Chemistry, 28(S19):61–68, March 18–23, 1985. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Fedyanin:1982:PEP

[FY82] V. K. Fedyanin and L. V. Yakushevich. Particlelike excitations in the polypeptide chain model. International Journal of Quantum Chemistry, 21(6):1019–1028, June 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Fink:1986:CDM

[FZB86] M. Fink, Y. Zhang, and R. A. Bonham. Charge deformation maps, molecular moments, and high-energy electron scattering. International Journal of Quantum Chemistry, 30(S20): 613–625, March 10–15, 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Syposium on Atomic, Molecular, and Solid-State Theory, Scattering Problems, Many Body Phenomena, and Computational Quantum Chemistry.

Feng:1987:TAP

[FZZ87] Jikang Feng, Xuehe Zheng, and Michael C. Zerner. A theoretical analysis of photoaddition reactions of hydroxylazoaromatic compounds and the related thione analogs with olefins. International Journal of Quantum Chemistry, 32(S21):283–295, March 12, 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Chemistry, Solid-State Theory, and Computational Methods.

Grelland:1982:RLE

[GA82] Hans H. Grelland and Jan Almlöf. Reduced local energy as a criterion for the accuracy of approximate Hartree–Fock wave functions. Application to H_{2}O and CH_{4}. International Journal of Quantum Chemistry, 21(5):885–897, May 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Gallup:1982:TSP

[Gal82] G. A. Gallup. Transformation of spin projected Slater determinants under orbital permutation. International Journal of Quantum Chemistry, 21(6):1057–1060, June 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Gomes:1986:IME

[GAL86] Vivilí M. S. Gomes, Lucy V. C. Assali, and José R. Leite. Ab-initio MO electronic structure calculations of defect-pair complexes in silicon. International Journal of Quantum Chemistry, 30(S20):749–761, March 10–15, 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Syposium on Atomic, Molecular, and Solid-State Theory, Scattering Problems, Many Body Phenomena, and Computational Quantum Chemistry.

Ganas:1980:OOS

[Gan80] P. S. Ganas. Optical oscillator strengths for lithium. International Journal of Quantum Chemistry, 17(6):1179–1184, June 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Ganas:1981:EIE

[Gan81a] P. S. Ganas. Electron impact excitation cross sections for O II. International Journal of Quantum Chemistry, 20(5):1073– 1076, November 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Ganas:1981:OOS

[Gan81b] P. S. Ganas. Optical oscillator strengths for isoelectronic ions of nitrogen. International Journal of Quantum Chemistry, 19 (5):729–734, May 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Gascoyne:1980:ESR

[Gas80] Peter R. C. Gascoyne. Electron spin resonance and spectral studies of bovine serum albumin-methylglyoxal complexes. International Journal of Quantum Chemistry, 18(S7):93–100, March 5–8, 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Gascoyne:1981:ESR

[Gas81] Peter R. C. Gascoyne. An electron spin resonance investigation of amine models for the protein–methylglyoxal interaction. International Journal of Quantum Chemistry, 20(S8): 265–270, March 5–7, 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Gaspar:1983:SPT

[Gás83] R. Gáspár. Scaling in pseudopotential theory. International Journal of Quantum Chemistry, 23(3):1017–1024, March 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Gascoyne:1985:RMC

[Gas85] Peter Gascoyne. The redox-mediated control of enzyme function and cellular structure. International Journal of Quantum Chemistry, 28(S12):245–255, March 14–16, 1985. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Gascoyne:1986:TCP

[Gas86a] Peter R. C. Gascoyne. Theoretical criteria for the production of long-lived populations of cytotoxic free radicals. International Journal of Quantum Chemistry, 30(S13):195–207, March 6–8, 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Gaspar:1986:CRS

[Gás86b] R. Gáspár. Chairman’S remarks, session on systems containing heavy atoms: Paradoxes and problems of large-molecule and heavy-molecule calculations. International Journal of Quantum Chemistry, 29(5):1043–1045, May 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Gupta:1981:NPS

[GB81] Raj K. Gupta and Jeffrey L. Benovic. NMR probes of the spatial arrangement of substrates at the active site of creatine kinase. International Journal of Quantum Chemistry, 20(S8): 247–256, March 5–7, 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Gadre:1985:MAI

[GB85a] Shridhar R. Gadre and Rajeev D. Bendale. Maximization of atomic information-entropy sum in configuration and momentum spaces. International Journal of Quantum Chemistry, 28 (2):311–314, August 1985. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Golebiewski:1985:NCB

[GB85b] A. Gołebiewski and E. Broclawik. New concepts in bonded functions theory. I. alternative way of matrix elements evaluation. International Journal of Quantum Chemistry, 27(5):613– 623, May 1985. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Gritsenko:1986:MUC

[GBK86] O. V. Gritsenko, A. A. Bagaturjants, and V. B. Kazansky. Model universal Coulomb hole function for many-electron systems. International Journal of Quantum Chemistry, 29(6): 1799–1813, June 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Gutsev:1986:ESA

[GBO86a] G. L. Gutsev, A. I. Boldyrev, and A. A. Ovchinnikov. The electronic structure of all-trans-polyenes C_{2n}H_{2n+2}, n = 3−−7. International Journal of Quantum Chemistry, 30(5):647–661, November 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Gutsev:1986:EOE

[GBO86b] G. L. Gutsev, A. I. Boldyrev, and A. A. Ovchinnikov. Energies of optical excitations and ionization potentials for hydrocarbons with multiple bonds according to results of the DVM–X_{α} calculations. International Journal of Quantum Chemistry, 30 (5):625–631, November 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Golebiewski:1988:NCB

[GBW88] A. Gołebiewski, E. Broclawik, and M. Witko. New concepts in bonded functions theory II. Canonical set of high internal symmetry. International Journal of Quantum Chemistry, 33 (3):247–262, March 1988. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Golelbiewski:1989:NCB

[GBW89] A. Gołelbiewski, E. Brocławik, and M. Witko. New concepts in bonded functions theory. III. Orthogonalization problem. International Journal of Quantum Chemistry, 35(3):425–432, March 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Gritsenko:1986:CEP

[GBZ86] O. V. Gritsenko, A. A. Bagaturyants, and G. M. Zhidomirov. On the construction of the effective pair correlation function with the fixed zero value at the points where two electron positions coincide. International Journal of Quantum Chemistry, 30(6):791–797, December 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Goycoolea:1989:SIC

[GBZ89] Candelaria Goycoolea, Mauricio Barrera, and Fernando Zuloaga. Self-interaction-corrected electronegativities and hardness for atoms. International Journal of Quantum Chemistry, 36(4):455–472, October 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Gustav:1981:SSS

[GC81] Klaus Gustav and Roland Colditz. Suitability of a square step potential as an oscillator potential. International Journal of Quantum Chemistry, 19(3):379–381, March 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Gustav:1982:PAU

[GC82a] Klaus Gustav and Roland Colditz. Potential adaption: Use of polynomials as anharmonic oscillator potentials. International Journal of Quantum Chemistry, 21(6):1091–1095, June 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Gustav:1982:VCE

[GC82b] Klaus Gustav and Roland Colditz. Vibronic coupling of electronic states. II. mathematical description of the model. International Journal of Quantum Chemistry, 22(1):31–36, July 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Gustav:1983:VCE

[GC83] Klaus Gustav and Roland Colditz. Vibronic coupling of electronic states. III. The definitive breakdown of the CBO potential picture within the framework of our model. International Journal of Quantum Chemistry, 24(3):327–332, September 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Goscinski:1984:LL

[GC84a] Osvaldo Goscinski and Jean-Louis Calais. Löwdin lectures. International Journal of Quantum Chemistry, 26(2):295, August 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Gould:1984:SDU

[GC84b] M. D. Gould and G. S. Chandler. A spin-dependent unitary group approach to many-electron systems. International Journal of Quantum Chemistry, 26(4):441–455, October 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Gould:1984:UGAa

[GC84c] M. D. Gould and G. S. Chandler. Unitary group approach to the many-electron problem. I. Matrix element evaluation and shift operators. International Journal of Quantum Chemistry, 25(3):553–601, March 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Gould:1984:UGAb

[GC84d] M. D. Gould and G. S. Chandler. Unitary group approach to the many-electron problem. II. Adjoint tensor operators for U (n). International Journal of Quantum Chemistry, 25(3):603– 633, March 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Gould:1984:UGAc

[GC84e] M. D. Gould and G. S. Chandler. Unitary group approach to the many-electron problem. III. Matrix elements of spindependent Hamiltonians. International Journal of Quantum Chemistry, 25(6):1089–1109, June 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Gustav:1984:VCE

[GC84f] Klaus Gustav and Roland Colditz. Vibronic coupling of electronic states. IV. Harmonic CBO potentials with equal minima positions—the influence of different vibronic coupling strengths on the electronic absorption and emission spectra containing a longest-wavelength electronically forbidden transition. International Journal of Quantum Chemistry, 25(4): 687–697, April 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Ghose:1985:MOS

[GC85a] Arup K. Ghose and Gordon M. Crippen. Molecular orbital study of the ring nitrogen basicity of various dihydrofolate reductase inhibitors. International Journal of Quantum Chemistry, 28(3):315–334, September 1985. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Gould:1985:UGA

[GC85b] M. D. Gould and G. S. Chandler. Unitary group approach to the many-electron problem. I, II, III. International Journal of Quantum Chemistry, 27(6):787–801, June 1985. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Gould:1987:BIR

[GC87] M. D. Gould and G. S. Chandler. Bases for the irreducible representations of O(3) symmetry adapted to a crystallographic point group. International Journal of Quantum Chemistry, 31 (3):535–563, March 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Gresh:1982:CII

[GCP82] N. Gresh, P. Claverie, and A. Pullman. Computations of intermolecular interactions: Expansion of a charge-transfer energy contribution in the framework of an additive procedure. Applications to hydrogen-bonded systems. International Journal of Quantum Chemistry, 22(1):199–215, July 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Gresh:1986:IIE

[GCP86] N. Gresh, P. Claverie, and A. Pullman. Intermolecular interactions: Elaboration on an additive procedure including an explicit charge-transfer contribution. International Journal of Quantum Chemistry, 29(1):101–118, January 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Gao:1987:PGS

[GCP87] M. J. Gao, J. Q. Chen, and J. Paldus. Point group symmetry adaptation in Clifford algebra unitary group approach. International Journal of Quantum Chemistry, 32(1):133–147, July 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Gutowski:1984:EBS

[GCV84] Maciej Gutowski, Grzegorz Chałasiński, and Jeanne Van Duijneveldt-Van De Rijdt. Effective basis sets for calculations of exchange-repulsion energy. International Journal of Quantum Chemistry, 26(6):971–982, December 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Garland:1982:CSV

[GD82a] Gregory E. Garland and James W. Dufty. Computer simulation of the velocity autocorrelation function at low density. International Journal of Quantum Chemistry, 22(S16): 91–100, March 1–13, 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Chemistry, Theory of Condensed Matter, and Propagator Methods in the Quantum Theory of Matter.

Ghosh:1982:QFD

[GD82b] S. K. Ghosh and B. M. Deb. Quantum fluid dynamics of many-electron systems in three-dimensional space. International Journal of Quantum Chemistry, 22(5):871–888, November 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Goscinski:1989:HA

[GD89] Osvaldo Goscinski and Joseph Delhalle. Hyperpolarizability approximations. International Journal of Quantum Chemistry, 35(6):761–767, June 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Gianturco:1987:RPD

[GDBRV87] F. A. Gianturco, G. Delgado-Barrio, O. Roncero, and P. Villarreal. Rotational predissociation dynamics in weakly bound molecular systems: The ArN_{2} and ArO_{2} examples. International Journal of Quantum Chemistry, 32(S21):389–405, March 12, 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Chemistry, Solid-State Theory, and Computational Methods.

Gurnani:1981:ITL

[GDM^{+}81] S. Gurnani, G. Datta, N. B. Mulchandani, G. Sen, S. K. Mehta, and S. Sengupta. Interaction of tylophorinidine with lysozyme. International Journal of Quantum Chemistry, 20 (2):523–531, August 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Grossman:1985:GTA

[GDR85] Stuart C. Grossman, Borka Jerman-Blažić Džonova, and Milan Randić. A graph theoretical approach to quantitative structure-activity relationship. International Journal of Quantum Chemistry, 28(S12):123–139, March 14–16, 1985. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Guangju:1989:TSC

[GDX89] Chen Guangju, Fang Decai, and Fu Xiaoyuan. Theoretical studies on the cycloaddition reaction of ethylene and formaldehyde. International Journal of Quantum Chemistry, 36(S23): 501–508, April 1–8, 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Gresh:1981:TSS

[GEdP81] N. Gresh, C. Etchebest, O. de la Luz Rojas, and A. Pullman. A theoretical study of the selective alkali and alkaline-earth cation binding properties of valinomycin. International Journal of Quantum Chemistry, 20(S8):109–116, March 5–7, 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Geertsen:1988:ACC

[Gee88] Jan Geertsen. An approximate coupled cluster doubles polarization propagator approximation. International Journal of Quantum Chemistry, 34(S22):491–495, March 12–19, 1988. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Syposium on Quantum Chemistry, Solid-State Theory, and Computational Methods.

Geertsen:1989:BRB

[Gee89] Jan Geertsen. Book review: Quantum description of highresolution NMR in liquids. By Maurice Goldman, Clarendon Press, Oxford, 1988. International Journal of Quantum Chemistry, 36(4):533, October 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). See [Gol88].

Gerloch:1986:OTS

[Ger86] M. Gerloch. Orbitals, terms and states. John Wiley, New York, NY, USA, 1986. ISBN 0-471-90935-1, 0-471-90936-X (paperback). xii + 176 pp. LCCN QD461 .G37 1986.

Georgian:1983:HFC

[GF83] T. Georgian and G. L. Findley. Hamiltonian formulations of chemical kinetics. International Journal of Quantum Chemistry, 24(S10):331–338, March 14–16, 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Guimaraes:1988:IDO

[GF88] P. S. Guimarães and A. Ferreira Da Silva. Interacting donors in the optical absorption of semiconductors. International Journal of Quantum Chemistry, 34(S22):21–24, March 12–19, 1988. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Syposium on Quantum Chemistry, Solid-State Theory, and Computational Methods.

Gaspar:1981:IMF

[GG81] R. Gáspár, Jr. and R. Gáspáar. Ab initio molecular fragment calculations with pseudopotentials. Some fragments containing nitrogen and oxygen. International Journal of Quantum Chemistry, 19(4):501–503, April 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Gaspar:1982:IMCb

[GG82a] R. Gáspár and R. Gáspár, Jr. Ab initio molecular calculations with pseudopotentials: Higher quality calculations on Li_{2}, LiH, and BeH_{2}. International Journal of Quantum Chemistry, 22(2):421–427, August 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Gaspar:1982:IMCa

[GG82b] R. Gáspár, Jr. and R. Gáspáar. Ab initio molecular calculations with pseudopotentials: Calculations of double-zeta quality on ethylene, acetylene, and water. International Journal of Quantum Chemistry, 22(2):415–420, August 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Gaspar:1983:IMF

[GG83] R. Gáspár, Jr. and R. Gáspáar. Ab initio molecular fragment calculations with pseudopotentials: Model peptide studies. International Journal of Quantum Chemistry, 24(6):767–771, December 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Gaspar:1985:BSS

[GG85] R. Gáspár and R. Gáspár, Jr. Basis set study in pseudopotential method. International Journal of Quantum Chemistry, 28 (S19):279–284, March 18–23, 1985. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Georgiev:1986:QTD

[GGDG86] M. Georgiev, A. Gochev, and A. Diaz-Gongora. Quantum theory of dielectric relaxation in polar solids. International Journal of Quantum Chemistry, 29(5):1365–1372, May 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Georgiev:1987:QTD

[GGDG87] M. Georgiev, A. Gochev, and A. Diaz-Gongora. Quantum theory of dielectric relaxation in polar solids. International Journal of Quantum Chemistry, 32(3):405–406, September 1987.

CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Gupta:1981:SMP

[GGM81] V. P. Gupta, V. D. Gupta, and C. Mehrotra. Study of molecular polarizabilities of some amino and polyamino acids. International Journal of Quantum Chemistry, 20(2):373–376, August 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Georgian:1984:HDS

[GHF84] T. Georgian, J. M. Halpin, and G. L. Findley. Hamiltonian dynamics of simple chemical reactions. International Journal of Quantum Chemistry, 26(S11):347–353, March 12–15, 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Gill:1989:EAG

[GHGP89] Peter M. W. Gill, Martin Head-Gordon, and John A. Pople. An efficient algorithm for the generation of two-electron repulsion integrals over Gaussian basis functions. International Journal of Quantum Chemistry, 36(S23):269–280, April 1–8, 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Girardeau:1980:SQC

[Gir80] M. D. Girardeau. Second quantization for composite particles, reactive collisions, and unstable particles. International Journal of Quantum Chemistry, 17(1):25–35, January 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Girardeau:1986:THC

[Gir86] M. D. Girardeau. Theory of half-collision cross sections. International Journal of Quantum Chemistry, 29(5):1483–1492, May 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Gomez-Jeria:1983:SPQ

[GJ83] J. S. Gomez-Jeria. On some problems in quantum pharmacology I. The partition functions. International Journal of Quantum Chemistry, 23(6):1969–1972, June 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Gomez-jeria:1986:TSL

[GjC86] J. S. Gómez-jeria and Renato R. Contreras. Theoretical study of lithium-fluoride and lithium-chloride ion pairs in aqueous solution. An SCF-CNDO /2 approach including continuum solvent effects. International Journal of Quantum Chemistry, 30(5):581–590, November 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Gomez-Jeria:1985:QCS

[GJMLRGSA85] J. S. Gomez-Jeria, D. Morales-Lagos, J. I. Rodriguez-Gatica, and J. C. Saavedra-Aguilar. Quantum-chemical study of the relation between electronic structure and pA_{2} in a series of 5-substituted tryptamines. International Journal of Quantum Chemistry, 28(4):421–428, October 1985. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Gondo:1980:NSA

[GK80a] Yasuhiko Gondo and Yoshiya Kanda. Novel sign-alternation rule for the triplet state in aromatic systems. International Journal of Quantum Chemistry, 18(1):185–188, July 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Gortel:1980:QST

[GK80b] Z. W. Gortel and H. J. Kreuzer. Quantum statistical theory of localized physisorption. International Journal of Quantum Chemistry, 18(S14):617–628, March 10–20, 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory, Collision Phenomena, Quantum Statistics, and Computational Methods.

Gumbinger:1980:SCM

[GK80c] H. G. Gumbinger and P. M. Kaiser. SCF CI MO calculations on the base–base interactions in dinucleoside monophosphates. International Journal of Quantum Chemistry, 18(2):439–448, August 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Gutowski:1981:AEE

[GKAP81] Maciej Gutowski, Małgorzata Kakol, Jan Andzelm, and Lucjan Piela. Approximate exchange and electrostatic interaction energies of deformed ions. International Journal of Quantum Chemistry, 19(3):401–411, March 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Graovac:1980:SOM

[GKMV80] Ante Graovac, Krešimir Kovačević, Zvonimir B. Maksi, and Ahmet Veseli. Slater-orbital molecular integrals by numerical Fourier-transform methods. II. Four-center integrals over is orbitals. International Journal of Quantum Chemistry, 17 (4):747–757, April 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Gutowski:1983:IEE

[GKP83] Maciej Gutowski, Małgorzata Kalkol, and Lucjan Piela. Importance of exchange effects in the deformation of interacting ions. International Journal of Quantum Chemistry, 23 (5):1843–1853, May 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Glass:1985:TAE

[Gla85] R. Glass. Theoretical approach to the evaluation of spinonly magnetic form factors for many-electron atomic systems. International Journal of Quantum Chemistry, 28(5):649–660, November 1985. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Gianturco:1981:VIS

[GLB81] F. A. Gianturco, U. T. Lamanna, and F. Battaglia. Vibrationally inelastic scattering and relaxation times in gaseous HF. International Journal of Quantum Chemistry, 19(2): 217–236, February 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Gonzaalez-Lafont:1986:TSS

[GLLOB86a] A. Gonzáalez-Lafont, J. M. Lluch, A. Oliva, and J. Bertrágn. Theoretical study of several Fe(H_{2}O)_{n}^{2+} clusters at different temperatures. International Journal of Quantum Chemistry, 29(5):1373–1382, May 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Gonzalez-Lafont:1986:IPF

[GLLOB86b] A. Gonzàlez-Lafont, J. M. Lluch, A. Oliva, and J. Bertrán. An intermolecular potential function for the Fe(II)–H_{2}O system from ab initio calculations. International Journal of Quantum Chemistry, 30(5):663–670, November 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Gonzalez-Lafont:1988:API

[GLLOB88] A. González-Lafont, J. M. Lluch, A. Oliva, and J. Bertrán. Analytical potential from ab initio calculations for the Fe^{+}– H_{2}O and Fe^{0}-H_{2}O systems. International Journal of Quantum Chemistry, 33(2):77–85, February 1988. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Geller:1980:EPA

[GLP80] M. Geller, B. Lesyng, and A. Pohorille. Energy of planar autoassociation of xanthine and spatial structure of poly(riboxanthylic acid). International Journal of Quantum Chemistry, 17(2):293–298, February 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Gubanov:1983:CXA

[GLP83] V. A. Gubanov, A. I. Liechtenstein, and A. V. Postnikov. Cluster X_{α} approach to magnetic interactions and phase transitions in solids. International Journal of Quantum Chemistry, 23(4):1517–1528, April 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Gabbay:1980:MSI

[GM80] I. Gabbay and N. H. March. Metal surfaces in interaction with molecular adsorbates. International Journal of Quantum Chemistry, 18(S14):665–679, March 10–20, 1980. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory, Collision Phenomena, Quantum Statistics, and Computational Methods.

Ghosh:1981:EBS

[GM81a] Subhendu Ghosh and Anjali Mookerjee. Estimation of binding sites from the adsorption profile of complexed DNA. International Journal of Quantum Chemistry, 20(1):185–198, July 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Graf:1981:DDI

[GM81b] P. Graf and E. L. Mehler. Direct determination of improved contraction coefficients by a least-squares approach. International Journal of Quantum Chemistry, 20(S8):63–71, March 5–7, 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Graf:1981:ESG

[GM81c] P. Graf and E. L. Mehler. Evaluation of small Gaussian basis sets for ab-initio calculations on biologically active molecules. International Journal of Quantum Chemistry, 20(S8):49–61, March 5–7, 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Goldberg:1983:QMC

[GM83] M. J. Goldberg and L. J. Massa. Quantum model of coherent X-ray diffraction: Extension to Bloch orbitals. International Journal of Quantum Chemistry, 24(1):113–126, July 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Gazdy:1986:VIT

[GM86a] Bela Gazdy and David A. Micha. Variationally improved transition amplitudes from time-dependent Hartree–Fock wave functions: Application to He + He^{2+} collisions. International Journal of Quantum Chemistry, 30(S20):689–697, March 10– 15, 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Syposium on Atomic, Molecular, and Solid-State Theory, Scattering Problems, Many Body Phenomena, and Computational Quantum Chemistry.

Goscinski:1986:ACS

[GM86b] Osvaldo Goscinski and Vladimiro Mujica. Adiabatic coordinate separation and large N-dimensional limit in two-electron ions. International Journal of Quantum Chemistry, 29(4):897– 908, April 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Guha:1986:SEP

[GM86c] Sephali Guha and Devashis Majumdar. A study of the electrostatic potential maps for several GABA (γ-aminobutyric acid) inhibitors of restricted conformation. International Journal of Quantum Chemistry, 30(S13):19–37, March 6–8, 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Garbato:1981:BMS

[GMM81] L. Garbato, P. Manca, and G. Mula. Bond model for semiconducting materials. International Journal of Quantum Chemistry, 19(6):1185–1189, June 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Gadea:1984:EVB

[GMM84] Xavier Gadea, Daniel Maynau, and Jean-Paul Malrieu. Effective valence bond Hamiltonian for the cations of conjugated hydrocarbons. International Journal of Quantum Chemistry, 26(1):1–20, July 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Gayoso:1980:RCS

[GMOBA80] J. Gayoso, B. Maouche, O. Ouamerali, and A. Baba-Ahmed. “ring currents” static hypersusceptibilities. International Journal of Quantum Chemistry, 18(1):117–124, July 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Gascoyne:1984:QAF

[GMP^{+}84a] Peter R. C. Gascoyne, Jane A. McLaughlin, Ronald Pethig, Albert Szent-Györgyi, Chiu-Nan Lai, and Frederick F. Becker. Quenching of anionic free radicals by normal and transformed cells: A probe of phenotypic changes. International Journal of Quantum Chemistry, 26(S11):309–314, March 12–15, 1984.

CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Green:1984:MMC

[GMP84b] A. E. S. Green, R. R. Meyreddy, and K. M. Pamidimukkala. A molecular model of coal pyrolysis. International Journal of Quantum Chemistry, 26(S18):589–599, March 1–15, 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Atomic, Molecular and Solid-State Theory, and Computational Quantum Chemistry.

Garbato:1981:DSB

[GMS81] L. Garbato, P. Manca, and A. Spiga. Diamagnetism as a structure and bonding related property in semiconductors. International Journal of Quantum Chemistry, 19(6):1191–1196, June 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Grinberg:1983:SRH

[GMV83] Horacio Grinberg, Julio Marañon, and Hector Vucetich. Some remarks on the hydrogen atom. International Journal of Quantum Chemistry, 23(2):379–385, February 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Gaspar:1987:GHF

[GN87] R. Gáspár and Á. Nagy. Generalized Hellmann–Feynman theorem in the Xα method. International Journal of Quantum Chemistry, 31(4):639–647, April 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Golden:1980:GHS

[Gol80] Sidney Golden. Generalized Hartree-Slater theory: Lower bounds to Hartree–Fock energies. International Journal of Quantum Chemistry, 18(S14):135–142, March 10–20, 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory, Collision Phenomena, Quantum Statistics, and Computational Methods.

Golovko:1984:IDP

[Gol84] V. F. Golovko. Iterative degenerate perturbation theory. International Journal of Quantum Chemistry, 26(4):463–478, October 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Goldman:1988:QDH

[Gol88] M. (Maurice) Goldman. Quantum description of highresolution NMR in liquids, volume 15 of The International series of monographs on chemistry. Clarendon Press, New York, NY, USA, 1988. ISBN 0-19-855639-X. xiv + 268 pp. LCCN QD96.N8 G65 1988. US$70.00. URL http://www.loc.gov/

catdir/enhancements/fy0639/88001668-d.html;   http:
//www.loc.gov/catdir/enhancements/fy0639/88001668-
t.html.

Gomes:1987:ITE

[GOLC87] V. M. S. Gomes, G. M. G. Oliveira, J. R. Leite, and A. S. Chaves. Intersubband transition energies in quantum wells in n-type GaAs-Al_{x}Ga_{1−x}As heterostructures. International Journal of Quantum Chemistry, 32(6):655–661, December 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Goscinski:1980:BRB

[Gos80] Osvaldo Goscinski. Book review: Excited states in quantum chemistry. Edited by Cleantes Nicolaides and D. R. Beck. D. Reidel Publishing Co., Dordrecht, 1979. 576 pp. Price: $59.00. International Journal of Quantum Chemistry, 17(1):149, January 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). See [NB79].

Goscinski:1982:AGP

[Gos82a] Osvaldo Goscinski. Antisymmetrized geminal power (AGP) wave-functions in finite and extended systems. International Journal of Quantum Chemistry, 22(S16):591–603, March 1– 13, 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Chemistry, Theory of Condensed Matter, and Propagator Methods in the Quantum Theory of Matter.

Goscinski:1982:SPR

[Gos82b] Osvaldo Goscinski. Some properties of reduced density matrices, correlated and uncorrelated, for pure and mixed states. International Journal of Quantum Chemistry, 21(1):269–273, January 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Goscinski:1985:LBS

[Gos85] Osvaldo Goscinski. Are localized broken symmetry solutions acceptable in molecular calculations? International Journal of Quantum Chemistry, 28(S19):51–59, March 18–23, 1985. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Geertsen:1987:CSS

[GOS87a] Jan Geertsen, Jens Oddershede, and Gustavo E. Scuseria. Calculation of spectra and spin–spin coupling constants using a coupled–cluster polarization propagator method. International Journal of Quantum Chemistry, 32(S21):475–485, March 12, 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Chemistry, Solid-State Theory, and Computational Methods.

Goscinski:1987:I

[Gos87b] Osvaldo Goscinski. Introduction. International Journal of Quantum Chemistry, 31(5):697, May 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Gould:1986:UGAb

[Gou86] M. D. Gould. Unitary group approach to general system partitioning. II. U (n) matrix element evaluation in a composite basis. International Journal of Quantum Chemistry, 30(3): 365–389, September 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Gupta:1980:MAL

[GP80] Raj K. Gupta and John M. Pesando. Mechanism of activitylinked ionization in carbonic anhydrase. International Journal of Quantum Chemistry, 18(S7):369–373, March 5–8, 1980.

CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Gresh:1982:TSI

[GP82] Nohad Gresh and Alberte Pullman. Theoretical study of the interaction of nonactin with Na^{+}, K^{+}, and NH. International Journal of Quantum Chemistry, 22(4):709–716, October 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Gresh:1983:TSM

[GP83a] Nohad Gresh and Alberte Pullman. Theoretical study of the 2: 1 Mg^{2+} and Ca^{2+} complexes of ionophore A23187. International Journal of Quantum Chemistry, 24(S10):215–229, March 14–16, 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Gresh:1983:TSI

[GP83b] Nohad Gresh and Bernard Pullman. Theoretical study of the interaction of tetramethylammonium with double-stranded oligonucleotides. International Journal of Quantum Chemistry, 24(5):491–507, November 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Gresh:1985:TSS

[GP85] Nohad Gresh and Bernard Pullman. A theoretical study of the selective entrapment of alkali and ammonium cations between guanine tetramers. International Journal of Quantum Chemistry, 28(S12):49–56, March 14–16, 1985. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Gould:1986:UGAa

[GP86a] M. D. Gould and J. Paldus. Unitary group approach to general system partitioning. I. Calculation of U(n = n_{1} +n_{2}): U(n_{1})× u(n_{2}) reduced matrix elements and reduced Wigner coefficients. International Journal of Quantum Chemistry, 30(3): 327–363, September 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Greenwood:1986:ITH

[GP86b] H. H. Greenwood and J. S. Plant. Ab initio theory of hydrogen bonding in vitamin B_{6}. International Journal of Quantum Chemistry, 30(1):127–136, July 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Gresh:1985:CLI

[GPC85] N. Gresh, A. Pullman, and P. Claverie. Cation–ligand interactions: Reproduction of extended basis set Ab initio SCF computations by the SIBFA 2 additive procedure. International Journal of Quantum Chemistry, 28(6):757–771, December 1985. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Gianturco:1985:CTR

[GPLP85] F. A. Gianturco, A. Palma, V. T. Lamanna, and G. Petrella. Collisional transfer of rovibrational energy from quantum calculations. II. The case of LiH with He. International Journal of Quantum Chemistry, 28(4):499–516, October 1985. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Garcia-Prieto:1980:NEM

[GPN80] J. García-Prieto and O. Novaro. Nonadditive effects in metal clusters and chemisorption. Pseudopotential study of palladium clusters. International Journal of Quantum Chemistry, 18(2):595–600, August 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Geldart:1989:EIS

[GR89] D. J. W. Geldart and M. Rasolt. Effective interactions, structure factors, and nonlocal correlations in many-electron systems. International Journal of Quantum Chemistry, 36(S23): 631–639, April 1–8, 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Graffi:1982:MPT

[Gra82] S. Graffi. Meaning of the perturbation theory for a class of multiple-well anharmonic oscillators. International Journal of Quantum Chemistry, 21(1):195–197, January 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Grant:1984:RAS

[Gra84] I. P. Grant. Relativistic atomic structure theory: Some recent work. International Journal of Quantum Chemistry, 25(1):23– 46, January 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Guldberg:1987:NSG

[GRBP87] Annette Guldberg, Sten Rettrup, Gian Luigi Bendazzoli, and Paolo Palmieri. A new symmetric group program for direct configuration interaction studies of molecules. International Journal of Quantum Chemistry, 32(S21):513–521, March 12, 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Chemistry, Solid-State Theory, and Computational Methods.

Green:1980:SDM

[Gre80] H. S. Green. Semigroups and the density matrix formulation of quantum mechanics. International Journal of Quantum Chemistry, 17(1):121–132, January 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Grelland:1981:ERA

[Gre81a] Hans H. Grelland. An Einstein relativistic atomic and molecular model based on the Horwitz–Piron–Reuse theory. International Journal of Quantum Chemistry, 19(5):873–883, May 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Grelland:1981:LEA

[Gre81b] Hans H. Grelland. n-Local energies as accuracy measures on approximate wave functions. International Journal of Quantum Chemistry, 19(5):847–857, May 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Gresh:1982:CNT

[Gre82] Nohad Gresh. Carcinogenicity of naphthylamines: a theoretical study of the stacking interaction of 1- and 2naphthylnitrenium ions with guanine, adenine, and the dinucleoside monophosphate cytidyl-(3′-5′)-guanosine. International Journal of Quantum Chemistry, 22(1):37–48, July 1982.

CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Gready:1987:AIR

[Gre87] Jill E. Gready. Analysis of inter-ring coupling effects in Nheterobicyclic π-systems using a structural definition of aromaticity. International Journal of Quantum Chemistry, 31 (3):369–382, March 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Grishin:1981:SFB

[Gri81] E. V. Grishin. Structure and function of Buthus eupeus scorpion neurotoxins. International Journal of Quantum Chemistry, 20(2):291–298, August 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Grossman:1985:COM

[Gro85] Stuart C. Grossman. Chemical ordering of molecules: a graph theoretical approach to structure-property studies. International Journal of Quantum Chemistry, 28(1):1–16, July 1985. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Grigoras:1980:HHE

[GRP^{+}80] S. Grigoras, I. Rusu, E. Pausescu, A. Medesan, and S. Moldoveanu. Hydrophile–hydrophobe equilibrium of some prostaglandins — CNDO/SW study. International Journal of Quantum Chemistry, 18(2):501–508, August 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Green:1982:APS

[GRS^{+}82] A. E. S. Green, D. E. Rio, P. F. Schippnick, J. M. Schwartz, and P. S. Ganas. Analytic phase shifts for Yukawa potentials. International Journal of Quantum Chemistry, 22(S16):331– 343, March 1–13, 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Chemistry, Theory of Condensed Matter, and Propagator Methods in the Quantum Theory of Matter.

Grudzinski:1985:RCG

[Gru85] Hubert Grudziński. N-representability conditions generated by a one-particle Grassmann factor of an antisymmetric function. International Journal of Quantum Chemistry, 27(6):709– 730, June 1985. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Guo-Sen:1980:GTF

[GS80] Yan Guo-Sen. The graph theoretical formulas for determinant expansions. International Journal of Quantum Chemistry, 18 (S14):549–555, March 10–20, 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory, Collision Phenomena, Quantum Statistics, and Computational Methods.

Garcia-Sucre:1989:EMS

[GS89] M. García-Sucre. The energy of a molecule as a sum of terms related to its nuclei when the nuclear motion is quantum mechanically described*. International Journal of Quantum Chemistry, 35(3):433–439, March 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Garcia-Sucre:1989:ENM

[GSC89] M. García-Sucre and V. Castells. The effect of nuclear motion on Berlin surfaces of a diatomic fragment in a polyatomic molecule. International Journal of Quantum Chemistry, 36 (S23):79–88, April 1–8, 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Gascoyne:1984:AEE

[GSG84] Peter R. C. Gascoyne and Albert Szent-Györgyi. Anomalous electrochemical and electron spin resonance properties of some biologically relevant methoxyl-substituted quinones. International Journal of Quantum Chemistry, 26(S11):217–222, March 12–15, 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Garcia-Sucre:1986:MRQ

[GSL86] M. Garcia-Sucre and R. Lefebvre. Multichannel resonance quantization with an optical potential. International Journal of Quantum Chemistry, 30(S20):541–546, March 10–15, 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Syposium on Atomic, Molecular, and Solid-State Theory, Scattering Problems, Many Body Phenomena, and Computational Quantum Chemistry.

Gascoyne:1983:SET

[GSM83] Peter R. C. Gascoyne, Martyn C. R. Symons, and Jane A. Mclaughlin. Spontaneous electron transfer in the reaction between methylgiyoxal and methylamine. International Journal of Quantum Chemistry, 24(S10):123–132, March 14–16, 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Garcia-Sucre:1985:IEA

[GSMRM85] M. García-Sucre, R. Medina, I. Reif, and G. J. Mata. Ionization energies of atomic systems obtained from nuclear and electronic charge scalings. International Journal of Quantum Chemistry, 28(S19):349–361, March 18–23, 1985. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Gascoyne:1982:FRP

[GSMSG82] Peter R. C. Gascoyne, Martyn C. R. Symons, Jane A. Mclaughlin, and Albert Szent-Györgyi. Free radicals produced in the interaction of cysteine with carbonyls of biological relevance. International Journal of Quantum Chemistry, 22(S9): 137–143, March 4–6, 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Guner:1987:PEM

[GSO^{+}87] O. F. Guner, D. D. Shillady, R. M. Ottenbrite, B. K. Rao, and E. Yurtsever. Pair-excitation MCSCF treatment of small molecules in an optimized Slater–Transform–Preuss basis set. International Journal of Quantum Chemistry, 32(5):551–562, November 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Green:1983:RHO

[GSR83] A. E. S. Green, P. F. Schippnick, and D. E. Rio. Reactive hydrogen–oxygen collisions in combustion and radiolysis. International Journal of Quantum Chemistry, 24(S17):127–136, March 3–12, 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-state Theory, Collision Phenomena, and Computational Quantum Chemistry.

Green:1980:AME

[GSRG80] A. E. S. Green, P. F. Schippnick, D. E. Rio, and P. S. Ganas. An analytic model of electron-oxygen elastic scattering in the 1–500-eV range. International Journal of Quantum Chemistry, 18(S14):71–82, March 10–20, 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory, Collision Phenomena, Quantum Statistics, and Computational Methods.

Green:1985:QME

[GSS85a] A. E. S. Green, J. M. Schwartz, and S. T. Suh. Quantum mechanical engineering of short range potentials. International Journal of Quantum Chemistry, 28(S19):411–419, March 18– 23, 1985. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Gupta:1985:DSS

[GSS85b] Goutam Gupta, Mukti H. Sarma, and Ramaswamy H. Sarma. Determination of the solution structure of nucleic acid double helices using NOE difference NMR spectroscopy, chemical shift calculations, and computer modeling. International Journal of Quantum Chemistry, 28(S12):183–199, March 14–16, 1985. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Guo:1983:AR

[GT83] Hong Guo and Au-Chin Tang. Approach to N representability. International Journal of Quantum Chemistry, 23(2):399–404, February 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Garrett:1986:TSS

[GT86] Bruce C. Garrett and Donald G. Truhlar. Thermal and stateselected rate constant calculations for O(^{3}p) + H_{2} → OH + H and isotopic analogs. International Journal of Quantum Chemistry, 29(5):1463–1482, May 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Garrett:1987:RRH

[GT87a] Bruce C. Garrett and Donald G. Truhlar. Reaction rates for O + HD → OH + D and O + HD → OD + H. International Journal of Quantum Chemistry, 31(1):17–31, January 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Gordon:1987:CBB

[GT87b] Mark S. Gordon and Donald G. Truhlar. Correlation balance in basis sets for electronic structure calculations. International Journal of Quantum Chemistry, 31(1):81–90, January 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

George:1980:OEE

[GTB80] Philip George, Mendel Trachtman, and Charles W. Bock. π-orbital electron–energy contributions to chemical reaction heats, I. Reactions involving only linear molecules containing up to three first–row atoms. International Journal of Quantum Chemistry, 17(4):599–608, April 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Guberman:1982:DEA

[Gub82] Steven L. Guberman. The doubly excited autoionizing states of H_{2}. International Journal of Quantum Chemistry, 22(S16): 668, March 1–13, 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Chemistry, Theory of Condensed Matter, and Propagator Methods in the Quantum Theory of Matter.

Gupta:1980:SMI

[Gup80] Raj K. Gupta. On the state of the magnesium ion in intact cells as observed by noninvasive ^{31}P NMR spectroscopy. International Journal of Quantum Chemistry, 18(S7):67–73, March 5–8, 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Gupta:1981:DCP

[Gup81] V. D. Gupta. Dynamics and conformation of polypeptides. International Journal of Quantum Chemistry, 20(1):9–21, July 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Guseinov:1985:EME

[Gus85] I. I. Guseinov. Evaluation of multicenter electron-repulsion integrals over Slater-type orbitals. International Journal of Quantum Chemistry, 28(S19):149–155, March 18–23, 1985. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Guseinov:1987:UAE

[Gus87] I. I. Guseinov. Unified analytical evaluation of the two-center Coulomb and hybrid integrals over Slater-type orbitals. International Journal of Quantum Chemistry, 31(4):565–567, April 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Gupta:1989:DBS

[GUSS89] Goutam Gupta, Kimiko Umemoto, Mukti H. Sarma, and Ramaswamy H. Sarma. DNA bending in solution: NMR studies on the structural roles of A/T tracts and the sequences at the junction. International Journal of Quantum Chemistry, 36 (S16):17–33, April 1–8, 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Gazquez:1988:NAS

[GV88] José L. Gázquez and Alberto Vela. A new approach to second order corrections based on density functional theory. International Journal of Quantum Chemistry, 34(S22):71–76, March 12–19, 1988. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Syposium on Quantum Chemistry, Solid-State Theory, and Computational Methods.

Gyemant:1980:REE

[GVB80] I. Gyemant, Zs. Varga, and M. G. Benedict. Resonances in the elastic electron-SF_{6} molecule scattering. International Journal of Quantum Chemistry, 17(2):255–263, February 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Garcia-Vela:1989:QAS

[GVVDB89] A. García-Vela, P. Villarreal, and G. Delgado-Barrio. Quantal approach to study the structure of triatomic I_{2}⋯X (X = He, Ne) VDW clusters. International Journal of Quantum Chemistry, 36(5):633–646, November 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Graham:1987:EEO

[GY87] Richard L. Graham and Danny L. Yeager. Excitation energies, oscillator strengths, and frequency dependent polarizabilities of Be: Comparison of TDHF, EOM (second order), and MCTDHF. International Journal of Quantum Chemistry, 31 (1):99–112, January 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Golab:1983:PCM

[GYJ83] Joseph T. Golab, Danny L. Yeager, and Poul Jørgensen. Proper characterization of MCSCF stationary points. International Journal of Quantum Chemistry, 24(S17):645, March 3–12, 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-state Theory, Collision Phenomena, and Computational Quantum Chemistry.

Guan-Zhi:1988:TKT

[GZWYCHC88] Ju Guan-Zhi, Ma Wan-Yong, Deng Cong-Hao, and Kurt Christoffel. Theoretical kinetic and thermodynamic study on the disilene-silylsilylene isomerization. International Journal of Quantum Chemistry, 34(S22):343–347, March 12–19, 1988.

CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Syposium on Quantum Chemistry, Solid-State Theory, and Computational Methods.

Hosoya:1990:QCL

[H^{+}90] H. Hosoya et al., editors. Quantum chemistry literature data base: Supplement 9: bibliography of ab initio calculations for 1989. Elsevier, Amsterdam, The Netherlands, 1990. 276 pp. LCCN ????

Helgaker:1984:SQA

[HA84] Trygve U. Helgaker and Jan Almlöf. A second-quantization approach to the analytical evaluation of response properties for perturbation-dependent basis sets. International Journal of Quantum Chemistry, 26(2):275–291, August 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Hoinkis:1983:STI

[HAB83] J. Hoinkis, R. Ahlrichs, and H.-J. Böhm. A simple treatment of intermolecular interactions: Synthesis of ab initio calculations and combination rules. International Journal of Quantum Chemistry, 23(3):821–834, March 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Halpern:1980:CLN

[Hal80] Francis R. Halpern. Classical limit of a nonseparable problem with exchange: Comparison of classical and quantum mechanics. International Journal of Quantum Chemistry, 17(1): 139–143, January 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Hall:1983:RLP

[Hal83] G. G. Hall. Reactions in the liquid phase. International Journal of Quantum Chemistry, 24(S17):425–429, March 3– 12, 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-state Theory, Collision Phenomena, and Computational Quantum Chemistry.

Hall:1987:DSW

[Hal87] G. G. Hall. Diabatic surfaces which permute into one another. International Journal of Quantum Chemistry, 31(3):383–391, March 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Hall:1988:CCS

[Hal88] George G. Hall. C. A. Coulson and the surface energy of metals: a further comment. International Journal of Quantum Chemistry, 34(3):301–304, September 1988. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Hamano:1982:VSF

[Ham82] Hidekazu Hamano. Variational solutions of a first-order perturbation equation. International Journal of Quantum Chemistry, 21(3):611–616, March 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Huai:1986:RML

[HAmGS86] Sun Huai, Tian An-min, and Yan Guo-Sen. Research in the method of large molecular calculations utilizing transferability of LMO. International Journal of Quantum Chemistry, 29(5): 1303–1324, May 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Hanson:1986:RSS

[Han86] Harold Hanson. Remarks at the Sanibel Symposium, Marineland, Florida, March 12, 1986. International Journal of Quantum Chemistry, 30(S20):13–19, March 10–15, 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Syposium on Atomic, Molecular, and Solid-State Theory, Scattering Problems, Many Body Phenomena, and Computational Quantum Chemistry.

Handler:1988:EHE

[Han88] George S. Handler. The energy of the homogeneous electron gas to second order. International Journal of Quantum Chemistry, 33(3):173–175, March 1988. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Harriman:1980:UHS

[Har80] John E. Harriman. Use of the Hiller–Sucher–Feinberg identity for spin-density calculations. International Journal of Quantum Chemistry, 17(4):689–695, April 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Harrell:1982:ETP

[Har82] Evans M. Harrell II. Estimating tunneling phenomena. International Journal of Quantum Chemistry, 21(1):199–207, January 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Harriman:1983:DMD

[Har83a] John E. Harriman. Density matrix and density functional theory with a Gaussian wavepacket basis. International Journal of Quantum Chemistry, 24(S17):501–515, March 3–12, 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-state Theory, Collision Phenomena, and Computational Quantum Chemistry.

Harris:1983:EGM

[Har83b] Frank E. Harris. Evaluation of GTO molecular integrals. International Journal of Quantum Chemistry, 23(4):1469–1478, April 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Harcourt:1987:BOT

[Har87] Richard D. Harcourt. Bohr orbit theory revisited. II. Energies for 1S, 2P, 3D, and 4F states of helium. International Journal of Quantum Chemistry, 31(3):445–453, March 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Hashimoto:1982:CSC

[Has82a] Katsufumi Hashimoto. Charge and spin correlation structures of conjugated π systems: Analysis of full CI wave functions of the PPP Hamiltonian. International Journal of Quantum Chemistry, 21(5):861–874, May 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Hashimoto:1982:RCS

[Has82b] Katsufumi Hashimoto. Relation of the charge and spin correlation functions to the structures of π electron wave functions in alternant hydrocarbons. International Journal of Quantum Chemistry, 22(4):851–862, October 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Hashimoto:1982:UHF

[Has82c] Katsufumi Hashimoto. Unrestricted-hartree–fock wave functions approximating low-lying covalent states of ring π systems. International Journal of Quantum Chemistry, 22(2): 397–414, August 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Hashimoto:1985:LDO

[Has85] Katsufumi Hashimoto. Length dependence of the optical transition energies of the exactly solvable Hubbard model for linear polyenes. International Journal of Quantum Chemistry, 28(5): 581–591, November 1985. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Hashimoto:1986:ERH

[Has86] Katsufumi Hashimoto. Exact results on homopolar singlet excitation energies of the Hubbard model for linear polyenes. International Journal of Quantum Chemistry, 30(5):633–646, November 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Hayns:1982:RTQ

[Hay82] M. R. Hayns. Relevance of the techniques of quantum chemistry in materials science and related areas. International Journal of Quantum Chemistry, 21(1):217–229, January 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Hingerty:1982:BDA

[HB82] B. E. Hingerty and S. Broyde. Base displacement in AAF and AF modified dCpdG: syn and anti guanine. International Journal of Quantum Chemistry, 22(S9):125–136, March 4–6, 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Harrison:1986:MBP

[HB86] Robert J. Harrison and Rodney J. Bartlett. A many-body perturbation theory and coupled cluster study of the water dimer. International Journal of Quantum Chemistry, 30(S20):437– 443, March 10–15, 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Syposium on Atomic, Molecular, and Solid-State Theory, Scattering Problems, Many Body Phenomena, and Computational Quantum Chemistry.

Hess:1986:TVT

[HBC86] Bernd A. Hess, Robert J. Buenker, and Praphull Chandra. Toward the variational treatment of spin-orbit and other relativistic effects for heavy atoms and molecules. International Journal of Quantum Chemistry, 29(4):737–753, April 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Habitz:1983:ECC

[HBSC83] P. Habitz, P. Bagus, P. Siegbahn, and E. Clementi. Electronic correlation contribution to the three-body potentials for water trimers. International Journal of Quantum Chemistry, 23(5): 1803–1806, May 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Hartmann:1980:CNF

[HC80] Hermann Hartmann and Kyu-Myung Chung. Classical nonlinear field theory of chemical bonding. International Journal of Quantum Chemistry, 18(6):1491–1503, December 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Hinze:1981:DCE

[HC81] Juergen Hinze and Fak. F. Chemie. Developments in the calculation of electronic wavefunctions for molecules: MCSCF, CI, and numerical SCF for molecules. International Journal of Quantum Chemistry, 20(S15):69–90, March 8–14, 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory, Collision Phenomena, and Computational Quantum Chemistry.

Hubac:1983:PTO

[HČ83] Ivan Hubač and Petr Čársky. Perturbation theory for openshell systems: Simulation of the UHF states by means of the perturbed RHF wave functions. International Journal of Quantum Chemistry, 24(2):141–148, August 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Harsanyi:1986:IVS

[HCCB86] L. Harsányi, P. Császár, A. Császár, and J. E. Boggs. Interpretation of the vibrational spectra of matrix-isolated uracil from scaled ab initio quantum mechanical force fields. International Journal of Quantum Chemistry, 29(4):799–815, April 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Hashimoto:1988:CRP

[HČP88] Katsufumi Hashimoto, Jiří Čížek, and Josef Paldus. Convergence radii of the perturbation expansions for the ground-state energies of finite Hubbard models. International Journal of Quantum Chemistry, 34(5):407–415, November 1988. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Hermann:1984:SOM

[HDFL84] M. R. Hermann, G. H. F. Diercksen, B. W. Fatyga, and P. W. Langhoff. Stieltjes orbitals for molecular photoexcitation and ionization spectra: N → V_{σ} and N → V_{π} resonance features in CO and H_{2}CO cross sections. International Journal of Quantum Chemistry, 26(S18):719–734, March 1–15, 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Atomic, Molecular and Solid-State Theory, and Computational Quantum Chemistry.

Holas:1988:IES

[HDO88] A. Holas, H. Działak, and S. Olszewski. Interaction energy in small N-particle systems and the permutation group. International Journal of Quantum Chemistry, 34(3):247–266, September 1988. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Holzwarth:1983:EBS

[HDTR83] N. A. W. Holzwarth, D. P. Divincenzo, R. C. Tatar, and S. Rabii. Energy-band structure and charge distribution for BaC_{6}. International Journal of Quantum Chemistry, 23(4): 1223–1230, April 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Helgaker:1982:SDP

[Hel82] Trygve Ulf Helgaker. Simple derivation of the potential energy gradient for an arbitrary electronic wave function. International Journal of Quantum Chemistry, 21(5):939–940, May 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Herman:1983:RAS

[Her83] Zelek S. Herman. Recent advances in simple valence-bond theory and the theory of hybrid bond orbitals. International Journal of Quantum Chemistry, 23(3):921–943, March 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Herzig:1984:CGC

[Her84] P. Herzig. Clebsch–Gordan coefficients for the group chains O ⊃ T ⊃ D_{2} and T_{d} ⊃ T ⊃ D_{2} by the method of projective representations. International Journal of Quantum Chemistry, 26(3):297–312, September 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Herman:1985:ESS

[Her85] Frank Herman. Electronic structure of semiconductor interfaces. International Journal of Quantum Chemistry, 28(S19): 547–557, March 18–23, 1985. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Heaton:1986:DRC

[HET86] M. M. Heaton and M. R. El-Talbi. The difluoroethylenes: a reconsideration of the “Cis effect”. International Journal of Quantum Chemistry, 29(1):61–72, January 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Haberland:1986:GKS

[HF86] Ralf Haberland and Lothar Fritsche. A generalized Kohn– Sham theory for low-energy electron scattering by atoms including relativistic effects. International Journal of Quantum Chemistry, 30(S20):289–297, March 10–15, 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Syposium on Atomic, Molecular, and Solid-State Theory, Scattering Problems, Many Body Phenomena, and Computational Quantum Chemistry.

Hurst:1986:IFV

[HFSB86] G. J. B. Hurst, P. W. Fowler, A. J. Stone, and A. D. Buckingham. Intermolecular forces in van der Waals dimers. International Journal of Quantum Chemistry, 29(5):1223–1239, May 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Hohlneicher:1986:TPT

[HG86] Georg Hohlneicher and Michael Gutmann. Two particle transition density and two particle bond order and their applications to electronic energy shifts caused by two particle interactions via geometric deformations. International Journal of Quantum Chemistry, 29(5):1291–1302, May 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Head-Gordon:1989:TSA

[HGHGF^{+}89] Teresa Head-Gordon, Martin Head-Gordon, Michael J. Frisch, Charles Brooks III, and John Pople. A theoretical study of alanine dipeptide and analogs. International Journal of Quantum Chemistry, 36(S16):311–322, April 1–8, 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Head-Gordon:1989:QCS

[HGPF89] Martin Head-Gordon, John A. Pople, and Michael J. Frisch. Quadratically convergent simultaneous optimization of wavefunction and geometry. International Journal of Quantum Chemistry, 36(S23):291–303, April 1–8, 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Hurley:1987:EGR

[HH87] A. C. Hurley and A. K. Head. Explicit Galois resolvents for sextic equations. International Journal of Quantum Chemistry, 31(3):345–359, March 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Hobza:1989:EPI

[HH89] P. Hobza and Z. Havlas. An empirical potential for interactions of large molecules: Application to binding of dipeptides to DNA. International Journal of Quantum Chemistry, 36(3): 287–298, September 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Hehenberger:1980:ASM

[HHL80] Michael Hehenberger, Ulf Hålenius, and Sven Larsson. Absorption spectrum of manganian andalusite: Cluster calculation by an ab initio method. International Journal of Quantum Chemistry, 18(6):1473–1477, December 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Hecht:1982:PTD

[HHT82] C. E. Hecht, D. Henderson, and A. Tani. Perturbation theory for dipolar hard spheres. International Journal of Quantum Chemistry, 22(S16):159–165, March 1–13, 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Chemistry, Theory of Condensed Matter, and Propagator Methods in the Quantum Theory of Matter.

Hiberty:1981:AMO

[Hib81] P. C. Hiberty. Analysis of molecular orbital wave functions in terms of valence bond functions for molecular fragments. I. Theory. International Journal of Quantum Chemistry, 19(2): 259–269, February 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Hickman:1983:CPI

[Hic83] Albert P. Hickman. Collisional phenomena involving Rydberg atoms. International Journal of Quantum Chemistry, 24 (S17):65–76, March 3–12, 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-state Theory, Collision Phenomena, and Computational Quantum Chemistry.

Hilal:1981:TDDa

[Hil81a] Rifaat Hilal. Thermal dissociation of 1,2-dioxethane. I. Charge density distribution along the reaction path and in different HF solutions. International Journal of Quantum Chemistry, 19(5):805–819, May 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Hilal:1981:TDDb

[Hil81b] Rifaat Hilal. Thermal dissociation of 1,2-dioxethane. II. Quantum topology of the charge distributions. International Journal of Quantum Chemistry, 19(5):821–831, May 1981. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Hilal:1981:TDDc

[Hil81c] Rifaat Hilal. Thermal dissociation of 1,2-dioxethane. III. Localized molecular-orbital study. International Journal of Quantum Chemistry, 19(5):833–845, May 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Hilal:1985:EST

[Hil85] Rifaat Hilal. On the electronic structure of transition-metal complexes II. Bonding characteristics in titanium chloride systems. International Journal of Quantum Chemistry, 28(6): 715–721, December 1985. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Hirayama:1980:TEV

[Hir80a] Satoshi Hirayama. Temperature effect on the vapor-phase exciplexes of cyano-substituted anthracenes. International Journal of Quantum Chemistry, 18(1):257–261, July 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Hiroike:1980:MVO

[Hir80b] Eiko Hiroike. Modified virtual orbitals of Hartree–Fock method in calculation of second-order perturbation energy. International Journal of Quantum Chemistry, 18(1):37–42, July 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Hirokawa:1980:IAN

[Hir80c] Shoji Hirokawa. Infrared absorption of nonpolar molecules in rattling motion. International Journal of Quantum Chemistry, 18(2):533–537, August 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Hirschfelder:1986:LMR

[Hir86] Joseph O. Hirschfelder. Laser and magnetic resonance propagators. International Journal of Quantum Chemistry, 29(5): 1139–1146, May 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Hosur:1980:TDF

[HJ80] Ramakrishna V. Hosur and Anjan S. Joshipura. Theoretical description of fluidity in biological membranes: Rotational motion in lipid hydrocarbon chains. International Journal of Quantum Chemistry, 17(5):983–993, May 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Haque:1986:OSC

[HK86] Azizul Haque and Uzi Kaldor. Open-shell coupled-cluster method: Variational and nonvariational calculation of ionization potentials. International Journal of Quantum Chemistry, 29(3):425–433, March 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Hori:1987:PMC

[HKFW87] Kenzi Hori, Joseph N. Kushick, Alan Factor, and Harel Weinstein. Parameters and mechanisms of calcium binding to peptides and proteins. International Journal of Quantum Chemistry, 32(S14):341–345, March 14, 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Hosur:1981:PNA

[HKG81] Ramakrishna V. Hosur, N. Vasanth Kumar, and Girjesh Govil. Protein–nucleic acid interactions: Investigations on the peptide backbone interaction with polynucleotides. International Journal of Quantum Chemistry, 20(1):23–32, July 1981. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Hariharan:1982:IMV

[HKK^{+}82] P. C. Hariharan, Walter S. Koski, Joyce J. Kaufman, Richard S. Miller, and Alfred H. Lowrey. Ab Initio MODPOT/ VRDDO/MERGE calculations on energetic compounds. II. Nitroexplosives: RDX and α-, β- and δ-HMX. International Journal of Quantum Chemistry, 22(S16):363–375, March 1– 13, 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Chemistry, Theory of Condensed Matter, and Propagator Methods in the Quantum Theory of Matter.

Hariharan:1983:IMS

[HKKM83] P. C. Hariharan, Walter S. Koski, Joyce J. Kaufman, and Richard S. Miller. Ab initio MODPOT /VRDDO /MERGE calculations on energetic compounds. III. Nitroexplosives: Polyaminopolynitrobenzenes (including DATB, TATB, and tetryl). International Journal of Quantum Chemistry, 23(4): 1493–1504, April 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Hariharan:1985:IMS

[HKLM85] P. C. Hariharan, Joyce J. Kaufman, Alfred H. Lowrey, and Richard S. Miller. Ab initio MODPOT /VRDDO /MERGE calculations on energetic compounds. IV. Nitrocubanes: Mononitro to octanitro quantum chemical calculations and electrostatic molecular potential contour maps. International Journal of Quantum Chemistry, 28(1):39–59, July 1985. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Hariharan:1981:EMP

[HKP81] P. C. Hariharan, Joyce J. Kaufman, and Carlo Petrongolo. Electrostatic molecular potential contour maps generated from ab initio MODPOT/VRDDO/MERGE wave functions of carcinogenic 3-methylcholanthrene and its metabolites. International Journal of Quantum Chemistry, 20(5): 1083–1095, November 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Heidrich:1985:ICH

[HKV85] D. Heidrich, H.-J. Köhler, and D. Volkmann. Ab initio calculations on H transfer in the HF trimer. International Journal of Quantum Chemistry, 27(6):781–786, June 1985. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Hopkinson:1980:TSS

[HL80] A. C. Hopkinson and M. H. Lien. Theoretical study of substituent effects on the acidity of the methyl group: Structure of anions CH_{2}X^{−}. International Journal of Quantum Chemistry, 18(6):1371–1391, December 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Heffernan:1981:ESI

[HL81] Daniel M. Heffernan and Richard L. Liboff. Exact solutions for interacting finite potential wells. International Journal of Quantum Chemistry, 19(5):745–753, May 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Hermann:1983:LPL

[HL83] M. R. Hermann and P. W. Langhoff. Löwdin partitioning in a Lanczos basis: Applications to scattering states. International Journal of Quantum Chemistry, 23(1):135–146, January 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Holme:1988:AHE

[HL88] T. A. Holme and R. D. Levine. An algebraic Hamiltonian for electronic and nuclear degrees of freedom based on the vector model. International Journal of Quantum Chemistry, 34(5): 457–470, November 1988. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Hariharan:1982:QCC

[HLKK82] P. C. Hariharan, Victor Lewchenko, Walter S. Koski, and Joyce J. Kaufman. Quantum chemical calculations for understanding and predicting toxicity. I. Ab initio MODPOT/ VRDDO calculations and electrostatic molecular potential contour maps of organophosphorus anticholinesterases. International Journal of Quantum Chemistry, 22(S9):259–274, March 4–6, 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Hotokka:1983:REH

[HLPV83] M. Hotokka, L. Laaksonen, P. Pyykkö, and A. Viste. Relativistic extended Hückel studies of cluster models of solid HgTe, CdTe, and PbTe: Densities of states and nuclear spin– spin coupling. International Journal of Quantum Chemistry, 23(4):1685, April 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Herman:1980:CGS

[HLR80] Zelek S. Herman, Gilda H. Loew, and Marie-Madeleine Rohmer. Calculated ground-state properties and optical spectrum of model carbonylheme complexes. International Journal of Quantum Chemistry, 18(S7):137–153, March 5–8, 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Heeger:1981:SMP

[HM81] A. J. Heeger and A. G. Macdiarmid. Semiconducting and metallic polymers: From quantum chemistry to new technology. International Journal of Quantum Chemistry, 20(S15): 243–249, March 8–14, 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory, Collision Phenomena, and Computational Quantum Chemistry.

Hilal:1984:ESP

[HM84] Rifaat Hilal and Hussein Moustafa. Electronic structure of the peptide linkage. I. Equilibrium geometry and electronic properties of formhydroxamic acid. International Journal of Quantum Chemistry, 26(2):183–196, August 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Heaton:1985:IQM

[HM85a] M. M. Heaton and D. Mills. Ab initio quantum-mechanical calculations on trifluoromethylamine. International Journal of Quantum Chemistry, 28(2):163–180, August 1985. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Hinchliffe:1985:ME

[HM85b] Alan Hinchliffe and Robert W. Munn. Molecular electromagnetism. John Wiley, New York, NY, USA, 1985. ISBN 0-471-10292-X (hardcover), 0-471-90721-9 (paperback). x + 252 pp. LCCN QC760 .H55 1985. UK £23.00, UK£12.50.

Hohlneicher:1986:FSE

[HM86] Georg Hohlneicher and Bernd Marquardt. Final state effects in the photoionization process. International Journal of Quantum Chemistry, 29(5):1437–1455, May 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Hu:1989:SSB

[HM89a] Chengzheng Hu and Hendrik J. Monkhorst. Symmetric and symmetry-broken Hartree–Fock solutions of one-dimensional models. International Journal of Quantum Chemistry, 36(5): 659–667, November 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Hunter:1989:EDP

[HM89b] G. Hunter and N. H. March. Explicit density–potential relation for 4-electron atoms in 2-orbital S states. International Journal of Quantum Chemistry, 35(5):649–663, May 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Hochmann:1982:IEC

[HMF82] P. Hochmann, S. P. Mcglynn, and G. L. Findley. Ionization energy correlations for alkyl-substituted chromophores. International Journal of Quantum Chemistry, 22(S9):173–179, March 4–6, 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Ho:1981:IHC

[Ho81] Y. K. Ho. Improved Hylleraas calculations for ground state energies of lithium ISO–electronic sequence. International Journal of Quantum Chemistry, 20(5):1077–1082, November 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Hiberty:1985:VBDa

[HO85a] P. C. Hiberty and G. Ohanessian. The valence bond description of conjugated molecules i. SCF level. International Journal of Quantum Chemistry, 27(3):245–257, March 1985.

CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Hiberty:1985:VBDb

[HO85b] P. C. Hiberty and G. Ohanessian. The valence bond description of conjugated molecules. II. A very simple method to approximate the structural weights of a fully correlated valence bond wavefunction. International Journal of Quantum Chemistry, 27(3):259–272, March 1985. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Hosokawa:1980:TMP

[Hos80] Masahiko Hosokawa. Theoretical model for the phase transition of a membrane as a multicomponent system. International Journal of Quantum Chemistry, 18(6):1517–1522, December 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Hall:1987:SPFa

[HOT87a] Lowell H. Hall, Bradley J. Orchard, and Sukant K. Tripathy. The structure and properties of flavins: Molecular orbital study based on totally optimized geometries. I. Molecular geometry investigations. International Journal of Quantum Chemistry, 31(2):195–216, February 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Hall:1987:SPFb

[HOT87b] Lowell H. Hall, Bradley J. Orchard, and Sukant K. Tripathy. The structure and properties of flavins: Molecular orbital study based on totally optimized geometries. II. Molecular orbital structure and electron distribution. International Journal of Quantum Chemistry, 31(2):217–242, February 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Hosur:1981:PBI

[HP81] Ramakrishna V. Hosur and A. Pohorille. Protein backbone interaction with nucleic acid base pairs: a theoretical investigation. International Journal of Quantum Chemistry, 20(1): 33–40, July 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Hariharan:1981:MCN

[HPK81] P. C. Hariharan, Herbert E. Popkie, and Joyce J. Kaufman. Molecular calculations with the nonempirical ab initio MODPOT, VRDDO, and MODPOT/VRDDO procedures. XII. Carcinogenic 3-methylcholanthrene and its metabolites using a MERGE technique. International Journal of Quantum Chemistry, 20(3):645–653, September 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Hofmann:1981:CHS

[HPKW81] Hans-Jörg Hofmann, Gustav Peinel, Christine Krebs, and Cornelius Weiss. Construction of hydration shells for the tautomeric forms of heterocycles based on electrostatic potentials. International Journal of Quantum Chemistry, 20(4):785–792, October 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Hron:1983:SAG

[HR83] M. Hron and M. Razavy. Some applications of the Gel’fand– Levitan inverse method in atomic and molecular physics. International Journal of Quantum Chemistry, 24(1):97–111, July 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Hale:1984:IET

[HR84] Paul D. Hale and Mark A. Ratner. Intramolecular electron transfer, with bridge assistance and without, in molecules and models. International Journal of Quantum Chemistry, 26(S18):195–211, March 1–15, 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Atomic, Molecular and Solid-State Theory, and Computational Quantum Chemistry.

Hodgkin:1987:MSB

[HR87] Edward E. Hodgkin and W. Graham Richards. Molecular similarity based on electrostatic potential and electric field. International Journal of Quantum Chemistry, 32(S14):105–110, March 14, 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Hirata:1988:VBM

[HRL88] Fumio Hirata, Paul Redfern, and Ronald M. Levy. Viewing the born model for ion hydration through a microscope. International Journal of Quantum Chemistry, 34(S15):179– 190, March 12–19, 1988. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Syposium on Quantum Biology and Quantum Pharmacology.

Hartmann:1980:SOC

[HS80a] Hermann Hartmann and Dieter Schuch. Spin–orbit coupling for the motion of a particle in a ring-shaped potential. International Journal of Quantum Chemistry, 18(1):125–141, July 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Ho:1980:SPC

[HS80b] Fat D. Ho and Frank C. Sanders. Study of pair correlation in three-electron atoms. International Journal of Quantum Chemistry, 17(4):619–629, April 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Hall:1984:FED

[HS84a] G. G. Hall and C. M. Smith. Fitting electron densities of molecules. International Journal of Quantum Chemistry, 25 (5):881–890, May 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Hameka:1984:FDD

[HS84b] Hendrik F. Hameka and Erik Nørby Svendsen. Frequency dependence of dc electric-field-induced second-harmonic generation. International Journal of Quantum Chemistry, 26(S18): 525–536, March 1–15, 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Atomic, Molecular and Solid-State Theory, and Computational Quantum Chemistry.

Hubac:1988:QCV

[HS88] Ivan Hubač and Michal Svrček. The quasiparticle concept in vibrational–electronic problems in molecules. I. Partitioning of the vibrational–electronic Hamiltonian. International Journal of Quantum Chemistry, 33(5):403–443, May 1988. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Head:1989:GTE

[HS89] John D. Head and Susil J. Silva. A guide to two-electron integral approximations through semiorthogonal basis sets. International Journal of Quantum Chemistry, 36(S23):231–244, April 1–8, 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Holleboom:1988:NEO

[HSB88] L. J. Holleboom, J. G. Snijders, and E. J. Baerends. Natural energy orbitals and the one-particle Green’s function. International Journal of Quantum Chemistry, 34(3):289–300, September 1988. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Hess:1986:CCP

[HSR86] B. Andes Hess, Jr., L. J. Schaad, and D. N. Reinhoudt. [2 + 2] Cycloadditions. A concerted pathway for acetylene plus ethylene. International Journal of Quantum Chemistry, 29 (3):345–350, March 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Harmalkar:1983:MSP

[HSSW83] Anjali Harmalkar, Alfredo M. Simas, Vedene H. Smith, Jr., and William M. Westgate. Momentum space properties of atoms. International Journal of Quantum Chemistry, 23(3): 811–820, March 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Hunter:1982:VMA

[HT82] Geoffrey Hunter and Chin Chiu Tai. Variational marginal amplitudes. International Journal of Quantum Chemistry, 21 (6):1041–1050, June 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Hunter:1985:EOE

[HT85] Geoffrey Hunter and Chin Chiu Tai. An exact one-electron model of the helium atom ground state. International Journal of Quantum Chemistry, 28(S19):173–180, March 18–23, 1985. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Harris:1980:FMM

[HTD80] Frank E. Harris, A. Trautwein, and J. Delhalle. The FAKE method of molecular orbital calculation. International Journal of Quantum Chemistry, 18(S14):355–361, March 10–20, 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory, Collision Phenomena, Quantum Statistics, and Computational Methods.

Hubac:1980:CEP

[Hub80] Ivan Hubač. Comments on the exclusion-principle-violating diagrams cancellation in correlations energy calculation by many-body Rayleigh–Schrödinger perturbation theory. International Journal of Quantum Chemistry, 17(2):195–209, February 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Huber:1985:EVR

[Hub85] Daniel Huber. Energies of vibrating and rotating molecules by ladder operators. International Journal of Quantum Chemistry, 28(2):245–267, August 1985. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Hunter:1980:NWF

[Hun80] Geoffrey Hunter. Nodeless wave function quantum theory. International Journal of Quantum Chemistry, 17(1):133–137, January 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Hunter:1981:NWF

[Hun81] Geoffrey Hunter. Nodeless wave functions and spiky potentials. International Journal of Quantum Chemistry, 19(5): 755–761, May 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Hunter:1986:EOE

[Hun86] Geoffrey Hunter. The exact one-electron model of molecular structure. International Journal of Quantum Chemistry, 29 (2):197–204, February 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Hurley:1982:ACB

[Hur82] A. C. Hurley. Analysis of the covalent bond: One center floating functions for the hydrogen molecule ion. International Journal of Quantum Chemistry, 22(2):241–251, August 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Hegarty:1983:IEA

[HV83] Dermot Hegarty and G. Van Der Velde. Integral evaluation algorithms and their implementation. International Journal of Quantum Chemistry, 23(4):1135–1153, April 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Head:1988:SOP

[HWZ88] John D. Head, Brain Weiner, and Michael C. Zerner. A survey of optimization procedures for stable structures and transition states. International Journal of Quantum Chemistry, 33(3): 177–186, March 1988. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Hong-Yi:1989:NUT

[HY89] Fan Hong-Yi. New unitary transformation for the three coupled oscillators. International Journal of Quantum Chemistry, 35(5):585–592, May 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Hui-Yun:1989:VMV

[HYS89] Pan Hui-Yun and Zhao Zu Sen. On the validity of the mass– velocity operator in quantum chemistry. International Journal of Quantum Chemistry, 36(1):15–18, July 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Hong-Yi:1989:NOF

[HYZ89] Fan Hong-Yi and H. R. Zaidi. Normally ordered Franck– Condon operator: a new approach for the calculation of Frank–Condon factors. International Journal of Quantum Chemistry, 35(2):277–282, February 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Hobza:1983:VWM

[HZ83] P. Hobza and R. Zahradník. Van der Waals molecules: Quantum chemistry, physical properties, and reactivity. International Journal of Quantum Chemistry, 23(1):325–338, January 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Havlas:1984:TSR

[HZ84] Zdeněk Havlas and Rudolf Zahradník. Theoretical studies of reaction mechanism in chemistry. International Journal of Quantum Chemistry, 26(5):607–619, November 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Hidalgo:1989:PVS

[HZBR89] A. Hidalgo, J. Zúñiga, A. Bastida, and A. Requena. Perturbative virtual SCF CI treatment for energy levels of coupled oscillator systems. International Journal of Quantum Chemistry, 36(1):49–60, July 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Hobza:1984:OHE

[HZL84] Pavel Hobza, Rudolf Zahradník, and Janos Ladik. Origin of high efficiency and specificity of biochemical reactions. International Journal of Quantum Chemistry, 26(5):857–872, November 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Imamura:1987:PAA

[IA87] Akira Imamura and Yuriko Aoki. Perturbational approach to aperiodicity of polymer systems. International Journal of Quantum Chemistry, 32(S21):137–151, March 12, 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Chemistry, Solid-State Theory, and Computational Methods.

Ishikawa:1989:NOC

[IAAA89] Yasuyuki Ishikawa, I. L. Aponte-Avellanet, and S. A. Alexander. Numerical orbital calculations using nonuniform grids. International Journal of Quantum Chemistry, 36(S23):209–215, April 1–8, 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Ishikawa:1985:GBD

[IBB85] Yasuyuki Ishikawa, Reinaldo Baretty, and Robert C. Binning, Jr. Gaussian basis for the Dirac-Fock discrete basis expansion calculations. International Journal of Quantum Chemistry, 28 (S19):285–295, March 18–23, 1985. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Ishikawa:1987:GBD

[IBB87] Yasuyuki Ishikawa, Reinaldo Baretty, and Robert C. Binning, Jr. Gaussian basis for the Dirac–Fock discrete basis expansion calculation. International Journal of Quantum Chemistry, 31 (2):329, February 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Inokuti:1987:TED

[IDK87] Mitio Inokuti, Michael A. Dillon, and Mineo Kimura. Theory of electron degradation and yields of initial molecular species produced by ionizing radiation. International Journal of Quantum Chemistry, 32(S21):251–266, March 12, 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Chemistry, Solid-State Theory, and Computational Methods.

Igawa:1985:MSP

[Iga85] Akira Igawa. An MC-SCF procedure using an orthogonal transformation represented by a direct sum of two-by-two rotations. International Journal of Quantum Chemistry, 28(2): 203–211, August 1985. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Ikuta:1980:DEP

[IHI80] Shigeru Ikuta, Shuichi Hashimoto, and Masashi Imamura. β^{−}- Decay effect on the proton movement in the pyridine–pyrrole hydrogen-bonded system. International Journal of Quantum Chemistry, 18(2):515–519, August 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Isihara:1982:CIP

[II82] A. Isihara and L. Ioriatti. Correlation-induced phase transitions in two-dimensional electron systems. International Journal of Quantum Chemistry, 22(S16):149–158, March 1–13, 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Chemistry, Theory of Condensed Matter, and Propagator Methods in the Quantum Theory of Matter.

Iachello:1983:AAM

[IL83] F. Iachello and R. D. Levine. Algebraic approach to molecular rotation–vibration spectra. International Journal of Quantum Chemistry, 23(4):1679–1681, April 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Iadonisi:1984:EHP

[IL84] Giuseppe Iadonisi and Andrea C. Levi. Electron-hole pair excitation in atom-surface scattering. International Journal of Quantum Chemistry, 26(5):823–830, November 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Ibantez-Mengual:1980:QMT

[IMTG80] J. A. Ibáñtez-Mengual and A. F. Tejerina-García. Quantum model for transport through membranes. International Journal of Quantum Chemistry, 17(6):1191–1200, June 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Ihaya:1984:ICO

[INFU84] Y. J. I’haya, S. Narita, Y. Fujita, and H. Ujino. Ab initio crystal orbital calculations on (CH)_{n} and (HF)_{n} with extended basis sets. International Journal of Quantum Chemistry, 26(S18):153–159, March 1–15, 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Atomic, Molecular and Solid-State Theory, and Computational Quantum Chemistry.

Ingolfsson:1980:NAS

[Ing80] K. Ingólfsson. Notes on asymptotic series by H^{0}-strictly singular perturbations. International Journal of Quantum Chemistry, 17(1):99–106, January 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Iwaki:1988:ICR

[INO88] Hiroaki Iwaki, Takeshi Noro, and Kimio Ohno. Ab-initio calculation of resonance states of H_{2}O. International Journal of Quantum Chemistry, 33(6):567–576, June 1988. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Ishida:1989:IBS

[IO89] Toshimasa Ishida and Koichi Ohno. The influence of basis sets on wave function tails. International Journal of Quantum Chemistry, 35(2):257–266, February 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Ischenko:1981:NFD

[IPV81] S. Ya. Ischenko and V. B. Pavlov-Verevkin. New finitedifference method for solving the pair radial equations of the many electron perturbation theory. International Journal of Quantum Chemistry, 19(2):361–372, February 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Ishikawa:1987:UEN

[IQ87] Yasuyuki Ishikawa and H. M. Quiney. On the use of an extended nucleus in Dirac–Fock Gaussian basis set calculations. International Journal of Quantum Chemistry, 32(S21):523– 532, March 12, 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Chemistry, Solid-State Theory, and Computational Methods.

Ishikawa:1987:SID

[IRA87] Yasuyuki Ishikawa, Wilfredo Rodriguez, and S. A. Alexander. Solution of the integral Dirac equation in momentum space. International Journal of Quantum Chemistry, 32(S21): 417–423, March 12, 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Chemistry, Solid-State Theory, and Computational Methods.

Iadonisi:1981:PAC

[IRP81] G. Iadonisi, V. Marigliano Ramaglia, and B. Preziosi. Physical approach to chemisorption through Green’s function method. International Journal of Quantum Chemistry, 19(6):1043– 1056, June 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Itagaki:1980:MBC

[IS80] T. Itagaki and A. Saika. Many-body calculation of the electric field gradient in the hydrogen molecule. International Journal of Quantum Chemistry, 18(6):1505–1510, December 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Iwai:1983:ECE

[IS83] Masahiro Iwai and A. Saika. Electron correlation effects on magnetic properties of BH. International Journal of Quantum Chemistry, 24(6):623–626, December 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Iijima:1985:NCM

[IS85] Naoto Iijima and A. Saika. A note on the convergence of multiconfigurational many-body perturbation theory. International Journal of Quantum Chemistry, 27(4):481–493, April 1985. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Ishikawa:1988:UGT

[IS88] Yasuyuki Ishikawa and Hideo Sekino. On the use of Gaussiantype functions in Dirac–Fock basis set expansion calculations. International Journal of Quantum Chemistry, 34(S22):457– 463, March 12–19, 1988. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Syposium on Quantum Chemistry, Solid-State Theory, and Computational Methods.

Ishida:1983:CSI

[Ish83] Kazuhiro Ishida. Calculation in k space of integrals arising in the theory of van der Waals forces. II. Formulas for arbitrary atoms. International Journal of Quantum Chemistry, 24(6): 633–649, December 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Ishida:1984:MIA

[Ish84a] Kazuhiro Ishida. Molecular integrals arising in the linear response theory of van der Waals forces. International Journal of Quantum Chemistry, 26(4):535–559, October 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Ishikawa:1984:BSE

[Ish84b] Yasuyuki Ishikawa. Basis-set expansion calculations with the Dirac Hamiltonian. International Journal of Quantum Chemistry, 26(S18):375–383, March 1–15, 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Atomic, Molecular and Solid-State Theory, and Computational Quantum Chemistry.

Ishida:1985:MIA

[Ish85] Kazuhiro Ishida. Molecular integrals arising in the use of the Hiller–Sucher–Feinberg–Harriman identity for describing the Fermi-contact interaction. International Journal of Quantum Chemistry, 28(3):349–373, September 1985. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Ishida:1986:SCD

[Ish86] Kazuhiro Ishida. Spin and charge densities from Hiller– Sucher–Feinberg identities. Test calculations for He, B, C, N, O, and F atoms and H_{2}O, N_{2}, and CH_{3} molecules. International Journal of Quantum Chemistry, 30(4):543–562, October 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Ishida:1988:E

[Ish88a] Kazuhiro Ishida. Erratum. International Journal of Quantum Chemistry, 34(1):89–90, July 1988. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Ishida:1988:SCD

[Ish88b] Kazuhiro Ishida. Spin and charge densities from Hiller– Sucher–Feinberg identities. Test calculations for He, B, C, N, O, and F atoms and H_{2}O, N_{2}, and CH_{3} molecules. International Journal of Quantum Chemistry, 34(2):195–197, August 1988. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Ishida:1989:CDU

[Ish89] Kazuhiro Ishida. Charge density by the use of the Hiller– Sucher–Feinberg identity. Perturbative corrections and configuration interaction effects. International Journal of Quantum Chemistry, 36(3):213–223, September 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Iyakutti:1989:ESH

[IST89] K. Iyakutti, S. Sankar, and M. Thiagarajan. Electronic structure of helium-trapped vacancy in aluminum using multiple scattering X_{α} method. International Journal of Quantum Chemistry, 35(5):679–685, May 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

IlIcheva:1980:CSD

[ITKD80] I. A. Il’Icheva, V. G. Tumanyan, A. E. Kister, and V. G. Dashevsky. Conformational studies of double-helical polynucleotides by the method of pair-wise potential functions. International Journal of Quantum Chemistry, 17(2):321–326, February 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Julg:1982:TSA

[JA82] André Julg and Alain Allouche. Theoretical study of the adsorption of silver atoms on the (111) face of silicon. International Journal of Quantum Chemistry, 22(4):739–752, October 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Jackiw:1980:TDV

[Jac80] R. Jackiw. Time-dependent variational principle in quantum field theory. International Journal of Quantum Chemistry, 17 (1):41–46, January 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Jacchieri:1989:CBT

[Jac89] S. G. Jacchieri. Cooperative binding of transition states in enzymatic catalysis. International Journal of Quantum Chemistry, 36(S16):301–309, April 1–8, 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Jarlskog:1989:PAM

[Jar89] C. Jarlskog. On particle–antiparticle mirror symmetry. International Journal of Quantum Chemistry, 35(6):769–778, June 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Jasselette:1980:IPD

[Jas80] P. Jasselette. Interventions in the panel discussion on measurement process and interpretation of quantum mechanics. International Journal of Quantum Chemistry, 17(1):83–84, January 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Jones:1987:MLF

[JBW87] H. W. Jones, B. Bussery, and C. A. Weatherford. An “E matrix” for the Löwdin α function, expanded in a Taylor series: an analytic treatment of molecular charge density near the origin. International Journal of Quantum Chemistry, 32 (S21):693–698, March 12, 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Chemistry, Solid-State Theory, and Computational Methods.

Jasien:1983:SDS

[JD83] Paul G. Jasien and Clifford E. Dykstra. The significance of double substitutions in well-correlated electronic wave functions. International Journal of Quantum Chemistry, 24(S17): 289–296, March 3–12, 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-state Theory, Collision Phenomena, and Computational Quantum Chemistry.

Jasien:1985:FDC

[JD85] Paul G. Jasien and Clifford E. Dykstra. First derivatives of correlated wave functions by a matrix-oriented method: Preliminary application to molecular dipole and quadrupole moments. International Journal of Quantum Chemistry, 28(3): 411–417, September 1985. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Jones:1988:AEE

[JEW88] H. W. Jones, B. Etemadi, and C. A. Weatherford. Analytical evaluation of the electrostatic potential for diatomic molecules. International Journal of Quantum Chemistry, 34 (S22):497–502, March 12–19, 1988. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Syposium on Quantum Chemistry, Solid-State Theory, and Computational Methods.

Jain:1981:BLD

[JG81] D. V. S. Jain and Vijay Gombar. Binding of leucomycin and its derivatives to E. coli ribosomes. International Journal of Quantum Chemistry, 20(2):419–427, August 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Jayasuriya:1988:CAO

[JI88] Keerthi Jayasuriya and Sury Iyer. A computational analysis of ortho-lithiation in aromatic systems. International Journal of Quantum Chemistry, 34(3):199–206, September 1988. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Jug:1987:DPS

[JIS87] Karl Jug, Rüdiger Iffert, and Joachim Schulz. Development and parametrization of SINDO1 for second-row elements. International Journal of Quantum Chemistry, 32(2):265–277, August 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Julg:1983:TNI

[JJ83] A. Julg and P. Julg. Towards a new interpretation of quantum mechanics? International Journal of Quantum Chemistry, 23 (2):369–377, February 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Jeziorska:1987:DCH

[JJČ87] Małgorzata Jeziorska, Bogumił Jeziorski, and Jiří Čížek. Direct calculation of the Hartree–Fock interaction energy via exchange–perturbation expansion. The He⋯He interaction. International Journal of Quantum Chemistry, 32(2):149–164, August 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Jiang:1988:SPG

[JJZ^{+}88a] X. Q. Jiang, B. Y. Jin, Z. H. Zeng, C. C. Sun, and A. J. Coleman. The sequential product of geminals and the ground states of some atoms. International Journal of Quantum Chemistry, 34(1):9–14, July 1988. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Jin:1988:SPG

[JJZ^{+}88b] B. Y. Jin, X. Q. Jiang, Z. H. Zeng, C. C. Sun, and A. J. Coleman. A sequential product of geminals and its applications. International Journal of Quantum Chemistry, 34(1): 15–23, July 1988. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Jaworski:1985:WII

[JKL85] Andrzej Jaworski, Józef S. Kwiatkowski, and Bogdan Lesyng. Why isoguanine and isocytosine are not the components of the genetic code. International Journal of Quantum Chemistry, 28(S12):209–216, March 14–16, 1985. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Jin:1984:QCS

[JL84] Suqian Jin and Ruozhuang Liu. Quantum-chemical study of electrophilic addition reaction of iodine to ethylene. International Journal of Quantum Chemistry, 25(4):699–705, April 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Jaiswal:1985:EIO

[JL85] P. S. Jaiswal and S. Lahiry. On the evaluation of the integral < p_{u}|∂/∂v|s > occurring in covalency reduction factor. International Journal of Quantum Chemistry, 28(6):723–729, December 1985. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Jhon:1987:SRC

[JL87] Mu Shik Jhon and Per-Olov Löwdin. Some remarks on certain magnetic properties of water in the study of cancer. International Journal of Quantum Chemistry, 32(S14):9–14, March 14, 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Jeffrey:1981:HBG

[JM81] G. A. Jeffrey and H. Maluszynska. Hydrogen-bonding geometry and patterns in carbohydrates and amino acids. International Journal of Quantum Chemistry, 20(S8):231–239, March 5–7, 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Jug:1983:SVD

[JM83] Karl Jug and P. C. Mishra. Single versus double insertion in a simple intermolecular carbene reaction. International Journal of Quantum Chemistry, 23(3):887–890, March 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Jones:1988:GOU

[JMD88] R. S. Jones, J. W. Mintmire, and B. I. Dunlap. Geometry optimization using local-density functional methods: Numerical aspects. International Journal of Quantum Chemistry, 34(S22):77–84, March 12–19, 1988. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Syposium on Quantum Chemistry, Solid-State Theory, and Computational Methods.

Jankowski:1985:ECE

[JMS^{+}85] K. Jankowski, P. Malinowski, A. Sokołowski, I. Lindgren, and A.-M. Mårtensson-Pendrill. Electron correlation effects in the 4 f^{14} shell. International Journal of Quantum Chemistry, 27 (6):665–675, June 1985. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Jankowski:1988:IEC

[JMS88] K. Jankowski, P. Malinowski, and A. Sokołowski. Intershell nl4f electron correlation effects. International Journal of Quantum Chemistry, 34(3):279–288, September 1988. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Janssen:1988:ICS

[JN88] G. J. M. Janssen and W. C. Nieuwpoort. On the ab-initio calculation of d-d spectra in transition metal compounds: The importance of relaxed charge transfer states. International Journal of Quantum Chemistry, 34(S22):679–696, March 12– 19, 1988. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Syposium on Quantum Chemistry, Solid-State Theory, and Computational Methods.

Jones:1980:CGF

[Jon80] Herbert W. Jones. Computer-generated formulas for overlap integrals of Slater-type orbitals. International Journal of Quantum Chemistry, 18(3):709–713, September 1980. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Jones:1981:CGFb

[Jon81a] Herbert W. Jones. Computer-generated formulas for hybrid integrals over Slater-type orbitals. International Journal of Quantum Chemistry, 20(S15):287–291, March 8–14, 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory, Collision Phenomena, and Computational Quantum Chemistry.

Jones:1981:CGFa

[Jon81b] Herbert W. Jones. Computer-generated formulas for twocenter Coulomb integrals over Slater-type orbitals. International Journal of Quantum Chemistry, 20(6):1217–1224, December 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Jones:1981:EFO

[Jon81c] Herbert W. Jones. Exact formulas for overlap integrals of Slater-type orbitals with equal screening constants. International Journal of Quantum Chemistry, 19(4):567–574, April 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Jones:1982:ETC

[Jon82] Herbert W. Jones. Evaluation of two-center overlap and Coulomb integrals derived from Slater-type orbitals. International Journal of Quantum Chemistry, 21(6):1079–1089, June 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Jones:1983:CGF

[Jon83] Herbert W. Jones. Computer-generated formulas for some three-center molecular integrals over Slater-type orbitals. International Journal of Quantum Chemistry, 23(3):953–957, March 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Jones:1984:FTC

[Jon84] Herbert W. Jones. Formulas for three-center exchange integrals by computer implementation of the Löwdin α-function method. International Journal of Quantum Chemistry, 26 (S18):61–66, March 1–15, 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Atomic, Molecular and Solid-State Theory, and Computational Quantum Chemistry.

Jones:1985:CGF

[Jon85] Herbert W. Jones. Computer-generated formulas for exchange integrals over Slater-type orbitals. International Journal of Quantum Chemistry, 28(S19):157–163, March 18–23, 1985. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Jones:1986:CGF

[Jon86] Herbert W. Jones. Computer-generated formulas for fourcenter integrals over Slater-type orbitals. International Journal of Quantum Chemistry, 29(2):177–183, February 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Jordan:1981:RDE

[Jor81] Kenneth D. Jordan. Recent developments in electron-molecule scattering. International Journal of Quantum Chemistry, 20 (S15):331–340, March 8–14, 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory, Collision Phenomena, and Computational Quantum Chemistry.

Jortner:1986:A

[Jor86] Joshua Jortner. Announcement. International Journal of Quantum Chemistry, 29(6):1849, June 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Jortner:1989:OPP

[Jor89] Joshua Jortner. Organizer’s preface to the proceedings of the sixth international congress on quantum chemistry. International Journal of Quantum Chemistry, 36(3):187–188, September 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Jankowski:1980:ACP

[JP80] K. Jankowski and J. Paldus. Applicability of coupled-pair theories to quasidegenerate electronic states: a model study. International Journal of Quantum Chemistry, 18(5):1243–1269, November 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Jaworski:1987:SCD

[JPAB87] Andrzej Jaworski, Willis B. Person, Ludwik Adamowicz, and Rodney J. Bartlett. Study of the conformation of the dilithioacetylene molecule. International Journal of Quantum Chemistry, 32(S21):613–621, March 12, 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Chemistry, Solid-State Theory, and Computational Methods.

Jorg:1985:ALX

[JR85] H. Jörg and N. Rösch. Application of the LCGTO–Xα method to chemisorption studies: The cluster Ni_{2} CO. International Journal of Quantum Chemistry, 28(S19):747, March 18–23, 1985. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Jonsson:1984:QMS

[JRK84] B. Jönsson, S. Romano, and G. Karlström. Quantummechanical and statistical mechanical studies of the torsional barrier of H_{2}O_{2} in aqueous solution. International Journal of Quantum Chemistry, 25(3):503–514, March 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Jorgensen:1986:GDE

[JS86] Poul Jørgensen and Jack Simons, editors. Geometrical derivatives of energy surfaces and molecular properties, volume 166 of NATO ASI series. Series C, Mathematical and physical sciences. D. Reidel, Dordrecht, Boston, Lancaster, Tokyo, 1986. ISBN 90-277-2176-9. LCCN QD461.5 .N35 1985.

Jasien:1988:GPP

[JS88] Paul G. Jasien and Ron Shepard. A general polyatomic potential energy surface fitting method. International Journal of Quantum Chemistry, 34(S22):183–198, March 12–19, 1988. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Syposium on Quantum Chemistry, Solid-State Theory, and Computational Methods.

Jones:1989:MMO

[JSDK89] L. F. Jones, D. D. Shillady, A. Duong, and L. B. Kier. A multiconfigurational molecular orbital study of N-methyl oxatriazole. International Journal of Quantum Chemistry, 36(S23): 465–471, April 1–8, 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Jolly:1980:EPM

[JSS80] Pratibha Jolly, Kamalni Sane, and K. V. Sane. Effective potential model for calculating the energy shift in the ground state of two-electron atoms. International Journal of Quantum Chemistry, 18(5):1223–1228, November 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Jolly:1980:VFP

[JSSS80] P. Jolly, R. P. Saxena, P. K. Srivastava, and K. V. Sane. Variational functions of the Padé type—application to the ground state of the helium atom. International Journal of Quantum Chemistry, 18(3):689–694, September 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Jolly:1982:SRU

[JSSS82] P. Jolly, K. V. Sane, R. P. Saxena, and P. K. Srivastava. Some remarks on the use of rational functions as variational functions. International Journal of Quantum Chemistry, 21(6): 1051–1055, June 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Jorgensen:1983:MHF

[JSYO83] Poul Jørgensen, Peter Swanstrøm, Danny L. Yeager, and Jeppe Olsen. Multiconfigurational Hartree–Fock response functions. International Journal of Quantum Chemistry, 23 (3):959–971, March 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Jones:1985:AOP

[JT85] R. S. Jones and S. B. Trickey. Avoiding orthogonality problems in the application of the alternant molecular orbital method to solids. International Journal of Quantum Chemistry, 28(S19):669–673, March 18–23, 1985. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Jain:1987:SFE

[JTST87] Arvind Kumar Jain, A. N. Tripathi, Vedene H. Smith, Jr., and Ajit J. Thakkar. Scattering of fast electrons and X-rays from CO_{2} molecules. International Journal of Quantum Chemistry, 32(S21):217–227, March 12, 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Chemistry, Solid-State Theory, and Computational Methods.

Julg:1984:CCB

[Jul84] A. Julg. The concept of the chemical bond. International Journal of Quantum Chemistry, 26(5):709–715, November 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Junker:1980:CSM

[Jun80] B. R. Junker. The complex stabilization method: Application to atomic resonances. International Journal of Quantum Chemistry, 18(S14):53–66, March 10–20, 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory, Collision Phenomena, Quantum Statistics, and Computational Methods.

Jayasuriya:1986:MEP

[JW86] Keerthi Jayasuriya and David J. Williams. Molecular electrostatic potential studies on antiepileptic drug: Troxidone. International Journal of Quantum Chemistry, 30(2):265–273, August 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Jensen:1983:APC

[JWÖ83] Hans Jørgen Aa. Jensen, Brian Weiner, and Yngve Öhrn. AGP Propagator calculations. International Journal of Quantum Chemistry, 23(1):65–70, January 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Jensen:1982:PPO

[JWOÖ82] H. J. A. A. Jensen, B. Weiner, J. V. Ortiz, and Y. Öhrn. A powerful procedure for optimizing AGP states. International Journal of Quantum Chemistry, 22(S16):615–631, March 1– 13, 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Chemistry, Theory of Condensed Matter, and Propagator Methods in the Quantum Theory of Matter.

Jayaraman:1981:PPL

[JY81] S. Jayaraman and N. Yathindra. Probing possible left- and right-handed poly(dinucleotide) helical conformations from (n–h) plots. models for polysequential nucleotides. International Journal of Quantum Chemistry, 20(1):211–230, July 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Jeziorek:1985:TIH

[JŻ85] Danuta Jeziorek and Bronisław Żurawski. Theoretical investigation of the Na^{+} + H_{2} system. I. Ab initio calculations including electron correlations of potential energy hypersurface. International Journal of Quantum Chemistry, 28(2):297–309, August 1985. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Kotzev:1985:IFT

[KAA85] J. N. Kotzev, M. I. Aroyo, and M. N. Angelova. Isoscalar factors for tetragonal, orthorhombic, monoclinic and triclinic Shubnikov point groups. International Journal of Quantum Chemistry, 27(1):45–58, January 1985. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Katriel:1981:MBL

[KAD81] J. Katriel, C. J. Appellof, and E. R. Davidson. Mapping between local potentials and ground state densities. International Journal of Quantum Chemistry, 19(2):293–300, February 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Kagawa:1983:GVM

[Kag83] Takashi Kagawa. General variation method for the relativistic calculations of atoms and molecules. International Journal of Quantum Chemistry, 23(3):973–979, March 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Kattija-Ari:1980:MSC

[KAH80] Mano Kattija-Ari and Marlin D. Harmony. The microwave spectrum and conformation of hydroxyacetone: The influence of hydrogen bonding on the barrier to internal rotation of the methyl group. International Journal of Quantum Chemistry, 18(S14):443–453, March 10–20, 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory, Collision Phenomena, Quantum Statistics, and Computational Methods.

Kahn:1984:ESM

[Kah84] Luis R. Kahn. Electronic structure of molecules using onecomponent wave functions and relativistic effective core potentials. International Journal of Quantum Chemistry, 25(1): 149–183, January 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Kaldor:1985:CNM

[Kal85] Uzi Kaldor. Can nondegenerate many-body perturbation theory Be applied to quasidegenerate electronic states? International Journal of Quantum Chemistry, 28(1):103–108, July 1985. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Kaldor:1986:DCE

[Kal86] Uzi Kaldor. Direct calculation of excitation energies by the coupled-cluster method: Mg and Ar atoms. International Journal of Quantum Chemistry, 30(S20):445–453, March 10– 15, 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Syposium on Atomic, Molecular, and Solid-State Theory, Scattering Problems, Many Body Phenomena, and Computational Quantum Chemistry.

Kulkarni:1985:EHI

[KAM85] Anil D. Kulkarni, S. A. Alexander, and F. A. Matsen. Effect of Hubbard interaction on solitonic excitations in polyacetylene. International Journal of Quantum Chemistry, 28(4):451–457, October 1985. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Karle:1980:SDA

[Kar80] Jerome Karle. Some developments in anomalous dispersion for the structural investigation of macromolecular systems in biology. International Journal of Quantum Chemistry, 18(S7): 357–367, March 5–8, 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Karpfen:1981:CPI

[Kar81] Alfred Karpfen. Convergence problem in ab initio crystal orbital calculations. International Journal of Quantum Chemistry, 19(6):1207–1214, June 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Karafiloglou:1984:AMS

[Kar84a] Padeleimon Karafiloglou. Analytical MO study of the singlet– triplet coupling in xylylenes. International Journal of Quantum Chemistry, 25(2):293–308, February 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Kari:1984:PCA

[Kar84b] Roy Kari. Parametrization and comparative analysis of the BFGS optimization algorithm for the determination of optimum linear coefficients. International Journal of Quantum Chemistry, 25(2):321–329, February 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Karadakov:1985:EPT

[Kar85] Peter Karadakov. An extension of the pairing theorem. International Journal of Quantum Chemistry, 27(6):699–707, June 1985. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Karadakov:1986:SPH

[Kar86] Peter Karadakov. The spin-projected Hartree–Fock method: Direct optimization schemes and stability conditions. International Journal of Quantum Chemistry, 30(2):239–264, August 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Karlsson:1989:RPU

[Kar89] Erik B. Karlsson. Recent progress in the understanding of light particle tunneling in metals. International Journal of Quantum Chemistry, 35(6):779–791, June 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Kato:1980:LIF

[Kat80] Hajime Katô. Laser-induced fluorescence of Cs_{2} and role of kinetic energy in the Franck–Condon principle. International Journal of Quantum Chemistry, 18(1):287–291, July 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Katriel:1983:RES

[Kat83] Jacob Katriel. Reduction of the excited state into the ground state of a super-Hamiltonian. International Journal of Quantum Chemistry, 23(5):1767–1780, May 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Katriel:1989:PCO

[Kat89] Jacob Katriel. Products of class operators of the symmetric group. International Journal of Quantum Chemistry, 35 (4):461–470, April 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Kaufman:1981:SCG

[Kau81] Joyce J. Kaufman. Strategy for computer-generated theoretical and quantum chemical prediction of toxicity and toxicology (and pharmacology in general). International Journal of Quantum Chemistry, 20(S8):419–439, March 5–7, 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Kaufman:1982:CSU

[Kau82a] Joyce J. Kaufman. Computational strategy using improved virtual orbitals to ensure orthogonality of excited states to the ground state and to each other for calculation of transition matrix elements. International Journal of Quantum Chemistry, 22(S16):649–651, March 1–13, 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Chemistry, Theory of Condensed Matter, and Propagator Methods in the Quantum Theory of Matter.

Kaur:1982:DSC

[Kau82b] Livjeet Kaur. Diamagnetic shielding constants of firstrow atoms: Studies with a piecewise exponentially decaying electron density function. International Journal of Quantum Chemistry, 22(5):1037–1039, November 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Kaufman:1983:IPE

[Kau83] Joyce J. Kaufman. Ab-initio potential energy functions for scattering from energy-partitioned ab-initio SCF intermolecular energies plus dispersion. International Journal of Quantum Chemistry, 24(S17):197–198, March 3–12, 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-state Theory, Collision Phenomena, and Computational Quantum Chemistry.

Kaufman:1987:SNM

[Kau87] Joyce J. Kaufman. Symposium note: More new desirable computational strategies for ab initio calculations on large molecules, clusters, solids, and crystals. International Journal of Quantum Chemistry, 31(1):179–184, January 1987. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Keller:1980:AKS

[KB80] Jaime Keller and Walter Baltensperger. The analysis of the Knight shift in the μ^{+} SR technique. International Journal of Quantum Chemistry, 18(S14):123–129, March 10–20, 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory, Collision Phenomena, Quantum Statistics, and Computational Methods.

Ketkar:1986:XRI

[KB86] S. N. Ketkar and R. A. Bonham. The x-ray incoherent scattering factor for molecular nitrogen as determined by high energy electron scattering. International Journal of Quantum Chemistry, 30(S20):627–633, March 10–15, 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Syposium on Atomic, Molecular, and Solid-State Theory, Scattering Problems, Many Body Phenomena, and Computational Quantum Chemistry.

Kumicaak:1987:SAA

[KB87a] Juraj Kumičáak and Erkki Brändas. Some aspects of approach to equilibrium in classical and quantum systems. International Journal of Quantum Chemistry, 32(6):669–683, December 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Kumicak:1987:AEL

[KB87b] Juraj Kumičák and Erkki Brändas. Approach to equilibrium and Lyapunov converters. International Journal of Quantum Chemistry, 32(S21):749–750, March 12, 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Chemistry, Solid-State Theory, and Computational Methods.

Kucharski:1988:MMB

[KB88] Stanislaw A. Kucharski and Rodney J. Bartlett. Multireference many-body perturbation theory. International Journal of Quantum Chemistry, 34(S22):383–405, March 12–19, 1988. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Syposium on Quantum Chemistry, Solid-State Theory, and Computational Methods.

Kothekar:1983:PIP

[KBR83] V. Kothekar, G. Bolis, and Robert Rein. Possible incorporation of the purine–purine mispairs in the DNA helix and the interpretation of the transversion-type point mutations. International Journal of Quantum Chemistry, 23(4):1295–1303, April 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Karpfen:1981:HBC

[KBS81] A. Karpfen, A. Beyer, and P. Schuster. Hydrogen bonding in clusters and molecular crystals. International Journal of Quantum Chemistry, 19(6):1113–1119, June 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Kumar:1981:EMS

[KBSG81] Kamal Kumar, Mahesh C. Bindal, Prithvi Singh, and Satya P. Gupta. Effect of molecular size on carbonic anhydrase inhibition by sulfonamides. International Journal of Quantum Chemistry, 20(1):123–129, July 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Kojima:1980:ETT

[KBT80] N. Kojima, T. Ban, and I. Tsujikawa. Exciton transfer in the two-dimensional antiferromagnet (C_{2}H_{5}NH_{3})_{2}MnCl_{4}. International Journal of Quantum Chemistry, 18(2):619–623, August 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Karadakov:1982:ISA

[KC82a] P. Karadakov and O. Castaño. Instabilities of the symmetryadapted Restricted-Hartree–Fock ground state in infinite polyenes. I. Singlet instabilities. International Journal of Quantum Chemistry, 22(2):351–365, August 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Karadakov:1982:NDS

[KC82b] P. Karadakov and O. Castaño. New derivation of (in)stability conditions for projected Hartree–Fock solutions. International Journal of Quantum Chemistry, 22(5):1033–1036, November 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Karunakaran:1982:ECCa

[KC82c] K. M. Karunakaran and Ralph E. Christoffersen. Explicitly correlated configuration interaction studies using spherical Gaussians. I. Theoretical considerations. International Journal of Quantum Chemistry, 22(1):1–9, July 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Karunakaran:1982:ECCb

[KC82d] K. M. Karunakaran and Ralph E. Christoffersen. Explicitly correlated configuration interaction studies using spherical Gaussians. II. Exploratory studies on LiH. International Journal of Quantum Chemistry, 22(1):11–21, July 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Karadakov:1983:SPC

[KC83] Peter Karadakov and Obis Castaño. Stability properties of closed-shell restricted Hartree–Fock solutions for electronic systems in the framework of the projected Hartree–Fock method and their utilization. International Journal of Quantum Chemistry, 24(5):453–477, November 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Kolos:1980:TSS

[KCC80] W. Kołos, G. Corongiu, and E. Clementi. Theoretical study of some van der Waals molecules. International Journal of Quantum Chemistry, 17(4):775–798, April 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Kapuy:1983:AMB

[KCK83] Ede Kapuy, Zoltán Csépes, and Cornelia Kozmutza. Application of the many-body perturbation theory by using localized orbitals. International Journal of Quantum Chemistry, 23(3):

981–990, March 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Kovarskii:1980:EMN

[KCP80] V. A. Kovarskii, I. A. Chaikovskii, and E. A. Popov. Electron mechanism of nonequilibrium molecular vibration excitation in biological systems. International Journal of Quantum Chemistry, 17(2):273–283, February 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Katriel:1980:SVS

[KD80a] J. Katriel and E. R. Davidson. Shellwise virial scaling: Approximation for atomic hole states. International Journal of Quantum Chemistry, 18(4):1049–1055, October 1980. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Kothekar:1980:RCI

[KD80b] V. Kothekar and S. Dutta. Role of CA^{2+} ion in the abortifacient action of prostaglandins. II. Molecular orbital and conformation energy calculations of PGA_{1}. International Journal of Quantum Chemistry, 18(3):891–904, September 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Kirtman:1986:ALS

[KD86] Bernard Kirtman and Celso P. Demelo. Accurate local-space treatment of hydrogen bonding in large systems. International Journal of Quantum Chemistry, 29(5):1209–1222, May 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Kobeissi:1981:EEF

[KDA81] Hafez Kobeissi, Mounzer Dagher, and Mohamad Adel Alameddine. Explicit expression of the Franck–Condon factors in terms of the potentials of the two states. International Journal of Quantum Chemistry, 20(3):633–644, September 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Kurtz:1982:CGA

[KE82] Henry A. Kurtz and Nils Elander. On the calculation of generalized antisymmetric geminal power (GAGP) wavefunctions.

International Journal of Quantum Chemistry, 22(S16):605– 614, March 1–13, 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Chemistry, Theory of Condensed Matter, and Propagator Methods in the Quantum Theory of Matter.

Krylstedt:1986:CSL

[KEB86] Peter Krylstedt, Nils Elander, and Erkki Brändas. Complex scaled local density calculations: A review of recent developments. International Journal of Quantum Chemistry, 30 (S20):253–269, March 10–15, 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Syposium on Atomic, Molecular, and Solid-State Theory, Scattering Problems, Many Body Phenomena, and Computational Quantum Chemistry.

Krylstedt:1987:SRE

[KEB87] Peter Krylstedt, Nils Elander, and Erkki Brändas. Shape resonances in electron–molecule scattering. A proposed model in terms of S matrix poles and associated generalized quantum numbers. International Journal of Quantum Chemistry, 31 (5):755–794, May 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Kurtz:1981:EOA

[KEGS81] Henry A. Kurtz, Nils Elander, Osvaldo Goscinski, and Erland Sangfelt. On the energy optimization of the AGP wavefunction. The ground state of the Be atom. International Journal of Quantum Chemistry, 20(S15):143–149, March 8–14, 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory, Collision Phenomena, and Computational Quantum Chemistry.

Katrib:1986:MBR

[KEIG86] A. Katrib, B. D. El-Issa, and R. Ghodsian. Multiple bonds and re(4 f) binding energies of some rhenium halides. International Journal of Quantum Chemistry, 29(4):993–999, April 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Kelm:1978:HPC

[Kel78] H. Kelm, editor. High pressure chemistry: proceedings of the NATO Advanced Study Institute held in Corfu, Greece, September 24–October 8, 1977, volume 41 of NATO advanced study institutes series: Series C, Mathematical and physical sciences. D. Reidel, Dordrecht, Boston, Lancaster, Tokyo, 1978. ISBN 90-277-0935-1. LCCN QD455 .H48.

Keller:1986:FHK

[Kel86] Jaime Keller. On the formulation of the Hohenberg–Kohn– Sham theory. International Journal of Quantum Chemistry, 30(S20):767–768, March 10–15, 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Syposium on Atomic, Molecular, and Solid-State Theory, Scattering Problems, Many Body Phenomena, and Computational Quantum Chemistry.

Kern:1981:EEC

[Ker81] R. Kern. Experimental evidence on the chemisorption bond of two- and one-dimensional boundaries. International Journal of Quantum Chemistry, 19(6):1017–1025, June 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Kertesz:1986:ESH

[Ker86] Miklos Kertesz. Electronic structure of highly doped conducting polymers. International Journal of Quantum Chemistry, 29(5):1165–1176, May 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Katriel:1984:WTM

[KFD84] J. Katriel, D. Feller, and E. R. Davidson. Why is there a molecular relativistic effect? International Journal of Quantum Chemistry, 26(4):489–495, October 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Kinoshita:1985:VCM

[KFD85] Dina L. Kinoshita, Luiz G. Ferreira, and Manoel L. De Siqueira. Variational cellular method for polyatomic molecules: SF_{6}. International Journal of Quantum Chemistry, 28(1):85–101, July 1985. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Kono:1980:FCE

[KFN80] H. Kôno, Y. Fujimura, and T. Nakajima. Field correlation effects on the resonant light scattering. International Journal of Quantum Chemistry, 18(1):293–300, July 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Kumar:1980:CFE

[KG80] N. Vasanth Kumar and Girjesh Govil. Conformation and flexibility of the 3′-end of tRNA. International Journal of Quantum Chemistry, 18(S7):101–113, March 5–8, 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Kibler:1985:ELP

[KG85] Maurice Kibler and Geneviève Grenet. Energy levels of paramagnetic ions: Algebra. III. The case of d^{N} ions in cubical symmetry. International Journal of Quantum Chemistry, 28 (2):213–232, August 1985. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Karna:1986:ESS

[KG86a] S. P. Karna and F. Grein. Erratum: Semidiffuse states of diatomic molecules: Configuration interaction studies of the lowest ^{2}Σ^{+} and ^{2}Δ states of NS, SiF, and CCI. International Journal of Quantum Chemistry, 30(3):449, September 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). See [KG86b].

Karna:1986:SSD

[KG86b] S. P. Karna and F. Grein. Semidiffuse states of diatomic molecules: Configuration interaction studies of the lowest^{2}Σ^{+} and ^{2}Δ states of NS, SiF, and CCl. International Journal of Quantum Chemistry, 29(4):755–766, April 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). See erratum [KG86a].

Kibler:1986:ELPa

[KG86c] Maurice Kibler and Geneviève Grenet. Energy levels of paramagnetic ions: Algebra. IV. The case of f^{N} ions in cubical symmetry. International Journal of Quantum Chemistry, 29 (1):11–22, January 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Kibler:1986:ELPb

[KG86d] Maurice Kibler and Geneviève Grenet. Energy levels of paramagnetic ions: Algebra. V. Weak-field models. International Journal of Quantum Chemistry, 29(3):485–495, March 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Karna:1989:IMC

[KG89] S. P. Karna and F. Grein. Ab initio MO calculations of hyperfine coupling constants. A basis set study for ^{19}F, ^{35}Cl, ^{19}F, and^{35}Cl. International Journal of Quantum Chemistry, 36(3): 265–275, September 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Kovacevic:1987:HSA

[KGB87] K. Kovačević, A. Graovac, and D. Babić. On the hybridization in some Archimedean carbon clusters. International Journal of Quantum Chemistry, 32(S21):589–593, March 12, 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Chemistry, Solid-State Theory, and Computational Methods.

Karna:1989:HCC

[KGEP89] S. P. Karna, F. Grein, B. Engels, and S. D. Peyerimhoff. The hyperfine coupling constants of ^{19}F: An ab initio MRD-CI basis set study. International Journal of Quantum Chemistry, 36(3):255–263, September 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Kohn:1989:DFT

[KGO89] W. Kohn, E. K. U. Gross, and L. N. Oliveira. Density functional theory for superconductors. International Journal of Quantum Chemistry, 36(S23):611–615, April 1–8, 1989. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Karo:1980:SRD

[KH80] Arnold M. Karo and John R. Hardy. Shock and reaction dynamics in condensed materials. International Journal of Quantum Chemistry, 18(S14):695, March 10–20, 1980. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory, Collision Phenomena, Quantum Statistics, and Computational Methods.

Karo:1985:GRV

[KH85] A. M. Karo and J. R. Hiskes. Generation and relaxation of vibrationally-excited H_{2} molecules by wall collisions. International Journal of Quantum Chemistry, 28(S19):729, March 18–23, 1985. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Karo:1986:MDS

[KH86] A. M. Karo and J. R. Hardy. Molecular dynamics simulations of energy transfer in shocked molecular systems. International Journal of Quantum Chemistry, 30(S20):763–765, March 10–15, 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Syposium on Atomic, Molecular, and Solid-State Theory, Scattering Problems, Many Body Phenomena, and Computational Quantum Chemistry.

Khait:1985:NKM

[Kha85] Yury L. Khait. A new kinetic many-body theory of rate processes in and on solids and its applications. International Journal of Quantum Chemistry, 28(S19):749–750, March 18–23, 1985. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Kaufman:1985:MDC

[KHCH85] Joyce J. Kaufman, P. C. Hariharan, C. Chabalowski, and Matti Hotokka. Multireference determinant CI calculations and CASSCF calculations on the CH_{3} − NO_{2} decomposition pathway of nitromethane. International Journal of Quantum Chemistry, 28(S19):221–235, March 18–23, 1985. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Khetrapal:1981:ANO

[Khe81] C. L. Khetrapal. Application of NMR of oriented systems in biochemistry and biophysics. International Journal of Quantum Chemistry, 20(2):485–494, August 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Kaufman:1987:IMCb

[KHK87] Joyce J. Kaufman, P. C. Hariharan, and Phillip B. Keegstra. Ab initio MRD–CI calculations for the propagation step of cationic polymerization of oxetanes based on localized orbitals. International Journal of Quantum Chemistry, 32(S21):623– 643, March 12, 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Chemistry, Solid-State Theory, and Computational Methods.

Kashiwagi:1986:IMC

[KHNT86] Hiroshi Kashiwagi, Fumihiko Hirota, Umpei Nagashima, and Toshikazu Takada. Ab initio MO calculations on the chlorophyll–water system and estimation of the structure of the special pair. International Journal of Quantum Chemistry, 30(3):311–326, September 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Kaufman:1981:ANC

[KHP81] Joyce J. Kaufman, P. C. Hariharan, and Herbert E. Popkie. Additional new computational strategies for ab initio calculations on large molecules. International Journal of Quantum Chemistry, 20(S15):199–201, March 8–14, 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory, Collision Phenomena, and Computational Quantum Chemistry.

Kaufman:1987:IMCa

[KHRK87] Joyce J. Kaufman, P. C. Hariharan, Szczepan Roszak, and P. B. Keegstra. Ab initio MRD-CI calculations on protonated cyclic ethers. I. Protonation pathways involve multipotential surfaces (protonation of oxetane). II. Differences from SCF in dominant configurations upon opening non-protonated oxirane rings (epoxides). International Journal of Quantum Chemistry, 32(S14):37–46, March 14, 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Khuri:1982:FSA

[Khu82] N. N. Khuri. Failure of strong asymptotic conditions in fourdimensional field theories. International Journal of Quantum Chemistry, 21(1):173–177, January 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Kiang:1980:DAM

[Kia80] Yuan-Sun Kiang. Determinant of adjacency matrix and Kekulé structures. International Journal of Quantum Chemistry, 18(S14):541–547, March 10–20, 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory, Collision Phenomena, Quantum Statistics, and Computational Methods.

Kiang:1981:PTM

[Kia81] Yuan-Sun Kiang. Partition technique and molecular graph theory. International Journal of Quantum Chemistry, 20(S15): 293–304, March 8–14, 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory, Collision Phenomena, and Computational Quantum Chemistry.

Kibler:1983:CGC

[Kib83] M. R. Kibler. Clebsch–Gordan coefficients for chains of groups of interest in quantum chemistry. III. The point symmetry groups. International Journal of Quantum Chemistry, 23(1): 115–125, January 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Kim:1981:GTM

[Kim81] Shoon K. Kim. A general theory of matrix transformation and its application to quantum mechanical problems. International Journal of Quantum Chemistry, 20(S15):321–329, March 8–14, 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory, Collision Phenomena, and Computational Quantum Chemistry.

King:1986:GTS

[Kin86] R. Bruce King. Graph theory in the study of metal cluster bonding topology: Applications to metal clusters having fused polyhedra. International Journal of Quantum Chemistry, 30 (S20):227–238, March 10–15, 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Syposium on Atomic, Molecular, and Solid-State Theory, Scattering Problems, Many Body Phenomena, and Computational Quantum Chemistry.

King:1989:WIC

[Kin89] James W. King. A Z-weighted information content index. International Journal of Quantum Chemistry, 36(S16):165–170, April 1–8, 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Kirtman:1989:FCA

[Kir89] Bernard Kirtman. Finite chain approximation for linear and nonlinear polarizabilities of polydiacetylene and polybutatriene. International Journal of Quantum Chemistry, 36(2): 119–125, August 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Kong:1987:SPT

[KJL87] Young Shik Kong, Mu Shik Jhon, and Per-Olov Löwdin. Studies on proton transfers in water clusters and DNA base pairs. International Journal of Quantum Chemistry, 32(S14):189– 209, March 14, 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Kolos:1985:QCC

[KJMS85] W. Kołos, B. Jeziorski, H. J. Monkhorst, and K. Szalewicz. Quantum chemical contribution to electron neutrino mass determination. International Journal of Quantum Chemistry, 28 (S19):421–441, March 18–23, 1985. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Krivka:1985:CCP

[KJT85] P. Křivka, Ž. Jeričević, and N. Trinajstić. On the computation of characteristic polynomial of a chemical graph. International Journal of Quantum Chemistry, 28(S19):129–147, March 18–23, 1985. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Kobeissi:1982:AER

[KK82a] Hafez Kobeissi and Mahmoud Korek. Analytic expression of the rotation harmonics in the vibration–rotation wave function of a diatomic molecule. International Journal of Quantum Chemistry, 22(1):23–29, July 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Krechl:1982:MOS

[KK82b] Jiří Krechl and Josef Kuthan. Molecular orbital study on the active center of lactate dehydrogenase. International Journal of Quantum Chemistry, 21(6):1029–1040, June 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Krechl:1983:SMS

[KK83] Jiří Krechl and Josef Kuthan. Semiempirical MO study of pyruvate substrate behavior in the active center of lactate dehydrogenase. International Journal of Quantum Chemistry, 24 (5):479–489, November 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Karwowski:1985:QM

[KK85] J. Karwowski and J. Kobus. Quasirelativistic methods. International Journal of Quantum Chemistry, 28(6):741–756, December 1985. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Karwowski:1986:DSO

[KK86] J. Karwowski and J. Kobus. The Dirac second-order equation and an improved quasirelativistic theory of atoms. International Journal of Quantum Chemistry, 30(6):809–819, December 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Kilzer:1987:ESP

[KK87] Patricia A. Kilzer and Henry A. Kurtz. Excited state properties utilizing effective core potentials. International Journal of Quantum Chemistry, 32(S21):539–545, March 12, 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Chemistry, Solid-State Theory, and Computational Methods.

Kerteasz:1980:ESP

[KKA80a] Miklóas Kertéasz, Jožke Koller, and Andrej Ažkman. On the electronic structure of periodic polyglycine. International Journal of Quantum Chemistry, 18(S7):177–179, March 5–8, 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Kertesz:1980:NDE

[KKA80b] M. Kertész, J. Koller, and A. Ažman. Nuclear distortion of the equidistant arrangement in trans-polyacetylene, (CH)_{x}. International Journal of Quantum Chemistry, 18(2):645–650, August 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Kertesz:1980:NEC

[KKA80c] Miklos Kertesz, Joze Koller, and Andrej Azman. Need for electronic correlation calculations in polymers. International Journal of Quantum Chemistry, 18(S14):463–466, March 10– 20, 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory, Collision Phenomena, Quantum Statistics, and Computational Methods.

Kluge:1983:PSH

[KKA^{+}83] G. Kluge, G. Kania, F. Achenbach, E. Lippmann, I. Novak, and L. Klasinc. Photoelectron spectroscopy of heterocycles: Ortho-disubstituted quinoxalines. International Journal of Quantum Chemistry, 24(S10):339–347, March 14–16, 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Kluge:1984:PSG

[KKA^{+}84] Gert Kluge, Gabriele Kania, Frank Achenbach, Horst Wilde, Igor Novak, and Leo Klasinc. Photoelectron spectra and gas phase tautomerism of some pyrazolones. International Journal of Quantum Chemistry, 26(S11):237–248, March 12–15, 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Kumar:1986:PLO

[KKC^{+}86] D. Kumar, L. Klasinc, P. L. Clancy, R. V. Nauman, and S. P. Mcglynn. Pulsed laser optogalvanic spectroscopy of nitrogen in a radiofrequency discharge. International Journal of Quantum Chemistry, 30(S20):635–645, March 10–15, 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Syposium on Atomic, Molecular, and Solid-State Theory, Scattering Problems, Many Body Phenomena, and Computational Quantum Chemistry.

Kumar:1985:PLO

[KKCM85] D. Kumar, L. Klasinc, P. L. Clancy, and S. P. McGlynn. Pulsed laser optogalvanic spectroscopy of xenon in RF discharge. International Journal of Quantum Chemistry, 28 (S19):403–410, March 18–23, 1985. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Kumar:1987:PLO

[KKCM87] D. Kumar, L. Klasinc, P. L. Clancy, and S. P. McGlynn. Pulsed laser optogalvanic spectroscopy of xenon in RF discharge. International Journal of Quantum Chemistry, 31 (2):327, February 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Kryachko:1980:NEF

[KKD80] Eugene S. Kryachko, Yuri A. Kruglyak, and Bryan J. Dalton. Nonrigidity effects of fragments of membrane-active complexons. International Journal of Quantum Chemistry, 17(6): 1229–1235, June 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Kaufman:1983:PTP

[KKH^{+}83] Joyce J. Kaufman, Walter S. Koski, P. C. Hariharan, Jack Crawford, David M. Garmer, and Louis Chan-Lizardo. Prediction of toxicology and pharmacology based on model toxicophores and pharmacophores using the new TOX-MATCH– PHARM-MATCH program. International Journal of Quantum Chemistry, 24(S10):375–415, March 14–16, 1983. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Klasinc:1987:PSB

[KKSM87] L. Klasinc, B. Kovač, A. Sabljić, and S. P. McGlynn. Photoelectron spectroscopy of biologically active molecules. 14. Some analgesic-antipyretic and anti-inflammatory agents. International Journal of Quantum Chemistry, 32(S14):317–324, March 14, 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Kubota:1980:ETP

[KKTE80] R. Kubota, H. Kobayashi, I. Tsujikawa, and T. Enoki. Electrical and thermal properties of the partially oxidized salt of the Magnus green salt with the formula Pt_{6}(NH_{3})_{10}Cl_{10}(HSO_{4})_{4}. International Journal of Quantum Chemistry, 18(6):1533– 1538, December 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Karelson:1986:RFE

[KKZ86] Mati M. Karelson, Alan R. Katritzky, and Michael C. Zerner. Reaction field effects on the electron distribution and chemical reactivity of molecules. International Journal of Quantum Chemistry, 30(S20):521–527, March 10–15, 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Syposium on Atomic, Molecular, and Solid-State Theory, Scattering Problems, Many Body Phenomena, and Computational Quantum Chemistry.

Kertesz:1980:ENS

[KL80] Judith C. Kertesz and Koloman Laki. Effect of neutral salts and methylglyoxal on the cis-trans equilibrium of nmethylacetamide. A 13-carbon nuclear magnetic resonance study. International Journal of Quantum Chemistry, 18(S7): 75–81, March 5–8, 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Kresin:1981:NAA

[KL81] Vladimir Z. Kresin and William A. Lester, Jr. A new adiabatic approach to the photodissociation of polyatomic molecules. International Journal of Quantum Chemistry, 20(S15):703– 714, March 8–14, 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory, Collision Phenomena, and Computational Quantum Chemistry.

Keller:1987:DFT

[KL87] Jaime Keller and Eduardo Ludeña. Density functional theory formalism. International Journal of Quantum Chemistry, 32 (S21):171–180, March 12, 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Chemistry, Solid-State Theory, and Computational Methods.

Kresin:1989:NHE

[KL89] Vladimir Z. Kresin and William A. Lester, Jr. Nonadiabatic Hamiltonian for electronic–vibrational coupling. International Journal of Quantum Chemistry, 36(S23):17–26, April 1–8, 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Kresin:1984:CRQ

[KLDD84] V. Z. Kresin, W. A. Lester, Jr., M. Dupuis, and C. E. Dateo. Chemical reaction as a quantum transition. International Journal of Quantum Chemistry, 26(S18):691–700, March 1– 15, 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Atomic, Molecular and Solid-State Theory, and Computational Quantum Chemistry.

Klessinger:1983:FMM

[Kle83] Martin Klessinger. The fragments-in-molecules method. II. Fim method for σ-bonded systems. International Journal of Quantum Chemistry, 23(2):535–542, February 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Klein:1986:CGT

[Kle86] D. J. Klein. Chemical graph-theoretic cluster expansions. International Journal of Quantum Chemistry, 30(S20):153–171, March 10–15, 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Syposium on Atomic, Molecular, and Solid-State Theory, Scattering Problems, Many Body Phenomena, and Computational Quantum Chemistry.

Klein:1987:EPT

[Kle87] D. J. Klein. Exchange perturbation theories. International Journal of Quantum Chemistry, 32(3):377–396, September 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Kertesz:1989:SES

[KLS89] Miklos Kertesz, Yong S. Lee, and James J. P. Stewart. Structure and electronic structure of polyacene. International Journal of Quantum Chemistry, 35(2):305–313, February 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Komiyama:1980:SEC

[KM80] Tomoko Komiyama and Makoto Miwa. Solvent effects on CD and MCD spectra of phenylalanine and its related compounds. International Journal of Quantum Chemistry, 18(2):527–531, August 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Kemister:1985:IPM

[KM85] G. Kemister and N. H. March. Ionization potentials of molecular systems with many electrons. International Journal of Quantum Chemistry, 28(S19):257–265, March 18–23, 1985. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Kozlowski:1989:ADE

[KM89a] P. M. Kozlowski and N. H. March. Approximate density– external potential relation and the Pauli potential for systems with Coulombic interaction. International Journal of Quantum Chemistry, 36(6):741–748, December 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Kuruoglu:1989:CRH

[KM89b] Zeki C. Kuruoglu and David A. Micha. Calculation of resonances in the H + H_{2} reaction using the Faddeev–AGS method. International Journal of Quantum Chemistry, 36 (S23):103–114, April 1–8, 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Kothekar:1988:PMR

[KMKC88] V. Kothekar, Mrigank, A. Kotwal, and B. Chandrashekhar. On the possible mode of regulation of DNA transcription by steroid hormones: Glucocorticoids. International Journal of Quantum Chemistry, 34(S15):167–178, March 12–19, 1988. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Syposium on Quantum Biology and Quantum Pharmacology.

Klasinc:1989:PSB

[KMPTK89] L. Klasinc, S. P. Mcglynn, L. J. Paša-Tolić, and B. Kovač. Photoelectron spectroscopy of biologically active molecules.

17. unsaturated steroids. International Journal of Quantum Chemistry, 36(S16):331–341, April 1–8, 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Kothekar:1986:ILE

[KMRK86] V. Kothekar, Mrigank, A. K. Royyuru, and A. Kalia. Influence of local excitations in DNA conformation on binding of nonintercalating antitumor antibiotic in the minor groove. International Journal of Quantum Chemistry, 30(S13):175–183, March 6–8, 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Klimko:1980:FCF

[KMW80] G. T. Klimko, M. M. Mestechkin, and G. E. Whyman. Fock coordinate function method for separation of spin variables in transition density matrices. International Journal of Quantum Chemistry, 17(3):415–428, March 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Kai:1980:EST

[KN80] Eiko Kai and Kichisuke Nishimoto. Electronic structure of transition-metal complexes. I. Parametrization for CNDO/2 method. International Journal of Quantum Chemistry, 18(2): 403–408, August 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Kibler:1984:MPR

[KN84] Maurice Kibler and Tidjani Negadi. Motion of a particle in a ring-shaped potential: An approach via a nonbijective canonical transformation. International Journal of Quantum Chemistry, 26(3):405–410, September 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Kozlowski:1986:GAG

[KN86] Pawel M. Kozłowski and Roman F. Nalewajski. A graph approach to the gradient expansion of density functionals. International Journal of Quantum Chemistry, 30(S20):219–226, March 10–15, 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Syposium on Atomic, Molecular, and Solid-State Theory, Scattering Problems, Many Body Phenomena, and Computational Quantum Chemistry.

Kestner:1983:PEC

[KNM83] N. R. Kestner, M. D. Newton, and T. L. Mathers. Potential energy calculations for various water dimer configurations. International Journal of Quantum Chemistry, 24(S17):431–439, March 3–12, 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-state Theory, Collision Phenomena, and Computational Quantum Chemistry.

Klasinc:1986:PSB

[KNS^{+}86] L. Klasinc, I. Novak, A. Sabljić, S. P. Mcglynn, and L. Klasinc. Photoelectron spectroscopy of biologically active molecules.

12. benzene-containing amides. International Journal of Quantum Chemistry, 30(S13):251–260, March 6–8, 1986. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Klasinc:1988:PSB

[KNSM88] L. Klasinc, I. Novak, A. Sabljić, and S. P. McGlynn. Photoelectron spectroscopy of biologically active molecules. XVI. Benzophenone derivatives. International Journal of Quantum Chemistry, 34(S15):259–266, March 12–19, 1988. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Syposium on Quantum Biology and Quantum Pharmacology.

Kitagawa:1980:PST

[KO80] M. Kitagawa and Y. H. Ohtsuki. Plasma solitons trailing moving ions. International Journal of Quantum Chemistry, 18 (S14):637–644, March 10–20, 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory, Collision Phenomena, Quantum Statistics, and Computational Methods.

Kashiwagi:1981:IMO

[KO81] Hiroshi Kashiwagi and Shigeru Obara. Ab initio molecular orbital calculation of Fe-porphine with a double zeta basis set. International Journal of Quantum Chemistry, 20(4):843– 859, October 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Kobeissi:1982:SEE

[Kob82] Hafez Kobeissi. Separate eigenvalue equations for the rotational constants for any potential in diatomic molecules. International Journal of Quantum Chemistry, 22(S16):293–300, March 1–13, 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461x (electronic). Supplement: Proceedings of the International Symposium on Quantum Chemistry, Theory of Condensed Matter, and Propagator Methods in the Quantum Theory of Matter.

Kobeissi:1986:DVR

[Kob86] H. Kobeissi. On the diatomic vibration–rotation eigenvalue problem. International Journal of Quantum Chemistry, 29 (1):163–164, January 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Koelling:1986:BDM

[Koe86] D. D. Koelling. The band description of materials with localizing orbitals. International Journal of Quantum Chemistry, 30(S20):377–392, March 10–15, 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Syposium on Atomic, Molecular, and Solid-State Theory, Scattering Problems, Many Body Phenomena, and Computational Quantum Chemistry.

Koga:1984:MDI

[Kog84] Toshikatsu Koga. Momentum density interpretation of the H(1s)–H^{+} long-range force. International Journal of Quantum Chemistry, 25(2):347–354, February 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Kohonen:1986:SOM

[Koh86] Teuvo Kohonen. Self-organization, memorization, and associative recall of sensory information by brain-like adaptive networks. International Journal of Quantum Chemistry, 30(S13): 209–221, March 6–8, 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Kohler:1989:MMH

[Köh89] H. S. Köhler. Microscopic model for heavy-ion nonrelativistic collisions. International Journal of Quantum Chemistry, 35 (6):793–799, June 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Konaka:1983:EFS

[Kon83] Shigehiro Konaka. Elastic forward scattering of high-energy electrons and molecular electron density. International Journal of Quantum Chemistry, 23(6):1891–1901, June 1983. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Kothekar:1981:SMM

[Kot81] V. Kothekar. Specificity and molecular mechanism of abortificient action of prostaglandins. International Journal of Quantum Chemistry, 20(1):167–178, July 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Kothekar:1982:CQC

[Kot82] V. Kothekar. Conformational and quantum chemical studies on prostaglandins. International Journal of Quantum Chemistry, 22(S9):281–288, March 4–6, 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Kotani:1984:SRM

[Kot84] Masao Kotani. Scientific reminiscence—my pilgrimage through quantum molecular sciences. International Journal of Quantum Chemistry, 26(S18):11–20, March 1–15, 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Atomic, Molecular and Solid-State Theory, and Computational Quantum Chemistry.

Koutecky:1980:SPS

[Kou80] J. Koutecký. Some problems of surface quantum chemistry. International Journal of Quantum Chemistry, 18(S14):696, March 10–20, 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory, Collision Phenomena, Quantum Statistics, and Computational Methods.

Kucar:1987:MCD

[KPH87] Jasna Kučar, Milica Pavlović, and Josip Hendeković. A method for constructing diabatic states by interpolation on the reduced density matrix. International Journal of Quantum Chemistry, 32(6):699–704, December 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Kahn:1988:MCR

[KPH88] S. D. Kahn, C. F. Pau, and W. J. Hehre. Models for chemical reactivity: Mapping of intermolecular potentials onto electron density surfaces. International Journal of Quantum Chemistry, 34(S22):575–591, March 12–19, 1988. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Syposium on Quantum Chemistry, Solid-State Theory, and Computational Methods.

Katriel:1983:SCI

[KPM83] J. Katriel, R. Pauncz, and J. J. C. Mulder. Studies in the configuration interaction method. II. Generating functions and recurrence relations for the number of many-particle configurations. International Journal of Quantum Chemistry, 23 (5):1855–1867, May 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Katriel:1985:GDI

[KPP85] J. Katriel, J. Paldus, and R. Pauncz. Generalized Dirac identities and explicit relations between the permutational symmetry and the spin operators for systems of identical particles. International Journal of Quantum Chemistry, 28(2):181–202, August 1985. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Katriel:1986:GDI

[KPP86] J. Katriel, J. Paldus, and R. Pauncz. Generalized Dirac identities and explicit relations between the permutational symmetry and the spin operators for systems of identical particles. International Journal of Quantum Chemistry, 29(1):171, January 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Kryachko:1987:MLSa

[KPS87a] Eugene S. Kryachko, Ivan Zh. Petkov, and Mario V. Stoitsov. Method of local-scaling transformations and density functional theory in quantum chemistry. II. The procedure for reproducing a many-electron wave function from x-ray diffraction data on one-electron density. International Journal of Quantum Chemistry, 32(4):467–472, October 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Kryachko:1987:MLSb

[KPS87b] Eugene S. Kryachko, Ivan Zh. Petkov, and Mario V. Stoitsov. Method of local-scaling transformations and density functional theory in quantum chemistry. III. The energy density functional: Spin-restricted approach. International Journal of Quantum Chemistry, 32(4):473–489, October 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Kryachko:1988:MLS

[KPS88] Eugene S. Kryachko, Ivan Zh. Petkov, and Mario V. Stoitsov. Method of local-scaling transformations and density functional theory in quantum chemistry. III. The energy density functional: Spin-restricted approach. International Journal of Quantum Chemistry, 34(3):307, September 1988. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Kuropteva:1980:EIM

[KPSK80] Z. V. Kuropteva, M. K. Pulatova, M. A. Smotryaeva, and K. E. Kruglyakova. Electronic interactions of N-methyl-N-nitrosourea with DNA and its chemical modifications. International Journal of Quantum Chemistry, 17(6):1221–1227, June 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Kwiatkowski:1987:ETS

[KPSS87] Józef S. Kwiatkowski, Willis B. Person, Krystyna Szczepaniak, and Marian Szcześniak. Experimental and theoretical studies of the stabilities and infrared spectra of the cytosine tautomers. International Journal of Quantum Chemistry, 32 (S21):763–764, March 12, 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Chemistry, Solid-State Theory, and Computational Methods.

Krasnogolovets:1988:MBFb

[KPT88] V. V. Krasnogolovets, N. A. Protsenko, and P. M. Tomchuk. The mechanism of bacteriorhodopsin functioning. II. Ejection and injection of proton. International Journal of Quantum Chemistry, 33(4):349–367, April 1988. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Krasnogolovets:1988:MBFa

[KPTG88] V. V. Krasnogolovets, N. A. Protsenko, P. M. Tomchuk, and V. S. Guriev. The mechanism of bacteriorhodopsin functioning. I. The light-induced proton throw-over by retinal. International Journal of Quantum Chemistry, 33(4):327–347, April 1988. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Kannan:1981:SRF

[KR81] K. K. Kannan and M. Ramanadham. Structure, refinement, and function of human carbonic anhydrase-B. International Journal of Quantum Chemistry, 20(1):199–209, July 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Klasinc:1985:APS

[KRBM85] L. Klasinc, B. Ruščić, N. S. Bhacca, and S. P. Mcglynn. Application of photoelectron spectroscopy to biologically active molecules and their constituent parts XI: Steroids. International Journal of Quantum Chemistry, 28(S12):161–167, March 14–16, 1985. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Kolos:1980:IMM

[KRCN80] W. Kołos, G. Ranghino, E. Clementi, and O. Novaro. Interaction of methane molecules. International Journal of Quantum Chemistry, 17(3):429–448, March 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Kreuzer:1981:KEP

[Kre81] H. J. Kreuzer. Kinetic equations for physisorption. International Journal of Quantum Chemistry, 20(S15):683–694, March 8–14, 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory, Collision Phenomena, and Computational Quantum Chemistry.

Kalia:1986:SSR

[KRMK86] A. Kalia, A. K. Royyuru, Mrigank, and V. Kothekar. Sequence-specific recognition of DNA by lac-repressor headpiece. International Journal of Quantum Chemistry, 30(S13): 185–194, March 6–8, 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Klasinc:1983:APS

[KRST83] L. Klasinc, B. Ruščić, A. Sabljić, and N. Trinajstić. Application of photoelectron spectroscopy to biologically active molecules and their constituent parts. IX. 1,4-Benzodiazepin-2-ones. International Journal of Quantum Chemistry, 23(4): 1667–1676, April 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Kauder:1983:GCT

[KRT83] L. R. Kauder, A. K. Ray, and S. B. Trickey. Generator coordinate treatment of some model periodic systems. International Journal of Quantum Chemistry, 23(4):1355–1361, April 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Kryachko:1980:HKT

[Kry80] Eugene S. Kryachko. Hohenberg–Kohn theorem. International Journal of Quantum Chemistry, 18(4):1029–1035, October 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Kryachko:1982:CMS

[Kry82a] Eugene S. Kryachko. Classification of molecular symmetry by framework groups. International Journal of Quantum Chemistry, 22(3):657–659, September 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Kryachko:1982:HCM

[Kry82b] Eugene S. Kryachko. Homogeneity characteristic of molecular electronic energies. International Journal of Quantum Chemistry, 22(3):661–662, September 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Kryachko:1982:MAR

[Kry82c] Eugene S. Kryachko. Matrix analysis of N-representability necessary conditions. International Journal of Quantum Chemistry, 21(6):1123–1125, June 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Kryachko:1984:FDD

[Kry84a] Eugene S. Kryachko. Four-dimensional density and energy density functional. International Journal of Quantum Chemistry, 25(4):637–644, April 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Kryachko:1984:CDP

[Kry84b] Eugene S. Kryachko. On construction of diatomic potential energy functions. International Journal of Quantum Chemistry, 25(2):277–291, February 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Kryachko:1986:RSO

[Kry86] Eugene S. Kryachko. On the red shift of OH stretching region vibrations in ice and water. International Journal of Quantum Chemistry, 30(4):495–508, October 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Kluge:1981:CES

[KS81] G. Kluge and M. Scholz. Chemistry of excited states. XI. Interpretation of photoelectron spectra by final state single CI method within semiempirical formalisms. International Journal of Quantum Chemistry, 20(3):669–682, September 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Kent:1982:TBO

[KS82] R. D. Kent and M. Schlesinger. Two-body operator matrix elements for pure and mixed orbital configurations in the unitary group approach. International Journal of Quantum Chemistry, 22(2):223–239, August 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Kent:1984:UGA

[KS84] R. D. Kent and M. Schlesinger. Unitary group approach to crystal field theory: Projection operator approach. International Journal of Quantum Chemistry, 25(6):1111–1121, June 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Karwowski:1985:GSE

[KS85] J. Karwowski and J. Styszyński. Ground-state energies of closed-shell atoms. International Journal of Quantum Chemistry, 28(1):27–37, July 1985. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Kent:1986:GUG

[KS86a] R. D. Kent and M. Schlesinger. Graphical unitary group approach to arbitrary spin representations. International Journal of Quantum Chemistry, 30(6):737–750, December 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Klei:1986:CPE

[KS86b] Herbert E. Klei and James J. P. Stewart. Calculation of polymer elastic moduli using semiempirical methods. International Journal of Quantum Chemistry, 30(S20):529–540, March 10– 15, 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Syposium on Atomic, Molecular, and Solid-State Theory, Scattering Problems, Many Body Phenomena, and Computational Quantum Chemistry.

Kurnig:1986:AEE

[KS86c] Ingrid J. Kurnig and Steve Scheiner. Additivity of the effects of external ions and dipoles upon the energetics of proton transfer. International Journal of Quantum Chemistry, 30(S13):71–79, March 6–8, 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Koch:1987:SOO

[KS87a] Wolfhard Koch and Friedrich Franz Seelig. Symmetry orbitals for one-electron band-structure computations of onedimensional crystals. International Journal of Quantum Chemistry, 32(2):249–264, August 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Kurnig:1987:IIS

[KS87b] Ingrid J. Kurnig and Steve Scheiner. Ab initio investigation of the structure of hydrogen halide-amine complexes in the gas phase and in a polarizable medium. International Journal of Quantum Chemistry, 32(S14):47–56, March 14, 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Klein:1989:EGS

[KS89a] D. J. Klein and T. G. Schmalz. Exact ground state for a Herndon–Simpson model via resonance–theoretic cluster expansion. International Journal of Quantum Chemistry, 35(3): 373–383, March 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Koizumi:1989:TSR

[KS89b] Hiroyasu Koizumi and George C. Schatz. Transition state resonances in collinear O(^{3}P) + HCl → OH + Cl. International Journal of Quantum Chemistry, 36(S23):137–145, April 1–8, 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Klein:1983:VLS

[KSGB^{+}83] D. J. Klein, W. A. Seitz, M. A. Garcia-Bach, J. M. Picone, and D. C. Foyt. Variational localized-site cluster expansions. XI. Coherent state derivation of effective Heisenberg Hamiltonians. International Journal of Quantum Chemistry, 24(S17): 555–571, March 3–12, 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-state Theory, Collision Phenomena, and Computational Quantum Chemistry.

Kocjan:1983:CMS

[KŠHH83] D. Kocjan, T. Šolmajer, M. Hodošček, and D. Hadži. Conformational and MO studies of hydroxy-2-aminotetralins. International Journal of Quantum Chemistry, 23(4):1121–1133, April 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Knop:1983:CGI

[KSJT83] J. V. Knop, K. Szymanski, Ž. Jeričević, and N. Trinajstić. Computer generation and identification of carcinogenic bay regions in benzenoid hydrocarbons. International Journal of Quantum Chemistry, 23(2):713–722, February 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Kohda-Sudoh:1986:PAS

[KSK86] S. Kohda-Sudoh and S. Katagiri. The proton affinity and the structure of protonated species of methylsilane. International Journal of Quantum Chemistry, 29(5):1591–1598, May 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Kalcher:1984:IPC

[KSO84] Josef Kalcher, Alexander Sax, and Gottfried Olbrich. Ab initio and pseudopotential calculations on the singlet and triplet states of the disilyne isomers. International Journal of Quantum Chemistry, 25(3):543–552, March 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Kuivalainen:1981:MPM

[KSS81] P. Kuivalainen, J. Sinkkonen, and T. Stubb. Magnetic polarons in magnetic semiconductors. International Journal of Quantum Chemistry, 20(S15):507–519, March 8–14, 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory, Collision Phenomena, and Computational Quantum Chemistry.

Klein:1985:OHR

[KSSH85] D. J. Klein, T. G. Schmalz, W. A. Seitz, and G. E. Hite. Overview of Hückel- and resonance-theoretic approaches to πnetwork polymers. International Journal of Quantum Chemistry, 28(S19):707–718, March 18–23, 1985. CODEN IJQCB2.

ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Knop:1983:CT

[KST83] J. V. Knop, K. Szymanski, and N. Trinajstić. Computers of tomorrow. International Journal of Quantum Chemistry, 24 (S17):629–642, March 3–12, 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-state Theory, Collision Phenomena, and Computational Quantum Chemistry.

Knop:1984:PPR

[KST84] J. V. Knop, K. Szymanski, and N. Trinajstić. Price– performance relationship of microcomputers and main frames for calculations in theoretical chemistry. International Journal of Quantum Chemistry, 26(S18):639–645, March 1–15, 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Atomic, Molecular and Solid-State Theory, and Computational Quantum Chemistry.

Katsuki:1980:MPX

[KT80a] Shinichi Katsuki and Hiroshi Taketa. Model potential Xα method for the electronic structure calculations. International Journal of Quantum Chemistry, 18(1):25–29, July 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Knop:1980:CGT

[KT80b] J. V. Knop and N. Trinajstić. Chemical graph theory. II. On the graph theoretical polynomials of conjugated structures. International Journal of Quantum Chemistry, 18(S14):503– 520, March 10–20, 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory, Collision Phenomena, Quantum Statistics, and Computational Methods.

Klimo:1984:SPC

[KT84] Villiam Klimo and Josef Tiňo. Study of potential curves by UHF-type methods. VI. Spin-extended Hartree–Fock method. International Journal of Quantum Chemistry, 25(4):733–742, April 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Kiang:1986:GEC

[KT86] Yuan-Sun Kiang and Au-Chin Tang. A graphical evaluation of characteristic polynomials of Hückel trees. International Journal of Quantum Chemistry, 29(2):229–240, February 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Koga:1988:LIW

[KT88] Toshikatsu Koga and Ajit J. Thakkar. Linear integrability of wave functions. International Journal of Quantum Chemistry, 34(2):103–106, August 1988. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Konschin:1989:MSC

[KTGR89] Henrik Konschin, Henrik Tylli, Jukka Gynther, and Juha Rouvinen. The molecular structure and conformational characteristics of some specific benzodiazepine receptor ligands: a molecular orbital study of C3-substituted betacarboline derivatives. International Journal of Quantum Chemistry, 36 (S16):245–259, April 1–8, 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Klasinc:1980:APS

[KTK80] L. Klasinc, N. Trinajstića, and J. V. Knop. Application of photoelectron spectroscopy to biologically active molecules and their constituent parts. VIII. Thalidomide. International Journal of Quantum Chemistry, 18(S7):403–409, March 5–8, 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Kumar:1986:SHE

[KTS86] Mukesh Kumar, A. N. Tripathi, and Vedene H. Smith, Jr. Scattering of high-energy electrons and x-rays from molecules: The 10-electon series Ne, HF, H_{2}O, NH_{3}, and CH_{4}. International Journal of Quantum Chemistry, 29(5):1339–1349, May 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Kello:1980:FOD

[KU80] Vladimir Kellö and Miroslav Urban. Fourth-order diagrammatic MB–RSPT calculations of the correlation energy: N_{2}, CO, F_{2} and the reaction energy of the process \frac{1}{2}F_{2} + \frac{1}{2}H_{2} = HF. International Journal of Quantum Chemistry, 18(6): 1431–1448, December 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Kubo:1982:FRR

[Kub82] Ryogo Kubo. Fluctuation, response, and relaxation: The linear response theory revisited. International Journal of Quantum Chemistry, 22(S16):25–32, March 1–13, 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Chemistry, Theory of Condensed Matter, and Propagator Methods in the Quantum Theory of Matter.

Kummel:1983:EVD

[Küm83] H. G. Kümmel. Expectation values and density matrices in the Coupled-Cluster theory. International Journal of Quantum Chemistry, 24(1):79–84, July 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Kumar:1987:BRB

[Kum87] Pradeep Kumar. Book review: Quantum theory of collective phenomena. By G. L. Sewell, Oxford University Press, 1986. International Journal of Quantum Chemistry, 32(3):399, September 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). See [Sew86].

Kunz:1981:PNB

[Kun81] A. Barry Kunz. Properties of narrow band insulators. International Journal of Quantum Chemistry, 20(S15):487–498, March 8–14, 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory, Collision Phenomena, and Computational Quantum Chemistry.

Kunz:1983:STS

[Kun83] A. Barry Kunz. A series of tests of small- and mediumscale computers commonly used for computations by solidstate theorists and quantum chemists. International Journal of Quantum Chemistry, 24(S17):623–627, March 3–12, 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-state Theory, Collision Phenomena, and Computational Quantum Chemistry.

Kuntz:1986:FIA

[Kun86] P. J. Kuntz. The formation of Na^{+} and Na ions from Na*(^{2}P) atoms near a tungsten surface. II. Extended diatomics-in-molecules models. International Journal of Quantum Chemistry, 29(5):1105–1116, May 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Kutzelnigg:1980:NDP

[Kut80] Werner Kutzelnigg. New derivation and a k-particle generalization of SCF-type theories. International Journal of Quantum Chemistry, 18(1):3–9, July 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Kutzelnigg:1984:BSE

[Kut84] Werner Kutzelnigg. Basis set expansion of the Dirac operator without variational collapse. International Journal of Quantum Chemistry, 25(1):107–129, January 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Kvasnicka:1982:PGC

[Kva82] V. Kvasnička. Permutation-group classification and description of the Feynman–Goldstone diagrams. International Journal of Quantum Chemistry, 21(6):1003–1017, June 1982. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Kvasnicka:1983:QCC

[Kva83] Vladimír Kvasnička. Quasidegenerate coupled-cluster approach with Hermitian model Hamiltonian. International Journal of Quantum Chemistry, 24(4):335–344, October 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Kventsel:1982:PMT

[Kve82] G. F. Kventsel. Peierls- and Mott-type instabilities in onedimensional chains—coexistence or contradiction. International Journal of Quantum Chemistry, 22(4):825–835, October 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Kaliannan:1985:AEP

[KVR85] P. Kaliannan, S. Vishveshwara, and V. S. R. Rao. Anomeric effect in phosphate moieties of carbohydrates: An Ab initio study. International Journal of Quantum Chemistry, 27(2): 181–194, February 1985. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Kunz:1980:TES

[KWBC80] A. Barry Kunz, R. S. Weidman, J. Boettger, and G. Cochran. Theory of electronic states in narrow-band materials: Applications to solids, argon, CuCl, and CdS. International Journal of Quantum Chemistry, 18(S14):585–596, March 10–20, 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory, Collision Phenomena, Quantum Statistics, and Computational Methods.

Kurtz:1983:PPB

[KWJ83] Henry A. Kurtz, Brian Weiner, and Hans Jørgen Aa. Jensen. The polarization propagator based on an AGP state: Theory and application to the helium atom. International Journal of Quantum Chemistry, 24(S17):415–423, March 3–12, 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-state Theory, Collision Phenomena, and Computational Quantum Chemistry.

Kutzler:1986:SDP

[KWM86] F. W. Kutzler, C. T. White, and J. W. Mintmire. Soliton defects in polyacetylene: Local-density functional results. International Journal of Quantum Chemistry, 29(4):793–797, April 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Ladik:1983:TAC

[Lad83] János J. Ladik. Treatment of aperiodicity and correlation in polymers. International Journal of Quantum Chemistry, 23 (3):1073–1081, March 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Ladik:1986:ICC

[Lad86] János J. Ladik. The initiation of carcinogenesis in the cell. International Journal of Quantum Chemistry, 30(S13):307–310, March 6–8, 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Ladik:1987:ASG

[Lad87] János J. Ladik. Albert Szent-Györgyi’s impact on theoretical biophysics. International Journal of Quantum Chemistry, 32 (S14):3–7, March 14, 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Lafourcade:1981:ICB

[Laf81] L. Lafourcade. Importance of the concept of bond and binding in the physical chemistry of solids. International Journal of Quantum Chemistry, 19(6):991–1000, June 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

LaFemina:1989:UPA

[LaF89] John P. LaFemina. Ultraviolet photoelectron and absorption spectra of substituted phenols. International Journal of Quantum Chemistry, 36(5):563–573, November 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Lefebvre:1989:ABQ

[LAGS89] R. Lefebvre, O. Atabek, and M. Garcia-Sucre. Application of box quantization to the determination of the dissociation rates of a system of two coupled Morse oscillators. International Journal of Quantum Chemistry, 36(5):647–657, November 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Lala:1981:CPM

[Lal81] Anil K. Lala. Cholesterol and phospholipids in membranes: The hydrogen bonding problem. International Journal of Quantum Chemistry, 20(1):93–97, July 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Landsberg:1984:TGP

[Lan84] P. T. Landsberg. Two general problems in quantum biology. International Journal of Quantum Chemistry, 26(S11):55–61, March 12–15, 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Lang:1986:A

[Lan86] E. A. Lang. Announcement. International Journal of Quantum Chemistry, 29(6):1847, June 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Larsson:1982:ETB

[Lar82] Sven Larsson. Electron transfer in biological systems. International Journal of Quantum Chemistry, 22(S9):385–397, March 4–6, 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Larson:1983:TDD

[Lar83] Everett G. Larson. Time-dependent density operator in the random impulse approximation: I. Connections with the “generalized sudden approximation” for photoionization and electron impact ionization; and II. Relevance to reduced density operators, particle–hole matrices, and propagators for the initial state. International Journal of Quantum Chemistry, 23 (2):639–652, February 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Larson:1986:SME

[Lar86a] Everett G. Larson. A solvable model with an extreme AGP ground state: Relationships among fermion pairs, pairons, and natural spin geminals. International Journal of Quantum Chemistry, 30(S20):95–107, March 10–15, 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Syposium on Atomic, Molecular, and Solid-State Theory, Scattering Problems, Many Body Phenomena, and Computational Quantum Chemistry.

Larsson:1986:ETC

[Lar86b] Sven Larsson. Electron transport in condensed matter. A novel approach using quantum chemical methods. International Journal of Quantum Chemistry, 30(1):31–49, July 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Laszlo:1982:CCR

[Lás82] I. László. CNDO/2 calculations of relaxation and reconstruction of diamond and silicon [111] surfaces. International Journal of Quantum Chemistry, 21(5):813–822, May 1982. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Laszlo:1985:OPD

[Lás85] I. László. One-parameter double-zeta atomic functions for the Hartree–Fock energies of helium isoelectronic sequence. International Journal of Quantum Chemistry, 27(5):559–565, May 1985. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Lauher:1988:WWS

[Lau88] Joseph W. Lauher. What is wrong with a square metal cluster? International Journal of Quantum Chemistry, 34(S22): 309–317, March 12–19, 1988. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Syposium on Quantum Chemistry, Solid-State Theory, and Computational Methods.

Lawley:1987:IMQ

[Law87] K. P. Lawley, editor. Ab initio methods in quantum chemistry, volume 67, 69 of Advances in chemical physics. John Wiley, New York, NY, USA, 1987. ISBN 0-471-90900-9 (v. 1), 0-471-90901-7 (v. 2). 566 (vol. 1), 588 (vol. 2) pp. LCCN QD453 .A27 vol. 67.

Les:1981:E

[LB81] A. Leś and I. Ortega Blake. Erratum. International Journal of Quantum Chemistry, 19(4):719–720, April 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Lipkowitz:1982:A

[LB82] K. B. Lipkowitz and D. B. Boyd. Announcement. International Journal of Quantum Chemistry, 22(6):1307, December 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Lee:1983:MMB

[LB83] Yoon S. Lee and Rodney J. Bartlett. A multireference manybody perturbation theory study of Be + H_{2} → BeH_{2}. International Journal of Quantum Chemistry, 24(S17):347–356, March 3–12, 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-state Theory, Collision Phenomena, and Computational Quantum Chemistry.

Last:1986:AFP

[LB86] I. Last and M. Baer. Analytical fitting of potential energy surfaces for the Hxy systems. International Journal of Quantum Chemistry, 29(5):1067–1076, May 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Landman:1987:ELC

[LBC^{+}87] Uzi Landman, R. N. Barnett, C. L. Cleveland, Dafna Scharf, and Joshua Jortner. Electron localization in clusters. International Journal of Quantum Chemistry, 32(S21):573–587, March 12, 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Chemistry, Solid-State Theory, and Computational Methods.

Larsson:1988:LDE

[LBKV88] Sven Larsson, Anders Broo, Bruno Källebring, and Andrey Volosov. Long distance electron transfer. International Journal of Quantum Chemistry, 34(S15):1–22, March 12–19, 1988. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Syposium on Quantum Biology and Quantum Pharmacology.

Langhoff:1984:VLI

[LBP84] Stephen R. Langhoff, Charles W. Bauschlicher, Jr., and Harry Partridge. On the validity of the Landé interval rule in the alkaline earth atoms. International Journal of Quantum Chemistry, 26(S18):457–466, March 1–15, 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Atomic, Molecular and Solid-State Theory, and Computational Quantum Chemistry.

Lledos:1986:WCI

[LBV86] A. Lledós, J. Bertrán, and Oscar N. Ventura. Water-chain intervention in the ketonization of vinyl alcohol. An ab initio study. International Journal of Quantum Chemistry, 30(4): 467–477, October 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Lee:1980:CBS

[LC80] Tieh-Sheng Lee and San-Yan Chu. Comparison between the symmetry-correlation examination and the perturbation method. International Journal of Quantum Chemistry, 18(1): 79–87, July 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Lipinski:1981:GOS

[LC81] Józef Lipiński and Henryk Chojnacki. Ghost orbitals in semiempirical methods. Estimation of basis set superposition error. International Journal of Quantum Chemistry, 19 (5):891–900, May 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Larsson:1982:CSL

[LC82a] Sven Larsson and Richard Crossley. Calculation of the ^{4}S 4P^{0} spectrum of lithiumlike ions by the Hylleraas method. International Journal of Quantum Chemistry, 22(4):837–849, October 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Luken:1982:MFH

[LC82b] William L. Luken and John C. Culberson. Mobility of the Fermi hole in a single-determinant wavefunction. International Journal of Quantum Chemistry, 22(S16):265–276, March 1– 13, 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Chemistry, Theory of Condensed Matter, and Propagator Methods in the Quantum Theory of Matter.

Lee:1983:ACI

[LC83a] Tieh-Sheng Lee and San-Yan Chu. Avoided crossing and intermediate states in certain photochemical reactions. International Journal of Quantum Chemistry, 23(6):1923–1930, June 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Lee:1983:RCO

[LC83b] Tieh-Sheng Lee and San-Yan Chu. Reactivity comparison of one-electron and two-electron processes. International Journal of Quantum Chemistry, 23(2):625–631, February 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Ladik:1984:PPM

[LČ84] János Ladik and Jiří Čížek. Probable physical mechanisms of the activation of oncogenes through carcinogens. International Journal of Quantum Chemistry, 26(5):955–964, November 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Lima:1988:CDM

[LC88a] Eneida G. Lima and Sylvio Canuto. A comparison of different many-body perturbation theory calculations of the ground state of SiS. International Journal of Quantum Chemistry, 33 (5):395–401, May 1988. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Lima:1988:CCE

[LC88b] Eneida G. Lima and Sylvio Canuto. Correlated calculations of the electron affinity of HC_{2}. International Journal of Quantum Chemistry, 34(S22):199–205, March 12–19, 1988. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Syposium on Quantum Chemistry, Solid-State Theory, and Computational Methods.

Loew:1989:TIP

[LCA89] Gilda H. Loew, Jack R. Collins, and Frank U. Axe. Theoretical investigation of a proposed [HIS 175-ASP 235-TRP 191] proton transfer system in cytochrome c peroxidase. International Journal of Quantum Chemistry, 36(S16):199–209, April 1–8, 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Largo-Cabrerizo:1980:OCS

[LCC80] J. Largo-Cabrerizo and F. M. Duque Calvo. Orbital contributions to some two-electron properties. International Journal of Quantum Chemistry, 18(6):1465–1472, December 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Laaksonen:1984:TMS

[LCC84] Aatto Laaksonen, Giorgina Corongiu, and Enrico Clementi. Theoretical modeling of sodium chloride and related systems. I. The interaction energies. International Journal of Quantum Chemistry, 26(S18):131–140, March 1–15, 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Atomic, Molecular and Solid-State Theory, and Computational Quantum Chemistry.

Langlet:1981:IBN

[LCCB81] J. Langlet, P. Claverie, F. Caron, and J. C. Boeuve. Interactions between nucleic acid bases in hydrogen bonded and stacked configurations: The role of the molecular charge distribution. International Journal of Quantum Chemistry, 20 (2):299–338, August 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Loew:1987:SDD

[LCCW87] Gilda Loew, Jack Collins, Lek Chantranupong, and Ahmad Waleh. Spin density distribution in oxygen-liganded model heme proteins: Predictions of ^{17}O hyperfine broadening of ESR spectra of metmyoglobin, cytochrome C peroxidase, catalase, and cytochrome P450. International Journal of Quantum Chemistry, 32(S14):75–83, March 14, 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement:

Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Lippert:1982:QSE

[LCD82] E. Lippert and C. A. Chatzidimitriou-Dreismann. A quantum statistical estimation of the amount of energy dissipated by molecular relaxation processes in liquids. International Journal of Quantum Chemistry, 22(S16):183–194, March 1– 13, 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Chemistry, Theory of Condensed Matter, and Propagator Methods in the Quantum Theory of Matter.

Lavery:1980:MEP

[LCDP80] Richard Lavery, Didier Cauchy, Olivia De La Luz Rojas, and Alberte Pullman. Molecular electrostatic potential of the B-DNA helix. VII. effect of screening by monovalent cations. International Journal of Quantum Chemistry, 18(S7):323–330, March 5–8, 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Largo-Cabrerizo:1989:TSP

[LCF89] A. Largo-Cabrerizo and J. R. Flores. A theoretical study of protonation of triatomic silicon–carbon compounds. International Journal of Quantum Chemistry, 36(3):241–253, September 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Lowdin:1982:DFT

[LCG82a] Per-Olov Löwdin, Jean-Louis Calais, and Osvaldo Goscinski. Density functional theory. International Journal of Quantum Chemistry, 21(2):357, February 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Lowdin:1982:P

[LCG82b] Per-Olov Löwdin, Jean-Louis Calais, and Osvaldo Goscinski. Preface. International Journal of Quantum Chemistry, 21(1): 215, January 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Lowdin:1983:P

[LCG83] Pre-Olov Löwdin, Jean-Louis Calais, and Osvaldo Goscinski. Preface. International Journal of Quantum Chemistry, 23(1): v, January 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Largo-Cabrerizo:1986:CEL

[LCLC86] A. Largo-Cabrerizo and J. Largo-Cabrerizo. The correlation effects at a local level. CI partitions. International Journal of Quantum Chemistry, 30(6):845–852, December 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Largo-Cabrerizo:1988:CEL

[LCLC88] A. Largo-Cabrerizo and J. Largo-Cabrerizo. Correlation effects at a local level and CI partitions: 2^{3} S State of helium. International Journal of Quantum Chemistry, 33(4):323–326, April 1988. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Lowdin:1984:EN

[LCÖ84] Per-Olov Löwdin, Jean-Louis Calais, and Yngve Öhrn. Editor’ Note. International Journal of Quantum Chemistry, 26(4):i, October 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Lavery:1981:EMP

[LCP81] Richard Lavery, Sylvie Corbin, and Alberte Pullman. The electrostatic molecular potential of tRNA^{P he}. V. The influence of counterion binding on the potential and the steric accessibility. International Journal of Quantum Chemistry, 20(S8):171–183, March 5–7, 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Loew:1988:ECS

[LCP^{+}88] Gilda Loew, Jack Collins, Philip Payne, Amrit K. Judd, and Kevin H. Wacknow. Energy-conformational studies of β-endorphins: Identification of plausible folded conformers. International Journal of Quantum Chemistry, 34(S15):55–66, March 12–19, 1988. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Syposium on Quantum Biology and Quantum Pharmacology.

Largo-Cabrerizo:1987:HCI

[LCULC87] A. Largo-Cabrerizo, C. Urdaneta, G. C. Lie, and E. Clementi. The Hylleraas–CI integrals in molecular, calculations. II. Three-and four-electron integrals and tests for two-electron many-center integrals. International Journal of Quantum Chemistry, 32(S21):677–692, March 12, 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Chemistry, Solid-State Theory, and Computational Methods.

Larsson:1980:ICE

[LdMH80] S. Larsson, P. Corrêa de Mello, and M. Hehenberger. Ab initio calculation of ESCA satellite intensities for a copper–chlorine complex. International Journal of Quantum Chemistry, 17 (4):719–724, April 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

LiberatoDeBrito:1980:TUI

[LE80] Adelsindo Liberato De Brito and Tomas P. Eggarter. Two useful identities involving spherical harmonics. International Journal of Quantum Chemistry, 17(3):459–465, March 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Lee:1983:SSS

[Lee83] D. M. Lee. Spectroscopy with sound on superfluid ^{3}He B. International Journal of Quantum Chemistry, 23(4):1191–1199, April 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Lefebvre:1983:SBS

[Lef83] R. Lefebvre. Symmetry broken self-consistent field approximation for coupled oscillators. International Journal of Quantum Chemistry, 23(2):543–550, February 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Leroy:1983:SCS

[Ler83] G. Leroy. Stability of chemical species. International Journal of Quantum Chemistry, 23(1):271–308, January 1983. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Leroy:1985:BRB

[Ler85] Georges Leroy. Book review: Information theoretic indices for characterization of chemical structures. By Danail Bonchev, Wiley, New York, 1983. International Journal of Quantum Chemistry, 27(1):103, January 1985. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). See [Bon83].

Lam:1987:GFA

[LER87] Miu-To Brenda Lam, Stephen T. Elbert, and Klaus Ruedenberg. Generation of a full active configuration space basis in terms of symmetry- and spin-adapted antisymmetrized orbital products. International Journal of Quantum Chemistry, 31 (3):489–505, March 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Leubner:1983:CFA

[Leu83] C. Leubner. The canonical function approach and diatomic molecules. International Journal of Quantum Chemistry, 24 (1):127–130, July 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Levy:1989:ACS

[Lev89] Mel Levy. Asymptotic coordinate scaling bound for exchangecorrelation energy in density-functional theory. International Journal of Quantum Chemistry, 36(S23):617–619, April 1–8, 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Lewis:1988:MOC

[Lew88] David F. V. Lewis. Molecular orbital calculations on tumorinhibitory nitrosoureas: QSARs. International Journal of Quantum Chemistry, 33(4):305–321, April 1988. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Leite:1981:VCM

[LFL^{+}81] José R. Leite, Adalberto Fazzio, Marco A. P. Lima, A. M. Dias, Abel Rosato, and E. Segre. The variational cellular method for quantum mechanical applications: Calculations of the ground and excited states of F_{2} and Ne_{2} molecules. International Journal of Quantum Chemistry, 20(S15):401–408, March 8–14, 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory, Collision Phenomena, and Computational Quantum Chemistry.

Lima:1989:MST

[LFM89a] Gilson A. R. Lima, Adalberto Fazzio, and Ronaldo Mota. Metal-semiconductor transition in cerium hydrides. International Journal of Quantum Chemistry, 36(S23):709–716, April 1–8, 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Lowdin:1989:SPB

[LFM89b] Per-Olov Löwdin, Piotr Froelich, and Manoj Mishra. Some properties of the bivariational Hartree–Fock scheme for complex symmetric many-particle operators. International Journal of Quantum Chemistry, 36(2):93–103, August 1989. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Les:1984:HGG

[LG84] Andrzej Leś and Sławomir Grabowski. Hydration of GABA and of the GABA·ZN^{2+} complex: A Monte Carlo simulation study. International Journal of Quantum Chemistry, 26(2): 167–181, August 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Lowdin:1989:PIW

[LGJLCB89] Per-Olov Löwdin, Osvaldo Goscinski, Jean-Louis-Calais, and Erkki Brändas. Professor Ivar Waller 90-years anniversary June 11, 1988. International Journal of Quantum Chemistry, 35(6):703, June 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Luke:1984:TIR

[LGL^{+}84] Brian T. Luke, Abha G. Gupta, Gilda H. Loew, James G. Lawless, and David H. White. Theoretical investigation of the role of clay edges in prebiotic peptide bond formation. I. Structures of acetic acid, glycine, H_{2}SO_{4}, H_{3}PO_{4}, Si(OH)_{4}, and Al(OH). International Journal of Quantum Chemistry, 26 (S11):117–135, March 12–15, 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Langlet:1980:PWI

[LGPP^{+}80] J. Langlet, C. Giessner-Prettre, B. Pullman, P. Claverie, and D. Piazzola. Purine–water interactions in base stacking. International Journal of Quantum Chemistry, 18(2):421–437, August 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Lefebvre:1987:CPD

[LGS87] R. Lefebvre and M. Garcia-Sucre. Complex perturbation and deperturbation procedures applied to photofragmentation cross-section calculations. International Journal of Quantum Chemistry, 32(S21):377–387, March 12, 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Chemistry, Solid-State Theory, and Computational Methods.

Lien:1985:TSC

[LH85] Min H. Lien and Alan C. Hopkinson. A theoretical study of C_{4}H_{5}N isomers: Comparison of cyano and isocyano as substituents. International Journal of Quantum Chemistry, 27 (1):1–10, January 1985. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Louie:1987:TQE

[LH87] Steven G. Louie and Mark S. Hybertsen. Theory of quasiparticle energies: Band gaps and excitation spectra in solids. International Journal of Quantum Chemistry, 32(S21):31–44, March 12, 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Chemistry, Solid-State Theory, and Computational Methods.

Larsson:1980:IHF

[LHC80] Sven Larsson, Michael Hehenberger, and Paulo Corrêa De Mello. Ab initio Hartree–Fock and multiple-scattering Xα calculation of the g factors for CuF_{2}. International Journal of Quantum Chemistry, 18(5):1271–1278, November 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Lundqvist:1983:TSM

[LHH^{+}83] B. I. Lundqvist, B. Hellsing, S. Holmström, P. Nordlander, M. Persson, and J. K. Nørskov. Theoretical studies of molecular adsorption on metal surfaces. International Journal of Quantum Chemistry, 23(3):1083–1090, March 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Lowrey:1981:MCN

[LHK81] A. H. Lowrey, P. C. Hariharan, and Joyce J. Kaufman. Molecular calculations with the nonempirical ab-initio MODPOT /VRDDO /MERGE procedures. XIV. 2,6-dimethyl-N-nitroso morpholine and its α-OH isomers: Conformations and electrostatic molecular potential contour maps. International Journal of Quantum Chemistry, 20(S8):149–160, March 5–7, 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Lewchenko:1982:QCC

[LHKK82] Victor Lewchenko, P. C. Hariharan, Walter S. Koski, and Joyce J. Kaufman. Quantum chemical calculations for understanding and predicting toxicity. II. The phosphorylation step in the inhibition of ache by organophosphorus anticholinesterases. International Journal of Quantum Chemistry, 22(S9):275–280, March 4–6, 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Lohr:1980:RPE

[LHP80] L. L. Lohr, Jr., M. Hotokka, and P. Pyykkö. Relativistically parametrized extended-Hückel calculations. II. Orbital energies of group-IV tetrahalides and tetramethyls. International Journal of Quantum Chemistry, 18(2):347–355, August 1980.

CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Loew:1980:COS

[LHZ80] Gilda H. Loew, Zelek S. Herman, and Michael C. Zerner. Calculated optical spectrum of model oxyheme complex. International Journal of Quantum Chemistry, 18(2):481–492, August 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Liberman:1984:USO

[Lib84] David A. Liberman. The use of scaled orbital functions in selfconsistent field calculations for atoms. International Journal of Quantum Chemistry, 26(S18):359–368, March 1–15, 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Atomic, Molecular and Solid-State Theory, and Computational Quantum Chemistry.

Lieb:1983:DFC

[Lie83] Elliott H. Lieb. Density functionals for Coulomb systems. International Journal of Quantum Chemistry, 24(3):243–277, September 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Lim:1980:SSE

[Lim80] T. K. Lim. Solving the Schrödinger equation without integration. International Journal of Quantum Chemistry, 18 (S14):263–268, March 10–20, 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory, Collision Phenomena, Quantum Statistics, and Computational Methods.

Lim:1982:EFE

[Lim82] T. K. Lim. Excess free energy changes for binary mixed electrolytes at vanishing ionic strengths. International Journal of Quantum Chemistry, 22(S16):247–257, March 1–13, 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Chemistry, Theory of Condensed Matter, and Propagator Methods in the Quantum Theory of Matter.

Lim:1986:DSL

[Lim86] T. K. Lim. Double symmetrization and the linearly independent spin eigenfunctions. International Journal of Quantum Chemistry, 29(6):1687–1695, June 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Lim:1987:OWH

[Lim87] Tiong-Koon Lim. Orthogonal Waller–Hartree spin eigenfunctions. International Journal of Quantum Chemistry, 31(2): 187–194, February 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Lim:1988:RDD

[Lim88] Tiong-Koon Lim. On the resolution of a double determinant. International Journal of Quantum Chemistry, 33(6):483–489, June 1988. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Lim:1989:EFH

[Lim89] Tiong-Koon Lim. Explicit formulas for the Hamiltonian matrix elements. International Journal of Quantum Chemistry, 35(4):551–566, April 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Linderberg:1981:ARR

[Lin81a] Jan Linderberg. Algebraic reductions of the real symmetric 4 × 4 secular problem. International Journal of Quantum Chemistry, 19(2):237–249, February 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Linderberg:1981:MSM

[Lin81b] Jan Linderberg. Multiple surface molecular motion: A tmatrix approach. International Journal of Quantum Chemistry, 20(S15):559–581, March 8–14, 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory, Collision Phenomena, and Computational Quantum Chemistry.

Linderberg:1985:BCC

[Lin85] Jan Linderberg. Basis for coupled channel approach to reactive scattering. International Journal of Quantum Chemistry, 28(S19):467–476, March 18–23, 1985. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Linderberg:1989:AMC

[Lin89] Jan Linderberg. An algorithm for R matrix calculations for atom–diatom reactive scattering. International Journal of Quantum Chemistry, 35(6):801–811, June 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Lipinski:1983:ASW

[Lip83] Józef Lipiński. On the artificial structure of the water dimer in the CNDO/2 method. International Journal of Quantum Chemistry, 24(1):131–133, July 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Lipinski:1988:MAV

[Lip88] Jóazef Lipiński. Modified all-valence INDO/spd method for ground and excited state properties of isolated molecules and molecular complexes. International Journal of Quantum Chemistry, 34(5):423–435, November 1988. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Little:1981:CDE

[Lit81] W. A. Little. Criteria for the design of an excitonic superconductor. International Journal of Quantum Chemistry, 20 (S15):545–554, March 8–14, 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory, Collision Phenomena, and Computational Quantum Chemistry.

Lim:1982:WHC

[LJ82] T. K. Lim and C. G. Jesudason. Waller–Hartree CI calculations. International Journal of Quantum Chemistry, 22(S16): 259–264, March 1–13, 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Chemistry, Theory of Condensed Matter, and Propagator Methods in the Quantum Theory of Matter.

Langhoff:1983:UEC

[LJC83] Stephen R. Langhoff, Richard L. Jaffe, and Delano P. Chong. Use of effective core potentials in perturbation corrections to the Koopmans theorem: Vertical ionization potentials of Cl_{2}, ClN_{3}, and ClNCO. International Journal of Quantum Chemistry, 23(3):875–886, March 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Luan:1989:SEE

[LJHU89] Chi-Hao Luan, John Jaggard, R. Dean Harris, and Dan W. Urry. On the source of entropic elastomeric force in polypeptides and proteins: Backbone configurational vs. side-chain solvational entropy. International Journal of Quantum Chemistry, 36(S16):235–244, April 1–8, 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Laskowski:1986:ICP

[LJK86] Bernard C. Laskowski, Richard L. Jaffe, and Andrew Komornicki. Ab initio calculations of phenylene ring motions in polyphenylene oxide. International Journal of Quantum Chemistry, 29(3):563–578, March 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

LHuillier:1987:MIM

[LJW87] Anne L’Huillier, Lars Jönsson, and Göran Wendin. Multiphoton ionization of many-electron atoms. International Journal of Quantum Chemistry, 31(5):833–840, May 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Laforgue-Kantzer:1981:SIN

[LK81] D. Laforgue-Kantzer. State of ions near a cathode: Adions or adatoms? International Journal of Quantum Chemistry, 19 (6):1031–1041, June 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Lee:1987:EAF

[LK87] Yong S. Lee and Miklos Kertesz. The effect of additional fused rings on the stabilities and the band gaps of heteroconjugated polymers. International Journal of Quantum Chemistry, 32 (S21):163–170, March 12, 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Chemistry, Solid-State Theory, and Computational Methods.

Ley-Koo:1989:UTT

[LKB89] E. Ley-Koo and Carlos F. Bunge. Use of truncated Taylor series for r_{12}^{−1} in electronic structure calculations. International Journal of Quantum Chemistry, 36(S23):263–268, April 1–8, 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Laurenzi:1982:IME

[LL82a] B. J. Laurenzi and C. Litto. Isoelectronic molecules: The 13-, 14-, and 22- electron diatomic sequences. International Journal of Quantum Chemistry, 22(S16):215–226, March 1– 13, 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Chemistry, Theory of Condensed Matter, and Propagator Methods in the Quantum Theory of Matter.

Lipinski:1982:DTF

[LL82b] Józef Lipiński and Jerzy Leszczyński. Derivation of theoretical formulas for resonance integrals from Heisenberg equation of motion. International Journal of Quantum Chemistry, 22(2): 253–263, August 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Leonard:1984:CLM

[LL84] Joseph M. Leonard and William L. Luken. Calculation of localized molecular orbitals. International Journal of Quantum Chemistry, 25(2):355–365, February 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Luke:1985:TIF

[LL85] Brian T. Luke and Gilda H. Loew. A theoretical investigation of the first step in the metabolic reduction of halogenated methanes by cytochrome P-450. International Journal of Quantum Chemistry, 28(S12):99–112, March 14–16, 1985. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Lafemina:1986:DLG

[LL86] John P. Lafemina and John P. Lowe. Degeneracy lifting in graphite and boron nitride: a perturbational approach. International Journal of Quantum Chemistry, 30(6):769–782, December 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Liang:1989:TSC

[LL89] Jiin-Yun Liang and William N. Lipscomb. Theoretical study of carbonic anhydrase-catalyzed hydration of CO_{2}: a brief review. International Journal of Quantum Chemistry, 36(3): 299–312, September 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Lino:1989:ESM

[LLAG89] A. T. Lino, J. R. Leite, L. V. C. Assali, and V. M. S. Gomes. Electronic states of Mn_{4} complex cluster in silicon. International Journal of Quantum Chemistry, 36(S23):701–708, April 1–8, 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Luken:1983:ASP

[LLC83] William L. Luken, Joseph M. Leonard, and John C. Culberson. Application of the saddle-point variational method to self-consistent-field wave functions. International Journal of Quantum Chemistry, 24(S17):401–405, March 3–12, 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-state Theory, Collision Phenomena, and Computational Quantum Chemistry.

Luke:1986:TES

[LLM86] Brian T. Luke, Gilda H. Loew, and A. D. McLean. A theoretical examination of substituent effects on the detoxification reaction between glutathione and halogenated methanes. International Journal of Quantum Chemistry, 29(4):883–896, April 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Laskowski:1983:TDX

[LLS83] Bernard C. Laskowski, Stephen R. Langhoff, and Per E. M. Siegbahn. Theoretical determination of the X^{2}Σ^{+} and A^{2}Π potentials of CsO using relativistic effective core potentials. International Journal of Quantum Chemistry, 23(2):483–490, February 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Larsson:1981:BRB

[LM81] Seven Larsson and Olle Mårtensson. Book review: Catalysis in chemistry and biochemistry. Theory and experiment. Jerusalem symposia on quantum chemistry and biochemistry. Vol. 12. Edited by B. Pullman. Reidel, Dordrecht, 1979. Price (cloth): $85.00. ISBN 90-277-1039-2. International Journal of Quantum Chemistry, 20(6):1331, December 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). See [Pul79].

Labbe:1982:ATS

[LM82a] A. Toro Labbe and J. Maruani. Application of the theory of symmetry of nonrigid molecules to the calculation of the conformational dependences of the torsional potential and dipole moment of acetone-related molecules. International Journal of Quantum Chemistry, 22(1):115–125, July 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Ludena:1982:VFH

[LM82b] Eduardo V. Ludeña and Vladimiro Mujica. Virial fragments and the Hohenberg–Kohn functional. International Journal of Quantum Chemistry, 21(5):927–935, May 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Laurenzi:1986:CPI

[LM86] Bernard J. Laurenzi and Madeline F. Mall. Critical phenomena in isoelectronic diatomic sequences—the 14-electron sequence. International Journal of Quantum Chemistry, 30(1): 51–75, July 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Lea:1989:QMW

[LM89] M. J. Lea and N. H. March. Quantum-mechanical Wigner electron crystallization with and without magnetic fields. International Journal of Quantum Chemistry, 36(S23):717–729, April 1–8, 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Lesyng:1985:ICB

[LMG85] Bogdan Lesyng, Christian Marck, and Wilhelm Guschlbauer. Ab initio calculations on the barrier to pseudorotation of model 2′-deoxyfuranose and 2′-deoxy-2′-fluorofuranose rings. International Journal of Quantum Chemistry, 28(4):517–523, October 1985. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Le-min:1983:ISE

[LmJQHXZ83] Li Le-min, Ren Jing-Qing, Xu Guang-Xian (K. H. Hsu), and Wang Xiu-Zhen. INDO Studies on the electronic structure of lanthanoid compounds. International Journal of Quantum Chemistry, 23(4):1305–1316, April 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Lundstrom:1983:EGE

[LN83] Ingemar Lundström and Claes Nylander. Electrostatics and gating of excitable membranes. International Journal of Quantum Chemistry, 23(4):1269–1282, April 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Lowdin:1989:SRP

[LN89a] Per-Olov Löwdin and Bengt Nagel. Some remarks on the perturbation and variation–perturbation calculations of the free energy for quantum systems. International Journal of Quantum Chemistry, 35(6):839–850, June 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Lunell:1989:CSS

[LN89b] S. Lunell and Y. Nakamura. Conformational studies on Lserine phosphate and hydrated L-serine phosphate using ab initio SCF calculations. International Journal of Quantum Chemistry, 35(3):395–407, March 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Logan:1984:FGE

[LNN84] J. Logan, M. D. Newton, and J. O. Noell. Factors governing electronic localization in transition metal clusters and complexes. International Journal of Quantum Chemistry, 26(S18): 213–235, March 1–15, 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Atomic, Molecular and Solid-State Theory, and Computational Quantum Chemistry.

Lieb:1981:ILB

[LO81] Elliott H. Lieb and Stephen Oxford. Improved lower bound on the indirect Coulomb energy. International Journal of Quantum Chemistry, 19(3):427–439, March 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Lowdin:1982:I

[LÖ82] Per-Olov Löwdin and Yngve Öhrn. Introduction. International Journal of Quantum Chemistry, 21(1):1–2, January 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Linderberg:1985:KEF

[LÖ85a] Jan Linderberg and Yngve Öhrn. Kinetic energy functional in hyperspherical coordinates. International Journal of Quantum Chemistry, 27(3):273–280, March 1985. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Lunell:1985:HFC

[LÖ85b] Sten Lunell and Per Öster. Hartree–Fock calculations on quartet s, p, d, and f states of boron III. International Journal of Quantum Chemistry, 27(4):475–479, April 1985. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Les:1985:TST

[LOB85] A. Leś and I. Ortega-Blake. A theoretical study of tautomerism: 2- and 4-oxopyrimidine and some of their derivatives. International Journal of Quantum Chemistry, 27(5): 567–583, May 1985. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Les:1986:TUC

[LOB86] Andrzej Leś and Iván Ortega-Blake. Tautomerism of uracil, cytosine, isocytosine, and some of their thio-derivatives. International Journal of Quantum Chemistry, 30(2):225–237, August 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Ladik:1986:QMT

[LOBS86] J. Ladik, P. Otto, A. K. Bakhshi, and M. Seel. Quantum mechanical treatment of biopolymers as solids: Possible implications for carcinogenesis. International Journal of Quantum Chemistry, 29(4):597–617, April 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Ladik:1983:SNR

[LOF83] J. Ladik, P. Otto, and W. Förner. Some new results in the quantum mechanical investigation of DNA. International Journal of Quantum Chemistry, 24(S10):73–80, March 14–16, 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Leite:1986:BCQ

[LOGC86] José R. Leite, Glaucia M. G. Oliveira, Vivilí M. S. Gomes, and Alaor S. Chaves. The behavior of carriers in quantum wells in GaAs-A_{x}Ga_{1−x}As superlattices under in-plane magnetic fields. International Journal of Quantum Chemistry, 30 (S20):335–346, March 10–15, 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Syposium on Atomic, Molecular, and Solid-State Theory, Scattering Problems, Many Body Phenomena, and Computational Quantum Chemistry.

Lohr:1984:EEE

[Loh84] Lawrence L. Lohr, Jr. Electronegativity equalization and the electronic structure of polyhedral clusters of main-group atoms. International Journal of Quantum Chemistry, 25(1): 211–221, January 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Lohr:1985:PCE

[Loh85] Lawrence L. Lohr. The protonic counterpart of electronegativity. International Journal of Quantum Chemistry, 28(S19): 731–732, March 18–23, 1985. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Lohr:1987:CDM

[Loh87] Lawrence L. Lohr. Centrifugal distortions in molecules: An ab initio approach with application to water. International Journal of Quantum Chemistry, 32(S21):407–415, March 12, 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Chemistry, Solid-State Theory, and Computational Methods.

Larsson:1984:ESP

[LOR84] Sven Larsson, Lars-Fride Olsson, and Arne Rosén. Electronic structure of the PdCl_{4}^{2−} and PtCl_{4}^{2−} ions. International Journal of Quantum Chemistry, 25(1):201–209, January 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Lowdin:1985:Ia

[LÖSZ85a] P. O. Löwdin, N. Y. Öhrn, J. R. Sabin, and M. C. Zerner. Introduction. International Journal of Quantum Chemistry, 28 (S12):v, March 14–16, 1985. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Lowdin:1985:Ib

[LÖSZ85b] P. O. Löwdin, N. Y. Öhrn, J. R. Sabin, and M. C. Zerner. Introduction. International Journal of Quantum Chemistry, 28(S19):1–2, March 18–23, 1985. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Lowdin:1986:Ia

[LÖSZ86a] P. O. Löwdin, N. Y. Öhrn, J. R. Sabin, and M. C. Zerner. Introduction. International Journal of Quantum Chemistry, 30 (S13):v, March 6–8, 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Lowdin:1986:Ib

[LÖSZ86b] P. O. Löwdin, N. Y. Öhrn, J. R. Sabin, and M. C. Zerner. Introduction. International Journal of Quantum Chemistry, 30(S20):1–2, March 10–15, 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Syposium on Atomic, Molecular, and Solid-State Theory, Scattering Problems, Many Body Phenomena, and Computational Quantum Chemistry.

Lowdin:1987:I

[LÖSZ87] P. O. Löwdin, N. Y. Öhrn, J. R. Sabin, and M. C. Zerner. Introduction. International Journal of Quantum Chemistry, 32(S21):1, March 12, 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Chemistry, Solid-State Theory, and Computational Methods.

Lowdin:1988:Ia

[LÖSZ88a] P. O. Löwdin, N. Y. Öhrn, J. R. Sabin, and M. C. Zerner. Introduction. International Journal of Quantum Chemistry, 34 (S15):v, March 12–19, 1988. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Syposium on Quantum Biology and Quantum Pharmacology.

Lowdin:1988:Ib

[LÖSZ88b] P. O. Löwdin, N. Y. Öhrn, J. R. Sabin, and M. C. Zerner. Introduction. International Journal of Quantum Chemistry, 34 (S22):ix, March 12–19, 1988. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Syposium on Quantum Chemistry, Solid-State Theory, and Computational Methods.

Lowdin:1980:P

[Löw80] Per-Olov Löwdin. Preface. International Journal of Quantum Chemistry, 18(1):1, July 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Lowdin:1981:AQT

[Löw81a] Per-Olov Löwdin. Addendum to “Quantum theory as a trace algebra”. International Journal of Quantum Chemistry, 20(3): 775–776, September 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Lowdin:1981:P

[Löw81b] Per-Olov Löwdin. Preface. International Journal of Quantum Chemistry, 19(6):973, June 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Lowdin:1982:OSH

[Löw82a] Per-Olov Löwdin. On operators, superoperators, Hamiltonians, and Liouvillians. International Journal of Quantum Chemistry, 22(S16):485–560, March 1–13, 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Chemistry, Theory of Condensed Matter, and Propagator Methods in the Quantum Theory of Matter.

Lowdin:1982:AQT

[Löw82b] Per-Olov Löwdin. On the axioms of quantum theory formulated as a trace algebra. International Journal of Quantum Chemistry, 21(1):275–353, January 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Lowdin:1982:PTP

[Löw82c] Per-Olov Löwdin. Partitioning technique, perturbation theory, and rational approximations. International Journal of Quantum Chemistry, 21(1):69–92, January 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Lowe:1982:TAK

[Low82d] John P. Lowe. Theoretical analysis of K-region and bay-region indicators of carcinogenicity. International Journal of Quantum Chemistry, 22(6):1221–1247, December 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Lowe:1982:TSR

[Low82e] John P. Lowe. Theoretical study of relations between bayregion and K-region indices of carcinogenicity. International Journal of Quantum Chemistry, 22(S9):5–13, March 4–6, 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Lowdin:1983:SCE

[Löw83] Per-Olov Löwdin. On the Sanibel coefficients in the expansion of spin-projected Slater determinants. International Journal of Quantum Chemistry, 24(6):729–745, December 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Lowdin:1984:SAR

[Löw84] Per-Olov Löwdin. Some aspects on the relation between the natural sciences and the human mind: An introduction to a panel discussion on the “origin of life and mind”. International Journal of Quantum Chemistry, 26(S11):31–43, March 12–15, 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Lowdin:1985:ACL

[Löw85a] Per-Olov Löwdin. Approximate calculation of lifetimes of resonance states in the continuum from real stabilization graphs. International Journal of Quantum Chemistry, 27(4):495–500, April 1985. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Lowdin:1985:TFY

[Löw85b] Per-Olov Löwdin. Twenty-five years of Sanibel Symposia: A brief historic and scientific survey. International Journal of Quantum Chemistry, 28(S19):19–37, March 18–23, 1985. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Lowdin:1986:A

[Löw86a] Per-Olov Löwdin. Announcement. International Journal of Quantum Chemistry, 29(2):vii–viii, February 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Lowdin:1986:CSM

[Löw86b] Per-Olov Löwdin. Comments on some models of carcinogenesis. International Journal of Quantum Chemistry, 30(S13): 311–317, March 6–8, 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Lowdin:1986:SAQ

[Löw86c] Per-Olov Löwdin. On the state of the art of quantum chemistry. International Journal of Quantum Chemistry, 29(5): 1651–1683, May 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Lowdin:1986:RAW

[Löw86d] Per-Olov Löwdin. Remarks about the wave-function representability of the first-order reduced density matrices. International Journal of Quantum Chemistry, 30(S20):769–771, March 10–15, 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Syposium on Atomic, Molecular, and Solid-State Theory, Scattering Problems, Many Body Phenomena, and Computational Quantum Chemistry.

Lowdin:1986:SCM

[Löw86e] Per-Olov Löwdin. Some comments on the method of complex scaling to find physical resonance states. International Journal of Quantum Chemistry, 30(S20):743–748, March 10–15, 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Syposium on Atomic, Molecular, and Solid-State Theory, Scattering Problems, Many Body Phenomena, and Computational Quantum Chemistry.

Lowdin:1988:STI

[Löw88] Per-Olov Löwdin. On some thermodynamical inequalities. International Journal of Quantum Chemistry, 34(S22):337–342, March 12–19, 1988. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Syposium on Quantum Chemistry, Solid-State Theory, and Computational Methods.

Lowdin:1989:PSQ

[Löw89] Per-Olov Löwdin. On the present situation of quantum chemistry: Chairman’s introductory remarks at the 6th ICQC. International Journal of Quantum Chemistry, 36(3):191–199, September 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Lumbroso:1980:IMO

[LP80] Henri Lumbroso and Giuseppe Concetto Pappalardo. Ab initio molecular orbital study of the hydrogen-bonded pyrrole · acetonitrile complex. International Journal of Quantum Chemistry, 18(1):89–99, July 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Lavery:1981:MEP

[LP81a] Richard Lavery and Bernard Pullman. Molecular electrostatic potential on the surface envelopes of macromolecules: B-DNA. International Journal of Quantum Chemistry, 20(1):259–272, July 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Loew:1981:QCS

[LP81b] Gilda H. Loew and Michael T. Poulsen. Quantum chemical studies of aflatoxins: Metabolism and carcinogenic activity. International Journal of Quantum Chemistry, 20(S8):95–107, March 5–7, 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Levy:1982:DHK

[LP82] M. Levy and J. P. Perdew. In defense of the Hohenberg–Kohn theorem and density functional theory. International Journal of Quantum Chemistry, 21(2):511–513, February 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Laurence:1984:SRP

[LP84] Patricia R. Laurence and Peter Politzer. Some reactive properties of chlorooxirane, a likely carcinogenic metabolite of vinyl chloride. International Journal of Quantum Chemistry, 25 (3):493–502, March 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Levy:1985:PEU

[LP85] Mel Levy and John P. Perdew. Properties of the exact universal density and one-matrix functionals. International Journal of Quantum Chemistry, 28(S19):743–744, March 18–23, 1985. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Luana:1987:CPE

[LP87] V. Luaña and L. Pueyo. Core projection effects in atomic frozen-core calculations: a numerical analysis. International Journal of Quantum Chemistry, 31(6):975–988, June 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Li:1989:ERG

[LP89] Xiangzhu Li and Josef Paldus. Explicit representation of Gel’fand–Tsetlin states in Clifford algebra unitary group approach. International Journal of Quantum Chemistry, 36(2): 127–140, August 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Laidig:1982:LOC

[LPB82] William D. Laidig, George D. Purvis III, and Rodney J. Bartlett. Localized orbitals in the coupled cluster singles and doubles model. International Journal of Quantum Chemistry, 22(S16):561–573, March 1–13, 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Chemistry, Theory of Condensed Matter, and Propagator Methods in the Quantum Theory of Matter.

Langhoff:1980:TSP

[LPC^{+}80] P. W. Langhoff, N. Padial, G. Csanak, T. N. Rescigno, and B. V. Mckoy. Theoretical studies of photoionization in diatomic and polyatomic molecules. International Journal of Quantum Chemistry, 18(S14):285–304, March 10–20, 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory, Collision Phenomena, Quantum Statistics, and Computational Methods.

Loew:1980:MSA

[LPCF80] Gilda H. Loew, Andrew T. Pudzianowski, Alfredo Czerwinski, and James E. Ferrell, Jr. Mechanistic studies of addition of nucleophiles to arene oxides and diol epoxides: Candidate ultimate carcinogens. International Journal of Quantum Chemistry, 18(S7):223–244, March 5–8, 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Lavery:1981:SAR

[LPP81] Richard Lavery, Alberte Pullman, and Bernard Pullman. Steric accessibility of reactive centers in B-DNA. International Journal of Quantum Chemistry, 20(1):49–62, July 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Laurence:1984:RPT

[LPP84] Patricia R. Laurence, Timothy R. Proctor, and Peter Politzer. Reactive properties of trans-dichlorooxirane in relation to the contrasting carcinogenicities of vinyl chloride and transdichloroethylen. International Journal of Quantum Chemistry, 26(3):425–438, September 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Laaksonen:1983:TDFa

[LPS83a] Leif Laaksonen, Pekka Pyykkö, and Dage Sundholm. Twodimensional fully numerical solutions of molecular Schrödinger equations. I. One-electron molecules. International Journal of Quantum Chemistry, 23(1):309–317, January 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Laaksonen:1983:TDFb

[LPS83b] Leif Laaksonen, Pekka Pyykkö, and Dage Sundholm. Twodimensional fully numerical solutions of molecular Schrödinger equations. II. Solution of the Poisson equation and results for singlet states of H_{2} and HeH^{+}. International Journal of Quantum Chemistry, 23(1):319–323, January 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Loew:1983:MSA

[LPSK83] Gilda H. Loew, Michael T. Poulsen, Dale Spangler, and Elizabeth Kirkjian. Mechanistic structure–activity studies of carcinogenic dialkylnitrosamines. International Journal of Quantum Chemistry, 24(S10):201–213, March 14–16, 1983. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Lochmann:1981:MEP

[LPW81] R. Lochmann, G. Peinel, and T. Weller. Molecular electrostatic potentials from PCILO wave functions. International Journal of Quantum Chemistry, 20(5):1097–1105, November 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Lavery:1983:EPP

[LPW83] R. Lavery, A. Pullman, and Y. K. Wen. On the electrostatic properties of papain in relation to its enzymatic activity. International Journal of Quantum Chemistry, 24(4):353–371, October 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Lindh:1989:TSD

[LR89] Roland Lindh and Björn O. Roos. A theoretical study of the diffuseness of the V (^{1}B_{1u}) state of planar ethylene. International Journal of Quantum Chemistry, 35(6):813–825, June 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Larson:1988:SGA

[LRT88] Everett G. Larson, Kenneth G. Richards, and Darlene E. Tolman. Some geometrical algorithms for generation of field lines of a vector field — designed for high accuracy with calculations at only a few field points. International Journal of Quantum Chemistry, 34(S22):515–535, March 12–19, 1988. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Syposium on Quantum Chemistry, Solid-State Theory, and Computational Methods.

Lazzeretti:1984:MPI

[LRZ84a] P. Lazzeretti, E. Rossi, and R. Zanasi. Magnetic properties and induced current density in acetylene. International Journal of Quantum Chemistry, 25(6):1123–1134, June 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Lazzeretti:1984:SMF

[LRZ84b] P. Lazzeretti, E. Rossi, and R. Zanasi. Singularities of magnetic-field induced electron current density: a study of the ethylene molecule. International Journal of Quantum Chemistry, 25(6):929–940, June 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Ladik:1980:IBS

[LS80] Jáanos Ladik and Sáandor Suhai. Ab-initio band structure of polycytidine; internal charge transfer in DNA. International Journal of Quantum Chemistry, 18(S7):181–186, March 5–8, 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Lozes:1982:LEE

[LS82] Richard L. Lozes and John R. Sabin. Localized electronic excitation in a hydrogen bond. International Journal of Quantum Chemistry, 22(S9):289–296, March 4–6, 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Ladikm:1985:PCS

[LS85] J. J. Ladikm and M. Seel. Perturbation of cell-selfregulation through oncoproteins. International Journal of Quantum Chemistry, 28(S12):235–244, March 14–16, 1985. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Latajka:1986:EPT

[LS86] Zdzisław Latajka and Steve Scheiner. Energetics of proton transfer between carbon atoms (H_{3}CHCH_{3})^{−}. International Journal of Quantum Chemistry, 29(2):285–292, February 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Lischka:1981:NIG

[LSBS81] Hans Lischka, Ron Shepard, Franklin B. Brown, and Isaiah Shavitt. New implementation of the graphical unitary group approach for multireference direct configuration interaction calculations. International Journal of Quantum Chemistry, 20(S15):91–100, March 8–14, 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory, Collision Phenomena, and Computational Quantum Chemistry.

LopezDeCompadre:1988:RHA

[LSH88] Rosa L. Lopez De Compadre, Alan J. Shusterman, and Corwin Hansch. The role of hydrophobicity in the Ames test. The correlation of the mutagenicity of nitropolycyclic hydrocarbons with partition coefficients and molecular orbital indices. International Journal of Quantum Chemistry, 34(2):91– 101, August 1988. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Laaksonen:1985:TDF

[LSP85] Leif Laaksonen, Dage Sundholm, and Pekka Pyykkö. Two-dimensional, fully numerical molecular calculations. IV. Hartree–Fock–Slater results on second-row diatomic molecules. International Journal of Quantum Chemistry, 27 (5):601–612, May 1985. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Lobatch:1984:CSC

[LSR^{+}84] V. A. Lobatch, A. B. Sobolev, I. R. Rubik, O. E. Taurian, and J.-L. Calais. A critical study of the cluster model for ionic crystals. International Journal of Quantum Chemistry, 26(S18): 121–130, March 1–15, 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Atomic, Molecular and Solid-State Theory, and Computational Quantum Chemistry.

Levitt:1983:NMP

[LSS83] Michael Levitt, Christian Sander, and Peter S. Stern. The normal modes of a protein: Native bovine pancreatic trypsin inhibitor. International Journal of Quantum Chemistry, 24 (S10):181–199, March 14–16, 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Ladik:1989:TCH

[LSW89] János J. Ladik, Peter Saalfrank, and Richard F. Wood. Theoretical considerations on high T_{c} superconductivity. International Journal of Quantum Chemistry, 36(S23):621–630, April 1–8, 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Lowdin:1983:ECb

[LSZ83a] Per-Olov Löwdin, John R. Sabin, and Michael Zerner. Editorial comments. International Journal of Quantum Chemistry, 23(2):vii, February 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Lowdin:1983:ECc

[LSZ83b] Per-Olov Löwdin, John R. Sabin, and Michael Zerner. Editorial comments. International Journal of Quantum Chemistry, 23(3):v, March 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Lowdin:1983:ECd

[LSZ83c] Per-Olov Löwdin, John R. Sabin, and Michael Zerner. Editorial comments. International Journal of Quantum Chemistry, 23(4):xi, April 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Lowdin:1983:ECa

[LSZ83d] Pre-Olov Löwdin, John R. Sabin, and Michael Zerner. Editorial comments. International Journal of Quantum Chemistry, 23(1):vii, January 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Laurence:1981:ICC

[LT81] Patricia R. Laurence and Colin Thomson. Ab-initio calculations of charge transfer interactions involving ascorbic acid and its metabolites. International Journal of Quantum Chemistry, 20(S8):81–93, March 5–7, 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Lee:1989:DDQ

[LT89] Timothy J. Lee and Peter R. Taylor. A diagnostic for determining the quality of single-reference electron correlation methods. International Journal of Quantum Chemistry, 36 (S23):199–207, April 1–8, 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Lino:1987:BSS

[LTLF87] A. T. Lino, E. K. Takahashi, J. R. Leite, and A. C. Ferraz. Band structure of silicon by the self-consistent variational cellular method. International Journal of Quantum Chemistry, 32(S21):105–113, March 12, 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Chemistry, Solid-State Theory, and Computational Methods.

Ludenta:1983:TFT

[Lud83] Eduardo V. Ludeñta. Thomas–Fermi term as the simplest correction to the Weizsacker term. International Journal of Quantum Chemistry, 23(1):127–133, January 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Lukovits:1981:CSD

[Luk81] I. Lukovits. Comparative studies on drug-receptor binding of phenylethylamine derivatives using calculated group contribution constants. International Journal of Quantum Chemistry, 20(2):429–438, August 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Luken:1982:ESO

[Luk82] William L. Luken. Evaluation of the second-order reduced density matrix for correlated electronic wave functions. International Journal of Quantum Chemistry, 22(5):889–902, November 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Lunell:1982:BRB

[Lun82] Sten Lunell. Book review: Understanding physical chemistry. By Arthur W. Adamson, Benjamin/Cummings, 1980. Price: $9.95. International Journal of Quantum Chemistry, 21(5): 941, May 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). See [Ada64, Ada69, Ada80a].

Lundqvist:1989:NNE

[Lun89a] S. Lundqvist. Nonlocal and nonlinear electromagnetic effects at surfaces. International Journal of Quantum Chemistry, 35 (6):827–838, June 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Lundqvist:1989:IWY

[Lun89b] Stig Lundqvist. Ivar Waller 90 years on June 11, 1988. International Journal of Quantum Chemistry, 35(6):705–707, June 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Luo:1987:CDM

[Luo87] Liaofu Luo. Conformation dynamics of macromolecules. International Journal of Quantum Chemistry, 32(4):435–450, October 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Linderberg:1987:RSH

[LV87] Jan Linderberg and Behnam Vessal. Reactive scattering in hyperspherical coordinates. International Journal of Quantum Chemistry, 31(1):65–71, January 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Lucas:1985:EEL

[LVL^{+}85] A. A. Lucas, J. P. Vigneron, P. H. Lambin, P. A. Thiry, M. Liehr, J. J. Pireaux, and R. Caudano. Electron energy loss spectroscopy of surface and interface phonons of insulators, semiconductors, and superlattices. International Journal of Quantum Chemistry, 28(S19):687–705, March 18–23, 1985.

CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Luzanov:1981:SSP

[LW81] A. V. Luzanov and G. E. Whyman. Structure and spin-purity conditions for reduced density matrices of arbitrary order. International Journal of Quantum Chemistry, 20(6):1179–1199, December 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Lochmann:1982:CPM

[LW82] R. Lochmann and Th. Weller. Calculation of periodic molecular systems with perturbed periodicity within the PCILO framework. I. Method. International Journal of Quantum Chemistry, 21(4):711–716, April 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Lochmann:1984:CPM

[LW84a] R. Lochmann and Th. Weller. Calculation of periodic molecular systems with perturbed periodicity within the PCILO framework. II. Model systems. International Journal of Quantum Chemistry, 25(6):1061–1067, June 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Lochmann:1984:CSI

[LW84b] R. Lochmann and Th. Weller. Calculation of systems involving strong delocalized bonds within a modified PCILO framework. I. Valence-bond PCILO method. International Journal of Quantum Chemistry, 25(3):475–481, March 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Lochmann:1984:MPF

[LWK84] R. Lochmann, Th. Weller, and H.-J. Köhler. A modified PCILO framework using the CNDO /Boyd–Whitehead parametrization. International Journal of Quantum Chemistry, 25(6):1055–1060, June 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Lozes:1981:ITA

[LWÖ81] Richard L. Lozes, Brian Weiner, and Yngve Öhrn. Invariance transformations and AGP optimization. International Journal of Quantum Chemistry, 20(S15):129–131, March 8–14, 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory, Collision Phenomena, and Computational Quantum Chemistry.

Leszczynski:1988:TIM

[LWZ88] Jerzy Leszczynski, Floyd Wiseman, and Michael C. Zerner. Toward an ionic mechanism of soot particle formation: Reactions between acetylene and tautomeric forms of the C_{3}H_{3}^{+} ions. International Journal of Quantum Chemistry, 34(S22): 117–125, March 12–19, 1988. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Syposium on Quantum Chemistry, Solid-State Theory, and Computational Methods.

Liu:1983:TSE

[LY83] Ruo-Zhuang Liu and Jian-Guo Yu. Theoretical study of the effect of substituents on electrophilic addition to olefins: Ab initio MO approach with energy decomposition. International Journal of Quantum Chemistry, 23(2):491–495, February 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Lowdin:1983:I

[LZ83] Per-Olov Löwdin and Michael Zerner. Introduction. International Journal of Quantum Chemistry, 24(S10):vii, March 14–16, 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Liu:1987:MOM

[LZ87] Frank Liu and Chang-Guo Zhan. Maximum overlap method and the bond strength. International Journal of Quantum Chemistry, 32(1):1–11, July 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Li:1989:BTU

[LZ89] Xiangzhu Li and Qianer Zhang. Bonded tableau unitary group approach to the many-electron correlation problem. International Journal of Quantum Chemistry, 36(5):599–632, November 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Liu:1983:MOS

[LZY83] Ruo-Zhuang Liu, Shi-Jin Zheng, and Jian-Guo Yu. Molecular orbital study of the relationship between electronic structure and toxicity of silatranes. International Journal of Quantum Chemistry, 23(4):1609–1610, April 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Mayer:1980:EDS

[MA80a] I. Mayer and S. A. Angelov. EHF Description of superexchange: Relations to Anderson’s model. International Journal of Quantum Chemistry, 18(3):783–796, September 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Murgich:1980:CBC

[MA80b] J. Murgich and Y. Aray. Correlation between the CN stretching force constant and the electric field gradient at the N nucleus in some nitriles. International Journal of Quantum Chemistry, 18(S14):461, March 10–20, 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory, Collision Phenomena, Quantum Statistics, and Computational Methods.

Maksic:1980:CQM

[Mak80] Z. B. Maksić. Comment on quadrupole moments, dipole quadrupole A and quadrupole C polarizabilities by means of perturbation theory. International Journal of Quantum Chemistry, 18(6):1483–1484, December 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Malykhanov:1982:VPT

[Mal82] Yu. B. Malykhanov. Variational-perturbation theory in the LCAO-MO method for closed-shell molecules. International Journal of Quantum Chemistry, 21(2):399–410, February 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Malmqvist:1986:CTD

[Mal86] Per Åke Malmqvist. Calculation of transition density matrices by nonunitary orbital transformations. International Journal of Quantum Chemistry, 30(4):479–494, October 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Malinowski:1991:FAC

[Mal91] Edmund R. Malinowski. Factor analysis in chemistry. John Wiley, New York, NY, USA, second edition, 1991. ISBN 0-471-53009-3. xii + 350 pp. LCCN QD39.3.F33 M35

1991. URL http://www.loc.gov/catdir/description/
wiley034/90024193.htm;   http://www.loc.gov/catdir/
enhancements/fy0607/90024193-b.htm;   http://www.loc.
gov/catdir/toc/onix05/90024193.htm.

Malinowski:2002:FAC

[Mal02] Edmund R. Malinowski. Factor analysis in chemistry. John Wiley, New York, NY, USA, third edition, 2002. ISBN 0-471-13479-1 (hardcover). xviii + 414 pp. LCCN

QD39.3.F33 M35 2002. URL http://www.loc.gov/catdir/
bios/wiley043/2001046660.htm;   http://www.loc.gov/
catdir/description/wiley036/2001046660.htm;   http:
//www.loc.gov/catdir/toc/wiley021/2001046660.htm.

Mangoni:1981:OS

[Man81] Lorenzo Mangoni. Opening statement. International Journal of Quantum Chemistry, 19(6):977, June 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Maartensson:1980:BRB

[Mår80] Olle Mårtensson. Book review: Dielectric and electronic properties of biological materials. R. Pethig. John Wiley & Sons, New York, 1979. 376 pp. Price: $15.00 (cloth). International Journal of Quantum Chemistry, 17(3):595, March 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). See [Pet79].

Martensson:1981:BRB

[Mar81a] Olle Martensson. Book review: Carcinogenesis: Fundamental Mechanisms and Environmental Effects. Edited by B. Pullman, P. O. P. Ts’o, and H. Gelboin. The Jerusalem Symposia on Quantum Chemistry and Biochemistry. Vol. 13. Reidel, Dordrecht, 1980. ISBN 90-277-1171-2. International Journal of Quantum Chemistry, 20(6):1334, December 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). See [PTG80].

Martin:1981:MIC

[Mar81b] T. P. Martin. Matrix isolated clusters and microcrystals. International Journal of Quantum Chemistry, 19(6):1099–1103, June 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Maartensson:1982:BRB

[Mår82] Olle Mårtensson. Book review: Applications of photochemistry in probing biological targets. Edited by A. M. Tometsko and F. W. Richards, Ann. New York Academy of Sciences, New York, NY, 1980. Vol. 346. Price: $88.00. International Journal of Quantum Chemistry, 22(1):217, July 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). See [TR80].

March:1985:ECE

[Mar85a] N. H. March. Electron correlation and electron momenta in polyacetylene. International Journal of Quantum Chemistry, 28(S19):719–727, March 18–23, 1985. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

March:1985:RTE

[Mar85b] N. H. March. Relativistic total energy of heavy atomic ions: Dimensionality dependent scaling. International Journal of Quantum Chemistry, 27(5):595–599, May 1985. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

March:1986:DEE

[Mar86a] N. H. March. Differential equation for the electron density in large molecules. International Journal of Quantum Chemistry, 30(S13):3–8, March 6–8, 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

March:1986:EMM

[Mar86b] N. H. March. Electron momenta in molecules and lowdimensional solids. International Journal of Quantum Chemistry, 30(S20):367–376, March 10–15, 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Syposium on Atomic, Molecular, and Solid-State Theory, Scattering Problems, Many Body Phenomena, and Computational Quantum Chemistry.

Maranon:1988:PIF

[Mar88a] Julio Marañón. Path integral formulation of Roothaan’s equations. International Journal of Quantum Chemistry, 33(6): 491–496, June 1988. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

March:1988:SCE

[Mar88b] N. H. March. Spatial correlation of electrons in metals. International Journal of Quantum Chemistry, 34(S22):655–664, March 12–19, 1988. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Syposium on Quantum Chemistry, Solid-State Theory, and Computational Methods.

Maroulis:1988:CTD

[Mar88c] George Maroulis. Control over theoretical descriptions of atoms and molecules. An information theoretic approach to the automatic classification and systematic improvement of approximate wave functions. International Journal of Quantum Chemistry, 34(2):185–190, August 1988. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Marconi:1989:QMC

[Mar89] Giancarlo Marconi. Quantum mechanical calculations on vibronic activity in the MCD spectra of carbonyl compounds. International Journal of Quantum Chemistry, 36(3):417–425, September 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Maslen:1987:CCH

[Mas87a] V. W. Maslen. On the contribution of A. C. Hurley to theoretical chemistry. International Journal of Quantum Chemistry, 31(3):339–343, March 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Masson:1987:SEC

[Mas87b] David R. Masson. Schrödinger’s equation and continued fractions. International Journal of Quantum Chemistry, 32(S21): 699–712, March 12, 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Chemistry, Solid-State Theory, and Computational Methods.

Moscardo:1988:EMS

[MASF88] Federico Moscardó, Ramón Andarias, and Emilio San-Fabián. Estimation of minima in ^{1}A_{1} states of the ozone molecule. International Journal of Quantum Chemistry, 34(4):375–383, October 1988. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Matthias:1980:SE

[Mat80] Bernd T. Matthias. Superconductivity and eutectics. International Journal of Quantum Chemistry, 18(S14):729–730, March 10–20, 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory, Collision Phenomena, Quantum Statistics, and Computational Methods.

Matsen:1981:PHG

[Mat81a] F. A. Matsen. Particle-hole Gel’fand states and quasidegenerate perturbation theory. International Journal of Quantum Chemistry, 20(S15):163–175, March 8–14, 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory, Collision Phenomena, and Computational Quantum Chemistry.

Matsuoka:1981:MCM

[Mat81b] Osamu Matsuoka. Methods of composite molecular wave functions. I. Variational principle and multiconfiguration SCF theory. International Journal of Quantum Chemistry, 20(4):897– 908, October 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Matsuno:1982:TCP

[Mat82] Koichiro Matsuno. A theoretical construction of protobiological synthesis: From amino acids to functional protocells. International Journal of Quantum Chemistry, 22(S9):181–193, March 4–6, 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Matsen:1984:CGS

[Mat84] F. A. Matsen. Canonical generator states and their symmetry adaptation. International Journal of Quantum Chemistry, 26 (S18):43–56, March 1–15, 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Atomic, Molecular and Solid-State Theory, and Computational Quantum Chemistry.

Matsen:1987:FUG

[Mat87a] F. A. Matsen. The freeon unitary group procedure and the structure of matter. International Journal of Quantum Chemistry, 32(S21):713–728, March 12, 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Chemistry, Solid-State Theory, and Computational Methods.

Matsen:1987:SFQa

[Mat87b] F. A. Matsen. Spin-free quantum chemistry. XXIII. The generator-state approach. International Journal of Quantum Chemistry, 32(1):71–86, July 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Matsen:1987:SFQb

[Mat87c] F. A. Matsen. Spin-free quantum chemistry. XXIV. Freeon many-body theory. International Journal of Quantum Chemistry, 32(1):87–104, July 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Matsen:1987:SFQc

[Mat87d] F. A. Matsen. Spin-free quantum theory. XXV. The unitarygroup formulation of fermion many-body theory. International Journal of Quantum Chemistry, 32(1):105–122, July 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Mayer:1983:TCH

[May83] I. Mayer. Towards a “Chemical” Hamiltonian. International Journal of Quantum Chemistry, 23(2):341–363, February 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Mayer:1984:BOV

[May84] I. Mayer. Bond order and valence: Relations to Mulliken’s population analysis. International Journal of Quantum Chemistry, 26(1):151–154, July 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Mayer:1985:BOV

[May85] I. Mayer. Bond order and valence: Relations to Mulliken’s population analysis. International Journal of Quantum Chemistry, 28(3):419, September 1985. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Mayer:1986:BOVb

[May86a] I. Mayer. Bond orders and valences from ab initio wave functions. International Journal of Quantum Chemistry, 29(3): 477–483, March 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Mayer:1986:BOVa

[May86b] I. Mayer. On bond orders and valences in the Ab initio quantum chemical theory. International Journal of Quantum Chemistry, 29(1):73–84, January 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Mayer:1986:SCP

[May86c] I. Mayer. Simple constructive proof of Karadakov’s extended pairing theorem. International Journal of Quantum Chemistry, 29(1):31–34, January 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Mollmann:1980:THG

[MB80a] James C. Mollmann and Libero J. Bartolotti. Third-harmonic generation and intensity-dependent refractive index of atomic hydrogen. International Journal of Quantum Chemistry, 18 (S14):31–41, March 10–20, 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory, Collision Phenomena, Quantum Statistics, and Computational Methods.

Morton-Blake:1980:ESL

[MB80b] D. A. Morton-Blake. Effect of symmetry lowering on the band structure of polyacetylene. International Journal of Quantum Chemistry, 18(4):937–951, October 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Mondragon:1982:PTH

[MB82] A. Mondragón and I. Ortega Blake. Proton transference in hydrogen bonded systems. I. Uracil tautomerism. International Journal of Quantum Chemistry, 22(1):89–100, July 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Mishra:1983:TLS

[MB83] R. K. Mishra and K. Bhowmik. Theory of living State. VII. Bose–Einstein-like ordering in temperature and time domain. International Journal of Quantum Chemistry, 23(4): 1579–1587, April 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Migdalek:1984:CPR

[MB84] J. Migdalek and W. E. Baylis. Core polarization and relativistic effects competition in the first ionization potentials for some systems in the Cu, Ag, and Au isoelectronic sequences. International Journal of Quantum Chemistry, 25(1):77, January 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Muller:1985:EFS

[MB85] Wolfgang Müller and Paul S. Bagus. Electric field and surface charge effects on the vibrational frequency of CO chemisorbed on metal surfaces: Cluster model studies for CO/Cu(100). International Journal of Quantum Chemistry, 28(S19):737–738, March 18–23, 1985. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Marchese:1986:TMM

[MB86] Francis T. Marchese and David L. Beveridge. Thermal motion from Monte Carlo simulations of aqueous ionic solutions. International Journal of Quantum Chemistry, 29(4):619–625, April 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Mola:1989:AIM

[MB89] E. E. Mola and L. Blum. The adsorption of incommensurate monolayers on an hexagonal substrate: Lead underpotentially deposited on silver (111). International Journal of Quantum Chemistry, 36(S23):687–692, April 1–8, 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Montella:1984:CTA

[MBD84] Nicola Montella, Vincenzo Barone, and Giuseppe Del Re. Charge transfer in adsorbate-substrate systems: Extension of the Anderson–Newns treatment to two- and three-dimensional substrates. International Journal of Quantum Chemistry, 26 (5):769–774, November 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Meghezzi:1983:PAF

[MBKH83] H. Meghezzi, A. Boucekkine, B. Kolli, and M. Hamdi. Protonation of aminopyrone: a four centers problem, CNDO/2 versus MNDO study. International Journal of Quantum Chemistry, 24(4):425–428, October 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Morton-Blake:1980:MOT

[MBR80] D. A. Morton-Blake and N. R. Russell. Molecular-orbital treatment of complex cations of magnesium and beryllium with acetonitrile. International Journal of Quantum Chemistry, 18(6):1405–1413, December 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Mishra:1981:VWF

[MBSC81] R. K. Mishra, K. Bhaumik, A. Srivastava, and S. S. Chaudhary. Van der Waals force between two localized clusters of bosons. International Journal of Quantum Chemistry, 20(2):

377–383, August 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Mekenyan:1980:CGT

[MBT80] O. Mekenyan, D. Bonchev, and N. Trinajstić. Chemical graph theory: Modeling the thermodynamic properties of molecules. International Journal of Quantum Chemistry, 18(2):369–380, August 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Mekenyan:1981:ACB

[MBT81] Ov. Mekenyan, D. Bonchev, and N. Trinajstić. Algebraic characterization of bridged polycyclic compounds. International Journal of Quantum Chemistry, 19(5):929–955, May 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Massa:1980:LEM

[MC80] L. J. Massa and L. Cohen. A local energy method based on the reduced density matrix equations. International Journal of Quantum Chemistry, 18(S14):167–173, March 10–20, 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory, Collision Phenomena, Quantum Statistics, and Computational Methods.

March:1988:DDT

[MC88] N. H. March and J. Cizek. Dimensionality dependence of total energy of closed shells in a bare Coulomb field for large atomic number. International Journal of Quantum Chemistry, 33(4):301–304, April 1988. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Misiak:1989:PTS

[MC89] P. Misiak and H. Chojnacki. Proton transfer and solitonic excitations in model one-dimensional hydrogen-bonded systems. International Journal of Quantum Chemistry, 36(3):319–323, September 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Mccarroll:1983:MME

[Mcc83] R. Mccarroll. The molecular model of electronic transitions in atom–atom collisions: Successes and shortcomings. International Journal of Quantum Chemistry, 24(S17):33–35, March 3–12, 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-state Theory, Collision Phenomena, and Computational Quantum Chemistry.

McDowell:1980:IBS

[McD80] Keith McDowell. Incomplete basis set problem. IV. Perturbative corrections to pair energies. International Journal of Quantum Chemistry, 17(5):895–905, May 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Mcdowell:1981:AQW

[Mcd81a] Keith Mcdowell. Assessing the quality of a wavefunction using quantum Monte Carlo. International Journal of Quantum Chemistry, 20(S15):177–181, March 8–14, 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory, Collision Phenomena, and Computational Quantum Chemistry.

McDowell:1981:UNO

[McD81b] Keith McDowell. Use of noncanonical orbitals in manybody perturbation theory. International Journal of Quantum Chemistry, 19(2):271–282, February 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

McMahan:1986:NMH

[McM86] A. K. McMahan. New materials at high pressure. International Journal of Quantum Chemistry, 30(S20):393–408, March 10–15, 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Syposium on Atomic, Molecular, and Solid-State Theory, Scattering Problems, Many Body Phenomena, and Computational Quantum Chemistry.

McWeeny:1983:SRM

[McW83] R. McWeeny. Some remarks on multiconfiguration timedependent Hartree–Fock theory. International Journal of Quantum Chemistry, 23(2):405–416, February 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

McWeeny:1984:SML

[McW84] R. McWeeny. Simple models for large-molecule calculations. International Journal of Quantum Chemistry, 26(5):693–708, November 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

McWeeny:1988:SFF

[McW88] Roy McWeeny. A spin-free form of valence bond theory. International Journal of Quantum Chemistry, 34(1):25–36, July 1988. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Marchetti:1982:NIF

[MD82] M. Christina Marchetti and James W. Dufty. Nonlinear irreversible fluxes in a fluid with large shear rate. International Journal of Quantum Chemistry, 22(S16):83–89, March 1–13, 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Chemistry, Theory of Condensed Matter, and Propagator Methods in the Quantum Theory of Matter.

Matos:1987:EKT

[MD87] J. Mauricio O. Matos and Orville W. Day. The extended Koopmans’ theorem Fock operator. International Journal of Quantum Chemistry, 31(6):871–892, June 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Mezey:1982:CTI

[MDO^{+}82] P. G. Mezey, G. Del Re, P. Otto, S. Suhai, and J. Ladik. Charge transfer and induced polarization in model peptide– ion complexes. International Journal of Quantum Chemistry, 21(4):677–697, April 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Mishra:1985:IBP

[MDT85] R. K. Mishra, S. K. Dubey, and R. S. Tyagi. Interaction between photons, “trapped electron” and vibronic excitation. International Journal of Quantum Chemistry, 27(3):303–321, March 1985. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Mohallem:1986:GHW

[MDT86] J. R. Mohallem, R. M. Dreizler, and M. Trsic. A Griffin–Hill– Wheeler version of the Hartree–Fock equations. International Journal of Quantum Chemistry, 30(S20):45–55, March 10–15, 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Syposium on Atomic, Molecular, and Solid-State Theory, Scattering Problems, Many Body Phenomena, and Computational Quantum Chemistry.

Mohallem:1987:GHW

[MDT87] J. R. Mohallem, R. M. Dreizler, and M. Trsic. A Griffin–Hill– Wheeler version of the Hartree–Fock equations. International Journal of Quantum Chemistry, 32(3):409, September 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Mulliken:1981:PMR

[ME81] Robert Sanderson Mulliken and Walter C. Ermler, editors. Polyatomic molecules: results of ab initio calculations. Academic Press, New York, USA, 1981. ISBN 0-12-509860-X. xvi + 431 pp. LCCN QD461 .M79 1981.

Melin:1987:IMT

[Mel87] Anders Melin. Intertwining methods in the theory of inverse scattering. International Journal of Quantum Chemistry, 31 (5):739–746, May 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Mestechkin:1980:OSR

[Mes80] M. M. Mestechkin. Open-shell restricted Hartree–Fock density matrix direct calculations. International Journal of Quantum Chemistry, 18(1):19–23, July 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Mestechkin:1981:E

[Mes81] M. M. Mestechkin. Erratum. International Journal of Quantum Chemistry, 19(6):1221, June 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Mestechkin:1985:NMO

[Mes85] M. M. Mestechkin. Numerical method for the open-shell restricted Hartree–Fock density-matrix direct calculation. International Journal of Quantum Chemistry, 27(6):653–664, June 1985. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Meyer:1988:CLI

[Mey88] J. Meyer. Construction of linearly independent relativistic symmetry orbitals for finite double-point groups including time reversal symmetry. International Journal of Quantum Chemistry, 33(5):445–465, May 1988. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Mezey:1981:IIE

[Mez81a] Paul G. Mezey. The isoelectronic and isoprotonic energy hypersurface and the topology of the nuclear charge space. International Journal of Quantum Chemistry, 20(S15):279–285, March 8–14, 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory, Collision Phenomena, and Computational Quantum Chemistry.

Mezey:1981:MTM

[Mez81b] Paul G. Mezey. Manifold theory of multidimensional potential surfaces. International Journal of Quantum Chemistry, 20(S8):185–196, March 5–7, 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Mezey:1982:LST

[Mez82] Paul G. Mezey. Level set topology of the nuclear charge space and the electronic energy functional. International Journal of Quantum Chemistry, 22(1):101–114, July 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Mezey:1983:ACA

[Mez83a] Paul G. Mezey. An approach to conformation analysis on multidimensional potential surfaces. International Journal of Quantum Chemistry, 24(S10):153–160, March 14–16, 1983.

CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Mezey:1983:DMM

[Mez83b] Paul G. Mezey. The differentiable manifold model of quantumchemical reaction networks. International Journal of Quantum Chemistry, 24(S17):137–152, March 3–12, 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-state Theory, Collision Phenomena, and Computational Quantum Chemistry.

Mezey:1983:IHR

[Mez83c] Paul G. Mezey. Inequalities and homotopy relations in reaction topology. International Journal of Quantum Chemistry, 24(S17):453–460, March 3–12, 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-state Theory, Collision Phenomena, and Computational Quantum Chemistry.

Mezey:1983:MGI

[Mez83d] Paul G. Mezey. A molecular geometry invariant property of energy level set boundaries in z space. International Journal of Quantum Chemistry, 24(6):523–526, December 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Mezey:1984:ASQ

[Mez84a] Paul G. Mezey. The algebraic structure of quantum-chemical reaction mechanisms. International Journal of Quantum Chemistry, 26(S18):77–85, March 1–15, 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Atomic, Molecular and Solid-State Theory, and Computational Quantum Chemistry.

Mezey:1984:GFQ

[Mez84b] Paul G. Mezey. A general formulation of the “quantum chemical le Chatelier principle”. International Journal of Quantum Chemistry, 25(5):853–861, May 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Mezey:1984:HCR

[Mez84c] Paul G. Mezey. Hyperspherical coordinate representation of potential surfaces of large molecules. International Journal of Quantum Chemistry, 26(S11):267–272, March 12–15, 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Mezey:1984:MPR

[Mez84d] Paul G. Mezey. The metric properties of the reduced nuclear configuration space. International Journal of Quantum Chemistry, 26(6):983–985, December 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Mezey:1984:NRP

[Mez84e] Paul G. Mezey. Network relations on potential surfaces as aids to computer-based quantum-chemical synthesis planning. International Journal of Quantum Chemistry, 26(S18):675–681, March 1–15, 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Atomic, Molecular and Solid-State Theory, and Computational Quantum Chemistry.

Mezey:1985:CTG

[Mez85a] Paul G. Mezey. A comparison of two group theoretical models of reaction mechanisms on potential surfaces. International Journal of Quantum Chemistry, 28(3):387–398, September 1985. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Mezey:1985:GTE

[Mez85b] Paul G. Mezey. Group theory of electrostatic potentials: a tool for quantum chemical drug design. International Journal of Quantum Chemistry, 28(S12):113–122, March 14–16, 1985. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Mezey:1985:RPG

[Mez85c] Paul. G. Mezey. The reaction polyhedron and group theory of reaction mechanisms. International Journal of Quantum Chemistry, 28(S19):93–105, March 18–23, 1985. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Mezey:1986:NGC

[Mez86a] Paul G. Mezey. New global constraints on electronic energy hypersurfaces. International Journal of Quantum Chemistry, 29(1):85–99, January 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Mezey:1986:NCM

[Mez86b] Paul G. Mezey. Nuclear charges and molecular total energies: a rule on nested reaction globes. International Journal of Quantum Chemistry, 29(3):333–343, March 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Mezey:1987:GTS

[Mez87a] Paul G. Mezey. Group theory of shapes of asymmetric biomolecules. International Journal of Quantum Chemistry, 32(S14):127–132, March 14, 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Mezey:1987:RPR

[Mez87b] Paul G. Mezey. Reflection properties of reaction paths in the reduced nuclear configuration space. International Journal of Quantum Chemistry, 32(S21):191–198, March 12, 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Chemistry, Solid-State Theory, and Computational Methods.

Mota:1987:PNU

[MF87] Ronaldo Mota and Adalberto Fazzio. On the possibility of negative U systems for transition metals impurities in semiconductors. International Journal of Quantum Chemistry, 32 (S21):73–78, March 12, 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Chemistry, Solid-State Theory, and Computational Methods.

Miertus:1989:QMS

[MFM89] S. Miertuš, V. Frecer, and M. Májeková. QSAR and Mechanistic studies on the genotoxic compounds including environmental effects. International Journal of Quantum Chemistry, 35(1):153–165, January 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Moore:1980:EII

[MG80] Richard D. Moore and Raj K. Gupta. Effect of insulin on intracellular pH as observed by ^{31}P NMR spectroscopy. International Journal of Quantum Chemistry, 18(S7):83–92, March 5–8, 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Maouche:1983:QSG

[MG83] Boubekeur Maouche and José Gayoso. Quantum study of the ground state of a series of unsaturated boron–nitrogen compounds by the MNDO method. I. Geometries and stabilities. International Journal of Quantum Chemistry, 23(3):891–904, March 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Murphy:1985:OIA

[MG85] William C. Murphy and Thomas F. George. Overlap integrals for atom–metal surface interactions. International Journal of Quantum Chemistry, 28(5):631–639, November 1985. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Mehler:1989:IMA

[MG89] E. L. Mehler and J. Gerhards. Interaction model for the antiinflammatory action of benzoic and salicylic acids and phenols. International Journal of Quantum Chemistry, 35(1):205–214, January 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Montenay-Garestier:1981:CFB

[MGBH81] T. Montenay-Garestier, J. C. Brochon, and C. Hélène. Complex formation between tryptophan-containing peptides and nucleic acids: Fluorescence decay studies using synchrotron radiation. International Journal of Quantum Chemistry, 20 (1):41–48, July 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Maranon:1982:CTC

[MGN82] Julio Marañón, Horacio Grinberg, and Norma Sbarbati Nudelman. Charge-transfer character of the intermolecular hydrogen bond in singlet and triplet excited electronic states of the adenine–thymine base pair. International Journal of Quantum Chemistry, 22(1):69–72, July 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Maekawa:1980:MLM

[MH80a] Hiroyuki Maekawa and Masahiko Hosokawa. Model for the lipid monolayer-phase transitions. International Journal of Quantum Chemistry, 18(6):1511–1516, December 1980. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Malinowski:1980:FAC

[MH80b] Edmund R. Malinowski and Darryl G. Howery. Factor analysis in chemistry. John Wiley, New York, NY, USA, 1980. ISBN 0-471-05881-5. ix + 251 pp. LCCN QD39.3.M3 M34. US$23.50

(est.). URL http://www.loc.gov/catdir/enhancements/
fy0607/79027081-d.htm;   http://www.loc.gov/catdir/
enhancements/fy0607/79027081-t.htm.

Mohammad:1982:THB

[MH82] S. Noor Mohammad and A. J. Hopfinger. Treatment of hydrogen bonding within CNDO/2 and MINDO /3: CNDO /2H and MINDO /3H. International Journal of Quantum Chemistry, 22(6):1189–1207, December 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Malinowski:1989:FAC

[MH89] Edmund R. Malinowski and Darryl G. Howery. Factor analysis in chemistry. Robert E. Krieger Publishing Company, Huntington, NY, USA, 1989. ISBN 0-89464-343-6. ix + 251 pp. LCCN QD39.3.F33 M35 1989.

Michalska:1986:ECE

[MHAS86] D. Michalska, B. Hess, Andes, Jr., and L. J. Schaad. The effect of correlation energy (MP2) on computed vibrational frequencies. International Journal of Quantum Chemistry, 29 (5):1127–1137, May 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Micha:1981:CET

[Mic81a] David A. Micha. Cumulant expansion of time-correlation functions for collisional energy transfer. International Journal of Quantum Chemistry, 20(S15):643–652, March 8–14, 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory, Collision Phenomena, and Computational Quantum Chemistry.

Michels:1981:EAI

[Mic81b] M. A. J. Michels. Evaluation of angular integrals in rotational invariants. International Journal of Quantum Chemistry, 20 (4):951–952, October 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Micha:1983:STC

[Mic83a] David A. Micha. Semiclassical time-correlation function approach to collisional energy transfer into many-atom systems. International Journal of Quantum Chemistry, 23(2):551–560, February 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Micha:1983:VDS

[Mic83b] David A. Micha. Variational determination of self-consistent interactions in atomic collisions. International Journal of Quantum Chemistry, 24(S17):153–161, March 3–12, 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-state Theory, Collision Phenomena, and Computational Quantum Chemistry.

Micha:1985:CTC

[Mic85] David A. Micha. Collisional time-correlation functions for molecular interactions. International Journal of Quantum Chemistry, 28(S19):443–455, March 18–23, 1985. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Micha:1986:CTC

[Mic86] David A. Micha. Collisional time-correlation functions for molecular interactions. International Journal of Quantum Chemistry, 30(S20):773–785, March 10–15, 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Syposium on Atomic, Molecular, and Solid-State Theory, Scattering Problems, Many Body Phenomena, and Computational Quantum Chemistry.

Micha:1987:BRB

[Mic87] David A. Micha. Book review: The theory of chemical reaction dynamics. Edited by D. C. Clary, D. Reidel Publishing Company, Dordrecht, Holland, 1986. International Journal of Quantum Chemistry, 32(3):397, September 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). See [Cla86].

Miertus:1989:I

[Mie89] Stanislav Miertuš. Introduction. International Journal of Quantum Chemistry, 35(1):1, January 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Millen:1986:GFG

[Mil86] D. J. Millen. General features of geometries and potential constants of hydrogen bonded dimers. International Journal of Quantum Chemistry, 29(2):191–196, February 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Minaev:1980:ISF

[Min80] B. F. Minaev. Intensities of spin-forbidden transitions in molecular oxygen and selective heavy-atom effects. International Journal of Quantum Chemistry, 17(2):367–374, February 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Mintmire:1987:PRL

[Min87] J. W. Mintmire. Preliminary report on a local-density functional study of polysilane chains. International Journal of Quantum Chemistry, 32(S21):761–762, March 12, 1987. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Chemistry, Solid-State Theory, and Computational Methods.

Mishra:1984:LSX

[Mis84] R. K. Mishra. The living state. XIII. Does matter become “mind”? International Journal of Quantum Chemistry, 26 (S11):45–53, March 12–15, 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Meissner:1989:SEM

[MJ89] L. Meissner and K. Jankowski. Size-extensivity in multireference many-body perturbation theories: a direct comparison between single-reference and multireference perturbation theories in the nondegenerate case. International Journal of Quantum Chemistry, 36(6):705–726, December 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Meissner:1988:CCM

[MJW88] L. Meissner, K. Jankowski, and J. Wasilewski. A coupledcluster method for quasidegenerate states. International Journal of Quantum Chemistry, 34(6):535–557, December 1988. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Morita:1980:LEL

[MK80] Makoto Morita and Yoshiaki Kato. N_{2} Laser-excited luminescence of antiferromagnetically coupled trinuclear chromium(III) complexes. International Journal of Quantum Chemistry, 18(2):625–631, August 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Moskowitz:1981:NLC

[MK81] Jules W. Moskowitz and M. H. Kalos. A new look at correlations in atomic and molecular systems. I. Application of fermion Monte Carlo variational method. International Journal of Quantum Chemistry, 20(5):1107–1119, November 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Miyoshi:1983:TSE

[MK83] Eisaku Miyoshi and Hiroshi Kashiwagi. Theoretical study of equilibrium Co F bond distance of CoF_{6}^{n−} complexes (n = 4, 3, and 2) in crystals. International Journal of Quantum Chemistry, 24(1):85–96, July 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Mathers:1985:CBS

[MK85] T. L. Mathers and N. R. Kestner. Correlation and basis set superposition effects on three body ion-water-water interactions. International Journal of Quantum Chemistry, 28(S19): 297–311, March 18–23, 1985. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Mladenov:1980:MCC

[MKVP80] Ivailo M. Mladenov, Marin D. Kotarov, and Julia G. Vassileva-Popova. Method for computing the characteristic polynomial. International Journal of Quantum Chemistry, 18(1):339–341, July 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Mintmire:1985:TPS

[MKW85] J. W. Mintmire, F. W. Kutzler, and C. T. White. Theoretical photoelectron spectra using local-density function results. International Journal of Quantum Chemistry, 28(S19): 745, March 18–23, 1985. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Mckeen:1981:SDR

[ML81] Leighton E. Mckeen and Everett G. Larson. Symmetrydecomposed reduced transition operators for degenerate atomic states — their generation and their utility. International Journal of Quantum Chemistry, 20(S15):183–188, March 8–14, 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory, Collision Phenomena, and Computational Quantum Chemistry.

Micha:1983:IP

[ML83] David A. Micha and Per-Olov Löwdin. Introduction, Part I. International Journal of Quantum Chemistry, 24(S17):1, March 3–12, 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-state Theory, Collision Phenomena, and Computational Quantum Chemistry.

Miller:1980:IMN

[MLA80] Kenneth J. Miller, Michael Lauer, and Sydney Archer. Interactions of molecules with nucleic acids. V. Intercalation of thioxanthenones into DNA. International Journal of Quantum Chemistry, 18(S7):11–34, March 5–8, 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Marchenko:1981:DDR

[MLC^{+}81] T. V. Marchenko, V. M. Lipkin, O. Yu. Chertov, E. D. Sverdlov, and Yu. A. Ovchinnikov. DNA-dependent RNA polymerase from E. coli: Structure investigation. International Journal of Quantum Chemistry, 20(2):283–289, August 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

McKoy:1983:DMP

[MLL83] Vincent McKoy, Diane Lynch, and Robert R. Lucchese. Dynamics of molecular photoionization processes. International Journal of Quantum Chemistry, 24(S17):89–100, March 3–12, 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-state Theory, Collision Phenomena, and Computational Quantum Chemistry.

Mccaskill:1980:DAI

[MM80] J. S. Mccaskill and N. H. March. Dynamics of adatoms interacting with metals and metal surfaces: The validity of the description given by Kramers’ equation. International Journal of Quantum Chemistry, 18(S14):681–693, March 10–20, 1980.

CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory, Collision Phenomena, Quantum Statistics, and Computational Methods.

Mukhopadhyay:1981:SCM

[MM81] A. K. Mukhopadhyay and N. G. Mukherjee. Self-consistent methods in Hückel and extended Hückel theories. International Journal of Quantum Chemistry, 19(4):515–519, April 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Magnoli:1982:OSC

[MM82] Douglas E. Magnoli and Joseph R. Murdoch. Obtaining selfconsistent wave functions which satisfy the virial theorem. International Journal of Quantum Chemistry, 22(6):1249–1262, December 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Manson:1987:EEC

[MM87a] Steven T. Manson and John H. Miller. Electron ejection cross sections in electron and ion impact ionization: Ab initio and semiempirical calculations. International Journal of Quantum Chemistry, 32(S21):297–306, March 12, 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Chemistry, Solid-State Theory, and Computational Methods.

McIsaac:1987:EWF

[MM87b] K. McIsaac and E. N. Maslen. Exact wave functions for fewparticle systems: The choice of expansion for Coulomb potentials. International Journal of Quantum Chemistry, 31(3): 361–368, March 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Mehrotra:1984:MCD

[MMB84] P. K. Mehrotra, M. Mezei, and D. L. Beveridge. Monte Carlo determination of the internal energies of hydration for the Ala dipeptide in the C_{7}, C_{5}, α_{R}, and P_{II} conformations. International Journal of Quantum Chemistry, 26(S11):301–308, March 12–15, 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Mukherjee:1983:FFC

[MMC83] Prasanta K. Mukherjee, Tsutomu Minato, and Delano P. Chong. Finite-field calculations of molecular polarizabilities using field-induced polarization functions. II. Second- and third-order perturbation correlation corrections to the coupled Hartree–Fock polarizability of HF. International Journal of Quantum Chemistry, 23(2):447–457, February 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Minaev:1982:QCC

[MMI^{+}82] B. F. Minaev, Z. M. Muldahmetov, I. S. Irgibaeva, T. O. Tlepbergenov, and D. M. Kizhner. Quantum chemical calculation of phosphorescence microwave double resonance spectra. International Journal of Quantum Chemistry, 22(5):863–869, November 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Marini:1981:RTB

[MMM81] U. Marini, Bettolo Marconi, and N. H. March. Relativistic theory of binding energies of heavy positive ions. International Journal of Quantum Chemistry, 20(3):693–704, September 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Mancini:1987:PGF

[MMN87] F. Mancini, M. Marinaro, and Y. Nakano. n-Point Green’s functions in the Anderson model. International Journal of Quantum Chemistry, 32(S21):55–64, March 12, 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Chemistry, Solid-State Theory, and Computational Methods.

Macias:1988:PSU

[MMRY88] A. Macías, F. Martín, A. Riera, and M. Yáañez. A practical solution to the “unknown normalization” problem. International Journal of Quantum Chemistry, 33(4):279–300, April 1988. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Matsumoto:1986:ICQ

[MMS^{+}86] Shiro Matsumoto, Kimiko Mizutani, Atsushi Sekiguchi, Takakazu Yano, and Masayuki Toyama. Ab initio calculations on the quenching of excited lithium atom by molecular hydrogen. I. International Journal of Quantum Chemistry, 29 (4):689–699, April 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Matsen:1981:SFQ

[MN81] F. A. Matsen and C. J. Nelin. Spin-free quantum chemistry. XXII. Multiconfiguration self-consistent field theory. International Journal of Quantum Chemistry, 20(4):861–878, October 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Murty:1985:AMA

[MNS85] J. S. Murty, R. S. Nikam, and C. R. Sarma. Angular momentum adaptation of spinor basis. International Journal of Quantum Chemistry, 28(S19):69–77, March 18–23, 1985. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Mishra:1980:CTP

[MÖ80] Manoj Mishra and Yngve Öhrn. Correlated treatment of photoionization cross sections in the orthogonalized plane wave approximation for the ejected electron. International Journal of Quantum Chemistry, 18(S14):335–348, March 10–20, 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory, Collision Phenomena, Quantum Statistics, and Computational Methods.

Modrak:1981:CDS

[Mod81] P. Modrak. Calculation of the density of states of tightbinding d band for simple cubic lattice. International Journal of Quantum Chemistry, 20(3):591–600, September 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Moffat:1981:SSE

[Mof81] J. B. Moffat. Structures, stabilities, and energies of isomerization of some alkyl cyanides and isocyanides and their corresponding α-cyanocarbonium and α-isocyanocarbonium ions. International Journal of Quantum Chemistry, 19(5):771–779, May 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Moffat:1982:SSE

[Mof82] J. B. Moffat. Structures, stabilities, and energies of isomerization of α-cyano and α-isocyano carbanions. International Journal of Quantum Chemistry, 22(2):299–306, August 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Mohammad:1980:HAWb

[Moh80a] S. Noor Mohammad. High accuracy wave functions and eigenenergies for the vibration–rotation states of diatomic molecules by factorization method. International Journal of Quantum Chemistry, 18(5):1133–1142, November 1980. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Mohammad:1980:HAWa

[Moh80b]  S. Noor Mohammad.  High accuracy wave functions

and eigenenergies for vibration–rotation states of diatomic molecules. International Journal of Quantum Chemistry, 17 (5):943–954, May 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Moiseyev:1981:SPR

[Moi81] Nimrod Moiseyev. Study of predissociation resonances by the complex coordinate method. International Journal of Quantum Chemistry, 20(4):835–842, October 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Moore:1981:ITS

[Moo81] Richard D. Moore. The insulin transduction system: A biophysical model for mitogenesis. International Journal of Quantum Chemistry, 20(S8):365–371, March 5–7, 1981. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Moriarty:1983:GPT

[Mor83] John A. Moriarty. Generalized pseudopotential theory of dband metals. International Journal of Quantum Chemistry, 24(S17):541–554, March 3–12, 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-state Theory, Collision Phenomena, and Computational Quantum Chemistry.

Morrison:1988:DDM

[Mor88] Robert C. Morrison. Density and density matrix from optimized linearly independent product basis functions for Be. International Journal of Quantum Chemistry, 34(S22):43–49, March 12–19, 1988. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Syposium on Quantum Chemistry, Solid-State Theory, and Computational Methods.

Morrison:1989:EDM

[Mor89] Robert C. Morrison. An explicit density matrix functional of the (N − 1)-particle system when the N-particle system is known. International Journal of Quantum Chemistry, 36 (S23):583–590, April 1–8, 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Miyoshi:1983:TSD

[MOT^{+}83] Eisaku Miyoshi, Shigeru Obara, Toshikazu Takada, Hiroshi Kashiwagi, and Kimio Ohno. Theoretical study of degree of covalency in some CoF_{6}(n−) complexes (n = 4, 3, and 2). International Journal of Quantum Chemistry, 23(5):1753–1765, May 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Mouritsen:1989:BRB

[Mou89] Ole G. Mouritsen. Book review: Introduction to Phase Transitions and Critical Phenomena. By H. Eugene Stanley, Oxford University Press, Oxford, 1987. International Journal of Quantum Chemistry, 35(4):583–584, April 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). See [Sta71, Sta87].

Morgan:1984:ESP

[MP84a] Hywel Morgan and Ronald Pethig. Experimental studies of proton transport in hydrated lysozyme and α-cyclodextrin. International Journal of Quantum Chemistry, 26(S11):209– 216, March 12–15, 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Mukhopadhyay:1984:UGF

[MP84b] Atri Mukhopadhyay and Barry T. Pickup. A unitary group formulation of many-body theory: Diagram systematics and use of the spin shifts. International Journal of Quantum Chemistry, 26(1):125–143, July 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Murray:1987:CSS

[MP87] Jane S. Murray and Peter Politzer. A computational study of some isomerization equilibria and their possible relation to vinyl chloride carcinogenicity. International Journal of Quantum Chemistry, 31(4):569–579, April 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Morales:1984:AAO

[MPB84] J. Morales, A. Palma, and M. Berrondo. Algebraic approach to one- and two-center harmonic oscillator integrals. International Journal of Quantum Chemistry, 26(S18):57–60, March 1–15, 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Atomic, Molecular and Solid-State Theory, and Computational Quantum Chemistry.

Maric:1981:GRM

[MPBN81] Z. Marić, M. Popović-Božić, and Z. Nikolovski. Generalized Racah method for the evaluation of CFP for mixed electronic configurations. International Journal of Quantum Chemistry, 20(3):763–774, September 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Muller-Plathe:1987:PBM

[MPD87] Florian Müller-Plathe and Geerd H. F. Diercksen. Properties of the BCl molecule studied by fourth-order many-body perturbation theory. International Journal of Quantum Chemistry, 32(S21):595–602, March 12, 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Chemistry, Solid-State Theory, and Computational Methods.

Myer:1981:CCA

[MPP^{+}81] Yash P. Myer, A. Pande, J. Pande, K. K. Thallam, A. F. Saturno, and Balbir C. Verma. Cytochrome c: Ascorbate reduction site and possible electron-transfer path. International Journal of Quantum Chemistry, 20(2):513–521, August 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Morales:1983:EPI

[MPP83] J. Morales, A. Palma, and E. Poulain. ESCA Peak intensities in the dipole and generalized sudden approximations using Fourier transforms of GTO’s. International Journal of Quantum Chemistry, 23(2):417–423, February 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Muszynska:1982:BAP

[MPPW82] J. Muszyńska, D. Papierowska, J. Pipin, and W. Woźnicki. Bounds to atomic properties: Electric polarizabilities of lithium isoelectronic sequence ions. International Journal of Quantum Chemistry, 22(6):1153–1175, December 1982. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Morales:1986:HTL

[MPS86] J. Morales, A. Palma, and L. Sandoval. Hypervirial theorem and ladder operators. Recurrence relations for harmonic oscillator integrals. International Journal of Quantum Chemistry, 29(2):211–219, February 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Macias:1980:UMP

[MR80a] A. Macías and A. Riera. Use of molecular properties in the one-dimensional model of nonadiabatic transitions. International Journal of Quantum Chemistry, 17(2):181–194, February 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Mukherjee:1980:SEM

[MR80b] Nirad Gopal Mukherjee and Pratip Ray. Simplification of an earlier MCSCF procedure for CI expansions involving orbital excitations. International Journal of Quantum Chemistry, 18 (4):1045–1047, October 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Murthy:1980:SMO

[MR80c] A. S. N. Murthy and Shoba Ranganathan. Semiempirical molecular-orbital studies on the problem of nonclassical resonance in the homoallylic cation. International Journal of Quantum Chemistry, 18(6):1479–1482, December 1980. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Mukerjee:1981:SEM

[MR81] Nirad Gopal Mukerjee and Pratip Ray. Simplification of an earlier MCSCF procedure for CI expansions involving orbital excitations. International Journal of Quantum Chemistry, 20 (4):953, October 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Murthy:1985:CFM

[MR85] A. S. N. Murthy and Shoba Ranganathan. Compliant fields for molecular interactions: Water dimer and formic acid dimer. International Journal of Quantum Chemistry, 27(5):547–557, May 1985. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Mikkelsen:1987:ETR

[MR87] Kurt V. Mikkelsen and Mark A. Ratner. Electron transfer reactions dynamically coupled to a dielectric medium: Orientational effects and bridge assistance. International Journal of Quantum Chemistry, 32(S21):341–354, March 12, 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Chemistry, Solid-State Theory, and Computational Methods.

Mikkelsen:1988:SSE

[MR88] Kurt V. Mikkelsen and Mark A. Ratner. Specific solvation effects on electron transfer. International Journal of Quantum Chemistry, 34(S22):707–720, March 12–19, 1988. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Syposium on Quantum Chemistry, Solid-State Theory, and Computational Methods.

Mcdonald:1989:MMC

[MR89] Joseph J. Mcdonald and Robert Rein. Molecular modeling of calmodulin: A comparison with crystallographic data. International Journal of Quantum Chemistry, 36(S16):57–72, April 1–8, 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Muller-Remmers:1985:OCDa

[MRJ85a] Peter L. Müller-Remmers and Karl Jug. Orbital correlation diagrams based on multiconfigurational variation of moments I. Theory. International Journal of Quantum Chemistry, 28 (6):691–701, December 1985. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Muller-Remmers:1985:OCDb

[MRJ85b] Peter L. Müller-Remmers and Karl Jug. Orbital correlation diagrams based on multiconfigurational variation of moments II. Application. International Journal of Quantum Chemistry, 28(6):703–713, December 1985. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Mintmire:1980:LDF

[MS80a] J. W. Mintmire and J. R. Sabin. Local density functional methods in two-dimensionally periodic systems. I. The atomic hydrogen monolayer. International Journal of Quantum Chemistry, 18(S14):707–713, March 10–20, 1980. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory, Collision Phenomena, Quantum Statistics, and Computational Methods.

Morgan:1980:BMP

[MS80b] John D. Morgan III and Barry Simon. Behavior of molecular potential energy curves for large nuclear separations. International Journal of Quantum Chemistry, 17(6):1143–1166, June 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Mele:1984:NIT

[MS84] E. J. Mele and J. Socolar. Numerical integration of the time evolution operator: Excited-state dynamics in conjugated molecules. International Journal of Quantum Chemistry, 26(S18):347–358, March 1–15, 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Atomic, Molecular and Solid-State Theory, and Computational Quantum Chemistry.

Manmoto:1986:MPS

[MS86] Yoshinori Manmoto and Mitsuo Satoh. Multiphoton processes in a solid in a doubly rotating system by NMR. International Journal of Quantum Chemistry, 30(S20):647–655, March 10– 15, 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Syposium on Atomic, Molecular, and Solid-State Theory, Scattering Problems, Many Body Phenomena, and Computational Quantum Chemistry.

Maclagan:1987:VBC

[MS87] Robert G. A. R. Maclagan and Richard W. Simpson. Valencebond calculations on N_{2} and isoelectronic species. International Journal of Quantum Chemistry, 31(3):463–470, March 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Martin:1988:IES

[MS88] Richard L. Martin and Paul W. Saxe. Ab initio electronic structure theory for a cluster model of La_{2−x}Sr_{x}CuO_{4}. International Journal of Quantum Chemistry, 34(S22):237–244, March 12–19, 1988. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Syposium on Quantum Chemistry, Solid-State Theory, and Computational Methods.

Mayer:1989:IIS

[MS89] I. Mayer and P. R. Surján. Improved intermolecular SCF theory and the BSSE problem. International Journal of Quantum Chemistry, 36(3):225–240, September 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Martin:1983:QCSa

[MSC^{+}83a] M. Martin, F. Sanz, M. Campillo, L. Pardo, J. Pérez, and J. Turmo. Quantum chemical study of the molecular patterns of MAO inhibitors and substrates. International Journal of Quantum Chemistry, 23(4):1627–1641, April 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Martin:1983:QCSb

[MSC^{+}83b] M. Martin, F. Sanz, M. Campillo, L. Pardo, J. Perez, J. Turmo, and J. M. Aulló. Quantum chemical structure– activity relationships on β-carbolines as natural monoamine oxidase inhibitors. International Journal of Quantum Chemistry, 23(4):1643–1652, April 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Moscardo:1986:DAC

[MSF86] Federico Moscardó and Emilio San-Fabián. Dynamical analysis of correlated wavefunctions. International Journal of Quantum Chemistry, 30(6):853–865, December 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Mitra:1980:SSD

[MSGPP80] C. K. Mitra, Ramaswamy H. Sarma, C. Giessner-Prettre, and Bernard Pullman. Solution structure of DNA: The method of nuclear magnetic resonance spectroscopy. International Journal of Quantum Chemistry, 18(S7):39–66, March 5–8, 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Morita:1980:ILM

[MSI80] Keiko Morita, Hiroshi Sato, and Eiichi Ishiguro. Ab initio LCAO MO SCF calculation of potential surface of the NO_{2} molecule. International Journal of Quantum Chemistry, 18 (1):151–156, July 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Meshitsuka:1981:PNS

[MSM81] Shunsuke Meshitsuka, Gary M. Smith, and Albert S. Mildvan. Proton NMR studies of the histidine residues and of bound ATP on rabbit muscle pyruvate kinase. International Journal of Quantum Chemistry, 20(S8):241–245, March 5–7, 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Mitani:1980:SRS

[MST^{+}80] T. Mitani, S. Suga, Y. Tokura, K. Koyama, I. Nakada, and T. Koda. Synchrotron radiation spectroscopy of quasi-one-dimensional materials. International Journal of Quantum Chemistry, 18(2):655–659, August 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Mayer:1989:BFS

[MSV89] I. Mayer, P. R. Surján, and Á. VibÓk. BSSE-free SCF methods for intermolecular interactions. International Journal of Quantum Chemistry, 36(S23):281–290, April 1–8, 1989. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Moskowitz:1988:ASA

[MSWC88] Jules W. Moskowitz, K. E. Schmidt, S. R. Wilson, and W. Cui. The application of simulated annealing to problems of molecular mechanics. International Journal of Quantum Chemistry, 34(S22):611–617, March 12–19, 1988. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Syposium on Quantum Chemistry, Solid-State Theory, and Computational Methods.

Majumdar:1985:ESP

[MT85] Rabi Majumdar and Ashoke R. Thakur. Effect of structural perturbations on helix–coil transitions in covalently closed supercoiled DNA. International Journal of Quantum Chemistry, 28(S12):229–234, March 14–16, 1985. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Mishra:1987:FCD

[MT87] P. C. Mishra and R. D. Tewari. Fixed closest distance of approach dipole potential mapping: An effective alternative to the monopole isopotential approach. International Journal of Quantum Chemistry, 32(2):181–191, August 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Mohallem:1988:CGE

[MT88] José R. Mohallem and Milan Trsic. Considerations on Gaussian expanded generator coordinate wave functions. International Journal of Quantum Chemistry, 33(6):555–561, June 1988. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Maruyama:1980:FST

[MTI80] Yusei Maruyama and Keiko Takamiya-Ichikawa. Fluorescence spectra and their temperature dependence of amorphous anthracene films doped with naphthacene. International Journal of Quantum Chemistry, 18(2):587–593, August 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Malinowski:1982:CSI

[MTLL82] P. Malinowski, A. C. Tanner, K. F. Lee, and B. Linder. Calculation in K space of integrals arising in the theory of van der Waals forces. International Journal of Quantum Chemistry, 21(4):753–765, April 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Miyoshi:1983:ESS

[MTN83] Eisaku Miyoshi, Hiroshi Tatewaki, and Takashi Nakamura. Electronic structure of small copper clusters. International Journal of Quantum Chemistry, 23(4):1201–1207, April 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Miyoshi:1981:TSE

[MTO^{+}81] Eisaku Miyoshi, Toshikazu Takada, Shigeru Obara, Hiroshi Kashiwagi, and Kimio Ohno. Theoretical study of excitation energies of some CoF_{6}^{n−} complexes. International Journal of Quantum Chemistry, 19(3):451–461, March 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Mazurek:1983:TSA

[MTWO83] Aleksander P. Mazurek, Sid Topiol, Harel Weinstein, and Roman Osman. Theoretical studies on the activation mechanism of the histamine H_{2}-receptor: The guanidine substitution and its role in the partial agonism of N(alpha)-guanylhistamine. International Journal of Quantum Chemistry, 24(S10):293– 300, March 14–16, 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Mujica:1986:BRB

[Muj86] Vladimiro Mujica. Book review: Semi-Empirical methods of quantum chemistry. By Joanna Sadlej (1979), translated by Ian L. Cooper, Uppsala: Ellis Horwood Limited, 1985. International Journal of Quantum Chemistry, 30(3):437, September 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). See [Sad85].

Mukhopadhyay:1984:USG

[Muk84] Atri Mukhopadhyay. On the use of spin graphs for spin adapting many-body perturbation theory. International Journal of Quantum Chemistry, 25(6):965–1002, June 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Mukherjee:1986:ALC

[Muk86] Debashis Mukherjee. Aspects of linked cluster expansion in general model space many-body perturbation and coupledcluster theory. International Journal of Quantum Chemistry, 30(S20):409–435, March 10–15, 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Syposium on Atomic, Molecular, and Solid-State Theory, Scattering Problems, Many Body Phenomena, and Computational Quantum Chemistry.

Muller:1981:UIN

[Mül81] Jiri Müller. Use of iterative natural orbital method for calculating energy barrier to predissociation of the first excited states of NH_{3} (3s^{1,3} A′′_{2}). International Journal of Quantum Chemistry, 19(5):885–889, May 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Muller:1982:CIC

[Mül82] Jiri Müller. Configuration interaction calculations of the valence and the inner valence levels of the CS molecule. International Journal of Quantum Chemistry, 21(2):465–473, February 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Mola:1987:TCC

[MV87] E. E. Mola and J. L. Vicente. Theoretical calculation of the capacity of an electrode/liquid electrolyte interface. International Journal of Quantum Chemistry, 32(S21):355–362, March 12, 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Chemistry, Solid-State Theory, and Computational Methods.

Moiseyev:1980:CAR

[MW80] Nimrod Moiseyev and Frank Weinhold. Criteria of accuracy of resonance eigenvalues. International Journal of Quantum Chemistry, 17(6):1201–1211, June 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Mintmire:1983:TPC

[MW83] J. W. Mintmire and C. T. White. Theoretical photoelectron cross sections for chain polymers. International Journal of Quantum Chemistry, 24(S17):609–612, March 3–12, 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-state Theory, Collision Phenomena, and Computational Quantum Chemistry.

Mintmire:1987:BZT

[MW87] J. W. Mintmire and C. T. White. Brillouin zone treatment in total energy calculations of Peierls distorted chains. International Journal of Quantum Chemistry, 32(S21):131–136, March 12, 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Chemistry, Solid-State Theory, and Computational Methods.

Mazurek:1984:TES

[MWO^{+}84] Aleksander P. Mazurek, Harel Weinstein, Roman Osman, Sid Topiol, and Barbara Jones Ebersole. Theoretical and experimental studies of drug–receptor interactions: Determinants for recognition of 5-hydroxytryptamine analogs. International Journal of Quantum Chemistry, 26(S11):183–194, March 12– 15, 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Malathi:1981:VBP

[MY81] R. Malathi and N. Yathindra. Virtual bond probe to study ordered and random coil conformations of nucleic acids. International Journal of Quantum Chemistry, 20(1):241–257, July 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Manne:1985:MES

[MZ85] Rolf Manne and Michael C. Zerner. Matrix elements of spindependent one-electron operators between bonded functions. International Journal of Quantum Chemistry, 28(S19):165– 172, March 18–23, 1985. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Nicolaides:1983:PSC

[NA83] C. A. Nicolaides and A. N. Andriotis. Photoelectron spectroscopy of chemisorbed atoms. International Journal of Quantum Chemistry, 23(2):561–566, February 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Nagel:1982:SRR

[Nag82a] Bengt Nagel. Some results on the relation between two projectors in Hilbert space. International Journal of Quantum Chemistry, 22(S16):666, March 1–13, 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Chemistry, Theory of Condensed Matter, and Propagator Methods in the Quantum Theory of Matter.

Naghizadeh:1982:MSC

[Nag82b] J. Naghizadeh. Melting of short-chain polymers. International Journal of Quantum Chemistry, 22(S16):205–214, March 1– 13, 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Chemistry, Theory of Condensed Matter, and Propagator Methods in the Quantum Theory of Matter.

Nagel:1985:BBA

[Nag85] Bengt Nagel. Breakdown of the Born approximation in Aharonov–Bohm scattering. International Journal of Quantum Chemistry, 28(S19):741, March 18–23, 1985. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Nagy:1987:IEC

[Nag87] Á. Nagy. Ab initio exchange-correlation parameter in the Xα method. International Journal of Quantum Chemistry, 31(2): 269–278, February 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Nakatsuji:1983:CEW

[Nak83] Hiroshi Nakatsuji. Cluster expansion of the wave function. Ionization and excitation spectra of NO radical studied by the SAC and SAC–CI theory. International Journal of Quantum Chemistry, 24(S17):241–255, March 3–12, 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-state Theory, Collision Phenomena, and Computational Quantum Chemistry.

Nalewajski:1980:UVT

[Nal80] Roman F. Nalewajski. Use of the virial theorem in modeling the potential energy surfaces for triatomic collinear reactions. International Journal of Quantum Chemistry, 18 (S14):483–492, March 10–20, 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory, Collision Phenomena, Quantum Statistics, and Computational Methods.

Nalewajski:1981:E

[Nal81] Roman F. Nalewajski. Erratum. International Journal of Quantum Chemistry, 19(3):489, March 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Nagy:1987:MEF

[NÁNSP87] Péter Nagy, János G. Ángyán, Gábor Náray-Szabó, and Gustav Peinel. Molecular electrostatic fields from bond fragments. International Journal of Quantum Chemistry, 31(6):927–939, June 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Nascimento:1983:CBA

[Nas83] Marco Antonio Chaer Nascimento. Cumulative B_{K} approximation as a method to select configurations for CI calculations of transition energies. International Journal of Quantum Chemistry, 23(3):1011–1016, March 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Natiello:1986:BRB

[Nat86] Mario A. Natiello. Book review: Orbitals, terms and states. By Malcom Gerloch. Wiley, Chichester, 1986. International Journal of Quantum Chemistry, 30(6):873, December 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). See [Ger86].

Natiello:1988:BRB

[Nat88] Mario A. Natiello. Book review: Intermediate physical chemistry: Stationary properties of chemical systems. By Joseph B. Dence and J. Diestler, John Wiley & Sons, New York, 1987. International Journal of Quantum Chemistry, 33(3): 263, March 1988. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). See [DD87b].

Nicolaides:1979:ESQ

[NB79] Cleanthes A. Nicolaides and Donald R. Beck, editors. Excited states in quantum chemistry: theoretical and experimental aspects of the electronic structure and properties of the excited states in atoms, molecules and solids: proceedings of the NATO Advanced Study Institute held at Kos, Greece, June 4– 18, 1978, volume 46 of NATO advanced study institutes series: Series C, Mathematical and physical sciences. D. Reidel, Dordrecht, Boston, Lancaster, Tokyo, 1979. ISBN 90-277-0961-0. LCCN QD461.5 .N36 1978.

Natiello:1987:SCA

[NBE87] Mario A. Natiello, Erkki J. Braändas, and Alejandro R. Engelmann. Some comments on the alteration of the Coulomb bound states by a barrier perturbation. International Journal of Quantum Chemistry, 32(S21):555–562, March 12, 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Chemistry, Solid-State Theory, and Computational Methods.

Novaro:1984:MAP

[NC84] O. Novaro and S. Castillo. Multibody analysis of potential energy surfaces for first- and second-row tetramers. I. The tetrahedral structures of P_{4} and N_{4}. International Journal of Quantum Chemistry, 26(3):411–423, September 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Nogueira:1987:SRT

[NC87] José Carlos Nogueira and José Eduardo Chaguri. Semiempirical relationship for the total cross section for electron scattering in atoms and molecules. International Journal of Quantum Chemistry, 32(S21):307–311, March 12, 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Chemistry, Solid-State Theory, and Computational Methods.

Niesar:1989:PON

[NCH^{+}89] U. Niesar, G. Corongiu, M.-J. Huang, M. Dupuis, and E. Clementi. Preliminary observations on a new water–water potential. International Journal of Quantum Chemistry, 36 (S23):421–443, April 1–8, 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Novaro:1981:TBP

[NCKL81] O. Novaro, S. Castillo, W. Kołos, and A. Leś. Three-body potential energy terms for methane trimers. International Journal of Quantum Chemistry, 19(4):637–648, April 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Nguyen-Dang:1982:SPL

[NDBE82] T. T. Nguyen-Dang, R. F. W. Bader, and H. Essén. Some properties of the Lagrange multiplier μ in density functional theory. International Journal of Quantum Chemistry, 22(5): 1049–1058, November 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Neckel:1983:RIE

[Nec83] A. Neckel. Recent investigations on the electronic structure of the fourth and fifth group transition metal monocarbides, mononitrides, and monoxides. International Journal of Quantum Chemistry, 23(4):1317–1353, April 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Nesbet:1985:ILE

[Nes85] R. K. Nesbet. The influence of localized excitations on spinselective electron scattering and Auger emission in transition metals. International Journal of Quantum Chemistry, 28 (S19):383–391, March 18–23, 1985. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Nesbet:1987:EPC

[Nes87] R. K. Nesbet. Electron pair correlation in atoms and molecules. International Journal of Quantum Chemistry, 31 (3):435–443, March 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Newton:1980:FEE

[New80] Marshall D. Newton. Formalisms for electron-exchange kinetics in aqueous solution and the role of ab initio techniques in their implementation. International Journal of Quantum Chemistry, 18(S14):363–391, March 10–20, 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory, Collision Phenomena, Quantum Statistics, and Computational Methods.

Nakatsuji:1981:CEW

[NH81] Hiroshi Nakatsuji and Kimihiko Hirao. Cluster expansion of the wave function. Electron correlations in singlet and triplet excited states, ionized states, and electron attached states by SAC and SAC–CI theories. International Journal of Quantum Chemistry, 20(6):1301–1313, December 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Nagakura:1984:EMF

[NH84] Saburo Nagakura and Hisaharu Hayashi. External magnetic field effects upon photochemical reactions. International Journal of Quantum Chemistry, 26(S18):571–578, March 1–15, 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Atomic, Molecular and Solid-State Theory, and Computational Quantum Chemistry.

Nakatsuji:1983:FST

[NHKY83] H. Nakatsuji, M. Hada, K. Kanda, and T. Yonezawa. Force in SCF Theories. First and second derivatives of the potential energy hypersurface of chemical reaction systems. International Journal of Quantum Chemistry, 23(2):387–397, February 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Niukkanen:1980:GME

[Niu80] A. W. Niukkanen. Gaussian matrix elements of the operator X: Reduction to confluent hypergeometric functions. International Journal of Quantum Chemistry, 18(3):695–708, September 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Niukkanen:1984:FTAa

[Niu84a] A. W. Niukkanen. Fourier transforms of atomic orbitals. I. Reduction to four-dimensional harmonics and quadratic transformations. International Journal of Quantum Chemistry, 25 (6):941–955, June 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Niukkanen:1984:FTAb

[Niu84b] A. W. Niukkanen. Fourier transforms of atomic orbitals. II. Convolution theorems. International Journal of Quantum Chemistry, 25(6):957–964, June 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Nicolaides:1981:TCR

[NKM81] C. A. Nicolaides, Y. Komninos, and Th. Mercouris. Theory and calculation of resonances using complex coordinates. International Journal of Quantum Chemistry, 20(S15):355–367, March 8–14, 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory, Collision Phenomena, and Computational Quantum Chemistry.

Novoselsky:1989:SMC

[NKV^{+}89] Akiva Novoselsky, Jacob Katriel, Michel Valliègres, Jin-Quan Chen, and Robert Gilmore. Shell model calculations: An efficient new algorithm. International Journal of Quantum Chemistry, 36(3):325–339, September 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Nalewajski:1988:MHS

[NKZ88] Roman F. Nalewajski, Jacek Korchowiec, and Zhongxiang Zhou. Molecular hardness and softness parameters and their use in chemistry. International Journal of Quantum Chemistry, 34(S22):349–366, March 12–19, 1988. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Syposium on Quantum Chemistry, Solid-State Theory, and Computational Methods.

Nakamura:1989:ISS

[NL89] Y. Nakamura and S. Lunell. Ab initio SCF study of the interaction between L-serine phosphate and ammonia. International Journal of Quantum Chemistry, 36(5):587–598, November 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Nicolaides:1984:MET

[NMK84] C. A. Nicolaides, Th. Mercouris, and Y. Komninos. Manyelectron theory of autoionizing states using complex coordinates: The position and the partial and total widths of the Ne^{+} 1s hole state. International Journal of Quantum Chemistry, 26(6):1017–1027, December 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Noack:1980:NMAa

[Noa80a] Wolf-Eckart Noack. New method for approximate Hartree– Fock calculations using density approximations and Coulomb field corrections. I. International Journal of Quantum Chemistry, 17(6):1111–1123, June 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Noack:1980:NMAb

[Noa80b] Wolf-Eckart Noack. New method for approximate Hartree– Fock calculations using density approximations and Coulomb field corrections. II. International Journal of Quantum Chemistry, 17(6):1125–1141, June 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Nomura:1980:PES

[Nom80] Okio Nomura. Potential-energy surfaces of low-lying states of HNO. International Journal of Quantum Chemistry, 18 (1):143–150, July 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Nose:1980:RTM

[NOMK80] Shuichi Nosé, Yoshiaki Ozaki, Kazuo Maki, and Yosuke Kataoka. Rotational tunneling of methane on the surface of graphite. International Journal of Quantum Chemistry, 18 (2):641–644, August 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Norcross:1983:REM

[Nor83] David W. Norcross. Review of electron-molecule scattering. International Journal of Quantum Chemistry, 24(S17):191, March 3–12, 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-state Theory, Collision Phenomena, and Computational Quantum Chemistry.

Novaro:1983:TBN

[Nov83a] O. Novaro. Three-body nonadditive effects for hydrogen bonded water molecules: SCF study of hydrated species including H_{2}S and various metal ions. International Journal of Quantum Chemistry, 23(4):1611–1625, April 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Novosadov:1983:HLA

[Nov83b] Boris K. Novosadov. Hydrogen-like atomic orbitals: Addition and expansion theorems, integrals. International Journal of Quantum Chemistry, 24(1):1–18, July 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Nyden:1980:PCG

[NP80] Marc R. Nyden and G. A. Petersson. Pairwise correlated generalized valence bond model of electronic structure. VII. The b^{3}Σ state of H_{2}. International Journal of Quantum Chemistry, 17(5):975–982, May 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Nalewajski:1981:NKF

[NP81] Roman F. Nalewajski and Roman Pastewski. Normalized kinetic field potentials for the atom-diatom reactions. Testing the collinear surfaces. International Journal of Quantum Chemistry, 20(S15):595–610, March 8–14, 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory, Collision Phenomena, and Computational Quantum Chemistry.

Neto:1989:HCS

[NPL89a] J. J. Soares Neto, Søren B. Padkjér, and Jan Linderberg. Hyperspherical coordinates in the study of rovibrational levels of H_{3}^{+} and its isotopomers. International Journal of Quantum Chemistry, 36(S23):127–136, April 1–8, 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Nevalainen:1989:DMA

[NPL89b] Vesa Nevalainen, Tapani A. Pakkanen, and Marina Lindblad. Density matrix approximation as a probe to the interaction range analysis of electronic structure. International Journal of Quantum Chemistry, 35(4):471–481, April 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Naray-Szabo:1982:EEC

[NS82a] G. Náray-Szabó. Electrostatic effect on catalytic rate enhancement in serine proteinases. International Journal of Quantum Chemistry, 22(3):575–582, September 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Nielsen:1982:CRT

[NS82b] Egon S. Nielsen and Jack Simons. Coordinate rotated TDHF excitation energies Li^{−} ^{1}S → ^{1}P. International Journal of Quantum Chemistry, 22(2):275–288, August 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Naray-Szabo:1983:ULE

[NS83] Gábor Náray-Szabó. Unusually large electrostatic field effect of the buried aspartate in serine proteinases: Source of catalytic power. International Journal of Quantum Chemistry, 23(2):723–728, February 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Natiello:1984:CPH

[NS84] Mario A. Natiello and Gustavo E. Scuseria. Convergence properties of Hartree–Fock SCF molecular calculations. International Journal of Quantum Chemistry, 26(6):1039–1049, December 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Naray-Szabo:1986:TCB

[NS86] Gábor Náray-Szabó, editor. Theoretical chemistry of biological systems, volume 41 of Studies in physical and theoretical chemistry. Elsevier, Amsterdam, The Netherlands, 1986. ISBN 0-444-42597-7 (US). xiii + 504 pp. LCCN QP517.M3 T47 1986.

Naray-SzabO:1989:ECA

[NS89] Gábor Náray-SzabÓ. Electrostatics in computer-aided drug design. International Journal of Quantum Chemistry, 36(S16): 87–99, April 1–8, 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Ni:1988:IPB

[NSL88] Xiangshan Ni, Xiufan Shi, and Lunjiang Ling. An interaction potential between an alanine zwitterion and a water molecule based on ab initio calculations. International Journal of Quantum Chemistry, 34(6):527–533, December 1988.

CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Naray-Szabo:1989:ECM

[NSN89] G. Náray-Szabó and P. Nagy. Electrostatic complementarity in molecular aggregates. 9: Protein-ligand complexes. International Journal of Quantum Chemistry, 35(1):215–221, January 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Naaray-Szaboa:1980:EEE

[NSP80] G. Náaray-Szabóa and L. Polgár. Effect of environment on enzymatic action: His-64⋯asp-32 proton transfer in subtilisin. International Journal of Quantum Chemistry, 18(S7): 397–402, March 5–8, 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Nath:1986:SBA

[NSS86] S. Nath, K. Shobha, and K. D. Sen. A simple bound to average electronic density and energy for atoms within an isoelectronic series. International Journal of Quantum Chemistry, 29(1): 165–167, January 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Nicolaides:1980:FAM

[NT80] C. A. Nicolaides and G. Theodorakopoulos. FOTOS applied to molecules: Oscillator strengths in H_{2}O. International Journal of Quantum Chemistry, 18(S14):315–322, March 10–20, 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory, Collision Phenomena, Quantum Statistics, and Computational Methods.

Nagamura:1980:WIP

[NTM80] Toshihiko Nagamura, Takashi Takayanagi, and Taku Matsuo. Weak interactions at the photoexcited states of partially oriented aromatic chromophores in man-made bilayer membranes. International Journal of Quantum Chemistry, 18(2): 509–513, August 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Nichols:1982:ESI

[NW82] David E. Nichols and Herschel J. R. Weintraub. Effects of steric interactions on models of charge-transfer complexes: A preliminary report. International Journal of Quantum Chemistry, 22(S9):205–211, March 4–6, 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Obcemea:1984:RTO

[Obc84] Ch. Obcemea. Response theory of open quantum systems. International Journal of Quantum Chemistry, 26(S18):33–41, March 1–15, 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Atomic, Molecular and Solid-State Theory, and Computational Quantum Chemistry.

Obcemea:1987:FED

[Obc87] Ceferino H. Obcemea. Free energy and dispersion forces via response functions. International Journal of Quantum Chemistry, 31(1):113–117, January 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Ovchinnikov:1982:PIV

[OBMU82] A. A. Ovchinnikov, A. E. Belinskii, I. A. Misurkin, and I. I. Ukrainskii. Peierls instability and vibrational spectra of polyacetylene. International Journal of Quantum Chemistry, 22 (4):761–774, October 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Oddershede:1982:RBP

[Odd82] Jens Oddershede. On the relations between the polarization propagator, the principal propagator, and the particle-hole propagator. International Journal of Quantum Chemistry, 22 (S16):583–589, March 1–13, 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Chemistry, Theory of Condensed Matter, and Propagator Methods in the Quantum Theory of Matter.

Odiot:1980:DMD

[OF80] J. Bridet Et S. Odiot and S. Fliszar. Dynamique moléculaire du cristal ammoniac. Calcul ab initio SCF de l’énergie electronique. (French) [molecular dynamics of the ammonium crystal. SCF ab initio calculation of the electronic energy]. International Journal of Quantum Chemistry, 18(3):675–688, September 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Obcemea:1981:GVR

[OFB81] Ch. Obcemea, P. Froelich, and E. J. Brändas. Generalized virial relations and the theory of subdynamics. International Journal of Quantum Chemistry, 20(S15):695–702, March 8– 14, 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory, Collision Phenomena, and Computational Quantum Chemistry.

Ouamerali:1986:EMC

[OG86] Ourida Ouamerali and Jose Gayoso. Étude MNDO des composés conjugués enneagonaux. (French) [MNDO study of n-ary conjugated composites]. International Journal of Quantum Chemistry, 29(5):1599–1624, May 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Ogawa:1980:NSC

[Oga80] Tohru Ogawa. New self-consistent theory of atoms: Toward a unified theory of atoms and atomic aggregates. International Journal of Quantum Chemistry, 18(1):31–35, July 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Osapay:1989:SBS

[ÖHD89] K. Ösapay, L. Hevesi, and J. Delhalle. A study of the basis set dependence of the geometry and proton affinity for a series of vinyl ethers and vinyl sulfides. International Journal of Quantum Chemistry, 36(S23):453–463, April 1–8, 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Ohno:1984:PMK

[Ohn84] Kimio Ohno. Professor Masao Kotani. International Journal of Quantum Chemistry, 26(S18):1–9, March 1–15, 1984.

CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Atomic, Molecular and Solid-State Theory, and Computational Quantum Chemistry.

Ohrn:1985:ABP

[Öhr85] Yngve Öhrn. The AGP-based polarization propagator. International Journal of Quantum Chemistry, 28(S19):39–50, March 18–23, 1985. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Ohrn:1986:ABP

[Öhr86] Yngve Öhrn. The agp-based polarization propagator. International Journal of Quantum Chemistry, 30(3):445–447, September 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Ohrn:1987:I

[Öhr87] Yngve Öhrn. Introduction. International Journal of Quantum Chemistry, 31(3):333, March 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Ohno:1986:BSD

[OI86] Koichi Ohno and Toshimasa Ishida. Basis-set dependence of exterior electron distributions of molecular orbitals. International Journal of Quantum Chemistry, 29(4):677–688, April 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Ojha:1989:ADS

[Ojh89] P. C. Ojha. Analysis of degeneracy in the spectrum of a class of molecular graphs. International Journal of Quantum Chemistry, 35(5):687–700, May 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Olsen:1983:SHO

[OJY83] Jeppe Olsen, Poul Jørgensen, and Danny L. Yeager. Secondand higher-order convergence in linear and nonlinear multiconfigurational Hartree–Fock theory. International Journal of Quantum Chemistry, 24(1):25–60, July 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Osanai:1980:IIA

[OK80] You Osanai and Hiroshi Kashiwagi. Improvement of an integral approximation scheme based on semiorthogonalized orbitals. International Journal of Quantum Chemistry, 17(5): 1031–1037, May 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Ondrechen:1985:MSB

[OKZ85] Mary Jo Ondrechen, Jaeju Ko, and Li-Tai Zhang. Models for the spectra of bridged mixed-valence dimers. International Journal of Quantum Chemistry, 28(S19):393–401, March 18– 23, 1985. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Otto:1980:CCT

[OL80] P. Otto and J. Ladik. Concept of charge transfer: Investigation of charge transfer in the glyoxal–formamide, water dimer, and water–lithium cation systems. International Journal of Quantum Chemistry, 18(5):1143–1155, November 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Otto:1982:III

[OL82] P. Otto and J. Ladik. Intermolecular and intramolecular interactions calculated with ab initio perturbative configuration interaction method using strongly localized orbitals. International Journal of Quantum Chemistry, 22(1):169–182, July 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Ohrn:1983:IPI

[ÖL83] Yngve Öhrn and Per-Olov Löwdin. Introduction, Part II. International Journal of Quantum Chemistry, 24(S17):199, March 3–12, 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-state Theory, Collision Phenomena, and Computational Quantum Chemistry.

Oie:1982:ISC

[OLB^{+}82] Tetsuro Oie, Gilda H. Loew, Stanley K. Burt, J. Stephen Binkley, and Robert D. MacElroy. Ab initio study of catalyzed and uncatalyzed amide bond formation as a model for peptide bond formation: Ammonia-formic acid and ammonia-glycine reactions. International Journal of Quantum Chemistry, 22 (S9):223–245, March 4–6, 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Lowdin:1981:SDC

[oLCC81] Per olov Löwdin, Jean-Louis Calais, and Jacques M. Calazans. Simple derivation of conditions for instability in the Hartree– Fock and projected Hartree–Fock schemes. International Journal of Quantum Chemistry, 20(6):1201–1215, December 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Olson:1983:ICC

[Ols83] R. E. Olson. Ionic–covalent collision interactions. International Journal of Quantum Chemistry, 24(S17):49–64, March 3–12, 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-state Theory, Collision Phenomena, and Computational Quantum Chemistry.

Olszewski:1986:AMQ

[Ols86] S. Olszewski. Angular-momentum quantization applied to the Thomas–Fermi atom. International Journal of Quantum Chemistry, 30(4):571, October 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Oliveira:1981:MXM

[OM81] P. Murilo Oliveira and B. Maffeo. MS X_{α} method and transition-state procedure. International Journal of Quantum Chemistry, 20(5):1127–1139, November 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Ohno:1982:QCLa

[OM82a] Kimio Ohno and Keiji Morokuma. Quantum chemistry literature data base: bibliography of ab initio calculations for 1978– 1980, volume 12 of Physical sciences data. Elsevier, Amsterdam, The Netherlands, 1982. ISBN 0-444-42074-6 (vol. 12), 0-444-41689-7 (ensemble). ix + 459 pp. LCCN Z5524.Q35 O38 1982.

Ohno:1982:QCLb

[OM82b] Kimio Ohno and Keiji Morokuma, editors. Quantum chemistry literature data base: supplement 1: bibliography of ab initio calculations for 1981. See: Journal of Molecular Structure. Theochem: Vol. 91, Nos. 1–2, 1982. Elsevier, Amsterdam, The Netherlands, 1982. ISSN 0022-2860. 252 pp. LCCN ????

Olson:1982:ESR

[OM82c] John A. Olson and David A. Micha. Electronic state representations at molecular potential pseudocrossings. International Journal of Quantum Chemistry, 22(5):971–988, November 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Ohno:1983:QCL

[OM83a] K. (Kimio) Ohno and K. (Keiji) Morokuma, editors. Quantum chemistry literature data base: Supplement 2, Bibliography of ab initio calculations for 1982, volume 15 of Theochem. Elsevier, Amsterdam, The Netherlands, 1983. ISSN 0166-1280 (print), 1872-7999 (electronic). 215 pp. LCCN Z5524.Q35 Q35 1983.

Ohno:1982:QCLc

[OM83b] Kimio Ohno and Keiji Morokuma, editors. Quantum chemistry literature data base. Supplement 1: bibliography of ab initio calculations for 1981, volume 91, no. 1/2; vol. 8 of Journal of molecular structure; Theochem, 0166-1280. Elsevier, Amsterdam, The Netherlands, 1982-1983. ISSN 0022-2860. 252 pp. LCCN ????

Ohno:1984:QCL

[OM84] Kimio Ohno and Keiji Morokuma, editors. Quantum chemistry literature data base: Supplement 3: bibliography of ab initio calculations for 1983. Elsevier, Amsterdam, The Netherlands, 1984. 229 pp. LCCN ????

Ohno:1985:QCL

[OM85] Kimio Ohno and Keiji Morokuma, editors. Quantum chemistry literature data base: Supplement 4: bibliography of ab initio calculations for 1984. Elsevier, Amsterdam, The Netherlands, 1985. 298 pp. LCCN ????

Ohno:1986:QCL

[OM86] Kimio Ohno and Keiji Morokuma, editors. Quantum chemistry literature data base: supplement 5: bibliography of ab initio calculations for 1985. Elsevier, Amsterdam, The Netherlands, 1986. 525 pp. LCCN ????

Oie:1982:SCS

[OMC82] Tetsuro Oie, Gerald M. Maggiora, and Ralph E. Christoffersen. Structural characterization of a special-pair chlorophyll dimer model of P700. International Journal of Quantum Chemistry, 22(S9):157–171, March 4–6, 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Oie:1981:DFI

[OMCD81] Tetsuro Oie, Gerald M. Maggiora, Ralph E. Christoffersen, and David J. Duchamp. Development of a flexible intra- and intermolecular empirical potential function for large molecular systems. International Journal of Quantum Chemistry, 20 (S8):1–47, March 5–7, 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Ohno:1989:QCL

[OMH89] K. Ohno, K. Morokuma, and H. Hosoya, editors. Quantum chemistry literature data base: bibliography of ab initio calculations for 1988. See: Journal of Molecular Structure. Theochem: V. 203, 1989. Special issue. ????, ????, 1989. LCCN ????

Odiot:1986:MTC

[OPSO86] S. Odiot, M. Peyrard, J. Schnur, and E. Oran. Molecular theory and cooperative mechanism of shock waves-induced detonations in energetic molecular crystals. International Journal of Quantum Chemistry, 29(5):1625–1634, May 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Odiot:1987:MTC

[OPSO87] S. Odiot, M. Peyrard, J. Schnur, and E. Oran. Molecular theory and cooperative mechanism of shock waves-induced detonations in energetic molecular crystals. International Journal of Quantum Chemistry, 31(2):331, February 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Oriade:1980:BQM

[Ori80] Joseph Oriade. Biquadratic quasianalytic methods for Brillouin zone integration. International Journal of Quantum Chemistry, 18(S14):597–605, March 10–20, 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory, Collision Phenomena, Quantum Statistics, and Computational Methods.

Oriade:1981:FSB

[Ori81a] Joseph Oriade. The Fermi surface of beryllium. International Journal of Quantum Chemistry, 19(5):721–727, May 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Oriade:1981:PBS

[Ori81b] Joseph Oladipo Oriade. Positron band structures for beryllium. International Journal of Quantum Chemistry, 20(4): 891–896, October 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Oriade:1982:EPA

[Ori82a] J. O. Oriade. Evaluation of positron annihilation in Beryllium. International Journal of Quantum Chemistry, 22(1): 49–67, July 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Oriade:1982:PRW

[Ori82b] J. O. Oriade. Product representation weights for Brillouin zone integrals in the biquadratic limit. International Journal of Quantum Chemistry, 22(S16):481–484, March 1–13, 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Chemistry, Theory of Condensed Matter, and Propagator Methods in the Quantum Theory of Matter.

Oriade:1982:ESB

[Ori82c] Joseph Oriade. Electronic structure of beryllium. International Journal of Quantum Chemistry, 21(2):359–367, February 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Oriade:1982:ESE

[Ori82d] Joseph Oriade. Evaluation of a self-energy graph for the homogeneous electron gas. International Journal of Quantum Chemistry, 22(S16):661–664, March 1–13, 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Chemistry, Theory of Condensed Matter, and Propagator Methods in the Quantum Theory of Matter.

Oriade:1984:BZD

[Ori84a] J. O. O. Oriade. Brillouin-zone double integrals. International Journal of Quantum Chemistry, 26(4):457–461, October 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Oriade:1984:SCF

[Ori84b] Joseph Oriade. Solution of a correlation function’s defining integral equation for complex domains. International Journal of Quantum Chemistry, 26(6):987–992, December 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Oriade:1985:BZD

[Ori85a] J. O. O. Oriade. Brillouin zone double integrals. International Journal of Quantum Chemistry, 27(2):109–113, February 1985. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Oriade:1985:WLA

[Ori85b] Joseph Oriade. Warped linear analytic tetrahedral scheme for Brillouin zone integrations. International Journal of Quantum Chemistry, 27(1):97–101, January 1985. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Ortiz:1987:AEP

[Ort87] J. V. Ortiz. Applying electron propagator theory to electron affinities. International Journal of Quantum Chemistry, 32 (S21):469–473, March 12, 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Chemistry, Solid-State Theory, and Computational Methods.

Ortiz:1988:PFO

[Ort88] J. V. Ortiz. Partial fourth order electron propagator theory. International Journal of Quantum Chemistry, 34(S22): 431–436, March 12–19, 1988. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Syposium on Quantum Chemistry, Solid-State Theory, and Computational Methods.

Ortiz:1989:EPC

[Ort89] J. V. Ortiz. Electron propagator calculations with nondiagonal partial fourth-order self-energies and unrestricted Hartree–Fock reference states. International Journal of Quantum Chemistry, 36(S23):321–332, April 1–8, 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Oddershede:1989:OIK

[OS89a] Jens Oddershede and John R. Sabin. On the orbital implementation of the kinetic theory of stopping. International Journal of Quantum Chemistry, 36(S23):557–563, April 1–8, 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Oneal:1989:ACL

[OS89b] Douglas W. O’neal and Jack Simons. Application of Choleskylike matrix decomposition methods to the evaluation of atomic orbital integrals and integral derivatives. International Journal of Quantum Chemistry, 36(6):673–688, December 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Ohrn:1989:Ia

[ÖSZ89a] N. Y. Öhrn, J. R. Sabin, and M. C. Zerner. Introduction. International Journal of Quantum Chemistry, 36(S16):v, April 1–8, 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Ohrn:1989:Ib

[ÖSZ89b] N. Y. Öhrn, J. R. Sabin, and M. C. Zerner. Introduction. International Journal of Quantum Chemistry, 36(S23):IX–X, April 1–8, 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Okido:1980:TBP

[OT80] Sadao Okido and Osamu Tanimoto. Theory of the Bose particle representation for geminals. International Journal of Quantum Chemistry, 18(1):211–217, July 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Otto:1986:SSD

[Ott86] P. Otto. On the stability of single- and double-stranded DNA helices: The application of the PPT-MCF method on large fragments of DNA. International Journal of Quantum Chemistry, 30(2):275–288, August 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Ovchinnikov:1981:IC

[Ovc81] Yu. A. Ovchinnikov. Ionophores and channels. International Journal of Quantum Chemistry, 20(2):461–478, August 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Ortiz:1981:AWR

[OWÖ81] J. V. Ortiz, Brian Weiner, and Yngve Öhrn. The AGP wavefunction and its relation to other descriptions of electronic structure. International Journal of Quantum Chemistry, 20 (S15):113–128, March 8–14, 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory, Collision Phenomena, and Computational Quantum Chemistry.

Olsen:1983:EHL

[OY83] Jeppe Olsen and Danny L. Yeager. Effective handling of a large configuration state function expansion for an MCSCF state and improved efficiency for two-electron integral transformations in MCSCF. International Journal of Quantum Chemistry, 23(3):789–809, March 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Osamura:1980:QQC

[OYH^{+}80] Y. Osamura, S. Yamabe, F. Hirota, H. Hosoya, S. Iwata, H. Kashiwagi, K. Morokuma, M. Togasi, S. Obara, K. Tanaka, and K. Ohno. QCLDB —Quantum chemistry literature data base—a trial. International Journal of Quantum Chemistry, 18(2):393–396, August 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Osamura:1980:MSP

[OYN80] Yoshihiro Osamura, Shinichi Yamabe, and Kichisuke Nishimoto. MO Study of the photochemical behavior of the imine bond. International Journal of Quantum Chemistry, 18(2): 457–462, August 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Ozkan:1984:GHO

[Özk84] İlker Özkan. General harmonic oscillator integrals by the operator method. International Journal of Quantum Chemistry, 26(3):339–347, September 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Piela:1980:LRI

[PABD80] Lucjan Piela, Jean-Marie André, Jean-Luc Brédas, and Joseph Delhalle. Long-range interactions in periodic helices. A method for accurate calculation at the Hartree–Fock level. International Journal of Quantum Chemistry, 18(S14):405–418, March 10–20, 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory, Collision Phenomena, Quantum Statistics, and Computational Methods.

Pernot:1988:RDV

[PABL88] P. Pernot, O. Atabek, J. A. Beswick, and B. Levy. Reduceddiabatic vibrational close coupled treatment of molecular dissociation dynamics. International Journal of Quantum Chemistry, 34(2):161–184, August 1988. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Pack:1982:TIE

[Pac82] George R. Pack. Theory of ion effects on DNA conformation: A high-salt model. International Journal of Quantum Chemistry, 22(S9):81–101, March 4–6, 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Pack:1983:DDE

[Pac83] George R. Pack. Determination of the distribution and electric fields of monovalent and divalent ions about DNA. International Journal of Quantum Chemistry, 24(S10):61–71, March 14–16, 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Paolillo:1984:SFP

[PAFW84] L. Paolillo, S. Andini, L. Ferrara, and S. Wurzburger. Structure and function of protamines: An investigation of clupeine properties by nuclear magnetic resonance. International Journal of Quantum Chemistry, 26(5):873–888, November 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Palma:1989:SSD

[Pal89] M. U. Palma. Solvent structure and dynamics: How relevant to molecular and quantum pharmacology? International Journal of Quantum Chemistry, 35(1):125–139, January 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Panda:1981:ATL

[Pan81] B. C. Panda. Active transport in the light of thermodynamics of open systems. International Journal of Quantum Chemistry, 20(2):567–571, August 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Panin:1982:MRCa

[Pan82a] A. I. Panin. Method of recurrent construction of Löwdin spinadapted wave functions. I. Addition and subtraction operators. International Journal of Quantum Chemistry, 22(3):

557–573, September 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Panin:1982:MRCb

[Pan82b] A. I. Panin. Method of recurrent construction of Löwdin spinadapted wave functions. II. Local representation of creation– annihilation operators. International Journal of Quantum Chemistry, 22(6):1177–1188, December 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Panin:1983:MRC

[Pan83] A. I. Panin. Method of recurrent construction of Löwdin spinadapted wave functions. III. Löwdin basis and its permutation symmetry: Evaluation of overlap integrals. International Journal of Quantum Chemistry, 24(3):279–306, September 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Panin:1984:MRC

[Pan84] A. I. Panin. Method of recurrent construction of Löwdin spinadapted wave functions. IV. Matrix element evaluation problem. International Journal of Quantum Chemistry, 25(5):767– 793, May 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Panin:1985:BLA

[Pan85] A. I. Panin. “basis” Lie algebra of electronic Fock space: Application to evaluation of matrix elements of spin tensor operators. International Journal of Quantum Chemistry, 27 (5):501–525, May 1985. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Panin:1989:COE

[Pan89] A. I. Panin. Contraction operator over electronic Fock space. II. Weak representability problem. International Journal of Quantum Chemistry, 35(5):629–647, May 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Parr:1984:RCA

[Par84] Robert G. Parr. Remarks on the concept of an atom in a molecule and on charge transfer between atoms on molecule formation. International Journal of Quantum Chemistry, 26 (5):687–692, November 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Pauncz:1979:SEC

[Pau79] Ruben Pauncz. Spin Eigenfunctions: Construction and Use. Plenum Press, New York, NY, USA; London, UK, 1979. ISBN 0-306-40141-X. xv + 370 pp. LCCN QD462.P38.

Pauncz:1981:SCI

[Pau81] Ruben Pauncz. Studies in the configuration interaction method. I. Combinatorial aspects of the downward-Robb algorithm. International Journal of Quantum Chemistry, 20(S15): 101–111, March 8–14, 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory, Collision Phenomena, and Computational Quantum Chemistry.

Pauncz:1984:NDW

[Pau84] Ruben Pauncz. New derivation of the Waller–Hartree–Fock spatial wave function. International Journal of Quantum Chemistry, 25(4):707–713, April 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Pauncz:1989:NDS

[Pau89a] Ruben Pauncz. New derivation of the Sasaki–Ohno formula of the Sanibel coefficients. International Journal of Quantum Chemistry, 36(5):669–670, November 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Pauncz:1989:WHD

[Pau89b] Ruben Pauncz. The Waller–Hartree double determinant in quantum chemistry. International Journal of Quantum Chemistry, 35(6):717–719, June 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Payne:1982:MEF

[Pay82] Philip W. Payne. Matrix element factorization in the unitary group approach for configuration interaction calculations. International Journal of Quantum Chemistry, 22(6):1085–1152, December 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Paldus:1982:CAF

[PB82] J. Paldus and M. J. Boyle. Cluster analysis of the full configuration interaction wave functions of cyclic polyene models. International Journal of Quantum Chemistry, 22(6):1281–1305, December 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Pluta:1988:NHF

[PBA88] Tadeusz Pluta, Rodney J. Bartlett, and Ludwik Adamowicz. Numerical Hartree–Fock characterization of metastable states of the He anion. International Journal of Quantum Chemistry, 34(S22):225–230, March 12–19, 1988. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Syposium on Quantum Chemistry, Solid-State Theory, and Computational Methods.

Pop:1989:TAC

[PBBK89] Emil Pop, Marcus E. Brewster, Nicholas Bodor, and J. J. Kaminski. Theoretical aspects of cephalosporin isomerism. International Journal of Quantum Chemistry, 36(S16):291–300, April 1–8, 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Pavato:1980:MSM

[PBFL80] A. C. Pavãto, M. Braga, A. Fazzio, and J. R. Leite. Multiple scattering mass operator method for molecular orbital calculations. International Journal of Quantum Chemistry, 18(5): 1165–1173, November 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Pucci:1980:TMD

[PBGT80] Renato Pucci, Marcello Baldo, Antonio Grassi, and Pasquale Tomasello. Toward a microscopic description of molecular electronic systems. International Journal of Quantum Chemistry, 18(1):11–18, July 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Pop:1989:TSS

[PBKB89] Emil Pop, Marcus E. Brewster, J. J. Kaminski, and Nicholas Bodor. Theoretical study of some heterocyclic amines with applications to the chemistry of 9-amino-1,2,3,4-tetrahydroacridine. International Journal of Quantum Chemistry, 35(2):315–324, February 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Petke:1985:IQM

[PBM85] J. D. Petke, Paul Butler, and Gerald M. Maggiora. Ab initio quantum-mechanical characterization of the electronic states of anthraquinone, quinizarin, and 1,4-diamino anthraquinone. International Journal of Quantum Chemistry, 27(1):71–87, January 1985. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Pucci:1984:DGE

[PBMR^{+}84] R. Pucci, M. Baldo, A. Martin-Rodero, G. Piccitto, and P. Tomasello. Dimensional and geometrical effects on the electronic structure of polycyclic hydrocarbons. International Journal of Quantum Chemistry, 26(5):783–791, November 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Paldus:1983:BLAa

[PC83a] J. Paldus and E. Chin. Bond length alternation in cyclic polyenes. I. Restricted Hartree–Fock method. International Journal of Quantum Chemistry, 24(4):373–394, October 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Pedrini:1983:ESL

[PC83b] C. Pedrini and H. Chermette. Electronic structure and luminescent properties of Cu^{+} and Ag^{+} impurity centers in NaCl. International Journal of Quantum Chemistry, 23(3): 1025–1032, March 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Purvis:1986:UDG

[PC86] George D. Purvis III and Chris Culberson. Using density gradients to map atomic structure on molecular isodensities. International Journal of Quantum Chemistry, 30(S13):261–265, March 6–8, 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Pacios:1988:CVP

[PC88] L. Fernandez Pacios and P. Gomez Calzada. Core/valence partition and relativistic effects in effective potentials for transition metals. International Journal of Quantum Chemistry, 34 (3):267–277, September 1988. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Padilla:1989:TSS

[PCA89] L. Padilla, R. Contreras, and A. Aizman. Theoretical study of solvent effects in the 1,3 rearrangement in the allyl–fluoride system in polar media. International Journal of Quantum Chemistry, 36(S23):483–488, April 1–8, 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Paldus:1983:BLAb

[PCG83] J. Paldus, E. Chin, and M. G. Grey. Bond length alternation in cyclic polyenes. II. Unrestricted Hartree–Fock method. International Journal of Quantum Chemistry, 24(4):395–409, October 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Pellegatti:1982:NEC

[PČP82] Alain Pellegatti, Jiří Čížek, and Josef Paldus. Numerical estimates of the convergence of the Rayleigh–Schrödinger perturbation expansions for the energy levels of various models of the benzene molecule. International Journal of Quantum Chemistry, 21(1):147–151, January 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Pisani:1980:EEH

[PD80] C. Pisani and R. Dovesi. Exact-exchange Hartree–Fock calculations for periodic systems. I. Illustration of the method. International Journal of Quantum Chemistry, 17(3):501–516, March 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Phillips:1983:CPH

[PD83] Philip Phillips and Ernest R. Davidson. Chemical potential for harmonically interacting particles in a harmonic potential. International Journal of Quantum Chemistry, 23(1):185–194, January 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Pelissier:1984:BAM

[PD84a] Michel Pelissier and Ernest R. Davidson. Bonding in alkali metal homonuclear diatomics. International Journal of Quantum Chemistry, 25(4):723–731, April 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Pelissier:1984:FOE

[PD84b] Michel Pelissier and Ernest R. Davidson. Frozen orbital effects in the computation of excitation energies of the iron atom. International Journal of Quantum Chemistry, 25(3):483–491, March 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Patnaik:1985:CPD

[PD85] Lalit N. Patnaik and Sarojini Das. Conformation of pdimethylaminobenzylidene-p-nitroaniline, p-nitrobenzylidene-p-dimethylaminoaniline, their stilbene and azobenzene derivatives. International Journal of Quantum Chemistry, 27(2): 135–144, February 1985. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Patnaik:1987:CPD

[PD87] Lalit N. Patnaik and Sarojini Das. Conformation of pdimethylamino aza styryl dyes. International Journal of Quantum Chemistry, 32(5):595–612, November 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Priyadarshy:1989:TEE

[PD89] Satyam Priyadarshy and Sambhu Nath Datta. Transport of excitation energy in a three-dimensional doped molecular crystal. International Journal of Quantum Chemistry, 36(2):141– 155, August 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Poirier:1980:UQGb

[PDC80] R. A. Poirier, R. Daudel, and I. G. Csizmadia. Uniform quality Gaussian basis sets for molecular calculations. II. C_{2} hydrocarbons. International Journal of Quantum Chemistry, 18(3): 727–733, September 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Poirier:1981:UQGa

[PDC81a] R. A. Poirier, R. Daudel, and I. G. Csizmadia. Uniform quality Gaussian basis sets for molecular calculations. III. Charge optimized basis sets. International Journal of Quantum Chemistry, 19(4):693–710, April 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Poirier:1981:UQGb

[PDC81b] Raymond A. Poirier, Raymond Daudel, and Imre G. Csizmadia. Uniform quality Gaussian basis sets for molecular calculations. IV. Gradient and charged optimized basis sets for CH_{4}. International Journal of Quantum Chemistry, 19(4):711–718, April 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Poirier:1984:ICS

[PDKC84] Raymond A. Poirier, Raymond Daudel, Roy E. Kari, and Imre G. Csizmadia. Ab initio calculations on sulfur-containing compounds. II. One-electron properties of H_{2}S. International Journal of Quantum Chemistry, 25(2):411–421, February 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Poirier:1980:UQGa

[PDMC80] R. A. Poirier, R. Daudel, P. G. Mezey, and I. G. Csizmadia. Uniform quality Gaussian basis sets for molecular calculations, I. C_{1} hydrocarbons. International Journal of Quantum Chemistry, 18(3):715–725, September 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Poirier:1982:ICS

[PDMC82] R. A. Poirier, R. Daudel, P. G. Mezey, and I. G. Csizmadia. Ab initio calculations on sulfur-containing compounds. I. Uniform quality basis sets for sulfur: Total energies and geometries of H_{2}S. International Journal of Quantum Chemistry, 21(5):799– 811, May 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Peel:1983:ERE

[Pee83] J. Barrie Peel. Electrostatic and relaxation effects in ionization of molecular dimers and intermolecular complexes. International Journal of Quantum Chemistry, 23(2):653–662, February 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Peek:1989:MEM

[Pee89] James M. Peek. Molecular effects on mean excitation energies. International Journal of Quantum Chemistry, 36(S23): 545–552, April 1–8, 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Percus:1982:NFG

[Per82] J. K. Percus. Nonuniform fluids in the grand canonical ensemble. International Journal of Quantum Chemistry, 22(S16): 33–48, March 1–13, 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Chemistry, Theory of Condensed Matter, and Propagator Methods in the Quantum Theory of Matter.

Perdew:1985:DFT

[Per85] John P. Perdew. Density functional theory and the band gap problem. International Journal of Quantum Chemistry, 28 (S19):497–523, March 18–23, 1985. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Perdew:1986:DFT

[Per86] John P. Perdew. Density functional theory and the band gap problem. International Journal of Quantum Chemistry, 30 (3):451, September 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Peshev:1983:PFD

[Pes83] O. Peshev. Photoreactions on finely divided semiconductors. International Journal of Quantum Chemistry, 23(4): 1529–1533, April 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Pethig:1979:DEP

[Pet79] Ronald Pethig. Dielectric and electronic properties of biological materials. John Wiley, New York, NY, USA, 1979. ISBN 0-471-99728-5. xiv + 376 pp. LCCN QH517 .P47.

Pattnaik:1981:UAE

[PFBF81] P. C. Pattnaik, J. L. Fry, N. E. Brener, and G. Fletcher. Use of ATM in evaluating spin-density. International Journal of Quantum Chemistry, 20(S15):499–505, March 8–14, 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory, Collision Phenomena, and Computational Quantum Chemistry.

Paul:1988:TIA

[PFC88] R. Paul, O. G. Fritz, and R. Chatterjee. Theoretical interpretation of the anomalous behavior of some macromolecules. International Journal of Quantum Chemistry, 34(S15):191– 199, March 12–19, 1988. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Syposium on Quantum Biology and Quantum Pharmacology.

Pople:1983:MPS

[PFLB83] John A. Pople, Michael J. Frisch, Brian T. Luke, and J. Stephen Binkley. A Møller–Plesset study of the energies of AH_{n} molecules (A = Li to F). International Journal of Quantum Chemistry, 24(S17):307–320, March 3–12, 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-state Theory, Collision Phenomena, and Computational Quantum Chemistry.

Pack:1980:MGO

[PG80] George R. Pack and Dudley V. Goetschel. Molecular geometry optimization by semiempirical methods using an optimized minimization of electronic and nuclear energies. International Journal of Quantum Chemistry, 18(S7):207–216, March 5–8, 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Plant:1989:HBV

[PG89] J. S. Plant and H. H. Greenwood. Hydrogen bonding in vitamin B_{6}. International Journal of Quantum Chemistry, 35 (3):385–393, March 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Palma:1980:EPI

[PGG80] A. Palma, E. González, and O. Goscinski. ESCA peak intensities for small molecules in the sudden approximation. International Journal of Quantum Chemistry, 18(1):237–241, July 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Pan:1986:SNA

[PGL^{+}86] Dao-Kai Pan, Jian-Nan Gao, Hai-Lun Lin, Ming-Bao Huang, and W. H. Eugen Schwarz. Some new aspects of bonding in cyclopropane. International Journal of Quantum Chemistry, 29(5):1147–1154, May 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Pettitt:1980:TCP

[PGM80] B. Montgomery Pettitt, Shridhar R. Gadre, and Robert L. Matcha. Theoretical Compton profile anisotropics in molecules and solids. VIII. Vibrational, rotational, and temperature-dependent diatomic alkali halide anisotropics. International Journal of Quantum Chemistry, 18(S14):697– 706, March 10–20, 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory, Collision Phenomena, Quantum Statistics, and Computational Methods.

Poulain:1986:STS

[PGPRN86] E. Poulain, J. Garcia-Prieto, M. E. Ruiz, and O. Novaro. Some theoretical studies on hydrogen molecule capture by platinum atoms. International Journal of Quantum Chemistry, 29(5): 1181–1190, May 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Politzer:1982:SRP

[PH82] Peter Politzer and Winston L. Hedges. A study of the reactive properties of the chlorinated ethylenes. International Journal of Quantum Chemistry, 22(S9):307–319, March 4–6, 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Porath:1987:TAN

[PH87] Jerker Porath and T. William Hutchens. Thiophilic adsorption: A new kind of molecular interaction revealed by chromatography. International Journal of Quantum Chemistry, 32(S14):297–315, March 14, 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Pople:1988:CCE

[PHGR88] J. A. Pople, M. Head-Gordon, and Krishnan Raghavachari. Corrections to correlations energies beyond fourth order Möller–Plesset (MP4) perturbation theory. Contributions of single, double, and triple substitutions. International Journal of Quantum Chemistry, 34(S22):377–382, March 12–19, 1988. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Syposium on Quantum Chemistry, Solid-State Theory, and Computational Methods.

Piacentini:1983:TFA

[Pia83] R. D. Piacentini. Translation factors in atomic collisions. International Journal of Quantum Chemistry, 23(2):567–575, February 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Pickup:1983:BRBb

[Pic83a] Barry T. Pickup. Book review: Electrons and valence: Development of the theory, 1900–25. By Anthony N. Stranges, Texas A&M University Press, College Station, 1982. Price: $28.50. International Journal of Quantum Chemistry, 24(1): 136, July 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). See [Str82].

Pickup:1983:BRBa

[Pic83b] Barry T. Pickup. Book review: Quantum chemistry literature data base bibliography of ab initio calculations for 1978–80.

By K. Ohno and K. Morokuma, Elsevier, New York, 1982. Vol. 12 in Physical Sciences Data. Price: $104.75. International Journal of Quantum Chemistry, 24(1):135, July 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). See [OM82a, OM82b, OM83b, OM83a, OM84, OM85, OM86, OMH89].

Piecuch:1982:IPM

[Pie82] Piotr Piecuch. Invariance properties of the multipole expansion. International Journal of Quantum Chemistry, 22(2):293– 298, August 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Piecuch:1984:STT

[Pie84] Piotr Piecuch. Spherical tensor theory of long-range interactions between two molecules. International Journal of Quantum Chemistry, 25(3):449–473, March 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Piecuch:1985:SST

[Pie85] Piotr Piecuch. Supplement to spherical tensor theory of longrange interactions between two molecules. International Journal of Quantum Chemistry, 28(3):375–386, September 1985. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Pincus:1988:CBP

[Pin88] Matthew R. Pincus. The chain build-up procedure in computing the structures of biologically active polypeptides and proteins. International Journal of Quantum Chemistry, 34 (S15):209–220, March 12–19, 1988. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Syposium on Quantum Biology and Quantum Pharmacology.

Pipek:1985:COL

[Pip85] J. Pipek. Controlled orthogonalization of localized orbitals. International Journal of Quantum Chemistry, 27(5):527–546, May 1985. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Pipek:1989:LMM

[Pip89] János Pipek. Localization measure and maximum delocalization in molecular systems. International Journal of Quantum Chemistry, 36(4):487–501, October 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Pitzer:1984:RCD

[Pit84] Kenneth S. Pitzer. Relativistic calculations of dissociation energies and related properties. International Journal of Quantum Chemistry, 25(1):131–148, January 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Purvis:1988:PTT

[PK88] George D. Purvis III and James W. King. Proton transfer and triol formation in the simulation of GABA solvation. International Journal of Quantum Chemistry, 34(S22):593–600, March 12–19, 1988. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Syposium on Quantum Chemistry, Solid-State Theory, and Computational Methods.

Poirier:1985:HGB

[PKC85] Raymond Poirier, Roy Kari, and I. G. Csizmadia. Handbook of Gaussian basis sets: a compendium for ab-initio molecular orbital calculations, volume 24 of Physical sciences data. Elsevier, Amsterdam, The Netherlands, 1985. ISBN 0-444-42493-8. xii + 674 pp. LCCN QD461 .P63 1985.

Panin:1986:SME

[PKG86] A. I. Panin, Yu. G. Khait, and N. N. Gorinchoy. Structure of MCSCF electronic energy domain. International Journal of Quantum Chemistry, 30(4):509–541, October 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Phadke:1981:SFP

[PKH^{+}81] Ratna S. Phadke, N. Vasanth Kumar, R. V. Hosur, Anil Saran, and Girjesh Govil. Structure and function of propranolol: a βadrenergic blocking drug. International Journal of Quantum Chemistry, 20(1):85–92, July 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Pavlovic:1987:VCD

[PKH87] Milica Pavlović, Jasna Kučar, and Josip Hendeković. Variational construction of diabatic states. International Journal of Quantum Chemistry, 32(6):705–710, December 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Psoda:1981:IMB

[PKP^{+}81] Anna Psoda, B. Kierdaszuk, A. Pohorille, M. Geller, J. T. Kusmierek, and D. Shugar. Interaction of the mutagenic base analogs O^{6}-methylguanine and N^{4}-hydroxycytosine with potentially complementary bases. International Journal of Quantum Chemistry, 20(2):543–554, August 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Pauncz:1987:SPO

[PKP87] R. Pauncz, B. Kirtman, and W. E. Palke. Studies in the paired orbital method. I. International Journal of Quantum Chemistry, 32(S21):533–538, March 12, 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Chemistry, Solid-State Theory, and Computational Methods.

Pakiari:1980:SDO

[PL80] A. H. Pakiari and J. W. Linnett. Single-determinantal open- and closed-shell FSGO calculations for some diatomic molecules. International Journal of Quantum Chemistry, 18 (3):661–673, September 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Pudzianowski:1983:MSO

[PL83] Andrew T. Pudzianowski and Gilda H. Loew. Mechanistic studies of oxene reactions with organic substrates: Reaction paths on MNDO enthalpy surfaces—models for cytochrome P450 oxidations. International Journal of Quantum Chemistry, 23(4):1257–1268, April 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Politzer:1984:HHE

[PL84a] Peter Politzer and Patricia R. Laurence. Halogenated hydrocarbon epoxides: Factors underlying biological activity. International Journal of Quantum Chemistry, 26(S11):155–166, March 12–15, 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Pudzianowski:1984:MSO

[PL84b] Andrew T. Pudzianowski and Gilda H. Loew. Mechanistic studies of oxene reactions with organic substrates: Reaction paths on MNDO enthalpy surfaces—models for cytochrome P450 oxidations. International Journal of Quantum Chemistry, 25(2):445, February 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Padkjer:1989:EHC

[PL89] Søren Berg Padkjér and Jan Linderberg. Electronic Hamiltonians for collision theory. International Journal of Quantum Chemistry, 36(S23):1–15, April 1–8, 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Pakkanen:1986:MVC

[PLL86] Tapani A. Pakkanen, Marina Lindblad, and Risto S. Laitinen. Molecular valence calculations based on stepwise approximations for orthogonality and core-valence interactions. International Journal of Quantum Chemistry, 29(6):1789–1797, June 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

PrasadBhattacharyya:1981:HTO

[PM81] Sankar PrasadBhattacharyya and Debashis Mukherjee. A hybrid technique of orthonormality constrained orbital optimization in SCF calculations. International Journal of Quantum Chemistry, 20(6):1165–1177, December 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Palma:1983:FCF

[PM83] A. Palma and J. Morales. Franck–Condon factors and ladder operators. I. harmonic oscillator. International Journal of Quantum Chemistry, 24(S17):393–400, March 3–12, 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-state Theory, Collision Phenomena, and Computational Quantum Chemistry.

Pickup:1984:SSP

[PM84a] B. T. Pickup and A. Mukhopadhyay. Spin-shift propagators. International Journal of Quantum Chemistry, 26(S18):309– 319, March 1–15, 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Atomic, Molecular and Solid-State Theory, and Computational Quantum Chemistry.

Pickup:1984:UGF

[PM84b] B. T. Pickup and A. Mukhopadhyay. A unitary group formulation of many-body theory: The spin-shift formalism. International Journal of Quantum Chemistry, 26(1):101–123, July 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Pucci:1986:GZE

[PM86] R. Pucci and N. H. March. Generalized 1/Z expansion for heteronuclear molecules. International Journal of Quantum Chemistry, 29(4):949–958, April 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Pipek:1988:DMS

[PM88] János Pipek and Paul G. Mezey. Dependence of MO shapes on a continuous measure of delocalization. International Journal of Quantum Chemistry, 34(S22):1–13, March 12–19, 1988. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Syposium on Quantum Chemistry, Solid-State Theory, and Computational Methods.

Parra-Mouchet:1988:SCR

[PMCA88] Julia Parra-Mouchet, Renato R. Contreras, and Arie Aizman. Self-consistent reaction field calculations on the proton transfer in ammonia-formic acid systems as a model for hydrogen bonding in amino acids in solution. International Journal of Quantum Chemistry, 33(1):41–52, January 1988. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Pardo:1989:TSA

[PMOW89] Leonardo Pardo, Aleksander P. Mazurek, Roman Osman, and Harel Weinstein. Theoretical studies of the activation mechanism of histamine H_{2}-receptors: Dimaprit and the receptor model. International Journal of Quantum Chemistry, 36(S16): 281–290, April 1–8, 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Paniagua:1984:LMO

[PMT84] Juan Carlos Paniagua, Albert Moyano, and Luis María Tel. Localized molecular orbitals for multiple bonds in the ZDO approximation: The σ-π vs. banana bonds dilemma. International Journal of Quantum Chemistry, 26(3):383–403, September 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Proinov:1986:CTA

[PNA86] E. Proinov, N. Neshev, and A. Andreev. The concept of the topological atom within the MO-LCAO approach. I. On the effective charge of an atom-in-molecule. International Journal of Quantum Chemistry, 30(1):119–126, July 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Proinov:1988:CTA

[PNA88] E. Proinov, N. Neshey, and A. Andreev. The concept of the topological atom within the MO–LCAO approach. II. On the nature of the nephelauxetic effect. International Journal of Quantum Chemistry, 34(1):1–8, July 1988. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Pan:1983:SNC

[PNGL83] Dao-Kai Pan, Ze-Fu Niu, Jian-Nan Gao, and San-Lin Liu. Some new concepts on the nature of the covalent bond based on ab initio quantum-chemical calculations. International Journal of Quantum Chemistry, 23(6):1945–1958, June 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Pilarski:1985:RBEb

[PO85] Bogusław Pilarski and Krzysztof OśMiałowski. The relationship between electron densities and the p K_{a} values in a series of methylpyrazines. International Journal of Quantum Chemistry, 28(2):239–244, August 1985. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Pohl:1980:OFA

[Poh80] Herbert A. Pohl. Oscillating fields about growing cells. International Journal of Quantum Chemistry, 18(S7):411–431, March 5–8, 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Pohl:1982:NCE

[Poh82] Herbert A. Pohl. Natural cellular electrical resonances. International Journal of Quantum Chemistry, 22(S9):399–409, March 4–6, 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Pohl:1983:CSR

[Poh83] Herbert A. Pohl. Cellular spin resonance: A new method for determining the dielectric properties of living cells. International Journal of Quantum Chemistry, 24(S10):161–174, March 14–16, 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Pohl:1984:NAF

[Poh84] Herbert A. Pohl. Natural alternating fields associated with living cells. International Journal of Quantum Chemistry, 26 (S11):367–368, March 12–15, 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Pilarski:1985:RBEa

[POK85] Bogusław Pilarski, Krzysztof Ośmiałowski, and Roman Kaliszan. The relationship between electron densities and the electronic substituent constants for substituted pyridines. International Journal of Quantum Chemistry, 28(2):233–237, August 1985. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Pedersen:1981:SST

[POS81] C. Th. Pedersen, J. Oddershede, and J. R. Sabin. Semiempirical studies on 1,2- and 1,3-tetrathiofulvalenes. International Journal of Quantum Chemistry, 20(2):369–371, August 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Poshusta:1983:NSP

[Pos83] R. D. Poshusta. Nonadiabatic Singer polymal wave functions for three-particle systems. International Journal of Quantum Chemistry, 24(1):65–77, July 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Polivtsev:1980:ITS

[POT80] O. F. Polivtsev, V. D. Osika, and D. Ya. Tsvankin. Investigation of tertiary structure of a series of DNA compact forms by the small-angle X-ray scattering method. International Journal of Quantum Chemistry, 17(4):813–817, April 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Pullman:1980:EEM

[PP80] Alberte Pullman and Bernard Pullman. Electrostatic effect of the macromolecular structure on the biochemical reactivity of the nucleic acids. Significance for chemical carcinogenesis. International Journal of Quantum Chemistry, 18(S7):245–259, March 5–8, 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Peters:1981:SRD

[PP81] David Peters and Jane Peters. Some recent developments in the quantum theory of hydrogen bonding and other effects in proteins. International Journal of Quantum Chemistry, 19 (6):1121–1132, June 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Politzer:1982:CPS

[PP82] Peter Politzer and Timothy R. Proctor. Calculated properties of some possible vinyl chloride metabolites. International Journal of Quantum Chemistry, 22(6):1271–1279, December 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Pauncz:1983:BLA

[PP83] R. Pauncz and J. Paldus. Bond length alternation in cyclic polyenes. III. Alternant molecular orbital method. International Journal of Quantum Chemistry, 24(4):411–423, October 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Pack:1984:NDC

[PP84] George R. Pack and C. V. Prasad. New developments in computational polyelectrolyte theory. International Journal of Quantum Chemistry, 26(S11):287–299, March 12–15, 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Pethig:1986:IEI

[PP86] Ronald Pethig and Jonathan A. R. Price. The influence of electrostatic interactions on the activity of lysozyme. International Journal of Quantum Chemistry, 30(S13):39–51, March 6–8, 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Piecuch:1989:OSA

[PP89] Piotr Piecuch and Josef Paldus. Orthogonally spin-adapted coupled-cluster equations involving singly and doubly excited clusters. Comparison of different procedures for spinadaptation. International Journal of Quantum Chemistry, 36 (4):429–453, October 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Padjen:1987:QCD

[PPB87] R. Padjen, D. Paquet, and F. Bonnouvrier. A quasiparticle calculation of the dispersion curves of gallium arsenide based on a full Hartree–Fock approximation. International Journal of Quantum Chemistry, 32(S21):45–53, March 12, 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Chemistry, Solid-State Theory, and Computational Methods.

Piccitto:1989:EEI

[PPMG89] G. Piccitto, R. Pucci, N. H. March, and A. Grassi. Electron–electron interaction energy in homonuclear diatomic molecules. International Journal of Quantum Chemistry, 36 (4):525–531, October 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Pollock:1987:CSR

[PPP87] J. Kent Pollock, Douglas G. Pohl, and Herbert A. Pohl. Cellular spin resonance inversion by ion-induced dipoles. International Journal of Quantum Chemistry, 32(1):19–29, July 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Pohl:1984:EC

[PPR84] Herbert A. Pohl, K. Pollock, and H. Rivera. The electrofusion of cells. International Journal of Quantum Chemistry, 26 (S11):327–345, March 12–15, 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Paul:1983:CCH

[PR83] J. Paul and A. Rosén. Cluster calculations of the H_{2}O/Pt(111) system. International Journal of Quantum Chemistry, 23(4): 1231–1238, April 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Prat:1983:PIA

[Pra83] R. F. Prat. A possible interpretation of atomic valency as a spontaneously broken symmetry response. International Journal of Quantum Chemistry, 24(6):697–706, December 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Petelenz:1980:BEI

[PS80a] Piotr Petelenz and Vedene H. Smith, Jr. Binding energies of ionized donor–exciton complexes in polar crystals. International Journal of Quantum Chemistry, 18(2):583–586, August 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Petelenz:1980:ITM

[PS80b] Piotr Petelenz and Vedene H. Smith, Jr. An integral transform method for solving a quantum mechanical three-body problem with exponential-type potentials. International Journal of Quantum Chemistry, 18(S14):83–92, March 10–20, 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory, Collision Phenomena, Quantum Statistics, and Computational Methods.

Prasad:1981:ILP

[PS81] C. V. Prasad and K. Sundaram. Impact of lone pairs on the conformation and configuration statistics of poly(lactic acid) chains. International Journal of Quantum Chemistry, 20(3): 613–621, September 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Pokleba:1984:PTC

[PS84] A. K. Pokleba and U. I. Safronova. 1/Z perturbation theory for calculation of line strengths in the neon isoelectronic sequence. International Journal of Quantum Chemistry, 25 (1):69–75, January 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Pavlides:1987:SPE

[PS87] P. Pavlides and K. D. Sales. Spin-projected EHF equations. International Journal of Quantum Chemistry, 32(6):711–722, December 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Palma:1988:NTD

[PS88] A. Palma and L. Sandoval. The nonabelian two-dimensional algebra and the Franck–Condon integral. International Journal of Quantum Chemistry, 34(S22):503–505, March 12–19, 1988. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Syposium on Quantum Chemistry, Solid-State Theory, and Computational Methods.

Purvis:1983:CIP

[PSBB83] George D. Purvis III, Ron Shepard, Franklin B. Brown, and Rodney J. Bartlett. C_{2V} insertion pathway for BeH_{2}: a test problem for the coupled-cluster single and double excitation model. International Journal of Quantum Chemistry, 23(3): 835–845, March 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Phadke:1983:EDD

[PSG83] Ratna S. Phadke, H. M. Sonawat, and Girjesh Govil. Electron delocalization during the oxidation-reduction cycle of FAD and NAD: A quantum chemical approach to the design of coenzyme-immobilized bioanode for biochemical fuel cells. International Journal of Quantum Chemistry, 24(S10):251–262, March 14–16, 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Polansky:1987:SPN

[PSIT87] O. E. Polansky, P. Schuster, C. I. Ivanov, and N. Tyutyulkov. Structure and properties of nonclassical polymers. V. On a class of nonalternant polymers with localized nonbonding bands. International Journal of Quantum Chemistry, 32(4): 491–499, October 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Pople:1981:MOS

[PSK^{+}81] J. A. Pople, H. B. Schlegel, R. Krishnan, D. J. Defrees, J. S. Binkley, M. J. Frisch, R. A. Whiteside, R. F. Hout, and W. J. Hehre. Molecular orbital studies of vibrational frequencies. International Journal of Quantum Chemistry, 20(S15):269–278, March 8–14, 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory, Collision Phenomena, and Computational Quantum Chemistry.

Petkov:1986:MLS

[PSK86] Ivan Zh. Petkov, Mario V. Stoitsov, and Eugene S. Kryachko. Method of local-scaling transformations and densityfunctional theory in quantum chemistry. International Journal of Quantum Chemistry, 29(1):149–161, January 1986. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Palma:1987:CFO

[PSM87] A. Palma, L. Sandoval, and J. Morales. Closed formulas for one- and two-center harmonic oscillator integrals. International Journal of Quantum Chemistry, 32(S21):729–735, March 12, 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Chemistry, Solid-State Theory, and Computational Methods.

Person:1986:KET

[PSS86] Willis B. Person, Krystyna Szczepaniak, and Marian Szczesniak. Keto-enol tautomerism of guanine and 9-methylguanine monomers. International Journal of Quantum Chemistry, 30 (S13):287–289, March 6–8, 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Person:1987:ETS

[PSS^{+}87] Willis B. Person, Krystyna Szczepaniak, Marian Szczesniak, Jozef S. Kwiatkowski, and Luis Hernandez. Experimental and theoretical studies of tautomers of nucleic acid bases. International Journal of Quantum Chemistry, 32(S21):765–766, March 12, 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Chemistry, Solid-State Theory, and Computational Methods.

Paldus:1984:BLA

[PT84] J. Paldus and M. Takahashi. Bond length alternation in cyclic polyenes. IV. Finite-order perturbation theory approach. International Journal of Quantum Chemistry, 25(2):423–443, February 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Pick:1989:SST

[PT89a] Štěpán Pick and Mojmír Tomášek. On surface states of transition metal carbides, nitrides, and oxides: (001), (011), and (111) surfaces. International Journal of Quantum Chemistry, 36(2):157–167, August 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Poltev:1989:CIS

[PT89b] V. I. Poltev and A. V. Teplukhin. Conformational implications of some nucleotide sequences. International Journal of Quantum Chemistry, 35(1):91–102, January 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Paldus:1984:DCC

[PTC84] J. Paldus, M. Takahashi, and B. W. H. Cho. Degeneracy and coupled-cluster approaches. International Journal of Quantum Chemistry, 26(S18):237–244, March 1–15, 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Atomic, Molecular and Solid-State Theory, and Computational Quantum Chemistry.

Pullman:1980:CFM

[PTG80] Bernard Pullman, Paul O. P. (Paul On Pong) Ts’o, and Harry Victor Gelboin, editors. Carcinogenesis, fundamental mechanisms and environmental effects: proceedings of the thirteenth Jerusalem Symposium on Quantum Chemistry and Biochemistry, held in Jerusalem, Israel, April 28–May 2, 1980, The Jerusalem symposia on quantum chemistry and biochemistry; v. 13. D. Reidel, Dordrecht, Boston, Lancaster, Tokyo, 1980. ISBN 90-277-1171-2. LCCN RC268.5 .J47 1980.

Pullman:1979:CCB

[Pul79] Bernard Pullman, editor. Catalysis in chemistry and biochemistry: theory and experiment: proceedings of the Twelfth Jerusalem Symposium on Quantum Chemistry and Biochemistry held in Jerusalem, Israel, April 2–4, 1979, volume 12 of The Jerusalem symposia on quantum chemistry and biochemistry. D. Reidel, Dordrecht, Boston, Lancaster, Tokyo, 1979. ISBN 90-277-1039-2. LCCN QD505 .J47 1979.

Pulay:1983:VFG

[Pul83] Peter Pulay. Variational formulation and gradient evaluation for coupled electron pair approximations: A model study. International Journal of Quantum Chemistry, 24(S17):257–263, March 3–12, 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-state Theory, Collision Phenomena, and Computational Quantum Chemistry.

Pullman:1986:FCQ

[Pul86] Bernard Pullman. A few considerations on quinones as antitumor agents. International Journal of Quantum Chemistry, 30(S13):95–105, March 6–8, 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Purdela:1988:RPT

[Pur88] D. Purdela. A regular periodic table of the elements and its quantum mechanical requirement. International Journal of Quantum Chemistry, 34(2):107–119, August 1988. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Palma-Vittorelli:1989:MIA

[PV89] M. B. Palma-Vittorelli. Molecular interactions in aqueous solutions: The role of random fluctuations of external parameters and of thermodynamic instabilities. International Journal of Quantum Chemistry, 35(1):113–124, January 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Pickett:1986:LDT

[PW86] W. E. Pickett and C. S. Wang. A local density theory for dynamical correlation corrections to single particle excitations in semiconductors. International Journal of Quantum Chemistry, 30(S20):299–311, March 10–15, 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Syposium on Atomic, Molecular, and Solid-State Theory, Scattering Problems, Many Body Phenomena, and Computational Quantum Chemistry.

Pack:1989:MSD

[PWL89] George R. Pack, Linda Wong, and Gene Lamm. Model systems for DNA and its environment: Suitability and accuracy in theoretical calculations. International Journal of Quantum Chemistry, 36(S16):1–15, April 1–8, 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Pan:1986:SPS

[PXLC86] Y. K. Pan, M. Xin, D. Li, and Lue-Yung Chow Chiu. Study of photoelectron spectra of some XPY_{3} molecules (X = O, S and Y = Cl, Br) by the SCF-X_{α}–SW method. International Journal of Quantum Chemistry, 29(3):393–397, March 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Pyykko:1984:I

[Pyy84] Pekka Pyykkö. Introduction. International Journal of Quantum Chemistry, 25(1):1–2, January 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Pyykko:1986:RTA

[Pyy86] Pekka Pyykkö. Relativistic theory of atoms and molecules: a bibliography, 1916–1985, volume 41 of Lecture notes in chemistry. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1986. ISBN 0-387-17167-3 (paperback). ix + 389 pp. LCCN Z5524.A6 P99 1986.

Pyykko:1993:RTA

[Pyy93] Pekka Pyykkö. Relativistic theory of atoms and molecules II: a bibliography, 1986–1992, volume 60 of Lecture notes in chemistry. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1993. ISBN 3-540-57219-8 (Berlin), 0-387-57219-8 (New York). viii + 479 pp. LCCN Z5524.A6 P99 1993.

Pyykko:2000:RTA

[Pyy00] Pekka Pyykkö. Relativistic theory of atoms and molecules III: a bibliography, 1993–1999, volume 76 of Lecture notes in chemistry, 0342-4901. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2000. ISBN 3-540-41398-7. x + 354 pp. LCCN Z5524.A6 P99 2000.

Pan:1984:TCV

[PZ84] Dai-Kai Pan and Bang-Rong Zhang. Theoretical calculation of vertical ionization potentials of B_{2}H_{6} by means of ΔE_{SCF} procedures. International Journal of Quantum Chemistry, 25 (3):535–542, March 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Preiskorn:1985:MEM

[PŻ85] Aleksandra Preiskorn and Bronislaw Żurawski. Manyelectron, multicenter integrals in superposition of correlated configurations method. I. one- and two-electron integrals. International Journal of Quantum Chemistry, 27(6):641–651, June 1985. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Prakash:1986:TRF

[PZPL86] A. S. Prakash, I. S. Zegar, H. L. Price, and P. R. Lebreton. Time-resolved fluorescence studies of the effects of metabolism on the in vitro reversible binding and accumulation of benzo[a]pyrene in DNA. International Journal of Quantum Chemistry, 30(S13):237–250, March 6–8, 1986. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Prakash:1983:SAI

[PZSL83] A. S. Prakash, I. S. Zegar, M. Shahbaz, and P. R. Lebreton. Stereoelectronic aspects of the intercalative binding of benz[a]anthracene metabolite models to DNA. International Journal of Quantum Chemistry, 24(S10):349–356, March 14– 16, 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Qianer:1983:SDO

[Qia83] Zhang Qianer. Symmetry determined orbital and group overlap. International Journal of Quantum Chemistry, 23(4): 1479–1492, April 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Qianer:1986:DCS

[QX86] Zhang Qianer and Li Xiangzhu. Double coset for symmetry orbitals. International Journal of Quantum Chemistry, 29(3): 293–303, March 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Qianer:1989:ITM

[QX89] Zhang Qianer and Li Xiangzhu. Improved treatment for matrix elements of spin-dependent operators. International Journal of Quantum Chemistry, 36(3):201–211, September 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Radzio-Andzelm:1981:III

[RA81] Elżbieta Radzio-Andzelm. Investigation of intermolecular interactions based on the atomic statistical model. International Journal of Quantum Chemistry, 20(3):601–612, September 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Rabinowitz:1984:MSE

[Rab84] James R. Rabinowitz. A model for some of the effects of an externally applied electric field on charged membrane constituents. International Journal of Quantum Chemistry, 26 (S11):249–256, March 12–15, 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Rico:1986:TAD

[RACPL86] J. Férnandez Rico, J. R. Alvarez-Collado, M. Paniagua, and R. López. Transferability of atomic descriptions in molecules. International Journal of Quantum Chemistry, 30(5):671–684, November 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Randic:1980:REV

[Ran80a] Milan Randić. Resonance energy of very large benzenoid hydrocarbons. International Journal of Quantum Chemistry, 17 (3):549–586, March 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Randic:1980:SPG

[Ran80b] Milan Randić. Symmetry properties of graphs of interest in chemistry. III. Homotetrahedryl rearrangement. International Journal of Quantum Chemistry, 18(S14):557–577, March 10– 20, 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory, Collision Phenomena, Quantum Statistics, and Computational Methods.

Randica:1980:GEC

[Ran80c] Milan Randića. Graphical enumeration of conformations of chains. International Journal of Quantum Chemistry, 18(S7): 187–197, March 5–8, 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Ransil:1980:HGG

[Ran80d] Bernard J. Ransil. How good is good agreement: Evaluating the reliability of quantum-mechanically calculated observables. International Journal of Quantum Chemistry, 17(3): 467–478, March 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Randic:1981:GTC

[Ran81] M. Randic. Graph theoretical characterization of the dendritic fields. International Journal of Quantum Chemistry, 20(S8): 463–479, March 5–7, 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Randic:1983:SSR

[Ran83] Milan Randić. Survey of structural regularities in molecular properties. I. Carbon-13 chemical shifts in alkanes. International Journal of Quantum Chemistry, 23(5):1707–1722, May 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Randic:1984:NAS

[Ran84] Milan Randić. Nonempirical approach to structure–activity studies. International Journal of Quantum Chemistry, 26 (S11):137–153, March 12–15, 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Randic:1988:CTD

[Ran88] Milan Randić. On characterization of three-dimensional structures. International Journal of Quantum Chemistry, 34(S15): 201–208, March 12–19, 1988. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Syposium on Quantum Biology and Quantum Pharmacology.

Rashin:1988:CEH

[Ras88] Alexander A. Rashin. Continuum electrostatics and hydration phenomena. International Journal of Quantum Chemistry, 34 (S15):103–118, March 12–19, 1988. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Syposium on Quantum Biology and Quantum Pharmacology.

Ratner:1987:BRB

[Rat87] M. A. Ratner. Book review: Chemical applications of ultrafast spectroscopy. By Graham R. Fleming, Oxford, New York, 1986. International Journal of Quantum Chemistry, 31(6): 989, June 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). See [Fle86].

Ratner:1988:BRB

[Rat88] Mark Ratner. Book review: Solid state chemistry—techniques. By A. K. Cheetham and P. Day, Editors, Clarendon, Oxford, 1987. International Journal of Quantum Chemistry, 34(1):87, July 1988. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). See [CD92].

Raveche:1982:DJM

[Rav82] Harold J. Raveché. Dedication to Joseph E. Mayer. International Journal of Quantum Chemistry, 22(S16):2–13, March 1–13, 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Chemistry, Theory of Condensed Matter, and Propagator Methods in the Quantum Theory of Matter.

Ramani:1981:IMO

[RB81] R. Ramani and Russell J. Boyd. Ab-initio molecular orbital study of the cis/trans conformations of the peptide bond. International Journal of Quantum Chemistry, 20(S8):117–127, March 5–7, 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Ramiarez:1986:SGG

[RB86] Rafael Ramíarez and Michael C. Böhm. Simple geometric generation of special points in brillouin-zone integrations. Twodimensional Bravais lattices. International Journal of Quantum Chemistry, 30(3):391–411, September 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Ramirez:1988:COA

[RB88a] Rafael Ramírez and Michael C. Böhm. A Crystal Orbital approach for two- and three-dimensional solids on the basis of CNDO /INDO Hamiltonians. Basis equations. International Journal of Quantum Chemistry, 34(1):47–71, July 1988. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Ramirez:1988:TDB

[RB88b] Rafael Ramírez and Michael C. Böhm. The two-dimensional band structure of (polyphthalocyaninato)Ni(II). International Journal of Quantum Chemistry, 34(1):73–84, July 1988. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Ramirez:1988:USR

[RB88c] Rafael Ramirez and Michael C. Böhm. The use of symmetry in reciprocal space integrations. Asymmetric units and weighting factors for numerical integration procedures in any crystal symmetry. International Journal of Quantum Chemistry, 34 (6):571–594, December 1988. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Rao:1988:CBR

[RB88d] Gita Subba Rao and J. S. Bajaj. Conformational basis of the receptor-binding potency of normal and mutant insulin molecules with relevance to the pathophysiology of noninsulin dependent diabetes mellitus (NIDDM). International Journal of Quantum Chemistry, 34(S15):95–101, March 12–19, 1988. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Syposium on Quantum Biology and Quantum Pharmacology.

Recamier:1989:VET

[RB89] José Récamier and Manuel Berrondo. VT energy transfer in molecular collisions: An algebraic approach. International Journal of Quantum Chemistry, 36(S23):27–29, April 1–8, 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Reynolds:1986:QCQ

[RBH^{+}86] P. J. Reynolds, R. N. Barnett, B. L. Hammond, R. M. Grimes, and W. A. Lester, Jr. Quantum chemistry by quantum Monte Carlo: Beyond ground-state energy calculations. International Journal of Quantum Chemistry, 29(4):589–596, April 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Randic:1985:FCP

[RBK85] M. Randić, B. Baker, and A. F. Kleiner. Factoring the characteristic polynomial. International Journal of Quantum Chemistry, 28(S19):107–127, March 18–23, 1985. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Reynolds:1984:QMC

[RBL84] P. J. Reynolds, R. N. Barnett, and W. A. Lester, Jr. Quantum Monte Carlo study of the classical barrier height for the H + H_{2} exchange reaction: Restricted versus unrestricted trial functions. International Journal of Quantum Chemistry, 26(S18):709–717, March 1–15, 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Atomic, Molecular and Solid-State Theory, and Computational Quantum Chemistry.

Rodwell:1980:ISS

[RBR80] William R. Rodwell, Willem J. Bouma, and Leo Radom. Ab initio studies of 1,3-sigmatropic rearrangements: Effect of basis set and electron correlation. International Journal of Quantum Chemistry, 18(1):107–116, July 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Ray:1984:PDP

[RBSR84] Naba K. Ray, Giorgio Bolis, Masayuki Shibata, and Robert Rein. Potential derived point charge model study of electrostatic interaction energies in some complexes of water with uracil, thymine, and cytosine. International Journal of Quantum Chemistry, 26(S11):257–266, March 12–15, 1984. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Romano:1980:MCS

[RC80a] Silvano Romano and Enrico Clementi. Monte Carlo simulation of water solvent with biomolecules: Serine and the corresponding zwitterion. International Journal of Quantum Chemistry, 17(5):1007–1029, May 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Roszak:1980:FLM

[RC80b] Szczepan Roszak and Henryk Chojnacki. Formulation of the LCAS MS SCF method within the Gaussian basis set. International Journal of Quantum Chemistry, 18(1):65–71, July 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Rentzepis:1986:ACR

[RC86] Peter M. Rentzepis and Christos Capellos, editors. Advances in chemical reaction dynamics, volume 184 of NATO ASI series. Series C, Mathematical and physical sciences. D. Reidel, Dordrecht, Boston, Lancaster, Tokyo, 1986. ISBN 90-277-2312-5. LCCN QD502 .N365 1985.

Randic:1984:SPC

[RD84] Milan Randić and Michael I. Davis. Symmetry properties of chemical graphs. VI. Isomerizations of octahedral complexes. International Journal of Quantum Chemistry, 26(1):69–89, July 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Rawlings:1984:TIEb

[RDG84a] Diane C. Rawlings, Ernest R. Davidson, and Martin Gouterman. Theoretical investigations of the electronic states of porphyrins. II. Normal and hyper phosphorus porphyrins. International Journal of Quantum Chemistry, 26(2):251–274, August 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Rawlings:1984:TIEa

[RDG84b] Diane C. Rawlings, Ernest R. Davidson, and Martin Gouterman. Theoretical invstigations of the electronic states of porphyrins. I. Basis set development and predicted spectrum of pyrrole. International Journal of Quantum Chemistry, 26(2): 237–250, August 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

RuizDeAzua:1986:PPA

[RDG^{+}86] M. C. Ruiz De Azua, A. C. Diz, C. G. Giribet, R. H. Contreras, and I. D. Rae. A polarization propagator analysis of through-space spin-spin coupling constants: ^{19}F −^{19}F couplings. International Journal of Quantum Chemistry, 30(S20): 585–601, March 10–15, 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Syposium on Atomic, Molecular, and Solid-State Theory, Scattering Problems, Many Body Phenomena, and Computational Quantum Chemistry.

Roos:1984:TCM

[RDH84] Björn O. Roos, Yuri Yu. Dmitriev, and Matti Hotokka. Theory and computational methods for studies of nonlinear phenomena in laser spectroscopy. II. Calculations of steady-state wave functions. International Journal of Quantum Chemistry, 26(1):51–68, July 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Rittby:1983:CRT

[REB83a] Magnus Rittby, Nils Elander, and Erkki Brändas. Complex rotated Titchmarsh–Weyl theory: A review and some recent results. International Journal of Quantum Chemistry, 24(S17):

117–126, March 3–12, 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-state Theory, Collision Phenomena, and Computational Quantum Chemistry.

Rittby:1983:SVT

[REB83b] Magnus Rittby, Nils Elander, and Erkki Brändas. Scattering in view of the Titchmarsh–Weyl theory. International Journal of Quantum Chemistry, 23(3):865–874, March 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Recami:1980:ELN

[Rec80] E. Recami. Explaining the “large numbers” via a unified (classical) theory of strong and gravitational interactions. International Journal of Quantum Chemistry, 17(1):37–40, January 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Reid:1980:ZQO

[Rei80] Charles E. Reid. Zeroes of the quartic oscillator eigenfunctions. International Journal of Quantum Chemistry, 17(5): 1039–1041, May 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Reinhardt:1982:PSR

[Rei82] William P. Reinhardt. Padé summations for the real and imaginary parts of atomic Stark eigenvalues. International Journal of Quantum Chemistry, 21(1):133–146, January 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Rerat:1989:EFV

[Rér89] M. Rérat. Electric field variant ket for the calculation of dynamic polarizabilities. Application to H_{2}O and N_{2}. International Journal of Quantum Chemistry, 36(2):169–178, August 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Rettrup:1986:DES

[Ret86] Sten Rettrup. Direct evaluation of spin representation matrices and ordering of permutation-group elements. International Journal of Quantum Chemistry, 29(1):119–128, January 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Regier:1985:GVS

[RFST85] Philip E. Regier, Jacob Fisher, B. S. Sharma, and Ajit J. Thakkar. Gaussian vs. Slater representations of d orbitals: An information theoretic appraisal based on both position and momentum space properties. International Journal of Quantum Chemistry, 28(4):429–449, October 1985. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Rawlings:1985:TIE

[RGDF85] Diane C. Rawlings, Martin Gouterman, Ernest R. Davidson, and David Feller. Theoretical investigations of the electronic states of porphyrins. III. Low-lying electronic states of porphinatoiron(II). International Journal of Quantum Chemistry, 28 (6):773–796, December 1985. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Radom:1988:SMD

[RGW88] Leo Radom, Peter M. W. Gill, and Ming Wah Wong. Simple models for describing the fragmentation behavior of multiply charged cations. International Journal of Quantum Chemistry, 34(S22):567–573, March 12–19, 1988. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Syposium on Quantum Chemistry, Solid-State Theory, and Computational Methods.

Randic:1988:CAM

[RHST88] Milan Randić, Larhee L. Henderson, Roland Stout, and Nenad Trinajstić. Conjugation and aromaticity of macrocyclic systems. International Journal of Quantum Chemistry, 34(S22): 127–141, March 12–19, 1988. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Syposium on Quantum Chemistry, Solid-State Theory, and Computational Methods.

Rodriguez:1988:FNS

[RI88] Wilfredo Rodríguez and Yasuyuki Ishikawa. Fully numerical solutions of the Hartree–Fock equation in momentum space: A numerical study of the He atom and H ion. International Journal of Quantum Chemistry, 34(S22):445–456, March 12– 19, 1988. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Syposium on Quantum Chemistry, Solid-State Theory, and Computational Methods.

Richards:1988:EMR

[Ric88] Nigel G. J. Richards. Electrostatics and molecular recognition. International Journal of Quantum Chemistry, 34(S15): 85–93, March 12–19, 1988. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Syposium on Quantum Biology and Quantum Pharmacology.

Rips:1987:ETM

[RJ87] Ilya Rips and Joshua Jortner. Electron transfer in a microscopically inhomogeneous medium. International Journal of Quantum Chemistry, 32(S21):313–319, March 12, 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Chemistry, Solid-State Theory, and Computational Methods.

Rao:1988:ELA

[RJ88] B. K. Rao and P. Jena. Electron localization in anionic and cationic clusters. International Journal of Quantum Chemistry, 34(S22):287–296, March 12–19, 1988. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Syposium on Quantum Chemistry, Solid-State Theory, and Computational Methods.

Randic:1987:SAS

[RJBR^{+}87] Milan Randić, Borka Jerman-Blažič, Dennis H. Rouvray, Paul G. Seybold, and Stuart C. Grossman. The search for active substructures in structure-activity studies. International Journal of Quantum Chemistry, 32(S14):245–260, March 14, 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Romano:1983:MCS

[RJK83] S. Romano, B. Jönsson, and G. Karlström. Monte Carlo simulations of solid nitrogen in the isothermal–isobaric ensemble with an ab initio SCF–CI potential. International Journal of Quantum Chemistry, 23(3):991–998, March 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Randic:1982:SPC

[RK82] Milan Randić and Vladimir Katović. Symmetry properties of chemical graphs. IV. Rearrangement of tetragonal-pyramidal complexes. International Journal of Quantum Chemistry, 21 (3):647–663, March 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Roszak:1988:IMC

[RKHK88] Szczepan Roszak, Phillip B. Keegstra, P. C. Hariharan, and Joyce J. Kaufman. Ab-initio MRD-CI calculations for breaking a chemical bond in a molecule in a crystal or other solid environment I. H_{3}C − NO_{2} decomposition in nitromethane. International Journal of Quantum Chemistry, 34(S22):619– 653, March 12–19, 1988. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Syposium on Quantum Chemistry, Solid-State Theory, and Computational Methods.

Roszak:1989:IMC

[RKO^{+}89] Szczepan Roszak, Phillip B. Keegstra, Douglas W. O’neal, P. C. Hariharan, and Joyce J. Kaufman. Ab-initio MRD-CI calculations for breaking a chemical bond in a molecule in a crystal or other solid environment. II. H_{3}C—NO_{2} decomposition of nitromethane in a nitromethane crystal with voids. International Journal of Quantum Chemistry, 36(3):353–368, September 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Rubin:1980:PEP

[RKV^{+}80] A. B. Rubin, A. A. Kononenko, P. S. Venediktov, Galina P. Borisevitch, P. P. Knox, and E. P. Lukashev. Polarization effects in photosynthetic membranes. International Journal of Quantum Chemistry, 17(3):587–593, March 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Rabinowitz:1986:MET

[RL86] James R. Rabinowitz and Stephen B. Little. Multipole expansion techniques for the calculation and characterization of molecular electrostatic potentials. International Journal of Quantum Chemistry, 30(S13):9–18, March 6–8, 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Rabinowitz:1988:DMC

[RL88] James R. Rabinowitz and Stephen B. Little. Development of a model that contains both multipole moments and Gaussians for the calculation of molecular electrostatic potentials. International Journal of Quantum Chemistry, 34(S22):721–733, March 12–19, 1988. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Syposium on Quantum Chemistry, Solid-State Theory, and Computational Methods.

Ruette:1986:TSH

[RLH86] Fernando Ruette, Eduardo V. Ludeñta, and Antonio Hernández. A theoretical study of hydrogen surface coverage over Ni(100). International Journal of Quantum Chemistry, 29(5):1351–1364, May 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Rico:1986:APT

[RLPFA86] Jaime Fernáandez Rico, Rafael Lóapez, Miguel Paniagua, and Jose Ignacio Fernáandez-Alonso. Atomic partitioning of twocenter potentials for Slater basis. International Journal of Quantum Chemistry, 29(5):1155–1164, May 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Rico:1988:SES

[RLR88] J. Fernández Rico, R. Lóapez, and G. Ramíarez. Simplified expansion of Slater orbitals about displaced centers. International Journal of Quantum Chemistry, 34(2):121–131, August 1988. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Rubin:1987:NMT

[RLS87] I. R. Rubin, V. A. Lobatch, and B. V. Shulgin. “non-muffin tin” corrections of the SCF X_{α}-SW method. I. International Journal of Quantum Chemistry, 32(2):193–205, August 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Rajabali:1982:GKH

[RM82] Gholomabbas Rajabali and Robert M. Mazo. Generalization of the Kikuchi–Hijmans–de Boer method for order-disorder problems to complex lattices. International Journal of Quantum Chemistry, 22(S16):117–123, March 1–13, 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Chemistry, Theory of Condensed Matter, and Propagator Methods in the Quantum Theory of Matter.

Riera:1984:CSC

[RM84] Josep M. Riera and Josep Mariné. Chemical shifts calculation in a CNDO framework. I. Formulation and analysis of the model. International Journal of Quantum Chemistry, 25 (4):715–721, April 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Ritchie:1987:LRI

[RM87] R. H. Ritchie and J. R. Manson. Long-range interactions between probes, particles, and surfaces. International Journal of Quantum Chemistry, 32(S21):363–375, March 12, 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Chemistry, Solid-State Theory, and Computational Methods.

Ritchie:1988:E

[RM88] R. H. Ritchie and J. R. Manson. Erratum. International Journal of Quantum Chemistry, 33(6):585, June 1988. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Rao:1983:CMEb

[RMM83] Gita Subba Rao, Sudha Mahajan, and R. K. Mishra. Calculation of the minimum energy conformation of biomolecules using a global optimization technique. IV. Preferred conformations and receptor maps of histamine. International Journal of Quantum Chemistry, 24(S10):5–9, March 14–16, 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Ricard:1984:DER

[RN84] Jacques Ricard and Georges Noat. Dynamics of enzyme reactions in solution and in heterogeneous media. International Journal of Quantum Chemistry, 26(5):943–953, November 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Randic:1986:SCH

[RN86] Milan Randić and Faith E. Nettleton. On the stability of conjugated hydrocarbon ions. International Journal of Quantum Chemistry, 30(S20):203–218, March 10–15, 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Syposium on Atomic, Molecular, and Solid-State Theory, Scattering Problems, Many Body Phenomena, and Computational Quantum Chemistry.

Rabinowitz:1986:FEM

[RNW86] James R. Rabinowitz, Krishnan Namboodiri, and Harel Weinstein. A finite expansion method for the calculation and interpretation of molecular electrostatic potentials. International Journal of Quantum Chemistry, 29(6):1697–1704, June 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Roeggen:1981:APG

[Røe81a] I. Røeggen. Antisymmetric product of geminals in the context of the method of moments. International Journal of Quantum Chemistry, 19(2):319–335, February 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Roeggen:1981:ECD

[Røe81b] I. Røeggen. Electron correlation described by an extended geminal model. International Journal of Quantum Chemistry, 20(4):817–833, October 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Roeggen:1982:ECD

[Røe82] I. Røeggen. Electron correlation described by extended geminal models: The EXGEM 2 and EXGEM 3 models. International Journal of Quantum Chemistry, 22(1):149–168, July 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Roeggen:1987:ECD

[Røe87] I. Røeggen. Electron correlation described by extended geminal models: The EXGEM 4 and EXGEM 5 models. International Journal of Quantum Chemistry, 31(6):951–974, June 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Romano:1981:MCS

[Rom81] Silvano Romano. Monte-Carlo simulation of water solvent with biomolecules: Serines with reaction-field correction. International Journal of Quantum Chemistry, 20(4):921–926, October 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Romelt:1983:HRB

[Röm83] Joachim Römelt. A Hermitean reformulation of the Born– Oppenheimer nonadiabatic coupling terms for diatomic molecules. International Journal of Quantum Chemistry, 24 (6):627–631, December 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Roos:1980:CAS

[Roo80] Björn O. Roos. The complete active space SCF method in a Fock-matrix-based super-CI formulation. International Journal of Quantum Chemistry, 18(S14):175–189, March 10– 20, 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory, Collision Phenomena, Quantum Statistics, and Computational Methods.

Rosch:1987:BRB

[Rös87] N. Rösch. Book review: Rotations, quaternions and double groups. By Simon L. Altmann, Clarendon Press, Oxford, 1986, 317 pp. International Journal of Quantum Chemistry, 32(3): 401, September 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). See [Alt86, Alt05].

Raghavachari:1981:COE

[RP81a] Krishnan Raghavachari and John A. Pople. Calculation of one-electron properties using limited configuration interaction techniques. International Journal of Quantum Chemistry, 20 (5):1067–1071, November 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Raghunathan:1981:SAR

[RP81b] S. Raghunathan and Vasantha Pattabhi. Structure–activity relationships in auxins. International Journal of Quantum Chemistry, 20(1):151–158, July 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Richards:1989:RME

[RP89] Nigel G. J. Richards and Sarah L. Price. The representation of molecular electrostatics using interactive graphics. International Journal of Quantum Chemistry, 36(S16):73–85, April 1–8, 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Requena:1980:ELP

[RPS80] A. Requena, R. Peña, and A. Serna. Energy levels for perturbed Morse oscillators. International Journal of Quantum Chemistry, 17(5):931–942, May 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Requena:1982:PRR

[RPS82] A. Requena, R. Peña, and A. Serna. Perturbation for a rigid rotator in an electric field. International Journal of Quantum Chemistry, 22(6):1263–1270, December 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Richard:1987:MMC

[RR87a] Ann M. Richard and James R. Rabinowitz. Modified molecular charge similarity indices for choosing molecular analogues. International Journal of Quantum Chemistry, 31(2):309–323, February 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Riggs:1987:BSE

[RR87b] Noel V. Riggs and Leo Radom. The 3-21G(N*) basis set: An economical polarized basis set for ab initio studies on nitrogencontaining molecules. International Journal of Quantum Chemistry, 31(3):393–403, March 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Richard:1988:FED

[RR88] Ann M. Richard and James R. Rabinowitz. Further evaluation and development of charge similarity indices for choosing molecular analogues. International Journal of Quantum Chemistry, 34(3):207–216, September 1988. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Roos:1980:DCM

[RS80] Björn O. Roos and Per E. M. Siegbahn. A direct CI method with a multiconfigurational reference state. International Journal of Quantum Chemistry, 17(3):485–500, March 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Ruelle:1982:EPA

[RS82] P. Ruelle and C. Sandorfy. Electrostatic potential around halogenated anesthetics. Acidic hydrogen; chlorine vs. fluorine. International Journal of Quantum Chemistry, 22(4):691–707, October 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Ramasesha:1984:DVB

[RS84] S. Ramasesha and Z. G. Soos. Diagrammatic valence-bond theory for finite model Hamiltonians. International Journal of Quantum Chemistry, 25(6):1003–1021, June 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Radzio:1986:LLC

[RS86] E. Radzio and D. R. Salahub. LCGTO-LSD calculations for CH_{2}. International Journal of Quantum Chemistry, 29(2): 241–247, February 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Roszak:1988:ECR

[RS88] S. Roszak and W. A. Sokalski. Effects of contraction and reduction of basis set size on the He_{2} interaction energy components. International Journal of Quantum Chemistry, 34(5): 437–444, November 1988. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Ray:1985:PDP

[RSBR85] Naba K. Ray, Masayuki Shibata, Giorgio Bolis, and Robert Rein. Potential-derived point-charge model study of electrostatic interaction energies in some hydrogen-bonded systems. International Journal of Quantum Chemistry, 27(4):427–437, April 1985. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Rettrup:1982:MPG

[RSD82] S. Rettrup, C. R. Sarma, and J. P. Dahl. Molecular point group adaptation of spin–free configurations. International Journal of Quantum Chemistry, 22(1):127–148, July 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Randic:1987:REL

[RSG^{+}87] Milan Randić, Vasanth Solomon, Stuart C. Grossman, Douglas J. Klein, and Nenad Trinajsti. Resonance energies of large conjugated hydrocarbons by a statistical method. International Journal of Quantum Chemistry, 32(1):35–59, July 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Rosch:1988:TCC

[RSGK88] N. Rösch, P. Sandl, A. Görling, and P. Knappe. Toward a chemisorption cluster model using the LCGTO–X_{α} method: Application to Ni(100)/Na. International Journal of Quantum Chemistry, 34(S22):275–285, March 12–19, 1988. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Syposium on Quantum Chemistry, Solid-State Theory, and Computational Methods.

Roy:1986:UCC

[RSMM86] S. Roy, S. Sengupta, D. Mukherjee, and P. K. Mukherjee. Use of coupled-cluster based linear response theory and multireference Hermitian MBPT to IP calculations of HF. International Journal of Quantum Chemistry, 29(2):205–210, February 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Randic:1988:RSG

[RSNT88] Milan Randić, Aleksandar Sabljić, Sonja Nikolić, and Nenad Trinajstić. A rational selection of graph-theoretical indices in the QSAR. International Journal of Quantum Chemistry, 34 (S15):267–285, March 12–19, 1988. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Syposium on Quantum Biology and Quantum Pharmacology.

Radhakrishnan:1980:NIM

[RSP^{+}80] R. Radhakrishnan, Subhashini Srinivasan, C. V. Prasad, S. R. Brinda, R. D. MacElroy, and K. Sundaram. Nonbonded interactions in membrane active cyclic biopolymers. IV. cation dependence. International Journal of Quantum Chemistry, 18(S7):315–322, March 5–8, 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Raina:1981:EDE

[RSSS81] V. Raina, K. V. Sane, R. P. Saxena, and P. K. Srivastava. Estimation of dissociation energy of alkali diatomics using an extended Linderberg relation. International Journal of Quantum Chemistry, 20(4):807–816, October 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Robertson:1986:QOD

[RSST86] Donald D. Robertson, Vedene H. Smith, Jr., Alfredo M. Simas, and Ajit J. Thakkar. The quality of s-orbitals determined by least-squares fitting and constrained variational methods. International Journal of Quantum Chemistry, 30(6):717–735, December 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Robertson:1987:MBS

[RST87] Donald D. Robertson, Vedene H. Smith, Jr., and Ajit J. Thakkar. More on basis set quality tests. International Journal of Quantum Chemistry, 32(4):427–434, October 1987.

CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Reynolds:1984:ICR

[RT84] Christopher Reynolds and Colin Thomson. Ab initio calculations relevant to the mechanism of chemical carcinogenesis by N-nitrosamines. I. The nitrosation of amines. International Journal of Quantum Chemistry, 26(S11):167–181, March 12– 15, 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Reynolds:1985:ICR

[RT85] C. A. Reynolds and C. Thomson. Ab initio calculations relevant to the mechanism of chemical carcinogenesis II: The nitrous acidium ion a powerful nitrosating agent. International Journal of Quantum Chemistry, 28(S12):263–279, March 14– 16, 1985. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Reynolds:1987:HCI

[RT87] Christopher A. Reynolds and Colin Thomson. Hydrated carbonium ions as possible nitrosamine metabolites: An ab initio study. International Journal of Quantum Chemistry, 32 (1):123–131, July 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Richards:1981:SOC

[RTC81] W. G. (William Graham) Richards, H. P. Trivedi, and D. L. Cooper. Spin-orbit coupling in molecules. The International series of monographs on chemistry. Clarendon Press, New York, NY, USA, 1981. ISBN 0-19-855614-4. viii + 105 pp. LCCN QD461 .R476 1981.

Rao:1981:CME

[RTM81] Gita Subba Rao, R. S. Tyagi, and R. K. Mishra. Calculation of the minimum energy conformation of acetylcholine using a global optimization technique. International Journal of Quantum Chemistry, 20(1):273–279, July 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Rao:1983:CMEa

[RTM83] Gita Subba Rao, R. S. Tyagi, and R. K. Mishra. Calculation of the minimum energy conformations of Di-, Tri-, and tetrapeptides of glycine using a global optimization technique. International Journal of Quantum Chemistry, 24(S10):1–4, March 14–16, 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Rao:1986:CME

[RTM86] Gita Subba Rao, R. S. Tyagi, and R. K. Mishra. Calculation of the minimum energy conformation of biomolecules by using a global optimization technique. V. Preferred conformations of the thyrotropin-releasing hormone. International Journal of Quantum Chemistry, 29(5):1177–1180, May 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Rudzikas:1984:RSA

[Rud84] Z. B. Rudzikas. Relativistic studies of atoms and ions with complex configurations. International Journal of Quantum Chemistry, 25(1):47–55, January 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Ruttink:1981:GGB

[Rut81] Paul J. A. Ruttink. Generation of generalized branching diagram spin functions by Schmidt orthogonalization of spinpaired functions. International Journal of Quantum Chemistry, 19(3):413–426, March 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Rossikhin:1980:DPE

[RVK80] V. V. Rossikhin, E. O. Voronkov, and Yu. A. Kruglyak. Determination of potential-energy surface of molecules in an applied field on the basis of virial relations. International Journal of Quantum Chemistry, 17(2):327–338, February 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Randic:1980:GTA

[RW80a] Milan Randić and Charles L. Wilkins. Graph-theoretical analysis of molecular properties. Isomeric variations in nonanes.

International Journal of Quantum Chemistry, 18(4):1005– 1027, October 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Rives:1980:NHO

[RW80b] A. B. Rives and F. Weinhold. Natural hybrid orbitals: Ab initio SCF and CI results for CO and NiCO. International Journal of Quantum Chemistry, 18(S14):201–209, March 10– 20, 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory, Collision Phenomena, Quantum Statistics, and Computational Methods.

Rives:1981:NHO

[RW81] A. B. Rives and F. Weinhold. Natural hybrid orbitals: ab initio SCF and CI results for CO and NiCO. International Journal of Quantum Chemistry, 20(S15):555, March 8–14, 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory, Collision Phenomena, and Computational Quantum Chemistry.

Randic:1983:URW

[RWG83] Milan Randic, Wayne L. Woodworth, and Ante Graovac. Unusual random walks. International Journal of Quantum Chemistry, 24(5):435–452, November 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Reggio:1981:MDB

[RWOT81] Patricia H. Reggio, Harel Weinstein, Roman Osman, and Sid Topiol. Molecular determinants for binding of methylenedioxytryptamines at 5-HT/LSD receptors. International Journal of Quantum Chemistry, 20(S8):373–384, March 5–7, 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Randic:1986:GTA

[RZ86] Milan Randić and Howard E. Zimmerman. A graph theoretical approach to Möbius systems in organic chemistry. International Journal of Quantum Chemistry, 30(S20):185–201, March 10–15, 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Syposium on Atomic, Molecular, and Solid-State Theory, Scattering Problems, Many Body Phenomena, and Computational Quantum Chemistry.

Requena:1984:VRL

[RZP84] A. Requena, J. Zuñiga, and R. Peña. Vibrational–rotational levels of diatomic RKR potentials. International Journal of Quantum Chemistry, 25(4):677–686, April 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Ruo-Zhuang:1980:LMT

[RZWL80] Liu Ruo-Zhuang and Feng Wen-Lin. LCAO–MO treatment for deep levels produced by ion implantation of oxygen in gallium arsenide. International Journal of Quantum Chemistry, 18 (2):601–605, August 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Sasaki:1980:CXS

[SA80] Teikichi Sasaki and Hirohiko Adachi. Calculations of XPS spectra for oxyanions and related compounds by the discrete variational-Xα method. International Journal of Quantum Chemistry, 18(1):227–235, July 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Savatinova:1986:DSA

[SA86a] I. Savatinova and E. Anachkova. Dynamical and structural aspects of water in potassium ferrocyanide trihydrate. International Journal of Quantum Chemistry, 29(5):1383–1396, May 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Sharma:1986:MSR

[SA86b] Krishan K. Sharma and Anil K. Aggarwal. An MNDO study of the reaction of methanol with fulminic acid and acetonitrile oxide. International Journal of Quantum Chemistry, 30(2): 213–224, August 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Sabin:1980:AE

[Sab80a] John R. Sabin. Addenda and errata. International Journal of Quantum Chemistry, 18(6):1489, December 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Sabin:1980:EN

[Sab80b] John R. Sabin. Editorial note. International Journal of Quantum Chemistry, 18(2):345, August 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Slater:1980:ETR

[SAB^{+}80c] Trevor F. Slater, Mustaq Ahmed, Chiara Benedetto, Kevin Cheeseman, John E. Packer, Robin L. Willson, and Mario U. Dianzani. Electron transfer reactions in endoplasmic reticulum: Free radical production, lipid peroxidation, covalent binding and cell division. International Journal of Quantum Chemistry, 18(S7):347–356, March 5–8, 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Sankey:1981:EEF

[SAD81] Otto F. Sankey, Roland E. Allen, and John D. Dow. Effects of the environment on Frenkel core-exciton energies. International Journal of Quantum Chemistry, 20(S15):469–476, March 8–14, 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory, Collision Phenomena, and Computational Quantum Chemistry.

Sadlej:1983:CEE

[Sad83] Andrzej J. Sadlej. Correlation effects in externally perturbed many-electron systems. International Journal of Quantum Chemistry, 23(1):147–167, January 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Sadlej:1985:SEM

[Sad85] Joanna Sadlej. Semi-empirical methods of quantum chemistry. Ellis Horwood series in chemical science. Ellis Horwood, New York, NY, USA, 1985. ISBN 0-85312-167-2, 0-470-27547-2 (Halsted Press). xiv + 386 pp. LCCN QD462 .S2313 1985.

Sadlej:1989:TDW

[Sad89] Andrzej J. Sadlej. Time-dependence of wave functions and the calculation of transition intensities. International Journal of Quantum Chemistry, 35(3):409–423, March 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Sabljic:1982:SRT

[SAM82] Aleksandar Sabljić, Abe Auerbach, and Ruth McDiarmid. Symmetry of Rydberg transitions in 1,4-cyclohexadiene: Multiphoton ionization investigation. International Journal of Quantum Chemistry, 22(S16):357–362, March 1–13, 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Chemistry, Theory of Condensed Matter, and Propagator Methods in the Quantum Theory of Matter.

Sandorfy:1981:RSC

[San81] C. Sandorfy. Rydberg states in chemistry. International Journal of Quantum Chemistry, 19(6):1147–1156, June 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Sandorfy:1984:PPS

[San84] C. Sandorfy. The photochemical primary step and the potential in the proton bridge of the chromophore of visual pigments and bacteriorhodopsin. International Journal of Quantum Chemistry, 26(5):907–915, November 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Sapse:1980:ESP

[Sap80] Anne-Marie Sapse. Electronic studies of the pyrrolidine molecule with aziridine ring. International Journal of Quantum Chemistry, 18(S7):155–162, March 5–8, 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Sbrana:1981:SPT

[SAP81] G. Sbrana, L. Angeloni, and S. Panerai. Spectroscopic properties of TCNQ–bifluorenylidene 1:1 complex. International Journal of Quantum Chemistry, 19(6):1157–1160, June 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Saran:1981:CNA

[Sar81] Anil Saran. Conformation of nucleoside antibiotics. International Journal of Quantum Chemistry, 20(2):439–447, August 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Sarma:1982:SDD

[Sar82] Ramaswamy H. Sarma. Structures of DNA double helix. International Journal of Quantum Chemistry, 22(S9):65–80, March 4–6, 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Saran:1989:CBC

[Sar89] Anil Saran. Correlation between the conformation of nucleoside antibiotics and their biological activity. International Journal of Quantum Chemistry, 35(1):193–203, January 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Selmani:1986:CAL

[SAS86] A. Selmani, J. Andzelm, and D. R. Salahub. Chemisorption of O and O_{2} on Ag(110): An LCGTO–LSD cluster study. International Journal of Quantum Chemistry, 29(4):829–842, April 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Savin:1988:CDF

[Sav88] A. Savin. A combined density functional and configuration interaction method. International Journal of Quantum Chemistry, 34(S22):59–69, March 12–19, 1988. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Syposium on Quantum Chemistry, Solid-State Theory, and Computational Methods.

Smeyers:1981:IDN

[SB81] Yves G. Smeyers and M. Nieves Bellido. Internal dynamics of nonrigid molecules. I. Application to acetone. International Journal of Quantum Chemistry, 19(4):553–565, April 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Segedy:1983:CCH

[SB83a] John J. Segedy and Edward A. Burke. Calculation of ns^{2} configurations for the helium isoelectronic series without matching. International Journal of Quantum Chemistry, 23(6): 1959–1968, June 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Smeyers:1983:PES

[SB83b] Yves G. Smeyers and M. Nieves Bellido. Potential energy surface determinations for nonrigid molecules: Application to acetone. International Journal of Quantum Chemistry, 23(2): 507–516, February 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Sekino:1984:LRC

[SB84] Hideo Sekino and Rodney J. Bartlett. A linear response, coupled-cluster theory for excitation energy. International Journal of Quantum Chemistry, 26(S18):255–265, March 1– 15, 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Atomic, Molecular and Solid-State Theory, and Computational Quantum Chemistry.

Sekino:1987:CCE

[SB87] Hideo Sekino and Rodney J. Bartlett. Coupled-cluster evaluation of geometrical derivatives of properties using nonrelaxed orbitals. International Journal of Quantum Chemistry, 32(S21):487–493, March 12, 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Chemistry, Solid-State Theory, and Computational Methods.

Sneddon:1988:IGF

[SB88] S. F. Sneddon and C. L. Brooks III. The influence of geometrical fluctuations on electron tunneling barriers in proteins. International Journal of Quantum Chemistry, 34(S15): 23–32, March 12–19, 1988. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Syposium on Quantum Biology and Quantum Pharmacology.

Sun:1989:DDR

[SB89] Qiyan Sun and Joel M. Bowman. Diatom-diatom reactive scattering in hypercylindrical coordinates. International Journal of Quantum Chemistry, 36(S23):115–126, April 1–8, 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Shavitt:1987:CSE

[SBB87] Isaiah Shavitt, Franklin B. Brown, and Peter G. Burton. Configuration selection and extrapolation in multireference configuration interaction calculations: The (H_{2})_{2} van der Waals complex as a benchmark example. International Journal of Quantum Chemistry, 31(3):507–520, March 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Shepard:1983:QCC

[SBE^{+}83] Ron Shepard, Raymond A. Bair, Robert A. Eades, Albert F. Wagner, Michael J. Davis, Lawrence B. Harding, and Thom H. Dunning, Jr. Quantum chemical calculations using the Floating Point Systems, Inc. Model 164 attached processor. International Journal of Quantum Chemistry, 24(S17):613–622, March 3–12, 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-state Theory, Collision Phenomena, and Computational Quantum Chemistry.

Schatz:1987:FHE

[SBE87] George C. Schatz, Jay K. Badenhoop, and Charles W. Eaker. The formation of highly excited H in the reaction H_{2}^{+}(v) + H_{2} → H + H. International Journal of Quantum Chemistry, 31 (1):57–63, January 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Seiders:1985:QMI

[SBH^{+}85] Barbara A. Seiders, Alan A. Brimfield, Kenneth Hunter, Michael C. Zerner, George D. Purvism III, and W. Daniel Edwards. A quantum mechanical investigation into the fine specificity of the antibodies to soman, BE2, and CC1. International Journal of Quantum Chemistry, 28(S12):89–98, March 14–16, 1985. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Seiders:1986:QMI

[SBH^{+}86] Barbara A. Seiders, Alan A. Brimfield, Kenneth Hunter, Michael C. Zerner, George D. Purvis III, and W. Daniel Edwards. A quantum mechanical investigation into the fine specificity of the antibodies to soman, BE2, and CC1. International Journal of Quantum Chemistry, 29(6):1845, June 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Santamaria:1989:GSA

[SBNK89] R. Santamaría, M. Berrondo, O. Novaro, and E. S. Kryachko. Generalized statistical approach to the study of interatomic interactions. International Journal of Quantum Chemistry, 35(2):267–275, February 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Smith:1986:HFH

[SBW86] J. L. Smith, A. Michael Boring, and P. Weinberger. Heavy fermions and heavy atoms. International Journal of Quantum Chemistry, 29(3):315–321, March 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Suzuki:1980:BED

[SBWW80] I. H. Suzuki, C. E. Brion, E. Weigold, and G. R. J. Williams. Binding energies at different momenta for the valence orbitals of HCl by the binary (e, 2 e) method. International Journal of Quantum Chemistry, 18(1):275–280, July 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Saran:1980:MOS

[SC80a] Anil Saran and C. L. Chatterjee. Molecular orbital studies on nucleoside antibiotics III. Conformation of toyocamycin and sangivamycin. International Journal of Quantum Chemistry, 18(S7):123–136, March 5–8, 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Spangler:1980:ICL

[SC80b] Dale Spangler and Ralph E. Christoffersen. Ab initio calculations on large molecules using molecular fragments. Generalization of the molecular fragment basis at the minimum basis set level. International Journal of Quantum Chemistry, 17 (6):1075–1097, June 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Saran:1984:MOS

[SC84a] Anil Saran and C. L. Chatterjee. Molecular orbital studies on the structure of nucleoside analogs. IV. Conformation of 3-deazapurine nucleosides. International Journal of Quantum Chemistry, 25(4):743–752, April 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Sykja:1984:GEM

[SC84b] Besnik Sykja and Jean-Louis Calais. Gap equations for molecular systems. International Journal of Quantum Chemistry, 25(2):331–346, February 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Schuch:1986:MIM

[SC86a] D. Schuch and K.-M. Chung. From macroscopic irreversibility to microscopic reversibility via a nonlinear Schrödinger-type field equation. International Journal of Quantum Chemistry, 29(5):1561–1573, May 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Scuseria:1986:ICT

[SC86b] Gustavo E. Scuseria and Rubén H. Contreras. Ab initio calculations of through-space nuclear spin–spin coupling constants with the IPPP method. International Journal of Quantum Chemistry, 30(S20):603–612, March 10–15, 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Syposium on Atomic, Molecular, and Solid-State Theory, Scattering Problems, Many Body Phenomena, and Computational Quantum Chemistry.

Schuch:1987:MIM

[SC87] D. Schuch and K.-M. Chung. From macroscopic irreversibility to microscopic reversibility via a nonlinear Schrödinger-type field equation. International Journal of Quantum Chemistry, 31(4):695–696, April 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Sim:1989:MLC

[SC89] Fiona Sim and C. R. A. Catlow. Mott–Littleton calculations on defects in α-quartz. International Journal of Quantum Chemistry, 36(S23):651–675, April 1–8, 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Sharpe:1988:ETR

[SCER88] Paul Sharpe, Charles S. Christ, John R. Eyler, and David E. Richardson. Electron-transfer reactions of organometallic and coordination compounds in the absence of solvent: Experimental results and theoretical approaches. International Journal of Quantum Chemistry, 34(S22):601–610, March 12–19, 1988. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Syposium on Quantum Chemistry, Solid-State Theory, and Computational Methods.

Scheiner:1980:PTP

[Sch80a] Steve Scheiner. Proton transfer potentials in hydrogen-bonded systems: (H_{5}O_{2})^{+}. International Journal of Quantum Chemistry, 18(S7):199–206, March 5–8, 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Schweitz:1980:HTO

[Sch80b] Jan-Åke Schweitz. Hypevirial theorem for open dynamical assemblies in the density matrix representation. International Journal of Quantum Chemistry, 18(3):811–817, September 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Scheiner:1981:PTH

[Sch81] Steve Scheiner. Proton transfers in hydrogen-bonded systems V. Analysis of electronic redistributions in (N_{2}H_{7})^{+}. International Journal of Quantum Chemistry, 20(S8):221–229, March 5–7, 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Schlegel:1982:IMO

[Sch82] H. Bernhard Schlegel. Ab initio molecular orbital study of the tautomerism of 4-hydroxy-2-pyridinone. International Journal of Quantum Chemistry, 22(5):1041–1047, November 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Scheiner:1983:EER

[Sch83a] Steve Scheiner. Energetics and electronic rearrangements of proton transfer in (H_{3}NHOH_{2})^{+}. International Journal of Quantum Chemistry, 23(2):753–764, February 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Schmidt:1983:CPP

[Sch83b] H. M. Schmidt. On the concept of “Partitions” in the perturbation theory of n-electron systems. International Journal of Quantum Chemistry, 24(6):551–570, December 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Schuch:1984:NST

[SCH84] D. Schuch, K.-M. Chung, and H. Hartmann. Nonlinear Schrödinger-type field equation for the description of dissipative systems. II. Frictionally damped motion in a magnetic field. International Journal of Quantum Chemistry, 25(2): 391–410, February 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Schluter:1988:ESS

[Sch88] M. Schluter. Electronic structure and superconductivity of high T_{c} Cu oxides. International Journal of Quantum Chemistry, 34(S22):735–736, March 12–19, 1988. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Syposium on Quantum Chemistry, Solid-State Theory, and Computational Methods.

Schuch:1989:CBN

[Sch89] D. Schuch. Connections between Newton- and Schrödingertype equations in the description of reversible and irreversible dynamics. International Journal of Quantum Chemistry, 36 (S23):59–72, April 1–8, 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Sannigrahi:1980:SVE

[SD80] A. B. Sannigrahi and B. R. De. Selected valence electron splitshell MO calculations on Cu_{2}, Ag_{2}, and Au_{2}. International Journal of Quantum Chemistry, 17(6):1237–1239, June 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Schulman:1981:TKA

[SD81] Joel N. Schulman and John D. Dow. On the theory of the KVV Auger spectrum of Li. International Journal of Quantum Chemistry, 20(S15):437–443, March 8–14, 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory, Collision Phenomena, and Computational Quantum Chemistry.

Saran:1982:SIS

[SD82] Anil Saran and M. M. Dhingra. Structure of the intermediate species of the photoreaction cycle of rhodopsin. International Journal of Quantum Chemistry, 22(S9):145–156, March 4–6, 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Seel:1986:ASC

[SD86] M. Seel and G. Del Re. Accurate SCF computations on hydrogen bonds: Role of polarization functions on the bridge hydrogen atom. International Journal of Quantum Chemistry, 30 (4):563–566, October 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Scheiner:1988:EEI

[SD88] Steve Scheiner and T. Das. Effects of external ions upon proton transfer reactions: H-bonded systems containing HCOOH. International Journal of Quantum Chemistry, 34(S15):137– 147, March 12–19, 1988. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Syposium on Quantum Biology and Quantum Pharmacology.

Smeyers:1986:OFL

[SDBDSMG86] Yves G. Smeyers, G. Delgado-Barrio, L. Doreste-Suárez, and J. M. Martin-González. The omega function for the lithium atom. International Journal of Quantum Chemistry, 29(6): 1705–1712, June 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Starodub:1983:EPC

[SDS83] V. A. Starodub, T. B. Drushlyak, and L. A. Sleta. Electron– phonon coupling detection in infrared spectra of ion–radical salts based on dibenztetrathiofulvalen. International Journal of Quantum Chemistry, 24(2):149–159, August 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Seel:1981:MCO

[See81] M. Seel. From molecular crystals and organic polymers to chemisorption processes: Importance of local interactions. International Journal of Quantum Chemistry, 19(6):1083–1098, June 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Seel:1984:EIQ

[See84] M. Seel. The effect of an impurity in quasi-one-dimensional systems: Green’s function and cluster model approach in comparison. International Journal of Quantum Chemistry, 26(5): 753–768, November 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Seel:1988:ACC

[See88] Max Seel. Atomic clusters and cluster models in solid state physics. International Journal of Quantum Chemistry, 34 (S22):265–274, March 12–19, 1988. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Syposium on Quantum Chemistry, Solid-State Theory, and Computational Methods.

Sellers:1986:AFCa

[Sel86a] Harrell Sellers. Analytical force constant calculation as a minimization problem. International Journal of Quantum Chemistry, 30(3):433–436, September 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Sellers:1986:AFCb

[Sel86b] Harrell Sellers. Analytical force constant calculation as a minimization problem. International Journal of Quantum Chemistry, 30(5):713, November 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Sellers:1988:VED

[Sel88] Harrell Sellers. Variational energy derivatives and perturbation theory. International Journal of Quantum Chemistry, 33 (4):271–277, April 1988. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Sen:1980:ARI

[Sen80a] K. D. Sen. Approximate relationship for isoelectronic changes in atomic average electron density. International Journal of Quantum Chemistry, 17(6):1241–1242, June 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Sen:1980:EES

[Sen80b] K. D. Sen. Energy-eigenvalue sum relationship in molecules. International Journal of Quantum Chemistry, 18(3):907–909, September 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Sen:1984:SES

[Sen84] K. D. Sen. Semiempirical estimates of static multipole polarizabilities for free atoms. International Journal of Quantum Chemistry, 26(6):1051–1052, December 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Sen:1985:EEE

[Sen85] K. D. Sen. Empirical estimates of electrostatic potentials at the nuclei in molecules. International Journal of Quantum Chemistry, 27(2):231–232, February 1985. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Serre:1981:MMB

[Ser81] Josiane Serre. Metal–metal bonds. International Journal of Quantum Chemistry, 19(6):1171–1183, June 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Serre:1984:TAC

[Ser84] Josiane Serre. Theoretical approach of the chemical reactivity: The S_{N} 2 reaction. International Journal of Quantum Chemistry, 26(5):593–605, November 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Setser:1979:RIG

[Set79] D. W. Setser, editor. Reactive intermediates in the gas phase: generation and monitoring. Academic Press, New York, USA, 1979. ISBN 0-12-637450-3. x + 358 pp. LCCN QD501 .R346.

Sewell:1986:QTC

[Sew86] G. L. Sewell. Quantum theory of collective phenomena. Monographs on the physics and chemistry of materials. Clarendon Press, New York, NY, USA, 1986. ISBN 0-19-851371-2. xiii + 229 pp. LCCN ????

Seybold:1980:ARF

[Sey80] Paul G. Seybold. Aryloxy radical formation as a deactivating process in chemical carcinogenesis: Phenolic derivatives of 7,12-dimethylbenz[a]anthracene and 5-methylchrysene. International Journal of Quantum Chemistry, 18(S7):271–276, March 5–8, 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Seydel:1981:MAQ

[Sey81] J. K. Seydel. Mode of action and quantitative structure– activity relationship of sulfonamides in biological systems of different complexity (enzymes, bacteria, rat, and human). International Journal of Quantum Chemistry, 20(1):131–150, July 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Seybold:1983:TDO

[Sey83a] Paul G. Seybold. Topological determinants in organic chemistry. International Journal of Quantum Chemistry, 23(4): 1683, April 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Seybold:1983:TICa

[Sey83b] Paul G. Seybold. Topological influences on the carcinogenicity of aromatic hydrocarbons. I. The bay region geometry. International Journal of Quantum Chemistry, 24(S10):95–101, March 14–16, 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Seybold:1983:TICb

[Sey83c] Paul G. Seybold. Topological influences on the carcinogenicity of aromatic hydrocarbons. II. Substituent effects. International Journal of Quantum Chemistry, 24(S10):103–108, March 14–16, 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Shibuya:1980:PEM

[SF80a] Tai-Ichi Shibuya and Tadashi Funada. Parametrization of the equations-of-motion method for conjugated hydrocarbons. International Journal of Quantum Chemistry, 18(1):205–209, July 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Steinborn:1980:EMI

[SF80b] E. Otto Steinborn and Eckhard Filter. Evaluation of multicenter integrals over Slater-type atomic orbitals by expansion in terms of complete sets. International Journal of Quantum Chemistry, 18(1):219–226, July 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Sheppard:1981:TOQ

[SF81a] Maurice G. Sheppard and Karl F. Freed. Third-order quasidegenerate many-body perturbation theory calculations for valence state correlation energies of the nitrogen and oxygen atoms and their ions. International Journal of Quantum Chemistry, 20(S15):21–31, March 8–14, 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory, Collision Phenomena, and Computational Quantum Chemistry.

Sudhindra:1981:IIP

[SF81b] B. S. Sudhindra and J.-H. Fuhrhop. Intermolecular interactions in porphin dimers: a quantum mechanical study. International Journal of Quantum Chemistry, 20(3):747–753, September 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Scolfaro:1989:NDT

[SF89] Luisa M. R. Scolfaro and A. Fazzio. Native defects and transition metal impurities at interstitial sites in GaAs. International Journal of Quantum Chemistry, 36(S23):677–685, April 1–8, 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Sapse:1986:ISI

[SFC86] A. M. Sapse, L. M. Fugler, and D. Cowburn. An ab initio study of intermolecular hydrogen bonding between small peptide fragments. International Journal of Quantum Chemistry, 29(5):1241–1251, May 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Seybold:1980:MOS

[SG80a] Paul G. Seybold and Astrid Gräuslund. A molecular orbital study of the metabolism and carcinogenicity of the phenols of benzo(a)pyrene. International Journal of Quantum Chemistry, 18(S7):261–270, March 5–8, 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Szent-Gyorgyi:1980:LSC

[SG80b] A. Szent-Györgyi. The living state and cancer. International Journal of Quantum Chemistry, 18(S7):217–222, March 5–8, 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Sheridan:1981:NRS

[SG81a] Robert P. Sheridan and Raj K. Gupta. ^{1}H nuclear relaxation study of the aniline-binding site in human hemoglobin. International Journal of Quantum Chemistry, 20(S8):257–264, March 5–7, 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Sheridan:1981:SSN

[SG81b] Robert P. Sheridan and Raj K. Gupta. Spectroscopic studies of neocarzinostatin and its chromophore: UV inactivation and mercaptan activation. International Journal of Quantum Chemistry, 20(S8):331–346, March 5–7, 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Szent-Gyorgyi:1982:BOC

[SG82] Albert Szent-Györgyi. Biological oxidation and cancer. International Journal of Quantum Chemistry, 22(S9):27–30, March 4–6, 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Szent-Gyorgyi:1984:P

[SG84] Albert Szent-Györgyi. Protometabolism. International Journal of Quantum Chemistry, 26(S11):63–67, March 12–15, 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Staemmler:1985:ASP

[SG85a] V. Staemmler and F. A. Gianturco. Adiabatic SCF potential energy curves relevant to proton–oxygen molecular collisions. International Journal of Quantum Chemistry, 28(5):553–564, November 1985. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Szent-Gyorgyi:1985:MC

[SG85b] Albert Szent-Gyorgyi. Metabolism and cancer. International Journal of Quantum Chemistry, 28(S12):257–261, March 14– 16, 1985. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Stamato:1986:CSA

[SG86] Fulvia M. L. G. Stamato and Julia M. Goodfellow. Computer simulations of amino acid zwitterions in solution. International Journal of Quantum Chemistry, 30(S13):277–286, March 6–8, 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Sinha:1987:GVC

[SG87] S. K. Sinha and R. Ghosh. A gauge-variational calculation of molecular diamagnetic susceptibility. International Journal of Quantum Chemistry, 32(5):575–589, November 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Shim:1989:CGI

[SG89] I. Shim and K. A. Gingerich. The chromium-group IV molecules CrC, CrSi, and CrGe: An all-electron ab initio HF– CI study. International Journal of Quantum Chemistry, 36 (S23):409–419, April 1–8, 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Sangfelt:1981:AAP

[SGEK81] Erland Sangfelt, Osvaldo Goscinski, Nils Elander, and Henry Kurtz. An analysis of the AGP and projected AGP wavefunctions. International Journal of Quantum Chemistry, 20(S15): 133–141, March 8–14, 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory, Collision Phenomena, and Computational Quantum Chemistry.

Sukhorukov:1980:ISO

[SGP^{+}80] B. I. Sukhorukov, I. Ya. Gukovsky, A. I. Petrov, A. S. Gukovskaya, A. A. Mayevsky, and N. M. Gusenkova. Interactions and self-organization of nucleic bases, nucleotides, and polynucleotides into ordered structures. Effect of ionization and solvent salt composition. International Journal of Quantum Chemistry, 17(2):339–359, February 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Smith:1987:OPA

[SH87a] C. M. Smith and G. G. Hall. Optimal population analysis. International Journal of Quantum Chemistry, 31(4):685–692, April 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Svrcek:1987:UTM

[SH87b] Michal Svrček and Ivan Hubač. Unitary transformation on the model space. Comparison of Brandow’s and Kirtman’s versions of quasidegenerate many body perturbation theory. International Journal of Quantum Chemistry, 31(4):625–638, April 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Saito:1989:CDP

[SH89] Nakaya Saito and George G. Hall. The change in the dressed potential of a polyatomic molecule in intense photon fields: Simple rules based on the nuclear charge-mass ratio. International Journal of Quantum Chemistry, 35(2):283–295, February 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Sham:1985:DFB

[Sha85] L. J. Sham. Density functionals beyond the local density approximation. International Journal of Quantum Chemistry, 28 (S19):491–495, March 18–23, 1985. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Shepard:1987:GED

[She87] Ron Shepard. Geometrical energy derivative evaluation with MRCI wave functions. International Journal of Quantum Chemistry, 31(1):33–44, January 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Scanlan:1983:PBL

[SHH^{+}83a] M. J. Scanlan, I. H. Hillier, E. E. Hodgkin, R. P. Sidebotham, C. M. Warwick, and R. H. Davies. Prediction of bond length variation in classical β-lactam structures and enzymatic ring instability. International Journal of Quantum Chemistry, 24 (S10):231–249, March 14–16, 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Solmajer:1983:IMI

[ŠHH83b] T. Šolmajer, M. Hodošček, and D. Hadži. Ab initio MO investigation of the ethanolamine–formic acid complex. International Journal of Quantum Chemistry, 23(3):945–952, March 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Shipsey:1984:AAE

[Shi84] Edward J. Shipsey. Asymptotic analysis of exchange interactions. International Journal of Quantum Chemistry, 25 (5):891–914, May 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Shimizu:1985:MEM

[Shi85] Mikio Shimizu. Molecular electric mechanism for discrimination of amino acid by cognate tRNA. International Journal of Quantum Chemistry, 28(S12):281–286, March 14–16, 1985. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Shirts:1987:RDV

[Shi87] Randall B. Shirts. Rotational decoupling of vibrational modes due to rotational excitation: Application to a model of the HDO stretching motions. International Journal of Quantum Chemistry, 31(1):119–131, January 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Sokalski:1987:CIM

[SHK87a] W. A. Sokalski, P. C. Hariharan, and Joyce J. Kaufman. Comparison of ab initio MODPOT interaction energy components against large basis set MBPT(4) calculations for nitromethane dimer. International Journal of Quantum Chemistry, 32(S21): 645–660, March 12, 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Chemistry, Solid-State Theory, and Computational Methods.

Sokalski:1987:LCA

[SHK87b] W. A. Sokalski, P. C. Hariharan, and Joyce J. Kaufman. Library of cumulative atomic multipole moments. I. Nucleic acid bases. International Journal of Quantum Chemistry, 32(S14): 111–126, March 14, 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Shibata:1980:EDW

[SHKM80] S. Shibata, F. Hirota, N. Kakuta, and T. Muramatsu. Electron distribution in water by high-energy electron scattering. International Journal of Quantum Chemistry, 18(1):281–285, July 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Sokalski:1980:BSS

[SHKP80] W. A. Sokalski, P. C. Hariharan, Joyce J. Kaufman, and Carlo Petrongolo. Basis set superposition effect on difference electrostatic molecular potential contour maps. International Journal of Quantum Chemistry, 18(1):165–171, July 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Smeyers:1982:FCI

[SHL82] Y. G. Smeyers and A. Hernández-Laguna. Further calculations on the internal rotation spectrum of phenol. International Journal of Quantum Chemistry, 22(4):681–690, October 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Smeyers:1986:CQM

[SHL86] Yves G. Smeyers and A. Hernández-Laguna. Classical and quantum mechanical calculations of conformational probability density distributions. Two-rotor molecules. International Journal of Quantum Chemistry, 29(3):553–562, March 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Sun:1983:LGU

[SHLL83] Chia-Chung Sun, Yan-De Han, Be-Fu Li, and Qian-Shu Li. Lie groups and unitary transformations in ligand field theory. International Journal of Quantum Chemistry, 23(1):169–183, January 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Sokalski:1980:MCN

[SHP^{+}80] W. A. Sokalski, P. C. Hariharan, Herbert E. Popkie, Joyce J. Kaufman, and Carlo Petrongolo. Molecular calculations with the nonempirical ab initio MODPOT, VRDDO, and MOD-POT /VRDDO procedures. XI. Theoretical study of the [C_{6}H_{5}OH·OC_{6}H_{5}]^{−} molecular complex: Ab initio MODPOT /VRDDO calculations and electrostatic molecular potential contour maps. International Journal of Quantum Chemistry, 18(1):173–184, July 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Sokalski:1980:EAS

[SHPK80] W. A. Sokalski, P. C. Hariharan, Herbert E. Popkie, and Joyce J. Kaufman. Explanation of the artifact structure predictions within the semiempirical ZDO SCF supermolecular approach. International Journal of Quantum Chemistry, 18 (1):189–191, July 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Sibert:1982:LME

[SHRH82] Edwin L. Sibert III, John S. Hutchinson, William P. Reinhardt, and James T. Hynes. Local mode energy transfer: Ebb and flow. International Journal of Quantum Chemistry, 22 (S9):375–383, March 4–6, 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Stroyer-Hansen:1984:FDN

[SHS84] T. Stroyer-Hansen and E. Nørby Svendsen. Frequencydependent nonresonance Raman scattering. International Journal of Quantum Chemistry, 26(S18):519–524, March 1– 15, 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Atomic, Molecular and Solid-State Theory, and Computational Quantum Chemistry.

Shull:1984:SIR

[Shu84] Harrison Shull. Science and international relations: Competitive cooperation. International Journal of Quantum Chemistry, 26(S18):25–32, March 1–15, 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Atomic, Molecular and Solid-State Theory, and Computational Quantum Chemistry.

Sekino:1989:RDP

[SI89a] Hideo Sekino and Yasuyuki Ishikawa. Relativistic diagrammatic perturbation theory calculations on neon atom. International Journal of Quantum Chemistry, 36(S23):339–344, April 1–8, 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Suzuki:1989:ARM

[SI89b] Masato Suzuki and Akira Imamura. Application of the recursion method to the study of the end effects on the bond length alternation in polyacetylene. International Journal of Quantum Chemistry, 35(5):613–628, May 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Siegbahn:1980:DCI

[Sie80] Per E. M. Siegbahn. Direct configuration interaction with a reference state composed of many reference configurations. International Journal of Quantum Chemistry, 18(5):1229–1242, November 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Siegbahn:1981:BRB

[Sie81] Hans Siegbahn. Book review: Reactive Intermediates in the Gas Phase, Generation and Monitoring. Edited by D. W. Setser. Academic, New York, 1979. International Journal of Quantum Chemistry, 20(6):1333–1334, December 1981. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). See [Set79].

Siegbahn:1983:BRB

[Sie83a] Per Siegbahn. Book review: Polyatomic molecules: Results of ab initio calculations. By Robert S. Mulliken and Walter C. Ermler, Academic Press, New York, 1981. Price: $49.00. International Journal of Quantum Chemistry, 24(1):135, July 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). See [ME81].

Siegbahn:1983:ECC

[Sie83b] Per E. M. Siegbahn. The externally contracted CI method applied to N_{2}. International Journal of Quantum Chemistry, 23(6):1869–1889, June 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Siedentop:1987:LR

[Sie87] Heinz K. H. Siedentop. On the localization of resonances. International Journal of Quantum Chemistry, 31(5):795–821, May 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Schulz:1985:RIF

[SIJ85] Joachim Schulz, Rüdiger Iffert, and Karl Jug. On the rotational invariance of the Fock equations in INDO methods using d functions. International Journal of Quantum Chemistry, 27 (4):461–464, April 1985. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Silverman:1980:VCP

[Sil80] Jeremiah N. Silverman. Variational constraints on perturbed eigenvalues and their perturbation expansions. II. Perturbational inequalities. International Journal of Quantum Chemistry, 18(3):867–890, September 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Silverstone:1982:UDP

[Sil82] Harris J. Silverstone. Unified derivation of the perturbation series for the real and imaginary parts of the energy of hydrogen in the Stark effect and of the negatively anharmonic oscillator. International Journal of Quantum Chemistry, 21(1): 125–131, January 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Silverman:1984:GHF

[Sil84] Jeremiah N. Silverman. Generalized Hellmann–Feynman and curvature theorems in classical and quantum mechanics. International Journal of Quantum Chemistry, 25(5):915–928, May 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Silverstone:1986:RCA

[Sil86] Harris J. Silverstone. Reality and complexity in asymptotic expansions for eigenvalues and eigenfunctions, with application to the JWKB connection-formula problem. International Journal of Quantum Chemistry, 29(2):261–272, February 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Simons:1980:CCR

[Sim80] Jack Simons. The complex coordinate rotation method and exterior scaling: A simple example. International Journal of Quantum Chemistry, 18(S14):113–121, March 10–20, 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory, Collision Phenomena, Quantum Statistics, and Computational Methods.

Simons:1981:SMR

[Sim81] Jack Simons. The Siegert methods in resonance scattering: Relation to L^{2} methods. International Journal of Quantum Chemistry, 20(3):779–780, September 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Simon:1982:LOS

[Sim82a] Barry Simon. Large orders and summability of eigenvalue perturbation theory: a mathematical overview. International Journal of Quantum Chemistry, 21(1):3–25, January 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Simons:1982:EPS

[Sim82b] Jack Simons. Electron propagator studies of molecular anions. International Journal of Quantum Chemistry, 22(S16): 575–581, March 1–13, 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Chemistry, Theory of Condensed Matter, and Propagator Methods in the Quantum Theory of Matter.

Simonetta:1986:TAS

[Sim86] Massimo Simonetta. The theoretical approach to surface chemistry and heterogeneous catalysis. International Journal of Quantum Chemistry, 29(5):1555–1560, May 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Shukla:1980:MPS

[SIN80a] G. C. Shukla, M. R. Issa, and M. S. Narayan. Multipole polarizabilities and shielding factors of the hydrogen atom from the hydrodynamic analogy to quantum mechanics. III. International Journal of Quantum Chemistry, 18(3):819–825, September 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Sinanoglu:1980:STP

[Sin80b] Oktay Sinanoğlu. The solvophobic theory for the prediction of molecular conformations and biopolymer bindings in solutions with recent direct experimental tests. International Journal of Quantum Chemistry, 18(2):381–392, August 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Singh:1981:DNJ

[Sin81] Jai Singh. Dynamic nonlinear Jahn–Teller effect of the type E ⊗ ϵ. International Journal of Quantum Chemistry, 19 (5):859–871, May 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Singh:1986:CEP

[Sin86] Jai Singh. Composite exciton-phonon states in molecular crystals. International Journal of Quantum Chemistry, 30(1):7– 18, July 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Sinanoglu:1988:TER

[Sin88] Oktay Sinanoğlu. Topological electronic rules for polycyclic hydrocarbons — quantum chemical deductions directly from structural formulas. International Journal of Quantum Chemistry, 34(S22):143–147, March 12–19, 1988. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Syposium on Quantum Chemistry, Solid-State Theory, and Computational Methods.

Sinanoglu:1989:ORT

[Sin89] Oktay Sinanoğlu. Are oxygen rings (O_{n}) and their negative ions (O) unstable? International Journal of Quantum Chemistry, 36(5):545–552, November 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Simons:1984:GDD

[SJ84] Jack Simons and Poul Jørgensen. Geometrical derivatives of dipole moments and polarizabilities. International Journal of Quantum Chemistry, 25(6):1135–1150, June 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Sapse:1985:ISF

[SJ85] A. M. Sapse and Duli C. Jain. Ab initio studies of F^{−} (H_{2}O)_{n} and Cl^{−} (H_{2}O)_{n} clusters for n = 1, 2. International Journal of Quantum Chemistry, 27(3):281–292, March 1985. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Sapse:1986:GAA

[SJ86] A. M. Sapse and Duli C. Jain. Guanine and adenine-amino acids interactions: An ab initio study. International Journal of Quantum Chemistry, 29(1):23–29, January 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Sapse:1988:CPA

[SJ88] A. M. Sapse and Duli C. Jain. The complexes of phenyl acetylene with HF, H_{2}O, and NH_{3}: An ab initio study. International Journal of Quantum Chemistry, 33(2):69–76, February 1988. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Shusterman:1989:CMD

[SJH89] Alan J. Shusterman, A. Sherwood Johnson, and Corwin Hansch. Correlation of mutagenicity of 1,1-dimethyl-3-(X-phenyl)- triazenes with molecular orbital energies and hydrophobicity. International Journal of Quantum Chemistry, 36(1):19– 33, July 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Sjostrand:1987:RMA

[Sjö87] Johannes Sjöstrand. Resonances and microlocal analysis. International Journal of Quantum Chemistry, 31(5):733–737, May 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Shillady:1988:TSG

[SJRP88] D. D. Shillady, P. Jena, B. K. Rao, and M. R. Press. Theoretical study of the geometry of Mn_{5}. International Journal of Quantum Chemistry, 34(S22):231–236, March 12–19, 1988. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Syposium on Quantum Chemistry, Solid-State Theory, and Computational Methods.

Seger:1980:ATD

[SK80] G. Seger and E. Kochanski. Approximate theoretical determination of molecular static polarizabilities. International Journal of Quantum Chemistry, 17(5):955–968, May 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Soni:1981:STP

[SK81] M. L. Soni and R. C. Kapoor. Some thermodynamic parameters for hydroxy amino acids: Bicine and tricine. International Journal of Quantum Chemistry, 20(2):385–391, August 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Saito:1987:IMS

[SK87] Minoru Saito and Hiroshi Kashiwagi. Ab initio MO study of Fe out-of-plane displacement and Fe-ligand vibration in five-coordinate Fe-porphyrin. International Journal of Quantum Chemistry, 32(S21):661–668, March 12, 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Chemistry, Solid-State Theory, and Computational Methods.

Skala:1983:CCM

[Ská83] L. Skála. Calculation of the coherent memory functions from the eigenvalues and eigenvectors of the corresponding stationary Schrödinger equation. International Journal of Quantum Chemistry, 23(2):497–505, February 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Skancke:1984:IBT

[Ska84] Per N. Skancke. On the interplay between theory and experiment in molecular structure problems: Some case studies. International Journal of Quantum Chemistry, 26(5):729–741, November 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Shibata:1981:CAS

[SKEDR81] Masayuki Shibata, Thomas Kieber-Emmons, Shibani Dutta, and Robert Rein. A comparative analysis of the stability of the chymotrypsin charge triad. International Journal of Quantum Chemistry, 20(S8):409–417, March 5–7, 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Shibata:1983:CTT

[SKER83] Masayuki Shibata, Thomas Kieber-Emmons, and Robert Rein. Comparison of theoretical and thermodynamic values of interconstituent interactions in DNA. International Journal of Quantum Chemistry, 23(4):1283–1293, April 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Solmajer:1981:CIA

[ŠKH81] T. Šolmajer, D. Kocjan, and D. Hadži. Catecholamine interaction with anionic sites—A model study. International Journal of Quantum Chemistry, 20(6):1225–1231, December 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Sasaki:1982:SAE

[SKIK82] Yorihiko Sasaki, Shigehiro Konaka, Takao Iijima, and Masao Kimura. Small-angle electron scattering and electron density in carbon dioxide. International Journal of Quantum Chemistry, 21(2):475–485, February 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Suhai:1980:ESC

[SKL80] Sándor Suhai, Johannes Kaspar, and János Ladik. Effect of side-chain disorder on the electronic structure of proteins. International Journal of Quantum Chemistry, 17(5):995–1006, May 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Sanyal:1981:IES

[SKR81] Nitish K. Sanyal, Upendra Kumar, and Mihir Roychoudhury. Interaction energy studies for the specificity of transcription. International Journal of Quantum Chemistry, 20(1):179–183, July 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Simonis:1984:IBA

[ŠKR84] V. Č. Šimonis, J. M. Kaniauskas, and Z. B. Rudzikas. Isospin basis for atoms in relativistic approximation. International Journal of Quantum Chemistry, 25(1):57–62, January 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Shinagawa:1983:MGB

[SKS83] Yasuko Shinagawa, Seiji Koyama, and Yoshiya Shinagawa. Molecular geometry of benzosemiquinone radicals: Radicals from resorcinol and its derivatives. International Journal of Quantum Chemistry, 23(2):477–481, February 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Spanget-Larsen:1980:AHC

[SL80] Jens Spanget-Larsen. Assignment of the hyperfine constants of the 1,2-benzosemiquinone radical anion: Comment on “ INDO studies on the structure of benzosemiquinone radicals”. International Journal of Quantum Chemistry, 18(2):365–367, August 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Seydel:1981:ALM

[SL81] U. Seydel and B. Lindner. Application of the laser microprobe mass analyzer (LAMMA) to qualitative and quantitative single cell analysis. International Journal of Quantum Chemistry, 20(2):505–512, August 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Sierraalta:1986:NCF

[SL86] Aníabal Sierraalta and Eduardo V. Ludeña. The nonlocal correlation function G(1, 2) in density functional theory. International Journal of Quantum Chemistry, 30(S20):277–287, March 10–15, 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Syposium on Atomic, Molecular, and Solid-State Theory, Scattering Problems, Many Body Phenomena, and Computational Quantum Chemistry.

Seybold:1987:PEA

[SL87] Paul G. Seybold and Kenny B. Lipkowitz. The peri effect in aromatic hydrocarbon carcinogenesis: An empirical force field examination. International Journal of Quantum Chemistry, 31(6):847–853, June 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Slanina:1983:ACI

[Sla83a] Zdeněk Slanina. Adsorption-complex isomerism and quantumchemical studies in heterogeneous catalysis. International Journal of Quantum Chemistry, 23(4):1563–1570, April 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Slanina:1983:SIA

[Sla83b] Zdeněk Slanina. Sequential isomerism of activated complexes: Interfering versus noninterfering intermediate. International Journal of Quantum Chemistry, 23(4):1553–1561, April 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Slanina:1984:ACI

[Sla84] Zdeněk Slanina. Adsorption-complex isomerism and quantumchemical studies in heterogeneous catalysis. International Journal of Quantum Chemistry, 26(3):439, September 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Slanina:1985:LBP

[Sla85] Zdeněk Slanina. Limiting behavior of partition functions of models with quadratically dependent energy spectra and consequences regarding theoretical study of chemical reactivity. International Journal of Quantum Chemistry, 27(6):691–697, June 1985. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Sun:1981:LAM

[SLT81] Chia-Chung Sun, Bei-Foo Li, and Au-Chin Tang. Lie algebra and molecular shell model. II. International Journal of Quantum Chemistry, 20(S15):305–320, March 8–14, 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory, Collision Phenomena, and Computational Quantum Chemistry.

Sun:1984:LAU

[SLT84a] Chia-Chung Sun, Xue-Qui Li, and Au-Chin Tang. Lie algebra and unitarily invariant decomposition of Hermitian operators. International Journal of Quantum Chemistry, 25(4):653–666, April 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Sun:1984:UID

[SLT84b] Chia-Chung Sun, Xue-Qui Li, and Au-Chin Tang. On the unitarily invariant decomposition of Hermitian operators. International Journal of Quantum Chemistry, 25(6):1045–1054, June 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Srivastava:1980:ESN

[SM80] S. K. Srivastava and P. C. Mishra. Electronic spectra of nucleic acid bases: Variable electronegativity calculations and an analysis of existing results. International Journal of Quantum Chemistry, 18(3):827–840, September 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Swaminathan:1982:SCE

[SM82] P. K. Swaminathan and David A. Micha. A self-consistent eikonal treatment of photodissociation by visible radiation. International Journal of Quantum Chemistry, 22(S16):377– 390, March 1–13, 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Chemistry, Theory of Condensed Matter, and Propagator Methods in the Quantum Theory of Matter.

Swaminathan:1983:MPV

[SM83] P. K. Swaminathan and David A. Micha. Molecular photodissociation by visible and ultraviolet radiation: Time evolution and state-to-state cross sections. International Journal of Quantum Chemistry, 24(S17):192, March 3–12, 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-state Theory, Collision Phenomena, and Computational Quantum Chemistry.

Stolarczyk:1984:CCM

[SM84] Leszek Z. Stolarczyk and Hendrik J. Monkhorst. Coupledcluster method with optimized reference state. International Journal of Quantum Chemistry, 26(S18):267–291, March 1– 15, 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Atomic, Molecular and Solid-State Theory, and Computational Quantum Chemistry.

Schmelzer:1985:GAE

[SM85] A. Schmelzer and J. N. Murrell. The general analytic expression for S_{4}-symmetry-invariant potential functions of tetraatomic homonuclear molecules. International Journal of Quantum Chemistry, 28(2):287–295, August 1985. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Srivastava:1987:CTC

[SM87a] Deepak Srivastava and David A. Micha. Collisional time correlation function approach to the interaction of light with a polyatomic system. International Journal of Quantum Chemistry, 32(S21):229–237, March 12, 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Chemistry, Solid-State Theory, and Computational Methods.

Stodden:1987:GWF

[SM87b] C. D. Stodden and D. A. Micha. Generating wave functions from classical trajectory calculations: The divergence of streamlines. International Journal of Quantum Chemistry, 32(S21):239–244, March 12, 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Chemistry, Solid-State Theory, and Computational Methods.

Sannigrahi:1983:CES

[SMD83] A. B. Sannigrahi, D. C. Mukherjee, and B. R. De. Calculation of excited-state properties of some small molecules: Comparative study of the CNDO/S and CNDO/2- v_{n−1} potential methods. International Journal of Quantum Chemistry, 23 (5):1695–1705, May 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Smeyers:1984:BRB

[Sme84] Yves G. Smeyers. Book review: ETO multicenter molecular integrals. Edited by Charles A. Weatherford and Herbert W. Jones, D. Reidel P., Dordrecht, Holland, 1982. ISBN 90-277-1488-6. International Journal of Quantum Chemistry, 26(2): 293, August 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). See [WJ82].

Smedarchina:1986:SNE

[Sme86] Z. Smedarchina. Strong nonadiabatic effects in reactions of photoisomerization. International Journal of Quantum Chemistry, 29(5):1429–1435, May 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Sokalski:1989:LCA

[SMHK89] W. A. Sokalski, K. Maruszewski, P. C. Hariharan, and Joyce J. Kaufman. Library of cumulative atomic multipole moments: II. neutral and charged amino acids. International Journal of Quantum Chemistry, 36(S16):119–164, April 1–8, 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Smith:1989:FRB

[Smi89] Vedene H. Smith, Jr. On a functional relationship between Compton profiles and stopping powers. International Journal of Quantum Chemistry, 36(S23):553–555, April 1–8, 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Smentek-Mielczarek:1983:EEC

[SMJ83] L. Smentek-Mielczarek and K. Jankowski. Effect of electron correlation on the forced electric-dipole transition probabilities in f^{N} systems. III. Some aspects of the mechanism of hypersensitive transitions. International Journal of Quantum Chemistry, 24(S17):339–346, March 3–12, 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-state Theory, Collision Phenomena, and Computational Quantum Chemistry.

Szymanski:1986:INC

[SMKT86] Klaus Szymanski, Wolfgang R. Müller, Jan V. Knop, and Nenad Trinajstić. On the identification numbers for chemical structures. International Journal of Quantum Chemistry, 30 (S20):173–183, March 10–15, 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Syposium on Atomic, Molecular, and Solid-State Theory, Scattering Problems, Many Body Phenomena, and Computational Quantum Chemistry.

Sanchez-Marin:1987:ASH

[SMMM87] J. Sanchez-Marin, J. P. Malrieu, and D. Maynau. Approximate solutions of Heisenberg Hamiltonians. International Journal of Quantum Chemistry, 31(6):903–925, June 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Simas:1984:ACH

[SMS84] Alfredo M. Simas, Janusz Mrozek, and Vedene H. Smith, Jr. Appraisal of computer hardware and software for quantumchemical calculations. International Journal of Quantum Chemistry, 26(S18):619–638, March 1–15, 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Atomic, Molecular and Solid-State Theory, and Computational Quantum Chemistry.

Sarma:1980:SDL

[SMSA80] M. H. Sarma, C. K. Mitra, Ramaswamy H. Sarma, and Sydney Archer. Structure of DNA-lucanthone intercalated complex. International Journal of Quantum Chemistry, 18(S7): 1–10, March 5–8, 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Schwarz:1986:CUDb

[SMVV86a] W. H. E. Schwarz, L. Mensching, P. Valtazanos, and W. Von Niessen. A chemically useful definition of electron difference densities. International Journal of Quantum Chemistry, 30(3):

439–444, September 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Schwarz:1986:CUDa

[SMVV86b] W. H. E. Schwarz, L. Mensching, P. Valtaznos, and W. Von Niessen. A chemically useful definition of electron difference densities. International Journal of Quantum Chemistry, 29 (4):909–914, April 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Shim:1986:ESC

[SNNG86] Irene Shim, H. M. Nagarathna-Naik, and Karl A. Gingerich. Electronic structure calculations for the molecules Si_{2} and Ge_{2} using all electron ab initio HF–CI methods. International Journal of Quantum Chemistry, 29(4):975–991, April 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Shoda:1986:ESL

[SNNO86] Takayuki Shoda, Takeshi Noro, Tsutomu Nomura, and Kimio Ohno. The electronic structure of low-lying excited states of two model systems of retinal. International Journal of Quantum Chemistry, 30(2):289–302, August 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Surjan:1982:CAB

[SNSM82] Peter R. Surján, Gábor Náray-Szabó, and István Mayer. Conformational analysis by bond orbitals with delocalization corrections: Rotation of the ser-195 side chain in α-chymotrypsin. International Journal of Quantum Chemistry, 22(5):929–938, November 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Sabin:1985:TLE

[SO85] John R. Sabin and J. Oddershede. Theoretical low energy proton stopping powers: Effects of sample phase. International Journal of Quantum Chemistry, 28(S19):733–734, March 18– 23, 1985. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Sokalski:1981:TME

[Sok81] W. Andrzej Sokalski. Theoretical model for exploration of catalytic activity of enzymes and design of new catalysts: CO_{2} hydration reaction. International Journal of Quantum Chemistry, 20(1):231–240, July 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Sapse:1984:ISNa

[SOS84a] Anne-Marie Sapse, Louis Osorio, and Grace Snyder. Ab initio studies of negative ion-molecule(s) clusters present in the atmosphere. II. OH^{−} (H_{2}O)_{n} for n = 0, 2, 3, 4. International Journal of Quantum Chemistry, 26(2):223–230, August 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Sapse:1984:ISNb

[SOS84b] Anne-Marie Sapse, Louis Osorio, and Grace Snyder. Ab initio studies of negative ion-molecule(s) clusters present in the atmosphere. III. OH^{−}(CO_{2})_{n} for n = 1, 2. International Journal of Quantum Chemistry, 26(2):231–236, August 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Saran:1981:MOS

[SP81] Anil Saran and Lalit N. Patnaik. Molecular orbital studies on nucleoside antibiotics. IV. conformation of 3’-deoxyadenosine (cordycepin) and 3’-amino-3’-deoxyadenosine. International Journal of Quantum Chemistry, 20(2):357–367, August 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Saran:1982:MOS

[SP82a] Anil Saran and Lalit N. Patnaik. Molecular orbital studies on nucleoside antibiotics. VI. Conformation of 5-vinyl- and 5-ethynyl-2′-deoxyuridines. International Journal of Quantum Chemistry, 22(S9):247–257, March 4–6, 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Stolarczyk:1982:DCL

[SP82b] Leszek Z. Stolarczyk and Lucjan Piela. Direct calculation of lattice sums. A method to account for the crystal field effects.

International Journal of Quantum Chemistry, 22(5):911–927, November 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Sordo:1985:QTH

[SP85] J. A. Sordo and L. Pueyo. A quality test for the Hartree– Fock–Roothaan SCF wave functions. International Journal of Quantum Chemistry, 28(6):687–690, December 1985. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Sannigrahi:1986:ISC

[SP86a] A. B. Sannigrahi and Sigrid D. Peyerimhoff. Ab initio SCF and CI study of the electronic structure of the trichlorine radical. International Journal of Quantum Chemistry, 30(3):413–420, September 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Saran:1986:MOS

[SP86b] Anil Saran and Lalit N. Patnaik. Molecular orbital studies on nucleoside antibiotics VIII. Conformation of 5-substituted uridines. International Journal of Quantum Chemistry, 30 (S13):121–131, March 6–8, 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Sarma:1989:SAA

[SPÖ89] C. R. Sarma, J. Paldus, and Y. Öhrn. Spin adaptation of antisymmetrized geminal product wave functions. International Journal of Quantum Chemistry, 36(1):35–48, July 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Sandoval:1989:OAR

[SPRS89] L. Sandoval, A. Palma, and F. Rivas-Silva. Operator algebra and recurrence relations for the Franck–Condon factors. International Journal of Quantum Chemistry, 36(S23):183–186, April 1–8, 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Srivastava:1985:QCM

[SPSG85] Sudha Srivastava, Ratna S. Phadke, Anil Saran, and Gerjesh Govil. Quantum chemical and magnetic resonance studies on interaction of α-tocopherol with phospholipids. International Journal of Quantum Chemistry, 28(S12):169–181, March 14– 16, 1985. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Szczesniak:1980:NSI

[SR80] Małgorzata M. Szcześniak and Henryk Ratajczak. Nonadditivity of the SCF interaction energy in the (LiH)_{3} complex. International Journal of Quantum Chemistry, 17(6):1069–1074, June 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Sanyal:1981:SSI

[SR81] Nitish K. Sanyal and Mihir Roychoudhury. Studies in some important aspects of RNA transcription on the basis of intermolecular energy calculations. International Journal of Quantum Chemistry, 20(1):63–69, July 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Sundaralingam:1983:WDP

[SR83] M. Sundaralingam and S. T. Rao. Why is DNA polymorphic while RNA is not? International Journal of Quantum Chemistry, 24(S10):301–308, March 14–16, 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Salahub:1984:ECA

[SR84a] D. R. Salahub and F. Raatz. Effects of chemisorption and alloying on the magnetism of nickel clusters. International Journal of Quantum Chemistry, 26(S18):173–182, March 1– 15, 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Atomic, Molecular and Solid-State Theory, and Computational Quantum Chemistry.

Sapse:1984:ICG

[SR84b] A. M. Sapse and C. S. Russell. Ab initio calculations of guanidinium–carboxylate interaction. International Journal of Quantum Chemistry, 26(1):91–99, July 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Sokalski:1987:BSE

[SR87] W. A. Sokalski and S. Roszak. Basis set extension effects on the He_{2} interaction energy components. International Journal of Quantum Chemistry, 32(3):279–293, September 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Stafstrom:1983:ISI

[SRC83] S. Stafström, R. Riklund, and K. A. Chao. Interacting stereoirregular chains: a model for conducting polymers. International Journal of Quantum Chemistry, 23(4):1363–1383, April 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Sundaralingam:1984:NHR

[SRC84] Muttaiya Sundaralingam, John R. Rubin, and John F. Cannon. Nonenzymatic hydrolysis of RNA: Pb(II)-catalyzed site specific hydrolysis of transfer RNA. The role of the tertiary folding of the polynucleotide chain. International Journal of Quantum Chemistry, 26(S11):355–366, March 12–15, 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Sundaralingam:1985:TDS

[SRDG85] M. Sundaralingam, S. T. Rao, W. Drendel, and M. L. Greaser. The three-dimensional structure of troponin-C: Supersecondary structural homology of calcium binding folds. International Journal of Quantum Chemistry, 28(S12):153– 160, March 14–16, 1985. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Scheiner:1986:FIP

[SRH86] Steve Scheiner, Paul Redfern, and Eric A. Hillenbrand. Factors influencing proton positions in biomolecules. International Journal of Quantum Chemistry, 29(4):817–827, April 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Sokalski:1983:ISI

[SRHK83] W. A. Sokalski, S. Roszak, P. C. Hariharan, and Joyce J. Kaufman. Improved SCF interaction energy decomposition scheme corrected for basis set superposition effect. International Journal of Quantum Chemistry, 23(3):847–854, March 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Srivastava:1986:CDT

[Sri86] Vipin Srivastava. The case of disordered two-dimensional electron systems. International Journal of Quantum Chemistry, 29(5):1525–1533, May 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Sokalski:1983:CSS

[SRL^{+}83] W. Andrzej Sokalski, Szczepan Roszak, Alfred H. Lowrey, P. C. Hariharan, Walter S. Koski, Joyce J. Kaufman, and Richard S. Miller. Crystal structure studies using ab-initio potential functions from partitioned ab-initio MODPOT/ VRDDO SCF energy calculations. I. N_{2} and CO_{2} test cases. II. Nitromethane, CH_{3}NO_{2}. International Journal of Quantum Chemistry, 24(S17):375–391, March 3–12, 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-state Theory, Collision Phenomena, and Computational Quantum Chemistry.

Sanyal:1981:BAN

[SRO81] Nitish K. Sanyal, Mihir Roychoudhury, and Rajendra P. Ojha. Biological activity of the nucleoside analogs: a theoretical study. International Journal of Quantum Chemistry, 20(1): 159–166, July 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Scheire:1981:MSX

[SRP81] L. Scheire, R. Rotthier, and P. Phariseau. MSW–SCF–Xα studies of some silicon compounds. I. Iminosilicon (HNSi). International Journal of Quantum Chemistry, 20(4):879–885, October 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Sapse:1981:SCF

[SRS81] A. M. Sapse, C. S. Russell, and G. Snyder. Self-consistent field calculations on 2-aminomethyl-3-carboxymethyl pyrrole, a substrate analog of PBG deaminase evidence of intramolecular hydrogen bonding. International Journal of Quantum Chemistry, 20(S8):73–80, March 5–7, 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Srinivasan:1985:MMM

[SRSR85] S. Srinivasan, G. Raghunathan, M. Shibata, and R. Rein. Multistep modeling (MSM) of biomolecular structure application to the A-G mispair in the B-DNA environment. International Journal of Quantum Chemistry, 28(S12):217–227, March 14– 16, 1985. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Simonetta:1981:NPI

[SRT81] Massimo Simonetta, Mario Raimondi, and Gian Franco Tantardini. New problems for ab-initio valence bond theory. International Journal of Quantum Chemistry, 20(S15):225–229, March 8–14, 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory, Collision Phenomena, and Computational Quantum Chemistry.

Scharfenberg:1980:BRF

[SS80a] Peter Scharfenberg and J. Sauer. Biological response as a function of conformation, chirality, and electronic characteristics: a catecholamine study. International Journal of Quantum Chemistry, 18(5):1309–1337, November 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Sears:1980:PCC

[SS80b] V. F. Sears and E. C. Svensson. Pair correlations and the condensate fraction in superfluid ^{4}He. International Journal of Quantum Chemistry, 18(S14):715–721, March 10–20, 1980.

CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory, Collision Phenomena, Quantum Statistics, and Computational Methods.

Shepard:1980:ESS

[SS80c] Ron Shepard and Jack Simons. The electronic structure of singlet cyclopentadienylidene. International Journal of Quantum Chemistry, 18(S14):349–353, March 10–20, 1980. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory, Collision Phenomena, Quantum Statistics, and Computational Methods.

Shepard:1980:MWO

[SS80d] Ron Shepard and Jack Simons. Multiconfigurational wavefunction optimization using the unitary group method. International Journal of Quantum Chemistry, 18(S14):211–228, March 10–20, 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory, Collision Phenomena, Quantum Statistics, and Computational Methods.

Shinagawa:1980:RAE

[SS80e] Yasuko Shinagawa and Yoshiya Shinagawa. Regression analysis of enzyme reaction. O-methylation of catechol derivatives by catechol-O-methyltransferase. International Journal of Quantum Chemistry, 18(2):521–526, August 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Spingarn:1980:MBI

[SS80f] Neil E. Spingarn and Alan C. Sartorelli. Mechanism of binding of iron to potential therapeutic chelating agents. International Journal of Quantum Chemistry, 18(2):493–500, August 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Sen:1981:SDQ

[SS81a] K. D. Sen and P. C. Schmidt. Static dipole and quadrupole polarizabilities of alkaline earth atoms. International Journal of Quantum Chemistry, 19(2):373–376, February 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Shugar:1981:TPP

[SS81b] David Shugar and Krystyna Szcześniak. Tautomerism of pyrimidines and purines in the gas phase and in lowtemperature matrices, and some biological implications. International Journal of Quantum Chemistry, 20(2):573–585, August 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Sapse:1982:ISR

[SS82] Anne Marie Sapse and Grace Snyder. An ab initio study of the reaction pathway between a mitomycin fragment model and ammonia. International Journal of Quantum Chemistry, 22(S9):297–305, March 4–6, 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Sapse:1983:TCP

[SS83] Anne-Marie Sapse and Grace Snyder. Theoretical calculations of properties of some nitrosoureas. International Journal of Quantum Chemistry, 24(S10):175–179, March 14–16, 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Sarai:1984:TSI

[SS84a] Akinori Sarai and Minoru Saito. Theoretical studies on the interaction of proteins with base pairs. I. Ab initio calculation for the effect of H-bonding interaction of proteins on the stability of adenine–uracil pair. International Journal of Quantum Chemistry, 25(3):527–533, March 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Scheiner:1984:IIA

[SS84b] Steve Scheiner and M. M. Szcześniak. Interactions involving aromatic systems: Benzene + acetylene. International Journal of Quantum Chemistry, 26(S11):201–208, March 12–15, 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Sarai:1985:TSI

[SS85] Akinori Sarai and Minoru Saito. Theoretical studies on the interaction of proteins with base pairs. II. Effect of external H-bond interactions on the stability of guanine–cytosine and non-Watson–Crick pairs. International Journal of Quantum Chemistry, 28(3):399–409, September 1985. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Shinagawa:1986:MEC

[SS86a] Y. Shinagawa and Yasuko Shinagawa. The minimal energy conformations of o-benzosemiquinone anion radicals in the solvents containing the alkali metals and the alkali earth metals. International Journal of Quantum Chemistry, 29(3):351–360, March 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Shinagawa:1986:CBR

[SS86b] Yasuko Shinagawa and Y. Shinagawa. The conformation of obenzosemiquinone radical and its assignment of the proton hyperfine coupling constants by using the INDO and the molecular geometry adjusting methods. International Journal of Quantum Chemistry, 29(3):361–371, March 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Sosa:1986:ICBa

[SS86c] Carlos Sosa and H. Bernhard Schlegel. Ab initio calculations on the barrier height for the hydrogen addition to ethylene and formaldehyde. The importance of spin projection. International Journal of Quantum Chemistry, 29(4):1001–1015, April 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Sosa:1986:ICBb

[SS86d] Carlos Sosa and H. Bernhard Schlegel. Ab initio calculations on the barrier height for the hydrogen addition to ethylene and formaldehyde. The importance of spin projection. International Journal of Quantum Chemistry, 30(1):155–156, July 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Sosa:1987:ICH

[SS87] Carlos Sosa and H. Bernhard Schlegel. Ab initio calculations on H + C_{2}H_{2} → C_{2}H_{3} using unrestricted Møller–Plesset perturbation theory with spin projection. International Journal of Quantum Chemistry, 32(S21):267–282, March 12, 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Chemistry, Solid-State Theory, and Computational Methods.

Szpunar:1988:IMH

[SS88] Barbara Szpunar and Vedene H. Smith, Jr. Itinerant magnetism in high T_{c} superconductors: YBa_{2}Cu_{3} oxides and related structures. International Journal of Quantum Chemistry, 34(S22):33–42, March 12–19, 1988. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Syposium on Quantum Chemistry, Solid-State Theory, and Computational Methods.

Shivaglal:1989:OSF

[SS89a] M. C. Shivaglal and Surjit Singh. OH stretching force constant in associated methanol species and the cooperativity effect. International Journal of Quantum Chemistry, 36(2):105–118, August 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Sjogren:1989:SAM

[SS89b] Lennart Sjögren and Alf Sjölander. Some applications of the mode–mode coupling approach. International Journal of Quantum Chemistry, 35(6):851–867, June 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Scheiner:1983:ISP

[SSB83] Steve Scheiner, Małgorzata M. Szczelśniak, and Larry D. Bigham. Ab initio study of proton transfers including effects of electron correlation. International Journal of Quantum Chemistry, 23(2):739–751, February 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Sokalski:1983:CAF

[SSC83] W. A. Sokalski, A. Sawaryn, and H. Chojnacki. Catalytic activity of the first solvation shell in the CO_{2} hydration reaction. International Journal of Quantum Chemistry, 24(S10):

321–329, March 14–16, 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Stoll:1984:PCI

[SSF^{+}84] Hermann Stoll, Lászlö V. Szentpály, Patricio Fuentealba, Jürgen Flad, Michael Dolg, Franz xaver Fraschio, Peter Schwerdtfeger, Gudrun Igel, and Heinzwerner Preuss. Pseudopotential calculations including core-valence correlation: Alkali and noble-metal compounds. International Journal of Quantum Chemistry, 26(5):725–727, November 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Simas:1984:NSM

[SSK84] Alfredo M. Simas, Vedene H. Smith, Jr., and Per Kaijser. The nodal structure of the momentum distributions of molecules. International Journal of Quantum Chemistry, 25(6):1035– 1044, June 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Szczesniak:1984:VSU

[SSN^{+}84] K. Szcześniak, M. Szczesniak, M. Nowak, I. Scott, S. Chin, and W. B. Person. Vibrational spectra of uracils: Experimental infrared matrix studies and ab initio quantummechanical calculations. International Journal of Quantum Chemistry, 26(S18):547–569, March 1–15, 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Atomic, Molecular and Solid-State Theory, and Computational Quantum Chemistry.

Shepard:1988:PRS

[SSP^{+}88] Ron Shepard, Isaiah Shavitt, Russell M. Pitzer, Donald C. Comeau, Melanie Pepper, Hans Lischka, Peter G. Szalay, Reinhart Ahlrichs, Franklin B. Brown, and Jian-Guo Zhao. A progress report on the status of the COLUMBUS MRCI program system. International Journal of Quantum Chemistry, 34(S22):149–165, March 12–19, 1988. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Syposium on Quantum Chemistry, Solid-State Theory, and Computational Methods.

Srinivasan:1986:MMP

[SSR86] Subhashini Srinivasan, Masayuki Shibata, and Robert Rein. Multistep modeling of protein structure: Application to bungarotoxin. International Journal of Quantum Chemistry, 30 (S13):167–174, March 6–8, 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Scuseria:1987:AEE

[SSR^{+}87] Gustavo E. Scuseria, Andrew C. Scheiner, Julia E. Rice, Timothy J. Lee, and Henry F. Schaefer III. Analytic evaluation of energy gradients for the single and double excitation coupled cluster (CCSD) wave function: A comparison with configuration interaction (CCSD, CISDT, and CISDTQ) results for the harmonic vibrational frequencies, infrared intensities, dipole moment, and inversion barrier of ammonia. International Journal of Quantum Chemistry, 32(S21):495–501, March 12, 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Chemistry, Solid-State Theory, and Computational Methods.

Srinivasan:1987:MMP

[SSRR87] Subhashini Srinivasan, Masayuki Shibata, Mihir Roychoudhury, and Robert Rein. Multistep modeling of protein structure: Application towards refinement of tyr-tRNA synthetase. International Journal of Quantum Chemistry, 32(S14):281– 288, March 14, 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Sapse:1981:ISS

[SSS81] Anne Marie Sapse, Grace Snyder, and Angelo V. Santoro. Ab initio SCF study of guanidine and substituted guanidines. International Journal of Quantum Chemistry, 20(3):755–762, September 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Sandorfy:1986:P

[SSS^{+}86] C. Sandorfy, W. Siebrand, V. H. Smith, Jr., J. Cizek, J. Paldus, and D. Salahub. Preface. International Journal of Quantum Chemistry, 29(2):v, February 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Sapse:1982:SCF

[SSSH82] A. M. Sapse, G. Snyder, A. V. Santoro, and L. Herzig. Self consistent field MO studies of the vinylguanidinium ion. International Journal of Quantum Chemistry, 22(3):649–655, September 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Staszewska:1983:OME

[SST83] Grazyna Staszewska, David W. Schwenke, and Donald G. Truhlar. Optical model for electron scattering by Ar at 30– 3000eV: Test of the adiabatic model for charge polarization and a quasi-free scattering model for inelastic effects. International Journal of Quantum Chemistry, 24(S17):163–176, March 3–12, 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-state Theory, Collision Phenomena, and Computational Quantum Chemistry.

Simas:1984:PWA

[SST84] Alfredo M. Simas, Vedene H. Smith, Jr., and Ajit J. Thakkar. Partial-wave analysis of the momentum densities of 14 electron diatomics. International Journal of Quantum Chemistry, 26(S18):385–392, March 1–15, 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Atomic, Molecular and Solid-State Theory, and Computational Quantum Chemistry.

Shinagawa:1980:ISS

[SSUF80] Yoshiya Shinagawa, Yasuko Shinagawa, Nobuhiro Uyesaka, and Kohji Fukuda. INDO Studies on the structure of benzosemiquinone radicals. International Journal of Quantum Chemistry, 18(2):357–363, August 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Sundaralingam:1987:SAH

[SSYR87] M. Sundaralingam, Y. C. Sekharudu, N. Yathindra, and V. Ravichandran. Stabilization of alpha helices by ion pairs. International Journal of Quantum Chemistry, 32(S14):289– 296, March 14, 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Sun:1982:SAS

[SSZJ82] Chia-Chung Sun, Wei-Guo Sun, Jing-Chang Zhang, and Guang-Zhu Jin. Symmetry analysis of sigmatropic rearrangements. International Journal of Quantum Chemistry, 22(6): 1209–1220, December 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Satoh:1980:ECG

[ST80a] Mitsuo Satoh and Toshihiko Taki. Electric center of gravity of overlap charge distribution in impurity-doped ionic crystals. International Journal of Quantum Chemistry, 18(S14):579– 583, March 10–20, 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory, Collision Phenomena, Quantum Statistics, and Computational Methods.

Satoh:1980:LRE

[ST80b] Mitsuo Satoh and Toshihiko Taki. Lattice relaxation and electric-field-gradient calculations in impurity-doped single ionic crystals. International Journal of Quantum Chemistry, 18(2):633–640, August 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Skancke:1982:ICE

[ST82] Anne Skancke and Colin Thomson. Ab initio calculations of the electronic structures of Dimethylnitrosamine (DMN) and its Metabolite Hydroxydimethylnitrosamine (HDMN): a possible intermediate in carcinogenesis by DMN. International Journal of Quantum Chemistry, 21(2):431–443, February 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Seibert:1983:CGT

[ST83] J. Seibert and N. Trinajstića. Chemical graph theory. IV. On the cyclic polynomial. International Journal of Quantum Chemistry, 23(5):1829–1841, May 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Safronova:1984:RCA

[ST84] U. I. Safronova and T. A. Tsirekidze. Relativistic calculation of autoionization widths of 1s2s2p_{j}J, 1s2s^{2}^{2}S_{1/2}, and 1s2p_{J}^{2}p′_{j}J levels. International Journal of Quantum Chemistry, 25(1):63–68, January 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Sharma:1986:WDK

[ST86a] B. S. Sharma and Ajit J. Thakkar. What do kinetic-energy anisotropies tell us about chemical bonding? I. Diatomic hydrides. International Journal of Quantum Chemistry, 29(3): 323–332, March 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Singh:1986:ETE

[ST86b] Jai Singh and A. Thilagam. Exciton trapping by emitting two phonons at impurities in molecular crystals. International Journal of Quantum Chemistry, 30(1):19–25, July 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Sutcliffe:1986:CFM

[ST86c] B. T. Sutcliffe and J. Tennyson. The construction and fitting of molecular potential energy surfaces and their use in vibration-rotation calculations. International Journal of Quantum Chemistry, 30(S20):507–520, March 10–15, 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Syposium on Atomic, Molecular, and Solid-State Theory, Scattering Problems, Many Body Phenomena, and Computational Quantum Chemistry.

Stamato:1988:ISC

[ST88] Fulvia M. L. G. Stamato and O. Tapia. Ab initio studies on the catalytic mechanism of ester hydroloysis by serine proteases. International Journal of Quantum Chemistry, 33(3):187–194, March 1988. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Sabio:1989:CCS

[ST89] Michael Sabio and Sid Topiol. Computational chemical studies of chiral stationary-phase models: The nature of the Pi interaction in complexes of methyl N-(2-naphthyl) alaninate with N-(3,5-dinitrobenzoyl)leucine n-propylamide. International Journal of Quantum Chemistry, 36(3):313–317, September 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Stanley:1971:IPT

[Sta71] H. Eugene (Harry Eugene) Stanley. Introduction to phase transitions and critical phenomena. The International series of monographs on physics. Clarendon Press, New York, NY, USA, 1971. ISBN 0-19-851257-0. xx + 308 + 3 pp. LCCN QC307 .S7 1971.

Stanley:1987:IPT

[Sta87] H. Eugene (Harry Eugene) Stanley. Introduction to phase transitions and critical phenomena. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 1987. ISBN 0-19-505316-8 (paperback). xx + 308 pp. LCCN QC175.16.P5 S72

1987. URL http://www.loc.gov/catdir/enhancements/
fy0603/87012357-d.htm;   http://www.loc.gov/catdir/
enhancements/fy0603/87012357-t.htm.

Stepanov:1981:UTU

[Ste81] N. F. Stepanov. Unitary transformations with the unitary operators depending on projection operators. International Journal of Quantum Chemistry, 19(5):793–803, May 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Stein:1984:MOB

[Ste84] D. L. Stein. A model for the origin of biological information. International Journal of Quantum Chemistry, 26(S11): 73–86, March 12–15, 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Steckler:1986:DCK

[STG86] Rozeanne Steckler, Donald G. Truhlar, and Bruce C. Garrett. Dynamics calculations of kinetic isotope effects for the reactions of muonium atoms with F_{2} and Cl_{2}. International Journal of Quantum Chemistry, 30(S20):495–506, March 10–15, 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Syposium on Atomic, Molecular, and Solid-State Theory, Scattering Problems, Many Body Phenomena, and Computational Quantum Chemistry.

Sakai:1980:CCG

[STH80] Yoshiko Sakai, Hiroshi Tatewaki, and Sigeru Huzinaga. Compact contracted Gaussian-type basis sets from Na to Ar: Application to molecular calculations. International Journal of Quantum Chemistry, 18(S14):107–112, March 10–20, 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory, Collision Phenomena, Quantum Statistics, and Computational Methods.

Stillinger:1982:SIB

[Sti82] F. H. Stillinger. Structure of the interface between coexisting fluid phases. International Journal of Quantum Chemistry, 22(S16):137–147, March 1–13, 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Chemistry, Theory of Condensed Matter, and Propagator Methods in the Quantum Theory of Matter.

Story:1980:GAC

[Sto80] Troy L. Story. Global analysis of composite particles. International Journal of Quantum Chemistry, 17(5):915–930, May 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Stolarczyk:1982:MTBa

[Sto82a] Leszek Z. Stolarczyk. Molecular tight-binding method. I. Many-electron method for hole bands of molecular crystals. International Journal of Quantum Chemistry, 21(6):967–991, June 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Stolarczyk:1982:MTBb

[Sto82b] Leszek Z. Stolarczyk. Molecular tight-binding method. II. One- and many-electron method for excess electron bands of molecular crystals. International Journal of Quantum Chemistry, 21(6):993–1001, June 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Stranges:1982:EVD

[Str82] Anthony N. (Anthony Nicholas) Stranges. Electrons and valence: development of the theory, 1900–1925. Texas A&M University Press, College Station, 1982. ISBN 0-89096-124-7. xii + 291 + 2 pp. LCCN QD469 .S73 1982. US$28.50.

URL http://www.loc.gov/catdir/enhancements/fy1003/
81048378-b.htm.

Simas:1982:MSP

[STS82] Alfredo M. Simas, Art J. Thakkar, and Vedene H. Smith, Jr. Momentum space properties of various orbital basis sets used in quantum chemical calculations. International Journal of Quantum Chemistry, 21(2):419–429, February 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Simas:1983:BSQ

[STS83] Alfredo M. Simas, Ajit J. Thakkar, and Vedene H. Smith, Jr. Basis set quality. II. Information theoretic appraisal of various s-orbitals. International Journal of Quantum Chemistry, 24 (6):527–550, December 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Smet:1980:EHF

[STV80] F. Smet, J. Tillieu, and Et. A. Van Groenendael. Equations de Hartree–Fock dependant du temps. International Journal of Quantum Chemistry, 17(3):531–547, March 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Suck:1981:KDT

[Suc81] Sung Ho Suck. The kernel of DWBA transition amplitude in atom–diatom reactive scattering. International Journal of Quantum Chemistry, 19(3):441–450, March 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Sucher:1984:FRT

[Suc84] J. Sucher. Foundations of the relativistic theory of manyelectron bound states. International Journal of Quantum Chemistry, 25(1):3–21, January 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Suhai:1983:PTI

[Suh83] Sándor Suhai. Perturbation theoretical investigation of electron correlation effects in infinite metallic and semiconducting polymers. International Journal of Quantum Chemistry, 23 (4):1239–1256, April 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Suhai:1984:ECE

[Suh84a] S. Suhai. Electron correlation effects on the mechanical and optical properties of polymers. International Journal of Quantum Chemistry, 26(S18):161–171, March 1–15, 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Atomic, Molecular and Solid-State Theory, and Computational Quantum Chemistry.

Suhai:1984:FPC

[Suh84b] S. Suhai. First principles charge transfer exciton theory of the UV spectrum of DNA. International Journal of Quantum Chemistry, 26(S11):223–235, March 12–15, 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Suhai:1986:ENF

[Suh86] S. Suhai. On the excitonic nature of the first UV absorption peak in polyene. International Journal of Quantum Chemistry, 29(3):469–476, March 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Smet:1982:EMH

[SV82] F. Smet and A. Van Groenendael. Équation matricielle de Hartree–Fock permettant de déterminer la nième harmonique. (French) [Hartree–Fock matrix equation allowing determination of the n-th harmonic]. International Journal of Quantum Chemistry, 21(6):1061–1078, June 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Smet:1985:TQEa

[SV85a] F. Smet and A. Van Groenendael. Sur la théorie quantique des équations de Hartree–Fock dépendant du temps. III. Détermination d’un ensemble unique de spin orbitales. (French) [On the quantum theory of time-dependent Hartree– Fock equations. III. determination of a unique set of spin orbitals]. International Journal of Quantum Chemistry, 27(2): 159–166, February 1985. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Smet:1985:TQEb

[SV85b] F. Smet and A. Van Groenendael. Sur la théorie quantique des équations de Hartree–Fock dépendant du temps. IV. Expression matricielle du tenseur d’hyperpolarisabilité (French) [On the quantum theory of time-dependent Hartree–Fock equations. IV. Matrix expression for the hyperpolarizabilty tensor]. International Journal of Quantum Chemistry, 27(2):167–171, February 1985. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Svendsen:1988:VMC

[Sve88] E. Norby Svendsen. A variational method for calculation of dynamic third-order susceptibilities. International Journal of Quantum Chemistry, 34(S22):477–480, March 12–19, 1988. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Syposium on Quantum Chemistry, Solid-State Theory, and Computational Methods.

Smet:1989:TEH

[SvG89] F. Smet and A. van Groenendael. Sur la théorie des équations de Hartree–Fock dépendent du temps. V. Étude d’un modèle sphérique à 2 électrons. Calcul du tenseur d’hyperpolarisabilite d’ordre 3 (French) [On the quantum theory of time-dependent Hartree–Fock equations. V. Study of a 2-electron spherical model]. International Journal of Quantum Chemistry, 36(1):67–91, July 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Seybold:1981:RBC

[SVS81] Paul G. Seybold, Richard Vestewig, and John D. Scribner. Relationships between carcinogenicity, mutagenicity, and theoretical reactivity indices for polycyclic aromatic hydrocarbons. International Journal of Quantum Chemistry, 20(S8): 401–408, March 5–7, 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Sundaralingam:1981:SMN

[SW81] M. Sundaralingam and E. Westhof. Structural motifs of the nucleotidyl unit and the handedness of polynucleotide helices. International Journal of Quantum Chemistry, 20(S8):

287–306, March 5–7, 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Spark:1982:CAF

[SWA82] M. J. Spark, D. A. Winkler, and P. R. Andrews. Conformational analysis of folates and folate analogues. International Journal of Quantum Chemistry, 22(S9):321–333, March 4–6, 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Swenson:1982:CHR

[Swe82] Robert J. Swenson. Comment on hypervirial relations for bounded systems. International Journal of Quantum Chemistry, 21(5):937–938, May 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Swenson:1983:LEC

[Swe83] Robert J. Swenson. Letter to the Editor: Constraints on boundary conditions for enclosed systems due to real energy eigenvalues. International Journal of Quantum Chemistry, 24 (3):333–334, September 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Snyder:1982:SBD

[SWM82] Lawrence C. Snyder, Zelda R. Wasserman, and Jules W. Moskowitz. Stability and bonding of disilyne and its isomers: a generalized valence bond-effective potential study. International Journal of Quantum Chemistry, 21(3):565–579, March 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Scheiner:1989:PPT

[SWW89] Steve Scheiner, Rongshun Wang, and Lan Wang. Perturbations of proton transfer potentials caused by polar molecules. International Journal of Quantum Chemistry, 36(S16):211– 217, April 1–8, 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Sawaryn:1982:ISN

[SY82] A. Sawaryn and J. S. Yadav. Ab initio studies on the nonplanarity of a peptide unit: Calculations on model compounds. International Journal of Quantum Chemistry, 22(3):547–556, September 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Singh:1986:STI

[SY86] O. P. Singh and J. S. Yadav. Similarity transformation of the ab initio density matrix. International Journal of Quantum Chemistry, 29(5):1283–1290, May 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Sauer:1984:QCS

[SZ84] Joachim Sauer and Rudolf Zahradník. Quantum chemical studies on zeolites and silica. International Journal of Quantum Chemistry, 26(5):793–822, November 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Szasz:1985:PTA

[Sza85] Levente Szasz. Pseudopotential theory of atoms and molecules. John Wiley, New York, NY, USA, 1985. ISBN 0-471-82417-8. xi + 309 pp. LCCN QD461 .S988196 1985. US$42.50 (est.).

Sun:1982:SAC

[SZJS82] Chia-Chung Sun, Jing-Chang Zhang, Guang-Zhu Jin, and Wei-Guo Sun. Symmetry analysis of cycloadditions and electrocyclic reactions. International Journal of Quantum Chemistry, 22(4):665–679, October 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Shibata:1980:NMV

[SZR80] Masayuki Shibata, Theresa Julia Zielinski, and Robert Rein. Normal-mode vibrational frequencies and thermodynamic properties of uracil and 4-hydroxyuracil. International Journal of Quantum Chemistry, 18(1):323–329, July 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Tabor:1982:ASI

[Tab82] Michael Tabor. Analytic structure and integrability of dynamical systems. International Journal of Quantum Chemistry, 22(S16):167–181, March 1–13, 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Chemistry, Theory of Condensed Matter, and Propagator Methods in the Quantum Theory of Matter.

Tachibana:1982:SQS

[Tac82] Akitomo Tachibana. “stable” quasistationary state: New solution of the time-dependent Schrödinger equation. International Journal of Quantum Chemistry, 22(1):191–197, July 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Tachibana:1983:CSQ

[Tac83] Akitomo Tachibana. Concept of the “stability” of the quantum mechanical state. International Journal of Quantum Chemistry, 23(1):195–215, January 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Tachibana:1987:DFR

[Tac87] Akitomo Tachibana. Density functional rationale of chemical reaction coordinate. International Journal of Quantum Chemistry, 32(S21):181–190, March 12, 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Chemistry, Solid-State Theory, and Computational Methods.

Tachibana:1988:SWD

[Tac88] Akitomo Tachibana. Shape wave in density functional theory. International Journal of Quantum Chemistry, 34(4):309–323, October 1988. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Tachibana:1989:AME

[Tac89] Akitomo Tachibana. Application of the Mermin entropy principle to the “apparatus” density functional theory. International Journal of Quantum Chemistry, 35(3):361–372, March 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Taijin:1989:SPG

[Tai89] Zhou Taijin. Simplification of point group projection operators and its application to symmetry adaptation of multishell electron configurations. International Journal of Quantum Chemistry, 35(5):593–612, May 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Takahata:1986:HSS

[Tak86] Y. Takahata. A HAM/3 study of substituted benzenes. International Journal of Quantum Chemistry, 30(4):453–465, October 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Tapia:1986:BRB

[Tap86] Orlando Tapia. Book review: Molecular Quantum Electrodynamics: An Introduction to Radiation-Molecular Interactions. By D. P. Craig and T. Thirunamachandran. Academic Press, London, 1984. International Journal of Quantum Chemistry, 29(6):1843–1844, June 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). See [CT84, CT98].

Tapia:1988:BRB

[Tap88] O. Tapia. Book review: Theoretical chemistry of biological systems. Edited by G. Náray-Szabó, Elsevier, Amsterdam, 1986. International Journal of Quantum Chemistry, 33(6):579, June 1988. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). See [NS86].

Tatake:1981:SET

[Tat81a] V. G. Tatake. Significance of the electron trap parameters of glow curves of photosynthetic membranes. International Journal of Quantum Chemistry, 20(2):533–541, August 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Tatum:1981:CCK

[Tat81b] J. Patrick Tatum. Calculations concerning Kato’s equations. International Journal of Quantum Chemistry, 20(3):717–745, September 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Taylor:1981:VTG

[Tay81] William J. Taylor. Variational treatment of the ground state of the hydrogenic atom using the trial function φ = A exp(−αr^{ν}). International Journal of Quantum Chemistry, 19(3):483–484, March 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Taylor:1983:RSP

[Tay83] William J. Taylor. Rayleigh–schrödinger perturbation theory for many-electron systems with partitioning of configurations by orders of excitation. International Journal of Quantum Chemistry, 23(4):1653–1665, April 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Taylor:1984:RSP

[Tay84] William J. Taylor. Rayleigh–schrödinger perturbation theory for many-electron systems with partitioning of configurations by orders of excitation. International Journal of Quantum Chemistry, 25(2):447, February 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Taylor:1985:SOM

[Tay85] Peter R. Taylor. Symmetrization of operator matrix elements. International Journal of Quantum Chemistry, 27(1):89–96, January 1985. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Taylor:1986:RSPa

[Tay86a] William J. Taylor. Rayleigh–schrödinger perturbation theory for many-electron systems with partitioning of configurations by orders of excitation. II. Analysis of the ground state wavefunction and energy. International Journal of Quantum Chemistry, 29(5):1263–1275, May 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Taylor:1986:RSPb

[Tay86b] William J. Taylor. Rayleigh–schrödinger perturbation theory for many-electron systems with partitioning of configurations by orders of excitation. II. Analysis of the ground-state wavefunction and energy. International Journal of Quantum Chemistry, 30(4):573, October 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Taylor:1987:SOS

[Tay87a] Howard S. Taylor. Spectroscopic order out of spectroscopic chaos. International Journal of Quantum Chemistry, 31(5): 747–753, May 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Taylor:1987:IPB

[Tay87b] Peter R. Taylor. Integral processing in beyond-Hartree–Fock calculations. International Journal of Quantum Chemistry, 31 (3):521–534, March 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Thuraisingham:1981:ECC

[TB81] R. A. Thuraisingham and G. Bandarage. Explicitly correlated configuration interaction wave functions for the He atom using correlated Gaussians. International Journal of Quantum Chemistry, 20(4):781–783, October 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Thomson:1983:TIS

[TB83] Colin Thomson and Richard Brandt. Theoretical investigations of the structure of potential inhibitors of the enzyme glyoxalase-I. International Journal of Quantum Chemistry, 24 (S10):357–373, March 14–16, 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Tellinghuisen:1988:NCC

[Tel88] Joel Tellinghuisen. Numerical comparison of the Cooley and Kobeissi methods for solving the one-dimensional Schrödinger equation. International Journal of Quantum Chemistry, 34 (4):401–406, October 1988. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

TenHoor:1980:GEH

[Ten80a] Marten J. Ten Hoor. Gaussian expansions of hydrogenic functions involving few variational parameters. International Journal of Quantum Chemistry, 17(6):1099–1109, June 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

TenHoor:1980:NHG

[Ten80b] Marten J. Ten Hoor. Notes on hydrogenic Gaussian lobe functions. International Journal of Quantum Chemistry, 17 (4):631–642, April 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

TenHoor:1986:ATF

[Ten86a] Marten J. Ten Hoor. Atomic Thomas–Fermi theory and electron subshell filling. International Journal of Quantum Chemistry, 30(6):831–844, December 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

TenHoor:1986:OPB

[Ten86b] Marten J. Ten Hoor. On optimization procedures based upon the atomic Thomas–Fermi energy. International Journal of Quantum Chemistry, 29(6):1713–1734, June 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

TenHoor:1988:HTC

[Ten88a] Marten J. Ten Hoor. Hall transformation of exp(−r): Correction of the results. International Journal of Quantum Chemistry, 33(6):563–566, June 1988. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

TenHoor:1988:HFG

[Ten88b] Marten J. Ten Hoor. The Hartree–Fock ground state of atomic two-electron systems and the Wilson–Silverstone 1s orbital. International Journal of Quantum Chemistry, 33(5):467–475, May 1988. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Tewari:1987:TSC

[Tew87] Ravindra Tewari. Theoretical studies on conformational preferences of modified nucleic acid base N^{6}-(N-glycylcarbonyl) adenine. International Journal of Quantum Chemistry, 31 (4):611–623, April 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Tewari:1988:CPM

[Tew88a] Ravindra Tewari. Conformational preferences of modified nucleic acid bases N^{6}-(Δ^{2}-isopentenyl) adenine and 2-methylthio-N^{6}-(Δ^{2}-isopentenyl) adenine by the quantum chemical PCILO calculations. International Journal of Quantum Chemistry, 34(2):133–142, August 1988. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Tewari:1988:TSC

[Tew88b] Ravindra Tewari. Theoretical studies on conformational preferences of model modified nucleic acid base N^{6}-(N-alanylcarbonyl) adenine. International Journal of Quantum Chemistry, 34(1):37–46, July 1988. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Teitell:1980:MSF

[TF80] Murray F. Teitell and J. Lawrence Fox. MO Study of flavin– protein interactions in flavodoxin catalysis. International Journal of Quantum Chemistry, 18(2):449–456, August 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Teitell:1982:MSF

[TF82] Murray F. Teitell and J. Lawrence Fox. MO study of flavin reduction. International Journal of Quantum Chemistry, 22 (3):583–594, September 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Takahashi:1983:PBT

[TF83a] Mitsuo Takahashi and Hideo Fukutome. Projected BCS Tamm-Dancoff method for molecular electronic structures. International Journal of Quantum Chemistry, 24(6):603–621, December 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Tang:1983:ISH

[TF83b] Ting-Hua Tang and Xiao-Yuan Fu. Ab initio studies of hydrogen bond formation in methyl cyanide or methyl isocyanide and methanol systems. International Journal of Quantum Chemistry, 24(3):317–325, September 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Tatewaki:1980:CCG

[TH80] Hiroshi Tatewaki and Sigeru Huzinaga. Compact contracted Gaussian-type basis sets from Li to Ne. International Journal of Quantum Chemistry, 18(S14):93–106, March 10–20, 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory, Collision Phenomena, Quantum Statistics, and Computational Methods.

Tai:1981:EOE

[TH81a] Chin Chui Tai and Geoffrey Hunter. An exact one-electron model of atomic structure: The helium atom ground state. International Journal of Quantum Chemistry, 20(S15):385, March 8–14, 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory, Collision Phenomena, and Computational Quantum Chemistry.

Tsukada:1981:ESV

[TH81b] Masaru Tsukada and Toshiharu Hoshino. Electronic structure of vacancy and chemisorptive bond on Si(111) surface by the DV–Xα cluster calculation. International Journal of Quantum Chemistry, 20(S15):445–455, March 8–14, 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory, Collision Phenomena, and Computational Quantum Chemistry.

Thomson:1988:EPM

[TH88] Colin Thomson and Derek Higgins. Electrostatic potential maps of flavone acetic acids: A new class of antitumor drugs. International Journal of Quantum Chemistry, 34(S22):697– 705, March 12–19, 1988. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Syposium on Quantum Chemistry, Solid-State Theory, and Computational Methods.

Thakkar:1983:IFC

[Tha83] Ajit J. Thakkar. Interatomic forces and Compton profiles. International Journal of Quantum Chemistry, 23(1):227–234, January 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Theophanides:1984:MIB

[The84] T. Theophanides. Metal ions in biological system. International Journal of Quantum Chemistry, 26(5):933–941, November 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Tobin:1981:IMV

[THKM81] Frank L. Tobin, P. C. Hariharan, Joyce J. Kaufman, and Richard S. Miller. Ab initio MODPOT/VRDDO/MERGE calculations on energetic compounds. I. Mechanism of initiation of cationic polymerization from electrostatic molecular potential contour maps. International Journal of Quantum Chemistry, 20(S15):203–216, March 8–14, 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory, Collision Phenomena, and Computational Quantum Chemistry.

Thomson:1980:ICE

[Tho80] Colin Thomson. Ab-initio calculations of the equilibrium geometry of β-propiolactone (2-oxetanone) by the force method. International Journal of Quantum Chemistry, 18(S7):163– 169, March 5–8, 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Thomas:1983:CLQ

[Tho83] L. H. Thomas. Classical limit of quantum mechanics in the large. International Journal of Quantum Chemistry, 23(1): 99–113, January 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Thole:1985:LSN

[Tho85] B. T. Thole. Least-squares numerical Rayleigh–Ritz and minimum-variance methods for molecular calculations. International Journal of Quantum Chemistry, 28(4):535–551, October 1985. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Thomas:1986:MGV

[Tho86a] Gerald F. Thomas. On the minimization of the global variance in the 1-Reduced local-energy matrix. International Journal of Quantum Chemistry, 29(4):867–873, April 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Thomson:1986:SRD

[Tho86b] Colin Thomson. Some recent developments concerning the mechanisms of tumor promotion. International Journal of Quantum Chemistry, 30(S13):297–305, March 6–8, 1986. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Thomson:1989:TSS

[Tho89] Colin Thomson. Theoretical studies of some problems in chemical carcinogenesis. International Journal of Quantum Chemistry, 36(S16):219–234, April 1–8, 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Timoneda:1986:ADP

[Tim86] J. Juanós I. Timoneda. Alternative derivation of perturbation expansions in the projection operator formalism. International Journal of Quantum Chemistry, 30(3):421–431, September 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Tindimubona:1980:OCE

[Tin80] Alex R. Tindimubona. One-center expansion method with model potentials. I. Formulation and test calculations. International Journal of Quantum Chemistry, 18(6):1415–1429, December 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Tyutyulkov:1988:SPN

[TIS^{+}88] N. N. Tyutyulkov, C. I. Ivanov, I. Schopov, O. E. Polansky, and G. Olbrich. Structure and properties of nonclassical polymers. VII. Polymers with heteroatoms. International Journal of Quantum Chemistry, 34(4):361–373, October 1988. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Tachibana:1986:VIO

[TIYH86] Akitomo Tachibana, Takayuki Inoue, Tokio Yamabe, and Kenzi Hori. Vibronic interaction in one-dimensional polymer. International Journal of Quantum Chemistry, 30(5):575–579, November 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Tvaroska:1983:TSC

[TK83] Igor Tvaroška and Tibor Kožár. Theoretical studies on the conformation of saccharides. V. Hydration of the acetal segment in glycosides. International Journal of Quantum Chemistry, 23(2):765–778, February 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Tyutyulkov:1986:SPN

[TK86] N. N. Tyutyulkov and S. C. Karabunarliev. Structure and properties of nonclassical polymers. III. Magnetic characteristics at finite temperatures. International Journal of Quantum Chemistry, 29(5):1325–1337, May 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Trinajstic:1986:SSU

[TKR86] Nenad Trinajstić, Douglas J. Klein, and Milan Randić. On some solved and unsolved problems of chemical graph theory. International Journal of Quantum Chemistry, 30(S20):699– 742, March 10–15, 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Syposium on Atomic, Molecular, and Solid-State Theory, Scattering Problems, Many Body Phenomena, and Computational Quantum Chemistry.

Tanaka:1989:ESR

[TKY89] Kazuyoshi Tanaka, Masahiro Kobashi, and Tokio Yamabe. Electronic structures and ring currents in planar carbon microcluster models. International Journal of Quantum Chemistry, 36(S23):641–649, April 1–8, 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Trsic:1980:IHF

[TL80] M. Trsic and W. G. Laidlaw. Ab initio Hartree–Fock–Slater calculations of polysulfanes H_{2}S_{n} (n = 1−−4) and the ions HS and S. International Journal of Quantum Chemistry, 17 (5):969–974, May 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Trsic:1982:IHF

[TL82] M. Trsic and W. G. Laidlaw. Ab initio Hartree–Fock–Slater calculation of Tetrathiafulvalene (TTF) and the TTF^{1+} and TTF^{2+} ions. International Journal of Quantum Chemistry, 21 (3):557–563, March 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Trsic:1983:UHF

[TL83] M. Trsic and W. G. Laidlaw. The use of the Hartree–Fock– Slater eigenvalues to estimate excitation and ionization energies in sulfur–nitrogen compounds. International Journal of Quantum Chemistry, 24(S17):367–374, March 3–12, 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-state Theory, Collision Phenomena, and Computational Quantum Chemistry.

Tang:1986:SRP

[TL86] Au-Chin Tang and Qian-Shu Li. A structural rule of polyhedral boranes and heteroboranes. International Journal of Quantum Chemistry, 29(4):579–587, April 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Toro-Labbe:1987:ICM

[TLC87] A. Toro-Labbe and C. Cardenas. The intramolecular conversion of monothioformic acid: An ab initio study. International Journal of Quantum Chemistry, 32(6):685–697, December 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Tamassy-Lentei:1989:PIS

[TLDK89] I. Tamássy-Lentei and Á. Derecskei-Kovács. Pseudopotential investigation of some alkali metal molecules. International Journal of Quantum Chemistry, 36(3):277–285, September 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Temussi:1984:SAR

[TLT^{+}84] P. A. Temussi, F. Lelj, T. Tancredi, M. A. Castiglione Morelli, and A. Pastore. Soft agonist receptor interactions: Theoretical and experimental simulation of the active site of the receptor of sweet molecules. International Journal of Quantum Chemistry, 26(5):889–906, November 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Thakkar:1981:EED

[TM81] Ajit J. Thakkar and Nancy J. Moore. Electronic extracule densities. International Journal of Quantum Chemistry, 20(S15): 393–400, March 8–14, 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory, Collision Phenomena, and Computational Quantum Chemistry.

Thulstrup:1983:OSP

[TM83] Erik W. Thulstrup and Josef Michl. Optical spectroscopy on partially oriented samples. International Journal of Quantum Chemistry, 24(S17):471–482, March 3–12, 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-state Theory, Collision Phenomena, and Computational Quantum Chemistry.

Tossell:1984:SMO

[TMC84] J. A. Tossell, J. H. Moore, and M. A. Coplan. Studies of molecular orbital momentum distributions by (e, 2e) spectroscopy. International Journal of Quantum Chemistry, 26(S18):483– 495, March 1–15, 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Atomic, Molecular and Solid-State Theory, and Computational Quantum Chemistry.

Tossell:1986:SUO

[TMO86] J. A. Tossell, J. H. Moore, and J. K. Olthoff. Studies of unoccupied orbitals of BF_{3} and BCL_{3} by electron transmission spectroscopy and multiple scattering Xα calculations. International Journal of Quantum Chemistry, 29(5):1117–1126, May 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Tanaka:1980:MSP

[TMON80] H. Tanaka, T. Matsushita, Y. Osamura, and K. Nishimoto. MO Study on the photochemical isomerization of isoxazole. International Journal of Quantum Chemistry, 18(2):463–468, August 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Thuraisingham:1980:SMO

[TN80] R. A. Thuraisingham and S. H. M. Nilar. Semiempirical molecular orbital calculations on π electron systems. International Journal of Quantum Chemistry, 18(2):397–401, August 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Tang:1989:PRP

[TN89] J. Tang and J. R. Norris. Phase refinement from a partially known structure using linear prediction filtering. International Journal of Quantum Chemistry, 36(S16):45–55, April 1–8, 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Tennyson:1986:CSH

[TNB86] Jonathan Tennyson, C. J. Noble, and P. G. Burke. Continuum states of the hydrogen molecule with the R-Matrix method. International Journal of Quantum Chemistry, 29(5): 1033–1042, May 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Takeuchi:1986:ETS

[TNT^{+}86] Munetaka Takeuchi, Makoto Nagashima, Kiyoshi Tanaka, Shigehiro Konaka, and Masao Kimura. Experimental and theoretical studies of small-angle electron scattering by the water molecule. International Journal of Quantum Chemistry, 30 (6):821–830, December 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Tanaka:1983:ELF

[TNY83] Kazuyoshi Tanaka, Masataka Nagaoka, and Tokio Yamabe. Electron localization in a finite one-dimensional chain. International Journal of Quantum Chemistry, 23(3):1101–1109, March 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Tachibana:1986:DPT

[TNY86] Akitomo Tachibana, Masataka Nagaoka, and Tokio Yamabe. Dynamic perturbation theory of energy transfer in nonrigid molecular systems: Vibrational predissociation of I_{2}He van der Waals molecule. International Journal of Quantum Chemistry, 29(5):1457–1462, May 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Tsuda:1980:PES

[TOK80] Minoru Tsuda, Setsuko Oikawa, and Katsuhiko Kimura. Potential-energy surfaces in the lower excited states of benzene ⇀ ↽ Dewar–benzene isomerization process. International Journal of Quantum Chemistry, 18(1):157–164, July 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Toyoda:1980:AEC

[Toy80] Kaoru Toyoda. Analyses of the electronic collective motions by means of a nonlinear dynamical equation. International Journal of Quantum Chemistry, 18(6):1523–1531, December 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Toyoda:1988:SCG

[Toy88] Kaoru Toyoda. Structure of consistent ground state: Re— derivation of RPA vacuum and inclusion of higher RPA effects. International Journal of Quantum Chemistry, 34(6):489–503, December 1988. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Tomasek:1980:PSE

[TP80] Mojmír Tomášek and Štěpán Pick. Projected surface energy bands, Shockley surface states, and bonding in transitionmetal monocompounds with rock-salt structure: Example of NbC. International Journal of Quantum Chemistry, 17(6): 1167–1177, June 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Takahashi:1984:BLA

[TP84] M. Takahashi and J. Paldus. Bond length alternation in cyclic polyenes. V. Local finite-order perturbation theory approach. International Journal of Quantum Chemistry, 26(3):349–371, September 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Takahashi:1985:BLA

[TP85] M. Takahashi and J. Paldus. Bond length alternation in cyclic polyenes. VI. Coupled cluster approach with Wannier orbital basis. International Journal of Quantum Chemistry, 28(4): 459–479, October 1985. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Tomasek:1987:SQR

[TP87] Mojmír Tomášek and Štěpán Pick. Simple qualitative rules to study the surface electronic structure of some binary crystals. International Journal of Quantum Chemistry, 32(4):501–509, October 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Thakkar:1989:MEE

[TP89] Ajit J. Thakkar and Ward A. Pedersen. Maclaurin expansions of the electron momentum densities of linear molecules. International Journal of Quantum Chemistry, 36(S23):245–253, April 1–8, 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Takahashi:1983:PTE

[TPČ83] M. Takahashi, J. Paldus, and J. Čížek. Perturbation theory and electron correlation in extended systems: Cyclic polyene model. International Journal of Quantum Chemistry, 24(6): 707–727, December 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Theodorakopoulos:1986:MCT

[TPN86] G. Theodorakopoulos, I. D. Petsalakis, and C. A. Nicolaides. A method for the calculation of transition moments between electronic states of molecules using a different one-electron basis set for each state. International Journal of Quantum Chemistry, 29(3):399–406, March 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Tometsko:1980:APP

[TR80] Andrew M. Tometsko and Frederic Middlebrook Richards, editors. Applications of photochemistry in probing biological targets, volume 346 of Annals of the New York Academy of Sciences. New York Academy of Sciences, New York, NY, USA, 1980. ISBN 0-89766-080-3, 0-89766-081-1 (paperback). LCCN Q11 .N5 vol. 346.

Tatar:1982:ACP

[TR82] Robert C. Tatar and Sohrab Rabii. Application of configuration propagator theory to graphite intercalation compounds. International Journal of Quantum Chemistry, 22(S16):669– 670, March 1–13, 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Chemistry, Theory of Condensed Matter, and Propagator Methods in the Quantum Theory of Matter.

Thomson:1986:TSN

[TR86] Colin Thomson and Christopher A. Reynolds. A theoretical study of N-nitrosamine metabolites: Possible alkylating species in carcinogenesis by N,N’-dimethyl nitrosamine. International Journal of Quantum Chemistry, 30(6):751–762, December 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Trinajstic:1986:MCC

[Tri86] Nenad Trinajstić, editor. Mathematical and computational concepts in chemistry. Ellis Horwood series, mathematics and its applications. Ellis Horwood, New York, NY, USA, 1986. ISBN 0-470-20289-0 (Halsted Press). 365 pp. LCCN QD39.3.M3 I57 1985.

Truax:1983:TDS

[Tru83a] D. Rodney Truax. Time-dependent Schrödinger equations: Symmetry breaking, separation of variables, and nonlinear effects. International Journal of Quantum Chemistry, 23(2): 663–678, February 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Truhlar:1983:RPA

[Tru83b] Donald G. Truhlar. Recent progress in atomic and molecular collisions and the interface with electronic structure theory. International Journal of Quantum Chemistry, 24(S17):77–87, March 3–12, 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-state Theory, Collision Phenomena, and Computational Quantum Chemistry.

Takada:1980:ITA

[TS80] Toshikazu Takada and Fukashi Sasaki. Improved transformation algorithm of two-electron integrals from atomic orbital basis to a symmetry orbital basis. International Journal of Quantum Chemistry, 18(5):1157–1163, November 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Tantardini:1981:IVB

[TS81] Gian Franco Tantardini and Massimo Simonetta. Ab initio valence bond calculations. XI. Pyridine: Ground and ionized states. International Journal of Quantum Chemistry, 20(3): 705–715, September 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Terasaka:1982:NCH

[TS82] Toshitaka Terasaka and Kazuyoshi Sakamoto. Note on calculations by hypervirial relations and sufficient conditions for these relations: Box potential and harmonic oscillator models. International Journal of Quantum Chemistry, 21(3):543–556, March 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Turner:1986:PDM

[TSA86] Ralph Eric Turner, Masayoshi Senba, and Donald J. Arseneau. On the pressure dependence of the muon spin polarization in mixtures of noble gases. International Journal of Quantum Chemistry, 29(5):1493–1502, May 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Tang:1980:LAM

[TSL80] Au-Chin Tang, Chia-Chung Sun, and Bei-Foo Li. Lie algebra and molecular shell model. I. International Journal of Quantum Chemistry, 18(S14):521–540, March 10–20, 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory, Collision Phenomena, Quantum Statistics, and Computational Methods.

Tripathi:1989:SFE

[TST89] A. N. Tripathi, Vedene H. Smith, Jr., and Ajit J. Thakkar. Scattering of fast electrons and X-rays from molecules: CH_{4} and C_{2}H_{2}. International Journal of Quantum Chemistry, 35 (6):869–885, June 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Talekar:1981:MCB

[TT81] Manjubala S. Talekar and Subhash V. Talekar. Mechanism of chromosome banding: a phenomenological approach. International Journal of Quantum Chemistry, 20(2):555–566, August 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Tanner:1983:DCC

[TT83] A. C. Tanner and Ajit J. Thakkar. Discrete and continuum contributions to multipole polarizabilities and shielding factors of hydrogen. International Journal of Quantum Chemistry, 24(4):345–352, October 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Tostes:1988:RPF

[TTRL88] J. G. R. Tostes, C. A. Taft, M. N. Ramos, and W. A. Lester, Jr. Role of polarization functions on the bridge hydrogen atom in HCN·HCN. International Journal of Quantum Chemistry, 34(1):85–86, July 1988. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Thakkar:1984:AEI

[TTS84] Ajit J. Thakkar, A. N. Tripathi, and Vedene H. Smith, Jr. Anisotropic electronic intracule densities for diatomics. International Journal of Quantum Chemistry, 26(2):157–166, August 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Tuszynski:1986:CDH

[Tus86] J. A. Tuszyński. Comments on the Davydov Hamiltonian for quasi-one-dimensional molecular chains. International Journal of Quantum Chemistry, 29(3):379–391, March 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Tillieu:1983:UALa

[TV83a] Jacques Tillieu and Augustin Van Groenedael. Usage de l’algègbre de Lie su(n) dans l’étude des systèmes quantiques à n états. I. Quelques conséquences générales. (French) [Use of Lie algebra su(n) in the study of quantum systems of n states. I. Some general consequences]. International Journal of Quantum Chemistry, 23(5):1807–1816, May 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Tillieu:1983:UALb

[TV83b] Jacques Tillieu and Augustin Van Groenendael. Usage de l’algègbre de Lie su(n) dans l’étude des systèmes quantiques à n états. II. Transformation de l’espace des observables, problèmes d’évolution. (French) [Use of Lie algebra su(n) in the study of quantum systems of n states. II. Transformation of the space of observables, problems of evolution]. International Journal of Quantum Chemistry, 23(5):1817–1827, May 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Tillieu:1984:UAL

[TV84] Jacques Tillieu and Augustin Van Groenendael. Usage de l’algègbre de Lie su(n) dans l’étude des systèmes quantiques à n états. III. Définition et obtention des constantes du mouvement conjoncturelles. (French) [Use of Lie algebra su(n) in the study of quantum systems of n states. III. Definition and determination of conjectured constants of movement]. International Journal of Quantum Chemistry, 26(4):479–488, October 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Tillieu:1989:UAL

[TV89] Jacques Tillieu and Augustin Van Groenendael. Usage de l’algèbre de Lie su(n) dans l’étude des systèmes quantiques à n etats. VII: Sur le “problème inverse”. (French) [Use of Lie algebra su(n) in the study of quantum systems of n states. VII. On the “inverse problem”]. International Journal of Quantum Chemistry, 35(2):331–337, February 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Tvaroska:1989:CAO

[Tva89] Igor Tvaroška. Conformational analysis of oligosaccharides in solution. International Journal of Quantum Chemistry, 35(1):

141–151, January 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Tian:1988:ECP

[TW88] Bogang Tian and Guozhen Wu. EHMO Calculations for a pyridine molecule adsorbed on a Ag adatom: Implications for surface enhanced Raman spectroscopy. International Journal of Quantum Chemistry, 33(6):529–542, June 1988. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Trsic:1982:ESB

[TWL82] Milan Trsic, Kim Wagstaff, and William G. Laidlaw. Electronic structure and bonding of the pentasulfur hexanitride (S_{5}N_{6}) molecule. International Journal of Quantum Chemistry, 22(5):903–910, November 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Tyrrell:1981:BAA

[TWW81] J. Tyrrell, R. B. Weinstock, and F. Weinhold. Bond–antibond analysis of internal rotation barriers in glyoxal and related molecules: Where INDO fails. International Journal of Quantum Chemistry, 19(5):781–791, May 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Ugalde:1985:AAE

[UB85] J. M. Ugalde and Russell J. Boyd. Angular aspects of exchange correlation and the Fermi hole. International Journal of Quantum Chemistry, 27(4):439–449, April 1985. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Ugalde:1986:RBE

[UB86] J. M. Ugalde and Russell J. Boyd. On the relationship between the electron-pair distribution function and the correlation energy of an atom. International Journal of Quantum Chemistry, 29(1):1–9, January 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Uehara:1980:VIP

[UI80] Hiroshi Uehara and Takuma Itagaki. Vertical ionization potential in hydrogen molecule with many-body Green’s function method. International Journal of Quantum Chemistry, 18(1): 73–78, July 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Ulmer:1981:ESM

[Ulm81] W. Ulmer. Electronic structure of the metabolites of cyclophosphamide. International Journal of Quantum Chemistry, 19(2):337–359, February 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Ulmer:1983:UTD

[Ulm83] W. Ulmer. On a unified treatment of diffusion and kinetic processes. International Journal of Quantum Chemistry, 23 (6):1931–1944, June 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Ulmer:1985:UTK

[Ulm85] W. Ulmer. On a unified treatment of kinetics and diffusion, and a connection to a nonlocal Boltzmann equation. International Journal of Quantum Chemistry, 27(2):203–218, February 1985. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Umrigar:1989:TAQ

[Umr89] C. J. Umrigar. Two aspects of quantum Monte Carlo: Determination of accurate wavefunctions and determination of potential energy surfaces of molecules. International Journal of Quantum Chemistry, 36(S23):217–230, April 1–8, 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Urry:1982:HEC

[Urr82] Dan W. Urry. Henry Eyring (1901–1981): A 20th-century architect of cathedrals of science. International Journal of Quantum Chemistry, 22(S9):1–3, March 4–6, 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Urry:1987:EEF

[Urr87] Dan W. Urry. Entropic elastomeric force in protein structure/function. International Journal of Quantum Chemistry, 32(S14):261–280, March 14, 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Urry:1988:FEC

[Urr88] Dan W. Urry. Free energy (chemomechanical) transduction in elastomeric polypeptides by chemical potential modulation of an inverse temperature transition. International Journal of Quantum Chemistry, 34(S15):235–245, March 12–19, 1988. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Syposium on Quantum Biology and Quantum Pharmacology.

Uchida:1980:EME

[UT80] Kenji Uchida and Yoshihiro Takahashi. Excimer and monomer emission of pyrene crystals associated with the phase transition. International Journal of Quantum Chemistry, 18(1):301– 305, July 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Urry:1982:MSI

[UTP82] Dan W. Urry, Tina L. Trapane, and Kari U. Prasad. Molecular structure and ionic mechanisms of an ion-selective transmembrane channel: Monovalent versus divalent cation selectivity. International Journal of Quantum Chemistry, 22(S9):31–40, March 4–6, 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Urry:1985:EPS

[UTVP85] Dan W. Urry, Tina L. Trapane, C. M. Venkatachalam, and Kari U. Prasad. Energy profiles for sodium ion passage through the single-filing gramicidin transmembrane channel. International Journal of Quantum Chemistry, 28(S12):1–13, March 14–16, 1985. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Urry:1983:LEM

[UV83] Dan W. Urry and Cherayathumadom M. Venkatachalam. A librational entropy mechanism for elastomers with repeating peptide sequences in helical array. International Journal of Quantum Chemistry, 24(S10):81–93, March 14–16, 1983.

CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Urry:1981:CPG

[UVP^{+}81] D. W. Urry, C. M. Venkatachalam, K. U. Prasad, R. J. Bradley, G. Parenti-Castelli, and G. Lenaz. Conduction processes of the gramicidin channel. International Journal of Quantum Chemistry, 20(S8):385–399, March 5–7, 1981. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

VanKampen:1982:CED

[Van82] N. G. Van Kampen. Cluster expansions for diffusion-controlled reactions. International Journal of Quantum Chemistry, 22 (S16):101–115, March 1–13, 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Chemistry, Theory of Condensed Matter, and Propagator Methods in the Quantum Theory of Matter.

VanGroenendael:1985:UALa

[Van85] Augustin Van Groenendael. Usage de l’algègbre de Lie su(n) dans l’étude des systèmes quantiques à n états. IV. Geometrie du domaine des vecteurs de cohérence. (French) [Use of Lie algebra su(n) in the study of quantum systems of n states. IV. Geometry of the domain of coherence vectors]. International Journal of Quantum Chemistry, 27(4):417–425, April 1985. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

VanGroenendael:1986:UAL

[Van86] A. Van Groenendael. Usage de l’algègbre de Lie su(n) dans l’étude des systèmes quantiques à n états. VI. Équations linéaires d’évolution positive. (French) [Use of Lie algebra su(n) in the study of quantum systems of n states. VI. Positive linear equations of evolution]. International Journal of Quantum Chemistry, 30(1):143–153, July 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Varma:1981:SLE

[Var81a] Shambhu D. Varma. Superoxide and lens of the eye: a new theory of cataractogenesis. International Journal of Quantum Chemistry, 20(2):479–484, August 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Varracchio:1981:EPT

[Var81b] E. Ficocelli Varracchio. Effective potential theory of atom– diatom scattering in the presence of “intermediate” reactive channels. International Journal of Quantum Chemistry, 19 (4):655–671, April 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Varracchio:1985:FTC

[Var85] E. Ficocelli Varracchio. Field theory of chemical reactions. I. Ideal space representation of A+BC → AC +B processes. International Journal of Quantum Chemistry, 27(1):11–44, January 1985. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Varma:1986:RRC

[Var86] Matesh N. Varma. Radiobiology and radiological and chemical physics research as an aid in development of carcinogenesis models. International Journal of Quantum Chemistry, 30(S13):291–295, March 6–8, 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Varandas:1987:DMB

[Var87] A. J. C. Varandas. The double many-body expansion of potential energy surfaces from interacting ^{2}S atoms. International Journal of Quantum Chemistry, 32(5):563–574, November 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Venkatachalam:1984:LGA

[VARPU84] C. M. Venkatachalam, S. Alonso-Romanowski, K. U. Prasad, and D. W. Urry. The Leu^{5} gramicidin A analog: Molecular mechanics calculations and analysis of single channel steps related to multiplicity of conducting states. International Journal of Quantum Chemistry, 26(S11):315–326, March 12–15, 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Vassileva:1982:FIC

[Vas82] Julia G. Vassileva. The fourth international colloquium on physical and chemical information transfer in the regulation of reproduction and ageing. International Journal of Quantum Chemistry, 22(3):663, September 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Venanzi:1985:MRA

[VB85] Carol A. Venanzi and Jeffrey D. Bunce. Molecular recognition by artificial enzymes: Cyclic urea mimetics of alpha chymotrypsin. International Journal of Quantum Chemistry, 28 (S12):69–87, March 14–16, 1985. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Ventura:1985:CBH

[VBS85] Oscar N. Ventura, J. Paolo Bartolucci, and Ramón M. Sosa. He_{2}^{2+}: a comparison between Hartree–Fock and density functional techniques. International Journal of Quantum Chemistry, 27(5):625–635, May 1985. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Vercauteren:1983:PSB

[VC83] D. P. Vercauteren and E. Clementi. Photoionization spectra of B-DNA and Z-DNA. International Journal of Quantum Chemistry, 24(S10):11–30, March 14–16, 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Vasilescu:1983:CAA

[VCB83] D. Vasilescu, D. Cabrol, and H. Broch. Conformons in amino acid sequences typical of collagen: A quantum mechanical study. International Journal of Quantum Chemistry, 24(S10): 109–122, March 14–16, 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Vela:1986:IID

[VCG86] Alberto Vela, Andrés Cedillo, and José L. Gázquez. Interatomic interactions in density functional theory. International Journal of Quantum Chemistry, 29(4):937–948, April 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Vinette:1987:RIP

[VCV87] Francine Vinette, Jiri Cizek, and Edward R. Vrscay. Renormalized inner projection—a case study: The octic oscillator. International Journal of Quantum Chemistry, 32(6):663–667, December 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

vanDuijnen:1980:ASH

[vDTBN80] P. Th. van Duijnen, B. Th. Thole, Ria Broer, and W. C. Nieuwpoort. Active-site α-helix in papain and the stability of the ion pair RS^{−}⋯ImH^{+}. Ab initio molecular orbital study. International Journal of Quantum Chemistry, 17(4):651–671, April 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Verwoerd:1980:VCS

[Ver80] W. S. Verwoerd. Variation of calculated single bond lengths in saturated hydrocarbon and silicon hydride molecules. International Journal of Quantum Chemistry, 18(6):1449–1463, December 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Vercauteren:1981:SDO

[VFA81] D. P. Vercauteren, J. G. Fripiat, and J. M. André. Symmetry dilemmas in oxocarbons. International Journal of Quantum Chemistry, 20(S15):219–224, March 8–14, 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory, Collision Phenomena, and Computational Quantum Chemistry.

Vela:1980:ESD

[VG80] Alberto Vela and Andoni Garritz. Electronic structure of dichloro-bridged ruthenium(II) dimers. International Journal of Quantum Chemistry, 18(S14):191–200, March 10–20, 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory, Collision Phenomena, Quantum Statistics, and Computational Methods.

Vela:1988:OBC

[VGG88] Alberto Vela, Marcelo Galván, and José Luis Gázquez. On the oscillatory behavior of the chemical potential of neutral atoms. International Journal of Quantum Chemistry, 34(S22): 329–335, March 12–19, 1988. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Syposium on Quantum Chemistry, Solid-State Theory, and Computational Methods.

VonBoehm:1984:RCV

[VI84] J. Von Boehm and H. M. Isomäki. Relativistic corrections to the valence and conduction band edges of anisotropic semiconductors. International Journal of Quantum Chemistry, 25(1): 263, January 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Villar:1989:MMB

[VL89] H. O. Villar and G. H. Loew. Molecular modulators of benzodiazepine receptor ligand binding. International Journal of Quantum Chemistry, 36(S16):261–271, April 1–8, 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Vilallonga:1981:TCF

[VM81] Eduardo Vilallonga and David A. Micha. Time-correlation function approach to molecular anharmonicity in hyperthermal atom-molecule collisions. International Journal of Quantum Chemistry, 20(S15):653–661, March 8–14, 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory, Collision Phenomena, and Computational Quantum Chemistry.

Vilallonga:1983:AET

[VM83] Eduardo Vilallonga and David A. Micha. Asymptotic expansions of time-correlation functions for energy transfer in molecular collisions. International Journal of Quantum Chemistry, 24(S17):194, March 3–12, 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-state Theory, Collision Phenomena, and Computational Quantum Chemistry.

Vitkovskaya:1980:QCI

[VMMV80] N. M. Vitkovskaya, V. B. Mantsivoda, T. E. Moskovskaya, and M. G. Voronkov. Quantum-chemical investigation of the silicon and carbon coordination bond in their isostructural compounds. International Journal of Quantum Chemistry, 17 (2):299–305, February 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Vogler:1986:SCS

[Vog86] H. Vogler. Structures and ^{1}H-chemical shifts of conjugation deficient hydrocarbons. International Journal of Quantum Chemistry, 30(1):97–107, July 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Vojtik:1985:SAD

[Voj85] Jan Vojtík. Symmetry aspects of diatomics-in-molecules (DIM) calculations: Construction of spin-adapted bases. International Journal of Quantum Chemistry, 28(5):593–601, November 1985. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Volosov:1988:OAS

[Vol88] Andrey Volosov. Optical activity of small peptide chains with right-handed α-helix structure. International Journal of Quantum Chemistry, 34(5):471–481, November 1988. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Volosov:1989:E

[Vol89a] Andrey Volosov. Erratum. International Journal of Quantum Chemistry, 36(4):543, October 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Volosov:1989:RDE

[Vol89b] Andrey Volosov. Role of doubly excited states in calculations of optical activity. International Journal of Quantum Chemistry, 36(4):473–486, October 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Viste:1984:SOE

[VP84] A. Viste and P. Pyykkö. Spin-orbit excitation in the system I + I_{2}: Relativistically parameterized extended-Hückel calculations. International Journal of Quantum Chemistry, 25(1): 223–231, January 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

VanZandt:1980:MAD

[VPK80] L. L. Van Zandt, E. W. Prohofsky, and M. Kohli. Microwave absorption by double-helical DNA. International Journal of Quantum Chemistry, 18(S7):35–38, March 5–8, 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Vrbik:1986:OSW

[VR86] Jan Vrbik and Stuart M. Rothstein. Optimal spacing and weights in diffusion Monte Carlo. International Journal of Quantum Chemistry, 29(3):461–468, March 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Vrscay:1987:RSPa

[Vrs87a] Edward R. Vrscay. Rayleigh–Schrödinger perturbation expansions for a hydrogen atom in a polynomial perturbation. International Journal of Quantum Chemistry, 32(5):613–620, November 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Vrscay:1987:RSPb

[Vrs87b] Edward R. Vrscay. Rayleigh–Schrödinger perturbation expansions for a hydrogen atom in a polynomial perturbation. International Journal of Quantum Chemistry, 32(S21):751, March 12, 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Chemistry, Solid-State Theory, and Computational Methods.

Viana:1982:PSC

[VT82] Luci Martins Viana and Yuji Takahata. PCILO study of the conformation of 20-oxopregnane hormones. International Journal of Quantum Chemistry, 22(2):265–274, August 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

VanGroenendael:1985:UALb

[VT85] Augustin Van Groenendael and Jacques Tillieu. Usage de l’algègbre de Lie su(n) dans l’étude des systèmes quantiques à n états. V. Application aux systèmes composés. (French) [Use of Lie algebra su(n) in the study of quantum systems of n states. V. Application to composite systems]. International Journal of Quantum Chemistry, 28(2):269–286, August 1985. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Valone:1981:DEA

[VTT81] Steven M. Valone, Devarajan Thirumalai, and Donald G. Truhlar. Dispersion-equation approach to obtaining complex optical potentials for electron scattering. International Journal of Quantum Chemistry, 20(S15):341–353, March 8–14, 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory, Collision Phenomena, and Computational Quantum Chemistry.

Venkatachalam:1985:CDM

[VU85] C. M. Venkatachalam and D. W. Urry. Calculation of dipole moment changes due to peptide librations in the dynamic βspiral of the polypentapeptide of elastin. International Journal of Quantum Chemistry, 28(S12):15–24, March 14–16, 1985. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Vekantchalam:1988:E

[VU88] C. M. Vekantchalam and D. W. Urry. Erratum. International Journal of Quantum Chemistry, 33(2):171, February 1988. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Vinszky:1983:GEC

[VV83] P. Csa Vinszky and F. Vosman. Gradient expansion correction to the Dirac exchange term in statistical models for the Na atom with shell structure. International Journal of Quantum Chemistry, 23(6):1973–1978, June 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Vasilescu:1987:MSA

[VV87] D. Vasilescu and R. Viani. Molecular similarity in aminothiol radioprotectors: a Randič graph approach. International Journal of Quantum Chemistry, 32(S14):149–165, March 14, 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

VanCamp:1980:VSM

[VVD80a] P. E. Van Camp, V. E. Van Doren, and J. T. Devreese. Vibrational spectra of molecules in the Hartree–Fock dielectric screening approach. International Journal of Quantum Chemistry, 18(1):311–315, July 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

VanDoren:1980:MTA

[VVD80b] V. E. Van Doren, P. E. Van Camp, and J. T. Devreese. Microscopic theory of the average band gap of semiconductors. International Journal of Quantum Chemistry, 18(2):615–618, August 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Venanzi:1982:SEE

[VWCC82] Carol A. Venanzi, Harel Weinstein, Georgina Corongiu, and Enrico Clementi. The solvent effect in enzyme–substrate interactions: Models of carboxypeptidase. International Journal of Quantum Chemistry, 22(S9):355–365, March 4–6, 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

vanZandt:1981:DDV

[vZ81] L. L. van Zandt. Damping of DNA vibration modes by viscous solvents. International Journal of Quantum Chemistry, 20(S8):271–276, March 5–7, 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Volkmann:1982:TIE

[VŻH82] D. Volkmann, B. Żrawski, and D. Heidrich. Theoretical investigation of the electrophilic attack. IX. Ab initio study of the C_{2}H_{4}·HF molecular complex. International Journal of Quantum Chemistry, 22(3):631–637, September 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Wagner:1983:CSH

[Wag83] A. J. Wagner. Comparative study of the Hartree–Fock and the Hartree–Fock–Slater method. International Journal of Quantum Chemistry, 23(1):235–248, January 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Wald:1984:LMU

[Wal84] George Wald. Life and mind in the universe. International Journal of Quantum Chemistry, 26(S11):1–15, March 12–15, 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Walnut:1988:PEC

[Wal88] Thomas H. Walnut. Properties of electronic current induced by nuclear motion: Interaction with radiation. International Journal of Quantum Chemistry, 34(S22):99–116, March 12– 19, 1988. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Syposium on Quantum Chemistry, Solid-State Theory, and Computational Methods.

Wasilewski:1989:GTC

[Was89] J. Wasilewski. Graphical techniques in the configuration interaction approach based on pure Slater determinants. International Journal of Quantum Chemistry, 36(4):503–524, October 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Wardlaw:1981:SDT

[WB81] David Wardlaw and Paul Brumer. State densities and time delay in molecular collisions. International Journal of Quantum Chemistry, 20(S15):583–594, March 8–14, 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory, Collision Phenomena, and Computational Quantum Chemistry.

Wierzbicki:1989:NNA

[WB89] Andrzej Wierzbicki and Joel M. Bowman. New numerical approaches to the solution of the N-well Schrödinger equation. International Journal of Quantum Chemistry, 35(2):297–303, February 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Warner:1980:ECH

[WBB80] J. W. Warner, L. S. Bartell, and S. M. Blinder. Electron correlation and Hund’s rule. International Journal of Quantum Chemistry, 18(4):921–935, October 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Witko:1986:PCS

[WBK86] Małgorzata Witko and Vlasta Bonačić-Koutecký. The pseudopotential-CI study of the interaction between ethylene and metal atoms, metal Oxides, and their cations (Me = Be, Mg, and Zn). International Journal of Quantum Chemistry, 29(5):1535–1554, May 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Woodward:1988:ESZ

[WBM88] C. Woodward, R. Benedek, and B. I. Min. Electronic structure of Zintl phase compounds by nonlinear optimization. International Journal of Quantum Chemistry, 34(S22):219–223, March 12–19, 1988. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Syposium on Quantum Chemistry, Solid-State Theory, and Computational Methods.

Wheeler:1989:TSP

[WCE89] Dale E. Wheeler, Leszek Czuchajowski, and W. Daniel Edwards. Theoretical studies of phenol disubstituted porphyrins.

International Journal of Quantum Chemistry, 36(S23):473– 482, April 1–8, 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Waleh:1986:CES

[WCLZ86] Ahmad Waleh, Jack R. Collins, Gilda H. Loew, and M. C. Zerner. Calculated electronic spectra of model ferric cytochrome P450 complexes. International Journal of Quantum Chemistry, 29(5):1575–1589, May 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Wedler:1981:IAH

[WD81] G. Wedler and W. Draxler. Influence of the adsorption of hydrogen on the transport properties of thin nickel films. International Journal of Quantum Chemistry, 19(6):1027–1030, June 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Weatherford:1982:KVP

[Wea82] Charles A. Weatherford. Kohn variational principle for a sphere of arbitrary radius. International Journal of Quantum Chemistry, 22(S16):277–284, March 1–13, 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Chemistry, Theory of Condensed Matter, and Propagator Methods in the Quantum Theory of Matter.

Weatherford:1983:VRF

[Wea83] Charles A. Weatherford. Vibrational resonances in the frametransformation theory of electron-H_{2} scattering. International Journal of Quantum Chemistry, 24(S17):195, March 3–12, 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-state Theory, Collision Phenomena, and Computational Quantum Chemistry.

Weatherford:1984:EMP

[Wea84] Charles A. Weatherford. Electron–molecule potentials using the Gegenbauer addition theorem. International Journal of Quantum Chemistry, 26(S18):735, March 1–15, 1984.

CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Atomic, Molecular and Solid-State Theory, and Computational Quantum Chemistry.

Weatherford:1988:AMT

[Wea88] Charles A. Weatherford. An analytic method for three-center nuclear attraction integrals: a generalization of the Gegenbauer addition theorem. International Journal of Quantum Chemistry, 33(1):19–26, January 1988. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Wedler:1981:EEC

[Wed81] G. Wedler. Experimental evidence on the chemisorption bond with the chemisorption on metals. International Journal of Quantum Chemistry, 19(6):1011–1016, June 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Weiner:1982:GLG

[Wei82] B. Weiner. General linear group approach to bivariational problems. International Journal of Quantum Chemistry, 21 (3):591–610, March 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Weiner:1985:CPA

[Wei85] Brian Weiner. Consistent propagator approximations. International Journal of Quantum Chemistry, 28(1):61–84, July 1985. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Weiner:1988:LAS

[Wei88] Brian Weiner. The Lie algebraic structure of the set of oneparticle fermion operators. International Journal of Quantum Chemistry, 34(S22):407–424, March 12–19, 1988. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Syposium on Quantum Chemistry, Solid-State Theory, and Computational Methods.

Wen:1986:CSC

[Wen86] Zhenyi Wen. Calculation of S_{N1+N2} ⊃ S_{N1} ⊗ s_{n2} and U(n_{1} + n_{2}) ⊃ u(n_{1}) ⊗ u(n_{2}) subduction coefficients by using spin graph. International Journal of Quantum Chemistry, 29 (6):1779–1787, June 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Westhaus:1981:CBP

[Wes81] Paul Westhaus. Connections between perturbation theory and the unitary transformation methods of deriving effective Hamiltonians. International Journal of Quantum Chemistry, 20(6):1243–1253, December 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Westhaus:1982:CBP

[Wes82] Paul Westhaus. Connections between perturbation theory and the Van Vleck transformation: Illustrative calculations on the perturbed harmonic oscillator. International Journal of Quantum Chemistry, 21(3):617–632, March 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Weiner:1980:SCA

[WG80] Brian Weiner and Osvaldo Goscinski. Self-consistent approximation to the polarization propagator. International Journal of Quantum Chemistry, 18(4):1109–1131, October 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Weiner:1982:SAP

[WG82] Brian Weiner and Osvaldo Goscinski. Superoperator approach to propagator approximations. International Journal of Quantum Chemistry, 21(2):369–398, February 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Weniger:1985:SAN

[WGS85] E. J. Weniger, J. Grotendorst, and E. O. Steinborn. Some applications of nonlinear convergence accelerators. International Journal of Quantum Chemistry, 28(S19):181–191, March 18– 23, 1985. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Williamson:1987:GOO

[WH87] Rodney L. Williamson and Michael B. Hall. Geometry optimization of organometallic complexes: A study of basis sets.

International Journal of Quantum Chemistry, 32(S21):503– 512, March 12, 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Chemistry, Solid-State Theory, and Computational Methods.

Wojcik:1986:ICC

[WHT86] Marek J. Wójcik, Akiko Y. Hirakawa, and Masamichi Tsuboi. Ab initio calculation of cubic force constants in formamide monomer, formamide dimer, formic acid dimer and their D, D_{2} and D_{4} derivatives. An estimation of Fermi resonance and intra–intermolecular coupling. International Journal of Quantum Chemistry, 30(S13):133–142, March 6–8, 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Wijkman:1983:OA

[Wij83a] Gunnar Wijkman. Opening address. International Journal of Quantum Chemistry, 24(S17):3–4, March 3–12, 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-state Theory, Collision Phenomena, and Computational Quantum Chemistry.

Wijkman:1983:OR

[Wij83b] Gunnar Wijkman. Opening remarks. International Journal of Quantum Chemistry, 24(S10):IX–X, March 14–16, 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Wilson:1980:SPS

[Wil80a] E. Bright Wilson. Some personal scientific reminiscences. International Journal of Quantum Chemistry, 18(S14):17–29, March 10–20, 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory, Collision Phenomena, Quantum Statistics, and Computational Methods.

Wilson:1980:FOI

[Wil80b]  Stephen Wilson.  Fourth-order invariant in Rayleigh–

Schrödinger perturbation theory. International Journal of Quantum Chemistry, 18(3):905–906, September 1980. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Wilson:1981:JHV

[Wil81] E. Bright Wilson. John Hasbrouck Van Vleck (1899–1980). International Journal of Quantum Chemistry, 20(S15):1–3, March 8–14, 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory, Collision Phenomena, and Computational Quantum Chemistry.

Wilhelmsson:1989:SFE

[Wil89] H. Wilhelmsson. Self-formation and evolution of singletons. International Journal of Quantum Chemistry, 35(6):887–893, June 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Winkler:1982:PVD

[Win82] Peter Winkler. Propagator version of the direct Siegert method. International Journal of Quantum Chemistry, 22 (S16):641–648, March 1–13, 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Chemistry, Theory of Condensed Matter, and Propagator Methods in the Quantum Theory of Matter.

Winkler:1983:UVP

[Win83] Peter Winkler. Use of V^{N−1} potentials in Green’s function calculations. International Journal of Quantum Chemistry, 23(4):1571–1577, April 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Winkler:1985:UMB

[Win85] Peter Winkler. A unified many-body theory of dielectronic recombination and autoionization. International Journal of Quantum Chemistry, 28(S19):201–216, March 18–23, 1985.

CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Winkler:1986:SRR

[Win86] Peter Winkler. Sum-rules in resonance calculations with complex coordinates. International Journal of Quantum Chemistry, 30(S20):147–151, March 10–15, 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Syposium on Atomic, Molecular, and Solid-State Theory, Scattering Problems, Many Body Phenomena, and Computational Quantum Chemistry.

Witschel:1981:OPE

[Wit81] W. Witschel. Ordered products of exponential operators by similarity transformations. International Journal of Quantum Chemistry, 20(6):1233–1241, December 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Weatherford:1980:AAE

[WJ80] C. A. Weatherford and B. L. Jain. An analytic approach to electron-molecule potentials using STOs. International Journal of Quantum Chemistry, 18(S14):493–502, March 10– 20, 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory, Collision Phenomena, Quantum Statistics, and Computational Methods.

Weatherford:1982:EMM

[WJ82] Charles A. Weatherford and Herbert W. Jones, editors. ETO Multicenter Molecular Integrals: Proceedings of the International Conference held at Florida A&M University, Tallahassee, Florida, USA, August 3–6, 1981. D. Reidel, Dordrecht, Boston, Lancaster, Tokyo, 1982. ISBN 90-277-1488-6. LCCN

QD461 E86 1982. URL http://www.gbv.de/dms/tib-ub-
hannover/243261756.pdf.

Wilson:1983:AND

[WJP83] S. Wilson, K. Janmkowski, and J. Paldus. Applicability of non-degenerate many-body perturbation theory to quasidegenerate electronic states: a model study. International Journal of Quantum Chemistry, 23(5):1781–1802, May 1983. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Wilson:1985:ANM

[WJP85] S. Wilson, K. Jankowski, and J. Paldus. Applicability of nondegenerate many-body perturbation theory to quasidegenerate electronic states. II. A two-state model. International Journal of Quantum Chemistry, 28(4):525–534, October 1985. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Watanabe:1983:IMC

[WK83] Yoshitaka Watanabe and Hiroshi Kashiwagi. Ab initio MO Calculations of benzene + TCNE and naphthalene + TCNE complexes with STO-3G π-split basis set. International Journal of Quantum Chemistry, 23(5):1739–1752, May 1983. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Weiner:1985:AGP

[WK85] Brian Weiner and Henry A. Kurtz. The antisymmetrized geminal power approximation to the excitation propagator. International Journal of Quantum Chemistry, 27(6):769–780, June 1985. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Waltz:1980:CME

[WKH80] Carl E. Waltz, Hae-Won Kim, and Hendrik F. Hameka. Calculation of molecular electric polarizabilities and dipole moments. III. BH molecule. International Journal of Quantum Chemistry, 17(3):479–483, March 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Westbrook:1988:IAE

[WKJ88] John D. Westbrook and Karsten Krogh-Jespersen. Ab initio all-electron and effective core potential calculations on CuCl. International Journal of Quantum Chemistry, 34(S22):245– 255, March 12–19, 1988. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Syposium on Quantum Chemistry, Solid-State Theory, and Computational Methods.

Weinstein:1983:MDR

[WMP^{+}83] Harel Weinstein, Saul Maayani, Barak Pazhenchevsky, Carol Venanzi, and Roman Osman. Molecular determinants for recognition of phencyclidine derivatives at muscarinic cholinergic receptors. International Journal of Quantum Chemistry, 24(S10):309–320, March 14–16, 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Wagner:1981:SDO

[WMS81] Belinda J. Wagner, Chanchal K. Mitra, and Ramaswamy H. Sarma. Structure of DNA oligonucleotides. International Journal of Quantum Chemistry, 20(S8):277–285, March 5–7, 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Wojcik:1986:NTI

[Wój86a] Marek J. Wójcik. Note on theoretical interpretation of infrared spectra of Cl H and Cl D in gaseous (CH_{3})_{2}O · HCl and (CH_{3})_{2}O · DCl complexes. International Journal of Quantum Chemistry, 30(4):567–569, October 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Wojcik:1986:TII

[Wój86b] Marek J. Wójcik. Theoretical interpretation of infrared spectra of the Cl H stretching vibration in the gaseous (Ch_{3})_{2}O⋯HCl complex. International Journal of Quantum Chemistry, 29(4):855–865, April 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Wold:1981:BRB

[Wol81] Svante Wold. Book review: Factor analysis in chemistry. E. R. Malinowski and D. G. Howery. Wiley-Interscience, New York, 1980. 250 pp., Price: $26.25. International Journal of Quantum Chemistry, 19(5):969, May 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). See [MH80b, MH89, Mal91, Mal02].

Wolfskill:1988:PIP

[Wol88] Troy Wolfskill. Preliminary investigations of puckering in cyclooctatetraene: Effects of spin multiplicity and oxidation state. International Journal of Quantum Chemistry, 34(S22): 739–740, March 12–19, 1988. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Syposium on Quantum Chemistry, Solid-State Theory, and Computational Methods.

Woolley:1980:QDM

[Woo80] R. G. Woolley, editor. Quantum dynamics of molecules: the new experimental challenge to theorists, volume 57 of Nato advanced study institutes series: Series B, Physics. Plenum Press, New York, NY, USA; London, UK, 1980. ISBN 0-306-40462-1. LCCN QD462.A1 N17 1979.

Woodruff:1982:ECL

[Woo82] Truman O. Woodruff. Excitation of core levels in nonmetallic solids: Some theoretical problems. International Journal of Quantum Chemistry, 22(S16):451–456, March 1–13, 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Chemistry, Theory of Condensed Matter, and Propagator Methods in the Quantum Theory of Matter.

Wormer:1981:NMP

[Wor81] P. E. S. Wormer. Note on the Matsen and Pauncz proof of the S_{n}-U(n) duality. International Journal of Quantum Chemistry, 20(3):777–778, September 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Wormer:1985:MEB

[Wor85] Paul E. S. Wormer. Matrix elements between spin-bonded functions in a hole-particle formalism. International Journal of Quantum Chemistry, 28(5):609–630, November 1985. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Weatherford:1989:SPA

[WOT89] Charles A. Weatherford, Gregory Odom, and Roy Tucker. A study of pseudoresonances in the application of the Schwinger variational principle to electron scattering from atomic hydrogen. International Journal of Quantum Chemistry, 36(S23): 95–101, April 1–8, 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Wormer:1980:CIM

[WP80] P. E. S. Wormer and J. Paldus. Configuration interaction matrix elements. III. Spin functions relating the unitary and symmetric group approaches. International Journal of Quantum Chemistry, 18(3):841–866, September 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Wieting:1979:EPL

[WS79] T. J. Wieting and Michael Schlüter, editors. Electrons and phonons in layered crystal structures, volume 3 of Physics and chemistry of materials with layered structures. D. Reidel, Dordrecht, Boston, Lancaster, Tokyo, 1979. ISBN 90-277-0841-X, 90-277-0897-5. x + 474 pp. LCCN QD921 .E5.

Weintraub:1980:CAA

[WS80a] H. J. R. Weintraub and Paul G. Seybold. Conformational analysis of anilinonaphthalenesulfonate derivatives. I. 2-anilino-6-naphthalenesulfonate. International Journal of Quantum Chemistry, 18(S7):115–122, March 5–8, 1980. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Wolf:1980:NCD

[WS80b] Andreas Wolf and Hans-Herbert Schmidtke. Nonempirical calculations on diatomic transition metals. II. RHF investigation of lowest closed-shell states of homonuclear 3d transitionmetal dimers. International Journal of Quantum Chemistry, 18(5):1187–1205, November 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Wilson:1981:SLP

[WS81] R. G. Wilson and C. S. Sharma. Studies in Löwdin’s partitioning technique in quantum theory. I. Invertibility of the operator K. International Journal of Quantum Chemistry, 20 (4):927–932, October 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Wilson:1982:SLP

[WS82] R. G. Wilson and C. S. Sharma. Studies in Löwdin’s partitioning technique in quantum theory. II. Operators T and Ω. International Journal of Quantum Chemistry, 22(1):183– 190, July 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Wahl:1980:DCE

[WSD80] Arnold C. Wahl, Edward S. Sachs, and John H. Detrich. The differential correlation energy method for the calculation of electron affinities, ionization potentials, and vertical spectra. International Journal of Quantum Chemistry, 18(S14):261, March 10–20, 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory, Collision Phenomena, Quantum Statistics, and Computational Methods.

Wu:1982:LOP

[Wu82] Tai Tsun Wu. Large-order perturbation theory. International Journal of Quantum Chemistry, 21(1):105–118, January 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Wesenberg:1982:CIR

[WW82] G. Wesenberg and F. Weinhold. Coupling of internal rotations in propanelike molecules. International Journal of Quantum Chemistry, 21(2):487–509, February 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Wang:1989:ASR

[WW89] Zhengming Wang and Peter Winkler. Analytic solution to the resonance problem of screened Coulomb potentials. International Journal of Quantum Chemistry, 36(S23):89–94, April 1–8, 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Wilkinson:1988:EIC

[WWD88] A. J. Wilkinson, C. M. Warwick, and R. H. Davies. Electrostatic interactions and conformation of some thromboxane antagonists. International Journal of Quantum Chemistry, 34 (S15):67–84, March 12–19, 1988. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Syposium on Quantum Biology and Quantum Pharmacology.

Xu:1986:ESC

[XHR86] Guang-Xian Xu, K. H. Hsu, and Jingqing Ren. Electronic structure and chemical bonding of the dimer of bis(η^{5}- cyclopentadienyl)ytterbium methyl. International Journal of Quantum Chemistry, 29(4):1017–1024, April 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Xiao-Yuan:1981:CEP

[XYSJ81] Fu Xiao-Yuan and Su Shu-Jun. Calculation of energies of protonation and tautomerization of methyl histamines and their relationship with pK_{a} values and biological activities. International Journal of Quantum Chemistry, 20(S8):197–200, March 5–7, 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Yamabe:1987:EVC

[YA87] Tokio Yamabe and Yoshihiro Asai. Effect of vibronic coupling on the long range intermolecular interaction. International Journal of Quantum Chemistry, 32(S21):569–572, March 12, 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Chemistry, Solid-State Theory, and Computational Methods.

Yadav:1983:QCS

[Yad83] Janardan Singh Yadav. Quantum chemical studies on the conformational structure of bacterial peptidoglycan. V. PCILO Calculations on β(1 → 4) linked disaccharide of N-acetyl-glucosamine and N-acetyl muramic acid. International Journal of Quantum Chemistry, 23(4):1441–1450, April 1983. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Yamaguchi:1980:IUH

[Yam80a] Kizashi Yamaguchi. Ab initio unrestricted Hartree–Fock (UHF) and UHF-natural orbital CI studies of ozone. International Journal of Quantum Chemistry, 18(1):101–106, July 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Yamaguchi:1980:MMC

[Yam80b] Kizashi Yamaguchi. Multireference (MR) configuration interaction (CI) approach for quasidegenerate systems. International Journal of Quantum Chemistry, 18(S14):269–284, March 10–20, 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory, Collision Phenomena, Quantum Statistics, and Computational Methods.

Yamaguchi:1981:TSP

[Yam81] Kizashi Yamaguchi. Theoretical studies of photo-oxidative cleavage reactions of nitrogen-activated C = C double bonds of enamines, indoles, and tryptamines. International Journal of Quantum Chemistry, 20(2):393–406, August 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Yamaguchi:1982:OSO

[Yam82] Kizashi Yamaguchi. Orbital symmetry, orbital stability, and orbital pairing rules for organic reactions in the ground state. International Journal of Quantum Chemistry, 22(3):459–484, September 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Yang:1982:JMS

[Yan82] Chen-Ning Yang. Joseph Mayer and statistical mechanics. International Journal of Quantum Chemistry, 22(S16):21–24, March 1–13, 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Chemistry, Theory of Condensed Matter, and Propagator Methods in the Quantum Theory of Matter.

Yang:1985:ROT

[Yan85] Cao Yang. The role of d orbitals in tetrahedral hybrid orbitals of oxy-ions. International Journal of Quantum Chemistry, 28 (5):641–648, November 1985. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Yaris:1982:PMP

[Yar82] Robert Yaris. Propagators in a model of polymer dynamics: The inclusion of chain–chain interactions in a model of polymer local main chain motion. International Journal of Quantum Chemistry, 22(S16):665, March 1–13, 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Chemistry, Theory of Condensed Matter, and Propagator Methods in the Quantum Theory of Matter.

Yarkony:1987:QLH

[Yar87] David R. Yarkony. On the quenching of Li(^{2}p) by HCl: Nonadiabatic effects. International Journal of Quantum Chemistry, 31(1):91–97, January 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Yadav:1986:APT

[YG86] J. S. Yadav and J. D. Goddard. Acetaldehyde photochemistry: Testing the additivity assumption and polarized basis set effects. International Journal of Quantum Chemistry, 29 (3):519–526, March 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Yuan:1987:TGS

[YG87] J. M. Yuan and Yan Gu. Transition to global stochasticity in a driven Morse oscillator. International Journal of Quantum Chemistry, 32(S21):743–748, March 12, 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Chemistry, Solid-State Theory, and Computational Methods.

Yliniemela:1989:HAB

[YGK^{+}89] A. Yliniemelä, J. Gynther, H. Konschin, H. Tylli, and J. Rouvinen. High-affinity benzodiazepine receptor ligands among benzodiazepines and betacarbolines with different intrinsic activity. International Journal of Quantum Chemistry, 36(S16): 273–280, April 1–8, 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Yadav:1989:CBI

[YHG89] Janardan Singh Yadav, Mark Hermsmeier, and Tamara Gund. Comparison between ab initio and semiempirical net atomic charges of some nicotinic acetylcholine receptor agonists. International Journal of Quantum Chemistry, 36(S16):101–117, April 1–8, 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Yoon:1985:IWS

[YJ85] Chang No Yoon and Mu Shik Jhon. Intermediate water structures in solution of alanine dipeptide. International Journal of Quantum Chemistry, 28(S12):33–47, March 14–16, 1985. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Yoshikawa:1980:PPS

[YKTA80] Kenichi Yoshikawa, Noriyuki Kumazawa, Hiroshi Terada, and Kazuo Akagi. Physicochemical properties of SF 6847, a potent uncoupler of oxidative phosphorylation in mitochondria in relation to its activity. International Journal of Quantum Chemistry, 18(2):539–544, August 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Yu:1989:GNV

[YKZ^{+}89] Chin-Hui Yu, Donald J. Kouri, Meishan Zhao, Donald G. Truhlar, and David W. Schwenke. Generalized Newton variational principle — L^{2} amplitude density treatment of the 3-dimensional quantal reaction F + H_{2} → HF(v_{f}) + H: Comparison of reaction probabilities and state-to-state collisional delay times for zero and nonzero total angular momentum. International Journal of Quantum Chemistry, 36(S23):45–58, April 1–8, 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Yadav:1983:CSL

[YL83] Janardan Singh Yadav and Peter Luger. Conformational studies of the β(1 → 3) linked disaccharide component of the immunologically active pentasaccharide of the Forssman antigen. International Journal of Quantum Chemistry, 23(4): 1433–1439, April 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Yusaf:1980:AFA

[YLM80] M. S. Yusaf, G. P. Lawes, and N. H. March. Asymptotic forms of atomic scattering factors and momentum densities. International Journal of Quantum Chemistry, 17(5):833–844, May 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Yamabe:1980:TSP

[YMO80] Shinichi Yamabe, Tsutomu Minato, and Yoshihiro Osamura. Theoretical study of photochemical reactions: Electron assignment and the state correlation diagram. International Journal of Quantum Chemistry, 18(1):243–250, July 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Yeager:1981:EMM

[YOJ81] Danny L. Yeager, Jeppe Olsen, and Poul Jørgensen. The evaluation of MCRPA (MCTDHF) electronic excitation energies, oscillator strengths, and polarizabilities: Application to O_{2}. International Journal of Quantum Chemistry, 20(S15): 151–162, March 8–14, 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory, Collision Phenomena, and Computational Quantum Chemistry.

Yoneda:1985:CUS

[Yon85] Shigetaka Yoneda. A calculation by using semiempirical potentials of a complex of nucleotides and the cognate amino acid: A numerical approach to the amino acid recognition by tRNA for the ten typical cases. International Journal of Quantum Chemistry, 28(S12):201–208, March 14–16, 1985. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Young:1984:EEO

[You84] Richard S. Young. Extraterrestrial exploration and the origin of life. International Journal of Quantum Chemistry, 26(S11): 69–72, March 12–15, 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Yi:1985:ESF

[YRS85] Xiao Yi and Chen Run-Sheng. Effect of some factors on the energy band structure of proteins. International Journal of Quantum Chemistry, 28(5):573–579, November 1985. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Yousif:1980:SOD

[YS80a] F. N. Yousif and G. C. Shukla. Second-order dispersion energy between interacting hydrogen pairs by using the Borel integral. International Journal of Quantum Chemistry, 17(4):643–650, April 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Yuan-Sun:1980:CDAa

[YS80b] Kiang Yuan-Sun. Calculation of the determinant of the adjacency matrix and the stability of conjugated molecules. International Journal of Quantum Chemistry, 18(1):331–338, July 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Yuan-Sun:1980:CDAb

[YS80c] Kiang Yuan-Sun. Calculation of the determinant of the adjacency matrix and the stability of conjugated molecules. International Journal of Quantum Chemistry, 18(6):1539, December 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Yamanouch:1984:CIH

[YUKK84] Kaoru Yamanouch, Masashi Uematsu, Tamotsu Kondow, and Kozo Kuchitsu. Collisional ionization of a highly excited neon atom by vibrational deexcitation of tetramethylsilane and neopentane molecules. International Journal of Quantum Chemistry, 26(S18):497–505, March 1–15, 1984. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Supplement: Atomic, Molecular and Solid-State Theory, and Computational Quantum Chemistry.

Yasnikov:1980:HPP

[YVK^{+}80] A. A. Yasnikov, N. V. Volkova, N. P. Kanivets, L. S. Mushketik, L. I. Vasilenok, V. P. Kukhar’, and A. F. Pavlenko. Hydrophobic “pocket” on photosynthesizing membrane near photosystem II. International Journal of Quantum Chemistry, 17(2):361–366, February 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Yamashita:1983:IRC

[YY83] Koichi Yamashita and Tokio Yamabe. The intrinsic reaction coordinate approach in polyatomic reaction dynamics. International Journal of Quantum Chemistry, 24(S17):177–189, March 3–12, 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-state Theory, Collision Phenomena, and Computational Quantum Chemistry.

Yan:1985:EST

[YZ85] Ji-Min Yan and Jian-Guo Zhao. Electronic structures of the π-π type charge-transfer complexes of pyridine with boron trihalides. International Journal of Quantum Chemistry, 27 (4):465–473, April 1985. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Zahradnik:1989:TAI

[ZA89] Rudolf Zahradník and Frank Achenbach. Theoretical approaches to interactions between biomolecules: Legends and approximations to reality. International Journal of Quantum Chemistry, 35(1):167–180, January 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Zorita:1985:LBK

[ZAB85] M. L. Zorita, J. A. Alonso, and L. C. Balbas. Local behavior of the kinetic energy in density functional theory. International Journal of Quantum Chemistry, 27(4):393–406, April 1985. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Zhao:1989:REA

[ZCH89] Cheng-Da Zhao, Bin Chen, and Jing-An Huang. Reaction ergodography for the additions of HLi and its dimer to acetylene. International Journal of Quantum Chemistry, 36(1):5– 14, July 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Zerner:1987:BRB

[Zer87] Michael C. Zerner. Book review: Conceptual quantum chemistry: Models and applications. Special Subject Issues of Croatica Chemica Acta, Vol. 57, Nos. 5 and 6, pages 765– 1674, 1984 Volumes Edited by Zvonimier B. Maksić. International Journal of Quantum Chemistry, 31(4):693, April 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Zerner:1989:CAH

[Zer89] Michael C. Zerner. A configuration-averaged Hartree–Fock procedure. International Journal of Quantum Chemistry, 35 (4):567–575, April 1989. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Zahradnik:1986:WII

[ZH86] Rudolf Zahradník and Pavel Hobza. Weak intermolecular interactions: Statics and dynamics. International Journal of Quantum Chemistry, 29(4):663–676, April 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Zuniga:1987:HST

[ZHB^{+}87] J. Zúñiga, A. Hidalgo, F. Bernabé, A. Fuster, and A. Requena. Hypervirial SCF treatment for vibrational energy levels of triatomic molecules. International Journal of Quantum Chemistry, 32(4):511–516, October 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Zhenyi:1983:AUG

[Zhe83] Wen Zhenyi. Application of the unitary group approach to crystal field theory. International Journal of Quantum Chemistry, 23(3):999–1009, March 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Zhou:1988:GTM

[Zho88] Zhongxiang Zhou. Graph theory of molecular orbitals beyond tight-binding approximation. International Journal of Quantum Chemistry, 34(4):325–332, October 1988. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Zielinski:1982:SSU

[Zie82] Theresa Julia Zielinski. STO-3G study of uracil, thymine, 5fluorouracil, 5-methoxyuracil, and their 4-hydroxytautomeric forms. International Journal of Quantum Chemistry, 22(3): 639–647, September 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Zivkovic:1983:HAS

[Živ83] Tomislav P. Živković. Heats of atomization of some conjugated molecules containing nitrogen or oxygen by a novel semiempirical method. International Journal of Quantum Chemistry, 23(2):679–686, February 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Zivkovic:1986:BBE

[Živ86] Tomislav P. Živković. Butadiene and benzene π-electron SCF ground states in the bond orbital resonance theory approach. International Journal of Quantum Chemistry, 30(5):591–615, November 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Zivkovic:1987:AST

[Živ87a] T. P. Živković. Alternant systems and their properties. International Journal of Quantum Chemistry, 32(3):313–347, September 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Zivkovic:1987:SIA

[Živ87b] T. P. Živković. Spin-independent alternant systems. International Journal of Quantum Chemistry, 32(6):621–653, December 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Zivkovic:1987:BBE

[Živ87c] Tomislav P. Živković. Butadiene and benzene π-electron SCF ground states in the bond orbital resonance theory approach.

International Journal of Quantum Chemistry, 31(6):991, June 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Zivkovic:1987:SRA

[Živ87d] Tomislav P. Živković. Selection rules in alternant systems. International Journal of Quantum Chemistry, 31(4):591–603, April 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Zivkovic:1988:PHS

[Živ88] T. P. Živković. Particle-hole symmetry and complementary quantum systems. International Journal of Quantum Chemistry, 34(4):333–359, October 1988. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Zuhrt:1988:TIH

[ZJŻ88] Christian Zuhrt, Danuta Jeziorek, and Bronislav Żurawski. Theoretical investigation of the Na^{+} + H_{2} system. III. Trajectory calculations of vibrationally inelastic collision processes. International Journal of Quantum Chemistry, 34(6):559–569, December 1988. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Zhan:1987:BSB

[ZLH87] Chang-Guo Zhan, Frank Liu, and Zhen-Min Hu. Bond strength and bond angles for hybrid orbitals composed of arbitrary sets of orbital angular momentum quantum number. International Journal of Quantum Chemistry, 32(1):13–18, July 1987. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Zakrzewska:1981:EPM

[ZLP81] Krystyna Zakrzewska, Richard Lavery, and Bernard Pullman. The electrostatic potential of a model phospholipid monolayer. International Journal of Quantum Chemistry, 20(S8): 161–170, March 5–7, 1981. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Zaitsevskii:1980:PGO

[ZNS80] A. V. Zaitsevskii, A. V. Nemukhin, and N. F. Stepanov. Possible generalization of the optimized diatomics-in molecules theory: Further studies of the diatomic state mixing parameter adjustment. International Journal of Quantum Chemistry, 17(4):679–687, April 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Zakrzewska:1982:DEP

[ZP82] K. Zakrzewska and A. Pullman. The dependence of the electrostatic properties of model phospholipid monolayers on their molecular organization. International Journal of Quantum Chemistry, 22(S9):411–424, March 4–6, 1982. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Zhixing:1983:VNR

[ZRY83] Chen Zhixing, Hong Ruiyu, and Zhang Yala. Violations of the noncrossing rule due to special symmetry and alternance. International Journal of Quantum Chemistry, 24(5):509–519, November 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Zeng:1985:POS

[ZS85] Zong-Hao Zeng and Chia-Chung Sun. The projection operators of SU(2) and the Sanibel coefficients. International Journal of Quantum Chemistry, 28(S19):79–92, March 18–23, 1985. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Zeng:1986:ITO

[ZS86] Zonghao Zeng and Chia-Chung Sun. On the irreducible tensor operators and the unitarily invariant decomposition of Hermitian operators. International Journal of Quantum Chemistry, 29(2):273–283, February 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Ziegler:1980:REB

[ZSB80] T. Ziegler, J. G. Snijders, and E. J. Baerends. Relativistic effects on bonding. International Journal of Quantum Chemistry, 18(S14):393–394, March 10–20, 1980. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular, and Solid-State Theory, Collision Phenomena, Quantum Statistics, and Computational Methods.

Zunger:1985:TSO

[Zun85] Alex Zunger. Ternary semiconductors and ordered pseudobinary alloys: Electronic structure and predictions of new materials. International Journal of Quantum Chemistry, 28(S19): 629–653, March 18–23, 1985. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Quantum Biology and Quantum Pharmacology.

Zasukha:1985:TME

[ZV85] V. A. Zasukha and S. V. Volkov. Theory and mechanism of electron transfer in a condensed medium at high temperatures with allowance for change in vibrational frequencies. International Journal of Quantum Chemistry, 28(1):17–26, July 1985. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Zasukha:1986:TME

[ZV86] V. A. Zasukha and S. V. Volkov. Theory and mechanism of electron transfer at low temperatures. International Journal of Quantum Chemistry, 29(6):1815–1824, June 1986. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Zeiss:1983:SDF

[ZW83] G. D. Zeiss and M. A. Whitehead. Shaping the density to fit one-electron properties: Constrained RHF calculations on N_{2}, FH, CO, and LiH. International Journal of Quantum Chemistry, 24(6):651–685, December 1983. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Zhang:1985:SEE

[ZY85] Qi-Yuan Zhang and Ji-Min Yan. Structure of the electronic energy bands of molecular crystal NMP-TCNQ. International Journal of Quantum Chemistry, 27(4):407–416, April 1985. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).